A Complete Bibliography of Publications in Communications in Statistics: Theory and

Methods: 2020–2029

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

18 October 2024 Version 1.10

Title word cross-reference

(0, 1) [ZXSZ20]. (i, k) [PS22]. (λ, f) [TEÇK22]. (m + 1) [PK22a]. (s, S) [KBKK24]. 2 [KM23b, PZZ20, ZYNX20]. 3 [CTEB20]. 4^{1}2^{n} [ZZZ24]. ^{(a,Y)} [BS21a]. ^{+} [MN21b]. ^{a,b} [SB22]. ^{X} [GC22a, LYY24]. _{1} [Ohy21]. _{c} [SZGK21]. CEV [ZYZ24]. _{p} [Zen20b]. _{P K} [ARAJ22]. A [CLZ24b, OK20]. α [BG22c, EAB20, EG20, EBG22, RBD24, SAF24, Sla22, TEÇK22, ZYS21, ZSL24, ZX22]. X̄ [MSEB23, NSM21]. β [DCT24, PY20, QT22, ZFDT24]. C_{pk} [BAAJ20, BA24a, Mat23]. D [CLZ24b, GC24a, HTP22, LZ23a, ZWLF24, ZYC24, GZ23a]. δ [Ano23c, LNP20, MBLY21, MSW23, PS23c, Pou20, SS22d]. Δ_{λ} [AEY21]. Δ_{v}^{m} [EG20]. df [KÇ21a]. E [Han20]. ℓ_{1} [Bey23]. η [Ass21, BMH24]. F [RJ23, SPZ23, Bar24, GWD24]. G [AE23b, GDSC15, KN22, KC21b, LNXC24, NT20a, SPZ24, PW23, Sav21, SJF21]. G_{n}^{(a,b)} [PG22]. γ [NH20]. h [PW23, SMH22]. h + 1 [MMSC22]. H ∈ (0, 1/2) [CL21a]. K [CWM23, DR22, FZ22b, JZC24, MiT24, WZA21, WLZ^{+}24, ZWW24, AALE22, BNO22, BG22b, BR23, CL23, EJJ23, FZL24, HLPL22, HN21, Jas20, Jas21, JG23, KLS23, KWB23, KHF22, LHWK24, LNXC24, MR21, MBLB24, RG21b, YDHY24, ZQX23]. k/r [PS23b]. L [DZ20]. L^{2} [You23a]. l^{∞} [BA23]. L^{p} [LF23]. L^{p}(1 < p < 2) [ZT21]. L_{0} [ZLZZ24]. L_{1} [AEBK20, ZSNX23]. L_{2} [Cha24]. L_{r} [jWWDjW24]. λ [EB21, YGAK23]. M [BG22c, FZW23, Ril24, SL20, WTS24, ZCL24, ZLH20, BKM20, CLW^{+}22, EB21, GZ23b, WS23, jWWDjW24, WWWZ24, ZHQ23]. M^{X}/(\frac{G_{1}}{G_{2}} )/1 [KC20]. P(X < Y) [PT22]. μ [Gül23]. N [BC24, KC20, LWKQ20, MiT24, WLZ^{+}23a, WH23a, Yan22a, ZLCW20, BR23, CLWZ22, EJJ23, FZL24, KHF22, LHWK24, LNXC24, NSG23, PS23b, RG21b, XL24, Zha21a, ZYNX20]. n: G [HLPL22]. P [HY23, KERAE23, LF23, LWD^{+}21, New23, ZCE23]. P (X < Y) [MHS22]. Φ [CPS21, LDZ^{+}22, dS20]. Φ_{q} [LY24]. ψ [DCY20]. Q [ADW22, Mon24, Rao22, RG22, Cha21b, Cha22, Cha23c, Cha23d, IAM21, KV23, NA23, OMS24, Vam20]. R [LLZ24, LZ24, PS24]. R = P (X < Y) [Koh20]. R^{2} [BMR^{+}23]. ρ [KdXYC23, TZ21, YC23]. S [JRD24]. S_{n} [KA20]. S_{pk} [NN21]. T [ÇA20, Aga24, AKUT23, Cel22, Gau21a, GB23b, JWW24, MPS23, MN23c, Oga24, OV21, XL23a]. T^{2} [HS20, NKK^{+}24]. τ [SH23]. R̃ [KB22b]. ρ̃ [ZSY21]. U [BN24, CC23, Ril24]. W [YSF22, MMSC22, Tee22]. X [MHS22]. Y [MHS22]. Z [BEhF24, CZL24, YSF22].

-aberration [DCT24, QT22]. -ANA [WWWZ24]. -analog [KV23]. -arm [PK22a]. -Bayesian [Han20]. -boundedness [EB21]. -Cauchy [Vam20]. -combined [CWM23]. -component [CLWZ22]. -conditional [AEBK20]. -convergence [EB21]. -convex [BMH24]. -D [KM23b]. -dependent [GZ23b]. -dimensional [CTEB20, GZ23a]. -distributed [MN23c]. -distribution [Gau21a, IAM21, MPS23, Oga24]. -Distributions [Vam20, Cha23d, GB23b]. -divergence [NH20]. -efficiency [LLZ24]. -efficient [OK20]. -END [CLW^{+}22, ZHQ23]. -estimation [FZW23, SL20, ZSNX23]. -estimator [BG22c, ZLH20]. -estimators [BEhF24, Ril24, WTS24]. -expectation [SJF21]. -extended [jWWDjW24]. -Factorial [Cha23d]. -function [RG22]. -functions [Tee22]. -Gaussian [NA23, OMS24]. -generalized [BKM20]. -group [WZA21]. -harmonic [SMH22]. -independent [Rao22]. -Jensen [RBD24]. -means [DR22, FZ22b]. -mixed [WLZ^{+}24]. -mixing [BG22c, DCY20, KdXYC23, LDZ^{+}22, Sla22, TZ21, YC23, ZSY21, dS20]. -mixture [SAF24]. -model [PY20]. -moment [GWD24]. -Multinomial [Cha23c, Cha21b, Cha21b]. -NN [MBLB24]. -norm [Bey23, KB22b]. -of- [MMSC22]. -optimal [CLZ24b, GC24a, HTP22, JZC24, LY24, LZ23a, LZ24, PS24, ZWLF24, ZYC24]. -order [LF23]. -out-of- [BR23, FZL24, KHF22, LHWK24, LNXC24, MiT24, PS23b, RG21b, ZYNX20, HLPL22]. -player [WH23a]. -policy [KC20, LWKQ20, WLZ^{+}23a, Yan22a, ZLCW20]. -Pólya [Cha22]. -posterior [Han20]. -preserving [SH23]. -record [BG22b]. -regularization [ZLZZ24]. -rules [YSF22]. -sample [BNO22, CL23, KWB23]. -series [ZX22]. -shock [LNP20, MSW23, Pou20, SS22d]. -sorts [ADW22]. -stable [ZYS21, ZSL24]. -statistic [Aga24]. -statistical [Gül23, KÇ21a, TEÇK22, AEY21, EG20]. -statistics [Bar24, BN24, CC23, Ril24]. -symmetric [Mon24]. -test [ZCL24]. -th [EB21, Jas20, Jas21, PS22, YDHY24]. -th-to-default [ZQX23]. -Tuple [JG23]. -type [HS20]. -value [ZCE23]. -valued [BA23, CZL24]. -values [KERAE23, New23, HY23]. -weak [Ass21]. -Wijsman [YGAK23]. -WOD [WS23]. -wordlength [ZFDT24].

/1 [BS21a, PG22, SB22]. /G/1/N [GC22a]. /M/1 [LYY24].

03610926.2020.1737711 [Cha23c].

1 [BS21a, BS24, CVDY21, GC22a, Kia22, KG23, KJ23, LYY24, MR21, PG22, SB22, SACQ23, TZ24]. 10.1080 [Cha23c]. 10.1080/03610926.2019.1703135 [KS22]. 10.1080/03610926.2020.1737711 [Cha23c]. 11th [DNP21]. 13ARIMA [CGLP24].

2- [MMSC22]. 2-orthogonal [Moh23]. 2022 [AE23c]. 2072 [AE23a].

3-interactions [FBNP23].

4/2 [LJ24]. 4095 [Ano23b]. 4240 [AE23b]. 44 [Ano23b, AE23b]. 45 [AE23c]. 46 [AE23d]. 47 [AE23a]. 4755 [AE23d].

A-optimal [ZH24]. A.s [LH22]. Aalen [sS22j]. Abdalla [BAE23]. Abdelghaly [BAE23]. aberration [DZZ23a, DCT24, JZ24, QT22, ZZZ24]. abridged [MU22]. absence [DSB22]. Absolute [ZS23, CLC22, DZZ23b, FZW23, KM24b, WWWH20]. Absolutely [BBP22b, GT22, UET24]. Abundance [LZLL20]. AC [Ohy21]. academic [KM24a]. accelerated [BSN22, CW23, FTTY22, He23, HKL21, JB22c, sK23b, LGX^{+}22, NSG22, Nak21, SS23b, SS24a, SWHW22, WT20, YLLZ24]. Accelerating [AÜ20, MN23c]. acceptance [BNAK23, BAJ20, RL23]. access [CW23]. accident [Gar21]. accommodation [ZLLS21]. accomplishments [GC21]. accounting [sK24a]. accumulation [LGLL21]. accuracy [HM22, JC22, RIA20, SCAK23]. across [BH24]. actions [BFSN22]. active [SGZA23, YW22]. activities [Gos22]. Actual [DD23]. Adaptive [Haq22, KOTAG21, nLSZ24, LS23, Liu24b, NS22b, OO20, TBCB24, YM22, AAM22, BL20, GCFS23, KR24b, sK20c, Li21b, MLSS24, PKK22, QG23, QH21, SSA23, SGTG24, SMP^{+}24, TH23, Wan23c, YY21, YT20]. adapts [DR21]. added [PS23d, WRZL21]. additional [JTS22, Li22]. Additive [JBSK22, AFFM21, AFFM23, Ben22b, CFS20, CB24a, CB24b, DL22, FSS22, Han23b, HKL21, HSC22, KIA^{+}21, LY24, LMH21, LL23a, LX22b, MWLS24, Sha22a, SZ24, VCOB22, ZL20, ZP22, Zha24a]. additive-multiplicative [DL22]. address [LZ21]. Addressing [YHA20]. Adjusted [TZYY22, TQ24, BPZ24, DC24, LS22b, RJ23, Zhu22a, ZHA20b]. adjusting [CGLP24]. adjustment [CG21, KO23a, LWK20, PLSW23, Wu21]. Admissibility [LY22b, CPS21, SBZ21]. adults [KL22a]. advanced [KSM^{+}21]. advances [Ale23]. Adversarial [EJJ23]. AEP [WYJD21]. Aeppli [KLS23]. affect [KASK24]. affiliation [WZL22a]. after [Gar21, HO24, LW23]. against [DMD23, GM20a, IM21, JAE23, RSA20, YO23, ZZ21]. age [IM21, MU22, SAK21, ZGQN20]. age-based [ZGQN20]. age-replacement [SAK21]. ageing [BGM22, GND^{+}23, PK22c]. aggregate [GLK24, JC20, LGL20, WQZL21]. aggregates [AD21]. aggregation [CH20b, LTR24]. aging [EAAC22, LL23b]. agreement [Ohy20]. agricultural [VCOS24]. Ahmed [BAE23]. AIC [OMYF21, SZGK21]. aid [Bai21]. air [SRSB20]. algebraic [SG24]. algorithm [CX22a, FZ22b, Hu20b, HFJ21, KOTAG21, MNAM20, MAT24a, MKG^{+}21, PCdP^{+}20, PAC20, QSP^{+}24, RG23, SMZ23, SK20a, WW22, YZC22, YG20, ZLB20]. algorithms [Che23e, MN23c, UN23, WY23]. Aligned [JAE23]. Aljarrah [Ano23b]. allocating [sK23a]. allocation [CLWZ22, GKvZB21, GA22, MNAM20, MAT24a, New23, WH23b, WLL23, WDLW20]. allocations [MGK22]. alloy [YXS^{+}24]. Almost [Ass21, Aug21, DT21b, SST22, XWH21, YXZ23, dS20, dS24a, ARAK22, rChZhN^{+}24, CDL24b, DC22b, KHZ24, MWW23, NTDB23, PP22a, SW24, TTL20, Wu22a, ZLH20]. alpha [KR23b]. alternate [sK20b]. alternating [PP22b]. Alternative [Oga20a, Alt21, AE23a, Bro22, DG21a, ERAS22, Far24c, GH18, KK23c, SPSS20, SPS21, SMP^{+}24, SW21, ZZ21]. alternatives [GM20a]. ambiguity [LZL20]. American [HCCM21, LYLY21]. among [BCH24, HN21]. amounts [ADIK22, FZD24]. ANA [Ko23c, WWWZ24]. analog [CBJ20, KV23, SC20]. analyses [YHH24]. Analysis [BC21, BC24, BAJ20, GPD21, KKNN23, KG21, KG23, LX21, LYY24, TZ24, XLW23, dHGS21, AYP21, Ale23, AGHK20, AAGE23, BHL21, Bai21, BAY20a, BP24, BA24a, Bic22c, BN22, CSZ20, CX24, CMP21, Che20a, CLF21, CZX24, DDG23, DG21b, DK23, EJJ23, EIAZ23b, FHZC23, FZL24, FMZM22, FS22, Fus22, Gao23, GB23b, HSWZ21, HKNW20, He23, HCX20, HX22, HLPL22, HWW24, JRD24, JS23b, KH20a, KW22, KS24b, KR21, KP21, KB23b, sK20d, sK22a, KF24, KSK^{+}23, LWY^{+}23, LHWK24, LWK20, LHS20, LXWY21, LL23c, LXLL23, LWYJ24, LWS24, LZ23c, MN23b, MU22, MNA22, NSS24, NCM22, NS21b, NC23, NN24, PB24, PMMK24, QQS^{+}22, RSG21, RFVR20, SS22a, SG24, SB22, SYRV20, sSL20, SK20a, SY21, SK24b, SJT21, SWLZ23, TW20, Ünv21b, Ünv21c, Ünv21d, UMA22, VE21, WZTW20, WT20, WXY21, WLZ^{+}23b, WLZ^{+}24, WjWbK22]. analysis [dXG20, Y.24, Yan22a, YSS22a, YL23, YY24, YZT21, ZSCH22, ZXLT21]. Analytic [Kwo22]. Analytical [BS21a, Zen20a, Zen23]. analyze [ADW22]. Analyzing [KKS24b, LA23, Pri24, Wic20, SYT24, WW22]. Anderson [MKM24, TB22]. angles [Han23a]. Anisotropic [Ben22a, BL20, GK24]. announcements [MDZN20]. annuities [AÜ20]. anomalous [Gos22]. ANOVA [ADW23, Cel22, LaM24, RJ23, SKRR23, WD20, AR24]. ANova-based [AR24]. answers [SHC20]. ANVILS [AR24]. ANVILS-VOCE [AR24]. aperiodic [ZOD20]. APND [NSJ22]. Application [Bic22c, CMMD20, KLK24, ARGD21, AP24, BPZ22, BAY20a, BAE23, BA21, BBP22b, CZX24, CLS24, CLW^{+}22, CW23, EEAA22, Fu24, GCCJO^{+}21, GÖ24, HW22c, Hua20, Jan20, JL22b, KdXYC23, KTMKS20, KK22, KAJ^{+}22, KP21, KK23c, Lee21, LBF24, LW22a, LX23b, LCYL24, MA23, ML23, MAAZ22, NSG22, NMT22, PA21, PKSS22, PVC^{+}23, Qar22, QZ22, SLV^{+}20, SS22g, Sho24, SRSB20, TTL20, TTLT23, Tan24, TG24, Ünv21a, UMA22, VRB22, WYJD21, WZL23, zWZ20, YXS^{+}24, YL23, YN24, YLC^{+}22, ZXSZ20, ZZZ20, ZZZ22, ZYC23, ZHA20b]. Applications [HCLW22, KK21, AMH22, AAQ23, Alv23, AHO22, AP20, Ary23, AE23d, BMK21, BCG24, BD22, CH20a, Che21, CRKC21, DS24b, DCY20, FZ22b, FTAL23, HY23, bJwWyX23, KS24a, KIA^{+}21, KH20c, KR23a, KS24c, KJ24, KS24d, LS20a, Li21a, LL24c, LYD24, LW21a, MS21, Mat24b, MAK20, MRD20, Nag20, OTL21, PWLO17, RSR24, RKAA21, RBCG20, RBD24, RBS20, RTN24, SRA20, Shi23a, Sho23, SK20e, SK24c, SS24c, Tah20, TT22, The22, WSSY20, WC20b, WWWZ24, Wir23, Wu22b, XY21, YCY22, ZAL21, ZZ23, dS20]. applied [AF23, EM21, Nou22, OCSV23, ROSL21]. Applying [SK24b]. Approach [Muk24, AF23, And20, APV24, AKUT23, ARK22, BC24, BP20, CZ20a, Cha24, CRVK24, Dij23, Dün21b, EIAZ23a, Fow22, GX21, GCCJO^{+}21, GdNM20, GB23a, GCFS23, HM22, HKL21, JS23b, Kha24, KP21, sK20b, sK20c, KK24, LZ21, LG24a, LX23b, LWS24, LZ23c, Moh22b, MRD20, Mur22, NSSR23, NNS20, NuAMA22, ÖB21, ONM^{+}24, Pra21, QSP^{+}24, Sal24, SCAK23, SS22h, SBK24, SS20a, SS22i, SJ23, SCP23, SSZC22, SWYF20, SZL21, TLRB21, WC24, XS21, YHA20, Yu22a, ZL21, ZG23, ZWW24, ZZZ20, vdMGSM21]. approaches [GZ20, KH20a, KF24, RKS22, RBSNB20, RG23]. approximants [YY24]. Approximate [JWW24, LCWZ21, ZGQN20, FSB22, JWW22, LJL^{+}21, Wir23, ZPL23]. Approximating [PF20, JBJ23]. Approximation [NC23, NCNM23, STW23, WSSY20, BS20b, HTP22, KU22, LXLL23, Maa23, Mon24, MW24b, NMT22, NK22, SP22, SS22k, Tee22, WV24, Wyw24b, YNC22, YN24]. approximations [BLZ23, CBJ20, CMMD20, EAS20, KBKK24]. arbitrarily [qDSzW21]. arbitrary [FW23, FLW24, Kon22]. ARCH [Gon21b, LYC20, ZYZ20]. ARCH-M [ZYZ20]. Archimedean [BAA24, BBAA21, FZL24, ZLF24]. arctan [JV24]. area [AC23, CN23b, Jan20, KNCK24, SRSB20, Wu20, vRvR20]. areas [MN21a]. ARFIMA [PHBV23]. Arising [Del21, CZ21, DK24a]. arithmetic [Fro23]. arm [GWM21, PK22a, ZWCP22]. ARMA [SA22, HO24, SWLZ23]. array [HT24, HCCzW23, HFGZ24, OYA24, RZ23]. arrays [CCW^{+}20, CQT21, rChZhN^{+}24, qDSzW21, FWWH21, HW23b, LW20, LF20, LM22, LM23b, PZZ20, PYZS23, PCZ24, SAP20, SW24, Tsa22, WYP21, jWWDjW24, ZDY24, ZLPZ24, ZFDT24]. arrival [BC21, CY20, GPD21, KG21, Kum24, LWKQ20, NP23, PG22]. arrivals [AD21, KG21, KSK^{+}23, LCF24, SCGB24, YTT23]. ARTICLE [Ano21b, PS23c]. Artificial [AdMSK23, GA20]. ascent [Lee22b]. aspects [ASS24, CSBM24, KH20c, NSG21]. Assessing [HM22, OC22a, PKSS22, HHKC23, RKS22]. Assessment [EM22, GSED24, Ker20, Mod23, PK22a, Pra21]. asset [GKvZB21, PC22, SLS21, WDLW20, ZW22b]. asset-liability [PC22, SLS21, ZW22b]. assets [Raï24, RYZ24, YZSS23]. assignable [NSM21]. assignment [YD23]. Associate [Che23c]. associated [BHB22, BGT23, CDL24a, CC23, DC22b, Fak23b, GD21, HMZ20, HW22c, KB23a, Ko20, Ko23d, LM20, MWW23, MW24a, Mir23, Moh23, PZAH22, SW24, WW20, WF23, ZB23b, ZLH20, ZLF24, ZWWW24, ZSYM21]. association [Dou22, Ünv21d, ZLS23b]. assumption [DG24, KAJ^{+}22]. assumptions [DY21, Qu22]. Assurance [HW20b]. Asymmetric [BBM^{+}22, MvSK20, GG24a, JKP20, KR21, LC24, Sem20, SCPAV21, Swa24, WYP21, ZDY24, ZLPZ24]. asymmetrical [PZZ20, YDHY24]. Asymptoic [Yo22]. Asymptotic [AO24, Alm24, BABF22, BBM24, BS20a, BS22a, BKS23b, BN24, CW20a, CWCY23, CY23, CY20, CB24a, CC23, DP22, GG24a, GSH22, HMZ20, HS20, Jas20, Jas21, JLG21, JPJB22, KW22, iK22, Lee22a, Li24, LZG20b, LLWA22, MZ22, MBLB24, NSH23, NN24, ND21, Oga20b, SGW23, WWWH20, WDW24, WGX21, dXG20, ZB23b, ZL22b, AAE22, AB22, Aug21, CMO^{+}23, CSBM24, DG21b, Fre20b, Gho23, HB20, HO23, KK21, LMH24, LLH^{+}24, Lou20, LQWW21, LMZ23, MKM24, MGK23, MÁH22, Mir23, Mot23, Nes24, Oga20c, ST24, hSsXfX24, TH24, Ten24, WV24, YSF22, Yam24, YLJ23]. asymptotically [DC22b, Ko23d, MWW23, MW24a, SW24, WW20, ZLH20]. Asymptotics [Ano24, CY24, LK22a, LCF24, Mat21, PY20, WM21, WWL22, WTS24, Yos23, YJZZ23, AO22, JPZ23, Lin21, LG24b, SYL23, WQSG22, ZPJY24]. atmospheric [Gar21]. ATS [KRCS24]. ATS-unbiased [KRCS24]. attacks [FL20a]. Attainments [iK21a]. attitude [MAT24a]. attribute [Arn23b, LS22b, LHS22, NS21c, ZLZ20b, ZYZ24]. attributed [SS20a]. ATTRIVAR [CN22]. auctions [EJJ23]. augmentation [ZK24]. augmented [FL20b, LK22a, LOL21, LWYL23, RAA22]. auto [SA22]. Autocorrelated [CLPS24, BA24a, BA24b, MB21a, SMMC21]. autocorrelation [XL23b]. autocovariance [Cav20, Cav24]. autoregression [Kur22, LMH24, PWM20, TY23]. autoregressions [Spe24, Wic20]. Autoregressive [Ano23c, SB23b, AB24, Ami23, BM23, BK24, FP23, FW22, GW21, HWZX23, JAE23, KA22, LP20, LY21, Li23a, LL23a, LZ20b, Liu24b, LKK24, LW22c, LW21b, MPS23, MBA20, Nes24, PWM20, PS23c, SHA24, Sha23, SS24b, Tan22c, WWC24, XW22, XWCT21, YLJ23, YPD23, YK20, ZCT^{+}20, ZLS23a, ZSYD24, ZGCZ23, dCBV21]. Auxiliary [HA22, uAK22, AAHS21, CLF21, GA22, Haq20, HDZL21, KNCK24, Pri20, SPSS20, SV22, WKS21, YKW23, ZY24b]. Availability [Gao23, Y.24, ZOD20]. Average [Ano23c, WS22, ANO23a, ACK23, BMG22, BGM22, Che23a, CW24a, CW24b, Din24b, GW21, HCL21, Kit21, LQ21, LWYL23, MM20, Nko23, NuAMA22, PS23c, Qu22, Sha23, SA22, SLYF23, WS23, Wu23d, Xia20, YN21a]. averages [Bic22b, uA21]. averaging [HCZ20, Liu24a, SZ24]. aversion [CLC22, Li23b]. awareness [HT23]. Ayala [ZSL24].

B [AE23d]. backed [GQW21]. background [LJLW23]. backward [CM24, JL24a, ZT21]. bag [Che24b]. balance [GC24a]. balanced [ATJ^{+}24, BAGR22, CPS21, DAS^{+}20, Gon21a, HNR^{+}23, HAA^{+}20, HGAJ21, JBR^{+}24, JRB^{+}23, KG20, Kha20, LY22b, MO22, RASD23, RNA^{+}22, SS23a, SSS23, SKRR23, SS21a, TBCB24, VPM^{+}24, YW24]. balking [BS24, KLSZ22, KJ23]. Banach [JSB22]. Banach-valued [JSB22]. band [HJ22, JCY22]. banded [YK23]. bands [MZ23]. Bandwidth [BS20b, AGB21, CB24b, EW23, Ory20]. bandwidths [Ten24]. Barlow [YF22]. barrier [MB20, Zha21a, ZL22a]. barriers [ME21]. based [AS22a, AR24, AU22, AN23, AFO21, AAGN24, ARAK22, Akd22, AYS21, AMBK22, AZAZ24, AÖK^{+}24, AD21, AHS^{+}24, AAAJ24, ABJR22, AuAHS23, AA22c, BBCL24, BAAJ20, BNO22, BNAK23, BEAA21, BT22, BFSN22, BL24a, BFPM22, BC23a, BGM22, BN22, BD22, BS21b, BLZ23, Bul20, CMO^{+}23, CZ20a, CGW23, CMP21, CRVK24, CH20b, CF22b, CZZ^{+}23, CLLD23, Che24b, CS24, CC23, CMM20, DKD22, DR21, Dij23, DZ20, EA21, EN20, FSS22, FL20b, GX21, GvRS24, GS24, GG23b, GM20b, GDSJ22, GA23, GZ21, Han20, Han23b, Han24, HA22, HM24, HCCM21, He23, HHKC23, HTP22, HHW21, HXCW22, HL22a, HX22, HS20, JAJ21, JC22, JRM21, KBKK24, KWC21, KWB23, KTDW20, KF23, KTMKS20, KH21, sK23b, Koh20, KGS20, Kur24, LL20, LXZ22, LWY^{+}23, LNWA22, LVB23, LGLL21, LYLY21, LGX^{+}22, LSY22]. based [LG22, LX23b, LK24, LDZ^{+}22, MGK22, Mak24, MMG24, MAT24a, Mao20, MZ23, MPM20, ML21, MPS23, MN23c, Mod23, MBA20, MAT24c, MOAF23, MHS22, NMF23, NN21, New23, NC23, uAK22, Nou23, OHS21, OFH23, PK20a, PK22c, PCdP^{+}20, Psa22, QSCH22, QG23, QQS^{+}22, RR20, RSR24, RL23, RG21a, RZ23, RG22, Ros22, SCAK23, SLV^{+}20, SAH^{+}23, SANHS21, SCPAV21, SS22i, SS24a, SA22, SWK24, ySLZ22, SF24a, SWHW22, SY21, SCGB24, SL20, SZW21, SH23, TH23, TBCB24, Tri21, UC21, VCOB22, VE21, Ver20, VK23b, WWZ20, WT20, WJS22, WCT22, Wan23c, WS23, WLS23, WJL23, Wu20, Wu22a, Wu23a, WK23, XZZ22, XL22, XWYZ23, XZL20, YM24, YCS20, YHA20, Yin24, Yu22c, ZCT^{+}20, ZHQ21, ZL23a, ZHQ23, ZQH23, ZZK23, ZWWW24, ZGQN20, ZYZ20, ZL22c, ZGCZ23, ZWW24, ZOD20, ZLZ20b, ZXYJ24, ZYZ24]. based [ZLNS22, ZPL23, vHP23]. Basic [CWM23, Ham24, YD23]. basis [BD24a, ZG21]. basket [Ünv21d]. batch [BC21, GPD21, KG21, KSK^{+}23, LLXW24, NP23, PG22]. batch-size-dependent [NP23, PG22]. bathtub [LG22]. bathtub-shaped [LG22]. Bayes [Alv23, AC23, BC23a, Bic22a, CVDY21, CYW21, DLC22, Fow23, JWW22, KSA23, Lig23, MN21b, Mat23, ML23, MU22, SA21, TK24, YTT23, ZXSZ20, ZZW^{+}24]. Bayesian [SM20, AB24, ASG24, AFO21, AFG20, AYP21, AAR22, Ami23, AD21, And20, APV24, AA22a, Ary23, AKUT23, AWN21, BR22, BM21, BS24, CSZ20, CX24, CZ20a, Che20a, CN23b, CG23, CZX24, Dag23, DHJS23, DE22, DG21a, FMZM22, Fow22, FS22, FSB22, GLB22, GZ21, GdNM20, HK20, Ham23, Han20, Han21a, Han23a, He23, HZ21, HHLG22, HX23, HK21, JWW24, KLK22, KW22, KGG21, Kia22, KAC22, Koh20, iK21a, iK22, KA22, LY24, LH21, LWY^{+}23, LL23c, LZW23, LM23a, MN21a, MRDD22, Mur22, NSD22, NH20, NS21b, NMT22, NuAMA22, OV21, PB24, PKSS22, Pra21, RSY21, RGS24, RBC^{+}23, SHFD23, SBK24, Sha23, Sho24, SACQ23, SCGB24, SMV23, Spe24, SZW21, SWLZ23, Tan22b, TT23, TÖ23, Tsi24, Wan23c, WC24, XTXW20, XL23a, YPD23, Yao22, Yu22a, Yu22b, ZN24, Zha22, ZXLT21]. Bayesian [ZPL23, ZCE23, vdMGSM21]. Bayesian-deep [APV24]. Bayesian-frequentist [ZN24]. Bayesian-Frequentists [vdMGSM21]. be [HNR^{+}23]. BE2 [ZZK23]. bearing [Pra21]. becomes [KT24]. Begg [MG24]. Behavior [Dey24, AHO22, CWCY23, DG21b, GLW24, Jas20, JL22a, Ko23d, Lee22a, Lou20, MBLY21, NSH23, Pri24, WGX21, XFP23, ZB23b, ZL22b]. Behavioral [BL23, WCY23]. behaviors [LMH24, XM21, ZYJ20, ZQH23]. behaviour [Lin20]. Behnken [RAA22]. Behrens [CBJ20, CLLD23, YFW^{+}23, YKW23]. Belief [ZD23, KK22, KK23b, LCYL24]. beliefs [CLZL24, HW23a]. Bell [Akd22, CLMA20, Pui24]. Benchmark [Ker20]. benefit [SCAK23]. Benford [Cam24, KK20]. Bennett [Wu22d]. bent [Dag22, Dag23, LH21]. bent-cable [Dag22, Dag23]. Bermudan [Alg23]. Bernoulli [BN21, BC21, BC24, BCH24, EG23, HCL22, HHKL23, KNL24, Nah22]. Bernstein [BN21, Bah24, BL24a, DBBK21, Kha24, SA21, WHCC23]. Berry [CL21a, CDL24b, Dou22, ZAF24]. Berry-Esséen [CDL24b]. Bessel [MP21b, YD24b]. best [IK21b, MRDD22]. Beta [FMM^{+}22, Alt21, EN20, GDSJ22, HZZA23, Jan20, JL22a, JB22b, KD24, NSH23, PB22, VdWMB23, ZXSZ20]. beta-binomial [ZXSZ20]. beta-generated [VdWMB23]. beta-normal [JL22a]. beta-Wishart [NSH23]. better [Wu24a]. between [CP22, FW23, FLW24, JC22, KHF22, KRCS24, NS21a, NCC24, Ohy20, Psa22, RMS^{+}24, Ril24, WRZL21, WJL23, XFP23, ZW23, dSO22]. beyond [KLS23]. BHHJ [Kur24]. Bi [AG23, AFFM23, DLZN23, KR21, KJ23, SLF^{+}22, WZL^{+}22b]. Bi-additive [AFFM23]. bi-degree [WZL^{+}22b]. bi-infinite [DLZN23, SLF^{+}22]. bi-level [KJ23]. bi-modal [KR21]. Bi-seasonal [AG23]. Biadditive [AOGM24]. Bias [Oku23, HZZA23, Lab23, LTTL24, MG24, MLSS24, OMYF21, PAdlPK22, Ros22, TWZ22, Wir23]. bias-corrected [OMYF21]. biased [AAQ23, AW20, BY24, KCG21, KC23, OIF21, OCSV23, QMS22, RSR22, SLX22, SZ20b, WS20, Xu22, XH24, ZAF24, ZAG24]. bibliometric [NCM22]. BIC [Cha23f]. Bickel [Bic24]. bidders [EJJ23]. bidimensional [CWCY23, GLK24, SGW23, WQSG22]. bifurcating [SYT21, ZSYM21]. big [RZ23, SY21]. bilinear [ASS23b, BM20b, KB23b, MB21b]. bimodal [Bay20b]. bimodality [SG24]. binary [Elm22, EA22, IY23, KS21, MC23, Ohy20, Ohy21, Özk21, Pho24, SYT21, SPS23, SMP^{+}24, WZA21, Wu22c, ZSYM21]. Binomial [CLZ21, RG22, ASG24, ADD22, Alm24, Ary23, Baz24b, BA21, CAAARCTS23, FVCdC23, Ham23, Han21a, HCCM21, JB22b, KWY20, Kum24, LHS22, MMLBP20, OTL21, RN20, RS24, Sho23, Sil23, SK20e, TIIB22, Tee22, WJL23, Xia20, YSS22b, ZXSZ20, ZSYD24, ZHS^{+}24, IAM21]. binomial-generalized [Ary23]. binomial-Lindley [TIIB22]. biomarker [AMBK22]. biomarkers [AT22, sK20d]. biosimilarity [PK22a]. bipartite [PH24, VPM^{+}24]. Birnbaum [Cha23a, DGCL23, OCSV23, SLV^{+}20, VVS^{+}24, YF22]. birth [AFG20]. birthday [Con24, MW24b]. Bisymmetry [Yam20]. Bitcoin [Ünv21c]. Bivariate [ARGD21, Bay20b, JRN24, NSSR23, SRA20, AKST22, AAGN24, ABXE22, BP24, Bou21a, Cha23d, CZL24, CDSG20, Dag23, DBBK21, GS24, GH22, HP20, HABE24, KLK21, KLK23, KO23a, KAJ^{+}22, KH21, KHD24, KR23b, LWS24, NSN20, NCGB24, PB24, PA21, SS22c, SLV^{+}20, She22, SK20a, SK20e, Swa24, UET24, WWC24, WSL23, WGX21, vdHvDZ22]. blinded [WBM^{+}23]. block [BG22a, BD24b, GPM23, HM22, KG20, Kha20, Kha22, PPM21, SF24a, SX23, VPM^{+}24]. block-sum [Kha20, Kha22]. blocked [WLZ20]. blocks [FF23, Nie20, RNA^{+}22]. Blockwise [SZGK21]. BLUEs [HMP23]. Blum [Gai22]. BLUPs [GB21, GB23a]. BMAP [BS21a, KG23, PG22]. BMAP/ [PG22]. BMAP/G [BS21a]. BMAP/MSP/1 [KG23]. Bochner [Pos24]. bond [DLHZ21, SZL21]. boosting [Cha24, UC21]. boosting-based [UC21]. Bootstrap [ÖA23, YFW^{+}23, ACCM24, HS24a, HZ21, RKK^{+}21, RIA20, SS21a]. Bootstrapping [HO24, RO23, IW22]. bootstraps [Che24b]. both [BPZ24, BFSN22, CH21a, CMM24, SS21b, YZC22]. bound [CL21a, CQT21, KT21, LLW23, Tee22, ZAF24]. Boundary [FM24, ARF23]. Boundary-free [FM24]. bounded [DSY21, MQK23, SDK^{+}24, TEHEG22, TP24b]. boundedness [AEY21, EB21, EBG22, TEÇK22]. Bounds [Fre20a, LL23d, ADIK22, CDL24b, DM24, DM22, GZZ21, GA23, Goo24, iK21a, iK22, LWYL23, LLS20, Sho23, Tan24, WM22, WC20c, XXBD23, Yin24, ZLB20]. Box [FHZC23, JZWT20, RAA22]. Box-Behnken [RAA22]. branching [GZ23a, LH22, LTW24, RK21, RK22, SBHY24]. break [SWLZ23]. break-points [SWLZ23]. breakdown [KC20]. breakdowns [KLSZ22, LXWY21]. breaks [CS24, XP21]. breast [AWN21]. Bregman [DE22]. bridge [HKL21, LS23, TG24]. bridging [ZN24]. brief [GH22]. brittle [SS22g]. broad [GSED24]. Broken [Muk24]. Brownian [DDM24, El 23, FE23, HCLW22, LYLY21, LCF24, Rao23, RTT23, STW23, XY23, YZ20, YCS23]. BSDE [MGK22, SY23, SZL21]. BSDE-based [MGK22]. BSDEs [HWWY20, LF23, ME21]. Bühlmann [YW24]. building [WCT22]. Burr [AAM22, BmAB22, Koh20, LKK24, SP24a, ZLF24]. busy [BTH23]. by-claim [CW20a].

C [Ano23b, ARAJ22, AE23b, AE23c, Cha23c, ZHQ23]. cable [Dag22, Dag23]. calculation [CAAARCTS23, HHKC23, WZA20, Wu22b]. calculations [WZL23, ZGQN20]. calibrated [ASS23a, ASI^{+}24, Bic23, Bic24, SSS22]. Calibrating [HL23]. Calibration [ASA24, PKT21, SBB21, CRVK24, GP20, MRD20, MRDD22, OMYF21, SSNP20, SF24b, ZL22b]. calibration-based [CRVK24]. call [YN24]. Cambanis [HABE24]. can [HNR^{+}23]. cancer [AWN21, KM24b]. candlestick [HWW24]. canonical [BN22, LLX21, LXLL23]. Cantelli [Mot24]. capability [BBP22a, GG23a, Mat23, Oh23, OYA24, RKS22, RL23, RAA22, SWHW22, WWJ23]. Capital [WH23b, CZLY22, GK20b, MGK22, TLL23]. capture [LZLL20, PSLLN20]. capture-recapture [LZLL20, PSLLN20]. Carlo [JC22, KOTAG21, SM20, TG24]. carryover [Gon21a]. Case [PLSW23, AMM20, ATAAM20, Bul23, CLF21, EJJ23, Fre20b, GZ21, Ham24, KH20a, KT24, LWS24, MMLBP20, NS21a, NNS20, SP24b, Sha22a, SU22, Wu22c, ZWW24]. Case-cohort [PLSW23, LWS24, Sha22a]. case-control [Wu22c]. cash [GLLL23]. catastrophes [KG21, Kum24]. categorical [CSA23, Hol23, LZLL20, Oga20b, ZG23]. category [Zen23]. Cauchy [CRKC21, Ekh23, Fak23b, Far24b, NNS20, Vam20, YDHY24]. causal [LHS23, LTTL24, MWLS24, NK23]. Causality [Ünv21c, Tri24, Ünv21b]. causes [BP24, NSM21]. Cayley [ZYY20a]. CD4 [NSG22]. CDF [MMS23, SS24c]. CDF-quantile [SS24c]. CDO [YNC22]. ceiling [CM20]. cell [NSG22]. censored [AFO21, AZAZ24, ACCM24, BSN22, BBM24, Bou21a, BKS23b, CLF21, CH21b, DV22, EW24, EA23, ES22, FL21, GG24a, Han20, KM21, Koh20, KG22, Lem22, Leu21, LZ20a, LZG20b, LHS20, LC22b, LG22, LWS24, LZ23c, MBLY21, MSW23, MWLS24, ML21, ND21, OFH23, PKK20, PMMK24, QSP^{+}24, QG23, RG21a, SVR21, SD20, sS20b, sS22j, SLX22, SZ20b, Tri21, TPS20, WZTW20, WW22, Wan22a, WS20, XO21, YSS22a, ZBB24, ZG21, ZWW24]. censoring [AV20, AGHK20, AAM22, BS22a, BER23b, CG23, CG21, CGL21, DBBK21, Don23, FTTY22, JMSG24, MK20, MHS22, NBM24, PR21, RG23, STK24, hSsXfX24, Shi23b, TPM22, WJXF21, WYG24, WJC23, XLY^{+}20, YD22]. censorship [BSR22, GS24, Kha24, LKS22]. Central [BEhF24, GZ23b, Aug21, CDP20, DT21b, FM20, LTW24, OC22a, OC22b, OTL21, Onw21, SST22, XWH21, YXZ23]. centralized [ABT20]. ceramics [GCCJO^{+}21]. certain [Kia22, LM23a, Mir23, ZYS21]. CEV [CLC22, LLL22, RG23]. chain [LGLL21, MBLY21, Rao22, SS22d, eZN22]. chains [DSY21, DLZN23, Ekh23, GC22b, PX23, SYT21, SLF^{+}22, SXZ24, Xin23, YXZ23, dZySgYW21, ZSYM21, ZYY20a]. challenges [KPH^{+}24]. Change [ZL22d, ADW22, ARK22, AAGE23, CBLS21, CX22a, GX21, GD21, IW22, IM21, JL22b, MM20, Mef23, NS21a, PCZ21, Sal24, SS22i, SA22, Skr23, Ver20, WWZ20, WZA21, WLZ^{+}23b, Yan22b, YK20, YC23]. change-point [CBLS21, CX22a, GX21, IW22, Sal24, WWZ20, Yan22b, YC23]. changepoint [SSM21]. changes [CW20b, QH22, ZW20]. changing [LC22a]. channel [EIAZ23b, SS22a]. character [LHSK20]. characteristic [Bah24, Gau21a, HXCW22, SC22, YD24b]. characteristics [EM21, HCH20]. Characterization [Rao22, RG23, GBGPNR23, KB22b, LNWA22, UET24, ZL22c]. characterizations [AMH22, BHB22, LZH20, NSN20]. Charalambides [Cha23c]. Charlier [Kwo22]. Chart [NSM21, AKST22, ACK23, AAAJ24, ARA^{+}20, ARAJ22, AKSL22, BKK22, BS21b, Bro22, CKL23, Che23b, CKKY22, CN22, Haq20, HWW24, KR24b, KGG21, LC24, MG23, MSEB23, MSY23, NMF23, NKK^{+}24, uAK22, NuAMA22, Sha21, TH23, ZLZ20b, ZYZ24, vdHvDZ22]. charting [MMG24, PP22b]. charts [AAGN24, ACK22, ACKL23, Asl23, DPP24, GN24, HA22, Haq22, HS24b, HCL21, KRCS24, uATH21, PP22b, PW23, QSCH22, SMMCM20, Tri21]. Chebyshev [Oga20d, Oga20e, Oga21]. checker [CMMD20]. checker-type [CMMD20]. checking [DY21, IKT22, Sun23]. checkpointing [ZOD20]. Chen [KTDW20]. Chi [Del21, DPP24, HZ21, ST23a]. chi-sequences [ST23a]. Chi-square [Del21, DPP24]. chi-squared [HZ21]. China [Ünv21c]. Choice [MMLBP20, Gri24, Liu22, MNAM20, RFVR20]. choices [HT23, KASK24]. choose [Sal24]. choosing [CGLP24]. Chover [DZ24, Din24a]. Chover-type [Din24a]. CIR [LJ24, WC20c, YLL^{+}24]. circular [ATJ^{+}24, BAGR22, CK21, Cha23e, HNR^{+}23, HAA^{+}20, HGAJ21, JBR^{+}24, KR21, KP21, Mah24, QZ22, RASD23, RNA^{+}22, SYG21, SX23, Ten24]. circular-circular [Mah24]. claim [AMH22, AD21, ADIK22, ATB24, CW20a, FZD24, FW23, FLW24, KWC21, LYC20, NZZ20, XWW24]. claim-number [LYC20]. claims [CW20a, CWCY23, CX22b, GLK24, Li24, Lin21, LGL20, LHXZ22, RS24, SYL23, SGW23, WQZL21, WQSG22, ZLZ20a]. class [AMH22, AAO21, BMT20, BGM22, CB22, CFS20, CDL24b, CW20b, Dij23, DLZN23, DMD23, Fan23, HSG21, KH20c, LCWZ21, MM20, Mir23, PH24, Pri24, SHFD23, SGN22, SGM22, SKCE23, SGTG24, SZ20a, TZ22b, VdWMB23, Xia20, XY23, YWS22, YZ23, ZK21, ZC23, ZYZ20, dZySgYW21, ZZGL22, ZSYM21]. classes [AAO21, BAGR22, BK22, BKS23a, RNA^{+}22, ZSYD24]. Classical [CVDY21, SCGB24, TT23, BZLW24, CCG22, CG23, Koh20, MSM20, RSY21]. Classification [LX22c, LLX21, LVB23, MC23, MBSS23]. classifications [SKRR23]. classifier [Ish20]. Classifying [BD24a]. classroom [And20]. clause [YCWZ22]. clauses [CH22, NCC24]. Clayton [FL20b]. clinical [IY23, sK20c, QLZZ20, SCP23, WBM^{+}23, ZWCP22]. Clipped [Fli20]. close [HWW24]. closed [Fro23, SS24c]. closeness [VK23b]. CLT [CDL24b]. cluster [CSA23, FS22, KRS20, LTTL24, QH21, SPS23, SGTG24, SS21b, TBCB24, WZA20, YSS22a, YT20, YHA20, ZZY^{+}23]. cluster-based [YHA20]. clustered [CSA23, CY23, JC20, New23, QH21, Wu20, Wu21, Wu22c, YSS22a, Yua21]. Clustering [BFSN22, Tap22, DR22, FS22, Mod23, ÖB21, TZZ24, WWX22, ZY24a]. clusters [NAJ22a]. CMC [JPJB22]. CO [KKNN23, KS24d]. CO-inertia [KKNN23]. co-quantiles [KS24d]. Coarea [Neg24]. Cochran [RJ23]. Coefficient [PF20, YPXZ21, AKSL22, CKKY22, DLT21, GP20, HC22, HCZ20, HWZX23, JZKH22, KLK23, LXZ22, LSY22, LW21b, Mah24, Mao20, MC23, ME21, Naj22b, O’G23, Ohy20, Ohy21, Par23, RK24, RAS^{+}22, RMS^{+}24, RuAD22, SU22, SLYF23, TZYY22, TWZ22, WZD21, WL22a, WWZ22, WL23a, XZZ22, XL22, YMG24, YH24, ZB23b, ZCT^{+}20, ZYZ20, ZGC22, ZCZJ20, ZWFZ23]. coefficients [DK24b, EAMH22, HN20, LSSW22, LW22b, LW22c, MWW23, MW24a, MB21b, TZ22b, WC20a, YL23, Yos21, ZY24a, ZT21]. Coffin [WT20]. cognition [ASG24]. coherent [FAN21, Kel22, KS21, RG24, YW22]. cohort [LWS24, Nak21, PLSW23, Sha22a, XZLB24]. coincidence [KT24]. cointegrated [Kur22]. cold [RG21b, RG24, TT23]. collapsed [SYT24, SG22b, SG22a]. collateralized [YN24]. collect [YSS22b]. collective [CMO^{+}23]. collisions [Con24, LAAA^{+}22]. Colored [Lou23]. column [Kim21]. COM [LZH20]. combination [AT22, Hom21, LT24]. combinations [AMBK22, EG23, LY23]. combined [CWM23, JZ24, KR24b]. Combining [SHC20, ATAAM20, DR22, HY23, KKS24b, MKÖ23, YCW21]. Comments [vdHvDZ22]. Common [XP21, BB24, CPS21, Dur22, JT24, KLK23, KTP22, KJKC20, Li23b, Mur22, NT20b, SHXW23, ZC21]. Communications [KS22, Ano23b, AE23a, AE23b, AE23c, AE23d, Cha23c]. commutative [LXZ20]. Commutativity [AOGM24]. compact [AGR21]. company [WRZW21]. comparative [ÖA23, PB22]. compare [HM21]. Comparing [ASKB23, CP22, DJW^{+}20, BS22b, CAAARCTS23, CMM24, GC22c, JKD20, WZA21, Wu22d]. Comparison [AOS24, BR22, Bia20, JAJ21, VK23b, Wu22c, ZL23a, ACMC21, CCG22, CKM23a, DK24a, GB23a, HS20, Ima20, KAC22, iK22, KK23c, Kur24, MÁH22, QR24, RKS22, ROSL21, Shi23a, Wu23c, Wu23d, Wu24a, Yan23]. Comparisons [BD21, sK23a, BmAB22, BKB23, BFPM22, BD24b, DDL20, FZD24, GBH23, KN22, Kel22, LZW24, MBR22, NT22, PK22c, PZY24, SBB22, SPZ23, SPZ24, WF23, WZ21, ZLF24]. compartment [LK24]. compartmental [BMK21]. compensate [SK20f]. Competing [BP24, AV20, AGHK20, BSN22, DV22, HL22a, KM24a, QG23, SK20a, VE21, WYG24, WLZ^{+}23b, ZLNS22]. competition [WH23a]. competitive [NCC24, XL24, Yan24]. competitor [Cha23f]. complaints [AMAA^{+}22, SY21]. complementary [DZZ23a, ZZZ24]. Complete [CCW^{+}20, CW22, CLS24, CW24a, CLW^{+}22, Din24b, FWWH21, GWD24, HW23b, JW24a, JW24b, LW20, LF20, LW21a, LM22, LW22b, LM23b, LYZ^{+}23, MWW22, Qu22, SW24, TZ22b, WW23a, WS23, WW24, XY21, ZHQ21, ZYC23, CW24b, Han20, HYWZ24, JKD20, Ko20, Ko23b, MLZ21, dXY23, Yam24]. completion [KM24a]. Complex [PD24, Arn23a, Arn23b, CMP21, DTDB22, HAK23, JRB^{+}23, KL22a, LYL22, NTDB23, SB20, Zal22]. Complex-valued [PD24]. complier [MWLS24]. Component [YW22, BR23, CLWZ22, CWK23, Che24a, Dag22, Dur22, EIAZ23b, FAN21, GB23b, HKL21, HWW24, KHF22, RG24, Sul22, YZZ21, YLL21]. components [BD21, BmAB22, BKB22, BKHB23, BAHB23, CB24a, CB24b, CKM23a, DT24, FMHP24, GBH23, HL22b, KN22, Kel22, Kob24, LX21, LNXC24, Özk21, PD23, PJB24, RG21b, SBB22, SHB23, SGZA23, SPZ23, SPZ24, WC22, WLLB24, WLZ^{+}24, YW22, YXZ21, Zal22, vdMGSM21]. Composite [JC20, AB24, FM20, LA23, OC22a, OC22b, ON23, OYO24, OYA24, Onw21, WJXF21, WJS22, Wan23a, WC24, WTT20, YO23]. compositional [Ale23]. Compound [Han23b, RKAA21, CH20b, Che22, GML20, Han24, LS20a, LLWL21, LYL22, LL24a, LGL20, LG24b]. compounded [SK20a]. comprising [SHB23, WLLB24]. Compromised [AHS^{+}24]. Computation [SL24a, ZPL23]. Computational [Yan22a, BS21a, BP20, HM21, JC22, KF23]. computationally [SB22]. computer [WCX22, You24]. Computing [SV22, YF22]. concave [PAC20]. Concentration [Tan24, Wir23, YLW24]. concerning [HCCzW23]. concomitant [AS22a]. Concomitants [ABXE22, HABE24, Moh22a]. condition [BFSN22, BG22c, CH21a, CH22, GK24, LF23, Nes24, QZ22, dS22, dHGS21]. condition-based [BFSN22]. Conditional [AR24, LCW21, RB21, ATAAM20, AEBK20, BL24a, BBM24, BN24, CZL24, DCY20, GK20a, Goo24, HMZ20, HO23, HFJ21, KdXYC23, Kit21, LKS22, LKA^{+}23, Leu21, LZG20b, LNWA22, LMS20, MBA20, RT23, SF24a, Swa24, Tel23, XO21, YWL20]. conditionally [CX22b]. conditionals [SRA20]. conditioned [SNS20]. conditions [Don23, WWWZ24, YD22]. conducted [Ünv21b]. conference [DNP21]. Confidence [AT22, BMG22, Bic20, Bic22a, HL24a, KR24a, LJLW23, MZ23, SO22, SXS24, CXT^{+}22, HJ22, JCY22, KERAE23, LL20, LK22c, NA23, New23, ÖA23, RKS22, SSH23, sS20b, WV24, XXBD23, YP20, YCY22, Zha22]. Confidence-credible [SO22]. conforming [ABJR22]. conformity [KK20]. confounder [NK23]. Confounding [CZ23b, BHL21, CLZ24a, LLZ20, WLZ20, YCC23, ZYY20b]. confusion [Zen20a]. conjectures [CS23]. conjugate [ZZW^{+}24]. connecting [DT24]. connection [KÇ21a, LK22b]. connections [Gaj22]. Conover [KOTAG21]. consecutive [ZYNX20, TK24]. consecutive- [ZYNX20]. Conservative [McI23]. considerations [SPS23]. considering [sK22a, RL23, WIN20, YZC22, ZY20]. Consistency [Don23, Wan23c, You23b, ZHQ23, ABL23, BG22c, BSR22, BER23b, CB24b, DC22b, KH20d, LDZ^{+}22, LYZ^{+}23, Nko23, OIF21, SZGK21, dXY23, YZXY24, YLL^{+}24, You23a, ZLH20, ZHQ21, ZWWW24]. Consistent [FW22, Sir24, Sun23, HZ21, Li23b, NCC24, SJF21, XZR^{+}20, YCW21]. consistently [LCF20]. consisting [BKHB23]. constant [AFO21, BG22a, CW20a, FTTY22, Kha20, Kha22, WM21, Zha21a, ZLCW20]. constant-stress [AFO21, FTTY22]. Constrained [NSM21, ADD22, SJT21, TW20]. constraint [Das23, JWW22, LY22b, LMLS23, MN21b, OO20, TY21, TLL23, WX20, YH23]. constraints [ADW23, GK20b, HGY24, LZ21, Li22, Maa23, PF23, WC24, ZDW22]. construct [KL22a]. constructed [BBP22b, OK20]. Constructing [DZZ23a, SD20]. Construction [BG22a, DAS^{+}20, GPM23, Gzy21, HNR^{+}23, JZ24, KG20, KS21, LLSZ22, Swa24, WYP21, ZDY24, AN23, And20, Cha23e, CQT21, HAA^{+}20, Kha20, Kha22, QT22, SS22e, SXS24, Tsa22, WWJ23, ZLPZ24]. Consumption [ZS20a, CLC22, CLZL24, GK20b, HT23]. Consumption-leisure-investment [ZS20a]. contaminated [Sha21, Zey20]. content [PG22]. context [Ish20, MC23]. contiguous [TBCB24]. contingency [CN23b, Gzy21, SYT24]. contingent [ZLZ20a]. Continuous [DMR22, Hon20, BL23, BBP22b, BZ21, BM21, BD22, Che21, Dij23, DG21b, DH22, EM21, Fab24, FBNP23, GT22, KE20, LP20, LZLL20, MN21a, Mat21, MB21b, NSN20, RSG21, SSG21, SZW21, UET24, YG20, dZySgYW21]. continuous-state [dZySgYW21]. continuous-time [FBNP23, LP20, LZLL20, MB21b]. contoured [BPZ22, Shu21]. contours [YLC^{+}22]. contract [HW23a]. contraction [DG21b]. contracts [WDLW20]. contrast [PS24]. Contributions [VdWMB23]. Control [NSM21, TH23, AKST22, AAGN24, AWB21, ACK22, ACKL23, ACK23, AAAJ24, ARAJ22, Asl23, BKK22, BS21b, BLZ23, CKL23, Che23b, DPP24, GN24, HCL21, KBKK24, KR24b, KGG21, LC24, LZL20, MSEB23, MSY23, MAT24a, MDB^{+}21, NNS20, NMF23, NKK^{+}24, uATH21, uAK22, NuAMA22, OK20, PW23, Pen22, Sha21, Tri21, Wu22c, Wu23c, Wu24a, Yan22a, ZAL21, Zha21b, ZLZ20b, Zhu22a, Zhu22b, vdHvDZ22]. controlled [RK22, SMP^{+}24]. conventional [ASI^{+}24]. Convergence [HN20, HS21, LG24c, MWW23, MW24a, RN22, UN23, YGG24, Ass21, BM20a, CCW^{+}20, CW22, rChZhN^{+}24, CLS24, CW24a, CW24b, CLW^{+}22, CDMR24, DTDB22, Din24b, EAB20, EB21, EG20, FWWH21, FH22, GZ23a, GWD24, Gül23, GK24, HW23b, HYWZ24, JW24a, JW24b, KÇ21a, KAJ^{+}22, Ko20, Ko23b, Ko23c, KHZ24, LW20, LH22, LTW24, LF20, LMS20, LW21a, LM22, LDZ^{+}22, LW22b, LM23b, MWW22, Mot23, NTDB23, PCdP^{+}20, Qu22, RN20, Sav21, SW24, SZ20a, TTL20, TTLT23, TZ22b, TG24, WW20, WW23a, WS23, jWWDjW24, WW24, WWWZ24, Xin23, XY21, YY21, YGAK23, YC23, ZSY21, ZYC23, dS20, dS24a]. converging [KH20b]. conversion [BAGR22, JBR^{+}24]. converted [HNR^{+}23]. convex [BMH24, SMH24, UN23]. Convoluted [LSSW22]. convolution [Nah22, PD22, TP24b]. convolutions [AHO22]. Conway [FWX24, HK21, RTB20]. cookie [GRBL24]. cooperative [WLZ^{+}23a, Yan24, ZZZ22]. Coordinate [OFO^{+}21, Lee22b]. coordinates [HKNW20]. coordinatewise [Ko20, Ko23b, Ko23c, TTLT23]. Copula [AAGN24, LCNB21, Bah24, BAA24, BAE23, BBAA21, Bou21a, EEAA22, FZL24, FL20b, GS24, GM20b, KGS20, PRZ23, RSG21, VE21, WJS22, WF23, ZLF24, ZLNS22]. Copula-based [AAGN24, FL20b, GM20b, VE21, ZLNS22]. Copulas [CMMD20, BBP22b, GBGPNR23, KWC21, SS22c, She22, SH23, Swa24]. Cordeiro [AE23b, AE23c]. corrected [NNS20, OMYF21]. Correction [Ano20a, Ano20b, Ano21a, Ano23b, AE23a, AE23b, AE23c, AE23d, BAE23, Cha23c, KV23, Dün21b, GG24b, HZZA23, MG24, YZC22]. corrections [vdHvDZ22]. correlate [Sha22b]. Correlated [HP20, Han21b, Han24, BKS24, FVCdC23, FMZM22, Gau21b, GC24a, HM21, Kum24, LCNB21, LY22a, NSM21, PS21, Pri24, RSG21, SPSS20, SK22b, Tsi24, XZR^{+}20, ZZY^{+}23]. Correlation [LP20, AA23a, AS23b, AWL^{+}22, CSA23, CG20, DWZZ24, HW22b, He23, HC22, JZKH22, KLK23, LTW21, LZ23a, LZG20a, LB22, LXLL23, LCW21, LO23, Mah24, Mao20, Naj22b, Par23, RYZ24, SSH23, SP24b, SHXW23, YW24, YH24]. correlations [BH24, CYW21]. correlogram [BZ21, IKM21]. corresponding [AT22]. corrupted [ZLZZ24]. cosine [KK22, Yam22]. cost [BP20, BAJ20, KM24b, Y.24, ZZZ20]. cost-effective [BP20, BAJ20]. Costa [TH23]. count [Akd22, BA21, CM20, CW20b, DKA20, JB22b, NSG22, PK22b, RSG21, SSG21, SF24b, SS21b, TIIB22, WZA20, WSL23, dCBV21]. counter [GH22]. counterexamples [Hon20]. counterpart [QR24]. counterparty [ZQX23]. counting [CB22, CYB23, LS20a]. counts [HK21, LWY^{+}23, NSG22, SG23, Wan22a]. couple [CS23]. coupled [Lee21]. couplings [Wir23]. Covariance [Lem22, AdMSK23, AvdWvW24, BN22, Bul20, CZ20a, CL21b, GB21, HN21, IKT22, IKM21, KA20, LMS20, MSY23, MS23, NKK^{+}24, Pos24, QH23, QZ22, SX23, TZ23, TK20, Wic20, YK23, YKH20, Yu22c, ZB20, ZCCT22, ZC22]. covariances [DOV22]. Covariate [CG21, Zhu22a, CGW23, HHW21, LZLL20, PLSW23, Yua21]. Covariate-adjusted [Zhu22a]. covariates [BHL21, CLF21, CY23, JLG21, JS23a, KHD24, LHS20, LX22b, MLSS24, QMS22, TZYY22, TLLL23, WZD21, XW23, ZZY^{+}23, ZG23, ZZZ20, ZY24b, ZWFZ23]. coverages [ADIK22]. Cox [CDMR24, DMR22, DM22, FHZC23, JZWT20, LS23, SYRV20, sS22j, SE21, ZZZ20]. Cox-proportional [SYRV20]. CPE [YLW^{+}23]. crashes [ERAS22]. credibility [YW24, ZC21]. credible [And20, SO22]. credit [Ünv21a, Zen20a, ZQX23]. credit-linked [ZQX23]. crisis [Ünv21b]. criteria [Dün21b, KH20d, Kur24]. Criterion [CZ23b, Che20b, CP21, Dün21a, HCZ20, OFH23, WDH^{+}22, dRJO21]. critical [CZ21, Pos24]. cross [BZ21, Cha23e, KTMKS20, KHVBRm22, Lee22a, Oga20c, ST21, SMV23, WW23b, XS21, ZS24]. cross-correlogram [BZ21]. cross-data [Oga20c]. cross-efficiency [KHVBRm22]. cross-over [Cha23e]. cross-sectional [WW23b]. cross-validation [SMV23, XS21, ZS24]. crossed [SKRR23]. crosses [SAP20]. crossing [BHL21, ST23a, SLS21, WC24]. crossover [Gon21a]. crucial [ST24]. CSS [SYG21]. cubic [ZH24]. cumulants [Oga20b]. Cumulative [KB23a, NSG23, Nou22, RBD24, CG24b, FTAL23, GK20a, HM24, JS23b, Kun23, MP21b, Mat24b, Moh22a, Nou23, Par21, RSR24, sS20b, SKS24, Tah20, TT22, Tan22b, TP24a, Yu22b, ZHA20b]. cure [VRB22, YSS22a]. cured [LHS20]. current [FSS22, LWK20, LZW23, SPSS20]. current-status [LWK20]. curvature [EW23]. curve [HBOR21, KT24, LL20, SZ20b, WC24, Wu20, WWX22, WjWbK22, ZZ23]. curves [GDSJ22, LKA^{+}23, Nie20, ROSL21]. customer [AMAA^{+}22, BAY20a, KLSZ22]. customers [KG23, KSK^{+}23, LBG23, LXWY21, SK22b, TZ24]. CUSUM [BS21b, Bro22, HS24b, Ver20]. cut [AT22, LWYJ24, SCAK23]. cut-point [AT22, SCAK23]. cutoff [HM22, Syr23]. cyber [FL20a]. cycle [MNAM20, ZS20a]. Cyclic [NK23, OK20, RASD23]. cycling [WT20].

D [AE23c, CVDY21, SACQ23, KM23b]. da-Silva [AE23b]. Dagum [FCCD20]. Dalenius [Arn23a]. Darling [MKM24]. Data [HJ22, SY21, AMM20, ANO23a, ASG24, AFO21, AMH22, ACCM24, Ale23, AAKL21, AO22, AA22a, ARK22, AWN21, BHL21, BY24, BCA22, BNO22, BBM24, BG22c, BA24a, BA24b, BGT23, BS22b, BM21, Bou21a, BS20b, BKS23b, Bul20, Bul23, CSZ20, CBLS21, CWJC20, CMP21, CSA23, CLPS24, CP24b, CLF21, Che23c, CY23, CY24, CZX24, CH21b, CS22, Dag23, DV22, DG24, DWZZ24, Don23, DL22, DD23, EM21, EIAZ23b, EIAZ23a, EAMH22, ES22, FVCdC23, FZ22a, FM24, FP23, FL21, FSS22, FZ22b, FS22, Fus22, Gai22, GL20, GÖ24, GG24a, GRBL24, GZ21, GdNM20, HMZ20, HP20, HKNW20, HYZ24, Hol23, HL22a, HX22, HWW24, HS20, HN21, IHDR20, bJwWyX23, JC20, JWWW24, JB22b, JRB^{+}23, KNCK24, KERAE23, KM23a, KH20a, KM21, Ker20, KM24b, KAJ^{+}22, KL22a, sK20d, KCG21, KX23]. data [LKS22, LS22a, Leu21, Li21b, LXZ22, LLQ22, LZG20b, nLSZ24, LTR24, LVB23, LCNB21, LY22a, LHS20, LC22b, LG22, LA23, LZW23, Liu24b, LCW21, LZ23c, MWLS24, MN21a, MPM20, MAAZ22, Mor24, NSM21, NSG22, Naj21, Naj22b, NMT22, New23, Nie20, Oga20b, Oga20c, OCSV23, OFH23, Özk21, Pho24, PMMK24, QSP^{+}24, QG23, QG24, QMS22, RZ23, RMS^{+}24, RFVR20, RLQ23, RK20, SBC23, SSH23, SVR21, sS20b, sSL20, sS22j, SLX22, SWHW22, Shu21, SU22, SJP23, Sla22, Sla23, SZ20b, SXS24, STC23, SF24c, SL24b, SS21b, TIIB22, TZ23, TZZ24, Ten24, Tri21, VCOS24, VRB22, WZTW20, WW22, WWZ22, WJS22, Wan22b, WYW24a, WSL23, WG20, WZ21, WS20, Wu20, Wu21, Wu22c, WWX22, WjWbK22, WW23b, XP21, XZZ22, XL22, XLY^{+}20, XO21, XFW22, XH24, YM24, YZC22, YSS22a, YXS^{+}24, YLL^{+}24, YO23, Yao22, YG20, YSS22b]. data [YKH20, Yu22b, YCY22, YWL20, ZAF24, ZBB24, Zey20, ZL21, ZJF23, ZLZZ24, ZK24, ZZY^{+}23, ZG23, ZWW24, ZZZ20, ZZZ22, ZSCL22, ZLNS22, ZY24b]. data* [MBLB24]. data-driven [WZ21]. dataset [BA24c]. DATE [ZLLZ23]. DC [GKvZB21, MZR20, NCC24, WCY23, YCWZ22, ZDW22]. DDRCINAR [LWD^{+}21]. DEA [ÜFR21]. debiased [MZ23]. debt [YN24]. debugging [DKD22]. decision [KTMKS20, KK23b, LCYL24, RL23]. decision-making [KTMKS20, KK23b, RL23]. Decomposition [SYT24, LLXW24, VRB22]. Deconvolution [PT22, BL20, Ben22a]. decreasing [KM21]. deduction [Zen22]. Deep [Bic22b, APV24, Bic22c, YLLZ24]. default [GQW21, Ünv21a, YMZD23, ZZRH22, ZQX23]. defaultable [DLHZ21]. defense [NLL22]. deferred [EBG22, JB22a]. deficit [Psa22]. defined [BS20b, Rao22]. definite [Kwo22, Pos24]. definition [KTMKS20]. Definitions [PZY24]. definitive [ON23]. degeneracy [Lee22a]. degradation [He23, HCLW22, LZW24, LGX^{+}22, Pra21, SS23b, SS24a, WLZ^{+}23b, YHH24, ZXLT21]. degraded [DYC^{+}24]. degree [AOS24, DSY21, KM24a, LQWW21, LLWA22, LMZ23, NSN20, PY20, WZL^{+}22b]. Degrees [Del21, Oga24]. delay [HGY24, Li23b, ZW24a]. delayed [CF22a, DH22, KC20, LP20, Lin21, LHXZ22, PAAA20, SB23a, ZYY20a, ZPJY24]. delta [BN24]. demands [LLXW24]. Deming [BMR^{+}23]. Dempster [HWNN23, KTMKS20, LCYL24]. dense [YN21b]. densities [HM21, PAC20, TPV20]. density [AP24, Ass21, AGB21, BR22, BL24a, Ber23a, BS20b, CDL24a, DT21a, EW24, EN20, GS24, Ham23, HZZA23, KC23, LKS22, LKA^{+}23, Lee22a, LWT24, LZG20b, LZ20b, LX22c, LDZ^{+}22, Mod23, NCNM23, PZAH22, RT23, Shi23b, SF24a, TQ24, Ten24, TPV20, WHMA22, WG20, WC20c, XO21, Xu22, XH24, You23b, YWL20, ZCCT22]. department [DG21b]. departmentalized [ONM^{+}24]. departure [NGC22]. dependence [Ass21, CH20b, CZLY22, FZL24, FW23, FLW24, GLK24, Hon20, JPZ23, JPJB22, LG24b, MGK23, MBLB24, Raï24, RM21, WQZL21, WW23b, XP21, ZLW21, ZZZ22, ZSNX23, ZPJY24]. dependency [BABF22, Zal22]. dependent [ABL23, AO22, ASKB23, AA22b, ABJR22, BSN22, BAA24, BR23, BBAA21, BmAB22, BKHB23, BP24, BA24c, CCW^{+}20, CW22, rChZhN^{+}24, CLS24, CY20, CG21, CMMD20, qDSzW21, Din24b, Din24a, EG23, FWWH21, GZ23a, GWD24, GLW24, GBH23, GZ23b, HY23, HW23b, HCCzW23, JB22a, JW24b, JPZ23, KOTAG21, Ko23b, LKS22, Lef22, LPFL23, Li23b, Li24, LNXC24, LZG20b, LZW24, LGL20, LLL23, Lou20, LW21a, LM22, LW22b, LM23b, MWW22, MHS22, NBM22, NP23, NZZ20, New23, PKP23, PG22, RAS^{+}22, RT23, RTN24, SVR21, ST23a, SGZA23, Sho24, SZ20a, TTL20, Tan24, TT23, VK23a, WC20b, WM21, WW23a, WCY23, jWWDjW24, WW24, WGX21, WLLB24, Wu23b, XO21, XH24, YW22, YFW^{+}23, YNC22, ZX24, ZCL24, ZW22a, ZQH23, ZW24a, ZN24]. depending [ZL22a]. deposit [TLL23]. depth [KH21, NMF23, YLC^{+}22]. depth-based [NMF23]. derivative [LL20, YCS23]. derivatives [BD22, KCG21]. derived [Gzy22, TA22]. describing [FBNP23]. descriptive [KOTAG21]. Design [ARA^{+}20, KASK24, MMG24, NKK^{+}24, WWJ23, ASI^{+}24, Ben22a, BKS23b, CH22, FM20, HTP22, HW23a, KGG21, KRS20, sK24a, KRCS24, LC24, LHWK24, LLZ24, LB22, LGX^{+}22, NSM21, New23, Nov24, Oh23, OYO24, PK22a, Pop20, QSCH22, RB21, SPS23, SMP^{+}24, SZW21, TH23, WBM^{+}23, WJC23, You24, Zey20, ZWLF24, ZXYJ24, ZH24]. designed [Yu22b].

Designing [DYC^{+}24, NN21, JB22a, MJB24]. Designs [CZ23b, ATJ^{+}24, AYS21, Arn23a, BG22a, BD24b, CLZ24a, CK21, Cha23e, CLZ24b, CKKY22, DZZ23a, DAS^{+}20, DCT24, DMD23, FHHW20, Gon21a, GPM23, GC21, GC24a, GWM21, HNR^{+}23, HSV^{+}22, HA20, HAA^{+}20, HGAJ21, JBR^{+}24, JZ24, JZC24, KG20, KKS24b, Kha20, Kha22, Kim21, KBMD23, LY24, LLZ20, LLSZ22, LZ23a, LOL21, LGZ22, LWYL23, LEM24, LZ24, uATH21, OC22a, OC22b, ON23, OYO24, OYA24, Onw21, OK20, PM21, PS24, PPM21, PKK22, QT22, QLZZ20, RASD23, RNA^{+}22, RAA22, SMS21, SS22e, SS22f, SS23a, SSS23, SAP20, ST21, SS24a, SB20, SS21a, SG22b, SG22a, VPM^{+}24, VVHK24, WLZ20, YO23, ZWLF24, ZYC24, ZYY20b, ZZZ24]. Destructive [RK24, VRB22, ZY20]. detect [CTEB20, Dün21b]. Detecting [Hua23, KM23a, RM21, ZL22c, vdMGSM21]. detection [ARK22, BCA22, FL20a, HHW21, KSM^{+}21, LHJ21, NS21a, PK20a, QH22, SS22i, SSM21, Ver20, WWZ20, Wu24b, YZC22, ZLLS21]. determinant [BN22, Bul20]. Determination [WHW23, BNAK23, CMM20, sK20b, MAT24a, PK22a, RMS^{+}24]. determine [LJL^{+}21, Zha22]. Determining [TG24]. deterministic [CP21]. Developed [KK22]. Development [SSA23]. developments [HSV^{+}22]. deviation [BZLW24, qDSzW21, DLZN23, DH22, HCCzW23, LZ23b, ML23, SU22, WWWH20, dZySgYW21, ZSYM21]. Deviations [ZS23, BMP20, CZ21, DZZ23b, FW23, FLW24, GLK24, IKM21, LGL20, LD22, LCF20, WC20b, Wan23b, WTS24]. device [LS22b, LCNB21]. devices [BKB23, LHS22, WHW23, YSS22b]. DEWMA [ACK22, HA22]. diagnosis [KK22, SY21]. Diagnostic [AUQ22, FVCdC23, KR21, ACMC21, CMCA20, Wu22c, XL23b]. Diagnostics [Ema24, MO20, DE22, DKA20, GZ21, SCAK23]. diagrams [Bai21]. diallel [ST21]. difference [ARAK22, BMT20, Das23, EA21, JL24a, JCY22, MMZ20, Par23, PK20a, SSH23, SGN22, SLX22, WDW24, Wu22a, Wu23a, WK23, Yin24, ZHQ21, ZL23a]. difference-based [ARAK22, PK20a, Wu22a, Wu23a, WK23]. difference-sum [Par23]. difference-type [BMT20]. different [AYS21, AdMSK23, AAR22, BZLW24, BH24, DHH^{+}23, Han23a, HAA^{+}20, Kel22, uATH21, PKSS22, QLZZ20, RNA^{+}22]. differential [BMK21, BM20b, BBP22b, HGY24, MKG^{+}21, Rao23, STW23, Shi23a, WX20, YZSS23, ZT21]. Differentially [PH24, PY20, WZL^{+}22b]. difficulties [ZHS^{+}24]. Diffuse [PF23]. diffusion [APV24, CLZZ21, GKS21, HSWZ21, HW24, Lef22, LWT24, LSSW22, SYZS23, TWZ22, WM21, XWYZ23, YGG24, Yan20, YCW21, YD24a, YZXY24, YLL^{+}24]. diffusions [Sud22]. Dimension [GZZ21, ANO23a, AA23a, CL23, CG20, Ish20, SWYF20]. Dimension-free [GZZ21]. dimensional [AS23b, BZ23, BBM^{+}22, Bey23, BMP20, Bul20, Bul23, CTEB20, CY23, CY24, CY20, DWZZ24, FSB22, GL20, GZ23a, HWW24, HN21, HWNN23, IHDR20, INKS23, JBJ23, Li21b, Li22, Li24, LL24a, LVB23, LX22b, LC22b, LCW21, Maa23, Mao20, Naj21, Naj22b, OMYF21, Pen24, QH23, SE22, Shi23a, SX23, Tan22a, TZ22a, WC20a, WXW24, XZZ22, YSF22, Yam24, YY21, YZ20, YXZ21, Yu22c, ZLZZ24, ZS24, ZK24, ZWYL22, ZLLZ23, Zhu22b, ZW24b, ZC22]. Dini [Gül23]. Direct [QT22, ASA24, KSA23, SHC20]. directed [Fan23, LCWZ21, LMZ23, WWL22, WZL^{+}22b, XW23]. directional [TDLL20]. Dirichlet [Hom21, HZ21, MVLMA20]. disaster [Sha22b]. disasters [KG23]. discontinuities [ZMLH21]. discontinuous [XWY21]. discounted [MGK23]. discouraged [YTT23]. discovery [Kan20, PAdlPK22, Zhu22a, Zhu22b]. discrepancy [JBJ23, LWYL23, Mak24]. Discrete [PD22, AG23, ABRT23, CBLS21, CLMA20, Cha23d, Che23c, Che23d, Ekh23, GG23b, Jas21, JPJB22, JMM22, KU22, Lee23, LYC20, LHXZ22, MBLY21, NP23, Pui24, STW23, SAK21, TKP24, Vam20, VVS^{+}24, VK23a, Wan23b, Wei21, XZL20, Yam22, YG20, ZW23]. discrete-time [Ekh23, JPJB22, LYC20, NP23, Wan23b]. discretely [Wei23]. discretized [AO22, CDMR24]. Discriminant [HKNW20, MN23b, GvRS24, NC23, NN24]. discrimination [DD23]. discussion [ZIZ23]. discussions [CYB23]. disease [LWY^{+}23]. diseases [WZZ22]. disnormality [GMC^{+}23]. dispersed [NSH23]. Dispersion [BBCL24, BHB22, BLC22, Gar21, RO24, SMZ23, WSL23]. Dispersive [PKP23, BKHB23]. dissimilar [WXY21]. dissimilarity [LG24c]. distance [AWB21, Bul20, Cam24, CG20, Kim20, LCW21, PS21, YM24, dSO22]. distance-based [YM24]. distances [PZZ20, PCZ24]. distinct [DT24]. distorted [SS22c]. distortion [XXZ24, dXG20, ZC24, Zha24a, ZFZ24, ZY23]. distress [STC23]. Distributed [CX22a, KX23, AD21, Ang24, CKM23a, CGL21, FZD24, Gai22, GWM21, KJKC20, nLSZ24, MMZ20, MN23c, SBB22, SMOO23, TÖ23, ZSCL22]. Distribution [AFL15, Del21, HY23, KM24b, MMRMG20, TLRB21, AKST22, AFO21, Ahm21, AYS21, AKAB23, ATAAM20, AZAZ24, ABXE22, AAE22, ARGD21, Ang24, Ano23b, AKUT23, AIGB20, AAM22, AE23a, AE23c, AE23d, BQW22, BA21, Bro22, CSZ20, CSBM24, CLMA20, Che23c, CZZ^{+}23, CG23, CBM21, CRKC21, DV22, DGCL23, DBBK21, DY21, EAS20, FAN21, FHL^{+}24, FTTY22, FMM^{+}22, GCCJO^{+}21, Gau21b, Gau21a, GH18, GKV21, GG24b, GM21, HK20, Ham23, Han20, Han21a, HAK23, HS22, HXCW22, HL22b, HW22c, HMM20, INKS21, IAM21, IW22, JS21, JL22a, JWW24, JLG21, JA22, JB22b, JMM22, JB22c, JRM21, JL24b, JKP20, KLK21, KLK24, KAAH21, KS24b, KTDW20, KIA^{+}21, KO23a, Kha24, KAJ^{+}22, KJAZ24, KMH23, Koh20, KSY20, KH20c, KR23b, KR23a, KS24c, KJ24, LHSK20, LHS22, LC24, LL20, LSP22, LNWA22, LOT22, LG22, LL24b]. distribution [LNP20, LLS20, LY23, LLG24, MMG24, MPM20, MP21b, Mat24b, MPS23, Moh23, NBM24, Nah22, NSG23, NCGB24, NP23, NS21c, NS22a, NNS20, Neg24, Oga22, Oga24, OSX^{+}16, OTL21, OV21, Ory20, PS23a, PB24, PWLO17, PKSS22, PKK20, PG22, Pui24, QG23, RKS22, RKAA21, RSA20, RR23, RG21a, RBCG20, RYJ20, RTT23, RBSNB20, RTN24, RTB20, SG24, SAH21, SP24a, STK24, SS22g, SS22h, SP24b, SAH^{+}23, SCPAV21, SK20a, SNS20, SC20, Sil23, SK24c, Sul21, Sul22, SKP23, TIIB22, TH24, TZL24, TEHEG22, TS22, The22, TP24a, TPS20, TPM22, TK20, VCOB22, VVS^{+}24, WZL23, Wan23a, WYG24, WSL23, WLL^{+}20, Wic20, Wir23, WHW23, WJC23, Wyw24b, XJ21, XTXW20, XL23a, YT22, Yam24, YLJ23, YDHY24, YXS^{+}24, YS20, YSS22b, YZT21, ZAF24, ZCW23, ZG21, ZZW^{+}24, ZCZ20b, ZLS23b, dC22a]. Distribution-free [KM24b, MMRMG20, MMG24, SS22g]. distributional [HL24a, LXW20]. Distributions [Dey24, Ino24, Kon22, PF20, Vam20, AMH22, AO24, AAO21, AAR21, AP20, BPZ22, BM20a, Bal22, BG21, BHB22, BAHB23, Bay20b, Baz24a, Bic22a, Bic22c, BCG24, ÇA20, CP22, Cha21b, Cha22, Cha23c, Cha23d, CDSG20, DR21, Dij23, DZ20, DLC22, EK22, EBX20, Fab24, Fur23, GH22, GSH22, GDSJ22, GZ21, GG24b, GSED24, GC22c, GB23b, HBH24, Hom21, Jas21, KP24, KK20, KBMD23, KG22, KL20, LL23b, LQWW21, LMZ23, MvSK20, MS23, Mir23, MMS23, NT22, NA23, NSN20, NSH23, OMS24, PJB24, PT22, PZAH22, PZY24, RBS20, SSA23, SRA20, SLV^{+}20, SP23, SD20, SCPAV21, SJ23, Sil21, SK20e, SS24c, TH24, TA22, TKP22, TP24b, UET24, ÜFR21, VdWMB23, Wal23, WZTW20, Wu24a, Xia20, XM21, XFP23, YWS22, Yo22, YZ23, ZC23, ZCZJ20]. district [Tap22]. disturbance [Wyw24b]. disturbances [Aga24, PWM20]. divergence [AP24, BBCL24, CMP21, DE22, Gzy22, Kur24, NH20, Oga20b, RBD24]. divergence-based [Kur24]. divergences [Bou21a]. diverging [LS23, LW21b, ZZY^{+}23]. diverse [SSA23]. dividend [LL24a, LLL23, SYZS23, SS23c, TY21, XZ24, YCS20, Zha21a, ZDZ22]. dividends [CZLY22, LHXZ22, MB20, YD24a]. divisibility [She22]. divisible [Nah22, SS22f, Wal23]. DNA [AYP21]. Does [MDB^{+}21, Sha22b, TS22, MC20]. DOI [KS22, Cha23c]. dollar [Ünv21b]. dollar/TL [Ünv21b]. domain [YZ21]. dominance [EG23]. don’t [Lav24]. Double [PTYZ23, AuAH22, BAAJ20, BAJ20, rChZhN^{+}24, FMM^{+}22, GG24b, Hu22, KR24b, LMS22, MCL23, MSEB23, uATH21, RIA20, SP22, SK24c, WX20, ZL22a, ZCZJ20]. double-barrier [ZL22a]. double-parameter [ZCZJ20]. double-Pareto-like [FMM^{+}22]. double-sequential [Hu22]. Doubly [CDCC24, SDK^{+}24, BBM24, CM24, CH21b, PKK20, sS22j, ZT21]. down [YCY22]. drift [CDMR24, DMR22, DDM24, FE23, LSSW22, MO22]. driven [ABRT23, CLP23, CW20b, Rao23, Rao24, RTT23, Sha21, STW23, WSG21, WWC24, Wei21, Wei23, WZ21, XY23, Yu20, ZYS21, ZSL24, ZLZ20a]. dropout [KM24a]. drug [SZW21]. DSGE [Cav20]. Dual [BBP22a, Baz24b, CLZZ21, DHH^{+}23, FTTY22, HS24b, LLL23, LLH^{+}24, MLW24, SS23c, YZ21, ZCT^{+}20]. dual-domain [YZ21]. duality [CZ20b]. due [Nak21, NGC22]. duplicate [Ano24]. during [SCGB24]. Dutter [ZX24]. Dynamic [CM24, SKS24, CP23, DDG23, FP23, HM24, KB22b, Li22, MGK22, Moh22a, RLQ23, SR21, SN23, YZ21, ZM24].

E-Bayesian [PKSS22]. each [DKD22, HWH22]. earlier [vdHvDZ22]. easily [JL24b]. Ebaid [BAE23]. echelon [Oga20d]. Economic [CKKY22, FZ22b, JZWT20, LC24, TH23, Ünv21b, YL23]. economic-statistical [CKKY22, LC24]. economy [VCOS24, ZL22a]. EDCC [QQS^{+}22]. edges [HCX20]. Edgeworth [Aga24, SX23]. editor [GG24b]. Effect [RuAD22, Shu21, ACKL23, AvdWvW24, BH24, CW23, CDCC24, GG20, Han21a, sK23a, sK23b, LXZ22, MSEB23, Mor24, NK23, PKSS22, SV22, STC23, WH23a, WC20c, XZLB24, XL23b, YHA20, Yua21, Zha22]. effective [BAY20a, BP20, BAJ20, MRD20]. Effects [RAA22, BY24, BD24a, CM20, DAS^{+}20, DG24, Dur22, Gon21a, GPM23, HYZ24, LaM24, LHS23, LTTL24, LTY23, Liu24b, MWLS24, MO20, Rao23, SMS21, Sha21, SS24a, SL20, SL24b, TP24a, TPV20, Ünv21b, WD20, WSZ22, XLY^{+}20, YWL20, ZC21, ZJF23, ZXLT21]. efficacy [sK23b]. Efficiency [KHVBRm22, SS21a, Wu22a, Wu23a, CP24b, DM24, Fre20a, GPM23, LLZ24, Yo22, ZY24b]. Efficient [CH21b, DLC22, EW24, KSA23, uATH21, PHBV23, PPM21, ZAL21, ZSCL22, AU22, ATJ^{+}24, AW20, BKS23a, Gon21b, HX22, HFJ21, HGAJ21, KJAZ24, Li21b, LHS23, NS21c, OK20, SB22, SYRV20, SKP23, WJL23, YLW^{+}23, ZB23a, ZIZ23]. efficiently [JL24b]. eigenvalue [Ish20]. eigenvalues [MLWSPS24, Mon24, NSH23]. eigenvectors [Mon24]. EIV [ASW21]. elastic [ySLZ22, YY21]. Elbadawy [BAE23]. element [Pra21]. Elementary [Bia22, WC20b, BM20a, Bia20]. elements [GRBL24, JSB22]. elephant [BL24b, HF24]. elitist [PCdP^{+}20]. ellipsoidal [LY22b]. ellipsoids [Zen20b]. Elliptical [EA21, ALS21, DGCL23, EM22, PZY24, RBC^{+}23, EK22]. elliptically [BPZ22, Shu21, YWS22]. ELR [MSY23]. ELS [YZC22]. EM-based [MN23c]. embedding [Ekh23]. EMP [KK23c]. Empirical [Alv23, CL21b, Che22, GX21, HSWZ21, HL24b, IK21b, LQ21, LXZ22, LLQ22, Li23a, Qin21, RLQ23, sS20b, SRSB20, WL22a, ZYZ20, BC23a, Bic22a, DLC22, HXCW22, HL22a, HX22, IW22, LD22, LZW23, LWYJ24, Shi23b, TZYY22, TQ24, TLN24, Ver20, XL22, XFW22, ZCT^{+}20, ZZW^{+}24, ZWFZ23, vHP23]. emulation [WCX22]. encodings [Zen23]. END [CLW^{+}22, FH22, HN20, LZ20a, Qu22, XY21, ZHQ23]. endogenous [TZYY22]. endpoint [SMP^{+}24]. endpoints [BCH24, IY23, QLZZ20, SMP^{+}24, SZW21]. energy [LVB23]. engineering [Sho23]. enhance [RG24]. Enhanced [HS24b, Bic22b, HF24, LLX21]. enhancement [CG20, HC22]. enrichment [SMP^{+}24]. ensemble [NSH23, STC23]. entries [LMS20]. entropic [ZCZ20a]. entropies [CP23, Kit21, KB22b, Yin24]. Entropy [BTH23, CD24, AA22b, BKM20, Bic22b, FTAL23, Gzy21, HCL22, HHKL23, HCCD24, KTMKS20, KS20, LCYL24, Moh22a, NSG21, NSSR23, NSG23, Nou22, OIF21, PD24, Par21, RSR22, RSR24, SXZ24, SK24c, SKS24, Tah20, YD23]. environment [DSY21, DLZN23, Dur22, FMHP24, GZ23a, KHVBRm22, LH22, LTW24, TY21, WLLB24]. environmental [PD23, VVHK24]. environments [SLF^{+}22]. EPMCs [YSF22]. epsilon [HMM20]. equal [RK24, YXZ21]. equality [EAMH22, PP23, YKH20, ZCZJ20, ZC22]. equally [GRBL24]. equation [Mur22, Shi23a]. equations [BBP22b, CM24, HN23, JL24a, LX22a, Rao23, STW23, Spe24, ZT21]. equidispersion [dCBV21]. Equilibrium [GLLL23, LRWZ22, LXWY21, WCY23, XWY21, ZDZ22, ZLCW20, NSG23, XL24, ZW22b]. equipped [RG21b]. equireplicated [Kha22]. equivalence [Oga20c, Yan23]. Equivalent [SYT21, WWWZ24, HW20b]. equivariant [AAKL21]. ERGM [SS22i]. ergodic [AAKL21, BD22, SXZ24]. ergodicity [HFJ21]. Erich [Bal22]. Erlang [TKP22, Zha21a]. erosion [Pou21]. Errata [Bic24]. Error [WC20c, AALE22, Ano21b, BER23b, CWJC20, Cel22, Che23b, DLT21, DSB22, DKD22, DR21, DD23, EA21, EM22, Ema24, EGD23, GG20, GDK22, GZ20, HW22b, IK21b, KE24, KM24b, KS22, KSS21, LTW21, LY22a, LZ20b, LLG24, MLW24, MHZJ23, OO20, Ory20, PT22, PS23d, QG24, RuAD22, SSA23, Sir24, TGMS23, WLL^{+}20, XH23, XFW22, YS20, YRB20, YA22, YZT21, ZYC24, ZFZ24, ZWFZ23]. error-prone [EGD23]. Errors [NGC22, AB24, BA24a, BA24b, BKS24, Che23a, FZW23, FHHW20, FMZM22, GC24a, JWW24, Kim20, KR21, LAAA^{+}22, LY24, LMH21, LLQ22, Li23a, LL23a, MSEB23, MB21a, PS21, Pri24, RKS22, RL23, RLQ23, SS24b, SMMC21, SV22, SX23, TGMS23, Tsi24, WDW24, XW22, XFW22, ZX24, ZCL24, ZLH20, ZHQ21, ZHQ23, ZQH23, ZC24, Zha24a, ZFZ24, ZWWW24, ZGC22]. errors-in-variables [XFW22, ZGC22]. Essam [BAE23]. Esséen [CDL24b, Dou22, ZAF24, CL21a]. essential [ZLW21]. estimate [AMM20, BABF22, BL24b, CSA23, DCY20, EW23, HMZ20, Shi23b, WW23b, ZCCT22, Zha22, ZZZ20]. estimated [Bic23, Bic24, MSY23, ZYC24]. estimates [Bic20, CLLD23, DLC22, Gho23, JPJB22, LMH24, MC20]. Estimating [AA23b, CW23, HWH22, JT24, JL24b, KP24, KE20, KM23b, QSP^{+}24, STK24, SSW20, TPM22, XXBD23, YXZ21, AS23a, BN21, CBM21, CDCC24, GPM23, HA20, KTDW20, LX23b, NS21c, NS22a, SSS20, SYG21, YZC22, ZHS^{+}24]. Estimation [AS22a, AYS21, AA22c, BY24, Ben22b, BZ20, DSB22, DZ20, FZ22a, FE23, FSS22, HSG21, HAK23, HDZL21, HW20a, Hua20, HSC22, IAM21, JS23a, JMM22, KO23a, Kim21, KHD24, LMH21, LL24b, MWLS24, MQK23, MN23a, MHZJ23, MO22, Nak21, PS23a, RYJ20, SHA24, SVR21, SJ23, SU22, TZ23, TZZ24, TA22, TLLL23, TK20, WWZ22, WYW24a, WLT21, YPXZ21, YCC23, YW24, YS20, YRB20, ZS23, ZCW23, ZCZ20a, ZP22, ZLS23a, ZZY^{+}23, ZMLH21, AV20, AR24, ANO23a, AB24, AMBK22, AGP21, ACKL23, AAHS21, AEBK20, AAR22, Ami23, AD21, AC23, AP24, ATB24, ABRT23, APV24, AW20, ASA24, AT22, ASI^{+}24, BR20, BM22, BR22, BL24a, BC23a, BGT23, Bic23, Bic24, BSR22, BB24, BS20b, BS24, CZ20a, CX22a, CVDY21, CRVK24, CS20, CL21a, CLP23, CZL24, CB24a, CB24b, CMMD20]. estimation [DT21a, Das23, DMR22, DBBK21, EW24, Ema20, EGD23, EN20, FZW23, FHZC23, FW22, GX21, GP20, GS24, Gon21b, GZ21, GZ20, GK24, Ham23, Han20, Han21a, Han23a, HS24a, HYZ24, HBOR21, HXCW22, INKS21, INKS23, JRCB23, Jan20, JJ20, bJwWyX23, JL22b, JB22c, JRM21, Jun20, KNCK24, KE24, KKS24a, Kha24, KM24b, KTP22, Kia22, Kim20, Koç21, Koh20, KC23, KGS20, KHK20, KSY20, KJKC20, Kun24, LS22a, LHSK20, LS22b, LHS22, Leu21, LYL22, LWT24, LZG20b, LZLL20, LWD^{+}21, LSSW22, LSY22, Liu22, LL24c, LM23a, LG24c, LMS20, LM20, LCWZ21, MA23, MZ23, Mat21, MVLMA20, MdWG23, Mod23, MRD20, MMS23, MHS22, NBM24, NT20b, Nko23, OIF21, OV21, Özk21, PM21, PH24, PHBV23, PKK20, PSLLN20, PS21, Pop20, PTYZ23, PVC^{+}23, PKT21, QH21, RK24, RSR22, RK21, Rao23, Rao24, RT23, RG22]. estimation [RG23, SBC23, SE22, SS22h, SANHS21, SA22, SLX22, hSsXfX24, SF24a, SPSS20, SK20f, SA21, SBB21, SV21, SS22k, SDK^{+}24, SBB24, Sir24, Sla22, SRSB20, SZ20b, SB23a, SL20, TZ22a, TY23, TS22, TÖ23, TDLL20, TPS20, TPV20, WWWH20, WY20, WZA21, Wan22b, WZZ22, Wan23a, WHMA22, Wei21, Wei23, WG20, XL23a, XZL20, XO21, Xu22, XH24, YT22, YP24, YGG24, YN21a, YK23, YXS^{+}24, YLL^{+}24, YT20, YA22, YHA20, YZ21, ZBB24, ZX24, Zen24, ZYS21, ZS24, ZK24, ZYZ20, ZCZ20b, ZG23, ZGCZ23, ZL22d, ZSNX23, Zhu22a, ZW24b, ZY24b, vRvR20]. estimations [LKA^{+}23, SS24b, SCGB24, TP24a, ZWT^{+}24]. estimatiors [Ass21]. estimator [AU22, ARAK22, AAUA23, AAQ23, BL20, BG22c, BM20b, BN22, CY23, CWK23, CY24, DR21, DC22b, Don23, EA23, EA22, Far24a, FF23, HO23, HX23, IW22, IKM21, JWW22, JWW24, JLG21, JZKH22, KdXYC23, KAJ^{+}22, KD24, KHZ24, LS22a, Lee22a, LZ20a, Li20, LS23, LDZ^{+}22, Mat23, MZ22, MKÖ23, MBLB24, NSS24, Oga20b, OHS21, OO20, PB22, QH21, RTT23, Ros22, SS22b, SSS22, SAH21, SBK24, SMV23, Sla23, SWW24, TG24, Ten24, WSG21, Wan23c, WDW24, WKS21, Wu20, Wu22a, Wu23a, WK23, WL23b, dXY23, YLJ23, You23a, YC23, ZAF24, Zei23, ZB23b, ZLH20, ZHQ21, ZC21, ZHQ23, ZQH23, ZWWW24, ZZGL22]. Estimators [HR23, AOGM24, AAHS21, ABL23, AAKL21, AHS^{+}24, AuAHS23, ASI^{+}24, BMT20, BK22, BKS23a, BKS24, BMP20, BZ21, BD22, BEhF24, Bul20, CCG22, ÇA20, CPS21, CDL24a, CDMR24, DC24, DDM24, DKA20, DM22, EA21, FM24, Fus22, GG24a, GG21, HZZA23, HZ22, Hu24, JMK24, JV24, KKS24a, KSA23, KJAZ24, Koç21, KK24, KHSS20, KCG21, KSS21, LKS22, Lem22, LB22, LZ20b, LYZ^{+}23, MN21b, MKÖ23, MNA22, uA21, Oga20c, Ory20, PTM21, PAAA20, Pri20, PKT21, Pri24, QMS22, RO23, Ril24, SGN22, SGM22, SANHS21, SKCE23, SSNP20, SYG21, SGTG24, SF24c, SBZ21, ÜK21, Wan22a, WSZ22, WTS24, YMG24, YZXY24, Yos23, You23b, YTT23, YD24c, ZB20, ZK21, ZB23a, ZIZ23, ZAG24, ZL23a, ZC23, ZZW^{+}24]. etc [Bic20]. etimate [BBM24]. Euclidean [Cam24]. European [RN22]. EV [MZ22, ZYW23]. evaluating [OYO24]. Evaluation [AFG20, MJB24, AA22a, Baz24b, BKS24, KHVBRm22, MiT24, Nov24, Oh23, PTM21, PLLC23, SGZA23, XWYZ23]. event [DL22, sK23b, NSG22]. events [CG21, KM24a, KRCS24, Yua21, dSO22]. evidence [HHLG22, Lav24, TG24]. evidential [Bic22a]. evolution [FBNP23]. Evolutionary [MB21b, MKG^{+}21]. Ewens [Wir23]. EWMA [AN23, AAGN24, ACK22, ARA^{+}20, Asl23, AKSL22, GN24, Haq20, HA22, QSCH22]. Exact [GZ23a, GT22, GWM21, HW22a, LTW24, Ory20, SP24b, YDHY24, MAT24a, MAK20, Ten24, ZHS^{+}24]. example [AAO21, Bar22]. examples [GH22]. excess [CMM20, KL22b, LRWZ22, MLSS24, NSG22, TIIB22, YCW21]. excess-of-loss [KL22b, LRWZ22]. excessive [PK22b]. Exchange [QQS^{+}22, CL21c, GLB22, HW24, LJ24, XWY21, YMZD23]. exchangeable [Elm22, Ino24, NJMM24]. exciting [Li23a, SZL21]. excluding [TBCB24]. exclusion [CH22]. exhibiting [MM20, Vid22]. Existence [SY23, BG22a]. exit [HW22a, ST23a]. exogeneity [Kur22]. exogenous [SWLZ23]. exotic [ZAL21]. expansion [Aga24, LYL22, SSW20, SX23, XZ24, Yam24, You24, ZOD20]. expansions [LXW20, LLG24, Ril24]. expansive [PYZS23]. EXPAR [BM22, YM22]. expect [BCKS22]. Expectation [AR24, WZL23, Far24c, GA23, GZ23b, GLL23, Hu20a, HFGZ24, HW22d, IKT22, LNWA22, LZ23b, Mon24, Oga20a, Oga21, SJF21, SW21, YN24, Zha24b]. expectations [CW22, CW24a, CW24b, Din24b, DZ24, Far24b, FWWH21, FH22, HW23b, JKD20, JW24a, JW24b, LW20, LF20, LM22, LM23b, WW23a, WW24, Wu23b, ZW22a]. Expected [Kre24, MGK23, Pen22, CTEB20, Che21, Kia22, LM20, NS21d, NN24, QSCH22, Tap22, WGX21]. expectile [SZ24]. expecting [BCKS22]. expenditure [LLL22]. expense [LLL23]. Experimental [SS24a]. experiments [CLZ24b, GWM21, LZ23a, LLZ24, PM21, PS24, PPM21, Wan23c, WCX22, You24, Yu22b]. experts [PK22b]. expiry [LW23]. explanatory [SWK24, YPD23]. Explicit [Fak23a, Alg23, PZAH22, ZWT^{+}24]. explosive [KH20b]. Exponential [JS21, KAJ^{+}22, YP24, ZC24, AFO21, Alt21, AA23b, AIGB20, BHB22, BS20a, BS22a, ÇA20, CBB^{+}21, CMM24, DV22, DZ20, FH24, FHL^{+}24, GT22, GKV21, GSH22, GM21, Han20, HYZ24, JAE23, JT24, JMK24, JB22b, JRM21, KP24, KSS21, KS24c, KRCS24, Lem22, LBF24, LK22b, MCL23, ML21, MdWG23, MBA20, MAT24c, MMS23, Nak21, NGC22, Nie20, PR21, PKSS22, PKK20, PMMK24, QSCH22, RSA20, SMZ23, SANHS21, SC20, SDK^{+}24, TH23, TPS20, TPM22, ÜK21, WC22, WKS21, Wu23c, Wu23d, WJC23, Wu24a, Wu24b, XM21, ZK21, ZCZJ20, ZZ23]. exponential-Poisson [JRM21]. exponential-type [KSS21]. exponentiality [GM20a, IM21, ZHA20b]. Exponentially [KKS24a, CGL21, KJKC20, uA21, NuAMA22]. Exponentiated [AE23d, BBH20, GDSC15, PWLO17, ARGD21, AE23b, BD21, DK24a, FZD24, GSH22, LA23, SPZ23]. exponentiation [MW24b]. exposure [Ker20, Nak21]. expressing [WSL23]. expression [XH23, YD24b, YCY22]. expressions [Kwo22, ZWT^{+}24]. Extended [KWY20, MB21a, Mot24, SB22, BS21b, Bro22, CP24a, Che23c, Din24b, GWD24, HCLW22, HW23b, KAAH21, KS24c, LBF24, LW20, LPFL23, MWW22, ONM^{+}24, PVC^{+}23, RB21, SHB23, jWWDjW24, Wu23b, Zei23, ZW22a, ZQH23, dC22a]. Extending [FL20b]. extension [AA22b, BAE23, DM24, EEAA22, SP24a, TH23]. Extensions [FTAL23, MSM20, Tah20]. external [DYC^{+}24]. extremal [FF23]. Extreme [DK24b, WL23a, Yos21, ZCZ20b, CCG22, DT21b, Kun23, SP23, SAH^{+}23, TDLL20, WF23]. extremely [TH24]. extremes [AFFM23, BAA24, BBAA21, CP22, GdNM20, JL22a, KK23a, Mef23, PKP23, SG23, ZLF24]. extremum [Ros22]. extropies [BBL22, BG22b, SN23]. Extropy [CG24a, QE21, BG22b, CP24a, CG24b, GC24b, HM24, JS23b, Kun23, LX23a, Nou23, QR24, SR21, TT22, XD23, ZD23].

F [Ano23b]. factor [BZ23, CYW21, FHHW20, KL22b, LY24, Lig23, LJL^{+}21, ML23, TK24, WC20c, XL24, Yu22a]. Factorial [Cha23d, JZ24, KKS24b, LEM24, PPM21, WD20]. factorization [AK21]. factors [BH24, Che20b, CLZ24b, DZZ23a, Fow23, HN23, LGZ22, LJL^{+}21, PD23, PPM21, dRJO21]. Factorwise [BLC22]. facts [TW20]. fail [BAHB23]. fail-safe [BAHB23]. failed [SLS23]. Failure [HCX20, WLS23, APC23, BGM22, CP23, Che24a, CW23, CS22, HLPL22, HKL21, IM21, JA22, KM21, sK23b, LHWK24, LNXC24, LCNB21, LZW24, LHS20, LW23, MM20, NSG22, Nak21, SN23, SGZA23, SK20a, SAK21, TKP22, TP24b, WLZ^{+}23b, YSS22a, YXS^{+}24, YLLZ24, ZWW24, ZY24b]. failures [CF22a, Yan22a]. fallback [WZ21]. false [Kan20, PAdlPK22, Zhu22a, Zhu22b]. falsely [DT24]. families [Bal22, CMO^{+}23, Fab24, Fak23b, FH24, NT22, SCPAV21]. family [AO24, BP24, CBB^{+}21, CBM21, GSED24, HBH24, HW22c, HABE24, KM21, KSS21, LLL22, MS21, MMS23, PJB24, PZAH22, SANHS21, SD20, SYG21, SH23, TA22, ÜK21, WZTW20, Wu24b, YMG24, ZB23a]. Famoye [Ano23b]. Farlie [ABXE22]. fast [CX22a]. fat [TGMS23]. fatigue [YXS^{+}24]. Fault [LVB23, DKD22, YZC22]. feasibility [UN23]. feasible [WW23b]. Feature [DWZZ24, HLPL22, LX22b, LC22b]. features [LVB23]. feedback [KLSZ22, SK22b]. Fejér [BMH24]. Fejér-type [BMH24]. Feller [Bou22, NSJ22]. Feller-Jajte [Bou22]. fertility [MU22]. few [JV24]. FFT [LLWL21]. FGM [BAE23, CDSG20, EEAA22, WF23]. fidelity [WV24]. Fiducial [LK22c, LX22c]. field [BLZ23, WH23a]. fields [AH22, BA23, CWM23, DOV22, PP22a, SHA24, WS22, ZT24, ZSYM21]. filter [Lou23, SMZ23]. filtering [HN23, PF23]. filtration [SJF21]. filtration-consistent [SJF21]. finance [AP20]. financial [BPZ22, MBA20, NS21a, STC23, Ünv21b]. finding [Ima20]. Finite [Maa23, ASI^{+}24, BC23b, CS20, Che23c, CX22b, CGLP24, CYB23, Fre20b, HAK23, Jaf22, JPJB22, KE20, Lab23, LX22a, LK22c, Mot23, Mot24, NT22, NS22b, NCNM23, PS23a, QH22, RS24, SE22, SBB22, SGM22, SSNP20, SYG21, SPS21, SSZC22, TZ22a, WM21, WL23b, XZ24, XWW24, YW20, ZK21, ZCW23]. Finite-dimensional [Maa23]. finite-time [CX22b, JPJB22, SSZC22, WM21, XZ24, XWW24]. fires [CMO^{+}23]. First [Lef22, LW22c, ARF23, BM20b, CDP20, CC22, GL21, Gaj22, Gon21a, ST23a, UET24, ZCT^{+}20, ZSYD24]. First-order [LW22c, BM20b, UET24, ZCT^{+}20, ZSYD24]. First-passage [Lef22]. Firth [LWK20]. Fisher [CBJ20, CLLD23, LSP22, YFW^{+}23, YKW23]. Fisher-type [YFW^{+}23]. fit [ABT20, GW23, GSED24, HZ21, KR22, Le 21, LNWA22, Mir23, OFH23, Par21, RG21a, SH23, Wu22d, YG20]. fitted [Nie20]. Fitting [CZZ^{+}23, FVCdC23, Far23, GCCJO^{+}21, PAC20, WC24, ZHS^{+}24]. five [sK23a, LWYL23, SE21]. five-level [LWYL23]. five-parameter [SE21]. fixed [BD24a, DG24, Dur22, HYZ24, sK23a, LXZ22, RBC^{+}23, SMV23, XLY^{+}20, ZJF23]. flats [Tsa22]. Flexible [DT21a, SCPAV21, SS24c, DL22, KIA^{+}21, RBS20, WSL23]. floor [CM20]. flow [GLLL23]. fluid [LXWY21]. focus [ZIZ23]. follow [Nak21]. follow-up [Nak21]. following [LWKQ20, PKP23]. follows [Guo21, OV21]. force [CW20a, HT23]. forecast [AdMSK23, zWZ20]. Forecasting [WHCC23, BZ23, LGLL21, MBA20, Sha23, TPF20]. Foreign [LJ24, XWY21]. forest [CMO^{+}23, LLX21]. forests [Loe22]. form [Ahm21, DR21, Fro23, GQW21, ZQX23]. formation [AdMSK23]. forms [Ang24]. formula [Alg23, Ekh23, Far24c, LaM20, RN20, RN22, SW21, Vam20, Wij24, Neg24]. formulae [MP21b]. formulas [GB21, Oga20a, Oga22, TW20]. formulating [RBS20]. fortune [GRBL24]. forward [JL24a]. forward-backward [JL24a]. four [CK21, sK23a, LWYL23, TIIB22]. four- [LWYL23]. four-parameter [TIIB22]. Fourier [LYL22, TPV20]. Fourier-oscillating [TPV20]. Fraction [OYO24, YSS22a]. Fractional [KLS23, RBSNB20, ABRT23, APV24, CL21a, CZX24, DDM24, El 23, FE23, FTAL23, HL23, JZ24, LGLL21, Rao23, WSG21, XY23, YGG24, YZ20, YCS23, ZZ23]. fractionally [VSVVLRV22]. frailty [CBLS21, DDL20, HP20, Han21b, Han23b, Han24, ZWW24]. frailty-based [ZWW24]. frame [MLW24]. framework [CCMtH22, SS22i, SJF21, Wan21a, dXG20, Yan20, ZCZ20b, ZSNX23]. Frechet [SAH^{+}23, BD24a]. free [Cha21a, Fak23a, FM24, GZZ21, HCX20, KM24b, LC22b, MMRMG20, MMG24, SS22g]. free-replacement [Cha21a]. Freedom [Del21, Oga24]. frequencies [WjWbK22]. frequency [Lee20, Lee22a, MAAZ22, YLL^{+}24, You23b]. frequentist [CL21c, ZN24, SM20]. Frequentists [vdMGSM21]. fresh [Ang24]. Friedman [RJ23]. frontier [FP23]. Fukushima [Sha22b]. full [GPM23, HMP23]. function [ASS23a, AA23b, ASS23b, BN21, Bah24, BG22c, BA21, BD22, CPS21, CDL24a, CBM21, EAS20, FM24, FL20b, FMM^{+}22, GvRS24, Gau21a, GS24, GND^{+}23, HMZ20, Han21a, HAK23, HXCW22, IKT22, IKM21, KK22, KK23b, KB22a, Kha24, KAJ^{+}22, KHK20, Lab23, LL24b, MQK23, MN23a, MMLBP20, MZ23, MGK23, MP21b, Mat24b, MO22, MBSS23, NSSR23, Ory20, PS23a, PKK20, PRZ23, RSR22, RG22, RG21b, RG24, SBC23, sS20b, SLS21, Shi23b, SKS24, SSW20, TKP22, ÜK21, Wan23a, WS20, XO21, YLJ23, YS20, YD24b, YN24, ZAF24, ZXSZ20, ZZW^{+}24]. function-lognormal [Wan23a]. Functional [ASW21, BM23, WWX22, YPXZ21, AMM20, ANO23a, AEBK20, AAKL21, AALE22, AO22, BNO22, BABF22, BL20, Ben22a, BEhF24, CLPS24, CDP20, DE22, EIAZ23b, ES22, FZ22b, FS22, HKNW20, HWZX23, HL24b, JH24, Leu21, LY21, LTW21, LSY22, MHZJ23, MZ22, MBLB24, NMT22, Nes24, SBC23, Sla23, Sma21, Sma24, SF24c, TZ23, TZZ24, WjWbK22, XW22, YM24, YLL21, ZYZ20]. functionals [AO22, DOV22, XY23, YZ20]. Functions [LL23d, AGP21, AAR22, ADIK22, ASS24, BHL21, BQW22, Ber23a, BZ20, BZ21, Che23c, CDCC24, DKA20, Gzy22, Jaf22, JCY22, LC24, LX22c, MS21, MB21b, NSD22, NCNM23, PKSS22, PAAA20, PKK22, Pos24, PZAH22, Rao22, Ril24, SAF24, SST22, Swa24, TQ24, Tee22, TZ23, TP24a, XZ24, ZCCT22]. fundamental [TW20]. Further [CWK23, CW24b, HFGZ24, Nag20, OTL21, PK20b, SC22, Zha20a, Far24c, HSV^{+}22, LXZ20, SW21]. fused [OFO^{+}21, ZSCL22]. fusion [HHLG22]. future [AFO21, SJT21, VK23b]. FWER [Dey24, Zha21b].

G [AE23a, AE23c, AE23d, CKM23b, GC22a, MR21, Yan22a, BS21a, ZYNX20]. G.M [AE23b]. Galton [CZ21]. game [WX20]. games [HGY24, WH23a]. Gamma [AE23a, GH18, LA23, WLL^{+}20, Wu22b, AF23, AAQ23, GL21, HZZA23, JL24b, LLWL21, LY23, MN23a, SBK24, Wal23, WM22, ZZW^{+}24, XTXW20]. Gamma-Kumaraswamy [AE23a, GH18]. Gamma-Pareto [LA23]. GAMs [ORH22]. gap [CFS20, sSL20, XFP23, ZN24]. GARCH} [KB23b, AL24, GML20, KM23b, LK22a, QQS^{+}22]. GARCH-type [AL24]. GARMA [Ano23c, PS23c]. garrote [JBSK22]. gauge [MSEB23, RKS22, RL23]. Gauss [BA21, LY22b, WSZ22]. Gaussian [AAUA23, AUQ22, AD21, AL24, BB24, CLP23, CDL24b, DHJS23, DD23, EBX20, FSB22, Fu24, Fur23, GZZ21, HP20, JB22c, Kou23, LG24c, Maa23, Moh23, MO22, NA23, NAAQ22, OMS24, Rao24, RSG21, SST22, SB23b, TZZ24, WSSY20, WJS22, YHH24, Yu20, Zhu22a]. Gaver [VM22]. GBM [Bia22, ZWT^{+}24]. GEE [Akd22, CY23, CY24, ZWW24]. GEE-based [Akd22]. gene [XH23, YCY22]. General [CG24b, CZ23b, Sma21, WTT20, BY24, BPZ24, CPS21, CF22a, CLP23, CY24, EJJ23, FM24, GB21, GCFS23, GC24b, HBH24, JS23a, KR24a, KNL24, KH20c, Lee21, LF23, LY22b, LLG24, OA21, PJB24, Par21, Pri24, SP24b, SGM22, SMH22, SJF21, SKCE23, SJT21, TW20, WZTW20, WLZ20, WYJD21, Xin23, ZG21, ZYY20b, ZSNX23, ZCE23, ZX22]. generalisation [RR23]. Generalised [Ano23c]. Generalization [AK21, Far21, HMM20, KR23a, LHS23, RTB20, XJ21, YD23]. Generalizations [Kur24, MvSK20]. Generalized [ARAK22, BA24c, BB24, Cav24, CP23, Che21, CZ23b, Gaj22, Hol23, Nah22, PZY24, QH21, RAS^{+}22, RBCG20, Sha22a, TPF20, WKS21, Zha21b, dSO22, AF23, APC23, ALS21, ATJ^{+}24, AYS21, ABXE22, AAO21, Ary23, AIGB20, BEAA21, BP24, Bar24, BT22, Baz24b, BKM20, BC23a, BMH24, ÇA20, CL23, CWCY23, CZZ^{+}23, CDSG20, Dag22, DLHZ21, DY21, EAAC22, FH24, FMZM22, GK20a, GKV21, GCFS23, GA22, HCCD24, HDZL21, HPW24, JBR^{+}24, JMQ20, KWY20, KAAH21, KN22, KGG21, KMH23, KS20, LS22a, LLSZ22, LL23b, LD22, LG22, LKK24, MMRMG20, MMSC22, MMLBP20, MPM20, MN23c, Moh22a, Moh22b, MBA20, MDB^{+}21, uAK22, O’G23, OHS21, OFO^{+}21, OK20, PHBV23, PX23, PS23c, Pri20, PZAH22, RKS22, RSY21, SBC23, SE22, SS22b, SG24, STK24, SBB22, Sav21, Sem20, SMH24]. generalized [ySLZ22, SLF^{+}22, SXZ24, SYG21, SGTG24, SK24c, Tah20, The22, VCOB22, WHMA22, WLL^{+}20, Wu22a, Wu23a, WK23, XL23a, ZLS23a, ZSYD24, ZCZ20b, EK22, KKNN23]. Generalizing [BMR^{+}23]. Generally [HCL21]. generated [CB22, CW24a, Din24b, GZ23b, KG23, LZ23b, SP24a, SBHY24, TZ21, TZ22b, TZL24, VdWMB23, Zen20b, ZCCT22]. generating [BCG24, KB22a, KB23a, LW22a, SKS24]. generation [WWJ23]. generator [ZDY24]. generators [HGAJ21, LF23]. genes [XH23]. genetic [HY23, JZWT20, sK24a, PCdP^{+}20]. genomic [GÖ24]. Geo [GC22a]. geographical [AFG20]. geographically [PWM20]. Geometric [HFJ21, AZAZ24, BD21, Fro23, HCLW22, KG21, LHSK20, LW23, LLS20, RR23, Vog22, XTXW20, YP20]. geometry [LSP22]. geostatistical [RFVR20]. Gerber [SSW20]. GGM [DT24]. Ghosh [AE23a]. GI [LYY24, SB22]. GI/M [SB22]. Gibbs [Ami23, KU22, Lee22b]. Gilpin [ZSL24]. GINI [ARAJ22, KF23, LG24c, Yin24, ZLS23b]. give [MC20]. glass [GCCJO^{+}21]. Glivenko [Mot24]. GLM [RO23]. GLMs [ORH22]. global [VM22]. GMCOA [KKNN23]. GME [JRCB23]. GMM [Qin21, ZGCZ23]. Gomes [AE23b]. Gompertz [KAAH21, KCB23, SPZ24]. Gompertz- [SPZ24]. Goodness [ABT20, GSED24, Le 21, OFH23, RG21a, HZ21, LNWA22, Mir23, Par21, SH23, Wu22d]. goodness-of [SH23]. Goodness-of-fit [ABT20, GSED24, Le 21, OFH23, RG21a, HZ21, LNWA22, Mir23, Wu22d]. Google [Ünv21b]. government [WRZW21]. gradient [KHK20]. graduation [Yam20]. Gram [Kwo22]. Granger [Tri24]. Graph [MiT24, DWZZ24, Fan23, LLWA22, MAT24c, PH24, Wan21b]. graphic [Zen22]. Graphical [Oh23, DHJS23, GSED24, KS24a, Zhu22a]. graphs [BLC22, LD22]. gray [LGLL21, zWZ20]. Grey [LX23b, LW22a]. Grey-based [LX23b]. Group [Nov24, CF22b, CS23, GWM21, HW20a, HWH22, HKL21, JB22a, NP23, PM21, PP22b, Pen24, SS22f, SMP^{+}24, WZA21, ZIZ23, ZLW21]. group-arrival [NP23]. grouped [Wan22a]. grouping [DHJS23]. groups [HW20a, JKD20]. growth [SY23, UC21]. Grubbs [SC20]. guarantee [KJKC20, WDLW20]. Gumbel [ABXE22]. Gwet [Ohy21].

Hadamard [BMH24, LLSZ22, OA21, SMH22]. Haenszel [RJ23]. Half [SKP23, ARGD21, AH22, AE23c, GM21, OSX^{+}16, SPZ23, XJ21]. half-logistic [AE23c, OSX^{+}16, XJ21]. half-plane [AH22]. Hamedani [AE23a]. Hamming [PZZ20, PCZ24]. hand [CP24b]. handling [SBB24]. Hardy [GA23]. harmful [Moh22b]. harmonic [SMH22]. Harris [AKAB23, BAHB23]. Hartman [GLL23]. Hasse [Bai21]. Hastings [HFJ21]. having [BAHB23, CKM23a, HL22b]. Hawkes [FW23, LC20, Wan23b].

Hawkes-type [FW23]. hazard [BHL21, BBAA21, BKB22, CLWZ22, DM22, FAN21, HMZ20, Han24, JMK24, Lee23, LZ20a, LX22b, MA23, MZ23, NSG21, sS20b, SAF24, Swa24]. hazards [FSS22, Han23b, JS23a, MWLS24, PLSW23, SYRV20, Sha22a]. health [SY21]. heavy [CWCY23, Li23a, Lin20, Lin21, MPM20, PR21, TH24, ZC23]. heavy-tailed [CWCY23, Li23a, Lin20, Lin21, MPM20, TH24]. hedging [SHXW23]. height [DT21b, Gar21]. Heine [KV23]. Helmert [FC22]. hemodialysis [CW23]. Henderson [Yam20]. Hepatitis [PA21]. Hermite [BMH24, OA21, SMH22]. Hermite-Hadamard [SMH22]. Heston [Guo21, MZR20]. heterogeneity [sK22a, sK23b, sK24a, Liu22, LQWW21, LMZ23, ONM^{+}24, QG24]. Heterogeneous [ZW24b, APC23, BD21, BKB22, FCCD20, FZD24, HW23a, Hua23, sK20c, LWW24, MR21, PSLLN20, SHB23, SPZ23, SPZ24, YW22, ZLF24, ZWCP22]. Heteroscedastic [ADW23, GW23, LY24, Mef23, WSZ22]. heteroscedasticity [DC24, HS24a, MBA20, Tel23, Wu23c, Wu23d, Wu24a]. heteroscedasticity-adjusted [DC24]. Heteroskedastic [Cel22]. hidden [Gho23, Hua20, UE21]. Hierarchical [MN21a, PK22b, AC23, Che23e, Han23a, HX23, NMT22, PS23d, UE21, WG20]. High [AS23b, Bey23, Li22, SX23, YSF22, Yam24, ZLLZ23, AMM20, AHO22, BZ23, BBM^{+}22, Bul20, Bul23, CL23, CG20, CY23, CY24, DWZZ24, DT21b, FSB22, GL20, HWW24, HN21, HWNN23, IHDR20, INKS23, Ish20, Li21b, LC22b, LCW21, Mao20, Naj21, Naj22b, OMYF21, Pen24, QH23, RG23, SE22, Tan22a, TZ22a, WC20a, WYP21, WZL23, WXW24, WjWbK22, XZZ22, YY21, YLL^{+}24, YXZ21, Yu22c, ZDY24, ZLZZ24, ZK24, ZWYL22, Zhu22b, ZW24b, ZC22]. high-dimension [CG20, Ish20]. High-dimensional [AS23b, Bey23, Li22, SX23, YSF22, Yam24, ZLLZ23, BZ23, CY23, CY24, FSB22, HN21, HWNN23, INKS23, Li21b, Mao20, Naj21, Naj22b, OMYF21, Pen24, QH23, YXZ21, Yu22c, ZLZZ24, ZK24, ZWYL22, Zhu22b, ZW24b, ZC22]. high-frequency [YLL^{+}24]. high-order [WZL23]. high-performance [RG23]. Higher [LLG24, SSS22, YXP24, AOS24, El 23, MC20, NSN20, Ril24, YCS23]. higher-degree [NSN20]. Higher-order [LLG24, YXP24, El 23, Ril24]. Hilbert [TLRB21, Ko20, Ko23b, Ko23c, Ko23d, TTL20, TTLT23, UN23, WW20]. Hilbert-valued [TTLT23]. history [VK23a]. history-dependent [VK23a]. HIV [VRB22]. HNBUE [GM20a]. Hoeffding [Fro23, Gai22, Wir23]. Hogg [SSA23]. Hölder [GK24]. Holt [LW22a]. Holt-Winters [LW22a]. Homogeneity [QH23, RR20, KW22, LMLS23, WH23b]. homogeneous [FS22, KLS23, KH20a, Rao22, SG22a, WS22, Y.24, Yu22b, dZySgYW21]. homology [AA20]. homophily [LQWW21, LMZ23]. horizon [GLLL23, MLZ21, SSZC22, ZDZ22]. horseshoe [ZK24]. hospital [PLLC23]. Hotelling [Bul23, NKK^{+}24]. household [LLL22]. HR [BAA24]. Huber [YLJ23, ZX24, ZXYJ24]. Huber-Support [ZXYJ24]. human [HY23]. hurdle [MPM20]. Hurst [CL21a]. Hurwitz [LOT22]. Hybrid [JC22, MKG^{+}21, NuAMA22, PP22b, AFO21, AP24, APV24, ARK22, Koh20, KSK^{+}23, LG22, Mur22, PMMK24, QG23, STK24]. hydrology [KAJ^{+}22]. HYGARCH [BR20]. Hyper [KS24a, KR23a, KS24c]. hyper-Poisson [KR23a, KS24c]. hyperbolic [CLC22, XL23a]. hypercube [Onw21]. hypergeometric [BA21, KL20, LHS22, RTB20, SS22h, WZL23, YSS22b]. hypergraph [YN21b, Yua24]. hypersphere [NK22]. hypertension [KL22a]. hypotheses [CMP21, Lig23, Wij24, Zha21b, ZCE23]. Hypothesis [CZL23, KTP22, Zhu22b, Ano21b, Bic23, Bic24, CL23, CYW21, IKT22, KS22, LZ21, Nov24, PD22, Sma21, Sma24, SZW21, TG24, Yan22b, ZN24].

i.i.d [CC23]. ideal [CC23]. ideas [Ham24]. identical [HCH20, YZ23]. identically [Gai22, KMH23, MMZ20]. Identifiability [AvdWvW24, Wan22b, YD22, Don23, PSLLN20]. Identification [Özk21, XH23, GL20, Li22, LTTL24, PCZ21]. Identifying [Kou23, LY23]. identity [WZL23]. idiosyncratic [HT23]. ignorable [ASG24, WYW24a]. IHS [Gri24]. IHS-type [Gri24]. II [AFO21, ASW21, AZAZ24, AGHK20, AAM22, BS22a, CLF21, DV22, Gho23, JMSG24, KG22, MMG24, MNA22, MK20, MHS22, QG23, QLZZ20, RG21a, SMS21, SD20, SWHW22, TPM22, WYG24, YHA20, ZG21]. II/III/IIIb [QLZZ20]. III [QLZZ20, LaM20, The22]. IIIb [QLZZ20]. illiquid [Raï24]. illustrated [Sha22b]. illustration [Ham24]. Illustrations [KM24b]. IM [WJL23]. IM-based [WJL23]. image [EIAZ23b, EIAZ23a, LVB23, SMZ23]. imaging [NMT22]. Iman [KOTAG21]. imbalanced [ZSCL22]. immigration [CZ21, LH22, LTW24, RK22, SBHY24]. impact [HHW21]. Impacts [Ünv21c]. impatient [LBG23]. imperfect [BFSN22, CC22, FL20a]. implement [FL20b]. implementation [MMG24, SHC20]. Implementing [KNCK24]. Implication [ML23]. importance [Cha23a, CF22b, Loe22, YF22]. important [DZZ23a]. imprecise [Tan22c, Yao22]. Improved [BMT20, CWJC20, CBB^{+}21, CS22, DPP24, JMK24, LGLL21, MAAZ22, NT20b, PKK20, SV21, SJP23, ÜK21, YMG24, YSS22b, AW20, CXT^{+}22, Che23b, Fro23, JB22c, LLS20, MMSC22, MVLMA20, SGN22, SGM22, SPSS20, SPS21, zWZ20, YK23, ZLB20, ZPL23]. Improvement [KT21, KK23b, SH23]. Improvements [Oga21, SC22]. Improving [ZY24b, AAHS21, Koç21]. impulse [BZ20]. imputation [AHS^{+}24, Che23e, HHW21, SK20f, SPS21, SJP23, UMA22]. inaccuracy [BBCL24, GK20a, HM24, RSR24, SVR21]. Inactivity [MGC22, ASKB23, PK20b]. INAR [CH20b, CZL24, PA21, RBSNB20, SG23, Zen24, ZZK23]. incapability [BA24b, GG20]. incidence [ÜFR21]. including [BA21, YWS22]. Income [BP22, AG23, SSW20]. incomes [LLH^{+}24]. incomplete [BG22c, CS22, DG24, GPM23, KG20, Kha20, PPM21, ZK24]. Inconsistency [AL24]. inconsistent [ZS20a]. incorporated [XH23]. Incorporating [DHJS23, IW22]. increase [LMLS23]. increases [O’G23, Oh23]. increasing [Ano24, DOV22, Fan23, GC22b, JA22, Nie20, YJZZ23]. Incremental [ZXYJ24]. increments [BA23]. Independence [CG20, Gai22, GBGPNR23, GZ24, KX23, LCW21, Mao20, Naj21, NGC22, Yam24, ZW23]. independent [AV20, APC23, ATAAM20, BKB22, CW24a, CX22b, FCCD20, HM21, HY23, KMH23, LW20, PJB24, PP23, Rao22, Sla23, Tsi24, WSSY20, WJL23, ZL20, dC22a]. index [AT22, BNAK23, BBP22a, BA24b, CCG22, ES22, FF23, GG23a, GG20, HMZ20, HSG21, HL24b, JH24, KF23, LY22a, LL23c, LG24c, MB21a, Mat23, MHZJ23, MdWG23, MBLB24, NN21, SLYF23, SF24c, Syr23, WWZ22, WWJ23, WL23a, WYW24a, WJC23, ZG21, ZS24, ZLS23b]. indexed [DSY21, PX23, SYT21, SXZ24, ZYY20a, ZSYM21]. indexes [DDG23, GMC^{+}23, Ünv21c]. indicators [CCMtH22, WJXF21]. indices [BBCL24, Bia20, Bia22, KK23c, RKS22]. indirect [ASA24, KSA23, SHC20]. individual [HM22, HT23, sK23a, NZZ20, TH23]. individuals [BH24, CLZL24]. induced [CBLS21, ZZZ22]. industry [BAY20a]. inefficiency [FP23, ÜFR21]. Inequalities [Baz24a, BMH24, CH20a, DCY20, KdXYC23, LPFL23, Oga20d, Oga20e, Oga21, OA21, Pen24, SMH22, Wir23, YLW24]. Inequality [SBK24, AA22b, BC23b, Far24b, FSS22, Fro23, GA23, JWW22, LMLS23, Maa23, Tan24]. Inertia [GB23a, GB21, KKNN23]. Inertial [GN24]. infectious [WZZ22]. Inference [CGW23, Fan23, LL20, WZL22a, WjWbK22, ACMC21, Alm24, Ary23, BSN22, Ber23a, Bey23, CG23, CRKC21, CMCA20, FAN21, FL20b, FSB22, GT22, HCCM21, HX22, HWZX23, KLK22, KK21, KA22, Lee22b, LXZ22, LY22a, LG22, LTY23, MÁH22, ML21, Mor24, NSD22, NH20, NMT22, Ohy20, Ohy21, PR21, PLSW23, PKSS22, QG23, QG24, RSY21, RGS24, SLV^{+}20, SHFD23, ySLZ22, Sho24, SACQ23, Syr23, TT23, Tsi24, WWWH20, WYG24, WS20, XTXW20, XL22, YPD23, Yao22, YFW^{+}23, Yos21, Yu20, YLW^{+}23, ZG21, ZP22, ZYW23, ZWYL22, ZLLZ23, ZPL23]. Inferences [DZZ23b, CH21b, Ker20, MÁH22, TZYY22, WZD21, WM22]. inferential [CSBM24, Gho23]. inferiority [WJL23, ZWCP22]. Infinite [XZ24, AL24, CL21b, Che22, DSY21, DLZN23, MLZ21, She22, SLF^{+}22, WTS24]. infinitely [Nah22, Wal23]. inflated [ADD22, GdNM20, JB22b, KHD24, KR23a, MOAF23, ND21, Pho24, SSG21, TKP24, TLN24, XTXW20, ZBB24, vdHvDZ22]. inflation [GKvZB21]. Influence [PD23, BBM^{+}22, DE22, EM22, GZ21, MLSS24, Ril24, RG22, SL20]. influenced [KG21]. influences [sK20d]. Influential [Gos22, KM23a, Özk21]. informatics [Sho23]. Information [HABE24, KB22a, KK24, AAHS21, BT22, CCMtH22, Che20b, DHJS23, Dün21a, Dün21b, GS24, GKvZB21, Haq20, HA22, HDZL21, HHLG22, JTS22, JS23b, KR24a, KS24b, KE20, KB23a, iK21a, iK22, KGS20, LLZ20, uAK22, OFH23, PB24, Pen22, PC22, Pri20, QE21, SPS21, SV22, SDK^{+}24, SGTG24, WYJD21, ZHA20b]. informative [AKUT23, HSV^{+}22, KAC22, YSS22a]. informatively [ZWW24]. infusion [LK24]. Ingersoll [CDMR24, DMR22]. inhomogeneity [LTR24]. inhomogeneous [GL21]. injection [CZLY22]. INLA [RFVR20].

Inner [AR24, Zen20b]. Inner-Loop [AR24]. inner-product [Zen20b]. innovation [RBSNB20, XZLB24]. innovational [CLZ21]. innovations [CZL24, MN23c, ZB23b, ZZK23]. innovative [SCAK23]. input [WCX22]. inputs [SWLZ23]. Insensitivity [LB22]. insight [Asl23]. inspected [PS23b]. inspection [BA24c, JB22a, Lee20, ZY20, ZLZ20b, ZYZ24]. instrument [RK24]. instrumental [ZX23]. insurance [AMH22, AÜ20, CH22, GK20b, Lee21, LA23, Sho24, WQZL21, WRZW21, Yan20]. insurer [BPZ24, CP21, DLHZ21, LRWZ22, Pen22, RYZ24, YH23, ZZRH22, ZW24a]. insurers [CH21a, LZL20, LS20b, WRZL21, WDH^{+}22, XL24, YCW21, Yan24]. Integer [Sil21, Alm24, GZ23a, GML20, JL22b, LWY^{+}23, LW22c, YK20, ZCT^{+}20, dCBV21]. integer-valued [Alm24, GML20, JL22b, LWY^{+}23, LW22c, YK20, dCBV21]. Integral [PRZ23, CM24, CW24b, DS23, Ekh23, JAJ21, JW24a, LW20, LM22, LM23b, OA21, SMH22, WW23a, WW24, XY23, YZ20]. Integrated [BNAK23, NMT22, Ory20, Psa22, TWZ22, VSVVLRV22, YZXY24, YLL^{+}24, vHP23]. Integrating [SMZ23]. integration [YZ21]. intelligence [AdMSK23]. intensity [GND^{+}23, MCL23, QSP^{+}24, SL24a, XWY21, Yan22b]. inter [BH24, CY20]. inter-arrival [CY20]. inter-correlations [BH24]. interaction [Che24a, FMHP24, sK22a, sK23b, LHXZ22]. interactions [FBNP23, NS21a, WD20, WCT22]. interarrival [KV23]. interest [CW20a, Guo21, LJ24, LMS22, MCL23, SZL21, WM21, YMZD23, ZLZ20a]. interests [ZZRH22]. interim [sK20c]. intermediate [JA22]. intermittent [JC20, UMA22]. international [DNP21]. Interpoint [YM24]. interpretation [Wan21a]. intersection [YCS23]. Interval [Bic23, Bic24, Zey20, AT22, AKSL22, BCG24, CXT^{+}22, CLF21, CKKY22, FM24, KR24b, KTP22, KKKM21, KS20, LF23, Lig23, LHS20, LWS24, LZ23c, MWLS24, NA23, OFH23, ÖA23, QSP^{+}24, RKS22, RBS20, RK20, SS22g, SS22h, SS24c, SXS24, TH23, zWZ20, WJC23, YSS22a, YXS^{+}24, ZWW24, ZCE23]. interval-censored [CLF21, LHS20, LWS24, LZ23c, MWLS24, QSP^{+}24, YSS22a, ZWW24]. interval-valued [RK20]. intervals [And20, BEAA21, BMG22, Bic20, DG21a, HA22, HL24a, KR24a, KERAE23, KL20, LJLW23, LK22c, New23, ORH22, RBC^{+}23, SSH23, sS20b, SO22, Sud22, YP20, Zha22]. intervened [JS21]. intra [CSA23]. intra-cluster [CSA23]. intraclass [SSH23]. intravenous [LK24]. intrinsic [SJ23]. Introduction [DG21b]. invariably [KBMD23]. invariance [CCR21, HF24]. invariant [RSA20, Sav21, ZC23]. inventory [LLXW24, SLS23, Tas24]. inverse [AYS21, AAUA23, AZAZ24, AUQ22, CSZ20, CZ20a, Cha22, HP20, HZZA23, JB22c, KK23a, LP20, LLZ20, LA23, Moh23, NAAQ22, PKSS22, PH21, STK24, SB23b, YHH24, ZZW^{+}24, AD21]. Inverse-Gaussian [AD21]. inversions [KT24]. Inverted [Del21]. Investigating [KSS21]. investment [CZ20b, CP21, GK20b, GKS21, HGY24, KL22b, LRWZ22, LLL22, Li23b, LZL20, LLH^{+}24, LS20b, MZR20, MLZ21, Pen22, RYZ24, TLL23, WRZL21, WRZW21, WDH^{+}22, WH23a, XL24, YCWZ22, Yan20, YCW21, Yan24, YZW23, YH23, ZZRH22, ZW24a, ZS20a, ZDW22]. investment-consumption [GK20b]. investment-reinsurance [Li23b, XL24]. investments [LG24b]. involving [AA23b, LS20a, YS20]. IPLSL [Che23e]. IPLSQ [Che23e]. IRF [BZ21]. irregular [Ben22a, Pop20]. issue [DNP21, sK22a, sK23b]. issues [Gho23]. Itô [DS23]. Item [CM20, BH24, RK24]. items [Ano24, ABJR22, BH24, LWW24, SLS23, YJZZ23]. iterated [ABXE22, DZ24, Din24a, GLL23, LZ20b, YZ23]. iteration [Dij23]. iterative [GCFS23, UN23]. ITRT [PK22c].

J [AE23c]. Jackknife [KD24, SZ24, TLN24, WY20, XFW22, ZWFZ23, HCZ20]. Jacobians [Li21a]. Jajte [Bou22, NJMM24]. Japan [Ünv21c]. Jeffreys [Fow22, LSP22]. Jeffreys-Lindley [Fow22]. Jensen [RBD24]. Johnson [CX24]. Joining [WLZ^{+}23a]. Joint [AuAH22, DY21, KM24a, LXW20, NSG22, RK21, SSG21, ASG24, Cha23a, CKL23, Ker20, MK20, NP23, uAK22, PP22a, RSG21, RuAD22, SS21b, WLT21, XY20, YZW23, ZZRH22]. joint-exposure [Ker20]. Jointly [KG22, Ang24]. judge [CP24b]. Jump [YZ21, GKS21, HSWZ21, HW24, LSSW22, MCL23, SYZS23, SHXW23, XWYZ23, XWY21, Yan20, YCW21, ZMLH21]. jump-diffusion [GKS21, HSWZ21, HW24, LSSW22, SYZS23, Yan20, YCW21]. Jump-robust [YZ21]. jumps [CM24, Lef22, LMS22, ME21, SZL21, Zha21a]. juxtaposition [WYP21].

Kalman [Lou23, PF23]. Kaplan [LZ20a]. Karamata [YXP24]. Karhunen [You24]. KBER [EW23]. Kendall [LZG20a]. Kernel [BNO22, BGT23, KHK20, MHS22, RSR22, RT23, TLRB21, WG20, ZAG24, Ass21, AGB21, BBM24, BL20, BD22, Cha24, DR21, DKA20, EW23, EN20, Fak23b, GG24a, HZZA23, HBOR21, KAJ^{+}22, LKS22, LSSW22, LM20, Mod23, Nko23, Ory20, hSsXfX24, Sla22, WS22, WW23b, dXY23, YZXY24, You23a, Zen20b, ZWWW24]. Kernel-based [MHS22, DR21]. Kerridge [BBCL24]. Kibria [KD24]. Kiefer [Gai22]. Kim [KS22]. kind [JL24a, WX20]. kinds [HS21]. kink [LHLD22]. knockoffs [YLLZ24]. knots [ZYC24]. known [BM20a]. Kolmogorov [Fre20b]. Kolmogrov [NSJ22]. Korean [KL22a]. Kuk [LS22b, SHC20]. Kullback [BBCL24, ZHA20b]. Kumaraswamy [AE23a, CSBM24, GH18, KN22, KK23a, KC21b, PB22, SS22c, WYG24]. Kumaraswamy- [KN22, KC21b]. Kumaraswamy-distorted [SS22c]. kurtosis [MAAZ22, SMOO23].

l [Zen20b]. Lack [LL23b, GW23]. lack-of-fit [GW23]. Lack-of-partial-memory [LL23b]. lacunary [EAB20]. LAD [MLSS24]. LAD-lasso [MLSS24]. lag [EGD23, TÖ23]. lags [RLQ23]. Laguerre [Ben22b, BC23a, SSW20]. lambda [MPM20]. Lancaster [HY23]. Laplace [DMP21, GG23b, JRN24, JV24, NMT22, Sem20]. Laplacian [ZL23b]. Large [BS22b, BMP20, IKM21, LGL20, LD22, LZ23b, Wan23b, AA23a, Bou21b, Bou22, CCR21, DSY21, DOV22, Fak23a, FLW24, GLK24, HN20, Hu20a, HS21, HFGZ24, HW22d, JSB22, LF20, LJL^{+}21, MSM20, MAK20, NBM22, NJMM24, NSJ22, PX23, SYM24, Sud22, WC20b, YK23, YDHY24, Zen20b, ZW22a, ZL20, Zha24b, ZZY^{+}23, ZYY20a, dS22]. Large-sample [BS22b]. largest [ADIK22, FCCD20, FZD24, GZZ21, HBH24, KC21b, KCB23, SAH^{+}23, XFP23]. Laspeyres [MB21a]. Lasso [GÖ24, Pen24, AZL22, Liu24b, MLSS24, ySLZ22, NLL22, OFO^{+}21, SW23]. last [Cha23b, CC22, MC20, ST23a]. latent [CZX24, Che23e, ONM^{+}24, WZZ22, XTXW20]. Latin [SMS21, SS22e]. lattice [GZ23a]. Law [LZ20b, YZ23, Bou22, Cam24, CCR21, CS22, DSY21, DZ24, Din24a, DOV22, Fak23a, GLL23, HN20, HS21, HHLG22, HW22d, KS24a, LF20, MSM20, Mat21, NBM22, NJMM24, NSJ22, PX23, Yua24, ZL20, Zha24b, ZYY20a, KK20]. laws [Bou21b, Hu20a, HFGZ24, JSB22, MAK20, SYM24, Vid22, YDHY24, ZW22a, dS22]. layer [ADIK22]. lead [Cha23b]. learning [APV24, Bic22b, Bic22c, DT24, HW23a, KW22, LLW23, LG24a, SY21, SK24b, STC23, YLLZ24, ZSCL22]. Least [Pop20, WSG21, Wei23, YN21a, YD24c, ZS23, ZZGL22, BM22, CZZ^{+}23, CZL24, DZZ23b, FZW23, Far23, Far24a, Hu24, JV24, Kun24, LLX21, LL24b, LYZ^{+}23, MKÖ23, Nie20, RTT23, SWW24, Tsi24, WWWH20, WY20]. Least-squares [YN21a, Nie20]. leave [CGW23, SMV23]. leave-one-covariate-out [CGW23]. leave-one-out [SMV23]. Lebesgue [NCNM23, WS22]. Lee [Ano23b]. Leffler [Gaj22]. left [KERAE23, sS20b]. left-truncated [KERAE23, sS20b]. Legendre [LLH^{+}24]. Lehmann [Bal22, SBB22]. Leibler [BBCL24, ZHA20b]. leisure [ZS20a]. lemma [Fow23, ALS21]. length [Kia22, MMG24, OIF21, OCSV23, QSCH22, QMS22, RSR22, SLX22, SZ20b, WS20, WHW23, ZAF24, ZAG24]. length-biased [OIF21, OCSV23, QMS22, RSR22, SLX22, SZ20b, WS20, ZAF24, ZAG24]. Leone [AAO21]. Leptokurtic [BPZ22]. Lerch [LOT22, Xia20]. less [DCT24, QT22, ZFDT24]. Letter [GG24b, Pui24]. level [CLZ24a, CQT21, DG21b, EJJ23, JZ24, KR24b, KKS24b, KJ23, LLZ20, LOL21, LWYL23, MJB24, QT22, SSH23, WLZ20, YZZ21, YW22, ZCT^{+}20, ZC21, ZFDT24]. level- [EJJ23]. levels [FHHW20, Sha22b]. Lévy [Gaj22, WSSY20, Xia20, CZ20b, Le 21, LYD24, MGK23, WSG21, Wei21, ZLZ20a, ZDZ22]. liability [PC22, SLS21, ZW22b]. Libby [Ahm21]. Life [LNP20, AÜ20, ABJR22, BSN22, FM24, FTTY22, GK20b, JKD20, KIA^{+}21, LC22a, Lee21, PK20b, PK22c, RG24, Sho24, SWHW22, SD23, WS20, Yan20, Yu22a, ZS20a]. life-cycle [ZS20a]. Lifetime [MBLY21, BAY20a, BR23, BAHB23, BP24, CSZ20, CBLS21, FAN21, GSH22, HL22b, KBMD23, PJB24, PMMK24, SD20, SWHW22, WHW23, WJC23, ZAG24, ZG21]. lifetime-capability [SWHW22]. lifetimes [ASKB23, BR23, CLWZ22, CGL21, FAN21, KCB23, Wu24a]. light [Baz24a, KNCK24]. light-tailed [Baz24a]. like [FMM^{+}22]. Likelihood [BKHB23, ML21, Skr23, WC22, Alm24, AuAH22, BZLW24, Bic20, BB24, BS24, ÇA20, CWJC20, CL23, CL21b, Che22, Das23, Don23, EGD23, EAAC22, FHZC23, GX21, GvRS24, Gho23, HL22a, HX22, HZ22, HL24b, HS20, JC20, JLG21, KW22, KGG21, KM23b, Lem22, LQ21, LXZ22, LLQ22, Li23a, LWD^{+}21, LZW23, LL24c, LWYJ24, Mat21, uAK22, PSLLN20, Qin21, Rao23, RLQ23, SP23, SS20a, SA22, sS20b, TZYY22, TQ24, TY23, TLN24, Ver20, VK23b, WL22a, Wu22c, WL23b, Wu24b, XL22, XFW22, YLJ23, Zen24, ZCT^{+}20, ZYZ20, ZWFZ23]. Likelihood-based [ML21, GvRS24, Ver20, VK23b]. likely [GRBL24]. Likert [BS22b, Mur22]. Likert-type [BS22b]. Limit [EBX20, MMZ20, RTT23, TZ21, TZL24, XY23, XM21, YZ20, ZQH23, AAGE23, BEhF24, CDP20, DT21b, GZ23a, GZ23b, HHW21, KNL24, LTW24, PP22a, ST23a, SLF^{+}22, SST22, WYJD21, Wu23b, XWH21, YXZ23, Zha20a]. Limiting [XFP23, Ko23d]. Lindley [AE23d, AV20, AYS21, Ary23, CKM23a, Fow22, KMH23, KG22, KJ24, PWLO17, STK24, TIIB22, TKP24]. line [BCA22, Dij23, Far23, HHKC23, KK21, LH21, Ver20]. Linear [ASW21, HR23, INKS21, INKS23, JWW22, MP21a, NSS22, APC23, AU22, AMM20, Aga24, AB24, AS23b, AGP21, AEBK20, AvdWvW24, AH22, Bai21, BD24a, BABF22, BN22, BSR22, CPS21, CL23, CGW23, CMP21, CLPS24, CW22, CWK23, CZ23a, CW24a, CW24b, CH21b, CB24a, CB24b, DC24, DV22, Din24b, DZ24, DK23, DY21, Don23, DD23, EG23, Ema20, EA21, EM22, EA23, Ema24, FMZM22, FWWH21, FSS22, FH22, GG23a, GX21, GÖ24, GKV21, Gri24, GB21, GB23a, GZ23b, GCFS23, HS24a, Haq22, HMP23, HDZL21, HYZ24, Hol23, Hom21, HN20, HC22, HCZ20, Hu20a, HZ22, HW23b, HL24b, HFGZ24, Hu24, IK21b, JRCB23, JAE23, JW24b, JZWT20, JWW24, JH24, JV24, JRB^{+}23, KE24, Kim20, KP21, KHSS20, Lee20, LY24, Leu21, LW20, LC20, Li20, LMH21, LY21, LTW21, LZW24, LY22a, LSY22, LZW23, LZ23b, LTY23]. linear [LKK24, LM22, LW22b, LM23b, LY23, LZ23c, Maa23, MQK23, MMLBP20, MHZJ23, MNA22, MRD20, MRDD22, NC23, NN24, Nie20, O’G23, OO20, ÖA23, PS24, PS21, PF23, PS23d, PTYZ23, QAO20, RKK^{+}21, Rao22, SBC23, SE22, SS22b, SS24a, ySLZ22, Sma21, Sma24, SSM21, SJT21, SG22a, SBZ21, Tan22a, TZ21, TZYY22, TZ22b, TZL24, TP24a, TW20, TGMS23, TPV20, VCOB22, VCOS24, WC20a, WZD21, WCT22, WW23a, WCY23, WW24, WHMA22, WY23, Wu22a, Wu23a, WK23, Wu23b, Wyw24b, XW22, XFW22, YGG24, YRB20, YA22, YHA20, YD22, Yua21, Zei23, ZX24, ZCL24, ZW22a, ZW20, ZLH20, ZP22, ZL23a, ZYW23, Zha24a, Zha24b, ZFZ24, ZX23, ZGCZ23, ZWFZ23, ZPJY24]. linear-circular [KP21]. linearly [LPFL23]. link [HM22, MMLBP20]. linked [ZQX23]. Liouville [FE23]. Lipschitz [ME21]. liquid [YH23]. liquidity [GLB22]. list [SS23a]. literature [LEM24]. little [Che24b]. Liu [ARAK22, EA21, EA22, Li20, LK24, NAAQ22, PB22, QAO20]. Liu-type [EA22]. live [AFG20]. live-birth [AFG20]. lives [JRN24, Lee21, Sho24].

Liyanage [AE23d]. load [LGLL21, SGZA23, ZYJ20]. load-sharing [SGZA23, ZYJ20]. Local [AMM20, AB22, HB20, Leu21, SL20, SLYF23, AEBK20, BSR22, EM22, GZ23a, HCL22, HHKL23, KNL24, Nes24, PTYZ23, SBC23, SHFD23, Wal22, XWH21, YCS23]. Localized [Qar22]. Locally [ZYC24, LL23a, YNC22, ZT21]. located [CTEB20]. locating [CQT21]. location [AP24, AGB21, CX22a, DK24a, DMP21, DT21b, FZD24, GT22, JT24, KH21, KSY20, MG23, MMRMG20, NZZ20, NMF23, NCNM23, PZY24, Wu23c, Wu23d, YXS^{+}24, ZC23, ZSNX23]. location-scale [AGB21, DK24a, FZD24, GT22, NCNM23, PZY24, ZSNX23]. Loève [You24]. Log [BCKS22, CKM23a, HMM20, Alt21, CMC23, GBH23, Gri24, KERAE23, LC24, ML23, New23, PS24, PAC20, SP24a, VCOB22, ZHS^{+}24]. log-concave [PAC20]. Log-epsilon-skew [HMM20]. Log-Lindley [CKM23a]. log-linear [Gri24]. log-logistic [GBH23, SP24a, VCOB22]. log-normal [HMM20, LC24, VCOB22]. log-rank [CMC23, KERAE23, New23]. log-weighted [Alt21]. logarithm [DZ24, Din24a, GLL23, LZ20b, YZ23]. logarithmic [BKS24, KR23b]. logistic [ARGD21, AKUT23, AW20, AZL22, AE23c, CMP21, CZZ^{+}23, EA22, FHZC23, GBH23, GM21, HO23, HTP22, HHW21, LG24a, MBSS23, OSX^{+}16, Özk21, RSY21, SP24a, SPZ23, The22, TLLL23, VCOB22, WWC24, XJ21, YWS22, Zen22, Zen23, ZWLF24, dC22a]. logistic- [SPZ23]. logistic-truncated [GM21]. logit [HS22, Pho24]. logit-normal [HS22]. lognormal [Wan23a, GG24b]. Lomax [LK22b, KLK21, KLK24, QG23, VK23b]. Lomax-exponential [LK22b]. long [Aga24, AO22, AB22, CBLS21, HB20, RM21, Sha22b, YLL^{+}24]. long-memory [Aga24, HB20]. long-range [AO22, RM21]. long-span [YLL^{+}24]. long-term [CBLS21]. longitudinal [Akd22, AA22a, AWN21, CY24, Dag23, GL20, HX22, KM24a, sK20d, NSG22, QG24, RSG21, SSG21, SF24c, SL24b, TZ23, WJS22, XL22, YG20, YWL20, ZGC22]. look [Ang24, Han23a, Par23]. Loop [AR24, ZW22b]. Lorenz [GDSJ22]. loss [AO24, AAR22, AHO22, AOS24, CPS21, GZ21, Han21a, HYZ24, Jaf22, KM24b, KL22b, LC24, LRWZ22, LY22b, MQK23, MN23a, MO22, NSN20, PKSS22, WQZL21, WGX21, YCW21, ZXSZ20, ZZW^{+}24]. lost [CTEB20]. lot [BNAK23, MJB24]. Lotka [Wei21]. lots [MJB24]. low [HWW24, Ish20, JBJ23]. low-sample-size [Ish20]. Lower [ADIK22, CQT21, CZ23b, LLZ20, LWYL23, BS20a, BS22a, Goo24, LLW23, WLZ20, ZYY20b]. Lower-order [LLZ20, ZYY20b]. Lq [HZ22]. Lq-likelihood [HZ22]. LR [SX23]. LS [MZ22]. LSE [ZB23b]. LTRC [BGT23]. Lukman [KD24].

M [AE23c, AE23d, BS24, Kia22, KJ23, LYY24, TZ24, BS24, CVDY21, Kia22, KJ23, MR21, SB22, SACQ23, TZ24, Yan22a, ZYZ20]. M.A [Ano23b]. M.L.E. [Yo22]. M/D/1 [SACQ23]. M/G/1 [MR21]. M/G/1/K [Yan22a]. M/M/1 [BS24, Kia22, KJ23, TZ24]. machine [LG24a, SY21, SK24b]. MaCSim [HM22]. magical [KT24]. magnetic [NMT22]. Mahalanobis [Bul20]. mail [ASI^{+}24]. main [DG24, GPM23, sK23b, Lee23]. Maintenance [CF22b, BFSN22, CC22, LHWK24, LW23, Tas24]. Makeham [KCB23]. making [KTMKS20, KK23b, LCYL24, RL23]. Mallow [Liu24a]. management [Gar21, PC22, SLS21, ZW22b]. manifold [Sal21]. manifold-valued [Sal21]. Manifolds [HR23, WHMA22]. Mann [Che20a, UN23]. manpower [DG21b, ONM^{+}24, UE21]. Manson [WT20]. Mantel [RJ23]. manufacturing [BFSN22]. many [Con24, JB22b, KT24, WCX22]. many-input [WCX22]. MAP [KG23]. mapping [LHJ21]. maps [Tap22]. margin [PK22a, ZW22b]. marginal [ADD22, DY21, HDZL21, LX22b]. marginally [CB22]. marginals [Bay20b, JBJ23, NCGB24, RBC^{+}23]. margins [FL20b]. mark [KM24a]. marked [Wan23b]. Market [TS22, Ünv21d, GLLL23, Hua20, Ünv21c]. Markov [ARF23, BR20, Baz24b, BC23b, Cav20, CDP20, CYB23, DSY21, DLZN23, Ekh23, FW22, GLLL23, GM20b, GC22b, Hua20, Ino24, KS24a, KG21, LGLL21, LY22b, LW22c, LMS22, MBLY21, MCL23, Oga20d, Oga20e, Oga21, ONM^{+}24, PX23, Rao22, SYZS23, SGZA23, SYT21, SLF^{+}22, SXZ24, SBHY24, UE21, Xin23, YXZ23, eZN22, dZySgYW21, ZYY20a, ZSYM21]. Markov-modulated [SYZS23]. Markov-switching [FW22, LW22c, ONM^{+}24]. Markovian [KG21, Kum24, LBG23, LWKQ20, MB20, PG22, SCGB24, TY21, TY23, ZWT^{+}24, ZOD20]. Marshall [BAHB23, BP24, HL22b, LL23b, PB24, She22]. Marshall/Olkin [HL22b]. Marshall/Olkin-type [HL22b]. Martingale [CZ20b, MMZ20, WDW24, ZHQ21]. martingales [DS23]. masked [CSZ20, YZC22]. massive [JWWW24, KX23, nLSZ24, SXS24, ZZZ22]. material [YXS^{+}24]. materials [SS22g]. Mathematical [RIA20, DNP21]. matrices [AdMSK23, AWL^{+}22, Cav24, GC22b, GB21, Hu24, Li21a, LLSZ22, PH21, QH23, Wic20, YKH20, Yu22c, ZB20, Zen20b, ZCCT22, ZC22]. Matrix [Hu24, AK21, AvdWvW24, CL21b, DGCL23, Far21, GZZ21, HW22c, HN21, KF23, Lem22, Li21a, LMS20, MSY23, MKÖ23, Mon24, RO24, SP24b, TK20, Wic20, Yam20, YK23, Zen20a, ZDY24, ZWYL22]. matrix-variate [Wic20]. matter [TS22]. Matusita [AAE22]. max [SG23, CKL23]. max-INAR [SG23]. maxima [LXW20, Lin20, Lou20, PP22a, SST22, ZT24]. Maximal [KdXYC23, BMG22, Bou21b, HYWZ24, LW21a]. maximin [LTR24]. maximization [BTH23, Pen22]. Maximum [BS24, Das23, EGD23, LWD^{+}21, Rao23, SA22, TY23, Bic20, BB24, ÇA20, CG24b, Don23, FHZC23, Gzy21, HA22, Ima20, JLG21, KM23b, Ko23d, Lem22, LL24c, Mak24, Mat21, NS21d, PSLLN20, QE21, WL23b, YLJ23, ZYC23]. Maxwell [FWX24, HK21, KAC22, PVC^{+}23, RB21, RTB20]. McDonald [AE23c, OSX^{+}16]. McKay [MP21b]. McMillan [DSY21, PX23]. McNemar [Wu21]. Mean [BLZ23, Fab21, JL22b, PC22, AU22, AAGN24, ACK23, AAHS21, AHS^{+}24, And20, AuAH22, ASS23b, BL23, BKK22, CPS21, CKL23, CRVK24, CP21, Che23b, CMM20, CN22, CDCC24, DLHZ21, DJW^{+}20, DL22, FM24, Fro23, GP20, GD21, Gau21b, GLLL23, Gzy21, HN23, Haq20, HA22, HS24b, HS20, HN21, INKS21, IM21, JB22c, KM21, KGG21, KKS24a, KO23a, KM24b, KTP22, Koç21, MSY23, MN21b, Mat23, ML23, MDB^{+}21, NT20b, uATH21, uA21, Ory20, PP22b, PAAA20, Pri20, PKT21, Pri24, QH21, RK21, RJ23, RuAD22, Ros22, RK20, RG24, SSS20, SLV^{+}20, SGM22, SANHS21, SKCE23, SMMC21, Shu21, SPSS20, SSNP20, SK20f, SYG21, SV21, SPS21, SGTG24, SAK21, TK24, TH24, Tas24, TZ23, TLN24, ÜK21, Vog22, WDH^{+}22, WH23a, WKS21, WC20c, WS20, WLT21, Wu24a, XZR^{+}20, YCW21, Yin24, YXZ21, Yua21]. mean [ZK21, ZAG24, ZM24, ZW22b, ZLZ20b, ZYZ24]. mean-based [SLV^{+}20]. mean-field [WH23a]. mean-reversion [WC20c]. mean-risk [BL23]. mean-RVaR [ZM24]. Mean-variance [PC22, CP21, DLHZ21, GLLL23, WDH^{+}22, XZR^{+}20, YCW21, ZW22b]. means [AKST22, Ano21b, CL23, DR22, FZ22b, INKS23, Ima20, KS22, LC20, RG23, XP21, YKW23, ZL21]. Measure [ZD23, AAE22, BBCL24, BBM^{+}22, CH21a, CF22b, Fab24, GK20a, HCCD24, HM24, KK22, KU22, Lav24, LCYL24, Mod23, OC22b, QH22, RSR24, RBD24, SVR21, SYT24, XXZ24, dXG20, ZY23, ZCE23]. Measurement [TGMS23, BA24a, BA24b, BKS24, CWJC20, DSB22, EA21, EM22, Ema24, GG20, GDK22, GZ20, KE24, KSS21, LMH21, MSEB23, MHZJ23, Pri24, QG24, RK24, RKS22, RL23, RuAD22, SMMC21, SV22, Sir24, The22, XH23, YS20, YRB20, YA22, ZC24, ZYC24, Zha24a, ZFZ24]. measurements [ATJ^{+}24, CK21, DAS^{+}20, HAA^{+}20, HGAJ21, JBR^{+}24, MO20, NC23, NN24, RASD23, WZA21]. measures [BT22, BFPM22, CLZ24a, CMP21, Cha23a, CP24a, CM24, HABE24, JS23b, KB23a, Kur24, LD22, MGK22, MAAZ22, Rao22, RG22, SCAK23, SCP23, SCGB24, SDK^{+}24, SBHY24, YF22, YTT23, ZL21]. Measuring [KK20, SHFD23, BH24]. mechanism [AAR21, sK20d, NSS24]. Median [ZL21, AU22, BMT20, BS21b, Bro22, Fab21, QSCH22, SGN22, SANHS21, Tsi24]. median-based [AU22]. Median-of-means [ZL21]. Medical [SCAK23, AAQ23, KK22, SY21, ZZZ20]. Meier [LZ20a]. Memory [AuAHS23, uA21, Aga24, AB22, BL24b, HB20, LL23b, MLZ21, WH23a]. meningococcal [LWY^{+}23]. merged [KA22]. Merton [SYZS23]. mesh [RFVR20]. Meta [KP21, PMMK24, RBC^{+}23, ZLLS21]. Meta-analysis [PMMK24]. meta-elliptical [RBC^{+}23]. meta-regression [ZLLS21]. meteorological [OCSV23]. Method [MLW24, ZS23, AF23, AFG20, AP24, BNO22, BMK21, BMP20, BCG24, BS20b, CSA23, CLLD23, DR22, FL20b, GLB22, Gau21b, GA20, GCFS23, GZ24, HS24a, HM21, HJ22, HHKC23, HHW21, HX22, HWW24, JC22, JWWW24, KOTAG21, KF23, KK23b, Koç21, Kun24, Kur24, Lee20, LC20, LYLY21, LC22b, LYY24, Liu24b, LY23, NMT22, Nko23, PHBV23, PYZS23, PLLC23, RASD23, RL23, RIA20, SYRV20, SAH^{+}23, Sha22a, SD20, SF24a, SC20, SS21b, TH23, Ünv21c, UMA22, WY20, WCT22, WXW24, zWZ20, YP24, YZ21, Yu22c, Zen22, ZCL24, ZAL21, ZPL23, ZHS^{+}24]. method-of-quantiles [BMP20]. Methods [DNP21, BR22, CCMtH22, CN23a, CZ20b, CGLP24, DMR22, DY21, MÁH22, SK20f, SPS21, SJP23, SRSB20, WY23, WTT20, WWX22, YP24, YXS^{+}24, ZB20, Ano23b, AE23a, AE23b, AE23d, Cha23c, KS22, AE23c]. metric [CC23, KÇ21a]. metrics [Gzy22, SHFD23]. Metropolis [HFJ21]. MEWMA [Haq22]. microarray [Kan20]. Min [She22]. Min-infinite [She22]. minima [LXW20, NS21d, NSS22, PP22a]. Minimal [ATJ^{+}24, JBR^{+}24, SKRR23, Gon21a, HNR^{+}23, LWW24, RNA^{+}22, TA22]. Minimally [VVHK24]. Minimax [BL20, CN23a, HA20, YLL21, Ben22a, Ema20, Far23, MN21b]. Minimization [YZW23, AP24]. Minimizing [CTEB20, TB22, ASS23a, HHW21, LG24c]. Minimum [CZ23b, Kim20, SLS21, ZZZ24, BN22, Bul20, CG24b, CYW21, DZZ23a, JZ24, KM24b, KJKC20, MKÖ23, Oga20b, PS21, QE21, TLL23, WLZ20, WDLW20, Zha22, ZYY20b]. mining [CCMtH22]. Mis [YHH24]. Mis-specification [YHH24]. misclassification [NC23]. misclassifications [NN24]. MISE [KCG21]. Mises [SG24, SA21]. misestimation [Sha21]. mismeasured [LHS20, ZG23]. mispricing [MZR20, WRZW21, WDH^{+}22]. missing [ASG24, FM20, Fus22, HL24b, HS20, JLG21, JS23a, JRB^{+}23, sK20d, KHD24, LS22a, LZLL20, LL23c, Mor24, ON23, QMS22, RAA22, SSH23, Shu21, SPS21, SU22, SJP23, Sun23, TLLL23, WZD21, WJXF21, WL22a, Wan22b, WYW24a, XFW22, YO23, YKH20, ZT24, ZL22d, ZWFZ23]. missingness [Mor24, UMA22, YG20]. misspecification [Oga20b]. misspecified [MP21a, Oku23]. mitigating [MLSS24]. Mittag [Gaj22]. Mittag-Leffler [Gaj22]. mix [XY20]. mixed [ARAK22, AS23b, AvdWvW24, ATB24, BAAJ20, CQT21, DAZ23, FMZM22, GB21, GB23a, HS24a, HBOR21, JRCB23, JZ24, JRB^{+}23, KB22a, KHZ24, Li20, LWYL23, MMLBP20, MA23, MN23b, MO20, PS23d, Rao23, RSG21, SS24a, SS23c, SKRR23, SL20, SL24b, TP24a, TPV20, WD20, WS22, WLZ^{+}24, Wu23a, WK23, XS21, Yan20, YLJ23, YG20, YRB20, YA22, eZN22, ZSYD24, ZW24b]. mixed-effects [TP24a, TPV20]. mixed-level [CQT21, JZ24]. Mixing [GM20b, BG22c, DCY20, DDM24, HB20, KdXYC23, Kim20, KCG21, KC23, LDZ^{+}22, Sla22, TZ21, TQ24, You23b, YC23, ZSY21, dS20]. Mixture [BG21, Lee23, AAGN24, BHB22, CX24, CB22, CLZ24b, Fu24, JJ20, KW22, KE20, KSA23, Lab23, LZ23a, LLZ24, LWYL23, LMLS23, LG24c, LKK24, LZ24, MPS23, PM21, PS24, PK22b, PAC20, Qar22, RS24, SE22, SBB22, SAF24, Shu21, SK20e, TZ22a, WLL^{+}20, WLT21, Wu22b, YW20, ZCW23, ZP22, ZH24]. mixture-of-experts [PK22b]. Mixtures [El 23, PH21, AYP21, BP22, GZ21, Hua23, LM23a, NT22, NCNM23, PZY24, TKP22, WL23b]. MLE [Aga24, AO24, GSH22, Hu20b, YLW24]. MMAP} [JRD24]. MMR [DNP21]. modal [AP20, KR21, XZZ22]. mode [BBM24, BKS23b, Fab21, JWWW24, LKA^{+}23, RT23, You23b, ZCW23]. Model [AS22b, Bic22b, Cha23f, Fu24, GDSC15, HCZ20, RKK^{+}21, AV20, AFO21, AAGN24, Akd22, AAUA23, AAQ23, AS23b, AG23, ADD22, Alt21, AvdWvW24, AUQ22, ATB24, ABRT23, Ary23, AW20, AE23b, BCA22, BD24a, BR20, Baz24b, BM22, BL20, Ben22b, Bia22, BM20b, BSR22, BER23b, BKS23b, BS24, CBLS21, CMO^{+}23, CWJC20, CLC22, CLPS24, CW20a, CZ20b, CLZZ21, CWCY23, CN23b, CWK23, CG23, CLP23, CZ23a, Che23a, CZX24, CZL24, CY20, CX22b, Che23e, CDMR24, CB24a, CB24b, CW23, Dag23, DHJS23, DC24, DK24a, DMP21, DHH^{+}23, DG21b, DC22b, DKA20, DY21, Don23, DAZ23, DP22, DL22, Dün21a, Dur22, DM22, EBX20, EM21, ES22, FVCdC23, FAN21, FZW23, FL20a, FZ22a, FBNP23, FP23, FSS22, FTTY22, FW23, FLW24, GLK24, GPD21, GW21, GT22, Gho23, Gon21b, GG24b, GKS21, Guo21, GQW21, GA22, HO24]. model [HN23, HK20, HL23, HDZL21, HCCM21, HW22a, HYZ24, HBOR21, HTP22, HCLW22, HX23, HWZX23, HL24b, Hua20, HKL21, HA20, HCH20, HSC22, IHDR20, Ish20, IKM21, Jaf22, JRD24, Jan20, JAE23, JZWT20, JPZ23, JWWW24, JZC24, JS23a, JPJB22, JMQ20, KBKK24, KTDW20, Kha24, KAC22, KR21, KP21, sK23a, sK23b, Kob24, KD24, KHZ24, KG21, KA22, Kum24, KCB23, Kun24, KH20d, Kur24, Lee21, LK22a, LBF24, LYC20, Li20, LH21, LY21, LYL22, LHLD22, LLL22, LL23a, Li24, LL24a, LT24, Lin21, LY22a, LGL20, LXWY21, LGLL21, LWD^{+}21, LY22b, LHXZ22, LW22a, LX22b, LC22b, Liu22, LS23, LA23, LZW23, LX23b, LLL23, LTY23, Liu24b, LK24, Liu24a, LG24b, LG24c, LNP20, LMM20, Lou20, LQWW21, LMZ23, LZ23c, LJ24, LMS22, MZR20, MBLY21, MSW23, MWLS24, MQK23, MB21a, MPM20, MP21a, Mef23, MPS23, MB20, MHZJ23, MdWG23, MR21]. model [MBA20, MAT24c, MOAF23, NT20a, NK23, NSG22, Nak21, NZZ20, NS21c, NS22a, NMT22, O’G23, Oga20b, Oku23, OCSV23, OO20, OV21, ONM^{+}24, ÖA23, PY20, PLSW23, PS24, PKP23, PS22, Pen24, PSLLN20, Pho24, PS23c, PS21, Pou20, Pou21, Psa22, PAC20, QAO20, QZ22, QG24, RSY21, RGS24, RG22, RYZ24, Ros22, SE22, SSH23, Sal24, SYRV20, SS22d, SS22g, SHC20, Sha22a, sS22j, SWK24, SS24b, SE21, Shu21, SY21, SS22k, SDK^{+}24, SMV23, Sma21, Sma24, SZGK21, SS23c, SSW20, SW23, SAK21, SL20, SJT21, SG22b, SG22a, SGW23, SWLZ23, Sun23, SZ24, SF24b, TY21, Tan22c, TKP24, TH23, TZZ24, TP24a, TW20, TB22, TÖ23, TLRB21, TPV20, UE21, UC21, VRB22, Ver20, VK23a, WZTW20, WT20, WM21, WZD21, WWL22, WW22, WWZ22, WZL22a, WSZ22, WQSG22, WTS24, WDW24, WWC24, WYG24, Wei21]. model [WY23, WC20c, Wu22a, Wu23a, WK23, XZZ22, XTXW20, XL22, XZLB24, XL23a, XWY21, XS21, XW23, XWW24, YN21a, YLJ23, YPD23, YW24, YD24a, YW20, Yu20, YD22, Yua21, YJZZ23, ZBB24, Zei23, ZCL24, ZWLF24, ZXSZ20, ZLH20, ZHQ21, Zha21a, ZHQ23, ZQH23, ZZK23, ZYW23, ZL23b, ZSL24, ZFZ24, ZS24, ZS20a, ZL22b, ZZZ20, ZQX23, ZXLT21, ZL22d, Zhu22a, ZW24b, ZH24, ZHS^{+}24, ZWFZ23, ZPJY24, ZZ23, dCBV21, vdMGSM21, Ano23c, Ano24, NSM21]. model-free [LC22b]. Model-modified [Cha23f]. Modeling [AA20, BHL21, Elm22, Gar21, PS23b, WLZ^{+}24, ASG24, AKUT23, AWN21, BP22, BA21, CBLS21, DKD22, Dün21b, Gos22, GZ21, Gri24, KM24a, LNXC24, LKK24, Mur22, NSG22, Pra21, Sha23, SSG21, SCP23, Spe24, SB23b, SD23, SS21b, TIIB22, VVS^{+}24, ZLNS22]. Modelling [WQZL21, Lee21]. Models [ASW21, DNP21, HR23, SRA20, YPXZ21, ZS23, ZC24, AB24, ARAK22, AGP21, AOGM24, Alm24, Ami23, AFFM21, AFFM23, AL24, BZ23, Bai21, BBH20, BMK21, BK24, Bou21a, CX24, CMO^{+}23, CPS21, CBB^{+}21, Cav20, CF22a, CFS20, CLW^{+}22, CH21b, CMM24, CW20b, DDL20, Dag22, DLT21, DGCL23, Ema20, EA21, EM22, EA23, Ema24, EGD23, EA22, EAMH22, Fan23, FMZM22, FL21, FL20b, FW22, GL20, GX21, GND^{+}23, GDK22, GZ20, GB21, GB23a, GCFS23, HO24, HSWZ21, HSG21, HW24, HP20, Han21b, Han23b, Han24, HS24a, HMP23, He23, Hol23, HCZ20, HZ22, Hu24, HK21, HPW24, IK21b, JTS22, JRCB23, bJwWyX23, JWW24, JMSG24, JRB^{+}23, KR24a, KWC21, KLK22, KS24a, KM23a, KE24, KM23b, KR22, KC21b, LaM24, Lab23, Lem22, LY24, LQ21, LMH21, LTW21, LXZ22, LLQ22, Li23a, LWY^{+}23, LG24a]. models [LCNB21, LZW24, LZ20b, LYLY21, LSSW22, LSY22, LGZ22, LKK24, LJL^{+}21, LW21b, LLWA22, LZ24, LYZ^{+}23, MMLBP20, MN21a, McI23, ML21, MZ22, MN23c, MU22, MO20, NS21b, Oku23, ORH22, PH24, PHBV23, PCZ21, PK22b, PWM20, PB22, PS23d, Qar22, Qin21, QMS22, RSG21, RBC^{+}23, RLQ23, ST24, SM22, SHA24, SS22b, SSA23, SBB22, SS24a, SC22, SA22, SYL23, ySLZ22, SK20a, Sho24, Sir24, SRSB20, SL24a, SKRR23, SZL21, SLYF23, SF24c, SL24b, SBZ21, Tan22a, TZYY22, TZ22a, TY23, TGMS23, Tsi24, VE21, WC20a, Wan21b, WJS22, WCT22, WL23a, WLS23, WYW24a, WLT21, WW23b, XLY^{+}20, XWCT21, XFW22, YY21, YCC23, YRB20, YA22, YM22, YW20, Yu22b, YLLZ24, YWL20, ZCW23, ZX24, ZLLS21, ZP22, ZCCT22, ZL23a, ZLS23a, ZYC24, Zha24a, ZSYD24, ZFZ24, ZYZ20, ZGC22, ZZY^{+}23, ZX23, ZGCZ23, ZOD20, ZMLH21, ZSNX23]. models [vRvR20]. models* [CX22a]. Moderate [BZLW24, CZ21, FW23, LCF20, Wan23b, WTS24]. Moderating [Bic22c]. modes [HLPL22, LHWK24, LNXC24, LCNB21, LZW24, LW23]. Modification [BMK21, AAO21]. Modified [AAUA23, ACK22, AIGB20, SP23, Wan22a, ZB20, ZX23, APC23, BBAA21, BS21b, Cha23f, CWK23, CRKC21, Dün21a, GAG24, HXCW22, LL24c, MKM24, RKS22, RBCG20, SSA23, Wu23c, XLW23, ZZ23, SKP23]. Modified-Half-Normal [SKP23]. modifier [CDCC24]. modulated [Ber23a, SYZS23]. modulus [KÇ21a]. molecularly [sK20c]. Moment [KBKK24, BPZ22, CCW^{+}20, CLW^{+}22, Che22, DM22, FWWH21, GL21, GWD24, HW22d, HYWZ24, KdXYC23, Ko20, Ko23b, LW21a, LW22b, MWW22, Qu22, Sha22a, SW24, TZ22b, Wal22, WS23, XY21]. Moment-based [KBKK24]. moment-parameterized [BPZ22]. Moments [AZAZ24, GvRS24, HS22, AF23, Cha23d, DMP21, GM20b, GM21, Oga22, Psa22, WZL23, dSO22]. momentum [KS24d]. monitor [PP22b, SMMC21]. Monitoring [AMAA^{+}22, ABJR22, CN22, KHSS20, MAT24c, MOAF23, PA21, SAH^{+}23, AKST22, AAGN24, ACK22, ACK23, AuAH22, BKK22, CKL23, EIAZ23b, EIAZ23a, KR24b, KGG21, KRCS24, MMRMG20, MMSC22, MNA22, NS21a, NMF23, uATH21, uAK22, RuAD22, SS20a, SMMCM20, YHA20, ZLZ20b, ZYZ24]. monotone [DG24, HS20, KWB23, Nie20, SSH23, YKH20, ZT21]. monotonic [BGM22]. monotonically [GC22b]. monotonicity [CMC23, HW22c, LF23]. Monte [SM20, JC22, KOTAG21, TG24]. month [MC20]. Morgenstern [ABXE22]. mortality [GKvZB21, HT23, PLLC23, Sho24, WDLW20, WHCC23, XZLB24, ZLZ20a]. mortgage [GQW21, ZWLG22]. mortgage-backed [GQW21]. Mosaic [SEGMA21]. most [BB24, RSA20]. motion [FE23, HCLW22, Rao23, RTT23, Sem20, STW23, XY23, YZ20, YCS23]. motions [DDM24, El 23, ZYS21]. Moving [Ano23c, ACK23, Che23a, CW24a, CW24b, Din24b, GW21, HCL21, LQ21, uA21, NuAMA22, PS23c, Qu22, Sha23, SA22, WS23, Wu24b, YN21a]. moving-average [GW21, Qu22]. MRA [QQS^{+}22]. MRA-EDCC [QQS^{+}22]. MSP [KG23]. MsS [HBOR21]. MsS-K [HBOR21]. MT [SS22b]. MT-estimator [SS22b]. MTP2 [Fur23]. Multi [AP20, CGL21, ERAS22, LGX^{+}22, CLZ24b, Che24a, DR21, DG21b, EIAZ23b, FMHP24, GLLL23, HW24, HWW24, sK20b, sK23a, KS21, LBG23, LY24, LOL21, LGZ22, LJL^{+}21, MLSS24, Pou20, QG24, Shi23a, VVHK24, WLZ^{+}24, WC20c, XZR^{+}20, Yu22c, ZZ21, ZX22]. multi-channel [EIAZ23b]. multi-component [Che24a]. multi-dimensional [HWW24, Shi23a]. multi-environment [FMHP24]. multi-environmental [VVHK24]. multi-factor [LY24, WC20c]. multi-factors [LGZ22]. multi-level [DG21b, LOL21]. Multi-modal [AP20]. Multi-objective [LGX^{+}22]. multi-outcome [MLSS24]. multi-period [GLLL23, XZR^{+}20]. multi-regional [sK20b, sK23a]. multi-response [CLZ24b, LGZ22]. Multi-sample [CGL21, Yu22c]. multi-samples [ZZ21]. multi-scale [HW24]. multi-server [LBG23]. multi-state [KS21, Pou20, QG24, WLZ^{+}24, ZX22]. multi-step [DR21, LJL^{+}21]. multiauxiliary [SGTG24]. Multicollinearity [GDK22, GZ20, Moh22b, SBK24]. multicomponent [CF22b, KTDW20, RSY21]. multidimensional [FS22, OA21, SMH22, SYL23]. multidimensionality [AW24]. multilevel [SSA23]. Multinomial [Cha23c, AZL22, Cha21b, Liu22, Mir23, UE21, Yo22, ZG23, Cha21b]. multinormality [ZL22c]. Multiple [AFFM21, Ima20, KKNN23, Yan22b, Arn23b, ABJR22, BKHB23, BA24c, BN22, BS21b, Bro22, CG23, CS24, DR22, DKA20, HSG21, HHW21, HC22, IY23, JB22a, KKS24b, Lee20, LNXC24, LLXW24, LWKQ20, LW23, McI23, NSM21, Naj22b, NP23, Oga20e, Oga24, OHS21, Ohy21, RAS^{+}22, RGS24, RYZ24, SD23, SWLZ23, Tas24, UMA22, WWZ20, WLZ^{+}23a, WY23, Wic20, WZ21, Wu23c, Wu23d, Wu24a, WV24, XLW23, Yan23, YHA20, YC23, ZAL21, Zha21b]. multiple-outlier [BKHB23]. multiple-set [BN22]. Multiplicative [HZZA23, RK22, Ben22b, DL22, Fak23a, ZFZ24]. multiplicity [LZ21]. multiscale [DWZZ24]. multisets [Kon22]. multistable [Le 21]. multistate [Cha23a, Ino24, JC20, ZLB20]. multitest [Fur23]. multitype [SBHY24]. Multivariate [ABL23, Cha22, KC23, MSY23, YKW23, ZSCH22, AN23, AAGN24, AKSL22, Bay20b, CMO^{+}23, CWJC20, CPS21, Cav24, CB22, CLPS24, CBJ20, DT21a, FVCdC23, GH22, GBGPNR23, Gos22, GZ21, GC22c, GB23b, HKNW20, Haq22, Hu20b, IK21b, JAJ21, JWW24, JKP20, KLK24, KM23a, KWB23, KRS20, KHSS20, KX23, KSY20, LT24, LL23b, MVLMA20, MN23c, NBM24, Naj21, NMF23, OMYF21, Oga20e, Oga24, OV21, PP22b, RYJ20, RBC^{+}23, Sem20, SJT21, TZ23, TDLL20, TK20, Vam20, VSVVLRV22, Wal23, WH23b, WL23b, dXG20, YHA20, Zhu22b, ZY24b]. musings [Bal22]. mutation [PCdP^{+}20]. muth [JA22]. mutual [GZ24, HHKL23, KE20, KGS20, KX23, Mao20]. My [Bal22].

N [GC22a]. Nadaraya [LB22]. Nadaraya-Watson [LB22]. natural [FH24]. ND [CW24b]. near [DS23]. near-martingales [DS23]. nearest [AALE22, LDZ^{+}22]. nearly [BAGR22, CZ21]. necessary [dS22]. needed [SYRV20]. Negation [KS24b, ZD23]. Negative [RG22, Alm24, Ary23, BA21, Cha21b, Cha23c, FVCdC23, GG24a, JBSK22, JW24b, JB22b, KG23, LHS22, LPFL23, LM22, LM23b, Nie20, Nou23, OTL21, RS24, RTB20, SK20e, TT22, TIIB22, TZ24, WW23a, WW24, Xia20, YSS22b, ZDZ22]. negative-binomial [Xia20]. negatively [CCW^{+}20, CW22, CC23, DC22b, Din24b, Din24a, FWWH21, GWD24, HW23b, Ko20, Ko23b, Ko23d, LM20, MWW22, MWW23, MW24a, NBM22, SW24, SPSS20, TTL20, WW20, jWWDjW24, Wu23b, XH24, ZW22a, ZLH20, ZQH23, ZWWW24]. neighbor [HNR^{+}23, LDZ^{+}22, RNA^{+}22, SMS21]. neighbors [AALE22]. nested [IK21b, NMT22, SKRR23]. nested-error [IK21b]. net [SCAK23, ySLZ22, YY21]. network [FBNP23, HCX20, KJ23, LQWW21, LMZ23, NS21a, SS22i, SW23, XW23, ZZZ22]. network-induced [ZZZ22]. networks [BM21, CZL23, GA20, Gos22, HCX20, JC22, LCWZ21, MAT24c, MOAF23, NS21a, PY20, QH22, SS20a, WWL22, WZL22a, WZL^{+}22b, XW23, ZLB20]. neural [GA20, JC22]. newly [TA22]. Neyman [Fow23, NAJ22a]. NHPP [AD21, DKD22]. night [KNCK24]. NN [MBLB24]. no [CM20, Raï24]. nodes [Gos22, HCX20, NS21a]. noise [Ben22b, CLP23, IKM21, Kou23, KHZ24, Oh23, OV21, SMZ23, TPV20, ZCT^{+}20]. noises [HB20, HCL22, HHKL23, WSG21, WSZ22, Wei21, Wei23]. noisy [LLWA22, Maa23]. nominal [GRBL24]. Nomogram [HHKC23, KL22a]. Non [AGR21, BKK22, KGS20, LKA^{+}23, LWT24, LTR24, LX22b, NBM24, PCZ21, QZ22, TWZ22, XO21, ASG24, AGP21, AAKL21, ABJR22, ASI^{+}24, BD24a, BGM22, BP20, Bro22, CDL24a, CRVK24, CLW^{+}22, EW24, EJJ23, FLW24, GA20, GG24a, GM20b, HCH20, Ima20, JBSK22, JJ20, KLS23, KM21, KMH23, LY24, LCF24, LKK24, MLW24, MN23b, Nie20, Oga22, OTL21, PP22a, QH23, SAH21, SK20f, SBB21, TZYY22, Tee22, TZZ24, Tri24, ÜK21, Wal22, WZD21, WQSG22, WJL23, WDW24, WC24, WYW24a, WKS21, WHMA22, Y.24, YXZ23, YS20, Yu22b, YZ23, Yua21, Zey20, ZL20, ZHQ23, ZQH23, ZFZ24, dZySgYW21, ZL22c, ZOD20, ZT24, ZWCP22, ZSNX23]. non-additive [ZL20]. non-central [OTL21]. non-conforming [ABJR22]. non-crossing [WC24]. non-homogeneous [KLS23, Y.24, Yu22b, dZySgYW21]. non-identical [YZ23]. non-ignorable [ASG24]. non-increasing [Nie20]. non-linear [BD24a, Nie20, WHMA22]. non-local [Wal22]. Non-marginal [LX22b]. non-Markovian [ZOD20]. non-maximum [Ima20]. non-monotonic [BGM22]. non-multinormality [ZL22c]. non-negative [GG24a, Nie20]. non-normal [SAH21]. non-normality [QH23]. Non-parametric [KGS20, CLW^{+}22, KM21, WHMA22]. non-recursive [Oga22]. non-response [ASI^{+}24, MLW24, SK20f, ÜK21]. non-sensitive [WKS21]. non-stationary [GA20, GM20b, PP22a, YXZ23]. non-strategic [EJJ23]. non-uniform [Tee22]. nonadditive [CCR21]. noncentral [BG21]. Nonconcave [FZW23]. noncooperative [WLZ^{+}23a]. Nonexistence [DMD23]. nonhomogeneous [CX22b]. nonignorable [Wan22b, YG20]. Nonlinear [BM22, FP23, AS22b, CBB^{+}21, CDP20, FL21, GW23, IKM21, LZ20b, MZ22, MO20, PTYZ23, SHA24, SJF21, SL20, XL22, ZC24]. nonnegative [DH22]. nonnormality [Shu21]. Nonparametric [ACMC21, BC23a, CMCA20, DBBK21, EN20, FZ22b, HCCM21, LYL22, LM20, Rao24, sSL20, SLX22, SZ20b, TP24a, TPV20, WS20, ZZ21, ZCCT22, AÖK^{+}24, BEAA21, BR22, Ben22b, BSR22, BS20b, BKS23b, BS21b, DC22b, DZZ23b, EIAZ23a, GW23, GK24, Haq20, Hua20, HKL21, LL20, LWY^{+}23, LMS20, NMF23, NS22b, OIF21, Sla23, SSM21, TZ23, WZZ22, Wu20, YZXY24, YZT21, ZHQ21, ZL23b, ZWWW24]. nonstandard [LGL20, LG24b]. nonstationary [AC23, ZYS21]. nonsymmetrical [AH22]. norm [Bey23, KB22b]. Normal [Dey24, Sul22, AKST22, AKAB23, ATAAM20, Ano21b, Che23c, CRKC21, DMP21, EAS20, FMZM22, Gau21b, GH22, GZ21, HS22, HX23, HMM20, Ima20, JAJ21, JL22a, JWWW24, KLK22, KLK23, KW22, KGG21, KO23a, KTP22, KS22, KSY20, LC24, LXW20, LMM20, MN21b, MN23b, NT20b, NAJ22a, Naj21, Naj22b, O’G23, Oga22, RKAA21, RG23, SAH21, SCPAV21, SKCE23, SMV23, Sul21, TK20, VCOB22, Wic20, WL23b, XXBD23, Yam24, YFW^{+}23, YKH20, ZCW23, Zey20, ZZW^{+}24, ZL22b, ZL22c, HXCW22, SKP23]. normality [AB22, AT22, BKS23b, CB24a, CC23, HB20, HMZ20, KK23c, LZG20b, MBLB24, Mot23, Nes24, Par21, QH23, SEGMA21]. normalization [CS24]. normalized [DE22, DJW^{+}20, Zha20a]. Normalizing [HWNN23]. normally [Ang24, SMOO23]. norms [XH24]. note [Arn23b, BA24a, DG21a, EG23, Far21, FC22, Far24b, Fus22, Ko20, KHF22, LHJ21, LHLD22, LHS23, LTTL24, Lig23, LQWW21, LMZ23, MGK22, MLWSPS24, MW24b, NBM22, OIF21, PSLLN20, PKK22, RSA20, ST24, SS22e, Sem20, SWYF20, SBZ21, Wan21b, Wu23d, XY20, YT22, Yam20]. notes [Far23, Far24a, ZQX23]. Notice [Ano24, KV23]. notions [Hon20]. novel [BKS24, BCG24, GA20, JWWW24, KIA^{+}21, LT24, Pho24, SS21b, Wij24, ZZZ20]. Novick [Ahm21]. np [ZYZ24]. NQD [SYM24, TTLT23]. NSD [CH20a, HYWZ24, KAJ^{+}22, TZL24]. nuclear [Gar21, Sha22b]. nuisance [AO24, BEhF24]. null [Bic23, Bic24, Yam24, ZN24]. number [ABJR22, BEAA21, Che20b, Fan23, FW22, KT24, LYC20, LS23, LJL^{+}21, LW21b, Oh23, YXZ21, ZZY^{+}23, dHGS21, dRJO21]. numbers [Bou21b, Bou22, CCR21, DSY21, DOV22, Fak23a, HN20, Hu20a, HS21, HFGZ24, HW22d, JSB22, LF20, MSM20, MAK20, NBM22, NJMM24, NNS20, NSJ22, PX23, SYM24, YDHY24, ZW22a, ZL20, Zha24b, ZYY20a, dS22]. Numerical [GLB22, LC20, Che23c, ZHS^{+}24, ZCE23].

Objective [Fow22, He23, KLK22, KLK23, PB24, KLK21, KLK24, LGX^{+}22]. obligation [YN24]. observation [CW20b, FM20, JC20, LL23c, STW23]. observation-driven [CW20b]. observations [AFO21, CC23, DMR22, DHH^{+}23, EW24, GS24, HL24b, LZ20a, Maa23, ON23, RAA22, RT23, RTT23, SAH^{+}23, SWK24, SMMC21, SB23a, SJT21, Tan22c, WL22a, WF23, Wei21, XZL20, Yao22, ZT24]. observed [CDCC24, Wei23]. observing [Sha22b]. obstacle [HWWY20]. obstacles [WX20]. obtain [HGAJ21]. obtained [MvSK20, PYZS23]. obtaining [ZIZ23]. Occam [Bic20, Bic22b]. occasion [SPSS20, SK20f, SJP23]. occasional [ZCT^{+}20]. occasions [BP20]. occupancy [KH20a]. occupation [LYD24]. occurrence [DKD22]. ocean [BA24c]. Odd [Del21, SP24a, VCOB22]. odds [DS24b, PAdlPK22, PKP23, WW22, WJL23]. of}- [MMSC22]. off [KSK^{+}23]. offspring [RK21]. Oliveira [AE23c]. Olkin [BAHB23, BP24, LL23b, PB24, She22]. Olkin-type [HL22b]. OLS [Ros22]. Oluyede [AE23d]. on-line [Ver20]. One [LK22c, MMSC22, MÁH22, AO24, BD24a, BMP20, Bul23, CGW23, DMD23, JAE23, JL24a, JB22b, JBJ23, LCNB21, MMS23, Nie20, QSCH22, RKS22, SMV23, TLN24, TP24b, WHW23, Wu23c, Wu24a, XTXW20, YZ20, ZY20, ZPJY24, vdMGSM21]. One- [LK22c]. one-dimensional [BMP20, JBJ23]. one-parameter [BD24a, MMS23, Nie20]. one-shot [LCNB21, WHW23, ZY20]. One-sided [MMSC22, AO24, QSCH22, RKS22, ZPJY24]. One-tailed [MÁH22]. one-way [vdMGSM21]. ones [JB22b]. online [ZXYJ24]. Open [ZW22b, HWW24, YZC22]. open-high-low-close [HWW24]. Open-loop [ZW22b]. Operating [EM21, Cha21a]. operator [KWY20, LW22a, LYY24, Nag20]. optical [LHJ21]. Optimal [ASS23a, AGB21, CLZ24a, Cha21a, CLC22, Cha23b, CK21, CP21, CH21a, CLWZ22, CZLY22, CH22, Che23d, Che24a, CMM20, FHL^{+}24, Gon21a, Gri24, GK20b, GKvZB21, GKS21, GK24, HN23, HT23, JB22a, Jun20, KL22b, LLL22, Li23b, LL24a, LZL20, LLL23, LW23, MZR20, MLZ21, Nov24, Onw21, QSCH22, QH22, RYZ24, ST21, SS23b, SG22b, SG22a, TLL23, WIN20, WRZL21, WDH^{+}22, WSZ22, WH23a, WDLW20, Wu20, YP20, Yan20, YD24a, Yan24, YZSS23, YH23, ZZRH22, ZDW22, ZY23, AN23, BPZ24, BK22, CZ20b, Che23c, CLZ24b, DLHZ21, DESSM21, DYC^{+}24, FZ22a, GC24a, HTP22, JZC24, KRS20, KS21, LY24, LHJ21, LLSZ22, LZ23a, LLZ24, LWW24, LGZ22, LLH^{+}24, LWYJ24, LS20b, LZ24, MJB24, NSM21, Naj22b, PS24, PKK22, SB20, Ten24, WRZW21, XXZ24, Xu22, YCWZ22, ZWLF24, ZYC24, ZM24, ZW24a, ZH24]. optimality [Cha23e, RASD23]. Optimization [SS22a, ZYNX20, BT22, BBP22a, LHWK24, LGLL21, LGX^{+}22, MKG^{+}21, OFO^{+}21, TY21, Tas24, WLLB24, XZR^{+}20, Yan22a]. optimized [XWH21]. optimizing [MAT24a]. Optimum [CC22, GA22, PM21, AOGM24, YHH24]. Option [Guo21, HW24, MCL23, SYZS23, YCS20, Alg23, AW24, GLB22, HSWZ21, HCCM21, LLWL21, LJ24, LMS22, RN20, RN22, ZAL21, ZL22a, ZL22b]. optional [Gao23, LBG23, NS21b]. options [LYLY21, SHXW23, XWY21, YMZD23]. Oracally [Gon21b]. Oracle [Pen24]. Order [CG23, CZ23b, APC23, AZAZ24, ABXE22, AÖK^{+}24, AHO22, Aug21, BG22b, BEAA21, BKM20, BM20b, CP22, CBJ20, Cha23f, CN23b, CZZ^{+}23, CKM23b, DK24a, DT21b, EG23, El 23, EAB20, EB21, EAAC22, EG20, EBG22, FCCD20, GL21, GG23b, GM21, HBH24, Hu20b, HABE24, JT24, JZC24, KLS23, KP24, KB22a, KBMD23, KHF22, KC21b, Kun23, LLZ20, LF23, LWT24, LLXW24, Lin20, Lin21, LGLL21, LW22c, LLG24, MBR22, MGC22, Moh22a, NSG23, RK21, RIA20, Ril24, SSS22, TEÇK22, UET24, WLZ20, WZL23, WF23, XM21, YDHY24, YXP24, Yin24, YCS23, ZCT^{+}20, ZSYD24, ZYY20b, ZX22]. order-replacement [ZX22]. Ordered [JG23, Hol23, JMQ20, JS23b, NT20b]. Ordering [APC23, CKM23b, FCCD20, HBH24, KK23a, KC21b, KCB23, NZZ20, PJB24, EAAC22, GG23b, HL22b, JAJ21, PKP23, PS22, Vid22, WC22, ZZ21]. Orderings [BBAA21, BAHB23, SHB23, BBH20, BAA24, CWM23]. orders [BHB22, BKB22, BKHB23, LEM24, NT20a, ROSL21]. ordinal [CY23, GZ20, SYT24, Yu22b]. ordinary [Hu24, JV24, MKÖ23]. ordinates [RM21]. origin [WZL23]. Ornstein [CL21a, HN23, RYZ24, SB20, WSG21, Wei23, XZL20, YD24c]. Ornstein-Uhlenbech [HN23]. orthant [LGL20, LW21a, LW22b]. orthant-dependent [LW21a]. Orthogonal [CZ23b, SAP20, Far23, Moh23, OYO24, OYA24, PZZ20, PYZS23, PCZ24, Pop20, RZ23, RMS^{+}24, Tsa22, WYP21, YO23, ZDY24, ZLPZ24, ZFDT24]. Orthogonality [XL22, ZGCZ23, FH24, OC22b]. Orthogonality-based [XL22]. oscillating [TPV20]. other [Alv23]. outcome [GWM21, sK23b, MLSS24]. outcomes [Akd22, LWS24, Ohy20, Ohy21, SCP23, WZA20]. Outlier [ZLLS21, AA22c, BKHB23, DMD23, PK20a, Sha21]. outlier-driven [Sha21]. outliers [Bey23, CLZ21, Dün21b, KSM^{+}21, MDB^{+}21, PS21, SBB24, Tsi24, YLJ23, ZG23, vdMGSM21]. outlying [Özk21]. output [BZ20, BZ21]. outputs [KHVBRm22]. overdispersed [HK21]. overdispersion [CH20b, FVCdC23, dCBV21]. overlap [DG21a, MC23, ZCZJ20]. overlapping [AAE22]. overlooked [LK22b]. overview [Pos24].

P [Lav24, Fli20, YM22, ZHQ23]. P-C [ZHQ23]. P-values [Lav24]. Padé [YY24]. paired [ADW22, KAC22, uATH21, RKK^{+}21]. Pairwise [BD24b, AWB21, Ko23b, SYM24, TTLT23, dS22, dS24a]. panel [ARK22, EAMH22, FP23, HYZ24, bJwWyX23, KA22, LXZ22, LLQ22, LY22a, Liu24b, RLQ23, WWZ22, WW23b, XP21, XLY^{+}20, ZJF23]. paper [BAE23, GG24b, KS22, Pui24]. papers [ST24]. parabolic [ARF23]. paradigm [KAC22]. paradox [CL21c, Fow22, MW24b]. parallel [AGHK20, BBH20, BD21, BmAB22, BKB22, BKHB23, BKB23, CSZ20, CKM23a, GBH23, Goo24, HL22b, KN22, LX21, LXWY21, PJB24, PK22a, SS22a, SHB23, SPZ23, Tsa22, VM22, WC22, WLLB24, ZYJ20]. parallel-flats [Tsa22]. parallelism [SSH23]. Parameter [AV20, ABRT23, APV24, CLP23, GZ20, HXCW22, JRM21, Wan23a, Wei21, XZL20, YK20, ZYS21, AO24, AB22, AAR22, BD24a, CPS21, CLMA20, CL21a, Che23b, CG23, CS22, CW20b, GCCJO^{+}21, Ham23, Han20, JWW22, JWW24, Jun20, LG24c, LJLW23, MQK23, MN23a, MKG^{+}21, MMS23, NS22a, Nie20, Oh23, PM21, PKK20, Pui24, RG23, SD20, Sha21, SS24b, SE21, SS22k, TIIB22, TA22, TPM22, WY20, WXW24, WK23, YT22, YGG24, YP20, YXS^{+}24, YLL^{+}24, YHA20, Zei23, ZXSZ20, ZZW^{+}24, ZWT^{+}24, ZG23, ZCZJ20, ZXYJ24, ZPL23]. parameter-specific [Jun20]. parameterization [Mat21]. parameterized [BPZ22, Wan21b]. parameters [ACKL23, BMP20, BEhF24, ÇA20, CDMR24, CMM24, DMR22, DLC22, Fan23, GT22, GSH22, Han23a, IAM21, Jaf22, JT24, JL24b, KP24, KSY20, LS23, LW21b, MN23a, NAAQ22, PKSS22, PF23, QAO20, SHA24, STK24, SS22h, Sho24, TLLL23, Wu23c, Wu23d, YW24, YRB20, ZLF24]. Parametric [GDSJ22, VCOS24, VE21, AAKL21, AA23b, BCA22, BKK22, Bro22, CDL24a, CLW^{+}22, Don23, EW24, Fab24, GG24b, HK20, Jaf22, JJ20, KM21, KGS20, LKA^{+}23, LWT24, LTR24, Liu22, LWS24, MU22, NBM24, QSP^{+}24, QZ22, Sun23, TWZ22, VCOB22, WDW24, WHMA22, XO21, ZHQ23, ZQH23, ZC24, ZSNX23]. parametrization [SM22]. Parametrizations [DESSM21]. Pararai [AE23d]. Pareto [Ano23b, AFL15, ATB24, BPZ24, BQW22, DLC22, FMM^{+}22, Gho23, GG24b, LSP22, LA23, MdWG23, SJ23, VK23b, YT22, ZG21, ZCZ20b]. Pareto-optimal [BPZ24]. Pareto-type [MdWG23]. parity [Ünv21b]. part [CZX24, ASW21, SMS21]. Partial [XW22, Zha24a, AS23b, AWL^{+}22, BABF22, CCW^{+}20, CYW21, EA21, GKvZB21, HWZX23, KdXYC23, Ko23d, LLX21, LY21, LL23b, NS21c, Oga21, PB24, SAP20, ST21, ST23b, TZL24, VCOB22, WW20, XWH21, ZWLG22]. Partially [JG23, VPM^{+}24, CB24a, CB24b, CDCC24, Ema24, HYZ24, HCZ20, HL24b, JH24, KG20, Kha20, LMH21, LTW21, LY22a, LSY22, MHZJ23, SS23a, SSS23, Tan22a, TZYY22, VCOS24, WC20a, WZD21, Wu22a, Wu23a, WK23, XZZ22, XL22, ZP22, ZL23a, ZGCZ23, ZWFZ23]. participating [WDLW20]. particle [SMZ23]. partitioned [LTY23]. partitions [Sil21]. Partly [HR23, Gai22]. Parzen [Ten24]. pass [ZWLG22]. pass-through [ZWLG22]. passage [ARF23, CDP20, Gaj22, Lef22]. past [AH22, BG22b, CG24b, HM24, Kun23, NSG21, SVR21]. path [CGW23]. patient [SY21]. patients [AWN21]. pattern [KT24, MU22, Shu21, ZFDT24]. pattern-mixture [Shu21]. patterns [SPSS20, Yan24]. payments [YD24a]. PBGC [XWYZ23]. PBIB [BG22a, VVHK24]. PCA [Dur22, SS20a]. PDEs [HWWY20]. PDF [MMS23]. peaks [ZCZ20b]. peaks-over-threshold [ZCZ20b]. Pearson [Mao20, Fow23, PZAH22]. penalization [Bey23]. Penalized [HZ22, JJ20, Liu24a, SE21, TZ22a, WL23b, ZJF23, FZW23, GG21, JV24, KE20, SS22b, XS21, Yos23, ZGCZ23]. penalties [JBSK22]. penalty [MGK23, YD24a, ZW24b]. pension [MZR20, NCC24, SK24b, WCY23, YCWZ22, ZDW22]. percent [OC22a]. percentiles [MMG24]. perception [ADW22]. Perfect [eZN22, BFSN22, CG24a]. Performability [JRD24]. Performance [BKS24, QAO20, RKS22, AMBK22, ACKL23, CGLP24, KSS21, Nov24, RG23, Sha21, SCGB24, WIN20, WH23a, WJC23, YTT23, ZG21]. period [BC24, BTH23, GLLL23, KSK^{+}23, WZZ22, XZR^{+}20]. periodic [Alm24, BM22, BM20b, DKD22, SS23c, YD24a, YK20]. periodically [AB22, Ber23a, HM21, PS23b]. periodicity [YM22]. periodogram [RM21]. periods [CK21, HAA^{+}20]. Permutation [CD24, WV24, CZ20a, HCCD24, Yu22c, YH24]. permutation-based [CZ20a]. persistence [Skr23]. persistent [AA20]. Perspective [ASW21, HC22, ZCZ20a]. perspectives [CH21a]. Perturbation [DDG23, LCF24, WC20c]. perturbations [MLWSPS24]. perturbed [CLZZ21, WM21, WQSG22]. pessimistic [CLZL24]. PH} [JRD24, JRD24]. Pharmacokinetics [LK24]. Phase [JMSG24, MMG24, MNA22, Yua24, CLZZ21, JMM22, QLZZ20, SMP^{+}24, Y.24, YHA20, ZK21, ZOD20]. Phase-II [MMG24]. Phase-type [JMSG24, CLZZ21]. phases [QLZZ20]. phenomenon [WRZW21]. phi [CMP21, Oga20b]. phi-divergence [CMP21, Oga20b]. physical [WIN20]. piece [SCPAV21, Sul21]. piecewise [GX21, Nak21, WCY23]. piecewise-exponential [Nak21]. pilot [AMBK22]. pioneering [Bal22, Yam22]. Pitman [VK23b]. plan [AN23, BAAJ20, BA24c, BAJ20, GKvZB21, JB22a, MZR20, MJB24, SS23b, TBCB24, WCY23, YHH24, ZY20, ZDW22]. plane [AH22]. planning [DG21b, ZZZ22]. plans [BNAK23, KS21, NN21, NCC24, RAS^{+}22, WIN20, YCWZ22]. plasticizing [Sul22]. play [EJJ23]. player [WH23a]. pleiotropy [JZWT20]. Pliable [AZL22]. pliant [BCA22]. plot [CMM20, DZZ23a, GC22c, JZ24, ZL22c]. plot3 [Wij24]. plots [OYO24]. PLS [Che23e, Che24b]. PLS-SEM [Che24b]. Plug [PAAA20, Jun20]. Plug-in [PAAA20, Jun20]. plus [KM24b]. point [AMM20, ARK22, AT22, AAGE23, Bic23, Bic24, CBLS21, CX22a, FWX24, GL21, GX21, Ham23, IW22, JL22b, KM24b, LWYJ24, Mef23, New23, Sal24, SCAK23, ST23a, SA22, TÖ23, WWZ20, WLZ^{+}23b, Yan22b, YC23, ZN24, ZL22d]. points [JBJ23, SWLZ23, XFP23]. pointspread [WSZ22]. Poisson [AE23b, AE23d, KV23, And20, ABRT23, CB22, CH20b, Che22, CX22b, FWX24, GDSC15, GML20, GWM21, HK20, Han23b, Han24, HK21, JS21, JRM21, KLS23, Koç21, KHD24, KR23a, KS24c, LYC20, LZH20, LS20a, LYL22, LL24a, LOT22, LJLW23, MSW23, MOAF23, NS21c, NS22a, ND21, PWLO17, PKSS22, Pen24, RKAA21, RTB20, SS22k, SBHY24, Wan22a, Wu22b, XW23, Y.24, Yan22b, YNC22, YN24, YTT23, ZBB24, ZZK23, dSO22, vdHvDZ22]. Poisson-BE2 [ZZK23]. Poisson-Gamma [Wu22b]. Poisson-stopped [LOT22]. Poissonian [LYD24]. policies [CC22, Che23d, LW23, MDZN20, WDH^{+}22, ZM24]. policy [BC21, Cha21a, Cha23b, GPD21, KC20, KJ23, LLXW24, LWW24, LWKQ20, LYY24, WLZ^{+}23a, WL22b, Yan22a, ZLCW20, ZY23, ZX22, BC24]. Pólya [JTS22, Cha22, KLS23, XTXW20]. polychoric [LO23]. polygon [You23b]. polynomial [JZC24, Kha24, MMS23, Pop20]. polynomials [BN21, BL24a, DBBK21, Fak23b, Moh23, WHCC23]. poolability [GW21]. population [AS22a, AF23, AU22, AKST22, ASI^{+}24, BMT20, CRVK24, CS20, CZ21, DDL20, DSB22, Das23, GP20, GRBL24, HAK23, HSC22, Jaf22, JMK24, Kim21, KG21, Kum24, MN23a, Mot23, Mot24, NSH23, uA21, PS23a, Pri24, QH21, SGN22, SGM22, SPSS20, SSNP20, SK20f, SBB21, SYG21, SPS21, SMP^{+}24, TLN24, ÜK21, YMG24, YS20, YFW^{+}23, ZK21, ZB23a, ZCZ20a]. populations [AA23b, BD24a, CZL23, FCCD20, HN21, JT24, KLK23, KTP22, LHS23, LK22c, MA23, NT20b]. Portfolio [BT22, AdMSK23, BL23, GLLL23, HT23, WGX21, XZR^{+}20, YZSS23]. portfolios [CLC22, SL24a]. positions [HL24a]. Positive [Pos24, Han21b, HW20a, Kwo22, SNS20, WH23b, XXBD23]. positively [MN21a, SPSS20]. possible [CP24b, Oga20b, ZCE23]. possibly [WTS24]. Post [ySLZ22, HT23, LTTL24, YY21, ZPL23]. post-processing [ZPL23]. post-randomization [LTTL24]. post-retirement [HT23]. Post-selection [ySLZ22, YY21]. Posterior [DMP21, KLK21, Bic22a, Bic23, Bic24, Ham24, Han20, Han21a, JBJ23, Syr23, ÜFR21, Xin23]. potential [DG21b, MDB^{+}21]. poverty [DDG23, Jan20]. Power [JRB^{+}23, AP24, APV24, AA22b, AE23d, BMG22, ÇA20, CG20, CBM21, CS22, CGLP24, DJW^{+}20, HHLG22, HS20, KASK24, KK23c, LMLS23, O’G23, PWLO17, Sul21, Wan23a, XJ21, Yua24, YMZD23, Zha22]. power-comparison [KK23c]. power-law [CS22, Yua24]. powered [KM24b]. powerful [RSA20]. powers [JT24]. PQD [dS22, dS24a]. practice [AE23c, OSX^{+}16]. precedence [KHF22, MMRMG20, MMSC22, MGC22]. preceding [KT24]. Precise [FLW24, GLK24, HTP22, Kob24, SS22h, WC20b]. precision [ZWYL22]. predict [GA20]. Predicting [BAY20a, SCP23]. Prediction [AAM22, FM20, Gzy22, Jaf22, KL20, OYA24, ORH22, AFO21, AH22, AA22c, BEAA21, BR23, BER23b, DESSM21, LT24, LM23a, MP21a, NSS22, Oh23, Onw21, Qar22, SB20, SACQ23, STC23, SJT21, SZ24, XZLB24, YLL21]. predictive [ACMC21, CMCA20, HK20, Ham23, HCCM21, LX22c, Oga20c, RAA22, SV21, Wu22d]. predictor [MRDD22, Tri24]. predictors [IK21b, VK23b]. Preface [GKA21, ZS20b]. preference [CLC22, ZS20a]. preinvex [OA21, SMH22]. premium [BPZ24, DLHZ21, LZL20, Wu22b, XWYZ23, ZM24]. premiums [AÜ20, KWC21, NCC24, YCWZ22]. prepayment [ZWLG22]. prescribed [BBP22b]. presence [AO24, AA22c, ASI^{+}24, BBM24, BA24a, BA24b, CRVK24, CG21, DSB22, DMD23, EGD23, GW23, JTS22, KHVBRm22, KSS21, Lee22a, PS21, Pri24, Shi23b, SPSS20, SBB21, SPS21, ÜK21, WKS21, YS20]. present [LGL20]. Preservation [NT20a]. preserve [BBH20]. preserving [SH23]. pretest [SKCE23]. Preventive [MDZN20, Cha21a, Cha23b, CC22, Che23d, LHWK24, LW23]. price [Alg23, Bia20, Bia22, MB21a, TS22, YCW21]. prices [HSWZ21, ZL22b]. Pricing [GQW21, SHXW23, YMZD23, ZL22a, AW24, GLB22, Guo21, HW24, HCCM21, LLWL21, LJ24, LMS22, RN20, RN22, SYZS23, SZL21, XWY21, YCS20, ZAL21]. primary [SMP^{+}24]. primate [ASG24]. prime [EN20, HZZA23]. principal [CWK23, Dur22, EIAZ23b, GB23b, HWW24, Kob24, Özk21, WWX22, YLL21]. principle [BZLW24, CCR21, DLHZ21, HF24, LZ23b, LL24b, ZM24]. principles [BPZ24]. prior [AKUT23, DMP21, HZ21, HHLG22, KR24a, LSP22, MGK23, Mat23, Spe24, Wal22, Yo22, ZZW^{+}24, ZN24]. priority [WXY21, XLW23]. priors [HX23, KLK21, KLK23, KLK24, KSA23, Kia22, KAC22]. privacy [CN23a]. private [PY20, PH24, SK24b, WZL^{+}22b]. proactive [DYC^{+}24]. Probabilistic [PTM21, GB23b, LZG20a]. probabilities [Baz24b, CCR21, CLZZ21, CY20, JPZ23, KKKM21, Li24, Lin21, NC23, SYL23, VK23a, WQSG22, ZL20, ZPJY24]. Probability [CH20a, AD21, Baz24a, Bic22c, Bic23, Bic24, CS20, CW20a, CWCY23, CX22b, DT24, GAG24, Ham23, JPJB22, KS24b, LS22a, LS22b, LYL22, LCF24, Lou20, MvSK20, NN24, NCNM23, PCdP^{+}20, Rao22, RS24, RBS20, SLS21, Sil21, SGW23, TY21, TQ24, TB22, WM21, XWW24, XH24, Yu22b, YD23, YZW23]. probable [Ano21b, KS22]. probit [GZ20, Pho24, WWL22, WZL22a, JMQ20]. problem [AÖK^{+}24, BNO22, CBJ20, CZ20b, CLLD23, Con24, Ekh23, GRBL24, GKS21, Hua23, LMLS23, MLW24, MNAM20, MLZ21, RYZ24, TY21, TLL23, UN23, WRZL21, WX20, WL22b, Yan24, YFW^{+}23, YKW23, ZZRH22, ZW24a]. Problems [Muk24, BY24, DHH^{+}23, HW22a, HWWY20, KWB23, Lef22, LL24a, LK22c, MKG^{+}21, ZYNX20]. Procedure [YPXZ21, CBJ20, CP24b, EM21, Fur23, IY23, sK20c, Lee20, LJL^{+}21, RBS20, SPSS20, SSM21, SF24b, WZ21, Wu23c, Wu24a, Yan23, YY21, YW20]. procedures [Far23, Ima20, Sun23, TT23, Wu23d, XL23b, YCY22, Zha21b]. proceeding [GC22b]. Process [BA24b, ACK22, ACK23, AGR21, ABRT23, BNAK23, BBP22a, BM23, BKK22, BAJ20, CKL23, CDL24a, CH20b, CZ21, CDL24b, CW24a, CW24b, CS22, CN22, CYB23, DMR22, Din24b, FWX24, Fu24, GL21, Gaj22, GG23a, GG20, GM20b, HB20, HN23, Haq20, HA22, He23, HHLG22, IKT22, IW22, KGG21, KTMKS20, KL22b, KJ23, Kum24, KV23, LYC20, LS20a, LLWL21, LTW24, LWKQ20, LGX^{+}22, LK24, LW22c, MG23, Mat23, MO22, NAJ22a, NNS20, uATH21, OV21, PP22b, PA21, PG22, RL23, RK21, RK22, Rao24, RuAD22, SS22d, SG23, SGZA23, SMMC21, SB23b, Tan22b, TEHEG22, WSSY20, WS23, Wan23b, Wei23, Xia20, XWYZ23, XL24, Y.24, YCS20, YCW21, YZXY24, YLL^{+}24, YK20, Zen24, ZCT^{+}20, ZSL24, ZLZ20b, ZYZ24, ZX22, dSO22, vHP23]. process-waiting [KV23]. processes [Aga24, AAGN24, ABT20, ABL23, AP20, Aug21, BM20b, BMH24, BD22, BS21b, Bro22, CB22, CDP20, CLZ21, CL21a, Che22, CYB23, FSB22, GML20, GZ23b, HN20, JC20, JRB^{+}23, KLS23, KWY20, KH20b, KB23b, Le 21, Lef22, LP20, LC20, LH22, LWT24, LZ23b, LYD24, LW22b, Maa23, MGK23, MB21b, Nes24, OA21, PAAA20, QH23, Qu22, RBSNB20, RG23, Sha23, SMH22, SMH24, SMMCM20, SB20, SBHY24, TZ21, TZ22b, TWZ22, TZL24, VSVVLRV22, WWWH20, WSSY20, WSG21, Wu22b, XZL20, YZC22, Yan22b, YHH24, You23b, Yu20, YD24c, ZYS21, ZL22a, Zha24b, ZLZ20a, ZDZ22, ZZGL22]. processing [ZPL23]. product [AuAHS23, Gau21b, Nah22, RL23, SMOO23, XWH21, Zen20b]. production [NNS20, Tas24]. products [Cha21a, KR24b, LW23, MDZN20, TZL24, ZG21]. profile [GG23a, Ker20, MNA22]. profiles [Haq22, KHSS20, YHA20]. progressive [AZAZ24, AGHK20, AAM22, BS22a, CGL21, DG21a, JMSG24, Koh20, LG22, MK20, MHS22, QG23, TPM22, WYG24, WJC23, ZG21]. Progressively [DV22, BSN22, RG21a, Tri21, TPS20]. projected [Che20b]. Projection [AWB21, Sma24, WCX22, CZ23a, Naj22b, UN23, YLC^{+}22]. projections [HW22c]. prone [EGD23]. proof [Gau21a, Ham24, RIA20]. propensity [CG21, Wan21a]. proper [Bar22]. Properties [AMH22, ARGD21, Ano23c, AE23d, HMP23, Jan20, KR23a, KS24c, KJ24, PWLO17, PVC^{+}23, RBD24, SLV^{+}20, SKP23, Ary23, Aug21, BBM24, BS22a, BN24, ÇA20, CG24a, CP24a, CY23, CWM23, CKM23b, CRKC21, EAAC22, FCCD20, Fus22, GN24, GG24a, Gho23, GSH22, GM20b, HCL22, HBH24, Jas21, JRN24, KAAH21, KB23a, Ko23c, KR23b, KC21b, KCB23, Kur22, KS24d, Lee23, LP20, LZH20, LZG20a, LL23b, LNP20, MKM24, MZ22, Mir23, Moh23, NT20a, NSN20, OTL21, Onw21, OMS24, PJB24, PK22c, PS23c, RG23, ST23a, SYT21, hSsXfX24, SKS24, SZ20a, SZGK21, TKP22, Wan21a, WW20, WDW24, WWWZ24, Zen20a, Zen23, ZYY20b]. property [AGB21, Baz24b, CMC23, MR21, Shi23a, Wal22]. proportion [AS22a, CXT^{+}22, JJ20, KSA23, NSS24, PAdlPK22, Sho23]. proportional [BBAA21, CLWZ22, DS24b, DM22, FAN21, GAG24, GA22, PLSW23, PKP23, SYRV20, WW22, ZW24a]. proportions [AC23, CAAARCTS23, Das23, HW20a, HWH22, Jan20, Kim21, LK22c, WJL23]. proposal [GMC^{+}23]. proposals [Sud22]. proposed [AAR21]. propriety [KLK21]. PROS [BKK22]. Proschan [YF22]. protecting [CN23a]. proteomics [MMLBP20]. proving [BM20a]. pseudo [Hu20b, HWW24, JLG21]. psi [RT23]. psi-weakly [RT23]. pTQAI [CW20a]. publication [Ano24, MG24]. published [ST24]. pursuit [WCX22].

Q [ZL22c]. QMLE [AL24]. QSS [BNAK23]. QSS-RS [BNAK23]. quadrant [Ko23b, LPFL23, SNS20]. Quadratic [HWWY20, Ang24, CDL24b, Ros22]. quadruple [ACK23]. qualification [SWHW22]. qualitative [CLZ24b]. quality [BA24c, BKK22, PLLC23, SRSB20, Tas24]. quantal [Ker20]. Quantile [Che24b, DLT21, Mat24b, TPS20, XLY^{+}20, ZZZ22, AB24, AHS^{+}24, ASS23b, ASS24, BL23, CDL24a, CS20, FL21, GL20, HSG21, JH24, KLK22, KKS24a, Leu21, LH21, LMH24, LL23c, MvSK20, MS21, NSD22, NSSR23, PTM21, SLX22, Shi23b, Sho23, SS24c, TY23, Ünv21a, WJXF21, WL22a, WJS22, WL23a, WC24, WTT20, YP24, YPD23, Yos21, Yos23, You23a, ZJF23, ZZY^{+}23, ZMLH21]. Quantile-based [Che24b]. quantiles [AEBK20, BMP20, CMMD20, DMP21, Dou22, Kha24, KTP22, KS24d, Mot23, NT20b, SB23a, TDLL20]. quantitative [HSC22, Ker20, PKT21]. Quantum [HCL22, HHKL23, Sil23]. quartile [SU22]. quasi [Alm24, BB24, HMZ20, HNR^{+}23, KM23b, KJ24, SS22d]. quasi-associated [HMZ20]. quasi-likelihood [Alm24]. quasi-Lindley [KJ24]. quasi-maximum [BB24, KM23b]. question [NS21b, NS22a]. questions [SHC20]. queue [BS21a, BC21, BC24, BTH23, GPD21, GC22a, KC20, KLSZ22, Kia22, KG23, KJ23, LAAA^{+}22, LYY24, NP23, PG22, WLZ^{+}23a, Yan22a, YTT23, ZLCW20]. queueing [BLZ23, BS24, CVDY21, KSK^{+}23, LBG23, LLXW24, MR21, SS22a, SB22, SACQ23, SCGB24, TZ24, XLW23]. queueing-inventory [LLXW24]. queues [SA21]. queuing [ERAS22, SK22b]. quick [WWJ23]. quick-switching [WWJ23]. quota [YCW21]. quota-share [YCW21]. quotient [SNS20].

R [LGZ22, LBF24, RFVR20]. R-INLA [RFVR20]. R-optimal [LGZ22]. radiation [Sha22b]. radio [Lee20]. Radon [WHMA22]. Rahim [TH23]. raise [dHGS21]. Random [CD24, Del21, LG24a, LL23d, ZCL24, APC23, ALS21, ARF23, AvdWvW24, ASKB23, Ang24, AH22, AA22b, BA23, BEAA21, BMK21, BA24a, BL24b, Bou21b, BKS23b, CH20a, CP22, Cha23b, CCW^{+}20, Che21, CW22, CWM23, rChZhN^{+}24, CLS24, CW24a, CW24b, CL21b, CLW^{+}22, CKM23b, CMMD20, DSY21, DC22b, DLZN23, Din24b, Din24a, DK24b, DH22, DOV22, EG23, EBX20, Fan23, FWWH21, FH22, GZ23a, GD21, Gau21b, GPD21, GWD24, GLW24, GZ23b, HCCD24, HN20, HS21, HW23b, HCCzW23, HF24, HFGZ24, Hu24, HYWZ24, IKM21, JSB22, JAJ21, JW24b, JLG21, KNL24, KdXYC23, KK23a, KH20b, KMH23, Ko20, Ko23b, sK23a, Ko23c, Ko23d, KKKM21, KC21b, LKS22, LC22a, LW20, LZH20, LLX21, Li21a, LH22, LLL22, LPFL23, LTW24, LF20, LD22, LTY23, Liu24b, LCF24, LLH^{+}24, Loe22, LCF20, LW21a, LM22, LDZ^{+}22, LW22b, LM23b, LY23, LLWA22, MAK20, MWW22, MWW23, MW24a]. random [MAT24c, MiT24, Mon24, Mot24, MHS22, NBM22, NJMM24, New23, NSJ22, Oga20a, PJB24, PP22a, PKP23, PK20b, PS22, QR24, Rao22, Rao23, RTT23, RTN24, SHA24, Sal21, SP24a, SP22, SMOO23, SGM22, SANHS21, SW24, SLF^{+}22, hSsXfX24, SYM24, SNS20, SSNP20, SBB21, SBHY24, SZ20a, SD23, SBZ21, TTL20, TTLT23, TZ22b, Tan24, TP24a, TPV20, UN23, VK23a, WC20b, WW20, Wan21b, WJXF21, WS22, WW23a, WS23, WLS23, jWWDjW24, WW24, WLLB24, Wu23b, XY20, XLY^{+}20, XXBD23, YMG24, Yam24, YZ23, YWL20, ZB20, ZB23a, ZW23, Zen20b, ZW22a, ZCT^{+}20, ZL20, ZLW21, ZWWW24, ZGQN20, dZySgYW21, ZDZ22, ZT24, ZSYM21, ZSY21, ZXLT21, ZL22d, ZYC23, dS20, dS22, dS24a, dC22a]. random-effects [LTY23, YWL20]. randomization [KASK24, LS22b, LHS22, LTTL24, WZA20, YSS22b, ZIZ23]. randomized [AS23a, Arn23a, Arn23b, CN23a, CP24b, DHH^{+}23, EM21, HCH20, HSC22, LHS23, LTTL24, LLXW24, LHXZ22, NSS24, NS21b, NS21c, NS22a, New23, SHC20, SC22, SS22k, SPS23, ST23b, SMP^{+}24, SS21b, Yan22a]. randomly [CTEB20, CLS24, FL21, GWD24, GLW24, JW24a, Lin20, LW21a, RTN24, XY21]. range [AO22, DR22, LLG24, RM21, TEHEG22, XP21]. Rank [GB21, JG23, LSY22, WWZ20, ASG24, CMC23, CDSG20, DK23, FMM^{+}22, GB23a, JAE23, KERAE23, KWB23, New23, Tri21, ZLS23b, vRvR20]. Rank-based [LSY22, WWZ20, KWB23]. Rank-Ordered [JG23]. rank-sum [Tri21]. ranked [AS22a, AuAH22, AuAHS23, BK22, CG24a, CG24b, DZ20, Fre20a, GC24b, JRM21, KK24, uATH21, QE21, QR24, RR20, SV21]. ranked-based [AuAHS23]. ranked-set [Fre20a]. ranking [CG24a]. ranks [BD24b]. Rao [LSP22]. rare [NS21c, NS22a, SS22k]. Rasha [BAE23]. ratcheting [SS23c]. Rate [CDMR24, Mot23, RN20, APC23, AG23, ABL23, BG22c, BGM22, GZ23a, Guo21, HW24, Han24, JMK24, JA22, Kan20, KHZ24, Lee23, LH22, LTW24, LDZ^{+}22, LMS22, MM20, MA23, MZ23, NSG21, PCdP^{+}20, QSP^{+}24, QQS^{+}22, SGZA23, Sho24, SAF24, SWHW22, SZL21, Swa24, TKP22, TP24b, WM21, WLS23, Xin23, YLL21, YMZD23, ZLZ20a, Zhu22a, Zhu22b]. rate-based [WLS23]. raters [Ohy20, Ohy21]. Rates [LMS20, YY21, CLWZ22, DD23, GK24, HN20, HS21, Jan20, KAJ^{+}22, LJ24, MCL23, WZA21, YC23]. Ratio [Koç21, AKST22, AuAH22, AuAHS23, BM20a, BZLW24, BKHB23, CWJC20, CL23, CL21b, Che22, EAAC22, Far24a, HL22a, HS20, Jas20, KW22, KGG21, KKS24a, KK24, uAK22, OHS21, PAdlPK22, Par23, PLLC23, SAH21, SP23, SS20a, sS20b, Skr23, WC22, WJL23, Wu24b]. Ratio-type [Koç21, OHS21]. rationally [sK23a]. ratios [DJW^{+}20, KLK22, MAK20, RM21, Swa24, Wu22c, XXBD23, XM21]. Rayleigh [BQW22, CG23, LNWA22, RSA20, RG21a, WHW23]. razors [Bic20, Bic22b]. RCAR [SWK24]. rcll} [ME21]. re [BC24, GCFS23]. re-service [BC24]. re-weighting [GCFS23]. Real [FH24, Dij23, KIA^{+}21, LZ23b, MKG^{+}21, Oga20a, dZySgYW21]. real-life [KIA^{+}21]. real-parameter [MKG^{+}21]. real-valued [Oga20a]. recalibration [ZPL23]. recapture [LZLL20, PSLLN20]. recessions [KT24]. reconcilability [ZCE23]. record [AA22c, ASS24, BG22b, EBX20, PS22, Sha22b, VK23b]. records [Jas20, Jas21]. recovery [CLLD23, VM22]. rectangular [Kou23]. Recurrence [GM21, RS24]. recurrent [CFS20, CG21, DL22, sSL20, Yua21]. Recursive [Sla22, Ass21, BS20b, LKS22, Oga22, Sla23, WX20]. rediscoveries [Bal22]. reduced [GQW21, ZQX23, vRvR20]. reduced-form [GQW21, ZQX23]. reducible [Sud22]. Reducing [MLSS24, SYRV20]. reduction [ANO23a, CDSG20, Oku23, SSZC22, SWYF20, TWZ22]. redundancies [CLWZ22, YZZ21, YW22]. Redundancy [MAT24a, KF24, KHF22, MNAM20, SGZA23, WLL23, WLZ^{+}24]. reengineering [SS22d]. Rees [HNR^{+}23]. reference [SAH21, SJ23, Tri21]. reference-intrinsic [SJ23]. Refined [Fre20b, KOTAG21]. refinement [Sil23]. Refinements [ZHS^{+}24, Fro23]. refitted [ZS24]. Reflected [ME21, HWWY20, YD24c]. refracted [LYD24]. refurbishment [SLS23]. regime [GLLL23, GQW21, HSWZ21, HW24, LMS22, MCL23, SHXW23, YCS20, ZSL24]. regime-switching [GLLL23, GQW21, HSWZ21, HW24, LMS22, MCL23]. regimes [FZ22a, FW22]. region [OYA24, WV24]. regional [sK20b, sK23a, YL23]. regions [sK23a, Kwo22, LZ24]. registration [FS22]. registrations [AFG20]. Regression [CLF21, HR23, IHDR20, KO23a, LHS20, YSS22a, ZC21, AU22, Aga24, AB24, ARAK22, AAUA23, AAQ23, AYP21, AAHS21, ADD22, AAKL21, AALE22, Alt21, AUQ22, AHS^{+}24, ATB24, Ary23, AKUT23, AW20, AZL22, Ass21, BN21, BBM^{+}22, BABF22, Ben22b, BG22c, BGT23, BA21, BSR22, BKS23b, CGW23, CMP21, Cha24, Cha23f, Che23a, CLW^{+}22, DLT21, DT21a, DC24, DR21, DC22b, DZZ23b, DKA20, Don23, Dün21b, EA22, EIAZ23a, EAMH22, ES22, FHZC23, FL21, GL20, GÖ24, GW23, GCFS23, GK24, HSG21, HJ22, HO23, HBOR21, HHKC23, HHW21, HC22, HZ22, HK21, HKL21, HA20, IK21b, IKM21, Jan20, JBSK22, bJwWyX23, JWW22, JCY22, JWWW24, JH24, JS23a, JV24, Jun20, KR24a, KKS24a, Kim20, KK21, KR21, KP21, Koç21, KD24, KCG21, KHZ24, KHD24, KHK20, LS22a, Lee20, Lem22, Li20, LH21, LTW21, LHLD22, nLSZ24]. regression [LSY22, LGZ22, LX22a, LL23c, LZW23, LWS24, LMM20, LZ23c, LYZ^{+}23, MQK23, McI23, ML21, MdWG23, Moh22b, MBLB24, MOAF23, MLSS24, MBSS23, NSS24, NS22b, NAAQ22, ND21, OHS21, ORH22, OCSV23, OO20, Özk21, ÖA23, Pen24, PS21, PB22, Pop20, PVC^{+}23, QAO20, RKK^{+}21, RO23, RMS^{+}24, RG22, RBC^{+}23, Ros22, SBC23, SBK24, SS24b, SE21, Sla22, Sla23, SW23, SSM21, SLYF23, TZ22a, TKP24, TGMS23, TLLL23, TLRB21, Ünv21a, UMA22, VE21, VRB22, Wan21a, WC20a, WJXF21, WL22a, WJS22, WL23a, WDW24, WWC24, WXW24, WTT20, WV24, Wyw24b, XZZ22, XTXW20, XW22, XLY^{+}20, YP24, YLL21, YCC23, YL23, YY24, Yos21, Yos23, YW20, YD22, ZB20, ZBB24, Zei23, Zen22, Zen23, ZCW23, ZWLF24, ZLH20, ZHQ21, ZLLS21, ZHQ23, ZQH23, ZJF23, ZLZZ24, ZWWW24, ZZY^{+}23, ZX23, ZZZ22, ZMLH21, ZL22d, ZXYJ24, Zhu22b]. regression [ZLNS22, ZW24b, ZHS^{+}24, dHGS21, vRvR20]. regression-ratio-type [KKS24a]. regressions [BMR^{+}23, HW22b, JJ20, MVLMA20, Sun23, VCOB22, VCOS24, VSVVLRV22, ZW20]. regressive [SA22]. regressor [AALE22, LKA^{+}23]. regressors [AEBK20, Leu21]. regret [EJJ23]. Regular [GC24a, CLZ24a, CB22, SYL23, WLZ20, dXG20]. regularization [Tan22b, ZLZZ24]. regularized [Bul20, LXLL23]. regulatory [SLS21]. Reinsurance [HW23a, BPZ24, CZ20b, CP21, CH21a, CZLY22, GKS21, HGY24, KL22b, LRWZ22, Li23b, RYZ24, TLL23, WDH^{+}22, XXZ24, XL24, YCW21, Yan24, YZSS23, YZW23, ZZRH22, ZM24, ZW24a, ZY23]. reinsurance-investment [WDH^{+}22, Yan24]. reinsurer [BPZ24, LRWZ22, ZZRH22]. reinsurers [CH21a]. rejuvenation [ZOD20]. related [CYB23, DHH^{+}23, Gho23, HCL22, HWWY20, SJ23, WM22, XZ24, YZ20]. Relating [RJ23]. relational [KR22]. Relations [Psa22, GM21]. relationship [CLPS24, RMS^{+}24, Ril24, Swa24]. Relative [BER23b, AALE22, FZW23, KB23a, MÁH22, PAdlPK22, SCAK23, WH23a, ZHS^{+}24]. relative-error [AALE22]. release [Gar21]. relevant [XH23]. relevation [DS24b]. Reliability [DNP21, DAZ23, FTTY22, GA23, HLPL22, LHWK24, LNXC24, MSW23, RG21b, SGZA23, WZTW20, WXY21, WLLB24, WLZ^{+}23b, AYS21, ASS24, BAE23, BQW22, BC23b, Cha23a, CBM21, DKD22, DZ20, EEAA22, FZL24, Gao23, JMM22, JMSG24, KTDW20, KB22a, KS21, KJKC20, MNAM20, MiT24, Pra21, RSY21, RYJ20, UC21, Wan23c, XJ21, YZC22, ZLB20]. reliability-redundancy [MNAM20]. reliant [Che20b]. Remark [Far24c, LLW23, SW21]. remarks [TW20]. reneging [SK22b]. renewal [BC23b, CY20, JPZ23, Lin21, LGL20, LG24b, Lou20, PAAA20, Psa22, SS22d, SGW23, WC20b, ZPJY24]. renewing [Cha21a]. Renyi [LX23a, BKM20, Bic22b, HCCD24, LCYL24, PD24, YD23]. repair [ERAS22, Gao23, KC20]. repairable [HLPL22, LX21, LHWK24, LWKQ20, SS22d, WXY21, WLZ^{+}24, ZX22]. repairers [TZ24]. repairs [LWW24]. Reparameterized [PVC^{+}23]. repeatability [RK24]. repeated [ATJ^{+}24, CK21, DAS^{+}20, HAA^{+}20, HGAJ21, JBR^{+}24, KC20, MO20, NC23, NN24, RASD23, SCP23, WZA21, ZL21]. repeaters [Lee20]. repelling [YGG24]. repetitive [AN23, AMAA^{+}22, ACKL23, ABJR22, HCL21, JB22a]. replaceable [PD23]. replacement [Cha21a, Cha23b, Che23d, Che24a, DYC^{+}24, IM21, LS22a, MDZN20, Mot24, PYZS23, SLS23, SAK21, ZGQN20, ZX22]. replicated [CWJC20, Sir24, VVHK24]. report [KS22]. reporting [MC20]. Representation [SJF21, Cav20, KB23b, MGK22, PZAH22, YXP24]. representations [BG21]. representative [XFP23]. Reproducible [YLLZ24]. reproducing [WS22, TLRB21]. requirements [ZW22b]. resampling [WY20]. research [CWK23, CW24b, Zha20a, ZZZ22]. reserving [GZ21]. Residual [BR23, ASKB23, ASS23b, CG24b, DAS^{+}20, FM24, FTAL23, GK20a, JRN24, JWWW24, KB23a, Kun23, Moh22a, NSSR23, NSG23, Nou22, PK20b, PK22c, RG24, Sha21, SKS24, Tah20, WS20, ZAG24]. residuals [GX21]. resolution [CL21c, LaM24]. resolvable [GPM23, SS22f, SSS23]. resonance [NMT22]. respect [LY22b, dZySgYW21]. respondent [CN23a]. Response [HHW21, ASG24, AS23a, Arn23a, Arn23b, ASI^{+}24, BBP22a, BP20, BZ20, CN23a, CRVK24, CP24b, CLZ24b, EM21, ES22, FHHW20, HSV^{+}22, HCH20, HSC22, LKA^{+}23, LGZ22, MLW24, MMLBP20, NSS24, NS21b, NS21c, NS22a, OMYF21, SSS20, SHC20, SC22, SK20f, SBB21, SS22k, ST23b, Sma21, Sma24, Sun23, SF24b, ÜK21, Wan22b, YS20, ZG23, ZL22d]. Response-based [HHW21]. responses [KSA23, RSG21, SSG21, SPS23]. restart [Vid22]. restoration [SMZ23]. Restricted [Ema20, KE24, TÖ23, CG23, DMD23, GG21, Ham23, Hu20b, MKÖ23, SBK24, Wu22a, ZWLF24]. restriction [HK20, JT24, KP24]. restrictions [Baz24a, CN23b, FSS22, GÖ24, Hu20b, HW22d, YRB20, YA22]. restrictive [YM22]. resubmission [AAAJ24]. resubmission-based [AAAJ24]. resubmitted [MJB24]. result [AAGE23, HO23, SY23]. Results [Goo24, RSR24, APC23, BA23, BABF22, BS20a, BD22, CMO^{+}23, CB24b, EAAC22, GG23b, HFGZ24, JV24, KK23a, KB22b, LX22a, NT22, Nag20, NZZ20, NS21d, NN24, ND21, PK20b, Pos24, SB22, Tsa22, UN23, VVS^{+}24, WLL23, YSF22, vdHvDZ22]. Retention [LBG23]. retesting [HW20a, HWH22]. retirement [HT23]. RETRACTED [Ano21b, PS23c]. Retraction [Ano21c, Ano23c]. retrial [Gao23, HLPL22, JRD24, KLSZ22, LAAA^{+}22, LHWK24, WLZ^{+}23a, XLW23, ZLCW20]. return [Li24, LG24b, NCC24, SWHW22, XWW24, YCWZ22]. returns [BPZ22, JPZ23, QQS^{+}22, SYL23, XZR^{+}20, ZPJY24]. Revealing [Kur22]. reversed [BBAA21, BKB22, FAN21, Han24, Lee23, NSG21]. reversion [WC20c]. reverting [HN23]. review [Ale23, HSV^{+}22, KP21, Lee22b, LEM24, NCM22]. Revised [HT24]. Revisit [XXZ24, WD20, Wu22d]. revisited [GC22a]. reward [CDP20]. Reweighting [QMS22]. RHR [BAA24]. Ricci [EW23]. Ridge [JRCB23, ZS23, AAUA23, AAR22, CWK23, DC24, EA21, EA23, GCFS23, Moh22b, OO20, ÖA23, Sir24, Wu22a, Wu23a, YA22]. Ridge-GME [JRCB23]. ridge-type [AAUA23]. Riemann [FE23]. Riemannian [Gzy22, HR23]. Riesz [DS23]. right [AZAZ24, ACCM24, DBBK21, Don23, JMSG24, LC22b, NBM24, PR21, SVR21, sS20b, SLX22, hSsXfX24, SZ20b, WW22, Wan22a, WS20, XO21, YD22, ZBB24]. right-censored [ACCM24, sS20b, SLX22, SZ20b, WW22, Wan22a, WS20, XO21, ZBB24]. right-censoring [PR21]. Risk [CH20b, KWC21, XWYZ23, AV20, AMM20, AG23, BPZ24, Baz24b, BFPM22, BL23, CMO^{+}23, CLC22, CW20a, CH21a, CWCY23, CM24, CY20, CX22b, DHH^{+}23, DCY20, DLC22, EJJ23, FW23, FLW24, GLK24, GQW21, Han20, Han21a, HW24, HT23, JPZ23, JPJB22, KdXYC23, Ker20, KM24b, LLW23, LYC20, Li23b, Li24, LL24a, Lin21, LGL20, LHXZ22, LG24b, Lou20, MGK22, MGK23, MÁH22, MB20, MKÖ23, Nak21, NZZ20, NCC24, PAdlPK22, Pen22, Psa22, QQS^{+}22, RYZ24, SYL23, SK20a, SS23c, SSW20, SHXW23, SGW23, Tan24, TPF20, TS22, The22, Ünv21c, WM21, WQSG22, WCY23, WH23b, WDLW20, XXZ24, dXG20, XWW24, YW24, YMZD23, ZSCH22, Zha21a, ZZRH22, ZCZ20b, ZQX23, ZWLG22, ZY23, ZHS^{+}24, ZPJY24]. risk-adjusted [BPZ24]. Risk-based [XWYZ23]. risks [AGHK20, BSN22, BP24, DV22, GKvZB21, HL22a, QG23, VE21, WYG24, ZW24a, ZLNS22]. risky [RYZ24, YZSS23]. rival [CP24b]. RLM [LG24a]. RLRT [PK22c]. RMDs [BAGR22]. ROBEM [ÖB21]. Robust [AAHS21, AAKL21, AKUT23, BZ23, BN22, CSZ20, CLZL24, DLHZ21, EA23, HYZ24, HR23, HX23, bJwWyX23, KBMD23, KF24, LO23, Mah24, MdWG23, NH20, PK20a, PS21, SS22b, Sha21, SBB24, SF24c, SL24b, Syr23, WRZW21, XL24, YCWZ22, ZM24, ZW24a, ZC22, AHS^{+}24, AP24, Bey23, Bul23, DMD23, ES22, GB23b, HX22, KKS24a, KHSS20, Kun24, LC22b, MNA22, MBSS23, Oh23, ÖB21, SE22, SAH21, SRSB20, Wu21, YP24, YLJ23, YLW24, YZ21, YH24, ZB20, ZB23a, Zey20, ZG23, ZXYJ24, ZW24b]. Robust-regression-type [AAHS21]. robust-type [ZB23a]. robustified [Cha24]. Robustness [ON23, YO23, ZY24a, AA22a, ÇA20, FHHW20, Ham24, KH20d, Kur24]. ROC [ROSL21, SZ20b, Wu20]. rolling [Pra21]. root [DP22, KPH^{+}24, WTS24, XL24]. roots [TG24]. Rosenblatt [AGR21, Gai22, Ten24]. Ross [CDMR24, DMR22]. rotatability [OC22a]. rotatable [KBMD23]. rotation [RMS^{+}24, SPSS20]. row [HW23b, LW20, LOL21, LWYL23, LM22]. row-wise [HW23b, LW20]. rowwise [CCW^{+}20, FWWH21, LM23b, SW24, jWWDjW24]. RS [BNAK23]. Ruin [CLZZ21, DHH^{+}23, RS24, AD21, Baz24b, Baz24a, CW20a, CWCY23, CY20, CX22b, JPZ23, JPJB22, LYL22, Li24, Lin21, LCF24, MGK23, Psa22, SYL23, SGW23, TB22, WM21, WQSG22, XZ24, XWW24, YZW23, ZDZ22, ZPJY24]. Ruin-related [DHH^{+}23]. rule [Bar22, KE24, MMRMG20, ZXSZ20]. rules [DPP24, KRCS24, MMSC22, RGS24, SMMCM20, SMMC21, Ünv21d, YSF22]. Run [AKST22, LEM24, DPP24, MMG24, NNS20, Pou21, QSCH22]. run-length [MMG24]. runs [Ino24, KU22, KRCS24, MMRMG20, MMSC22, PP22b, SMMCM20, SMMC21]. runs-rule [MMRMG20]. runs-rules [MMSC22, SMMCM20, SMMC21]. RVaR [ZM24].

S [ZL22c]. Saddle [New23]. Saddle-point [New23]. Saddlepoint [KERAE23]. safe [BAHB23]. safety [VM22]. salary [NCC24, YCWZ22]. same [GRBL24]. Sample [AMBK22, DK23, Neg24, WZA20, WZA21, AWM21, AÖK^{+}24, Ano21b, AKUT23, BY24, BNO22, BEAA21, BHB22, BS22b, Bul23, CL23, CMP21, CG24a, Che22, CKKY22, CMM24, CGLP24, CGL21, Das23, Dou22, HHKC23, HL22a, Hua23, Ish20, KWB23, sK20b, sK22a, sK23a, KC23, KS22, Li21b, LX22a, LMLS23, LLG24, LK22c, Mak24, McI23, Mor24, Mot23, PR21, PKP23, PKK22, RG21a, RGS24, SYRV20, SP24b, SD20, SKRR23, TH24, Wan22b, Wij24, YM24, Yu22c, YLC^{+}22, ZG21, ZCCT22, ZN24]. sampled [RTT23]. sampler [Lee22b]. samples [BC23b, CG24b, DJW^{+}20, GRBL24, Han20, JCY22, JRM21, KK23a, Koh20, LXW20, LDZ^{+}22, PMMK24, QE21, RR20, Ros22, TQ24, TH23, YFW^{+}23, ZZ21, ZLF24]. Sampling [JG23, PF20, WJC23, AS22a, AN23, AYS21, AMAA^{+}22, ACKL23, Ami23, AuAH22, ABJR22, AuAHS23, AKSL22, BY24, BMT20, BCA22, BAAJ20, BA24c, BK22, BKS23a, BAJ20, CG24b, CS20, CKKY22, DSB22, DZ20, Fre20a, GP20, GAG24, GC24b, HA22, HAK23, Hu22, HCL21, JB22a, KR24b, KOTAG21, KRS20, KM24b, KL22a, KK24, LS22a, LS22b, LHJ21, MLW24, MSEB23, Mot24, MJB24, NSS24, NN21, uATH21, OIF21, Pri20, PKT21, QE21, QMS22, QR24, QH21, RSR22, RKS22, RAS^{+}22, SSS22, SYRV20, SP24b, SGN22, SGM22, SANHS21, SMMCM20, SSNP20, SK20f, SBB21, SYG21, SV21, SV22, SJP23, SGTG24, Sud22, STC23, SKP23, TH23, Tel23, TBCB24, WS22, WWJ23, YP24, YMG24, YT20, ZB20, ZK21, ZB23a, ZAG24, ZY20]. Santos [AE23c]. satellite [MAT24a]. satisficing [WL22b]. satisfying [LF23]. saturated [PZZ20, PYZS23, ZFDT24]. Saunders [DGCL23, OCSV23, SLV^{+}20, VVS^{+}24]. Scalable [YLW^{+}23, ZWYL22, FSB22]. scalar [MBSS23]. scalar-on-function [MBSS23]. scale [AP24, AGB21, BS22b, CS22, DK24a, FZD24, GT22, GZ21, GC22c, HW24, HCX20, JT24, KP24, MN23a, NZZ20, NCNM23, PKK20, PZY24, Sla23, TA22, TPM22, ZSNX23]. scale-free [HCX20]. scale-scale [GC22c]. scaled [MN23c, uA21]. scales [AEY21, Mur22, YGAK23]. scaling [Mur22]. scan [CS24]. scans [Ino24]. scenario [ZSCH22]. schedule [BC24, DKD22]. scheduled [Che23d]. scheduling [Che24a]. Scheffé [JZC24, SG22b]. Schematic [PYZS23, HT24]. scheme [AAM22, BLZ23, FTTY22, Hu22, MK20, SSA23, STK24, SANHS21, SKP23, ZPL23]. schemes [AuAHS23, MMRMG20, MMSC22, MMG24, SMMC21]. Schott [Yam24]. Schwarz [Far24b]. sciences [AAQ23]. score [CXT^{+}22, CBB^{+}21, CDCC24, Fab21, Lab23, RJ23, Wan21a]. scores [CG21, O’G23, ÜFR21]. scoring [Bar22, Zen20a]. Scott [NAJ22a]. scrambled [SSS20]. screening [DWZZ24, LWW24, LX22b, LC22b, LCW21, ON23, SZW21, Yu22a]. SDEs [Rao24]. Seamless [QLZZ20]. search [CTEB20]. searches [Ünv21b]. Searls [SV21]. Seasonal [HPW24, AG23, Ami23, TG24]. seasonality [BM23]. seasonally [CGLP24]. seat [Tap22]. seat-share [Tap22]. SEATS [CGLP24]. Second [CBJ20, Lin20, Lin21, Che22, DK24a, EG23, GL21, Gao23, Gon21a, HBH24, JZC24, KBMD23, LWT24, RK21, SS22a]. Second-order [CBJ20, DK24a, EG23, JZC24, KBMD23, LWT24]. secondary [ZZZ24]. secretary [WL22b]. sectional [WW23b]. sectoral [QQS^{+}22]. securities [GQW21, ZWLG22]. see [ZX23]. seemingly [HW22b, LX22a]. segmented [KK21]. selected [JMK24, MN23a]. Selecting [BCH24, OMYF21]. selection [AS23b, AS22b, AGB21, BL23, Cha23f, DHJS23, Dün21a, EM21, Fu24, GLLL23, HO24, HSG21, HZ22, HKL21, sK20c, KH20d, Kur24, LMH21, LTTL24, Liu24b, LW21b, Oku23, Ory20, RKK^{+}21, RO23, RZ23, RGS24, SS22b, SCAK23, SS22g, ySLZ22, SWHW22, SMP^{+}24, SZGK21, SL24b, WYW24a, WXW24, WY23, WLT21, XZZ22, XWCT21, XS21, YY21, YCC23, YW20, YWL20, ZK24, ZGC22, ZX23, ZGCZ23]. selector [BS20b]. self [Che20b, CS24, DJW^{+}20, Li23a, SZL21, YGG24, Yu20, YCS23, Zha20a]. self-exciting [Li23a, SZL21]. self-intersection [YCS23]. self-normalization [CS24]. self-normalized [DJW^{+}20, Zha20a]. self-reliant [Che20b]. self-repelling [YGG24]. self-similar [Yu20]. SEM [Che24b]. Semi [LKS22, LY21, Liu22, WBM^{+}23, BABF22, CDP20, Don23, HL22a, LK22a, LWS24, QSP^{+}24, ZLNS22]. Semi-blinded [WBM^{+}23]. semi-competing [HL22a, ZLNS22]. Semi-functional [LY21, BABF22]. semi-Markov [CDP20]. Semi-parametric [Liu22, Don23, LWS24, QSP^{+}24]. Semi-recursive [LKS22]. semi-strong [LK22a]. SEMIFAR [LBF24]. semimodules [LXZ20]. Semiparametric [BM21, DKA20, PP23, Yua21, ZLNS22, ARAK22, Ber23a, Bou21a, Ema20, EA23, FZ22a, HBH24, JLG21, JS23a, LHS23, Liu24b, LW21b, LYZ^{+}23, SWYF20, SL24b, SH23, VE21, WW22, YWL20]. semipartial [HC22]. semirings [LXZ20]. semivariograms [DD23]. semivarying [WJS22]. sensitive [AS23a, AAHS21, Arn23b, HCH20, KSA23, LHSK20, LS22b, LHS22, NS21c, NS22a, PS23a, PKT21, SMMCM20, SS22k, WKS21, YSS22b, ZIZ23]. sensitivity [HLPL22, ROSL21, SSS20, SHFD23, UMA22]. sentiment [Bic22c, CCMtH22]. separately [TLLL23]. sequence [DTDB22, qDSzW21, EBX20, HCCzW23, LLWA22, PY20, Raï24, RBSNB20, TZ21, TZL24, WZL^{+}22b, WWWZ24, YC23, dZySgYW21]. sequences [BN24, CP22, DCY20, Ino24, LZ23b, LM20, NTDB23, QH22, RS24, ST23a, XY21, ZSY21]. Sequential [Sal24, SB23a, Wu24b, BR23, CC22, FHL^{+}24, GWM21, HW20b, Hu22, JB22c, MR21, NS21d, NSS22, Nov24, QH22, SCP23, SMP^{+}24, Wan23c]. serially [LY22a, XZR^{+}20]. series [Alm24, AS22b, AGHK20, BZ23, BBH20, BD21, BmAB22, BKHB23, Ben22b, BC23a, BER23b, BB24, Cav24, CLPS24, Cha23f, CLWZ22, Che23a, CBM21, CGLP24, CW20b, GZZ21, GBH23, GA20, Gos22, HO24, HM21, JL22b, KR23b, KK23c, Kwo22, LYL22, LWY^{+}23, LL23a, LT24, LKK24, MBA20, NGC22, PJB24, PTYZ23, SHB23, SG23, Sha21, SPZ24, SS24b, SSW20, Tel23, WLL23, WLLB24, XZ24, YZZ21, Yan23, ZX22, dCBV21]. served [SK22b]. server [BC21, BTH23, ERAS22, GPD21, KC20, KLSZ22, LAAA^{+}22, LBG23, PG22, SA21, SCGB24, SK22b, WLZ^{+}23a]. service [BC24, Gao23, KC20, LBG23, MR21, NP23, PG22, SCGB24, WRZL21, XLW23]. services [BC21]. Set [CD24, JG23, AS22a, AuAH22, BC24, BK22, BN22, CG24a, CG24b, Che23c, DZ20, Fre20a, GC24b, HCCD24, JRM21, KK24, Lee22b, uATH21, QE21, QR24, RR20, SSG21, SV21]. set-inflated [SSG21]. set-theoretical [Lee22b]. set-up [BC24]. sets [RMS^{+}24, SXS24, Yan23, ZZZ24]. setting [DG21b]. settings [HWNN23]. setup [KC20, ZLCW20]. several [AA23b, CCG22, Han23a, KP24, QH23, Wu24a, Yan23]. severities [ATB24, FZD24, NZZ20]. severity [MGK23]. SGPII [RYJ20]. SH50ETF [HSWZ21]. Shafer [KTMKS20, LCYL24]. Shannon [DSY21, OIF21, PX23, RBD24]. Shannon-McMillan [PX23]. shape [AWL^{+}22, JBJ23, MN23a, PKK22, YT22, ZLF24]. shaped [LG22]. share [Tap22, YCW21]. shared [Han23b]. sharing [SGZA23, ZYJ20]. sharpened [Bic20, Bic22b]. sharpening [HJ22]. Sharpness [BC23b]. Shewhart [NMF23]. Shewhart-type [NMF23]. shift [LYY24, MMSC22, ZCT^{+}20]. shifted [CMM24, JT24, dCBV21]. shifts [RASD23, Tas24, ZLZ20b, ZYZ24]. Shiu [SSW20]. Shock [NSM21, DAZ23, Lee21, LNP20, MBLY21, MSW23, Pou20, Pou21, SS22d, SK20a, WLLB24]. shock-erosion [Pou21]. shocks [CC22, DYC^{+}24, KC21b, Li23b, PJB24, WLS23]. shoestring [Alv23]. Short [Ham24, NNS20, WW23b, WHCC23, XP21, ZL22b]. short-range [XP21]. short-term [WHCC23]. short-time [ZL22b]. shortfall [LM20, WGX21, ZSCH22]. shot [LCNB21, WHW23, ZY20]. showing [dCBV21]. Shrinkage [DM24, KSY20, ZBB24, INKS21, INKS23, Jun20, NAAQ22, Tan22b, ZL23b]. shrunken [SWW24]. Side [SMMCM20, HW22a]. Side-sensitive [SMMCM20]. sided [AO24, MMSC22, MGK23, QSCH22, RKS22, Zha21a, ZPJY24]. sign [ARA^{+}20, MG23, Zha22]. signal [Wu24b]. signals [Kou23]. signature [YF22]. signatures [NCM22]. Signed [GML20, JAE23]. signed-rank [JAE23]. significance [Bic20, Bic23, Bic24, O’G23, YL23]. Silva [AE23b]. similar [Yu20]. similarity [KK22, KK23b, YLC^{+}22]. Simon [sK24a]. Simple [Li21b, Mef23, Alg23, CXT^{+}22, GG23a, Gau21a, Ham24, Haq22, KSA23, KHSS20, Mot24, QR24, SGM22, SANHS21, SS21b, WXW24, WZ21, YMG24, ZB23a, ZZ21]. Simplex [Del21]. simplicial [KH21]. simulation [JC22, KOTAG21, ÖA23, WCT22, WL22b, eZN22]. simulation-based [WCT22]. simulations [JPJB22, KAAH21, SS21a]. Simultaneous [AWN21, CZ23a, HN21, JCY22, SMP^{+}24, Tas24, ZWCP22, UN23]. simultaneously [CMM24, TLLL23]. Single [JH24, BC21, BTH23, CXT^{+}22, ES22, GPD21, HMZ20, HL24b, KSK^{+}23, LAAA^{+}22, LY22a, LL23c, MHZJ23, MBLB24, NP23, SA21, SCGB24, SK22b, STC23, SLYF23, SF24c, WWZ22, WL23a, WLZ^{+}23a, WYW24a, ZS24, CMO^{+}23]. Single-index [JH24, HL24b, LY22a, LL23c, WWZ22, WL23a, WYW24a]. Singular [DGCL23, GZZ21, KW22, Li21a, Tan22b]. sinusoidal [Kun24]. situations [ASA24]. Size [AR24, AMBK22, AKUT23, CAAARCTS23, CKKY22, DK23, GAG24, GG24b, HHKC23, Ish20, sK20b, sK22a, sK23a, KC23, McI23, Mor24, NP23, PKK22, PG22, SYRV20, SKRR23, WZA20, WZA21, YSS22a, ZCZ20a, ZN24]. size-biased [KC23]. size-dependent [ZN24]. sizes [AMBK22, FW23, FLW24, HM22, HWH22, HAA^{+}20, KWC21, MMSC22, Nov24, PPM21, RNA^{+}22, WZA20, ZZY^{+}23]. sizes* [XWW24]. Skellam [CZL24]. Sketched [LXLL23]. skew [ALS21, Cel22, CRKC21, FMZM22, HMM20, JAJ21, JWWW24, KSY20, LXW20, NAJ22a, SCPAV21, TK20, WL23b, XL23a, XZL20, YFW^{+}23, ZCW23, EK22, HXCW22]. skew- [Cel22]. skew-elliptical [ALS21, EK22]. skew-normal [JAJ21, JWWW24, KSY20, LXW20, NAJ22a, TK20, YFW^{+}23, ZCW23, HXCW22]. skew-normal-Cauchy [CRKC21]. skewed [AKAB23, MN21a, The22, TGMS23]. skewing [AAR21]. skewness [AWM21, MAAZ22, SMOO23]. skip [MJB24]. skip-lot [MJB24]. skipping [SMMCM20]. slash [RBCG20, dC22a]. slashed [AIGB20]. sliced [Nko23, Tsa22]. slope [MQK23]. slopes [HA20]. Slow [KH20b]. Slow-explosive [KH20b]. Small [PR21, vRvR20, AMBK22, AC23, Ano21b, AKUT23, CN23b, DLZN23, HMP23, Jan20, KNCK24, KS22, Lig23, MN21a, MLWSPS24, Ros22, WSG21, Wei21, Wei23, ZLZ20b]. smallest [APC23, ADIK22, CKM23b, FZD24, HBH24, KCB23]. Smirnov [Fre20b]. smooth [BR20, FMM^{+}22, ZAF24]. Smoothed [ANO23a, ACCM24, LWYJ24, LSSW22, SLX22]. smoother [Yam20]. smoothing [AALE22, HBOR21, PTYZ23, SA22, hSsXfX24]. Smoothness [YCS23]. social [Gos22, MAT24c, MOAF23, SS20a, SS22i]. Software [DKD22, SS24c, UC21, Y.24, YZC22]. Solution [MR21, AW24, CGW23, LLH^{+}24, MAT24a, MB21b, SBK24, Sha21, WC20c]. solutions [HW22a, LF23, SSS23, ZT21]. Solvability [JL24a]. solve [ZHS^{+}24]. solving [MLW24, MNAM20]. Some [AA23a, AAQ23, ASS24, BA23, BAGR22, BKS23a, BD22, CSBM24, CP24a, CMM24, qDSzW21, EAAC22, HAA^{+}20, HGAJ21, Ko23d, LZH20, LX22a, LLS20, MSM20, Moh23, NSG21, NS21d, OA21, OMS24, PS23c, Pos24, RNA^{+}22, SS22f, SLF^{+}22, SSNP20, SK20f, SPS21, SPS23, SDK^{+}24, TW20, Tsa22, TKP22, WLL23, ZWT^{+}24, ZYY20b, AHO22, Bal22, BHB22, Baz24a, CF22a, CDL24a, CYB23, DCY20, GG23b, GH22, Gho23, GND^{+}23, Jaf22, KBMD23, iK22, KH20c, KB22b, MBR22, NT22, NSN20, ORH22, Onw21, QAO20, RO23, ST24, XZ24, XWH21, ZLPZ24, ZSYD24, Ano23c]. sorts [ADW22]. source [YZC22]. sources [WYJD21]. space [CYB23, DK24b, GW21, HN23, Ko23b, Ko23d, OYO24, QH22, UN23, WS22, eZN22, ZWLF24, TLRB21]. space-time [DK24b, GW21]. spaces [BZ21, DS23, Hu24, HW22d, KÇ21a, Ko20, Ko23c, NCNM23, TTL20, Tri24, WW20]. spacing [Kre24]. spacings [BHB22, BS20a, BS22a, Jas21]. span [YLL^{+}24]. Sparre [TB22]. Sparse [ZL23b, CGW23, Hua23, TZ23, YY21, YN21b]. Sparsely [GG21]. Sparsity [GL20]. Spatial [AMM20, AEBK20, AC23, EIAZ23a, BABF22, BS20b, DLT21, DD23, EM22, EGD23, FMZM22, HWZX23, LY21, LLQ22, Liu24b, LW21b, NMT22, PWM20, Qin21, RLQ23, SBC23, XWCT21, YLJ23, YCC23, ZJF23, ZGCZ23]. Spatial-nonparametric [EIAZ23a]. SPC [NNS20]. Spearman [YH24]. Special [DNP21, Li23a, SBZ21, ZH24]. specific [CP24b, Jun20, MU22]. specification [Sal24, YHH24]. specified [Mat23]. Spectral [Cav20, KB23b, EW24, HM21, Lee22a, ZCCT22]. spectrally [ZDZ22]. Spectrum [Zen20b]. speedy [RIA20]. spherical [OYA24]. spherically [OO20, PX23, SXZ24]. spiked [Ish20]. spillovers [KS24d]. splicing [MvSK20]. Spline [WW23b, Che23c, HBOR21, SE21, Yos23, ZYC24]. Spline-kernel [WW23b]. splines [JBSK22]. Split [AWM21, DZZ23a, JZ24]. split-plot [DZZ23a, JZ24]. splitting [CYB23]. spot [dXY23]. spring [GC24a]. SPRT [MG23]. Spurious [VSVVLRV22]. square [Che23b, DPP24, Del21, Hu24, KG20, RTT23, SMS21, SS22e, SYT24, XL24]. square-root [XL24]. squared [HYZ24, HZ21, Ory20]. squares [BM22, CZZ^{+}23, CZL24, Far23, JV24, Kun24, LaM20, LLX21, LL24b, LYZ^{+}23, MKÖ23, Nie20, Pop20, SWW24, Tsi24, WY20, WSG21, Wei23, YN21a, YD24c, ZZGL22]. SSEs [Par23]. Stability [BR20, Sud22, ZSL24]. stable [AP20, BK24, Han21b, LP20, LMM20, Mat21, Wei23, ZYS21, ZSL24, ZL22b]. Stage [LC22a, BL24a, CBJ20, HWH22, Kan20, KH20a, sK24a, LGLL21, LYY24, NS22a, PKK22, SSS22, SSNP20, SS22k, ST23b, SWW24, SZW21, Wu23c, Wu23d, Wu24a]. standard [DM22, EAS20, RTT23, WQSG22]. Standardized [SWHW22, PLLC23]. Standby [YZZ21, LWKQ20, RG21b, RG24, TT23, WXY21]. standbys [Gao23]. star [PKP23]. starting [BKB23, Yan22a]. State [CYB23, ABJR22, BA24c, HN23, JB22a, KS21, KSK^{+}23, Lef22, Li23b, PF23, Pou20, QG24, RAS^{+}22, SA22, WCY23, WLZ^{+}24, eZN22, dZySgYW21, ZX22]. state-dependent [Lef22, Li23b, WCY23]. Statement [Ano21c]. states [ZL22a]. statewide [ZZZ22]. station [SS22a]. Stationary [NP23, PG22, AH22, Aug21, BD22, CDL24a, EBX20, FLW24, GA20, GM20b, IW22, LL23a, LZ23b, LCF24, PCZ21, PP22a, Pos24, QH23, ST23a, SA22, YXZ23, ZT24]. statistic [Aga24, AuAH22, ARA^{+}20, Bul23, Cam24, Che20a, CKM23b, Gai22, HY23, HWNN23, uAK22, PP22b, PD22, SX23, TA22, Wir23, Wu22d, Wu24b, Yam24, ZL22c]. Statistical [AEY21, BSN22, DTDB22, EIAZ23b, FAN21, HWZX23, KRCS24, KS24d, LY22a, LG22, LTY23, Muk24, NS21a, NNS20, Ohy20, Ohy21, QG23, QG24, WT20, WZD21, WYG24, Yu20, ZG21, ZYW23, AK21, BA24a, CKKY22, EAB20, EG20, EBG22, Far21, GLB22, Gül23, KAAH21, KÇ21a, LC24, LG24a, RG23, Sav21, SS22i, TH23, TEÇK22, TW20, WWWZ24, XFP23, YGAK23, ZSYD24, ZLLZ23, DNP21]. Statistics [Ano23b, AE23a, AE23b, AE23c, AE23d, Cha23c, KS22, APC23, AZAZ24, ABXE22, AÖK^{+}24, AA22c, Aug21, BG22b, BEAA21, Bar24, BKM20, BN24, CP22, CZZ^{+}23, CC23, DK24a, DJW^{+}20, DT21b, EAAC22, FCCD20, GG23b, GM21, HBH24, HABE24, IHDR20, KB22a, KC21b, Kun23, LL24c, MKM24, MBR22, Moh22a, Ril24, WF23, WH23b, XM21, YDHY24, Yin24, ZSCH22]. Statistics-Theory [AE23a]. status [FSS22, LWK20, LZW23]. steady [KSK^{+}23, eZN22]. steady-state [KSK^{+}23, eZN22]. Stein [ALS21, DDM24, Gau21b, KE24, LY23, MN23a, UET24, ZZW^{+}24]. Stein-rule [KE24]. step [AV20, BSN22, CG23, CZL24, DR21, HS20, LJL^{+}21, SD23, WYG24, YCY22]. step-down [YCY22]. step-stress [BSN22, CG23, SD23, WYG24]. steps [MR21]. stepwise [GC22b]. Stick [Muk24]. Stieltjes [Fak23b]. stigmatized [HSC22]. Stochastic [BmAB22, BKB23, DK24a, FZD24, GBH23, HGY24, HW22c, KN22, LLXW24, LZW24, Mon24, PK22c, SBB22, SPZ23, SPZ24, WF23, ZLF24, AGP21, BBH20, BKB22, BFPM22, BM20b, BMH24, BS20b, CWM23, CM24, qDSzW21, DP22, EG23, FL20a, FP23, GLLL23, GÖ24, GG23b, Hom21, IKT22, JAJ21, JL24a, JPZ23, KBKK24, Kel22, KHVBRm22, KL22b, KHF22, Li20, LF23, Li24, LGL20, LG24b, LJ24, LMS22, MCL23, MBR22, MB21a, ME21, MB21b, MGC22, NT20a, NT22, OA21, PZY24, Rao23, Ril24, SM22, SMH22, SMH24, STW23, SYL23, Shi23a, SSW20, SB23b, UN23, Vid22, Wei21, WX20, XL23a, XWY21, XWW24, YCWZ22, YG20, YRB20, YA22, YMZD23, ZYJ20, ZZ21, ZSL24, ZLZ20a, ZT21, ZPJY24]. stochastic-monotonicity [LF23]. stock [BPZ22, GLB22, Hua20, Ünv21c]. stop [AHO22, AOS24, NSN20]. stop-loss [AHO22, AOS24, NSN20]. stopped [LOT22]. stopping [LL24a]. strategic [EJJ23]. strategies [DLHZ21, DYC^{+}24, ERAS22, LRWZ22, MNAM20, NCC24, SMMCM20, ST23b, WRZW21, WLZ^{+}23a, XZR^{+}20, Yan20, YZW23, YH23, ZS20a, ZDZ22, ZLCW20]. strategy [CF22b, GLLL23, LHJ21, LLL22, Li23b, LLL23, LLH^{+}24, MZR20, SGZA23, SBB24, SS23c, WCY23, WLZ^{+}24, XL24, YCWZ22, YCW21, YSS22b, ZN24, ZW22b]. stratification [GA22]. stratified [BKS23a, CRVK24, GP20, GAG24, HCH20, KC23, LHSK20, LS22b, LHS22, SSNP20, SBB21, SS22k, ST23b, ZB20, ZB23a]. strawberry [MKG^{+}21]. stream [BS21b, Bro22]. strength [JMM22, JMSG24, KTDW20, KJKC20, PD23, PZZ20, RSY21, RYJ20, ST24, SS22g, WZTW20, WYP21, XJ21, ZDY24]. strengths [Wan21b]. Stress [KJKC20, AV20, AFO21, BSN22, CG23, FTTY22, JMM22, JMSG24, KTDW20, PD23, RSY21, RYJ20, ST24, SD23, WZTW20, WYG24, XJ21]. Stress-strength [KJKC20, JMM22, JMSG24, KTDW20, PD23, RSY21, RYJ20, ST24, WZTW20, XJ21]. stripe [Oga22]. Strong [BG22c, BSR22, BER23b, DH22, DOV22, FH22, HW22d, JSB22, Nko23, hSsXfX24, WW20, Wu23b, YZXY24, YLL^{+}24, YC23, ZW22a, Zha24b, ZYY20a, ZSY21, dS22, ABL23, Bou22, CCR21, DSY21, qDSzW21, HB20, Hu20a, HCCzW23, HF24, Kim20, KCG21, KC23, LK22a, LF20, MAK20, OIF21, PX23, SLF^{+}22, SYM24, SZ20a, TQ24, WYJD21, ZL20, dZySgYW21, ZSYM21]. Strongly [NTDB23, ATJ^{+}24, BAGR22, DAS^{+}20, HNR^{+}23, Ish20, JBR^{+}24, KS21, SMH24, You23b]. structural [LHJ21, Mur22, NK23, PCZ21, YW24, ZW20]. structure [Cav20, CF22b, GL21, LY22a, LMLS23, MAT24c, Raï24, SX23, SBZ21, TT23, WGX21, WC20c]. structured [CMO^{+}23]. structures [GLK24, GC22b, JPZ23, JPJB22, LZ23a, LG24b, MS23, NKK^{+}24]. Stuart [RB21]. student [KM24a, AKUT23, Gau21a, MPS23, OV21, XL23a]. Student- [OV21]. studentized [Kob24]. studies [AMBK22, BS21a, HY23, Kan20, LZLL20, LWS24, WZA21, ZWCP22]. Study [Bou21a, Ben22a, Han21a, LXZ20, MMLBP20, NS21a, Nak21, NNS20, OHS21, ÖA23, PW23, PWM20, PB22, RG24, SN23, SRSB20, XJ21, Yam22, ZSYD24]. stutter [AYP21]. Sub [LS20b, CZ21, CW22, CW24a, CW24b, Din24b, DZ24, FWWH21, FH22, GZ23b, Hu20a, HW23b, HFGZ24, JW24b, KK24, LW20, LZ23b, LM22, LM23b, WW23a, WW24, Wu23b, ZW22a]. sub-critical [CZ21]. sub-linear [CW22, CW24a, CW24b, Din24b, DZ24, FWWH21, FH22, GZ23b, Hu20a, HW23b, HFGZ24, JW24b, LW20, LZ23b, LM22, LM23b, WW23a, WW24, Wu23b, ZW22a]. Sub-optimal [LS20b]. sub-ratio [KK24]. subadditivity [WH23b]. Subcritical [SBHY24]. Subdata [RZ23]. subdiffusion [ZZGL22]. subexponential [CX22b, FLW24, GLW24, SGW23, Tan24, WQSG22, XWW24]. subfractional [STW23]. subgraphs [YN21b]. Subgroup [YZT21, HDZL21, LHS20]. subgroups [FS22]. subject [CC22, DYC^{+}24, KC20, KLSZ22, sK20c, LZLL20, Tas24, WLS23, ZT24]. sublinear [GLL23, HW22d, JW24a, LF20, Zha24b]. subordinator [LP20]. subordinators [Gaj22]. subplot [DZZ23a]. subpopulation [AS23a, ASA24]. subpopulations [DDL20]. Subsample [CS24]. subsampling [JWWW24]. subset [EM21]. subspace [WS22]. subsystem [WLLB24]. subunits [New23]. successions [Kon22]. successive [Cha23b, DSB22, PKT21, SK20f, SBB21, SJP23]. Sudoku [KG20]. sufficiency [MP21a]. sufficient [TA22]. Sum [WX20, AKST22, BG22a, CWCY23, Kha20, Kha22, KMH23, Ko23d, LLS20, Par23, RTN24, SGW23, Tri21, WM22, dC22a]. sum-ruin [SGW23]. Summability [SP22]. sums [Bou21b, CH20a, CCW^{+}20, CLS24, CMMD20, DZ24, Din24a, DH22, Elm22, FWWH21, FH22, GWD24, GLW24, HYWZ24, JW24a, JW24b, KNL24, KdXYC23, Ko23c, LaM20, Lin20, LCF20, LW21a, LM23b, MWW22, NS21d, NSS22, SZ20a, TTL20, TTLT23, TZL24, Tan24, WW20, WW23a, WW24, Wu23b, XWH21, XY21, ZW22a, Zha20a, ZYY20a, ZYC23, dS20, dS24a]. Sungsu [KS22]. super [Kur22]. superadditive [CCW^{+}20, Din24a, NBM22, TTL20]. supercompensation [WIN20]. supercritical [LH22, LTW24]. superiority [SWW24, ZWCP22]. superpopulation [GA22]. superquadratic [SY23]. Supersaturated [DCT24, LLSZ22]. supervised [VM22]. Supplementary [Far23, Far24a]. suppliers [SWHW22]. supply [SS22d]. Support [ZXYJ24, Bic22a, BBP22b, Fre20b, Kou23, LZG20a, nLSZ24, NS22b, TP24b]. Suppression [HC22]. Sure [ZS23, Ass21, Aug21, rChZhN^{+}24, CDL24b, DT21b, KHZ24, LCW21, PP22a, SST22, TTL20, XWH21, YXZ23, dS20, dS24a]. surface [BBP22a, FHHW20, HSV^{+}22]. surplus [LS20a, LLL23, MGK23]. surplus-dependent [LLL23]. surrogate [IHDR20]. survey [Arn23a, ASI^{+}24, BMT20, CMP21, CP24b, GH22, HAK23, Kim21, MPM20, Pri20, SSS23, SGN22, Yo22]. surveys [Arn23b, uA21, ZIZ23]. Survival [CBLS21, SK20a, ADIK22, AWN21, BAY20a, CF22a, CP23, CP24a, CLZL24, FZ22a, FZD24, HP20, KM23a, sK20d, sK22a, KB22b, LZ20a, LWK20, LCW21, New23, ORH22, PR21, RO23, SR21, SN23, SAF24, SK24b, SS21b, TY21, WW22, XH23, ZL23b]. survivors [CBLS21]. Sv [Wij24]. Sv-plot3 [Wij24]. SVD [SA22]. SVD-based [SA22]. swap [Ünv21a]. switch [BR20, KSK^{+}23]. switch-off [KSK^{+}23]. switching [Cav20, CW23, FW22, GLLL23, GQW21, HSWZ21, HW24, LW22c, LMS22, MCL23, MLWSPS24, ONM^{+}24, SHXW23, SZW21, TY23, WWJ23, YCS20, ZWT^{+}24, ZSL24]. Symmetric [PF20, AAR21, AGR21, MvSK20, Mon24, OO20, PX23, PAC20, SCPAV21, SXZ24, Swa24, YWS22, YZT21]. Symmetry [Sal21, DG24, SYT24, SX23, ZFZ24, vHP23]. synergism [PM21]. synthesis [CCMtH22]. synthetic [LC24, PP22b, SMMCM20, SMMC21, ZLZ20b]. System [BQW22, NCM22, AYS21, BLZ23, CSZ20, CVDY21, Cha23b, CF22b, CLWZ22, CC22, Che23d, Che24a, DZ20, DAZ23, ERAS22, FAN21, Gao23, GDSJ22, Goo24, HLPL22, Kia22, KHF22, KJKC20, LBG23, LX21, LHWK24, LNXC24, LLXW24, LWKQ20, LX22a, MAT24a, NGC22, PS23b, Pou20, RG21b, RG24, SS22a, SB22, SLS23, SGZA23, SK24b, SACQ23, SCGB24, SK22b, TT23, VM22, WXY21, WWJ23, WLZ^{+}23b, WLZ^{+}24, XLW23, YF22, YZZ21, YW22, ZZZ22, ZYNX20, ZX22]. Systematic [ML23, SYG21, SV22, Tel23]. systems [AGHK20, BBH20, BR23, BD21, BmAB22, BKB22, BKHB23, BAHB23, BKB23, BFSN22, BZ20, BZ21, Cha23a, CKM23a, DYC^{+}24, FZL24, GBH23, HL22b, KN22, Kel22, KB22a, KHF22, KS21, KSK^{+}23, Li22, MiT24, ONM^{+}24, PD23, PJB24, SHB23, SPZ23, SPZ24, Tas24, TZ24, UE21, WC22, WLS23, WLL23, WLLB24, Y.24, YW22, Zal22, ZYJ20, ZY20].

table [MAAZ22]. tables [CN23b, Gzy21, SYT24]. tabular [Zen22]. tackling [GZ20]. Taguchi [LC24]. Tail [EK22, JL22a, KKKM21, WM22, BFPM22, GLW24, Lin20, Lou20, LLS20, MdWG23, WL23a, dXG20, ZCZ20b]. tailed [Baz24a, CWCY23, KJ24, Li23a, Lin20, Lin21, MPM20, MÁH22, TH24, ZC23]. tails [LCF20, Psa22, TGMS23]. Take [Han23a]. Tampered [SD23]. tandem [SS22a]. TAR [OV21]. target [CTEB20, LHS23]. targeting [JL22b]. Tauberian [MR21]. technical [FP23]. technique [AS23a, Arn23a, Arn23b, AW20, BM20a, CTEB20, CG20, CM20, GAG24, SV21, Ünv21a]. techniques [AUQ22, SC22, SK24b, STC23]. telecommunication [BAY20a]. temperature [WT20]. Tempered [BK24, AP20, LMM20, ZL22b]. temporal [ZZZ20]. term [CBLS21, WC20c, WHCC23]. terminal [Yua21]. terms [Cel22, DD23, HCL22, HHKL23, sS20b, VK23b]. Test [CAAARCTS23, CW20b, JKD20, Tan22a, WC20a, YN21b, AÖK^{+}24, Ano21b, AuAH22, BSN22, BS21b, Bro22, Bul23, CL23, CG20, CYW21, CZ23a, CLLD23, CS24, CS22, CMCA20, DJW^{+}20, FHL^{+}24, FTTY22, Gai22, GW23, GM20a, GRBL24, HZ21, HL22a, HW20b, HS20, IKT22, JL22b, KWB23, KH21, KS22, KS21, Le 21, LNWA22, LGX^{+}22, MKM24, Mak24, MG24, Moh22b, Naj22b, uAK22, O’G23, OFH23, QH23, RR20, RSA20, RJ23, RG21a, SEGMA21, SS20a, SD20, Skr23, SD23, WT20, WJL23, Wu21, Yam24, YHH24, YL23, YM22, YK20, Yu22c, YH24, ZCL24, ZZ21, ZHA20b]. Testing [CMP21, DG24, EAMH22, FMHP24, GD21, GW21, IM21, JZWT20, KR22, KK23c, LaM24, LTW21, Naj21, Nou23, RK20, SSH23, Tel23, Yan23, YKH20, ZW23, ZW20, ZFZ24, ZCZJ20, Bic23, Bic24, CZL23, CS23, Fur23, Gon21b, GZ24, HW20a, Hua20, HN21, IY23, KW22, KTP22, KX23, LC22a, LZ21, LCNB21, Nou22, PD22, SS23b, SWHW22, SM20, Sma21, SX23, Wij24, Yan22b, ZN24, Zha21b, Zhu22b]. Tests [Nes24, vHP23, AA23a, ATAAM20, ACMC21, ABT20, AA22a, ABJR22, BZLW24, BS22b, CWJC20, CBB^{+}21, CMM24, CMC23, DR22, DK23, Fre20b, JAE23, KERAE23, KM21, KSM^{+}21, KKS24b, KPH^{+}24, KASK24, KK23c, Li21b, LMLS23, Mao20, Mir23, New23, Nov24, PP23, QH22, RO24, RB21, Sal21, Sal24, SP23, SS24a, Shu21, Sma24, SZW21, Wu22c, YM24, Yu22a, ZWCP22, ZC22]. TEWMA [CKL23]. TEWMA-Max [CKL23]. text [CCMtH22]. th [EB21, Jas20, Jas21, PS22, YDHY24, ZQX23]. their [BZ23, BHB22, BAGR22, CP23, Das23, HABE24, JBR^{+}24, JRN24, KPH^{+}24, KS24d, Lin20, MAK20, Tah20]. theorem [CDP20, CW22, DT21b, DSY21, Dou22, GZ23a, Gül23, GZ23b, KNL24, LTW24, Mot24, PP22a, PX23, Pos24, Shi23a, SXZ24, Sil23, SA21, WS22, Xin23, XWH21, YXZ23, YXP24]. theorems [BEhF24, CW24a, qDSzW21, DLZN23, DH22, HW23b, HCCzW23, MMZ20, SP22, SLF^{+}22, SST22, TZ21, WC20b, WYJD21, Wu23b, XY23, YZ20, Zha20a, dZySgYW21, ZSYM21]. theoretic [BT22, KS24b]. Theoretical [HSV^{+}22, VVS^{+}24, JV24, Lee22b]. Theory [Ano23b, AE23a, AE23b, AE23c, AE23d, Cha23c, KS22, OSX^{+}16, WTT20, Ang24, BC23b, DK24b, DP22, HHLG22, JLG21, KW22, KTMKS20, KK22, KK23b, KK24, LCYL24, Neg24, WL22b, YP24]. there [Con24]. thinking [EJJ23]. thinning [CZLY22, KWY20]. third [RIA20, WWJ23]. third-generation [WWJ23]. thoughts [HSV^{+}22]. Three [WLZ20, ZFDT24, CAAARCTS23, CBJ20, GCCJO^{+}21, Hon20, KR24b, QT22, SS22k, TT23, ZWCP22]. three-arm [ZWCP22]. Three-level [WLZ20, ZFDT24, KR24b, QT22]. three-parameter [GCCJO^{+}21]. three-stage [CBJ20, SS22k]. threshold [BLZ23, CMM20, Fli20, Li23a, LWY^{+}23, SLS21, SF24a, SWLZ23, Tan22c, XL23a, ZLS23a, ZCZ20b]. threshold-based [BLZ23]. thresholds [CMCA20]. Thurstone [Ano24, YJZZ23]. ties [RJ23]. tight [PCZ24]. Time [LL23a, NCC24, XZR^{+}20, YCW21, AG23, ARF23, Alm24, AS22b, AEY21, AB22, ABRT23, BZ23, BL23, BM20b, BER23b, BB24, BD22, CTEB20, Cav24, Cha23b, CLPS24, CDP20, Cha23f, Che23a, CX22b, CW23, CGLP24, CW20b, DG21b, DK24b, Ekh23, FBNP23, GZ23a, GLLL23, GW21, GA20, Gos22, HO24, HM21, HW22a, HKL21, IM21, JC22, JPZ23, JPJB22, JL22b, KC20, KM21, sK23b, KK23c, KRCS24, KJKC20, LP20, LYC20, LZ20a, LWY^{+}23, Li23b, LF23, LT24, LZW24, LZLL20, LHS20, LCF24, Lou20, LKK24, MB21b, MLZ21, MBA20, NSG22, Nak21, NP23, uA21, PK20b, PTYZ23, SG23, ST23a, Sha21, SGZA23, SS23b, sSL20, STW23, SS24b, SK20a, Sla23, SSZC22, Sud22, SAK21, Tel23, VM22, WM21, Wan23b, XZ24, XWW24, Yan23, YSS22a, YGAK23, YCS23, YLLZ24, Yua21, ZW20, ZS20a, ZL22b, ZDZ22, ZWW24]. time [ZOD20, ZLCW20, ZY24b, dCBV21]. time-based [ZOD20]. Time-consistent [NCC24, XZR^{+}20, YCW21, Li23b]. time-dependent [GZ23a, JPZ23, Lou20, SGZA23]. time-horizon [GLLL23]. time-inconsistent [ZS20a]. time-to-event [sK23b, NSG22]. Time-varying [LL23a, BM20b, CW23, MB21b, ZW20]. times [ASKB23, CFS20, Che23d, CY20, FW23, FLW24, Gaj22, JRD24, KT24, KV23, LC22a, LYD24, PR21, RTT23, SCGB24, ZGQN20]. TL [Ünv21b]. Tobit [FL20b, Lou23]. tolerance [WCY23]. too [Con24]. tool [ZIZ23]. tools [FVCdC23]. topic [Gos22]. Topp [AAO21]. total [AS23a, ASA24, CZ21, HSC22, JBSK22, WY20]. Touchard [CLMA20, Pui24]. toxicology [BBP22b]. traces [PH21]. tracking [SS22g]. trade [Raï24]. trade/no [Raï24]. tradeoff [JC22]. training [WIN20]. transect [BCA22]. transfer [MB21b]. transform [AHO22, DS24b, GG23b, JRN24, LLH^{+}24, Yam22]. transform-dual-asymptotic [LLH^{+}24]. transformation [BMK21, CX24, CFS20, CZZ^{+}23, CH21b, CGLP24, DL22, FC22, HWNN23, Li21a, LZ23c, Mur22, NT20a, PRZ23, QMS22, SH23, ZL23b]. transformations [BCKS22, Gri24]. transformed [GB23a, XJ21]. Transforming [LG24a]. transforms [AOS24, NSN20, WHMA22]. Transient [KSK^{+}23, Wu24b]. transition [BR20, GC22b, LWT24, QSP^{+}24, QG24, VK23a, WC20c]. transitions [Yua24]. transitivity [LZG20a]. transmission [LAAA^{+}22]. transmuted [CP22, GKV21, NT20a, SRA20]. transmuted- [NT20a]. treatment [BY24, CK21, Cha23e, CW23, FZ22a, sK22a, MWLS24, OK20, Wu24a]. treatment-control [OK20]. treatments [BCH24]. tree [DSY21, HCCM21, SYT21, SXZ24, WCT22, ZYY20a, ZSYM21]. trees [KE20, PX23]. trend [AS22b, Gon21b, IM21, MM20, Rao24]. trends [KWB23, WHCC23]. trial [sK20b, sK23a, QLZZ20]. triallel [SAP20, ST21]. trials [FMHP24, HCH20, IY23, sK20c, LHS23, LTTL24, SPS23, SMP^{+}24, SZW21, VVHK24, WBM^{+}23, WZA20]. trigonometric [HA20]. trimmed [Tsi24]. Trindade [AE23c]. trinomial [RN22]. triple [Ami23, DTDB22]. trivariate [FL20b, GT22]. Truncated [The22, AO24, ABJR22, FHL^{+}24, Gho23, GM21, HW20b, KERAE23, LZG20b, RKAA21, RM21, sS20b, SK24c, TH24, WZTW20]. truncation [AO24, Oga22]. Tsallis [Moh22a, RSR24, XD23]. Tukey [LC24]. tuning [WXW24]. Tuple [JG23]. Turkey [SK24b, Ünv21c]. Two [BEAA21, BL24a, CZL24, Che23e, CDSG20, Dag22, DMR22, HL22a, Kan20, KH20a, Mak24, ROSL21, ST23b, Sla23, SWW24, Sul21, ACMC21, AG23, AÖK^{+}24, ADW23, BY24, BC21, BP20, BZ20, BZ21, BCH24, CLZ24a, CLMA20, CP22, CK21, Cha23e, CLWZ22, CG23, CZX24, CY20, Che22, CMM24, CKM23a, CGLP24, Das23, DT24, DG21a, DHH^{+}23, DJW^{+}20, DDM24, DK23, DYC^{+}24, FMHP24, GRBL24, GWM21, HN23, HW22a, HL22b, HLPL22, Hua23, HAA^{+}20, HS20, JKD20, JT24, JCY22, Kel22, KKS24b, KTP22, sK24a, KJ24, LLZ20, LX21, Li21b, Li24, LL24a, LHWK24, LVB23, LXWY21, LGLL21, LMLS23, LYY24, LK24, LZ24, LK22c, MGK23, ME21, MJB24, NT20b, NS22a, NCC24, Ohy20, PPM21, PP23, PKK22, Pui24, RK24, RNA^{+}22, RTN24, RG21b, SSS22, SMOO23, SCPAV21, SPSS20, SSNP20, SK20f, SJP23, SMP^{+}24]. two [SZW21, TP24b, WSSY20, WD20, WRZL21, WXY21, Wan22b, WJL23, Wic20, Wu22c, Wu22d, WK23, Wu23d, XFP23, YM24, Yan24, YZSS23, YKH20, YXZ21, ZK21, Zei23, ZWLF24, Zha21a, ZC21, ZCZJ20, ZDW22, ZC22, dC22a]. two-arm [GWM21]. two-compartment [LK24]. Two-component [Dag22]. two-dimensional [CY20, Li24, LL24a, LVB23]. two-dissimilar-unit [WXY21]. two-level [CLZ24a, KKS24b, LLZ20, MJB24, ZC21]. two-occasion [SPSS20, SK20f, SJP23]. two-output [BZ20, BZ21]. two-parameter [CLMA20, CG23, Pui24, WK23]. two-part [CZX24]. two-phase [ZK21]. Two-piece [Sul21, SCPAV21]. Two-sample [BEAA21, HL22a, Mak24, AÖK^{+}24, Che22, CMM24, Das23, DK23, Hua23, Li21b, LMLS23, LK22c, Wan22b, YM24]. two-side [HW22a]. two-sided [MGK23, Zha21a]. Two-stage [BL24a, Kan20, KH20a, ST23b, SWW24, sK24a, LGLL21, LYY24, NS22a, PKK22, SSS22, SSNP20, SZW21, Wu23d]. Two-step [CZL24, HS20]. two-tailed [KJ24]. Two-time-scale [Sla23]. two-treatment [Cha23e]. two-type [Zei23]. two-variable [ZWLF24]. two-way [ADW23, FMHP24, WD20]. type [AV20, AAUA23, AZAZ24, ABXE22, AAHS21, APV24, AL24, AuAHS23, AAM22, BMT20, BmAB22, BS22b, BM20b, BMH24, Cha23a, CLZZ21, CMMD20, Din24a, Don23, EA21, EA22, FW23, Gho23, Gri24, HL22b, HS20, HWNN23, JMM22, JMSG24, KBKK24, KKS24a, Kim21, Koç21, Koh20, KG22, KS21, KSS21, LM23a, MdWG23, Mur22, MHS22, NSS24, NMF23, NSJ22, uA21, OHS21, OA21, RO24, RG21a, RTB20, SHA24, SSA23, SS22e, SGN22, SMH22, SWHW22, Ten24, The22, Tri21, TPM22, Tsa22, WYG24, WKS21, WJC23, Y.24, YFW^{+}23, Yu20, ZB23a, Zei23, ZG21, ZLF24, AFO21, AGHK20, AAM22, BSN22, BS22a, CG23, CGL21, DV22, FTTY22, Han20, LaM20, MK20, PMMK24, QG23, SD20]. Type-I [BSN22, CGL21, FTTY22, PMMK24]. type-II [AZAZ24, AAM22, JMSG24, KG22, RG21a, WYG24, AFO21, AGHK20, BS22a, DV22, MK20, QG23, SD20]. Type-XII [AAM22]. types [BZLW24, BC21, Che23e, DYC^{+}24, Kel22, LXWY21, LZ24, RK24].

Uhlenbech [HN23]. Uhlenbeck [CL21a, RYZ24, SB20, WSG21, Wei23, XZL20, YD24c]. ultra [DWZZ24, GL20, LC22b, LCW21, TZ22a, YY21]. ultra-high [GL20, LC22b, LCW21, TZ22a, YY21]. ultra-high-dimensional [DWZZ24]. Ultrahigh [ZS24, LX22b]. ultrahigh-dimensional [LX22b]. ultrastructural [KE24, Sir24]. UMVUEs [YTT23]. unbalanced [Fre20a, LaM24]. Unbiased [JZKH22, Loe22, SMV23, ARAK22, IK21b, KRCS24, NSS24, Wu22a]. unboundedly [DOV22]. Uncertain [Che23a, FHZC23, Tan22c, YL23, DTDB22, DZZ23b, GLLL23, HX23, KR24a, Kia22, LL24c, MN21b, NTDB23, SS24b, WLZ^{+}23b, YN21a, Yao22]. Uncertainty [Gar21, Bic22c, HCCD24, LL24b, LCYL24, ZD23]. unconditional [MMG24, RB21]. underdispersed [HK21, TIIB22]. underdispersion [dCBV21]. undesirable [KHVBRm22]. undirected [KS24a, LLWA22, Wan21b]. unequal [CG24b, CS20, HWH22, LS22a, LS22b, MN23a, PPM21, QE21]. Unified [Oga22, Bic22a, LMS22]. Uniform [CB24b, DM22, JG23, JPZ23, KHZ24, LOL21, LG24b, Lou20, SYL23, WQSG22, ZPJY24, BG22c, BA21, BER23b, Gül23, HW22b, NCGB24, OYA24, Tee22, YO23, YN21b, Yua24]. Uniform-in-bandwidth [CB24b]. uniform-negative [BA21]. uniformity [Nou22, Nou23]. uniformly [DSY21]. Unimodal [LL23d, GSED24, Kwo22, YZT21]. unique [WSSY20]. Uniqueness [Nie20]. Unit [KPH^{+}24, BCG24, DP22, RBS20, RBC^{+}23, SS24c, TG24, TT23, VM22, WXY21, WTS24]. units [TBCB24]. univariate [AP24, Haq22, HXCW22, MvSK20, SKCE23]. Universal [NK22, RASD23, Vid22]. unknown [BL20, DR21, Fli20, INKS21, MMRMG20, MMSC22, PT22]. unnoticed [Kur22]. unobserved [Liu22, NK23]. unrecognized [sK22a, sK23b]. unrelated [HW22b, LX22a, NS21b, NS22a, SC22]. unreliable [BC21, BC24, LAAA^{+}22, TZ24]. unreplicated [CLPS24, KKS24b]. unrepresented [Bic22c]. unresolved [CS23]. unrestricted [FMZM22, ZZ21]. unstable [CZ21]. until [KT24]. update [MDZN20]. updating [CF22b]. upper [ADIK22, LGL20, Sha22b, SLS21, XXBD23]. urn [JTS22]. USA [Ünv21c]. Use [GP20, SSS20, AS23a, CP24b, Ekh23, sK20c, MC23, SV22, WXY21, You24]. useful [RNA^{+}22, VRB22]. Ushakov [KT21]. Using [AdMSK23, GS24, LMLS23, AS22a, AAGN24, AFG20, AMAA^{+}22, AAHS21, AGHK20, AHS^{+}24, Ano21b, AuAH22, ARA^{+}20, ARAJ22, Asl23, ASS24, ASI^{+}24, BN21, BP24, BQW22, Baz24b, BBP22a, BD24b, BN24, CTEB20, CCMtH22, CP22, CG20, CW23, DE22, DBBK21, EW23, EJJ23, EIAZ23b, Fro23, GBGPNR23, Gos22, GB23b, HM22, HBOR21, HZ21, HCL21, HCH20, Jaf22, JMM22, JBJ23, KOTAG21, KE20, KSA23, KAC22, KR21, KL22a, Koç21, KHSS20, KS22, Lee20, LHSK20, LS22b, LHS22, LHJ21, LYY24, MMSC22, MU22, MR21, MBA20, MNA22, NSS24, NSG22, Nak21, NS21c, Neg24, NKK^{+}24, uATH21, uA21, uAK22, NuAMA22, Oga21, PYZS23, PP22b, Pri20, QSP^{+}24, RKK^{+}21, RK24, RFVR20, SYRV20, SS22d, SHFD23, SAH^{+}23, SGZA23, SA22, Sil23, SV21, SGTG24, TA22, Tsi24, UN23, ÜK21, Ünv21a, WHMA22]. using [Wir23, WWX22, WHCC23, XJ21, XTXW20, YF22, YZ21, YSS22b, Yo22, YKW23, ZB20, ZCW23, ZK24, vRvR20]. Usual [BKB22]. utility [Pen22]. utilization [RG24]. utilize [KR24a]. utilizing [AKUT23, CC23].

vacation [BC21, GPD21, KJ23, LX21, LYY24, NP23]. vacations [LBG23, LLXW24, LWKQ20, MR21, WLZ^{+}23a, XLW23]. Vajda [CH21a, CH22]. validation [SMV23, XS21, ZS24]. validity [HHKC23]. Valuation [ZLZ20a, ZQX23, ZWLG22, MCL23]. Value [PS23d, BPZ24, BFPM22, CCG22, Che21, DCY20, DK24b, KdXYC23, LGL20, PAAA20, SP23, Sha22b, Tan22b, Tan24, TPF20, Ünv21c, WRZL21, Yos21, ZDZ22, ZCE23]. value-added [WRZL21]. value-at-risk [DCY20, KdXYC23, Tan24]. valued [Alm24, BA23, CZL24, GML20, JSB22, JL22b, LWY^{+}23, LW22c, MKÖ23, Oga20a, PD24, RK20, Sal21, TTLT23, YK20, ZCT^{+}20, dZySgYW21, dCBV21]. values [ASS24, BG22b, Bic20, DT21b, GZZ21, HM22, HY23, KERAE23, Lav24, New23, PS22, VK23b, Wu22d]. valuing [LYLY21]. VAR [Ünv21c, MN23c, CH21a, HGY24, SL24a, ZDW22]. Varentropy [MBR22]. varextropy [Goo24]. variability [BHB22, Fab24]. Variable [AKSL22, LL23d, LW21b, XZZ22, XWCT21, YWL20, ZK24, ZGC22, AAAJ24, Che23b, CZX24, Che23e, CKKY22, GA22, HSG21, HA22, HZ22, KNCK24, KR24b, LMH21, Liu24b, Loe22, Nov24, Oh23, RO23, SS22b, SSS20, SD23, SL24b, WYW24a, WKS21, WY23, WLT21, Yan22b, Zen22, ZWLF24, ZX23, ZGCZ23]. variable* [PS23a]. Variables [Del21, APC23, ACKL23, AAHS21, ASKB23, Ang24, AA22b, BAA24, BBAA21, Bou21b, CH20a, CP22, CCW^{+}20, Che21, CW22, rChZhN^{+}24, CLS24, CW24a, CW24b, CLW^{+}22, CKM23b, CMMD20, DC22b, Din24b, Din24a, DH22, EG23, Elm22, EGD23, FWWH21, FH22, GD21, Gau21b, GWD24, GLW24, GWM21, GZ23b, HS21, HW23b, HFGZ24, HYWZ24, JB22a, JW24b, KNL24, KdXYC23, KOTAG21, KMH23, LW20, LZH20, LPFL23, LF20, LCF20, LLS20, LW21a, LM22, LM23b, LY23, MAK20, MN23b, MWW22, MWW23, MW24a, MHS22, NBM22, NJMM24, NTDB23, NN21, NSJ22, OMYF21, PKP23, PK22b, PS22, PS23d, Rao22, RTN24, Sal21, SP24a, SP22, SMOO23, SW24, SWK24, hSsXfX24, SYM24, SNS20, SPSS20, SZ20a, TTLT23, TZ22b, Tan24, WC20b, WM22, WW23a, WS23, jWWDjW24, WW24, Wu23b, XTXW20, XXBD23, XFW22, Yam24, YPD23, YZ23, YKW23, ZW23, ZW22a, ZL20, dZySgYW21]. variables [ZGC22, ZYC23, dS20, dS22, dS24a, dC22a]. Variance [AR24, LL23d, SSZC22, YT20, ANO23a, ASS23a, AdMSK23, AuAH22, AL24, ASI^{+}24, Bai21, BLC22, CKL23, CP21, CLLD23, DSB22, DLHZ21, DJW^{+}20, EK22, Far24a, FM20, GLLL23, Goo24, HA22, INKS21, LS22a, LLWL21, Mak24, MMLBP20, NSS24, Nko23, OYA24, Onw21, PAdlPK22, PK22c, PAAA20, PC22, RK21, Ros22, ST24, SAH^{+}23, SBB21, SMV23, Tas24, VRB22, WDH^{+}22, WTS24, Wij24, WLT21, XZR^{+}20, XFW22, YCW21, ZB23a, ZW20, ZL23a, ZZW^{+}24, ZW22b, ZWCP22, ZWFZ23, vdMGSM21]. variance-covariance [AdMSK23]. variances [GW23, NT20b, PP23]. variant [LHJ21, YG20]. variants [Cha23f, Con24]. variate [DGCL23, Wic20, ZAL21]. Variation [PF20, AFG20, AKSL22, CDL24b, CKKY22, GP20, JBSK22, RK24, RAS^{+}22, RuAD22, SYL23, dXG20, YMG24]. Variational [ZXLT21, Lee22b]. various [LZ23a, Mur22, NKK^{+}24, YTT23, ZL22a]. VARMA [ZCCT22]. Varying [AR24, YPXZ21, AB22, BM20b, CW23, DLT21, HM22, HCZ20, HWZX23, LXZ22, LL23a, LSY22, LCF20, LW21b, MB21b, SLYF23, TZYY22, WZA20, WZD21, WL22a, WWZ22, WL23a, XZZ22, XL22, Yos21, ZW20, ZGC22, ZWFZ23]. varying-coefficient [HCZ20, HWZX23, LW21b, WL22a, WWZ22, WL23a, XL22, ZWFZ23]. vascular [CW23]. Vasicek [CLP23, Guo21, HL23, HW22a, Yu20]. Vasicek-type [Yu20]. vecd [Nag20]. vector [AAGN24, HS20, HN21, JWW24, Kur22, nLSZ24, MSY23, PP22b, Sha23, Shu21, Spe24, TK24, Wic20, ZXYJ24]. vectors [ALS21, AA23a, CL21b, JAJ21, Ko20, Ko23b, Ko23c, Ko23d, Oga20a, TTL20, WW20, XY20, YXZ21, ZLW21]. version [Moh22a, PP22a]. versions [CP23]. versus [CP24b, HMP23, YZZ21, YW22]. vertices [Wan21b]. via [Ben22b, BL23, Bic22c, Bou21a, Che23c, CZ23a, CM24, CG21, DZZ23a, DWZZ24, DJW^{+}20, HHLG22, LLSZ22, LVB23, LCYL24, LY23, MiT24, NH20, Nes24, NCNM23, ÖA23, PF20, RS24, SS22b, SF24a, SK20a, Spe24, Tan22b, WjWbK22, XS21, Y.24, YHA20, YLW^{+}23, YLLZ24, ZDY24, ZS24, ZLLZ23, ZCZJ20, ZZZ24]. Viability [Shi23a]. view [JC22, TÖ23]. viewpoint [HC22]. visual [Bai21]. Visualization [Pri20]. Vitale [Arn23a]. VOCE [AR24]. Volatility [MBA20, AGP21, LJ24, LMS22, MCL23, SM22, SSZC22, SB23b, WjWbK22, XL23a, dXY23, YZ21, YMZD23]. Volterra [CM24, Wei21].

W [ZL22c]. waiting [FW23, FLW24, KV23]. Wald [RO24]. Walid [BAE23]. walk [ARF23, BL24b, GZ23a, HF24, KH20b, KKKM21, VK23a]. Walker [BM20b, Spe24]. Walsh [SLYF23]. Walsh-average [SLYF23]. WAPND [JSB22]. Warahena [AE23d]. Warahena-Liyanage [AE23d]. Waring [VRB22]. warm [Gao23, LWKQ20, WXY21]. warranty [Cha21a, LW23, SWHW22, WHW23]. warranty-return-rate-based [SWHW22]. Water [MNAM20]. Watson [CZ21, LB22]. Wavelet [XH24, ZSNX23, ABL23, BL20, CB24a, GS24, GA20, KCG21, KC23, KHZ24, MBA20, QQS^{+}22, SF24a, WSZ22, WDW24, Xu22]. Wavelet- [ZSNX23]. wavelet-based [QQS^{+}22, SF24a]. way [ADW23, FMHP24, WD20, vdMGSM21]. Weak [Bou21b, Hu20a, ZLH20, ZWWW24, Ass21, DPP24, EAB20, Jas20, LF23, MBLB24, NBM22, NJMM24, NSJ22, WC20c]. weak-run [DPP24]. weakly [ABL23, AKUT23, HAA^{+}20, HGAJ21, RNA^{+}22, RT23]. weather [MAAZ22]. Weaver [SC20]. week [MC20]. Weibull [Ano23b, AZAZ24, AFL15, BP22, CSZ20, CKM23b, FTTY22, GCCJO^{+}21, GND^{+}23, JKP20, KIA^{+}21, KJAZ24, LWK20, LX23b, NSM21, PB24, SHB23, SWHW22, YXS^{+}24]. Weibull-G [CKM23b]. Weibull-geometric [AZAZ24]. Weibull-life [SWHW22]. weighing [GC21, GC24a]. weight [Sav21, XWH21]. Weighted [BG22b, KERAE23, Kun24, LLZ24, Par21, Tah20, WJXF21, WZL^{+}22b, YCY22, AB24, AFO21, ARAK22, Alt21, BBL22, Bou21b, CP23, CP24a, CG24b, CLS24, CMC23, CDCC24, DC22b, DZ24, Din24a, FWWH21, FL21, FH22, GLLL23, GWD24, GLW24, GC24b, HY23, HCL21, HYWZ24, JW24a, JW24b, KNL24, Ko23c, LNXC24, Lin20, LY22b, LCF20, LW21a, LM23b, LYZ^{+}23, MWW22, NBM22, uA21, NuAMA22, PWM20, Pen24, QR24, RTN24, SN23, SAH^{+}23, SK24c, SZ20a, TTLT23, WY20, WW23a, WW24, Wu23a, Wu23b, XY21, XW23, YGG24, YA22, ZW22a, ZYC23, dS20, dS24a]. weighted- [LNXC24]. weighted-fractional [YGG24]. weighting [GCFS23, ZCL24]. weights [ASS23a, DESSM21]. Welch [Cel22]. where [Guo21]. whether [GRBL24]. which [HNR^{+}23, TP24b]. while [ASS23a]. white [WSZ22]. Whitney [Che20a]. Whittaker [Yam20]. Whittle [Aga24, Zen24]. whole [Dij23]. widely [JW24b, LGL20, LW21a, LM22, LW22b, LM23b, WC20b, WW23a, WW24]. Wiener [He23, LGX^{+}22, TEHEG22, YHH24, ZXLT21]. Wijsman [YGAK23]. Wilcoxon [Tri21]. Wilcoxon-type [Tri21]. Winters [LW22a]. Wintner [GLL23]. wise [HW23b, LW20, LM22]. Wishart [ABT20, NSH23, PH21]. within [Fab24]. without [LS22a, Mot24]. WNOD [ZYC23]. WOD [WS23]. wordlength [ZFDT24]. work [Bal22]. working [LBG23]. workload [BLZ23]. works [Cha23b]. Wrong [ST24, RO24]. WUOD [LCF20].

X [CGLP24]. X-13ARIMA-SEATS [CGLP24]. Xavier [AE23c]. XII [AAM22, BmAB22, Koh20, SP24a, ZLF24].

Yates [LLZ20]. Yates-order [LLZ20]. Yeo [CX24]. Yeo-Johnson [CX24]. yield [AN23, BNAK23, BAJ20, KT24, NN21]. yield-based [AN23]. York [ZZZ22]. Youden [AT22, Syr23]. you’re [BCKS22]. Yule [BM20b, KH20c, Spe24].

Zero [JB22b, Pho24, WX20, ADD22, Gau21b, GdNM20, KHD24, KR23a, Lee22a, MOAF23, NSG22, Naj22b, ND21, RKAA21, TKP24, TLN24, Wir23, XTXW20, ZBB24, vdHvDZ22]. zero-and-one [TLN24]. zero-and-one-inflated [XTXW20]. Zero-inflated [Pho24, ADD22, GdNM20, KHD24, KR23a, MOAF23, ND21, TKP24, ZBB24, vdHvDZ22]. Zero-one [JB22b]. Zero-Sum [WX20]. zerotruncated [RKAA21]. zeros [JB22b, MLSS24, PK22b, TIIB22]. zeta [LOT22]. Zhang [ZXSZ20]. Zolotarev [CC23].

References

Agami:2020:MPH

[AA20] Sarit Agami and Robert J. Adler. Modeling of persistent homology. Communications in Statistics: Theory and Methods, 49(20):4871–4888, 2020. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

               URL http://www.tandfonline.com/doi/full/10.1080/
               03610926.2019.1615091.

Arnroth:2022:REB

[AA22a] Lukas Arnroth and Rauf Ahmad. A robustness evaluation of Bayesian tests for longitudinal data. Communications in Statistics: Theory and Methods, 51(24):8754–8775, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

               tronic). URL http://www.tandfonline.com/doi/full/10.
               1080/03610926.2021.1906432.

Asgari:2022:EEP

[AA22b] Fatemeh Asgari and Mohammad Hossein Alamatsaz. An extension of entropy power inequality for dependent random variables. Communications in Statistics: Theory and Methods, 51(13):4358–4369, 2022. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2020.1813305.

Astaneh:2022:EPB

[AA22c] Bahareh Khatib Astaneh and Jafar Ahmadi. Estimation and prediction based on record statistics in the presence of an outlier. Communications in Statistics: Theory and Methods, 51(20):7038–7055, 2022. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2020.1870141.

Ahmad:2023:SCT

[AA23a] M. Rauf Ahmad and S. Ejaz Ahmed. Some correlation tests for vectors of large dimension. Communications in Statistics: Theory and Methods, 52(7):2144–2160, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1945631.

Arshad:2023:EPF

[AA23b] Mohd Arshad and Omer Abdalghani. Estimating a parametric function involving several exponential populations. Communications in Statistics: Theory and Methods, 52(23):8351–8370, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2022.2061999.

Arshad:2024:RBV

[AAAJ24] Asma Arshad, Muhammad Azam, Muhammad Aslam, and Chi-Hyuck Jun. A resubmission-based variable control chart. Communications in Statistics: Theory and Methods, 53(2): 574–586, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2022.2087880.

Alodat:2022:ADM

[AAE22] M. T. Alodat, Moh’d Al Fayez, and Omer Eidous. On the asymptotic distribution of Matusita’s overlapping measure. Communications in Statistics: Theory and Methods, 51(20): 6963–6977, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2020.1869260.

Atya:2023:NLR

[AAGE23] Shrief Prince Atya, Abdalla Abdel-Ghaly, and Rasha Ebaid. A new limit result in change point analysis. Communications in Statistics: Theory and Methods, 52(1):196–207, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1910843.

Ahmad:2024:CBM

[AAGN24] Hussam Ahmad, Mohammad Amini, Bahram Sadeghpour Gildeh, and Adel Ahmadi Nadi. Copula-based multivariate EWMA control charts for monitoring the mean vector of bivariate processes using a mixture model. Communications in Statistics: Theory and Methods, 53(12):4211–4234, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2176717.

Ali:2021:RRT

[AAHS21] Nasir Ali, Ishfaq Ahmad, Muhammad Hanif, and Usman Shahzad. Robust-regression-type estimators for improving mean estimation of sensitive variables by using auxiliary information. Communications in Statistics: Theory and Methods, 50(4):979–992, 2021. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2019.1645857.

Almanjahie:2021:REN

[AAKL21] Ibrahim M. Almanjahie, Mohammed Kadi Attouch, Zoulikha Kaid, and Hayat Louhab. Robust equivariant non parametric regression estimators for functional ergodic data. Communications in Statistics: Theory and Methods, 50(15):3505–3521, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2019.1705980.

Almanjahie:2022:NNS

[AALE22] Ibrahim M. Almanjahie, Khlood A. Aissiri, Ali Laksaci, and Zouaoui Chikr Elmezouar. The k nearest neighbors smoothing of the relative-error regression with functional regressor. Communications in Statistics: Theory and Methods, 51(12):4196– 4209, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2020.1811870.

Ateya:2022:PUA

[AAM22] Saieed F. Ateya, M. M. Amein, and Heba S. Mohammed. Prediction under an adaptive progressive type-II censoring scheme for Burr Type-XII distribution. Communications in Statistics: Theory and Methods, 51(12):4029–4041, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1808685.

Ali:2021:MGC

[AAO21] Zeeshan Ali, Azeem Ali, and Gamze Ozel. A modification in generalized classes of distributions: a new Topp– Leone class as an example. Communications in Statistics: Theory and Methods, 50(19):4548–4570, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1719419.

Akram:2023:NBE

[AAQ23] Muhammad Nauman Akram, Muhammad Amin, and Muhammad Qasim. A new biased estimator for the gamma regression model: Some applications in medical sciences. Communications in Statistics: Theory and Methods, 52(11):3612–3632, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2021.1977958.

Alodat:2021:PMS

[AAR21] Moh’d T. Alodat and Mohammad Al-Rawwash. A proposed mechanism for skewing symmetric distributions. Communications in Statistics: Theory and Methods, 50(11):2674–2695, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2019.1676441.

Amin:2022:BER

[AAR22] Muhammad Amin, Muhammad Nauman Akram, and Qasim Ramzan. Bayesian estimation of ridge parameter under different loss functions. Communications in Statistics: Theory and Methods, 51(12):4055–4071, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1809675.

Akram:2023:MRT

[AAUA23] Muhammad Nauman Akram, Muhammad Amin, Muhammad Aman Ullah, and Saima Afzal. Modified ridge-type estimator for the inverse Gaussian regression model. Communications in Statistics: Theory and Methods, 52(10):3314–3332, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2021.1970773.

Amimour:2022:LAN

[AB22] Amine Amimour and Karima Belaide. Local asymptotic normality for a periodically time varying long memory parameter. Communications in Statistics: Theory and Methods, 51(9): 2936–2952, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2020.1784435.

Aghamohammadi:2024:BWC

[AB24] A. Aghamohammadi and M. Bahmani. Bayesian weighted composite quantile regression estimation for linear regression models with autoregressive errors. Communications in Statistics: Theory and Methods, 53(8):2888–2907, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2150054.

Aslam:2022:MNN

[ABJR22] Muhammad Aslam, S. Balamurali, P. Jeyadurga, and Muhammad Ali Raza. Monitoring number of non-conforming items based on multiple dependent state repetitive sampling under truncated life tests. Communications in Statistics: Theory and Methods, 51(17):5806–5825, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1847294.

Allaoui:2023:MWE

[ABL23] Soumaya Allaoui, Salim Bouzebda, and Jicheng Liu. Multivariate wavelet estimators for weakly dependent processes: strong consistency rate. Communications in Statistics: Theory and Methods, 52(23):8317–8350, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2061715.

Araya:2023:PED

[ABRT23] Héctor Araya, Natalia Bahamonde, Tania Roa, and Soledad Torres. Parameter estimation for a discrete time model driven by fractional Poisson process. Communications in Statistics: Theory and Methods, 52(10):3452–3477, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1973504.

Alfelt:2020:GFT

[ABT20] Gustav Alfelt, Taras Bodnar, and Joanna Tyrcha. Goodnessof-fit tests for centralized Wishart processes. Communications in Statistics: Theory and Methods, 49(20):5060–5090, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1612917.

Alawady:2022:CGO

[ABXE22] M. A. Alawady, H. M. Barakat, Shengwu Xiong, and M. A. Abd Elgawad. Concomitants of generalized order statistics from iterated Farlie–Gumbel–Morgenstern type bivariate distribution. Communications in Statistics: Theory and Methods, 51(16):5488–5504, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1842452.

Anjoy:2023:SNH

[AC23] Priyanka Anjoy and Hukum Chandra. Spatial nonstationary hierarchical Bayes estimation of small area proportions. Communications in Statistics: Theory and Methods, 52(7):2161– 2181, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2021.1945632.

AlLuhayb:2024:SBR

[ACCM24] Asamh Saleh M. Al Luhayb, Frank P. A. Coolen, and Tahani Coolen-Maturi. Smoothed bootstrap for right-censored data. Communications in Statistics: Theory and Methods, 53(11): 4037–4061, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2023.2171708.

Alevizakos:2022:MED

[ACK22] Vasileios Alevizakos, Kashinath Chatterjee, and Christos Koukouvinos. Modified EWMA and DEWMA control charts for process monitoring. Communications in Statistics: Theory and Methods, 51(21):7390–7412, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1872642.

Alevizakos:2023:QMA

[ACK23] Vasileios Alevizakos, Kashinath Chatterjee, and Christos Koukouvinos. The quadruple moving average control chart for monitoring the process mean. Communications in Statistics: Theory and Methods, 52(9):2882–2916, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1962351.

Alevizakos:2023:EPE

[ACKL23] Vasileios Alevizakos, Kashinath Chatterjee, Christos Koukouvinos, and Angeliki Lappa. The effect of parameters estimation on the performance of variables control charts under repetitive sampling. Communications in Statistics: Theory and Methods, 52(7):2379–2401, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1952269.

Alabdulhadi:2021:NPI

[ACMC21] Manal H. Alabdulhadi, Tahani Coolen-Maturi, and Frank P. A. Coolen. Nonparametric predictive inference for comparison of two diagnostic tests. Communications in Statistics: Theory and Methods, 50(19):4470–4486, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1719157.

Aminzadeh:2021:BER

[AD21] M. S. Aminzadeh and Min Deng. Bayesian estimation of ruin probability based on NHPP claim arrivals and Inverse-Gaussian distributed claim aggregates. Communications in Statistics: Theory and Methods, 50(17):4096–4118, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1710763.

Ali:2022:CMZ

[ADD22] Essoham Ali, Aliou Diop, and Jean-François Dupuy. A constrained marginal zero-inflated binomial regression model. Communications in Statistics: Theory and Methods, 51(18): 6396–6422, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2020.1861296.

Amiri:2022:LUB

[ADIK22] Masoud Amiri, Jan Dhaene, Muhyiddin Izadi, and Baha-Eldin Khaledi. Lower and upper bounds for survival functions of the smallest and largest claim amounts in layer coverages. Communications in Statistics: Theory and Methods, 51(18):6385– 6395, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2020.1861295.

Albuquerque:2023:AIP

[AdMSK23] Pedro Henrique Melo Albuquerque, João Gabriel de Moraes Souza, and Herbert Kimura. Artificial intelligence in portfolio formation and forecast: Using different variancecovariance matrices. Communications in Statistics: Theory and Methods, 52(12):4229–4246, 2023. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2021.1987472.

Akhtar-Danesh:2022:HAC

[ADW22] Noori Akhtar-Danesh and Stephen C. Wingreen. How to analyze change in perception from paired Q-sorts. Communications in Statistics: Theory and Methods, 51(16):5681–5691, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2020.1845734.

Ananda:2023:HTW

[ADW23] Malwane M. A. Ananda, Osman Dag, and Samaradasa Weerahandi. Heteroscedastic two-way ANOVA under constraints. Communications in Statistics: Theory and Methods, 52(22): 8207–8222, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2022.2059682.

Azimi:2023:CGK

[AE23a] Reza Azimi and Mahdy Esmailian. Correction to: Ghosh, I. K., and Hamedani, G. G. (2018). The Gamma-Kumaraswamy distribution: an alternative to Gamma distribution. Communications in Statistics-Theory and Methods, 47, 2056–2072. Communications in Statistics: Theory and Methods, 52(18): ??, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2021.1999476. See [GH18].

Azimi:2023:CGS

[AE23b] Reza Azimi and Mahdy Esmailian. Correction to: Gomes, A. E., da-Silva, C., Cordeiro, G.M. (2015). The exponentiated G Poisson model. Communications in Statistics — Theory and Methods, 44:20, 4217–4240. Communications in Statistics: Theory and Methods, 52(16):??, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.2018464. See [GDSC15].

Azimi:2023:COJ

[AE23c] Reza Azimi and Mahdy Esmailian. Correction to: Oliveira, J., J. Santos, C. Xavier, D. Trindade, and G. M. Cordeiro. 2016. The McDonald half-logistic distribution: Theory and practice. Communications in Statistics — Theory and methods 45 (7):2005–2022. Communications in Statistics: Theory and Methods, 52(17):??, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1931336. See [OSX^{+}16].

Azimi:2023:CPM

[AE23d] Reza Azimi and Mahdy Esmailian. Correction to: Pararai, M., Warahena-Liyanage, G., and Oluyede, B. (2017). Exponentiated power Lindley Poisson distribution: Properties and applications. Communications in Statistics — Theory and Methods, 46, 4726–4755. Communications in Statistics: Theory and Methods, 52(18):??, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.2009871. See [PWLO17].

Almanjahie:2020:SLL

[AEBK20] Ibrahim M. Almanjahie, Zouaoui Chikr Elmezouar, Bachir Ahmed Bachir, and Zoulikha Kaid. Spatial local linear estimation of the L_{1} -conditional quantiles for functional regressors. Communications in Statistics: Theory and Methods, 49(23):5666– 5685, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2019.1620781.

Altin:2021:SBT

[AEY21] Yavuz Altin, Büsra Nur Er, and Emrah Yilmaz. Δ_{λ}- statistical boundedness on time scales. Communications in Statistics: Theory and Methods, 50(3):738–746, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1640880.

Abbaszadehpeivasti:2023:MMA

[AF23] Hadi Abbaszadehpeivasti and J. B. G. Frenk. On the method of moments approach applied to a (generalized) gamma population. Communications in Statistics: Theory and Methods, 52(11):3685–3708, 2023. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2021.1979042.

Antunes:2021:MAM

[AFFM21] Patrícia Antunes, Sandra S. Ferreira, Dário Ferreira, and João T. Mexia. Multiple additive models. Communications in Statistics: Theory and Methods, 50(19):4649–4655, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1723636.

Antunes:2023:BAM

[AFFM23] Patrícia Antunes, Sandra S. Ferreira, Dário Ferreira, and João T. Mexia. Bi-additive models for extremes. Communications in Statistics: Theory and Methods, 52(21):7543–7554, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2022.2051053.

Akomolafe:2020:EGV

[AFG20] Abayomi Ayodele Akomolafe, Funmilayo Adenike Fadiji, and Ezra Gayawan. Evaluation of geographical variation in livebirth registrations using Bayesian method. Communications in Statistics: Theory and Methods, 49(24):5883–5896, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1625921.

Aljarrah:2015:NWP

[AFL15] Mohammad A. Aljarrah, Felix Famoye, and Carl Lee. A new Weibull–Pareto distribution. Communications in Statistics: Theory and Methods, 44(19):4077–4095, 2015. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic). See correction [Ano23b].

Ahmad:2021:BPF

[AFO21] Abd El-Baset A. Ahmad, Mohamad A. Fawzy, and Hosam Ouda. Bayesian prediction of future observations from weighted exponential distribution constant-stress model based on Type-II hybrid censored data. Communications in Statistics: Theory and Methods, 50(12):2732–2746, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1667394.

Alencenovic:2023:BSD

[AG23] Alina Alencenovic and Andrius Grigutis. Bi-seasonal discrete time risk model with income rate two. Communications in Statistics: Theory and Methods, 52(17):6161–6178, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2026962.

Aga:2024:EES

[Aga24] Mosisa Aga. Edgeworth expansion of the t-statistic of the Whittle MLE for linear regression processes with long-memory disturbances. Communications in Statistics: Theory and Methods, 53(5):1760–1776, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2111525.

Aydin:2021:OBS

[AGB21] Celal Aydin, Necla Gündüz, and Jale Balibeyoglu. Optimal bandwidth selection for a kernel density with a location-scale property. Communications in Statistics: Theory and Methods, 50(7):1671–1684, 2021. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2020.1804586.

Amiri:2020:CRA

[AGHK20] Leila Amiri, Mojtaba Ganjali, Reza Hashemi, and Mojtaba Khazaei. The competing risks analysis for parallel and series systems using Type-II progressive censoring. Communications in Statistics: Theory and Methods, 49(22):5598–5612, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2019.1620779.

Albano:2021:ENL

[AGP21] Giuseppina Albano, Francesco Giordano, and Cira Perna. On the estimation of non linear functions in stochastic volatility models. Communications in Statistics: Theory and Methods, 50(2):387–399, 2021. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2019.1635700.

Araya:2021:NSR

[AGR21] Héctor Araya, Johanna Garzón, and Tania Roa. Non symmetric Rosenblatt process over a compact. Communications in Statistics: Theory and Methods, 50(23):5517–5529, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1734830.

Arezki:2022:LPS

[AH22] Ouerdia Arezki and Abdelghani Hamaz. On linear prediction for stationary random fields with nonsymmetrical half-plane past. Communications in Statistics: Theory and Methods, 51(15):5298–5309, 2022. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2020.1837880.

Ahmed:2021:NFL

[Ahm21] Mohamed Ali Ahmed. The new form Libby–Novick distribution. Communications in Statistics: Theory and Methods, 50(3):540–559, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03610926.2019.1639744.

Arab:2022:BHO

[AHO22] Idir Arab, Milto Hadjikyriakou, and Paulo Eduardo Oliveira. On the behavior of the high order stop-loss transform for convolutions with some applications. Communications in Statistics: Theory and Methods, 51(14):4638–4652, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1818101.

Anas:2024:CIB

[AHS^{+}24] Malik Muhammad Anas, Zhensheng Huang, Usman Shahzad, Tolga Zaman, and Shabnam Shahzadi. Compromised imputation based mean estimators using robust quantile regression. Communications in Statistics: Theory and Methods, 53(5): 1700–1715, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2022.2108057.

Astorga:2020:MSG

[AIGB20] Juan M. Astorga, Yuri A. Iriarte, Héctor W. Gómez, and Heleno Bolfarine. Modified slashed generalized exponential distribution. Communications in Statistics: Theory and Methods, 49(19):4603–4617, 2020. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2019.1604959.

Akkus:2021:GSM

[AK21] Ilker Akkus and Gonca Kizilaslan. Generalization of a statistical matrix and its factorization. Communications in Statistics: Theory and Methods, 50(4):963–978, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1645854.

Al-Kandari:2023:HSN

[AKAB23] Noriah Al-Kandari, Emad-Eldin A. A. Aly, and Lakdere Benkherouf. Harris skewed normal distribution. Communications in Statistics: Theory and Methods, 52(18):6597–6615, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2022.2032169.

Akdur:2022:GBB

[Akd22] Hatice Tul Kubra Akdur. GEE-based Bell model for longitudinal count outcomes. Communications in Statistics: Theory and Methods, 51(13):4602–4616, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2056751.

Ayyoub:2022:VSI

[AKSL22] Heba N. Ayyoub, Michael B. C. Khoo, Sajal Saha, and Ming Ha Lee. Variable sampling interval EWMA chart for multivariate coefficient of variation. Communications in Statistics: Theory and Methods, 51(14):4617–4637, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1818100.

Abubakar:2022:RSC

[AKST22] Sani Salihu Abubakar, Michael B. C. Khoo, Sajal Saha, and Wei Lin Teoh. Run sum control chart for monitoring the ratio of population means of a bivariate normal distribution. Communications in Statistics: Theory and Methods, 51(13):4559– 4588, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2020.1818099.

Asanya:2023:RBA

[AKUT23] Kenneth Chukwuemeka Asanya, Mohamed Kharrat, Akaninyene Udo Udom, and Emmanuel Torsen. Robust Bayesian approach to logistic regression modeling in small sample size utilizing a weakly informative Student’s t prior distribution. Communications in Statistics: Theory and Methods, 52(2):283–293, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2021.1912767.

Arvanitis:2024:IGQ

[AL24] Stelios Arvanitis and Alexandros Louka. Inconsistency for the Gaussian QMLE in GARCH-type models with infinite variance. Communications in Statistics: Theory and Methods, 53(5):1684–1699, 2024. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2022.2107665.

Alenazi:2023:RCD

[Ale23] Abdulaziz Alenazi. A review of compositional data analysis and recent advances. Communications in Statistics: Theory and Methods, 52(16):5535–5567, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.2014890.

Alghalith:2023:PBO

[Alg23] Moawia Alghalith. The price of the Bermudan option: a simple, explicit formula. Communications in Statistics: Theory and Methods, 52(9):3174–3177, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1969407.

Almohaimeed:2024:ANB

[Alm24] Bader S. Almohaimeed. Asymptotic negative binomial quasilikelihood inference for periodic integer-valued time series models. Communications in Statistics: Theory and Methods, 53(2):587–606, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03610926.2022.2087881.

Adcock:2021:SLG

[ALS21] Chris Adcock, Zinoviy Landsman, and Tomer Shushi. Stein’s Lemma for generalized skew-elliptical random vectors. Communications in Statistics: Theory and Methods, 50(13):3014– 3029, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2019.1678642.

Altun:2021:LWE

[Alt21] Emrah Altun. The log-weighted exponential regression model: alternative to the beta regression model. Communications in Statistics: Theory and Methods, 50(10):2306–2321, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1664586.

Alvo:2023:EBS

[Alv23] Mayer Alvo. Empirical Bayes on a shoestring and other applications. Communications in Statistics: Theory and Methods, 52(7):2228–2239, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03610926.2021.1948061.

Al-Marshadi:2022:MCC

[AMAA^{+}22] Ali Hussein Al-Marshadi, Muhammad Aslam, Abdullah Hamoud Alharbey, Nasrullah Khan, and Liaquat Ahmad. Monitoring customer complaints using the repetitive sampling. Communications in Statistics: Theory and Methods, 51(2):313–327, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2020.1748199.

Al-Mekhlafi:2022:SSP

[AMBK22] Amani Al-Mekhlafi, Tobias Becker, and Frank Klawonn. Sample size and performance estimation for biomarker combinations based on pilot studies with small sample sizes. Communications in Statistics: Theory and Methods, 51(16):5534– 5548, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2020.1843053.

Ahmad:2022:CCD

[AMH22] Zubair Ahmad, Eisa Mahmoudi, and Gholamhossien Hamedani. A class of claim distributions: Properties, characterizations and applications to insurance claim data. Communications in Statistics: Theory and Methods, 51(7):2183–2208, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1772306.

Amin:2023:GSB

[Ami23] Ayman A. Amin. Gibbs sampling for Bayesian estimation of triple seasonal autoregressive models. Communications in Statistics: Theory and Methods, 52(20):7303–7322, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2043379.

Abeidallah:2020:LLE

[AMM20] M. Abeidallah, B. Mechab, and T. Merouan. Local linear estimate of the point at high risk: Spatial functional data case. Communications in Statistics: Theory and Methods, 49(11): 2561–2584, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2019.1580735.

Afshari:2023:OCY

[AN23] Robab Afshari and Adel Ahmadi Nadi. An optimal construction of yield-based EWMA repetitive multivariate sampling plan. Communications in Statistics: Theory and Methods, 52(22):7819–7839, 2023. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2022.2050401.

Andersson:2020:CAC

[And20] Per Gösta Andersson. A classroom approach to the construction of Bayesian credible intervals of a Poisson mean. Communications in Statistics: Theory and Methods, 49(22):5493– 5503, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2019.1619768.

Angus:2024:FLD

[Ang24] John E. Angus. A fresh look at distribution theory for quadratic forms in jointly normally distributed random variables. Communications in Statistics: Theory and Methods, 53(3):1135–1142, 2024. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2022.2106374.

Anonymous:2020:Ca

[Ano20a] Anonymous. Correction. Communications in Statistics: Theory and Methods, 49(7):1792, 2020. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2020.1723326.

Anonymous:2020:Cb

[Ano20b] Anonymous. Correction. Communications in Statistics: Theory and Methods, 49(20):5114, 2020. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2019.1708402.

Anonymous:2021:C

[Ano21a] Anonymous. Correction. Communications in Statistics: Theory and Methods, 50(14):3489, 2021. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2019.1705348.

Anonymous:2021:RAP

[Ano21b] Anonymous. RETRACTED ARTICLE: The probable error in the hypothesis test of normal means using a small sample. Communications in Statistics: Theory and Methods, 50(1):??, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/abs/
10.1080/03610926.2019.1703135.

Anonymous:2021:SR

[Ano21c] Anonymous. Statement of retraction. Communications in Statistics: Theory and Methods, 50(1):249, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1854014.

Affossogbe:2023:SAV

[ANO23a] Mètolidji Moquilas Raymond Affossogbe, Guy Martial Nkiet, and Carlos Ogouyandjou. Smoothed average variance estimation for dimension reduction with functional data. Communications in Statistics: Theory and Methods, 52(3):806–829, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2021.1931330.

Anonymous:2023:CAM

[Ano23b] Anonymous. Correction to: Aljarrah, M.A., Famoye, F. and Lee, C. (2015). A New Weibull–Pareto distribution. Communications in Statistics — Theory and Methods, 44:19, 4077–4095. Communications in Statistics: Theory and Methods, 52(16):??, 2023. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2020.1822408. See [AFL15].

Anonymous:2023:RSP

[Ano23c] Anonymous. Retraction: Some Properties of the Generalised Autoregressive Moving Average (GARMA(1, 2; δ, 1)) Model. Communications in Statistics: Theory and Methods, 52(14): 5114, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2022.2155435. See [PS23c].

Anonymous:2024:NDP

[Ano24] Anonymous. Notice of duplicate publication: Asymptotics in the Thurstone Model with an increasing items. Communications in Statistics: Theory and Methods, 53(9):3350–3364, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2024.2350917.

Alodat:2022:ADF

[AO22] Tareq Alodat and Andriy Olenko. On asymptotics of discretized functionals of long-range dependent functional data. Communications in Statistics: Theory and Methods, 51(2): 448–473, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2020.1750653.

Akahira:2024:ALM

[AO24] M. Akahira and N. Ohyauchi. Asymptotic loss of the MLE of a truncation parameter in the presence of a nuisance parameter for a one-sided truncated family of distributions. Communications in Statistics: Theory and Methods, 53(21):7526–7540, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2023.2269436.

Alexandre:2024:BMC

[AOGM24] Armando Alexandre, Manuela Oliveira, Eugénio Garção, and João Mexia. Biadditive models: Commutativity and optimum estimators. Communications in Statistics: Theory and Methods, 53(6):2021–2033, 2024. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2022.2117560.

Aliev:2024:NTT

[AÖK^{+}24] Fazil Aliev, Levent Özbek, Mehmet Fedai Kaya, Coskun Kus, Hon Keung Tony Ng, and Haikady N. Nagaraja. A nonparametric test for the two-sample problem based on order statistics. Communications in Statistics: Theory and Methods, 53(10):3688–3712, 2024. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2022.2161310.

Arab:2024:CHD

[AOS24] Idir Arab, Paulo Eduardo Oliveira, and Beatriz Santos. Comparison of higher degree stop-loss transforms. Communications in Statistics: Theory and Methods, 53(21):7642–7650, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2023.2269453.

Arefi:2020:MMT

[AP20] Ahmad Arefi and Reza Pourtaheri. Multi-modal tempered stable distributions and processes with applications to finance. Communications in Statistics: Theory and Methods, 49(17): 4133–4149, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2019.1594304.

Anum:2024:HMD

[AP24] Andrews T. Anum and Michael Pokojovy. A hybrid method for density power divergence minimization with application to robust univariate location and scale estimation. Communications in Statistics: Theory and Methods, 53(14):5186–5209, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2023.2209347.

Abdolahi:2023:ORS

[APC23] Molod Abdolahi, Gholam Ali Parham, and Rahim Chinipardaz. Ordering results of the smallest order statistics from independent heterogeneous new modified generalized linear failure rate random variables. Communications in Statistics: Theory and Methods, 52(16):5606–5639, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.2013501.

Araya:2024:PEF

[APV24] Héctor Araya and Francisco Plaza-Vega. Parameter estimation for fractional power type diffusion: a hybrid Bayesian-deep learning approach. Communications in Statistics: Theory and Methods, 53(22):8234–8254, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2280522.

Abdulla:2024:AVA

[AR24] Mohammed Shahid Abdulla and L. Ramprasath. ANVILS-VOCE: ANova-based varying inner-loop size estimation of variance of conditional expectation. Communications in Statistics: Theory and Methods, 53(12):4442–4449, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2182158.

Aslam:2020:DSC

[ARA^{+}20] Muhammad Aslam, Muhammad Ali Raza, Muhammad Azam, Liaquat Ahmad, and Chi-Hyuck Jun. Design of a sign chart using a new EWMA statistic. Communications in Statistics: Theory and Methods, 49(6):1299–1310, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1563163.

Aslam:2022:NCC

[ARAJ22] Muhammad Aslam, G. Srinivasa Rao, Liaquat Ahmad, and Chi-Hyuck Jun. A new control chart using GINI C_{P K}. Communications in Statistics: Theory and Methods, 51(1):197– 211, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2020.1746971.

Akdeniz:2022:GDB

[ARAK22] Fikri Akdeniz, Mahdi Roozbeh, Esra Akdeniz, and Naushad Mamode Khan. Generalized difference-based weighted mixed almost unbiased Liu estimator in semiparametric regression models. Communications in Statistics: Theory and Methods, 51(13): 4395–4416, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2020.1814340.

Aliyev:2023:FPT

[ARF23] Rovshan Telman Aliyev, Fada Rahimov, and Aynura Farhadova. On the first passage time of the parabolic boundary by the Markov random walk. Communications in Statistics: Theory and Methods, 52(17):6078–6087, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.2024852.

Alotaibi:2021:BEH

[ARGD21] Refah Mohammed Alotaibi, Hoda Ragab Rezk, Indranil Ghosh, and Sanku Dey. Bivariate exponentiated half logistic distribution: Properties and application. Communications in Statistics: Theory and Methods, 50(24):6099–6121, 2021.

CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1739310.

Atashgar:2022:NHA

[ARK22] Karim Atashgar, Naser Rafiee, and Mahdi Karbasian. A new hybrid approach to panel data change point detection. Communications in Statistics: Theory and Methods, 51(5):1318– 1329, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2020.1760298.

Arnab:2023:DVR

[Arn23a] Raghunath Arnab. Dalenius and Vitale randomized response technique for complex survey designs. Communications in Statistics: Theory and Methods, 52(3):684–692, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1921211.

Arnab:2023:NRR

[Arn23b] Raghunath Arnab. A note on randomized response technique for multiple sensitive attribute from complex surveys. Communications in Statistics: Theory and Methods, 52(1):94–103, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2021.1910301.

Aryuyuen:2023:BIN

[Ary23] Sirinapa Aryuyuen. Bayesian inference for the negative binomial-generalized Lindley regression model: properties and applications. Communications in Statistics: Theory and Methods, 52(13):4534–4552, 2023. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2021.1995434.

Abbasi:2022:EPP

[AS22a] Azhar Mehmood Abbasi and Muhammad Yousaf Shad. Estimation of population proportion using concomitant based ranked set sampling. Communications in Statistics: Theory and Methods, 51(9):2689–2709, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1916529.

Alraddadi:2022:MST

[AS22b] R. Alraddadi and Q. Shao. Model selection for time series with nonlinear trend. Communications in Statistics: Theory and Methods, 51(20):7208–7224, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1871628.

Ahmed:2023:URR

[AS23a] Shakeel Ahmed and Javid Shabbir. On use of randomized response technique for estimating sensitive subpopulation total. Communications in Statistics: Theory and Methods, 52(5): 1417–1430, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2021.1928199.

Alabiso:2023:HDL

[AS23b] Audry Alabiso and Junfeng Shang. High-dimensional linear mixed model selection by partial correlation. Communications in Statistics: Theory and Methods, 52(18):6355–6380, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2022.2028838.

Ashutosh:2024:CES

[ASA24] Ashutosh Ashutosh, Usman Shahzad, and Nadia H. Al Noor. Calibration estimation of subpopulation total for direct and indirect situations. Communications in Statistics: Theory and Methods, 53(19):6997–7012, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2256437.

Aghayerashti:2024:BJM

[ASG24] Maryam Aghayerashti, Ehsan Bahrami Samani, and Mojtaba Ganjali. Bayesian joint modeling of binomial and rank response with non-ignorable missing data for primate cognition. Communications in Statistics: Theory and Methods, 53(10): 3758–3778, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2022.2163367.

Audu:2024:EFP

[ASI^{+}24] Ahmed Audu, Ran Vijay Kumar Singh, Olatunji Olawoyin Ishaq, Supriya Khare, Rajesh Singh, and Adedayo Amos Adewara. On the estimation of finite population variance for a mail survey design in the presence of non-response using new conventional and calibrated estimators. Communications in Statistics: Theory and Methods, 53(3):848–864, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2093906.

Amini-Seresht:2023:CRL

[ASKB23] Ebrahim Amini-Seresht, Maryam Kelkinnama, and Narayanaswamy Balakrishnan. Comparing residual lifetimes and inactivity times of dependent random variables. Communications in Statistics: Theory and Methods, 52(24):8738–8748, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2070764.

Aslam:2023:ICC

[Asl23] Muhammad Aslam. An insight into control charts using EWMA. Communications in Statistics: Theory and Methods, 52(5):1507–1511, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03610926.2021.1928705.

Assia:2021:ASC

[Ass21] Mezhoud Kenza Assia. Almost sure convergence of recursive kernel estimatiors of the density and the regression under η-weak dependence. Communications in Statistics: Theory and Methods, 50(17):3913–3927, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1710751.

Alam:2023:OCW

[ASS23a] Shameem Alam, Sarjinder Singh, and Javid Shabbir. Optimal calibrated weights while minimizing a variance function. Communications in Statistics: Theory and Methods, 52(5):1634– 1651, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2021.1937649.

Aswin:2023:BMR

[ASS23b] I. C. Aswin, P. G. Sankaran, and S. M. Sunoj. The bilinear mean residual quantile function. Communications in Statistics: Theory and Methods, 52(19):7115–7129, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2042022.

Aswin:2024:SRA

[ASS24] I. C. Aswin, P. G. Sankaran, and S. M. Sunoj. Some reliability aspects of record values using quantile functions. Communications in Statistics: Theory and Methods, 53(13):4755–4775, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2023.2189499.

Al-Sharadqah:2021:NPF

[ASW21] Ali Al-Sharadqah and Nicholas Woolsey. A new perspective in functional EIV linear models: Part II. Communications in Statistics: Theory and Methods, 50(4):856–873, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1642493.

Attwood:2022:CIE

[AT22] Kristopher Attwood and Lili Tian. Confidence interval estimation of the Youden index and corresponding cut-point for a combination of biomarkers under normality. Communications in Statistics: Theory and Methods, 51(2):501–518, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1751852.

Al-Talib:2020:CIT

[ATAAM20] Mohammad Al-Talib, Mohammad Al Kadiri, and Abedel-Qader Al-Masri. On combining independent tests in case of conditional normal distribution. Communications in Statistics: Theory and Methods, 49(23):5627–5638, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1621343.

Aradhye:2024:EEM

[ATB24] Girish Aradhye, George Tzougas, and Deepesh Bhati. EM estimation for the mixed Pareto regression model for claim severities. Communications in Statistics: Theory and Methods, 53(15):5507–5523, 2024. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2023.2221358.

Ahmed:2024:MCE

[ATJ^{+}24] Rashid Ahmed, M. H. Tahir, Rida Jabeen, H. M. Kashif Rasheed, and Abid Khan. Minimal circular efficient generalized strongly balanced repeated measurements designs. Communications in Statistics: Theory and Methods, 53(12):4395– 4406, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2023.2179885.

Alpman:2020:APA

[AÜ20] Burcu Alpman and Deniz Ünal. Accelerating the premiums for annuities, life annuities and life insurance. Communications in Statistics: Theory and Methods, 49(7):1665–1694, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1564329.

Abdullahi:2022:EMB

[AU22] Umar Kabir Abdullahi and Fidelis Ifeanyi Ugwuowo. On efficient median-based linear regression estimator for population mean. Communications in Statistics: Theory and Methods, 51(15):5012–5024, 2022. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2020.1831540.

Arif:2022:JMM

[AuAH22] Farah Arif, Muhammad Noor ul Amin, and Muhammad Hanif. Joint monitoring of mean and variance under double ranked set sampling using likelihood ratio test statistic. Communications in Statistics: Theory and Methods, 51(17):6032– 6048, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2020.1851721.

Aslam:2023:MTR

[AuAHS23] Irfan Aslam, Muhammad Noor ul Amin, Muhammad Hanif, and Prayas Sharma. Memory type ratio and product estimators under ranked-based sampling schemes. Communications in Statistics: Theory and Methods, 52(4):1155–1177, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1924784.

Augustynowicz:2021:ASA

[Aug21] Aneta Augustynowicz. Almost sure asymptotic properties of central order statistics from stationary processes. Communications in Statistics: Theory and Methods, 50(24):6273–6285, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2020.1741629.

Amin:2022:DTI

[AUQ22] Muhammad Amin, Muhammad Aman Ullah, and Muhammad Qasim. Diagnostic techniques for the inverse Gaussian regression model. Communications in Statistics: Theory and Methods, 51(8):2552–2564, 2022. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2020.1777308.

A:2020:PEL

[AV20] Sharon Varghese A. and V. S. Vaidyanathan. Parameter estimation of Lindley step stress model with independent competing risk under type 1 censoring. Communications in Statistics: Theory and Methods, 49(12):3026–3043, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1584317.

Amestoy:2024:IRE

[AvdWvW24] Matteo Amestoy, Mark A. van de Wiel, and Wessel N. van Wieringen. Identifiability of the random effects’ covariance matrix of the linear mixed model. Communications in Statistics: Theory and Methods, 53(21):7711–7722, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2272003.

Asar:2020:IEB

[AW20] Yasin Asar and Jibo Wu. An improved and efficient biased estimation technique in logistic regression model. Communications in Statistics: Theory and Methods, 49(9):2237–2252, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2019.1568494.

Alghalith:2024:SMO

[AW24] Moawia Alghalith and Wing Keung Wong. A solution to the multidimensionality in option pricing. Communications in Statistics: Theory and Methods, 53(7):2477–2482, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2137680.

Alawieh:2021:PUP

[AWB21] Hiba Alawieh, Nicolas Wicker, and Christophe Biernacki. Projection under pairwise distance control. Communications in Statistics: Theory and Methods, 50(24):5692–5720, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1741626.

Artner:2022:SPC

[AWL^{+}22] Richard Artner, Paul P. Wellingerhof, Ginette Lafit, Tim Loossens, Wolf Vanpaemel, and Francis Tuerlinckx. The shape of partial correlation matrices. Communications in Statistics: Theory and Methods, 51(12):4133–4150, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1811338.

Adil:2021:SSS

[AWM21] Iftikhar Hussain Adil, Abdul Wahid, and Edmund H. Mantell. Split sample skewness. Communications in Statistics: Theory and Methods, 50(22):5171–5188, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1804588.

Azarbar:2021:SBM

[AWN21] Ali Azarbar, Yu Wang, and Saralees Nadarajah. Simultaneous Bayesian modeling of longitudinal and survival data in breast cancer patients. Communications in Statistics: Theory and Methods, 50(2):400–414, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1635701.

Alaeddini:2021:BRA

[AYP21] Reza Alaeddini, Mo Yang, and Borek Puza. Bayesian regression analysis of stutter in DNA mixtures. Communications in Statistics: Theory and Methods, 50(17):4066–4080, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1710760.

Akgul:2021:ESR

[AYS21] Fatma Gul Akgul, Keming Yu, and Birdal Senoglu. Estimation of the system reliability for generalized inverse Lindley distribution based on different sampling designs. Communications in Statistics: Theory and Methods, 50(7):1532–1546, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2019.1705977.

Al-Zaydi:2024:MIW

[AZAZ24] Areej M. Al-Zaydi and Bander Al-Zahrani. Moments of inverse Weibull-geometric distribution based on progressive type-II right censored order statistics. Communications in Statistics: Theory and Methods, 53(20):7352–7369, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2263117.

Asenso:2022:PLM

[AZL22] Theophilus Quachie Asenso, Hai Zhang, and Yong Liang. Pliable lasso for the multinomial logistic regression. Communications in Statistics: Theory and Methods, 51(11):3596–3611, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2020.1800041.

Bhati:2021:UNB

[BA21] Deepesh Bhati and Ishfaq S. Ahmed. On uniform-negative binomial distribution including Gauss hypergeometric function and its application in count regression modeling. Communications in Statistics: Theory and Methods, 50(13):3106–3122, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2019.1682163.

Bahram:2023:SRI

[BA23] A. Bahram and B. Almohaimeed. Some results on increments of l^{∞}-valued random fields. Communications in Statistics: Theory and Methods, 52(9):3124–3131, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1967398.

Bera:2024:NSA

[BA24a] Kuntal Bera and M. Z. Anis. A note on statistical analysis of C_{pk} for autocorrelated data in the presence of random measurement errors. Communications in Statistics: Theory and Methods, 53(20):7326–7331, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2263115.

Bera:2024:PII

[BA24b] Kuntal Bera and M. Z. Anis. Process incapability index for autocorrelated data in the presence of measurement errors. Communications in Statistics: Theory and Methods, 53(15):5439– 5459, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2023.2220921.

Bhattacharya:2024:GMD

[BA24c] Ritwik Bhattacharya and Muhammad Aslam. Generalized multiple dependent state sampling plan for quality inspection in ocean dataset. Communications in Statistics: Theory and Methods, 53(8):2857–2867, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2150052.

Balakrishnan:2024:HRO

[BAA24] Ghobad Saadat Kia (Barmalzan) Narayanaswamy Balakrishnan, Seyed Masih Ayat, and Abbas Akrami. HR and RHR orderings of extremes of dependent variables under Archimedean copula. Communications in Statistics: Theory and Methods, 53(13):4776–4789, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2189506.

Balamurali:2020:MDS

[BAAJ20] Saminathan Balamurali, Muhammad Aslam, Liaquat Ahmad, and Chi-Hyuck Jun. A mixed double sampling plan based on C_{pk}. Communications in Statistics: Theory and Methods, 49(8):1840–1857, 2020. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2019.1565836.

Benallou:2022:ARS

[BABF22] M. Benallou, M. K. Attouch, T. Benchikh, and O. Fetitah. Asymptotic results of semi-functional partial linear regression estimate under functional spatial dependency. Communications in Statistics: Theory and Methods, 51(20):7172–7192, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2020.1871021.

Barakat:2023:CPN

[BAE23] Haroon M. Barakat, Metwally A. Alawady, and M. A. Abd Elgawad. Correction to the paper “A new extension of the FGM copula with an application in reliability” by Rasha Ebaid, Walid Elbadawy, Essam Ahmed and Abdalla Abdelghaly. Communications in Statistics: Theory and Methods, 52(16): ??, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2021.1879864. See [EEAA22].

Bashir:2022:SCC

[BAGR22] Zahid Bashir, Rashid Ahmed, Jigneshkumar Gondaliya, and Kashif Rasheed. Some classes of circular balanced RMDs and their conversion into circular strongly and nearly strongly balanced RMDs. Communications in Statistics: Theory and Methods, 51(19):6573–6584, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1863989.

Bahraoui:2024:BCC

[Bah24] Tarik Bahraoui. Bernstein copula characteristic function. Communications in Statistics: Theory and Methods, 53(18): 6513–6526, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2023.2247107.

Barmalzan:2023:OFS

[BAHB23] Ghobad Barmalzan, Abbas Akrami, Ali Akbar Hosseinzadeh, and Narayanaswamy Balakrishnan. Orderings of fail-safe systems with components having Marshall–Olkin– Harris lifetime distributions. Communications in Statistics: Theory and Methods, 52(19):7022–7040, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2037648.

Bailey:2021:HDV

[Bai21] R. A. Bailey. Hasse diagrams as a visual aid for linear models and analysis of variance. Communications in Statistics: Theory and Methods, 50(21):5034–5067, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1676443.

Butt:2020:APY

[BAJ20] Kalsoom Afzal Butt, Muhammad Aslam, and Chi-Hyuck Jun. Analysis of process yield in a cost-effective double acceptance sampling plan. Communications in Statistics: Theory and Methods, 49(24):5975–5987, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1625922.

Balakrishnan:2022:MMP

[Bal22] Narayanaswamy Balakrishnan. My musings on a pioneering work of Erich Lehmann and its rediscoveries on some families of distributions. Communications in Statistics: Theory and Methods, 51(22):8066–8073, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1896003.

Barczy:2022:NEP

[Bar22] Mátyás Barczy. A new example for a proper scoring rule. Communications in Statistics: Theory and Methods, 51(11):3705– 3712, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2020.1801737.

Barthelemy:2024:GS

[Bar24] Quentin Barthélemy. On generalized f-statistics. Communications in Statistics: Theory and Methods, 53(20):7281–7297, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2023.2263112.

Bansal:2020:PEC

[BAY20a] Gunjan Bansal, Adarsh Anand, and V. S. S. Yadavalli. Predicting effective customer lifetime: an application of survival analysis for telecommunication industry. Communications in Statistics: Theory and Methods, 49(10):2305–2320, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1570264.

Bayramoglu:2020:BMD

[Bay20b] Ismihan Bayramoglu. Bivariate and multivariate distributions with bimodal marginals. Communications in Statistics: Theory and Methods, 49(2):361–384, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1543766.

Bazyari:2024:IRP

[Baz24a] Abouzar Bazyari. Inequalities on the ruin probability for lighttailed distributions with some restrictions. Communications in Statistics: Theory and Methods, 53(23):8312–8328, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2281273.

Bazyari:2024:ERP

[Baz24b] Abouzar Bazyari. On the evaluation of ruin probabilities in a generalized dual binomial risk model using Markov property. Communications in Statistics: Theory and Methods, 53(4): 1162–1187, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2022.2093910.

Boularouk:2024:GGQ

[BB24] Yakoub Boularouk and Jean-Marc Bardet. Generalized Gaussian quasi-maximum likelihood estimation for most common time series. Communications in Statistics: Theory and Methods, 53(4):1459–1478, 2024. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2022.2103148.

Barmalzan:2021:OED

[BBAA21] Ghobad Barmalzan, Narayanaswamy Balakrishnan, Seyed Masih Ayat, and Abbas Akrami. Orderings of extremes dependent modified proportional hazard and modified proportional reversed hazard variables under Archimedean copula. Communications in Statistics: Theory and Methods, 50(22):5358– 5379, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2020.1728331.

Balakrishnan:2024:DIB

[BBCL24] Narayanaswamy Balakrishnan, Francesco Buono, Camilla Calì, and Maria Longobardi. Dispersion indices based on kerridge inaccuracy measure and Kullback–Leibler divergence. Communications in Statistics: Theory and Methods, 53(15): 5574–5592, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2023.2222926.

Balakrishnan:2020:EMP

[BBH20] Narayanaswamy Balakrishnan, Ghobad Barmalzan, and Abedin Haidari. Exponentiated models preserve stochastic orderings of parallel and series systems. Communications in Statistics: Theory and Methods, 49(7):1592–1602, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1532007.

Balakrishnan:2022:WE

[BBL22] Narayanaswamy Balakrishnan, Francesco Buono, and Maria Longobardi. On weighted extropies. Communications in Statistics: Theory and Methods, 51(18):6250–6267, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1860222.

Barry:2022:AIM

[BBM^{+}22] Amadou Barry, Nikhil Bhagwat, Bratislav Misic, Jean-Baptiste Poline, and Celia M. T. Greenwood. Asymmetric influence measure for high dimensional regression. Communications in Statistics: Theory and Methods, 51(16):5461–5487, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2020.1841793.

Benchoulak:2024:APK

[BBM24] H. Benchoulak, M. Boukeloua, and F. Messaci. Asymptotic properties of a kernel etimate of the conditional mode in the presence of doubly censored data. Communications in Statistics: Theory and Methods, 53(4):1143–1161, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2155790.

Bendersky:2022:DRS

[BBP22a] Michael Bendersky, Noam Barak, and Yisrael Parmet. Dual response surface optimization using a process “capability” index. Communications in Statistics: Theory and Methods, 51(20):7006–7020, 2022. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2020.1869990.

Bjornham:2022:ACC

[BBP22b] Oscar Björnham, Niklas Brännström, and Leif Persson. Absolutely continuous copulas with prescribed support constructed by differential equations, with an application in toxicology. Communications in Statistics: Theory and Methods, 51(19):

6601–6625, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2020.1864825.

Begum:2021:AUS

[BC21] Anjana Begum and Gautam Choudhury. Analysis of an unreliable single server batch arrival queue with two types of services under Bernoulli vacation policy. Communications in Statistics: Theory and Methods, 50(9):2136–2160, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1659972.

Benhaddou:2023:NEB

[BC23a] Rida Benhaddou and Matthew A. Connell. Nonparametric empirical Bayes estimation based on generalized Laguerre series. Communications in Statistics: Theory and Methods, 52(19):6896–6915, 2023. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2022.2036346.

Brown:2023:MIS

[BC23b] Mark Brown and Joel E. Cohen. Markov’s inequality: Sharpness, renewal theory, finite samples, reliability theory. Communications in Statistics: Theory and Methods, 52(11):3652– 3660, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2021.1977960.

Begum:2024:AUP

[BC24] Anjana Begum and Gautam Choudhury. Analysis of an unreliable N-policy queue with set-up period under Bernoulli schedule and re-service approach. Communications in Statistics: Theory and Methods, 53(1):1–33, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2072515.

Bakouch:2022:PPD

[BCA22] Hassan S. Bakouch, Christophe Chesneau, and Rawda I. Abdullah. A pliant parametric detection model for line transect data sampling. Communications in Statistics: Theory and Methods, 51(21):7340–7353, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1872640.

Biswas:2024:NMG

[BCG24] Aniket Biswas, Subrata Chakraborty, and Indranil Ghosh. A novel method of generating distributions on the unit interval with applications. Communications in Statistics: Theory and Methods, 53(22):8210–8233, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2280506.

Buzaianu:2024:SAT

[BCH24] Elena M. Buzaianu, Pinyuen Chen, and Lifang Hsu. Selecting among treatments with two Bernoulli endpoints. Communications in Statistics: Theory and Methods, 53(6):1964–1984, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2022.2117557.

Bridges:2022:LTW

[BCKS22] William C. Bridges, Neil J. Calkin, Catherine M. Kenyon, and Matthew J. Saltzman. Log transformations: What not to expect when you’re expecting. Communications in Statistics: Theory and Methods, 51(5):1514–1521, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1771368.

Barmalzan:2021:CSP

[BD21] Ghobad Barmalzan and Somayeh Shahraki Dehsukhteh. Comparisons of series and parallel systems with heterogeneous exponentiated geometric components. Communications in Statistics: Theory and Methods, 50(18):4352–4366, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1716251.

Bouzebda:2022:SRA

[BD22] Salim Bouzebda and Sultana Didi. Some results about kernel estimators for function derivatives based on stationary and ergodic continuous time processes with applications. Communications in Statistics: Theory and Methods, 51(12):3886–3933, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2020.1805466.

Baratnia:2024:COP

[BD24a] Mohammad Baratnia and Mahdi Doostparast. Classifying one-parameter Fréchet populations on the basis of a non-linear fixed effects model. Communications in Statistics: Theory and Methods, 53(9):3246–3258, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2151309.

Boos:2024:PCU

[BD24b] Dennis Boos and Kaiyuan Duan. Pairwise comparisons using ranks in block designs. Communications in Statistics: Theory and Methods, 53(9):3259–3275, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2151310.

Barakat:2021:TSN

[BEAA21] H. M. Barakat, Magdy E. El-Adll, and Amany E. Aly. Twosample nonparametric prediction intervals based on random number of generalized order statistics. Communications in Statistics: Theory and Methods, 50(19):4571–4586, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1719421.

Bouzebda:2024:CLT

[BEhF24] Salim Bouzebda, Thouria El-hadjali, and Anouar Abdeldjaoued Ferfache. Central limit theorems for functional Zestimators with functional nuisance parameters. Communications in Statistics: Theory and Methods, 53(7):2535–2577, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2022.2138439.

Benhaddou:2022:AFD

[Ben22a] Rida Benhaddou. Anisotropic functional deconvolution for the irregular design: a minimax study. Communications in Statistics: Theory and Methods, 51(13):4589–4601, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1818783.

Benhaddou:2022:ENR

[Ben22b] Rida Benhaddou. Estimation in nonparametric regression model with additive and multiplicative noise via Laguerre series. Communications in Statistics: Theory and Methods, 51(20):7193–7207, 2022. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2020.1871490.

Beran:2023:SIP

[Ber23a] Jan Beran. On semiparametric inference for periodically modulated density functions. Communications in Statistics: Theory and Methods, 52(23):8478–8500, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2064501.

Bouhadjera:2023:REP

[BER23b] Feriel Bouhadjera, Ould Saïd Elias, and Remita Mohamed Riad. Relative error prediction: Strong uniform consistency for censoring time series model. Communications in Statistics: Theory and Methods, 52(11):3709–3729, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1979584.

Beyhum:2023:HDI

[Bey23] Jad Beyhum. High-dimensional inference robust to outliers with ℓ_{1}-norm penalization. Communications in Statistics: Theory and Methods, 52(16):5866–5876, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.2021239.

Belzunce:2022:NSC

[BFPM22] Félix Belzunce, Alba M. Franco-Pereira, and Julio Mulero. New stochastic comparisons based on tail value at risk measures. Communications in Statistics: Theory and Methods, 51(3):767–788, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03610926.2020.1754857.

Bazeli:2022:CCB

[BFSN22] Shakiba Bazeli, Mohammad Saber Fallahnezhad, Ahmad Sadegheih, and Hasan Hosseini Nasab. Clustering conditionbased maintenance for manufacturing systems with both perfect and imperfect maintenance actions. Communications in Statistics: Theory and Methods, 51(17):6109–6126, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1854303.

Baringhaus:2021:MRN

[BG21] Ludwig Baringhaus and Rudolf Grübel. Mixture representations of noncentral distributions. Communications in Statistics: Theory and Methods, 50(24):5997–6013, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1738487.

Bansal:2022:CEC

[BG22a] Namisha Bansal and Davinder Kumar Garg. Construction and existence of constant block sum PBIB designs. Communications in Statistics: Theory and Methods, 51(7):2231–2241, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2020.1772308.

Bansal:2022:WEP

[BG22b] Shilpa Bansal and Nitin Gupta. Weighted extropies and past extropy of order statistics and k-record values. Communications in Statistics: Theory and Methods, 51(17):6091–6108, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2020.1853773.

Benseradj:2022:SUC

[BG22c] Hassiba Benseradj and Zohra Guessoum. Strong uniform consistency rate of an M-estimator of regression function for incomplete data under α-mixing condition. Communications in Statistics: Theory and Methods, 51(7):2082–2115, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1764037.

Bhattacharyya:2022:NMA

[BGM22] Dhrubasish Bhattacharyya, Shyamal Ghosh, and Murari Mitra. On a non-monotonic ageing class based on the failure rate average. Communications in Statistics: Theory and Methods, 51(14):4807–4826, 2022. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2020.1824273.

Bey:2023:KRE

[BGT23] Siham Bey, Zohra Guessoum, and Abdelkader Tatachak. Kernel regression estimation for LTRC and associated data. Communications in Statistics: Theory and Methods, 52(18):6381– 6406, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2022.2028839.

Beauducel:2024:EIM

[BH24] André Beauducel and Norbert Hilger. On the effect of items measuring different factors across individuals on item inter-correlations. Communications in Statistics: Theory and Methods, 53(18):6362–6379, 2024. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2023.2244100.

Barmalzan:2022:DVO

[BHB22] Ghobad Barmalzan, Ali Akbar Hosseinzadeh, and Narayanaswamy Balakrishnan. Dispersion and variability orders of mixture exponential distributions and their sample spacings, and some associated characterizations. Communications in Statistics: Theory and Methods, 51(24):8657–8670, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1901925.

Bagdonavicius:2021:MAD

[BHL21] Vilijandas Bagdonavicius, Mohamed Ali Hafdi, and Ruta Levuliene. Modeling and analysis of data with confounding covariates and crossing of the hazard functions. Communications in Statistics: Theory and Methods, 50(22):5262–5284, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2020.1728330.

Bialek:2020:CEP

[Bia20] Jacek Białek. Comparison of elementary price indices. Communications in Statistics: Theory and Methods, 49(19):4787– 4803, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2019.1609035.

Bialek:2022:EPI

[Bia22] Jacek Białek. Elementary price indices under the GBM price model. Communications in Statistics: Theory and Methods, 51(5):1232–1251, 2022. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2021.1938127.

Bickel:2020:CIS

[Bic20] David R. Bickel. Confidence intervals, significance values, maximum likelihood estimates, etc. sharpened into Occam’s razors. Communications in Statistics: Theory and Methods, 49(11):2703–2712, 2020. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2019.1580739.

Bickel:2022:CDE

[Bic22a] David R. Bickel. Confidence distributions and empirical Bayes posterior distributions unified as distributions of evidential support. Communications in Statistics: Theory and Methods, 51(10):3142–3163, 2022. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2020.1790004.

Bickel:2022:MAS

[Bic22b] David R. Bickel. Model averages sharpened into Occam’s razors: Deep learning enhanced by Rényi entropy. Communications in Statistics: Theory and Methods, 51(23):8283–8295, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2021.1891438.

Bickel:2022:MPD

[Bic22c] David R. Bickel. Moderating probability distributions for unrepresented uncertainty: Application to sentiment analysis via deep learning. Communications in Statistics: Theory and Methods, 51(19):6559–6572, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1863988.

Bickel:2023:IEP

[Bic23] David R. Bickel. Interval estimation, point estimation, and null hypothesis significance testing calibrated by an estimated posterior probability of the null hypothesis. Communications in Statistics: Theory and Methods, 52(3):763–787, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1921805. See errata [Bic24].

Bickel:2024:EIE

[Bic24] David R. Bickel. Errata: Interval estimation, point estimation, and null hypothesis significance testing calibrated by an estimated posterior probability of the null hypothesis Bickel (2023). Communications in Statistics: Theory and Methods, 53(14):5297, 2024. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2023.2203788. See [Bic23].

Ji:2023:REP

[bJwWyX23] Ai bing Ji, Bo wen Wei, and Lan ying Xu. Robust estimation of panel data regression models and applications. Communications in Statistics: Theory and Methods, 52(21):7647–7659, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2022.2050403.

Bhushan:2022:OCE

[BK22] Shashi Bhushan and Anoop Kumar. On optimal classes of estimators under ranked set sampling. Communications in Statistics: Theory and Methods, 51(8):2610–2639, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1777431.

Bhootna:2024:TSA

[BK24] Niharika Bhootna and Arun Kumar. Tempered stable autoregressive models. Communications in Statistics: Theory and Methods, 53(2):765–785, 2024. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2022.2092145.

Barmalzan:2022:USR

[BKB22] Ghobad Barmalzan, Sajad Kosari, and Narayanaswamy Balakrishnan. Usual stochastic and reversed hazard orders of parallel systems with independent heterogeneous components. Communications in Statistics: Theory and Methods, 51(14): 4781–4806, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2020.1823415.

Barmalzan:2023:SCP

[BKB23] Ghobad Barmalzan, Sajad Kosari, and Narayanaswamy Balakrishnan. Stochastic comparisons of parallel systems with starting devices. Communications in Statistics: Theory and Methods, 52(1):170–182, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1910841.

Barmalzan:2023:LRD

[BKHB23] Ghobad Barmalzan, Sajad Kosari, Ali Akbar Hosseinzadeh, and Narayanaswamy Balakrishnan. Likelihood ratio and dispersive orders of parallel and series systems consisting of dependent multiple-outlier components. Communications in Statistics: Theory and Methods, 52(19):6695–6715, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2032751.

Boroomandi:2022:NPP

[BKK22] Fahimeh Boroomandi and Mahmood Kharrati-Kopaei. Non parametric PROS quality control chart for monitoring the process mean. Communications in Statistics: Theory and Methods, 51(6):1706–1723, 2022. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2020.1767139.

Bedbur:2020:REG

[BKM20] Stefan Bedbur, Udo Kamps, and Miriam Marner. Rényi entropy of m-generalized order statistics. Communications in Statistics: Theory and Methods, 49(14):3397–3406, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1588321.

Bhushan:2023:SEC

[BKS23a] Shashi Bhushan, Anoop Kumar, and Saurabh Singh. Some efficient classes of estimators under stratified sampling. Communications in Statistics: Theory and Methods, 52(6):1767–1796, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2021.1939052.

Bouzebda:2023:ANR

[BKS23b] Salim Bouzebda, Salah Khardani, and Yousri Slaoui. Asymptotic normality of the regression mode in the nonparametric random design model for censored data. Communications in Statistics: Theory and Methods, 52(19):7069–7093, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2039200.

Bhushan:2024:PEN

[BKS24] Shashi Bhushan, Anoop Kumar, and Shivam Shukla. Performance evaluation of novel logarithmic estimators under correlated measurement errors. Communications in Statistics: Theory and Methods, 53(15):5353–5363, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2219793.

Benhaddou:2020:MAW

[BL20] Rida Benhaddou and Qing Liu. Minimax adaptive wavelet estimator for the anisotropic functional deconvolution model with unknown kernel. Communications in Statistics: Theory and Methods, 49(21):5312–5331, 2020. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2019.1617880.

Bi:2023:BMR

[BL23] Junna Bi and Danping Li. Behavioral mean-risk portfolio selection in continuous time via quantile. Communications in Statistics: Theory and Methods, 52(14):4904–4933, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1990955.

Belalia:2024:TSC

[BL24a] Mohamed Belalia and Guanjie Lyu. Two-stage conditional density estimation based on Bernstein polynomials. Communications in Statistics: Theory and Methods, 53(11):4172– 4193, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2023.2176715.

Bercu:2024:HEM

[BL24b] Bernard Bercu and Lucile Laulin. How to estimate the memory of the elephant random walk. Communications in Statistics: Theory and Methods, 53(7):2578–2598, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2139149.

Borkowski:2022:FVD

[BLC22] John J. Borkowski, Wanida Limmun, and Boonorm Chomtee. Factorwise variance dispersion graphs. Communications in Statistics: Theory and Methods, 51(23):8427–8445, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1897144.

Bu:2023:MFA

[BLZ23] Qihui Bu, Liwei Liu, and Yiqiang Q. Zhao. Mean field approximations to a queueing system with thresholdbased workload control scheme. Communications in Statistics: Theory and Methods, 52(11):3960–3981, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1983601.

Bagui:2020:ERT

[BM20a] Subhash C. Bagui and K. L. Mehra. On an elementary ratio technique for proving convergence of known distributions. Communications in Statistics: Theory and Methods, 49(22): 5535–5552, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2019.1619771.

Bibi:2020:YWT

[BM20b] Abdelouahab Bibi and Fateh Merahi. Yule–Walker type estimator of first-order time-varying periodic bilinear differential model for stochastic processes. Communications in Statistics: Theory and Methods, 49(16):4046–4072, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1594300.

Boukabour:2021:SBN

[BM21] Seloua Boukabour and Afif Masmoudi. Semiparametric Bayesian networks for continuous data. Communications in Statistics: Theory and Methods, 50(24):5974–5996, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1738486.

Becila:2022:NLS

[BM22] S. Becila and M. Merzougui. Nonlinear least squares estimation of the periodic EXPAR(1) model. Communications in Statistics: Theory and Methods, 51(15):5369–5381, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1839099.

Bensaber:2023:FAP

[BM23] Fatna Bensaber and Tahar Mourid. Functional autoregressive process with seasonality. Communications in Statistics: Theory and Methods, 52(20):7131–7145, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2093907.

Barmalzan:2022:SCS

[BmAB22] Ghobad Barmalzan, Seyed masih Ayat, and Narayanaswamy Balakrishnan. Stochastic comparisons of series and parallel systems with dependent Burr type XII components. Communications in Statistics: Theory and Methods, 51(7):2209–2230, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2020.1772307.

Bartels:2022:CIM

[BMG22] Christian Bartels, Johanna Mielke, and Ekkehard Glimm. Confidence intervals with maximal average power. Communications in Statistics: Theory and Methods, 51(14):4940–4956, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2020.1828465.

Bisht:2024:HHF

[BMH24] Jaya Bisht, Rohan Mishra, and A. Hamdi. Hermite– Hadamard and Fejér-type inequalities for generalized ηconvex stochastic processes. Communications in Statistics: Theory and Methods, 53(15):5299–5310, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2218506.

Bekiryazici:2021:MRD

[BMK21] Zafer Bekiryazici, Mehmet Merdan, and Tulay Kesemen. Modification of the random differential transformation method and its applications to compartmental models. Communications in Statistics: Theory and Methods, 50(18):4271– 4292, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2020.1713372.

Bignozzi:2020:LDM

[BMP20] Valeria Bignozzi, Claudio Macci, and Lea Petrella. Large deviations for method-of-quantiles estimators of one-dimensional parameters. Communications in Statistics: Theory and Methods, 49(5):1132–1157, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03610926.2018.1554134.

Bosse:2023:GDR

[BMR^{+}23] Michael Bossé, Eric Marland, Gregory Rhoads, Jose Almer Sanqui, and Zack BeMent. Generalizing R^{2} for Deming regressions. Communications in Statistics: Theory and Methods, 52(21):7731–7743, 2023. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2022.2059678.

Baig:2020:ICD

[BMT20] Afifa Baig, Saadia Masood, and Tanveer Ahmed Tarray. Improved class of difference-type estimators for population median in survey sampling. Communications in Statistics: Theory and Methods, 49(23):5778–5793, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1622017.

Babilua:2021:EBR

[BN21] P. K. Babilua and E. A. Nadaraya. On estimating the Bernoulli regression function using Bernstein polynomials. Communications in Statistics: Theory and Methods, 50(17): 3928–3941, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2019.1709872.

Bivigou:2022:RMS

[BN22] Ulrich Djemby Bivigou and Guy Martial Nkiet. Robust multiple-set linear canonical analysis based on minimum covariance determinant estimator. Communications in Statistics: Theory and Methods, 51(22):7783–7800, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1880593.

Bouzebda:2024:APC

[BN24] Salim Bouzebda and Amel Nezzal. Asymptotic properties of conditional U-statistics using delta sequences. Communications in Statistics: Theory and Methods, 53(13):4602–4657, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2023.2179887.

Banihashemi:2023:IQR

[BNAK23] Atefe Banihashemi, Mohammad Saber Fallah Nezhad, Amirhossein Amiri, and Michael B. C. Khoo. Integrated QSS-RS plans based on the process yield index for lot acceptance determination. Communications in Statistics: Theory and Methods, 52(23):8584–8606, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2065303.

Balogoun:2022:KBM

[BNO22] Armando S. K. Balogoun, Guy M. Nkiet, and Carlos Ogouyandjou. Kernel based method for the k-sample problem with functional data. Communications in Statistics: Theory and Methods, 51(17):5826–5849, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1849719.

Boukeloua:2021:SSC

[Bou21a] Mohamed Boukeloua. Study of semiparametric copula models via divergences with bivariate censored data. Communications in Statistics: Theory and Methods, 50(23):5429–5452, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2020.1734834.

Boukhari:2021:WLL

[Bou21b] Fakhreddine Boukhari. Weak laws of large numbers for maximal weighted sums of random variables. Communications in Statistics: Theory and Methods, 50(1):105–115, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1630437.

Boukhari:2022:FJS

[Bou22] Fakhreddine Boukhari. On a Feller-jajte strong law of large numbers. Communications in Statistics: Theory and Methods, 51(18):6218–6226, 2022. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2020.1858102.

Bhushan:2020:CEC

[BP20] Shashi Bhushan and Shailja Pandey. A cost-effective computational approach with non response on two occasions. Communications in Statistics: Theory and Methods, 49(20):4951– 4973, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2019.1609518.

Bakar:2022:IMW

[BP22] Shaiful Anuar Abu Bakar and Dharini Pathmanathan. Income modeling with the Weibull mixtures. Communications in Statistics: Theory and Methods, 51(11):3612–3628, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1800737.

Barnwal:2024:CRA

[BP24] Vikas Barnwal and M. S. Panwar. Competing risks analysis for dependent causes using Marshall–Olkin bivariate generalized lifetime family. Communications in Statistics: Theory and Methods, 53(4):1212–1240, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2094412.

Bagnato:2022:LMP

[BPZ22] Luca Bagnato, Antonio Punzo, and Maria Grazia Zoia. Leptokurtic moment-parameterized elliptically contoured distributions with application to financial stock returns. Communications in Statistics: Theory and Methods, 51(2):486–500, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2020.1751202.

Bao:2024:POR

[BPZ24] Qian Bao, Jiangyan Peng, and Lei Zou. Pareto-optimal reinsurance for both the insurer and the reinsurer under the riskadjusted value and general premium principles. Communications in Statistics: Theory and Methods, 53(10):3616–3641, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2022.2158344.

Bashir:2022:SRF

[BQW22] Shakila Bashir, Ahmad Mahmood Qureshi, and Noor Waseem. System reliability of the functions using Pareto– Rayleigh distribution. Communications in Statistics: Theory and Methods, 51(23):8130–8148, 2022. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2021.1888124.

Basatini:2020:MSS

[BR20] Ferdous Mohammadi Basatini and Saeid Rezakhah. Markov switch smooth transition HYGARCH model: Stability and estimation. Communications in Statistics: Theory and Methods, 49(10):2384–2409, 2020. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2019.1576884.

Bedoui:2022:CBN

[BR22] Adel Bedoui and Ori Rosen. Comparison of Bayesian nonparametric density estimation methods. Communications in Statistics: Theory and Methods, 51(19):6667–6682, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1864828.

Baratnia:2023:RLP

[BR23] M. Baratnia and A. H. Rezaei Roknabadi. Residual lifetime prediction for sequential k-out-of-n systems with dependent component lifetimes. Communications in Statistics: Theory and Methods, 52(9):3178–3192, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1969408.

Brown:2022:ADN

[Bro22] Austin R. Brown. The alternative distribution of the non parametric extended median test CUSUM chart for multiple stream processes. Communications in Statistics: Theory and Methods, 51(16):5750–5761, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1850792.

Berred:2020:ARL

[BS20a] Alexandre Berred and Alexei Stepanov. Asymptotic results for lower exponential spacings. Communications in Statistics: Theory and Methods, 49(7):1730–1741, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1565781.

Bouzebda:2020:BSN

[BS20b] Salim Bouzebda and Yousri Slaoui. Bandwidth selector for nonparametric recursive density estimation for spatial data defined by stochastic approximation method. Communications in Statistics: Theory and Methods, 49(12):2942–2963, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2019.1584313.

Bank:2021:ACS

[BS21a] B. Bank and S. K. Samanta. Analytical and computational studies of the BMAP/G^{(a,Y)} /1 queue. Communications in Statistics: Theory and Methods, 50(15):3586–3614, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1708941.

Brown:2021:NCC

[BS21b] Austin R. Brown and Jay R. Schaffer. A nonparametric CUSUM control chart for multiple stream processes based on a modified extended median test. Communications in Statistics: Theory and Methods, 50(24):6067–6080, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1738492.

Berred:2022:APL

[BS22a] Alexandre Berred and Alexei Stepanov. Asymptotic properties of lower exponential spacings under Type-II progressive censoring. Communications in Statistics: Theory and Methods, 51(14):4841–4853, 2022. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2020.1824276.

Bhattacharya:2022:LST

[BS22b] Tuhinshubhra Bhattacharya and Arindam Sengupta. Largesample tests for comparing Likert-type scale data. Communications in Statistics: Theory and Methods, 51(5):1179–1196, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2021.1872634.

Bura:2024:MLB

[BS24] Gulab Singh Bura and Himanshi Sharma. Maximum likelihood and Bayesian estimation on M/M/1 queueing model with balking. Communications in Statistics: Theory and Methods, 53(14):5117–5145, 2024. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2023.2208695.

Bai:2022:SIT

[BSN22] Xuchao Bai, Yimin Shi, and Hon Keung Tony Ng. Statistical inference of Type-I progressively censored step-stress accelerated life test with dependent competing risks. Communications in Statistics: Theory and Methods, 51(10):3077–3103, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2020.1788081.

Bouhadjera:2022:SCN

[BSR22] Feriel Bouhadjera, Elias Ould Saïd, and Mohamed Riad Remita. Strong consistency of the nonparametric local linear regression estimation under censorship model. Communications in Statistics: Theory and Methods, 51(20):7056–7072, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2020.1870142.

Batra:2022:POB

[BT22] Luckshay Batra and H. C. Taneja. Portfolio optimization based on generalized information theoretic measures. Communications in Statistics: Theory and Methods, 51(18):6367– 6384, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2020.1861294.

Bounkhel:2023:EMB

[BTH23] Messaoud Bounkhel, Lotfi Tadj, and Ramdane Hedjar. Entropy maximization for the busy period of a single server queue. Communications in Statistics: Theory and Methods, 52(23):8555–8565, 2023. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2022.2065301.

Bulut:2020:MDB

[Bul20] Hasan Bulut. Mahalanobis distance based on minimum regularized covariance determinant estimators for high dimensional data. Communications in Statistics: Theory and Methods, 49(24):5897–5907, 2020. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2020.1719420.

Bulut:2023:RHT

[Bul23] Hasan Bulut. A robust Hotelling test statistic for one sample case in high dimensional data. Communications in Statistics: Theory and Methods, 52(13):4590–4604, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1996606.

Bai:2024:ETE

[BY24] Fangfang Bai and Ruiyu Yang. Estimation of treatment effects in two sample problems under general biased sampling data. Communications in Statistics: Theory and Methods, 53(22): 7829–7841, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2023.2273208.

Blazhievska:2020:EIR

[BZ20] Irina Blazhievska and Vladimir Zaiats. Estimation of impulse response functions in two-output systems. Communications in Statistics: Theory and Methods, 49(2):257–280, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1536210.

Blazhievska:2021:CCI

[BZ21] Irina Blazhievska and Vladimir Zaiats. On cross-correlogram IRF’s estimators of two-output systems in spaces of continuous functions. Communications in Statistics: Theory and Methods, 50(24):6024–6048, 2021. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2020.1738490.

Bai:2023:RFM

[BZ23] Xiaodong Bai and Li Zheng. Robust factor models for highdimensional time series and their forecasting. Communications in Statistics: Theory and Methods, 52(19):6806–6819, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2022.2033777.

Bai:2024:MDP

[BZLW24] Yansong Bai, Yong Zhang, Congmin Liu, and Zhiming Wang. Moderate deviation principle for different types of classical likelihood ratio tests. Communications in Statistics: Theory and Methods, 53(7):2599–2616, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2144376.

Cankaya:2020:RPM

[ÇA20] Mehmet Niyazi Çankaya and Olcay Arslan. On the robustness properties for maximum likelihood estimators of parameters in exponential power and generalized T distributions. Communications in Statistics: Theory and Methods, 49(3):607–630, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2018.1549243.

Castro-Alva:2023:TSC

[CAAARCTS23] José Juan Castro-Alva, Félix Almendra-Arao, Hortensia Josefina Reyes-Cervantes, and Francisco Solano Tajonar-Sanabria. Test size calculation by comparing three binomial proportions. Communications in Statistics: Theory and Methods, 52(13): 4578–4589, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2021.1996605.

Campanelli:2024:EDS

[Cam24] Leonardo Campanelli. On the Euclidean distance statistic of Benford’s law. Communications in Statistics: Theory and Methods, 53(2):451–474, 2024. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2022.2082480.

Cavicchioli:2020:SRA

[Cav20] Maddalena Cavicchioli. Spectral representation and autocovariance structure of Markov switching DSGE models. Communications in Statistics: Theory and Methods, 49(7):1635– 1652, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2018.1563184.

Cavicchioli:2024:GAM

[Cav24] Maddalena Cavicchioli. Generalized autocovariance matrices for multivariate time series. Communications in Statistics: Theory and Methods, 53(10):3797–3817, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2164465.

Cha:2022:NCM

[CB22] Ji Hwan Cha and F. G. Badía. A new class of marginally regular multivariate counting processes generated by the mixture of multivariate Poisson processes. Communications in Statistics: Theory and Methods, 51(13):4235–4251, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1812652.

Chokri:2024:ANW

[CB24a] Khalid Chokri and Salim Bouzebda. Asymptotic normality for the wavelet partially linear additive model components estimation. Communications in Statistics: Theory and Methods, 53(23):8376–8411, 2024. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2023.2286905.

Chokri:2024:UBC

[CB24b] Khalid Chokri and Salim Bouzebda. Uniform-in-bandwidth consistency results in the partially linear additive model components estimation. Communications in Statistics: Theory and Methods, 53(9):3383–3424, 2024. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2022.2153605.

Cavalcanti:2021:IST

[CBB^{+}21] Alexsandro B. Cavalcanti, Denise A. Botter, Lúcia P. Barroso, Manoel Santos-Neto, and Gauss M. Cordeiro. Improved score tests for exponential family nonlinear models. Communications in Statistics: Theory and Methods, 50(15):3731–3745, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2019.1710202.

Chaturvedi:2020:SOA

[CBJ20] Ajit Chaturvedi, Sudeep R. Bapat, and Neeraj Joshi. Secondorder approximations for a multivariate analog of Behrens– Fisher problem through three-stage procedure. Communications in Statistics: Theory and Methods, 49(14):3466–3480, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2019.1589517.

Cancho:2021:SMI

[CBLS21] Vicente G. Cancho, Gladys Barriga, Jeremias Leão, and Helton Saulo. Survival model induced by discrete frailty for modeling of lifetime data with long-term survivors and change-point. Communications in Statistics: Theory and Methods, 50(5):1161–1172, 2021. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2019.1648826.

Choudhury:2021:ERF

[CBM21] Mriganka Mouli Choudhury, Rahul Bhattacharya, and Sudhansu S. Maiti. On estimating reliability function for the family of power series distribution. Communications in Statistics: Theory and Methods, 50(12):2801–2830, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1676446.

Chen:2022:OSP

[CC22] Yen-Luan Chen and Chin-Chih Chang. Optimum sequential preventive maintenance first or last policies with imperfect maintenance for a system subject to shocks. Communications in Statistics: Theory and Methods, 51(1):162–178, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1746807.

Christofides:2023:ANS

[CC23] Tasos C. Christofides and Charalambos Charalambous. Asymptotic normality of U-statistics based on i.i.d. or negatively associated observations by utilizing Zolotarev’s ideal metric. Communications in Statistics: Theory and Methods, 52(12):4083–4102, 2023. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2021.1986533.

Cabral:2022:CSC

[CCG22] Ivanilda Cabral, Frederico Caeiro, and M. Ivette Gomes. On the comparison of several classical estimators of the extreme value index. Communications in Statistics: Theory and Methods, 51(1):179–196, 2022. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2020.1746970.

Casarin:2022:FIS

[CCMtH22] Roberto Casarin, Jorge E. Camargo, German Molina, and Enrique ter Horst. A framework for information synthesis into sentiment indicators using text mining methods. Communications in Statistics: Theory and Methods, 51(15):5265–5283, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2020.1837878.

Chen:2021:IPS

[CCR21] Xiaoyan Chen, Zengjing Chen, and Liying Ren. An invariance principle of strong law of large numbers under nonadditive probabilities. Communications in Statistics: Theory and Methods, 50(10):2398–2418, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1669805.

Chen:2020:CMC

[CCW^{+}20] Meiqian Chen, Kan Chen, Zijian Wang, Zhengliang Lu, and Xuejun Wang. Complete moment convergence for partial sums of arrays of rowwise negatively superadditive dependent random variables. Communications in Statistics: Theory and Methods, 49(5):1158–1173, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1554136.

Chen:2024:ERP

[CD24] Luyuan Chen and Yong Deng. Entropy of random permutation set. Communications in Statistics: Theory and Methods, 53(11):4127–4146, 2024. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2023.2173975.

Cuerden:2024:DWM

[CDCC24] Meaghan S. Cuerden, Liqun Diao, Cecilia A. Cotton, and Richard J. Cook. Doubly weighted mean score estimating functions with a partially observed effect modifier. Communications in Statistics: Theory and Methods, 53(11):3899–3919, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2023.2166790.

Chaubey:2024:SNP

[CDL24a] Yogendra P. Chaubey, Isha Dewan, and Jun Li. On some non parametric estimators of the quantile density function for a stationary associated process. Communications in Statistics: Theory and Methods, 53(15):5553–5573, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2222922.

Chen:2024:BEB

[CDL24b] Yong Chen, Zhen Ding, and Ying Li. Berry-esséen bounds and almost sure CLT for the quadratic variation of a class of Gaussian process. Communications in Statistics: Theory and Methods, 53(11):3920–3939, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2167055.

Chernova:2024:RCD

[CDMR24] Oksana Chernova, Olena Dehtiar, Yuliya Mishura, and Kostiantyn Ralchenko. Rate of convergence of discretized drift parameters estimators in the Cox–Ingersoll–Ross model. Communications in Statistics: Theory and Methods, 53(13):4857– 4879, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2023.2196591.

Chatrabgoun:2020:FCL

[CDP20] Omid Chatrabgoun, Alireza Daneshkhah, and Gholamali Parham. On the functional central limit theorem for first passage time of nonlinear semi-Markov reward processes. Communications in Statistics: Theory and Methods, 49(19):4737– 4750, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2019.1606917.

Cuadras:2020:TGB

[CDSG20] Carles M. Cuadras, Walter Diaz, and Sonia Salvo-Garrido. Two generalized bivariate FGM distributions and rank reduction. Communications in Statistics: Theory and Methods, 49(23):5639–5665, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03610926.2019.1620780.

Celik:2022:WAH

[Cel22] N. Celik. Welch’s ANOVA: Heteroskedastic skew-t error terms. Communications in Statistics: Theory and Methods, 51(9):3065–3076, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03610926.2020.1788084.

Cha:2022:SGS

[CF22a] Ji Hwan Cha and Maxim Finkelstein. On some general survival models with delayed failures. Communications in Statistics: Theory and Methods, 51(22):7911–7928, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1884261.

Chen:2022:MSM

[CF22b] Yijie Chen and Hailin Feng. Maintenance strategy of multicomponent system based on structure updating and group importance measure. Communications in Statistics: Theory and Methods, 51(9):2919–2935, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1784434.

Chen:2020:CAT

[CFS20] Ling Chen, Yanqin Feng, and Jianguo Sun. A class of additive transformation models for recurrent gap times. Communications in Statistics: Theory and Methods, 49(16):4030–4045, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2019.1594299.

Chen:2020:ITH

[CG20] Yongshuai Chen and Wenwen Guo. Independence test in high-dimension using distance correlation and power enhancement technique. Communications in Statistics: Theory and Methods, 49(17):4216–4233, 2020. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2019.1595656.

Cho:2021:CAP

[CG21] Youngjoo Cho and Debashis Ghosh. Covariate adjustment via propensity scores for recurrent events in the presence of dependent censoring. Communications in Statistics: Theory and Methods, 50(1):216–236, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1634208.

Chen:2023:ORC

[CG23] Siqi Chen and Wenhao Gui. Order restricted classical and Bayesian inference of a multiple step-stress model from twoparameter Rayleigh distribution under Type I censoring. Communications in Statistics: Theory and Methods, 52(14):

5082–5112, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2021.2012685.

Chacko:2024:EPR

[CG24a] Manoj Chacko and Varghese George. Extropy properties of ranked set sample when ranking is not perfect. Communications in Statistics: Theory and Methods, 53(9):3187–3210, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2022.2150825.

Chaudhary:2024:GWC

[CG24b] Santosh Kumar Chaudhary and Nitin Gupta. General weighted cumulative residual (past) extropy of minimum (maximum) ranked set sampling with unequal samples. Communications in Statistics: Theory and Methods, 53(22):8076– 8096, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2023.2279910.

Cramer:2021:MSP

[CGL21] Erhard Cramer, Julian Górny, and Benjamin Laumen. Multisample progressive Type-I censoring of exponentially distributed lifetimes. Communications in Statistics: Theory and Methods, 50(22):5285–5313, 2021. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2020.1728328.

Corona:2024:FSP

[CGLP24] Francisco Corona, Víctor M. Guerrero, and Jesús López-Pérez. The finite sample performance of two methods for choosing a power transformation when seasonally adjusting a time series with X-13ARIMA-SEATS. Communications in Statistics: Theory and Methods, 53(3):965–979, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2098334.

Cao:2023:ISL

[CGW23] Xiangyang Cao, Karl Gregory, and Dewei Wang. Inference for sparse linear regression based on the leave-one-covariate-out solution path. Communications in Statistics: Theory and Methods, 52(18):6640–6657, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2032171.

Cai:2020:PIS

[CH20a] Ting Cai and Hong Chang Hu. Probability inequalities for sums of NSD random variables and applications. Communications in Statistics: Theory and Methods, 49(2):281–306, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2018.1536211.

Chen:2020:RAD

[CH20b] Mi Chen and Xiang Hu. Risk aggregation with dependence and overdispersion based on the compound Poisson INAR(1) process. Communications in Statistics: Theory and Methods, 49(16):3985–4001, 2020. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2019.1594297.

Chen:2021:ORP

[CH21a] Yanhong Chen and Yijun Hu. Optimal reinsurance from the perspectives of both insurers and reinsurers under the VaR risk measure and Vajda condition. Communications in Statistics: Theory and Methods, 50(15):3677–3694, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1710197.

Choi:2021:EIL

[CH21b] Sangbum Choi and Xuelin Huang. Efficient inferences for linear transformation models with doubly censored data. Communications in Statistics: Theory and Methods, 50(9):2188– 2200, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2019.1662046.

Chen:2022:OID

[CH22] Yanhong Chen and Yijun Hu. Optimal insurance design under Vajda condition and exclusion clauses. Communications in Statistics: Theory and Methods, 51(18):6268–6295, 2022.

CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1860223.

Chang:2021:OPR

[Cha21a] Chin-Chih Chang. Optimal preventive replacement policy for operating products with renewing free-replacement warranty. Communications in Statistics: Theory and Methods, 50(18): 4255–4270, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2020.1713371.

Charalambides:2021:MNM

[Cha21b] Charalambos A. Charalambides. q-Multinomial and negative q-multinomial distributions. Communications in Statistics: Theory and Methods, 50(24):5873–5898, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1737711.

Charalambides:2022:MPI

[Cha22] Charalambos A. Charalambides. Multivariate q-Pólya and inverse q-Pólya distributions. Communications in Statistics: Theory and Methods, 51(14):4854–4876, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1825740.

Chacko:2023:BTJ

[Cha23a] V. M. Chacko. On Birnbaum type joint importance measures for multistate reliability systems. Communications in Statistics: Theory and Methods, 52(9):2799–2818, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1961000.

Chang:2023:OPR

[Cha23b] Chin-Chih Chang. Optimal preventive replacement last policy for a successive random works system with random lead time. Communications in Statistics: Theory and Methods, 52(4): 1202–1216, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2021.1926506.

Charalambides:2023:CCC

[Cha23c] Charalambos A. Charalambides. Correction to: Charalambides, C. A. (2021). q-Multinomial and negative q-multinomial distributions. Communications in Statistics — Theory and Methods doi:10.1080/03610926.2020.1737711. Communications in Statistics: Theory and Methods, 52(17):??, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1938129. See [Cha23d].

Charalambides:2023:FMB

[Cha23d] Charalambos A. Charalambides. q-factorial moments of bivariate discrete q-distributions. Communications in Statistics: Theory and Methods, 52(16):5677–5702, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.2016834. See correction [Cha23c].

Chasiotis:2023:OCT

[Cha23e] Vasilis Chasiotis. On optimality and construction of twotreatment circular cross-over designs. Communications in Statistics: Theory and Methods, 52(6):1692–1701, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1937654.

Chaurasia:2023:MMB

[Cha23f] Ashok Chaurasia. Model-modified BIC as a competitor of BIC variants for model selection in regression and order selection in time series. Communications in Statistics: Theory and Methods, 52(23):8425–8453, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2064497.

Chatla:2024:NKR

[Cha24] Suneel Babu Chatla. A new kernel regression approach for robustified L_{2} boosting. Communications in Statistics: Theory and Methods, 53(22):8186–8209, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2280497.

Chechile:2020:BAM

[Che20a] Richard A. Chechile. A Bayesian analysis for the Mann– Whitney statistic. Communications in Statistics: Theory and Methods, 49(3):670–696, 2020. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2018.1549247.

Chen:2020:SRP

[Che20b] Mingjing Chen. A self-reliant projected information criterion for the number of factors. Communications in Statistics: Theory and Methods, 49(10):2466–2484, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1576889.

Chen:2021:GEV

[Che21] Su Chen. Generalized expected value of continuous random variables and its applications. Communications in Statistics: Theory and Methods, 50(20):4844–4866, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1725823.

Cheng:2022:ELR

[Che22] Conghua Cheng. Empirical likelihood ratio for two-sample compound Poisson processes under infinite second moment. Communications in Statistics: Theory and Methods, 51(11): 3787–3798, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2020.1801741.

Chen:2023:URM

[Che23a] Dan Chen. Uncertain regression model with moving average time series errors. Communications in Statistics: Theory and Methods, 52(21):7632–7646, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2050402.

Chen:2023:IVP

[Che23b] Pei-Le Chen. An improved variable parameter mean square error control chart. Communications in Statistics: Theory and Methods, 52(16):5752–5766, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.2019769.

Chen:2023:AON

[Che23c] Ray-Ming Chen. Associate an optimal normal distribution with a finite numerical discrete data set via extended spline functions. Communications in Statistics: Theory and Methods, 52(10):3478–3491, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1974480.

Chen:2023:OPR

[Che23d] Yen-Luan Chen. Optimal preventive replacement policies for a system with discrete scheduled times. Communications in Statistics: Theory and Methods, 52(5):1356–1368, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1926513.

Cheng:2023:IIT

[Che23e] Hao Cheng. IPLSL and IPLSQ: Two types of imputation PLS algorithms for hierarchical latent variable model. Communications in Statistics: Theory and Methods, 52(8):2493–2513, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2021.1955382.

Chen:2024:ORS

[Che24a] Yen-Luan Chen. Optimal replacement scheduling for a multicomponent system with failure interaction. Communications in Statistics: Theory and Methods, 53(6):2092–2102, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2117988.

Cheng:2024:QBP

[Che24b] Hao Cheng. Quantile-based PLS-SEM with bag of little bootstraps. Communications in Statistics: Theory and Methods, 53(18):6561–6579, 2024. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2023.2248324.

Chasiotis:2021:OTT

[CK21] Vasilis Chasiotis and Stratis Kounias. Optimal two treatment circular repeated measurements designs up to four periods. Communications in Statistics: Theory and Methods, 50(20): 4867–4878, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2020.1725824.

Chew:2022:EES

[CKKY22] Yiying Chew, Michael B. C. Khoo, Khai Wah Khaw, and Wai Chung Yeong. Economic and economic-statistical designs of variable sample size and sampling interval coefficient of variation chart. Communications in Statistics: Theory and Methods, 51(6):1811–1835, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1769670.

Chatterjee:2023:JMP

[CKL23] Kashinath Chatterjee, Christos Koukouvinos, and Angeliki Lappa. A joint monitoring of the process mean and variance with a TEWMA-Max control chart. Communications in Statistics: Theory and Methods, 52(22):8069–8095, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2056748.

Chowdhury:2023:CTP

[CKM23a] Shovan Chowdhury, Amarjit Kundu, and Surja Kanta Mishra. On comparison of two parallel systems having log-Lindley distributed components. Communications in Statistics: Theory and Methods, 52(1):121–140, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1910838.

Chowdhury:2023:OPS

[CKM23b] Shovan Chowdhury, Amarjit Kundu, and Surja Kanta Mishra. Ordering properties of the smallest order statistic from Weibull-g random variables. Communications in Statistics: Theory and Methods, 52(18):6525–6541, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2031220.

Chen:2021:BEB

[CL21a] Yong Chen and Ying Li. Berry–Esséen bound for the parameter estimation of fractional Ornstein–Uhlenbeck processes with the Hurst parameter H ∈ (0, 1/2). Communications in Statistics: Theory and Methods, 50(13):2996–3013, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1678641.

Cheng:2021:ELR

[CL21b] Conghua Cheng and Zhi Liu. Empirical likelihood ratio under infinite covariance matrix of the random vectors. Communications in Statistics: Theory and Methods, 50(18):4300–4307, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2020.1713377.

Cho:2021:FRE

[CL21c] Seonghun Cho and Johan Lim. A frequentist’s resolution of the exchange paradox. Communications in Statistics: Theory and Methods, 50(20):4879–4889, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1725826.

Cao:2023:GLR

[CL23] Mingxiang Cao and Shiting Liang. A generalized likelihood ratio test for linear hypothesis of k-sample means in high dimension. Communications in Statistics: Theory and Methods, 52(24):8725–8737, 2023. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2022.2069820.

Chang:2022:OCP

[CLC22] Hao Chang, Xueyan Li, and Xingjiang Chen. Optimal consumption and portfolios with the hyperbolic absolute risk aversion preference under the CEV model. Communications in Statistics: Theory and Methods, 51(24):8799–8821, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1907411.

Chen:2021:RAC

[CLF21] Yurong Chen, Ji Luo, and Jie Feng. Regression analysis of case II interval-censored data with auxiliary covariates. Communications in Statistics: Theory and Methods, 50(17):4022–4038, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2019.1710755.

Chen:2023:TBF

[CLLD23] Chao Chen, Yilin Li, Keqing Liang, and Jinlin Du. A test for the Behrens–Fisher problem based on the method of variance estimates recovery. Communications in Statistics: Theory and Methods, 52(18):6444–6455, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2028842.

Castellares:2020:TPB

[CLMA20] Fredy Castellares, Artur J. Lemonte, and Germán Moreno-Arenas. On the two-parameter Bell–Touchard discrete distribution. Communications in Statistics: Theory and Methods, 49(19):4834–4852, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03610926.2019.1609515.

See letter [Pui24].

Chen:2023:PEV

[CLP23] Yong Chen, Ying Li, and Xingzhi Pei. Parameter estimation for Vasicek model driven by a general Gaussian noise. Communications in Statistics: Theory and Methods, 52(9):3132–3148, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2021.1967399.

Chang:2024:AUL

[CLPS24] Yun Fah Chang, Sing Yan Looi, Wei Yeing Pan, and Shin Zhu Sim. Autocorrelated unreplicated linear functional relationship model for multivariate time series data. Communications in Statistics: Theory and Methods, 53(20):7244–7261, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2263110.

Chen:2024:CCR

[CLS24] Pingyan Chen, Jingjing Luo, and Soo Hak Sung. Complete convergence for randomly weighted sums of dependent random variables and an application. Communications in Statistics: Theory and Methods, 53(20):7197–7215, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2262636.

Cheng:2022:CMC

[CLW^{+}22] Nan Cheng, Xiaoqin Li, Minghui Wang, Xuejun Wang, and Mengmei Xi. Complete moment convergence for m-END random variables with application to non-parametric regression models. Communications in Statistics: Theory and Methods, 51(11):3573–3595, 2022. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2020.1800040.

Chen:2022:OAT

[CLWZ22] Jianbin Chen, Xiaoying Lai, Mengmeng Wang, and Peng Zhao. Optimal allocation of two redundancies in a ncomponent series system with proportional hazard rates lifetimes. Communications in Statistics: Theory and Methods, 51(9):2748–2764, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03610926.2020.1779297.

Chen:2021:BAP

[CLZ21] Huaping Chen, Qi Li, and Fukang Zhu. Binomial AR(1) processes with innovational outliers. Communications in Statistics: Theory and Methods, 50(2):446–472, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1635704.

Can:2024:OCM

[CLZ24a] Peng Can, Zhi-Ming Li, and Li Zhi. Optimal confounding measures for two-level regular designs. Communications in Statistics: Theory and Methods, 53(16):5954–5971, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2238859.

Chen:2024:ODM

[CLZ24b] Jiali Chen, Ling Ling, and Chongqi Zhang. D- and A-optimal designs for multi-response mixture experiments with qualitative factors. Communications in Statistics: Theory and Methods, 53(15):5593–5611, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2223705.

Chen:2024:RCI

[CLZL24] Xuejiao Chen, Peng Li, Ming Zhou, and Bing Liu. Robust consumption for individuals with pessimistic survival beliefs. Communications in Statistics: Theory and Methods, 53(14): 4991–5002, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2023.2198659.

Chen:2021:RPP

[CLZZ21] Yu Chen, Yujie Liao, Qi Zhang, and Weiping Zhang. Ruin probabilities for the phase-type dual model perturbed by diffusion. Communications in Statistics: Theory and Methods, 50(23):5634–5651, 2021. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2020.1737126.

Christofides:2020:ICT

[CM20] Tasos C. Christofides and Eleni Manoli. Item count technique with no floor and ceiling effects. Communications in Statistics: Theory and Methods, 49(6):1330–1356, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1563165.

Chen:2024:DRM

[CM24] Yanhong Chen and Liangliang Miao. Dynamic risk measures via backward doubly stochastic Volterra integral equations with jumps. Communications in Statistics: Theory and Methods, 53(14):5092–5116, 2024. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2023.2206503.

Coolen-Maturi:2023:MPW

[CMC23] Tahani Coolen-Maturi and Frank P. A. Coolen. A monotonicity property of weighted log-rank tests. Communications in Statistics: Theory and Methods, 52(7):2402–2416, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1952270.

Coolen-Maturi:2020:NPI

[CMCA20] Tahani Coolen-Maturi, Frank P. A. Coolen, and Manal Alabdulhadi. Nonparametric predictive inference for diagnostic test thresholds. Communications in Statistics: Theory and Methods, 49(3):697–725, 2020. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2018.1549249.

Chukwudum:2020:OTD

[CMM20] Queensley C. Chukwudum, Peter Mwita, and Joseph K. Mung’atu. Optimal threshold determination based on the mean excess plot. Communications in Statistics: Theory and Methods, 49(24):5948–5963, 2020. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2019.1624772.

Chong:2024:STS

[CMM24] Zhi Lin Chong, Amitava Mukherjee, and Marco Marozzi. Some two-sample tests for simultaneously comparing both parameters of the shifted exponential models. Communications in Statistics: Theory and Methods, 53(2):524–556, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2085875.

Cuberos:2020:CCT

[CMMD20] A. Cuberos, E. Masiello, and V. Maume-Deschamps. Copulas checker-type approximations: Application to quantiles estimation of sums of dependent random variables. Communications in Statistics: Theory and Methods, 49(12):3044–3062, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2019.1586936.

Cantarinha:2023:RMF

[CMO^{+}23] Ana Cantarinha, Elsa Moreira, Manuela Oliveira, Susete Marques, and João T. Mexia. A risk model for forest fires based on asymptotic results for multivariate collective models. Single models and structured families of models. Communications in Statistics: Theory and Methods, 52(19):6857–6877, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2034867.

Castilla:2021:TLH

[CMP21] E. Castilla, N. Martín, and L. Pardo. Testing linear hypotheses in logistic regression analysis with complex sample survey data based on phi-divergence measures. Communications in Statistics: Theory and Methods, 50(22):5228–5247, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1746342.

Costa:2022:MPM

[CN22] Antonio Fernando Branco Costa and Antonio Faria Neto. Monitoring the process mean with an ATTRIVAR chart. Communications in Statistics: Theory and Methods, 51(14): 4903–4920, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2020.1828463.

Chai:2023:MRR

[CN23a] Jichong Chai and Tapan K. Nayak. Minimax randomized response methods for protecting respondent’s privacy. Communications in Statistics: Theory and Methods, 52(10):3429– 3451, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2021.1973503.

Chen:2023:BSA

[CN23b] Xinyu Chen and Balgobin Nandram. A Bayesian small area model with order restrictions for contingency tables. Communications in Statistics: Theory and Methods, 52(13):4636– 4658, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2021.1998534.

Connett:2024:WTT

[Con24] John E. Connett. When are there too many collisions? Variants of the birthday problem. Communications in Statistics: Theory and Methods, 53(12):4487–4497, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2184186.

Chen:2021:ODR

[CP21] Fenge Chen and Xingchun Peng. Optimal deterministic reinsurance and investment for an insurer under meanvariance criterion. Communications in Statistics: Theory and Methods, 50(13):3123–3136, 2021. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2019.1682165.

Catana:2022:CEO

[CP22] Luigi-Ionut Catana and Vasile Preda. Comparing the extremes order statistics between two random variables sequences using transmuted distributions. Communications in Statistics: Theory and Methods, 51(24):8499–8516, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1898641.

Chakraborty:2023:GWS

[CP23] Siddhartha Chakraborty and Biswabrata Pradhan. Generalized weighted survival and failure entropies and their dynamic versions. Communications in Statistics: Theory and Methods, 52(3):730–750, 2023. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2021.1921803.

Chakraborty:2024:SPW

[CP24a] Siddhartha Chakraborty and Biswabrata Pradhan. Some properties of weighted survival extropy and its extended measures. Communications in Statistics: Theory and Methods, 53 (1):66–89, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2022.2076118.

Chaudhuri:2024:HUR

[CP24b] Arijit Chaudhuri and Dipika Patra. How to use randomized response survey data at hand by a specific procedure to judge its efficiency versus a possible rival. Communications in Statistics: Theory and Methods, 53(19):6762–6776, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2250489.

Cao:2021:ALE

[CPS21] Mingxiang Cao, Junyong Park, and Guangjun Shen. Φ admissibility of linear estimators of common mean parameter in general multivariate linear models under a balanced loss function. Communications in Statistics: Theory and Methods, 50(17): 4050–4065, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2019.1710757.

Chen:2021:LBC

[CQT21] E. Chen, Zong-Feng Qi, and Yu Tang. Lower bound and construction of mixed-level locating arrays. Communications in Statistics: Theory and Methods, 50(19):4606–4616, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1722841.

Contreras-Reyes:2021:MSN

[CRKC21] Javier E. Contreras-Reyes, Fereshte Kahrari, and Daniel Devia Cortés. On the modified skew-normal-Cauchy distribution: properties, inference and applications. Communications in Statistics: Theory and Methods, 50(15):3615–3631, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1708942.

Chaudhary:2024:CBA

[CRVK24] Manoj K. Chaudhary, Basant K. Ray, Gautam K. Vishwakarma, and Cem Kadilar. A calibration-based approach on estimation of mean of a stratified population in the presence of non response. Communications in Statistics: Theory and Methods, 53(19):7054–7068, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2257818.

Chaudhuri:2020:FPQ

[CS20] Arijit Chaudhuri and Purnima Shaw. A finite population quantile estimation by unequal probability sampling. Communications in Statistics: Theory and Methods, 49(22):5419– 5426, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2019.1618475.

Chumnaul:2022:ITS

[CS22] J. Chumnaul and M. Sepehrifar. Improved test for the scale parameter of the power-law process with incomplete failure data. Communications in Statistics: Theory and Methods, 51(12):4221–4234, 2022. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2020.1813303.

Cizikoviene:2023:CUG

[CS23] Ugne Cizikoviene and Viktor Skorniakov. On a couple of unresolved group testing conjectures. Communications in Statistics: Theory and Methods, 52(8):2448–2460, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1953531.

Choi:2024:SST

[CS24] Ji-Eun Choi and Dong Wan Shin. Subsample scan test for multiple breaks based on self-normalization. Communications in Statistics: Theory and Methods, 53(2):627–640, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2087883.

Chakraborty:2023:MEI

[CSA23] Hrishikesh Chakraborty, Nicole Solomon, and Kevin J. Anstrom. A method to estimate intra-cluster correlation for clustered categorical data. Communications in Statistics: Theory and Methods, 52(2):429–444, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1914660.

Carneiro:2024:SAI

[CSBM24] Hérica P. A. Carneiro, Mônica C. Sandoval, Denise A. Botter, and Tiago M. Magalhães. Some asymptotic inferential aspects of the Kumaraswamy distribution. Communications in Statistics: Theory and Methods, 53(17):6160–6176, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2241091.

Cai:2020:RBA

[CSZ20] Jing Cai, Yimin Shi, and Yongjin Zhang. Robust Bayesian analysis for parallel system with masked data under inverse Weibull lifetime distribution. Communications in Statistics: Theory and Methods, 49(6):1422–1434, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1563173.

Caraballo:2020:MET

[CTEB20] Tomás Caraballo, Abd El-Moneim Anwar Teamah, and Abd Al-Aziz Hosni El-Bagoury. Minimizing the expected time to detect a randomly located lost target using 3dimensional search technique. Communications in Statistics: Theory and Methods, 49(13):3313–3328, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1588323.

Chandrasekhar:2021:CBE

[CVDY21] P. Chandrasekhar, V. S. Vaidyanathan, T. M. Durairajan, and V. S. S. Yadavalli. Classical and Bayes estimation in the M—D—1 queueing system. Communications in Statistics: Theory and Methods, 50(22):5411–5421, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1734833.

Chen:2020:ARP

[CW20a] Cen Chen and Shijie Wang. Asymptotic ruin probability for a by-claim risk model with pTQAI claims and constant interest force. Communications in Statistics: Theory and Methods, 49(18):4367–4377, 2020. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2019.1599022.

Cui:2020:TPC

[CW20b] Yunwei Cui and Rongning Wu. Test of parameter changes in a class of observation-driven models for count time series. Communications in Statistics: Theory and Methods, 49(8):1933– 1959, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2019.1565843.

Chen:2022:CCT

[CW22] Binxia Chen and Qunying Wu. Complete convergence theorem for negatively dependent random variables under sub-linear expectations. Communications in Statistics: Theory and Methods, 51(10):3202–3215, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1790603.

Chu:2023:ETV

[CW23] Fang-I Chu and Yuedong Wang. Estimating time-varying treatment switching effect using accelerated failure time model with application to vascular access for hemodialysis. Communications in Statistics: Theory and Methods, 52(15):5145– 5154, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2021.2004423.

Chen:2024:CCT

[CW24a] Xiaocong Chen and Qunying Wu. Complete convergence theorems for moving average process generated by independent random variables under sub-linear expectations. Communications in Statistics: Theory and Methods, 53(15):5378–5404, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2023.2220449.

Chen:2024:FRC

[CW24b] Xiaocong Chen and Qunying Wu. Further research on complete integral convergence for moving average process of ND random variables under sub-linear expectations. Communications in Statistics: Theory and Methods, 53(19):7102–7118, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2023.2258428.

Chen:2023:ABS

[CWCY23] Zhangting Chen, Bingjie Wang, Dongya Cheng, and Jigao Yan. Asymptotic behavior for sum ruin probability of a generalized bidimensional risk model with heavy-tailed claims. Communications in Statistics: Theory and Methods, 52(22): 8002–8017, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2022.2055072.

Cao:2020:ILR

[CWJC20] Chunzheng Cao, Yahui Wang, Shaobo Jin, and Yunjie Chen. Improved likelihood ratio tests in a measurement error model for multivariate replicated data. Communications in Statistics: Theory and Methods, 49(5):1025–1042, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1554125.

Chen:2023:FRM

[CWK23] Hongmei Chen, Jibo Wu, and B. M. Golam Kibria. Further research on the modified ridge principal component estimator in linear model. Communications in Statistics: Theory and Methods, 52(24):8894–8901, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2085876.

Chen:2023:CRF

[CWM23] Boming Chen, Fangfang Wang, and Chunsheng Ma. Kcombined random fields: Basic properties and stochastic orderings. Communications in Statistics: Theory and Methods, 52(2):409–428, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03610926.2021.1914100.

Cao:2022:DEF

[CX22a] Ping Cao and Zhiming Xia. Distributed estimation and its fast algorithm for change-point in location models*. Communications in Statistics: Theory and Methods, 51(23):8328–8348, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2021.1894447.

Cheng:2022:FTR

[CX22b] Fengyang Cheng and Hui Xu. The finite-time ruin probability of the nonhomogeneous Poisson risk model with conditionally independent subexponential claims. Communications in Statistics: Theory and Methods, 51(12):4119–4132, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1811337.

Cai:2024:BAM

[CX24] Jingheng Cai and Xiaoli Xu. Bayesian analysis of mixture models with yeo-Johnson transformation. Communications in Statistics: Theory and Methods, 53(18):6600–6613, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2248326.

Cao:2022:SIS

[CXT^{+}22] Yingshu Cao, Ying Xu, Ming T. Tan, Pingyan Chen, and Chongyang Duan. A simple and improved score confidence interval for a single proportion. Communications in Statistics: Theory and Methods, 51(8):2659–2675, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1779747.

Cheng:2020:ARP

[CY20] Dongya Cheng and Changjun Yu. Asymptotic ruin probabilities of a two-dimensional renewal risk model with dependent inter-arrival times. Communications in Statistics: Theory and Methods, 49(7):1742–1760, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1565782.

Chen:2023:APG

[CY23] Xianbin Chen and Juliang Yin. Asymptotic properties of GEE estimator for clustered ordinal data with high-dimensional covariates. Communications in Statistics: Theory and Methods, 52(4):1300–1317, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1934029.

Chen:2024:AGG

[CY24] Xianbin Chen and Juliang Yin. Asymptotics of the general GEE estimator for high-dimensional longitudinal data. Communications in Statistics: Theory and Methods, 53(14):5041– 5056, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2023.2205045.

Cui:2023:SSS

[CYB23] Lirong Cui, He Yi, and Narayanaswamy Balakrishnan. State space splitting of a finite Markov process and some discussions on related counting processes. Communications in Statistics: Theory and Methods, 52(11):4021–4052, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1986531.

Chen:2021:MBF

[CYW21] Fang Chen, Keying Ye, and Min Wang. The minimum Bayes factor hypothesis test for correlations and partial correlations. Communications in Statistics: Theory and Methods, 50(11): 2467–2480, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2019.1667397.

Cao:2020:PBB

[CZ20a] Xuan Cao and Shaojun Zhang. A permutation-based Bayesian approach for inverse covariance estimation. Communications in Statistics: Theory and Methods, 49(14):3557–3571, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1590601.

Chen:2020:MDM

[CZ20b] Xu Chen and WenYan Zhuo. Martingale and duality methods for optimal investment and reinsurance problem in a Lévy model. Communications in Statistics: Theory and Methods, 49(23):5738–5764, 2020. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2019.1620953.

Chen:2021:MDT

[CZ21] Haotian Chen and Yong Zhang. Moderate deviations for the total population arising from a nearly unstable sub-critical Galton–Watson process with immigration. Communications in Statistics: Theory and Methods, 50(2):432–445, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1635703.

Chen:2023:STL

[CZ23a] Hui Chen and Xuemin Zi. Simultaneous test for linear model via projection. Communications in Statistics: Theory and Methods, 52(7):2365–2378, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1952268.

Cheng:2023:GGM

[CZ23b] Yi Cheng and Runchu Zhang. A generalized general minimum lower order confounding criterion for general orthogonal designs. Communications in Statistics: Theory and Methods, 52(14):4799–4814, 2023. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2013.765474.

Chen:2023:HTP

[CZL23] Li Chen, Jie Zhou, and Lizhen Lin. Hypothesis testing for populations of networks. Communications in Statistics: Theory and Methods, 52(11):3661–3684, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1977961.

Chen:2024:TSC

[CZL24] Huaping Chen, Fukang Zhu, and Xiufang Liu. Two-step conditional least squares estimation for the bivariate Z-valued INAR(1) model with bivariate Skellam innovations. Communications in Statistics: Theory and Methods, 53(11):4085– 4106, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2023.2172587.

Chen:2022:ODR

[CZLY22] Mi Chen, Ming Zhou, Haiyan Liu, and Kam Chuen Yuen. Optimal dividends and reinsurance with capital injection under thinning dependence. Communications in Statistics: Theory and Methods, 51(16):5728–5749, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1845737.

Chen:2024:BAT

[CZX24] Jinye Chen, Linyi Zheng, and Yemao Xia. Bayesian analysis for two-part latent variable model with application to fractional data. Communications in Statistics: Theory and Methods, 53(21):7760–7788, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2273205.

Chen:2023:FGL

[CZZ^{+}23] Haiqing Chen, Xu Zhao, Leilei Zhu, Weihu Cheng, and Lu Xu. Fitting generalized logistic distribution by least squares based on the logistic transformation of order statistics. Communications in Statistics: Theory and Methods, 52(2):1–11, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1912353.

Dagne:2022:TCG

[Dag22] Getachew A. Dagne. Two-component generalized bentcable models. Communications in Statistics: Theory and Methods, 51(13):4464–4475, 2022. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2020.1815781.

Dagne:2023:BBB

[Dag23] Getachew A. Dagne. Bayesian bivariate bent-cable model for longitudinal data. Communications in Statistics: Theory and Methods, 52(21):7709–7717, 2023. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2022.2053544.

Daniyal:2020:CRM

[DAS^{+}20] Muhammad Daniyal, Rashid Ahmed, Farrukh Shehzad, M. H. Tahir, and Zafar Iqbal. Construction of repeated measurements designs strongly balanced for residual effects. Communications in Statistics: Theory and Methods, 49(17):4288– 4297, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2019.1599019.

Das:2023:MLE

[Das23] Shubhabrata Das. Maximum likelihood estimation of two-sample population proportions under constraint on their difference. Communications in Statistics: Theory and Methods, 52(9):2836–2851, 2023. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2021.1961152.

Doostmoradi:2023:RSU

[DAZ23] Ali Doostmoradi, Mohammad Reza Akhoond, and Mohammad Reza Zadkarami. Reliability of a system under a new mixed shock model. Communications in Statistics: Theory and Methods, 52(1):156–169, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1910840.

Dib:2021:NBD

[DBBK21] K. Dib, T. Bouezmarni, M. Belalia, and A. Kitouni. Nonparametric bivariate distribution estimation using Bernstein polynomials under right censoring. Communications in Statistics: Theory and Methods, 50(23):5574–5584, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1734832.

delCastillo:2022:ESD

[dC22a] Jose Maria del Castillo. The extended slash distribution of the sum of two independent logistic random variables. Communications in Statistics: Theory and Methods, 51(23):8110–8129, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2021.1888123.

Ding:2022:CWE

[DC22b] Liwang Ding and Ping Chen. On consistency of the weighted estimator in nonparametric regression model with asymptotically almost negatively associated random variables. Communications in Statistics: Theory and Methods, 51(20):7120– 7135, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2020.1871018.

Dar:2024:NHA

[DC24] Irum Sajjad Dar and Sohail Chand. New heteroscedasticityadjusted ridge estimators in linear regression model. Communications in Statistics: Theory and Methods, 53(19):7087– 7101, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2023.2258427.

daCunha:2021:SIV

[dCBV21] Enai Taveira da Cunha, Marcelo Bourguignon, and Klaus L. P. Vasconcellos. On shifted integer-valued autoregressive model for count time series showing equidispersion, underdispersion or overdispersion. Communications in Statistics: Theory and Methods, 50(20):4822–4843, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1725822.

Daraz:2024:SDL

[DCT24] Umer Daraz, E. Chen, and Yu Tang. Supersaturated designs with less β-aberration. Communications in Statistics: Theory and Methods, 53(9):3235–3245, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2150828.

Ding:2020:SIM

[DCY20] Liwang Ding, Ping Chen, and Li Yongming. On some inequalities for ψ-mixing sequences and its applications in conditional value-at-risk estimate. Communications in Statistics: Theory and Methods, 49(22):5455–5467, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1618872.

Ducinskas:2023:AER

[DD23] Kestutis Ducinskas and Lina Dreiziene. Actual error rates in linear discrimination of spatial Gaussian data in terms of semivariograms. Communications in Statistics: Theory and Methods, 52(9):3165–3173, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1968903.

DAmico:2023:PAD

[DDG23] Guglielmo D’Amico, Riccardo De Blasis, and Fulvio Gismondi. Perturbation analysis for dynamic poverty indexes. Communications in Statistics: Theory and Methods, 52(19): 6820–6839, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2022.2034018.

Da:2020:CPS

[DDL20] Gaofeng Da, Weiyong Ding, and Xiaohu Li. On comparisons of population and subpopulations in frailty models. Communications in Statistics: Theory and Methods, 49(15):3698–3711, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2019.1708401.

Djerfi:2024:SED

[DDM24] Kouider Djerfi, Ghaouti Djellouli, and Fethi Madani. Stein estimators for the drift of the mixing of two fractional Brownian motions. Communications in Statistics: Theory and Methods, 53(6):1891–1905, 2024. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2022.2122838.

Danilevicz:2022:BID

[DE22] Ian M. Danilevicz and Ricardo S. Ehlers. Bayesian influence diagnostics using normalized functional Bregman divergence. Communications in Statistics: Theory and Methods, 51(6): 1637–1652, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2020.1764583.

DelCastillo:2021:DSA

[Del21] Jose Maria Del Castillo. A distribution on the simplex arising from inverted chi-square random variables with odd degrees of freedom. Communications in Statistics: Theory and Methods, 50(4):890–909, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1643481.

Dermoune:2021:PWO

[DESSM21] Azzouz Dermoune, Khalifa Es-Sebaiy, Mohammed Es. Sebaiy, and Jabrane Moustaaid. Parametrizations, weights, and optimal prediction. Communications in Statistics: Theory and Methods, 50(4):815–836, 2021. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2019.1642489.

Dey:2024:BFN

[Dey24] Monitirtha Dey. Behavior of FWER in normal distributions. Communications in Statistics: Theory and Methods, 53(9): 3211–3225, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2022.2150826.

DeSantis:2021:NPO

[DG21a] Fulvio De Santis and Stefania Gubbiotti. A note on the progressive overlap of two alternative Bayesian intervals. Communications in Statistics: Theory and Methods, 50(14):3301– 3318, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2019.1692034.

Dimitriou:2021:IAA

[DG21b] V. A. Dimitriou and A. C. Georgiou. Introduction, analysis and asymptotic behavior of a multi-level manpower planning model in a continuous time setting under potential department contraction. Communications in Statistics: Theory and Methods, 50(5):1173–1199, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1648827.

Demircioglu:2024:TMF

[DG24] Sevgi Demircioglu and Bilgehan Güven. Testing for main fixed effects: The symmetry assumption and monotone incomplete data. Communications in Statistics: Theory and Methods, 53(8):2848–2856, 2024. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2022.2150051.

Diaz-Garcia:2023:SMV

[DGCL23] José A. Díaz-García and Francisco J. Caro-Lopera. Singular matrix variate Birnbaum–Saunders distribution under elliptical models. Communications in Statistics: Theory and Methods, 52(6):1653–1667, 2023. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2021.1937650.

Dong:2022:SDT

[DH22] Fengkai Dong and Huilin Huang. Strong deviation theorems for delayed sums of the nonnegative continuous random variables. Communications in Statistics: Theory and Methods, 51(5):1522–1530, 2022. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2020.1774059.

deHierro:2021:ACN

[dHGS21] A. F. Roldán López de Hierro, C. García, and R. Salmerón. Analysis of the condition number in the raise regression. Communications in Statistics: Theory and Methods, 50(24):6195– 6210, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2020.1740737.

Deng:2023:RRP

[DHH^{+}23] Yingchun Deng, Kang Hu, Ya Huang, Hui Ou, and Jieming Zhou. Ruin-related problems in the dual risk model under two different randomized observations. Communications in Statistics: Theory and Methods, 52(17):6241–6265, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2027450.

Dai:2023:IGI

[DHJS23] Wei Dai, Taizhong Hu, Baisuo Jin, and Xiaoping Shi. Incorporating grouping information into Bayesian Gaussian graphical model selection. Communications in Statistics: Theory and Methods, 52(22):7966–7983, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2053864.

Dijoux:2023:NCD

[Dij23] Yann Dijoux. A new class of distributions on the whole real line based on the continuous iteration approach. Communications in Statistics: Theory and Methods, 52(14):4934–4959, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2021.1991376.

Ding:2024:CTL

[Din24a] Xue Ding. The Chover-type law of iterated logarithm for the weighted sums of negatively superadditive dependent random variables. Communications in Statistics: Theory and Methods, 53(4):1277–1293, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2097696.

Ding:2024:CCM

[Din24b] Xue Ding. Complete convergence for moving average process generated by extended negatively dependent random variables under sub-linear expectations. Communications in Statistics: Theory and Methods, 53(22):8166–8185, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2279924.

Ding:2020:CRM

[DJW^{+}20] Shu Ding, Baisuo Jin, Yuehua Wu, Jing Li, and Baiqi Miao. Comparing ratios of the mean to a power of variance in two samples via self-normalized test statistics. Communications in Statistics: Theory and Methods, 49(11):2787–2799, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1584305.

Doll:2023:SSA

[DK23] Monika Doll and Ingo Klein. Sample size analysis for two-sample linear rank tests. Communications in Statistics: Theory and Methods, 52(24):8658–8676, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2068029.

Das:2024:SCS

[DK24a] Sangita Das and Suchandan Kayal. Stochastic comparison of the second-order statistics arising from exponentiated location-scale model. Communications in Statistics: Theory and Methods, 53(7):2430–2458, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2134974.

Diouf:2024:EVT

[DK24b] Saliou Diouf and Yentchabaré Kolani. Extreme value theory for space-time with random coefficients. Communications in Statistics: Theory and Methods, 53(5):1736–1744, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2109676.

Djerroud:2020:SMK

[DKA20] Lamia Djerroud, Tristan Senga Kiessé, and Smail Adjabi. Semiparametric multiple kernel estimators and model diagnostics for count regression functions. Communications in Statistics: Theory and Methods, 49(9):2131–2157, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1568488.

Das:2022:SRM

[DKD22] Sudipta Das, Damitri Kundu, and Anup Dewanji. Software reliability modeling based on NHPP for error occurrence in each fault with periodic debugging schedule. Communications in Statistics: Theory and Methods, 51(14):4890–4902, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1828462.

Du:2022:FAM

[DL22] Yanbin Du and Yuan Lv. A flexible additive-multiplicative transformation mean model for recurrent event data. Communications in Statistics: Theory and Methods, 51(2):328–339, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2020.1748654.

Du:2022:EEB

[DLC22] Yongmei Du, Zhouping Li, and Xiaosong Chen. Efficient empirical Bayes estimates for risk parameters of Pareto distributions. Communications in Statistics: Theory and Methods, 51(6):1674–1692, 2022. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2020.1766501.

Deng:2021:ROS

[DLHZ21] Yingchun Deng, Man Li, Ya Huang, and Jieming Zhou. Robust optimal strategies for an insurer under generalized meanvariance premium principle with defaultable bond. Communications in Statistics: Theory and Methods, 50(21):5126–5159, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2020.1726391.

Dai:2021:QRV

[DLT21] Xiaowen Dai, Erqian Li, and Maozai Tian. Quantile regression for varying coefficient spatial error models. Communications in Statistics: Theory and Methods, 50(10):2382–2397, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2019.1667396.

Ding:2023:CSD

[DLZN23] Chengjun Ding, Cong Liu, Qingpei Zang, and Yannan Niu. A class of small deviation theorems for Markov chains in bi-infinite random environment. Communications in Statistics: Theory and Methods, 52(3):693–701, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1921212.

Durot:2022:UMB

[DM22] Cécile Durot and Eni Musta. Uniform moment bounds for the standard estimators in the Cox proportional hazard model. Communications in Statistics: Theory and Methods, 51(21): 7452–7464, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2021.1873376.

DeLuca:2024:SEB

[DM24] Giuseppe De Luca and Jan R. Magnus. Shrinkage efficiency bounds: an extension. Communications in Statistics: Theory and Methods, 53(11):4147–4152, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2173976.

Dutta:2023:NRD

[DMD23] Ganesh Dutta, Nripes Kumar Mandal, and Premadhis Das. Nonexistence of robust designs against presence of more than one outlier in a restricted class. Communications in Statistics: Theory and Methods, 52(6):1668–1675, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1937651.

DeLuca:2021:PMQ

[DMP21] Giuseppe De Luca, Jan R. Magnus, and Franco Peracchi. Posterior moments and quantiles for the normal location model with Laplace prior. Communications in Statistics: Theory and Methods, 50(17):4039–4049, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1710756.

Dehtiar:2022:TME

[DMR22] Olena Dehtiar, Yuliya Mishura, and Kostiantyn Ralchenko. Two methods of estimation of the drift parameters of the Cox– Ingersoll–Ross process: Continuous observations. Communications in Statistics: Theory and Methods, 51(19):6818–6833, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2020.1866611.

Dewan:2021:SIS

[DNP21] Isha Dewan, Hon Keung Tony Ng, and Christian Paroissin. Special issue on Statistical Models and Methods in Reliability for the 11th international conference on mathematical methods in reliability (MMR 2019). Communications in Statistics: Theory and Methods, 50(16):3747–3748, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1952839.

Dong:2023:CSP

[Don23] Junyi Dong. Consistency of semi-parametric maximum likelihood estimator under identifiability conditions for the linear regression model with type I right censoring data. Communications in Statistics: Theory and Methods, 52(22):8152–8168, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2022.2059679.

Douge:2022:BET

[Dou22] Lahcen Douge. A Berry–Esseen theorem for sample quantiles under association. Communications in Statistics: Theory and Methods, 51(18):6515–6528, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1862871.

Donhauzer:2022:SLL

[DOV22] Illia Donhauzer, Andriy Olenko, and Andrei Volodin. Strong law of large numbers for functionals of random fields with unboundedly increasing covariances. Communications in Statistics: Theory and Methods, 51(20):6947–6962, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1868515.

Du:2022:ATS

[DP22] Lingjie Du and Tianxiao Pang. Asymptotic theory for a stochastic unit root model. Communications in Statistics: Theory and Methods, 51(5):1461–1487, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1764582.

Dafnis:2024:ICS

[DPP24] Spiros D. Dafnis, Theodoros Perdikis, and Georgios K. Papadopoulos. Improved chi-square control charts with weak-run rules. Communications in Statistics: Theory and Methods, 53(14):5248–5264, 2024. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2023.2215357.

DeGooijer:2021:MSK

[DR21] Jan G. De Gooijer and Hugo Reichardt. A multi-step kernel-based regression estimator that adapts to error distributions of unknown form. Communications in Statistics: Theory and Methods, 50(24):6211–6230, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1741625.

Devika:2022:CMC

[DR22] T. J. Devika and J. Ravichandran. A clustering method combining multiple range tests and K-means. Communications in Statistics: Theory and Methods, 51(21):7322–7339, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1872639.

denReijer:2021:CNF

[dRJO21] Ard H. J. den Reijer, Jan P. A. M. Jacobs, and Pieter W. Otter. A criterion for the number of factors. Communications in Statistics: Theory and Methods, 50(18):4293–4299, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1713376.

daSilva:2020:ASC

[dS20] João Lita da Silva. Almost sure convergence for weighted sums of ϕ-mixing random variables with applications. Communications in Statistics: Theory and Methods, 49(16):3882–3894, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2019.1594302.

daSilva:2022:SLL

[dS22] João Lita da Silva. Strong laws of large numbers for pairwise PQD random variables: a necessary condition. Communications in Statistics: Theory and Methods, 51(16):5455–5460, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2020.1841238.

Divandar:2023:IIN

[DS23] Mahin Sadat Divandar and Ghadir Sadeghi. The Itô integral and near-martingales in Riesz spaces. Communications in Statistics: Theory and Methods, 52(14):5068–5081, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.2003401.

daSilva:2024:ASC

[dS24a] João Lita da Silva. Almost sure convergence for weighted sums of pairwise PQD random variables. Communications in Statistics: Theory and Methods, 53(10):3435–3458, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2154129.

Dileepkumar:2024:POR

[DS24b] M. Dileepkumar and P. G. Sankaran. On proportional odds relevation transform and its applications. Communications in Statistics: Theory and Methods, 53(7):2369–2389, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2129994.

Das:2022:EPV

[DSB22] Pitambar Das, Garib Nath Singh, and Arnab Bandyopadhyay. Estimation of population variance in successive sampling in presence and absence of measurement error. Communications in Statistics: Theory and Methods, 51(14):4653–4666, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1818102.

daSilva:2022:GMD

[dSO22] Adolfo M. D. da Silva and C. E. G. Otiniano. Generalized moments of the distance between Poisson process events. Communications in Statistics: Theory and Methods, 51(8):2330– 2342, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2021.1968901.

Ding:2021:SLL

[DSY21] Chengjun Ding, Zhiyan Shi, and Weiguo Yang. The strong law of large numbers and Shannon–McMillan theorem for Markov chains indexed by an infinite tree with uniformly bounded degree in random environment. Communications in Statistics: Theory and Methods, 50(15):3573–3585, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1708398.

Dao:2021:FMR

[DT21a] Tung Dao and Minh-Ngoc Tran. Flexible multivariate regression density estimation. Communications in Statistics: Theory and Methods, 50(20):4703–4717, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1723633.

Ding:2021:ASC

[DT21b] Luyun Ding and Zhongquan Tan. Almost sure central limit theorem for the location and height of extreme order statistics and high values. Communications in Statistics: Theory and Methods, 50(8):1897–1910, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1656745.

DeCanditiis:2024:PFC

[DT24] Daniela De Canditiis and Marika Turdó. On the probability of (falsely) connecting two distinct components when learning a GGM. Communications in Statistics: Theory and Methods, 53(11):4107–4115, 2024. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2023.2173973.

Das:2022:SCC

[DTDB22] Birojit Das, Binod Chandra Tripathy, Piyali Debnath, and Baby Bhattacharya. Statistical convergence of complex uncertain triple sequence. Communications in Statistics: Theory and Methods, 51(20):7088–7100, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1871016.

Dunder:2021:MIC

[Dün21a]  Emre Dünder.  A modified information criterion for

model selection. Communications in Statistics: Theory and Methods, 50(11):2710–2721, 2021. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2019.1708395.

Dunder:2021:NCA

[Dün21b] Emre Dünder. A new correction approach for information criteria to detect outliers in regression modeling. Communications in Statistics: Theory and Methods, 50(10):2451–2465, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2020.1792497.

Duras:2022:FEP

[Dur22] Toni Duras. The fixed effects PCA model in a common principal component environment. Communications in Statistics: Theory and Methods, 51(6):1653–1673, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1765255.

Davies:2022:PTI

[DV22] Katherine F. Davies and William Volterman. Progressively Type-II censored competing risks data from the linear exponential distribution. Communications in Statistics: Theory and Methods, 51(5):1444–1460, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1764044.

Deng:2024:FSU

[DWZZ24] Luojia Deng, Jinhai Wu, Bin Zhang, and Yue Zhang. Feature screening for ultra-high-dimensional data via multiscale graph correlation. Communications in Statistics: Theory and Methods, 53(22):7942–7979, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2277130.

Xing:2020:AAT

[dXG20] Guo dong Xing and Xiaoli Gan. Asymptotic analysis of tail distortion risk measure under the framework of multivariate regular variation. Communications in Statistics: Theory and Methods, 49(12):2931–2941, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1584312.

Xing:2023:CCK

[dXY23] Guo dong Xing and Shanchao Yang. On the complete consistency of the kernel estimator of spot volatility. Communications in Statistics: Theory and Methods, 52(21):7576–7585, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2022.2049309.

Dong:2021:JDM

[DY21] Junyi Dong and Qiqing Yu. Joint distribution and marginal distribution methods for checking assumptions of generalized linear model. Communications in Statistics: Theory and Methods, 50(6):1456–1476, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1651860.

Dong:2024:DOP

[DYC^{+}24] Wenjie Dong, Yingjie Yang, Yingsai Cao, Jingru Zhang, and Sifeng Liu. Designing optimal proactive replacement strategies for degraded systems subject to two types of external shocks.

Communications in Statistics: Theory and Methods, 53(12): 4450–4468, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2023.2182179.

Dong:2020:ESR

[DZ20] Xiaofang Dong and Liangyong Zhang. Estimation of system reliability for exponential distributions based on L ranked set sampling. Communications in Statistics: Theory and Methods, 49(15):3650–3662, 2020. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2019.1691735.

Ding:2024:CLI

[DZ24] Xue Ding and Yong Zhang. Chover’s law of the iterated logarithm for weighted sums under sub-linear expectations. Communications in Statistics: Theory and Methods, 53(17):6055– 6075, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2023.2239399.

Zhao:2021:CSD

[dZySgYW21] Meng di Zhao, Zhi yan Shi, Wei guo Yang, and Bei Wang. A class of strong deviation theorems for the sequence of real valued random variables with respect to continuous-state non-homogeneous Markov chains. Communications in Statistics: Theory and Methods, 50(23):5475–5487, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1734838.

Dang:2023:CMA

[DZZ23a] Hui-Ping Dang, Sheng-Li Zhao, and Qian-Qian Zhao. Constructing minimum aberration split-plot designs via complementary designs when the subplot factors are more important. Communications in Statistics: Theory and Methods, 52(18): 6542–6560, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2022.2031221.

Ding:2023:IUN

[DZZ23b] Jianhua Ding, Hongyu Zhang, and Zhiqiang Zhang. Inferences for uncertain nonparametric regression by least absolute deviations. Communications in Statistics: Theory and Methods, 52(16):5640–5649, 2023. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2021.2016832.

Emami:2021:EDB

[EA21] Hadi Emami and Ali Aghamohammadi. Elliptical difference based ridge and Liu type estimators in partial linear measurement error models. Communications in Statistics: Theory and Methods, 50(21):4913–4933, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1472793.

Ertan:2022:NLT

[EA22] Esra Ertan and Kadri Ulas Akay. A new Liu-type estimator in binary logistic regression models. Communications in Statistics: Theory and Methods, 51(13):4370–4394, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1813777.

Emami:2023:RRE

[EA23] Hadi Emami and Korosh Arzideh. Robust ridge estimator in censored semiparametric linear models. Communications in Statistics: Theory and Methods, 52(17):5989–6007, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.2023573.

Esna-Ashari:2022:SNR

[EAAC22] Maryam Esna-Ashari, Mahdi Alimohammadi, and Erhard Cramer. Some new results on likelihood ratio ordering and aging properties of generalized order statistics. Communications in Statistics: Theory and Methods, 51(14):4667–4691, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2020.1818103.

Ercan:2020:LWS

[EAB20] Sinan Ercan, Yavuz Altin, and Çiğdem A. Bektaş. On lacunary weak statistical convergence of order α. Communications in Statistics: Theory and Methods, 49(7):1653–1664, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2018.1563185.

Esmaeli-Ayan:2022:TER

[EAMH22] Asghar Esmaeli-Ayan, Ahad Malekzadeh, and Farshin Hormozinejad. Testing equality of the regression coefficients in panel data models. Communications in Statistics: Theory and Methods, 51(20):7284–7296, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1872637.

Eidous:2020:NAS

[EAS20] Omar M. Eidous and Rima Abu-Shareefa. New approximations for standard normal distribution function. Communications in Statistics: Theory and Methods, 49(6):1357–1374, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2018.1563166.

Ercan:2021:CBO

[EB21] Sinan Ercan and Çigdem Asma Bektas. On λ-convergence and λ-boundedness of m-th order. Communications in Statistics: Theory and Methods, 50(14):3276–3285, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1692032.

Et:2022:DSB

[EBG22] Mikail Et, Vinod K. Bhardwaj, and Sandeep Gupta. On deferred statistical boundedness of order α. Communications in Statistics: Theory and Methods, 51(24):8786–8798, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1906434.

Elgawad:2020:LDR

[EBX20] M. A. Abd Elgawad, H. M. Barakat, and S. Xiong. Limit distributions of random record model in a stationary Gaussian sequence. Communications in Statistics: Theory and Methods, 49(5):1099–1119, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1554131.

Ebaid:2022:NEF

[EEAA22] Rasha Ebaid, Walid Elbadawy, Essam Ahmed, and Abdalla Abdelghaly. A new extension of the FGM copula with an application in reliability. Communications in Statistics: Theory and Methods, 51(9):2953–2961, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1785501. See correction [BAE23].

Et:2020:SCO

[EG20] Mikail Et and Hatice Gidemen. On Δ_{v}^{m}-statistical convergence of order α. Communications in Statistics: Theory and Methods, 49(14):3521–3529, 2020. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2019.1589520.

Egozcue:2023:NSO

[EG23] Martín Egozcue and Luis Fuentes García. A note on secondorder stochastic dominance for linear combinations of dependent Bernoulli random variables. Communications in Statistics: Theory and Methods, 52(10):3354–3360, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1971248.

Eralp:2023:MLE

[EGD23] Anil Eralp, Sahika Gokmen, and Rukiye Dagalp. Maximum likelihood estimation of spatial lag models in the presence of the error-prone variables. Communications in Statistics: Theory and Methods, 52(10):3229–3240, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2147795.

Eslami:2023:SNR

[EIAZ23a] Dariush Eslami, Hamidreza Izadbakhsh, Orod Ahmadi, and Marzieh Zarinbal. Spatial-nonparametric regression: an approach for monitoring image data. Communications in Statistics: Theory and Methods, 52(12):4114–4137, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1986535.

Eslami:2023:SMI

[EIAZ23b] Dariush Eslami, Hamidreza Izadbakhsh, Orod Ahmadi, and Marzieh Zarinbal. Statistical monitoring of image data using multi-channel functional principal component analysis. Communications in Statistics: Theory and Methods, 52(12):4165– 4182, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2021.1986539.

Ejaz:2023:ARA

[EJJ23] Muhammad Ejaz, Chaitanya Joshi, and Stephen Joe. Adversarial risk analysis for auctions using non-strategic play and level-k thinking: a general case of n bidders with regret. Communications in Statistics: Theory and Methods, 52(20):7146– 7164, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2022.2042023.

Eini:2022:TVG

[EK22] Esmat Jamshidi Eini and Hamid Khaloozadeh. Tail variance for Generalized Skew-Elliptical distributions. Communications in Statistics: Theory and Methods, 51(2):519–536, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1751853.

Ekhosuehi:2023:UCI

[Ekh23] Virtue U. Ekhosuehi. On the use of Cauchy integral formula for the embedding problem of discrete-time Markov chains. Communications in Statistics: Theory and Methods, 52(4): 973–987, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2021.1921806.

ElOmari:2023:MHO

[El 23] Mohamed El Omari. Mixtures of higher-order fractional Brownian motions. Communications in Statistics: Theory and Methods, 52(12):4200–4215, 2023. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2021.1986541.

Elmore:2022:MSE

[Elm22] Ryan Elmore. Modeling sums of exchangeable binary variables. Communications in Statistics: Theory and Methods, 51(18):6502–6514, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03610926.2020.1861467.

Elkadry:2021:OCS

[EM21] Alaa Elkadry and Gary C. McDonald. Operating characteristics of a subset selection procedure applied to a randomized response model for continuous data. Communications in Statistics: Theory and Methods, 50(1):161–179, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1633352.

Emami:2022:ALI

[EM22] Hadi Emami and Ali M. Mosammam. Assessment of local influence in spatial elliptical linear measurement error models. Communications in Statistics: Theory and Methods, 51(10): 3285–3300, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2020.1793204.

Emami:2020:RME

[Ema20] Hadi Emami. Restricted minimax estimation in semiparametric linear models. Communications in Statistics: Theory and Methods, 49(8):1793–1800, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1565783.

Emami:2024:DPL

[Ema24] Hadi Emami. Diagnostics for partially linear measurement error models. Communications in Statistics: Theory and Methods, 53(17):6224–6239, 2024. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2023.2242983.

Ercelik:2020:NDE

[EN20] Elif Erçelik and Mustafa Nadar. Nonparametric density estimation based on beta prime kernel. Communications in Statistics: Theory and Methods, 49(2):325–342, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1538458.

Elliriki:2022:MSQ

[ERAS22] Mamatha Elliriki, C. S. Reddy, Krishna Anand, and S. Saritha. Multi server queuing system with crashes and alternative repair strategies. Communications in Statistics: Theory and Methods, 51(23):8173–8185, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1889603.

Ezzahrioui:2022:RRC

[ES22] M’hamed Ezzahrioui and Elias Ould Saïd. On the robust regression for a censored response data in the single functional index model. Communications in Statistics: Theory and Methods, 51(15):5162–5186, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1836217.

Eid:2023:KKB

[EW23] Abdelrahman Eid and Nicolas Wicker. KBER: a kernel bandwidth estimate using the Ricci curvature. Communications in Statistics: Theory and Methods, 52(2):398–408, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1914099.

Efromovich:2024:ENP

[EW24] Sam Efromovich and Jiaju Wu. Efficient non parametric spectral density estimation with censored observations. Communications in Statistics: Theory and Methods, 53(18):6671–6694, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2023.2250484.

Zakrad:2022:PSS

[eZN22] Az eddine Zakrad and Abdelaziz Nasroallah. Perfect simulation of steady-state Markov chain on mixed state space. Communications in Statistics: Theory and Methods, 51(6):1569– 1587, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2021.1924783.

Fabian:2021:MMM

[Fab21] Zdenek Fabián. Mean, mode or median? The score mean. Communications in Statistics: Theory and Methods, 50(10): 2360–2370, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2019.1666142.

Fabian:2024:MVW

[Fab24] Zdenek Fabián. A measure of variability within parametric families of continuous distributions. Communications in Statistics: Theory and Methods, 53(10):3568–3580, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2155792.

Fakhfakh:2023:EFM

[Fak23a] Raouf Fakhfakh. Explicit free multiplicative law of large numbers. Communications in Statistics: Theory and Methods, 52(7):2031–2042, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03610926.2021.1944212.

Fakhfakh:2023:PAC

[Fak23b] Raouf Fakhfakh. On polynomials associated with Cauchy– Stieltjes kernel families. Communications in Statistics: Theory and Methods, 52(19):7009–7021, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2037647.

Fallah:2021:SIC

[FAN21] Adeleh Fallah, Akbar Asgharzadeh, and Hon Keung Tony Ng. Statistical inference for component lifetime distribution from coherent system lifetimes under a proportional reversed hazard model. Communications in Statistics: Theory and Methods, 50(16):3809–3833, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1824275.

Fan:2023:ICD

[Fan23] Yifan Fan. Inference in a class of directed random graph models with an increasing number of parameters. Communications in Statistics: Theory and Methods, 52(13):4491–4513, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1995432.

Farhadian:2021:NGS

[Far21] Reza Farhadian. A note on a generalization of a statistical matrix. Communications in Statistics: Theory and Methods, 50(12):2938–2946, 2021. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2019.1678646.

Farebrother:2023:SNM

[Far23] Richard William Farebrother. Supplementary notes on the minimax and orthogonal least squares line fitting procedures. Communications in Statistics: Theory and Methods, 52(10): 3350–3353, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2021.1971247.

Farebrother:2024:SNL

[Far24a] Richard William Farebrother. Supplementary notes on the least variance ratio estimator. Communications in Statistics: Theory and Methods, 53(4):1354–1357, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2100911.

Farhadian:2024:NCS

[Far24b] Reza Farhadian. A note on the Cauchy–Schwarz inequality for expectations. Communications in Statistics: Theory and Methods, 53(2):812–813, 2024. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2022.2092147.

Farhadian:2024:RFR

[Far24c] Reza Farhadian. Remark on “A further remark on the alternative expectation formula”. Communications in Statistics: Theory and Methods, 53(1):448–449, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2042028. See [SW21].

Fazekas:2023:CTN

[FBNP23] István Fazekas, Attila Barta, Csaba Noszály, and Bettina Porvázsnyik. A continuous-time network evolution model describing 3-interactions. Communications in Statistics: Theory and Methods, 52(11):4001–4020, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1985141.

Farhadian:2022:NHT

[FC22] Reza Farhadian and Brenton R. Clarke. A note on the helmert transformation. Communications in Statistics: Theory and Methods, 51(15):5258–5264, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1836223.

Fang:2020:OPL

[FCCD20] Longxiang Fang, Meifang Cheng, Daoxiang Cao, and Ying Ding. Ordering properties of largest order statistics from independent and heterogeneous Dagum populations. Communications in Statistics: Theory and Methods, 49(7):1768–1779, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2019.1565835.

Fatima-Ezzahra:2023:EDR

[FE23] Farah Fatima-Ezzahra. Estimation of the drift of Riemann– Liouville fractional Brownian motion. Communications in Statistics: Theory and Methods, 52(13):4719–4728, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1999475.

Ferreira:2023:NBE

[FF23] Helena Ferreira and Marta Ferreira. A new blocks estimator for the extremal index. Communications in Statistics: Theory and Methods, 52(21):7660–7668, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2050405.

Feng:2022:SCW

[FH22] Fengxiang Feng and Haiwu Huang. Strong convergence for weighted sums of END random variables under the sublinear expectations. Communications in Statistics: Theory and Methods, 51(22):7885–7896, 2022. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2021.1883654.

Fakhfakh:2024:RNE

[FH24] Raouf Fakhfakh and Marwa Hamza. Real natural exponential families and generalized orthogonality. Communications in Statistics: Theory and Methods, 53(16):5871–5889, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2235447.

Fang:2020:RRS

[FHHW20] Juntao Fang, Zhen He, Shuguang He, and Guodong Wang. The robustness of response surface designs with errors in factor levels. Communications in Statistics: Theory and Methods, 49(10):2365–2383, 2020. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2019.1576883.

Fang:2024:OTS

[FHL^{+}24] Maoda Fang, Sigui Hu, Qiude Li, Huijuan Chen, Rongjin Long, and Maoyue Ye. Optimal truncated sequential test for exponential distribution. Communications in Statistics: Theory and Methods, 53(22):7893–7907, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2274811.

Fang:2023:ULB

[FHZC23] Liang Fang, Yiping Hong, Zaiying Zhou, and Wenhui Chen. Uncertain logistic and Box–Cox regression analysis with maximum likelihood estimation. Communications in Statistics: Theory and Methods, 52(1):19–38, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1908562.

Fang:2020:SMC

[FL20a] Rui Fang and Xiaohu Li. A stochastic model of cyber attacks with imperfect detection. Communications in Statistics: Theory and Methods, 49(9):2158–2175, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1568489.

Ferreira:2020:EIF

[FL20b] Paulo H. Ferreira and Francisco Louzada. Extending the inference function for augmented margins method to implement trivariate Clayton copula-based SUR Tobit models. Communications in Statistics: Theory and Methods, 49(6):1375–1401, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2018.1563167.

Feng:2021:WQR

[FL21] Hailin Feng and Qianqian Luo. A weighted quantile regression for nonlinear models with randomly censored data. Communications in Statistics: Theory and Methods, 50(18):4167–4179, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2020.1713364.

Flimmel:2020:CAP

[Fli20] Samuel Flimmel. Clipped AR(p) with unknown threshold. Communications in Statistics: Theory and Methods, 49(21): 5115–5122, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2019.1614192.

Fu:2024:PLD

[FLW24] Ke-Ang Fu, Yang Liu, and Jiangfeng Wang. Precise large deviations in a non stationary risk model with arbitrary dependence between subexponential claim sizes and waiting times. Communications in Statistics: Theory and Methods, 53(11): 4116–4126, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2023.2173974.

Fujiwara:2020:PVC

[FM20] Kohei Fujiwara and Shun Matsuura. Prediction variance of a central composite design with missing observation. Communications in Statistics: Theory and Methods, 49(24):6016–6031, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2019.1625925.

Fauzi:2024:BFE

[FM24] Rizky Reza Fauzi and Yoshihiko Maesono. Boundaryfree estimators of the mean residual life function for data on general interval. Communications in Statistics: Theory and Methods, 53(11):3958–3972, 2024. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2023.2168484.

Forkman:2024:TCT

[FMHP24] Johannes Forkman, Waqas Ahmed Malik, Steffen Hadasch, and Hans-Peter Piepho. Testing components of two-way interaction in multi-environment trials. Communications in Statistics: Theory and Methods, 53(5):1716–1735, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2108058.

Fontanelli:2022:BRF

[FMM^{+}22] Oscar Fontanelli, Pedro Miramontes, Ricardo Mansilla, Germinal Cocho, and Wentian Li. Beta rank function: a smooth double-Pareto-like distribution. Communications in Statistics: Theory and Methods, 51(11):3645–3668, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1800739.

Farzammehr:2022:BAG

[FMZM22] Mohadeseh Alsadat Farzammehr, Mohsen Mohammadzadeh, Mohammad Reza Zadkarami, and Geoffrey J. McLachlan. Bayesian analysis of generalized linear mixed models with spatial correlated and unrestricted skew normal errors. Communications in Statistics: Theory and Methods, 51(24):8476–8498, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2021.1897843.

Fowlie:2022:OBA

[Fow22] Andrew Fowlie. Objective Bayesian approach to the Jeffreys-Lindley paradox. Communications in Statistics: Theory and Methods, 51(19):6760–6765, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1866206.

Fowlie:2023:NPL

[Fow23] Andrew Fowlie. Neyman–Pearson lemma for Bayes factors. Communications in Statistics: Theory and Methods, 52(15): 5379–5386, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2021.2007265.

Feizi:2023:NAS

[FP23] Bahareh Feizi and Ahmad Pourdarvish. Nonlinear autoregressive stochastic frontier model with dynamic technical inefficiency in panel data. Communications in Statistics: Theory and Methods, 52(4):1058–1075, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1923746.

Frey:2020:BEU

[Fre20a] Jesse Frey. Bounds on the efficiency of unbalanced ranked-set sampling. Communications in Statistics: Theory and Methods, 49(1):243–256, 2020. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2018.1543769.

Frey:2020:RAK

[Fre20b] Jesse Frey. Refined asymptotic Kolmogorov–Smirnov tests for the case of finite support. Communications in Statistics: Theory and Methods, 49(23):5829–5841, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1622726.

From:2023:IHI

[Fro23] Steven G. From. An improved Hoeffding’s inequality of closed form using refinements of the arithmetic mean– geometric mean inequality. Communications in Statistics: Theory and Methods, 52(14):4792–4798, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2012.756913.

Fradi:2022:BCA

[FS22] Anis Fradi and Chafik Samir. Bayesian cluster analysis for registration and clustering homogeneous subgroups in multidimensional functional data. Communications in Statistics: Theory and Methods, 51(7):2242–2258, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1772979.

Fradi:2022:SAB

[FSB22] Anis Fradi, Chafik Samir, and François Bachoc. A scalable approximate Bayesian inference for high-dimensional Gaussian processes. Communications in Statistics: Theory and Methods, 51(17):5937–5956, 2022. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2020.1850793.

Feng:2022:EAH

[FSS22] Yanqin Feng, Jianguo Sun, and Lingli Sun. Estimation of the additive hazards model with linear inequality restrictions based on current status data. Communications in Statistics: Theory and Methods, 51(1):68–81, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1742922.

Foroghi:2023:EFC

[FTAL23] Farid Foroghi, Saeid Tahmasebi, Mahmoud Afshari, and Fazlollah Lak. Extensions of fractional cumulative residual entropy with applications. Communications in Statistics: Theory and Methods, 52(20):7350–7369, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2044493.

Feng:2022:RMD

[FTTY22] Xuefeng Feng, Jiayin Tang, Qitao Tan, and Zekai Yin. Reliability model for dual constant-stress accelerated life test with Weibull distribution under Type-I censoring scheme. Communications in Statistics: Theory and Methods, 51(24):8579– 8597, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2021.1900868.

Fu:2024:MSG

[Fu24] Xinyu Fu. Model selection of Gaussian mixture process and its application. Communications in Statistics: Theory and Methods, 53(5):1576–1589, 2024. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2022.2104875.

Furmanczyk:2023:MPT

[Fur23] Konrad Furmańczyk. A multitest procedure for testing MTP2 for Gaussian distributions. Communications in Statistics: Theory and Methods, 52(24):8865–8874, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2081339.

Fushiki:2022:NPE

[Fus22] Tadayoshi Fushiki. A note on the properties of estimators in missing data analysis. Communications in Statistics: Theory and Methods, 51(17):6144–6149, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1854305.

Fabio:2023:DTM

[FVCdC23] Lizandra C. Fabio, Cristian Villegas, Jalmar M. F. Carrasco, and Mário de Castro. Diagnostic tools for a multivariate negative binomial model for fitting correlated data with overdispersion. Communications in Statistics: Theory and Methods, 52(6):1833–1853, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1939380.

Fu:2022:CEN

[FW22] Jingxue Fu and Lan Wu. Consistent estimation of the number of regimes in Markov-switching autoregressive models. Communications in Statistics: Theory and Methods, 51(8):2496– 2518, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2020.1777304.

Fu:2023:MDH

[FW23] Ke-Ang Fu and Jiangfeng Wang. Moderate deviations for a Hawkes-type risk model with arbitrary dependence between claim sizes and waiting times. Communications in Statistics: Theory and Methods, 52(17):6266–6274, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2027451.

Feng:2021:CCM

[FWWH21] Fengxiang Feng, Dingcheng Wang, Qunying Wu, and Haiwu Huang. Complete and complete moment convergence for weighted sums of arrays of rowwise negatively dependent random variables under the sub-linear expectations. Communications in Statistics: Theory and Methods, 50(3):594–608, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1639747.

Flint:2024:CMP

[FWX24] Ian Flint, Yan Wang, and Aihua Xia. On the Conway– Maxwell–Poisson point process. Communications in Statistics: Theory and Methods, 53(16):5687–5705, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2229028.

Fang:2022:EOT

[FZ22a] Yuexin Fang and Yong Zhou. Estimation for optimal treatment regimes with survival data under semiparametric model. Communications in Statistics: Theory and Methods, 51(4): 883–894, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2020.1808686.

Feng:2022:NMA

[FZ22b] Zhangmei Feng and Jiamin Zhang. Nonparametric K-means algorithm with applications in economic and functional data. Communications in Statistics: Theory and Methods, 51(2): 537–551, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2020.1752383.

Fang:2024:SCL

[FZD24] Longxiang Fang, Qi Zheng, and Ying Ding. Stochastic comparisons of the largest and smallest claim amounts with heterogeneous survival exponentiated location-scale distributed claim severities. Communications in Statistics: Theory and Methods, 53(21):7541–7559, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2269440.

Fang:2024:RAS

[FZL24] Longxiang Fang, Shuai Zhang, and Jinling Lu. On reliability analysis in k-out-of-n systems under Archimedean copula dependence. Communications in Statistics: Theory and Methods, 53(1):365–377, 2024. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2022.2078840.

Fan:2023:NPE

[FZW23] Ruiya Fan, Shuguang Zhang, and Yaohua Wu. Nonconcave penalized M-estimation for the least absolute relative errors model. Communications in Statistics: Theory and Methods, 52(4):1118–1135, 2023. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2021.1923749.

Ghanbarzadeh:2020:NWA

[GA20] Mitra Ghanbarzadeh and Mina Aminghafari. A novel wavelet artificial neural networks method to predict nonstationary time series. Communications in Statistics: Theory and Methods, 49(4):864–878, 2020. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2018.1549259.

Gupt:2022:OSG

[GA22] Bhuwaneshwar Kumar Gupt and Md. Irphan Ahamed. Optimum stratification for a generalized auxiliary variable proportional allocation under a superpopulation model. Communications in Statistics: Theory and Methods, 51(10):3269–3284, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2020.1793203.

Goodarzi:2023:REB

[GA23] F. Goodarzi and M. Amini. Reliability and expectation bounds based on Hardy’s inequality. Communications in Statistics: Theory and Methods, 52(9):2983–2997, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1966037.

Goyal:2024:SMP

[GAG24] Anupama Goyal, Sangeeta Arora, and Anju Goyal. A stratified modified probability proportional to size sampling technique. Communications in Statistics: Theory and Methods, 53(23):8525–8542, 2024. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2023.2292969.

Gaigall:2022:HBK

[Gai22] Daniel Gaigall. Hoeffding–Blum–Kiefer–Rosenblatt independence test statistic on partly not identically distributed data. Communications in Statistics: Theory and Methods, 51(12): 4006–4028, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2020.1805767.

Gajda:2022:GML

[Gaj22] Janusz Gajda. Generalized Mittag-Leffler lévy process and its connections to first passage times of lévy subordinators. Communications in Statistics: Theory and Methods, 51(13):4500– 4508, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2020.1817486.

Gao:2023:ARA

[Gao23] Shan Gao. Availability and reliability analysis of a retrial system with warm standbys and second optional repair service. Communications in Statistics: Theory and Methods, 52(4): 1039–1057, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2021.1922702.

Gargoum:2021:MAD

[Gar21] A. S. Gargoum. Modeling atmospheric dispersion: Uncertainty management of release height after a nuclear accident. Communications in Statistics: Theory and Methods, 50(19): 4639–4648, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2020.1722844.

Gaunt:2021:SPC

[Gau21a] Robert E. Gaunt. A simple proof of the characteristic function of Student’s t-distribution. Communications in Statistics: Theory and Methods, 50(14):3380–3383, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1702695.

Gaunt:2021:SMD

[Gau21b] Robert E. Gaunt. Stein’s method and the distribution of the product of zero mean correlated normal random variables. Communications in Statistics: Theory and Methods, 50(2): 280–285, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2019.1634210.

Guler:2021:RIF

[GB21] Nesrin Güler and Melek Eris Büyükkaya. Rank and inertia formulas for covariance matrices of BLUPs in general linear mixed models. Communications in Statistics: Theory and Methods, 50(21):4997–5012, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1599950.

Guler:2023:IRA

[GB23a] Nesrin Güler and Melek Eris Büyükkaya. Inertia and rank approach in transformed linear mixed models for comparison of BLUPs. Communications in Statistics: Theory and Methods, 52(9):3108–3123, 2023. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2021.1967397.

Guo:2023:RPP

[GB23b] Yiping Guo and Howard Bondell. On robust probabilistic principal component analysis using multivariate t-distributions. Communications in Statistics: Theory and Methods, 52(23):8261–8279, 2023. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2022.2060512.

Gonzalez-Barrios:2023:CMI

[GBGPNR23] José M. González-Barrios, Eduardo Gutiérrez-Peña, Juan D. Nieves, and Raúl Rueda. A characterization of multivariate independence using copulas. Communications in Statistics: Theory and Methods, 52(16):5716–5726, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.2017458.

Ghanbari:2023:SCS

[GBH23] Fariba Ghanbari, Ghobad Barmalzan, and Reza Hashemi. Stochastic comparisons of series and parallel systems with dependent log-logistic components. Communications in Statistics: Theory and Methods, 52(12):4259–4282, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1990951.

Graczyk:2021:RAW

[GC21] Małgorzata Graczyk and Bronisław Ceranka. Recent accomplishments in weighing designs. Communications in Statistics: Theory and Methods, 50(21):5013–5021, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1639748.

Goswami:2022:QGG

[GC22a] Veena Goswami and M. L. Chaudhry. The queue Geo^{X} /G/1/N+1 revisited. Communications in Statistics: Theory and Methods, 51(10):3181–3201, 2022. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2020.1790602.

Guerry:2022:MPS

[GC22b] Marie-Anne Guerry and Philippe Carette. On monotonically proceeding structures and stepwise increasing transition matrices of Markov chains. Communications in Statistics: Theory and Methods, 51(1):51–67, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1742921.

Guha:2022:SSP

[GC22c] Pritha Guha and Biman Chakraborty. On a scale-scale plot for comparing multivariate distributions. Communications in Statistics: Theory and Methods, 51(5):1270–1289, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1759639.

Graczyk:2024:ROS

[GC24a] Małgorzata Graczyk and Bronisław Ceranka. Regular Doptimal spring balance weighing designs with correlated errors. Communications in Statistics: Theory and Methods, 53(9):3425–3433, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03610926.2022.2154128.

Gupta:2024:GWE

[GC24b] Nitin Gupta and Santosh Kumar Chaudhary. On general weighted extropy of ranked set sampling. Communications in Statistics: Theory and Methods, 53(12):4428–4441, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2179888.

Garrido:2021:NAF

[GCCJO^{+}21] Arturo Garrido, Raquel Caro-Carretero, Jesús R. Jimenez-Octavio, Alberto Carnicero, and Miguel Such. A new approach to fitting the three-parameter Weibull distribution: an application to glass ceramics. Communications in Statistics: Theory and Methods, 50(14):3403–3420, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1702698.

Guo:2023:GAR

[GCFS23] Zijun Guo, Mengxing Chen, Yali Fan, and Yan Song. A general adaptive ridge regression method for generalized linear models: an iterative re-weighting approach. Communications in Statistics: Theory and Methods, 52(18):6420–6443, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2028841.

Garg:2021:TCM

[GD21] Mansi Garg and Isha Dewan. Testing for change in mean for associated random variables. Communications in Statistics: Theory and Methods, 50(16):3834–3850, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1873375.

Gokmen:2022:MME

[GDK22] Sahika Gokmen, Rukiye Dagalp, and Serdar Kilickaplan. Multicollinearity in measurement error models. Communications in Statistics: Theory and Methods, 51(2):474–485, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1750654.

Gramosa:2020:BAZ

[GdNM20] Alexandre Henrique Quadros Gramosa, Fernando Ferraz do Nascimento, and Fidel Ernesto Castro Morales. A Bayesian approach to zero-inflated data in extremes. Communications in Statistics: Theory and Methods, 49(17):4150–4161, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1594305.

Gomes:2015:EPM

[GDSC15] Antonio E. Gomes, Cibele Q. Da-Silva, and Gauss M. Cordeiro. The exponentiated G Poisson model. Communications in Statistics: Theory and Methods, 44(20):4217–4240, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). See correction [AE23b].

Gomez-Deniz:2022:PLC

[GDSJ22] Emilio Gómez-Déniz, José María Sarabia, and Vanesa Jordá. Parametric Lorenz curves based on the beta system of distributions. Communications in Statistics: Theory and Methods, 51(23):8371–8390, 2022. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2021.1894449.

Gildeh:2020:EME

[GG20] Bahram Sadeghpour Gildeh and Zainab Abbasi Ganji. The effect of measurement error on the process incapability index. Communications in Statistics: Theory and Methods, 49(3): 552–566, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2018.1543777.

Guler:2021:SRP

[GG21] Huseyin Guler and Ebru Ozgur Guler. Sparsely restricted penalized estimators. Communications in Statistics: Theory and Methods, 50(7):1656–1670, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1682164.

Ganji:2023:NPC

[GG23a] Zainab Abbasi Ganji and Bahram Sadeghpour Gildeh. A new process capability index for simple linear profile. Communications in Statistics: Theory and Methods, 52(11):3879–3894, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2021.1980804.

Gharari:2023:NSO

[GG23b] Fatemeh Gharari and Masoud Ganji. A new stochastic order based on discrete Laplace transform and some ordering results of the order statistics. Communications in Statistics: Theory and Methods, 52(6):1963–1980, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1943442.

Ghettab:2024:APA

[GG24a] Sarah Ghettab and Zohra Guessoum. Asymptotic properties of asymmetric kernel estimators for non-negative and censored data. Communications in Statistics: Theory and Methods, 53(8):2977–3004, 2024. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2022.2150059.

Grbac:2024:LEP

[GG24b] Neven Grbac and Tihana Galinac Grbac. Letter to the editor: on the paper “The double Pareto–Lognormal distribution — a new parametric model for size distributions” and its correction. Communications in Statistics: Theory and Methods, 53(11):4207–4209, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2174788.

Ghosh:2018:GKD

[GH18] Indranil Ghosh and G. G. Hamedani. The gamma-Kumaraswamy distribution: an alternative to gamma distribution. Communications in Statistics: Theory and Methods, 47(9):2056–2072, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03610926.2015.1122055.

See correction [AE23a].

Ghosh:2022:SCE

[GH22] Indranil Ghosh and G. G. Hamedani. On some counter examples of the bivariate and multivariate normal distributions: a brief survey. Communications in Statistics: Theory and Methods, 51(14):4957–4972, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1828466.

Ghosh:2023:APL

[Gho23] Indranil Ghosh. On the asymptotic properties of the likelihood estimates and some inferential issues related to hidden truncated Pareto (type II) model. Communications in Statistics: Theory and Methods, 52(15):5133–5144, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.2004422.

Ghosh:2020:GCC

[GK20a] Amit Ghosh and Chanchal Kundu. On generalized conditional cumulative residual inaccuracy measure. Communications in Statistics: Theory and Methods, 49(6):1402–1421, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1563168.

Guambe:2020:OIC

[GK20b] Calisto Guambe and Rodwell Kufakunesu. Optimal investmentconsumption and life insurance with capital constraints. Communications in Statistics: Theory and Methods, 49(3):648– 669, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2018.1549246.

Guo:2024:ORC

[GK24] Huijun Guo and Junke Kou. Optimal rates of convergence for nonparametric regression estimation under anisotropic hölder condition. Communications in Statistics: Theory and Methods, 53(2):687–699, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2091781.

Goktas:2021:P

[GKA21] Atila Göktas, Selahattin Kaçiranlar, and Yilmaz Akdi. Preface. Communications in Statistics: Theory and Methods, 50 (7):1531, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2021.1890521.

Guo:2021:OIR

[GKS21] Mengmeng Guo, Xiu Kan, and Huisheng Shu. Optimal investment and reinsurance problem with jump-diffusion model. Communications in Statistics: Theory and Methods, 50(5): 1082–1098, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2019.1646764.

Ghosh:2021:TGL

[GKV21] S. Ghosh, K. K. Kataria, and P. Vellaisamy. On transmuted generalized linear exponential distribution. Communications in Statistics: Theory and Methods, 50(9):1978–2000, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1655577.

Guambe:2021:OAA

[GKvZB21] Calisto Guambe, Rodwell Kufakunesu, Gusti van Zyl, and Conrad Beyers. Optimal asset allocation for a DC plan with partial information under inflation and mortality risks. Communications in Statistics: Theory and Methods, 50(9):2048– 2061, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2019.1657458.

Gao:2020:SIU

[GL20] Xianli Gao and Qiang Liu. Sparsity identification in ultra-high dimensional quantile regression models with longitudinal data. Communications in Statistics: Theory and Methods, 49(19): 4712–4736, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2019.1604966.

Gaitan:2021:FSO

[GL21] Rodrigo Saul Gaitan and Keh-Shin Lii. The first and second order moment structure of an inhomogeneous gamma point process. Communications in Statistics: Theory and Methods, 50(3):582–593, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03610926.2019.1639746.

Gao:2022:NPE

[GLB22] Rui Gao, Yaqiong Li, and Yanfei Bai. Numerical pricing of exchange option with stock liquidity under Bayesian statistical method. Communications in Statistics: Theory and Methods, 51(10):3312–3333, 2022. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2020.1793364.

Gao:2024:PLD

[GLK24] Qingwu Gao, Wen Li, and Linmin Kan. Precise large deviations of aggregate claims in bidimensional risk model with dependence structures. Communications in Statistics: Theory and Methods, 53(22):8062–8075, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2278430.

Guo:2023:HWL

[GLL23] Xiaofan Guo, Shan Li, and Xinpeng Li. On the Hartman– Wintner law of the iterated logarithm under sublinear expectation. Communications in Statistics: Theory and Methods, 52(17):6126–6135, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03610926.2022.2026394.

Ge:2023:ESM

[GLLL23] Hao Ge, Xingyi Li, Xun Li, and Zhongfei Li. Equilibrium strategy for a multi-period weighted mean-variance portfolio selection in a Markov regime-switching market with uncertain time-horizon and a stochastic cash flow. Communications in Statistics: Theory and Methods, 52(6):1797–1832, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1939379.

Geng:2024:TBR

[GLW24] Bingzhen Geng, Zaiming Liu, and Shijie Wang. On tail behavior of randomly weighted sums of dependent subexponential random variables. Communications in Statistics: Theory and Methods, 53(5):1653–1668, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2107224.

Ghosh:2020:NTE

[GM20a] Shyamal Ghosh and Murari Mitra. A new test for exponentiality against HNBUE alternatives. Communications in Statistics: Theory and Methods, 49(1):27–43, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1528370.

Gobbi:2020:MMP

[GM20b] Fabio Gobbi and Sabrina Mulinacci. Mixing and moments properties of a non-stationary copula-based Markov process. Communications in Statistics: Theory and Methods, 49(18): 4559–4570, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2019.1602653.

Gul:2021:RRM

[GM21] Ahtasham Gul and Muhammad Mohsin. Recurrence relations for moments of order statistics from half logistictruncated exponential distribution. Communications in Statistics: Theory and Methods, 50(17):3889–3902, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1710753.

Girone:2023:PND

[GMC^{+}23] Giovanni Girone, Antonella Massari, Francesco Campobasso, Angela Maria D’Uggento, Claudia Marin, and Fabio Manca. A proposal of new disnormality indexes. Communications in Statistics: Theory and Methods, 52(5):1548–1563, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1930053.

Goncalves:2020:SCP

[GML20] Esmeralda Gonçalves and Nazaré Mendes-Lopes. Signed compound Poisson integer-valued GARCH processes. Communications in Statistics: Theory and Methods, 49(22):5468–5492, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2019.1619767.

Ghasemian:2024:IPE

[GN24] Poune Ghasemian and Rassoul Noorossana. The inertial properties of EWMA control charts. Communications in Statistics: Theory and Methods, 53(12):4542–4555, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2184190.

Giri:2023:AIF

[GND^{+}23] Rajib Lochan Giri, Asok K. Nanda, Mahua Dasgupta, Satya Kr. Misra, and Subarna Bhattacharjee. On ageing intensity function of some Weibull models. Communications in Statistics: Theory and Methods, 52(1):227–262, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1910845.

Genc:2024:LRU

[GÖ24] Murat Genç and M. Revan Özkale. Lasso regression under stochastic restrictions in linear regression: an application to genomic data. Communications in Statistics: Theory and Methods, 53(8):2816–2839, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2149243.

Gondaliya:2021:OMB

[Gon21a] Jigneshkumar Gondaliya. Optimal minimal balanced crossover designs in first and second carryover effects. Communications in Statistics: Theory and Methods, 50(17):4119–4133, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1713370.

Gong:2021:OEE

[Gon21b] Huan Gong. Oracally efficient estimation and testing for an ARCH model with trend. Communications in Statistics: Theory and Methods, 50(14):3384–3402, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1702696.

Goodarzi:2024:RCV

[Goo24] F. Goodarzi. Results on conditional variance in parallel system and lower bounds for varextropy. Communications in Statistics: Theory and Methods, 53(5):1590–1610, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2105892.

Goswami:2022:INA

[Gos22] Suchismita Goswami. Influential nodes and anomalous topic activities in social networks using multivariate time series and topic modeling. Communications in Statistics: Theory and Methods, 51(10):3382–3407, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1821891.

Garg:2020:UCV

[GP20] Neha Garg and Menakshi Pachori. Use of coefficient of variation in calibration estimation of population mean in stratified sampling. Communications in Statistics: Theory and Methods, 49(23):5842–5852, 2020. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2019.1622729.

Gautam:2021:AMB

[GPD21] Choudhury Gautam, Kalita Priyanka, and Selvamuthu Dharmaraja. Analysis of a model of batch arrival single server queue with random vacation policy. Communications in Statistics: Theory and Methods, 50(22):5314–5357, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1728329.

Gopinath:2023:CRI

[GPM23] Pratheesh P. Gopinath, Rajender Parsad, and B. N. Mandal. Construction of resolvable incomplete block designs for estimating main effects with full efficiency. Communications in Statistics: Theory and Methods, 52(6):1923–1936, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1942047.

Guo:2021:PDR

[GQW21] Jie Guo, Xiaosong Qian, and Guojing Wang. Pricing default risk in mortgage-backed securities under a regimeswitching reduced-form model. Communications in Statistics: Theory and Methods, 50(9):2117–2135, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1659971.

Gou:2024:FCP

[GRBL24] Jiangtao Gou, Karen Ruth, Stanley Basickes, and Samuel Litwin. A fortune cookie problem: a test for nominal data whether two samples are from the same population of equally likely elements. Communications in Statistics: Theory and Methods, 53(9):3063–3077, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2150062.

Grimm:2024:OCI

[Gri24] Wolfgang M. Grimm. Optimal choice of IHS-type of transformations for log-linear modeling. Communications in Statistics: Theory and Methods, 53(22):7980–8008, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2277671.

Ghanbari:2024:UCI

[GS24] Bahareh Ghanbari and Esmaeil Shirazi. Using copula information in wavelet estimation of bivariate density function based on censorship observations. Communications in Statistics: Theory and Methods, 53(5):1810–1824, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2113798.

Grima:2024:GFG

[GSED24] Pere Grima, José A. Sánchez-Espigares, and Pedro Delicado. Goodness-of-fit graphical assessment for a broad family of unimodal distributions. Communications in Statistics: Theory and Methods, 53(1):396–418, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2082479.

Gleaton:2022:APM

[GSH22] James U. Gleaton, Ping Sa, and Sami Hamid. Asymptotic properties of MLE’s of parameters of exponentiated exponential lifetime distributions. Communications in Statistics: Theory and Methods, 51(22):7856–7870, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1882495.

George:2022:EIP

[GT22] Roshini George and S. Thobias. Exact inference for the parameters of absolutely continuous trivariate exponential location-scale model. Communications in Statistics: Theory and Methods, 51(17):6021–6031, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1851720.

Gulfirat:2023:SUC

[Gül23] Mustafa Gülfirat. On μ-statistical uniform convergence and Dini’s theorem. Communications in Statistics: Theory and Methods, 52(21):7744–7751, 2023. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2022.2059684.

Guo:2021:OPU

[Guo21] Zhidong Guo. Option pricing under the Heston model where the interest rate follows the Vasicek model. Communications in Statistics: Theory and Methods, 50(12):2930–2937, 2021.

CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1678643.

Gasana:2024:MLB

[GvRS24] Emelyne Umunoza Gasana, Dietrich von Rosen, and Martin Singull. Moments of the likelihood-based discriminant function. Communications in Statistics: Theory and Methods, 53(3):1122–1134, 2024. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2022.2100909.

Gehman:2021:TPS

[GW21] Andrew J. Gehman and William W. S. Wei. Testing for poolability of the space-time autoregressive moving-average model. Communications in Statistics: Theory and Methods, 50(20): 4787–4808, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2020.1725052.

Gharaibeh:2023:NNL

[GW23] Mohammed M. Gharaibeh and Haiyan Wang. A new nonparametric lack-of-fit test of nonlinear regression in presence of heteroscedastic variances. Communications in Statistics: Theory and Methods, 52(22):7886–7914, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2051051.

Ge:2024:CMC

[GWD24] Meimei Ge, Yongfeng Wu, and Xin Deng. Complete fmoment convergence for randomly weighted sums of extended negatively dependent random variables. Communications in Statistics: Theory and Methods, 53(20):7384–7404, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2264992.

Grayling:2021:EGS

[GWM21] Michael J. Grayling, James M. S. Wason, and Adrian P. Mander. Exact group sequential designs for two-arm experiments with Poisson distributed outcome variables. Communications in Statistics: Theory and Methods, 50(1):18–34, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1628273.

Gao:2021:ELA

[GX21] Pengli Gao and Zhiming Xia. Empirical likelihood approach for change-point estimation based on residuals in piecewise linear models. Communications in Statistics: Theory and Methods, 50(10):2371–2381, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1667395.

Guan:2020:PEA

[GZ20] Jing Guan and Yunfeng Zhao. Parameter estimation approaches to tackling measurement error and multicollinearity in ordinal probit models. Communications in Statistics: Theory and Methods, 49(16):3835–3859, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1592193.

Goudarzi:2021:BMM

[GZ21] Monir Goudarzi and Mohammad Zokaei. Bayesian modeling of multivariate loss reserving data based on scale mixtures of multivariate normal distributions: estimation and case influence diagnostics. Communications in Statistics: Theory and Methods, 50(21):4934–4962, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1520883.

Gao:2023:ECR

[GZ23a] Zhiqiang Gao and Xiaoyan Zhang. Exact convergence rate of the local limit theorem for a branching random walk in a time-dependent random environment on d-dimensional integer lattice. Communications in Statistics: Theory and Methods, 52(4):988–1011, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1921807.

Guo:2023:CLT

[GZ23b] Shuang Guo and Yong Zhang. Central limit theorem for linear processes generated by m-dependent random variables under the sub-linear expectation. Communications in Statistics: Theory and Methods, 52(18):6407–6419, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2028840.

Guo:2024:NMT

[GZ24] Xiangyu Guo and Fukang Zhu. A new method of testing mutual independence. Communications in Statistics: Theory and Methods, 53(17):6102–6115, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2239402.

Gzyl:2021:CCT

[Gzy21] Henryk Gzyl. Construction of contingency tables by maximum entropy in the mean. Communications in Statistics: Theory and Methods, 50(20):4778–4786, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1723639.

Gzyl:2022:PRM

[Gzy22] Henryk Gzyl. Prediction in Riemannian metrics derived from divergence functions. Communications in Statistics: Theory and Methods, 51(2):552–568, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1752384.

Gao:2021:DFB

[GZZ21] Xianjie Gao, Chao Zhang, and Hongwei Zhang. Dimensionfree bounds for largest singular values of matrix Gaussian series. Communications in Statistics: Theory and Methods, 50(10):2419–2428, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03610926.2019.1670846.

Huda:2020:MDE

[HA20] S. Huda and Fatemah Alqallaf. Minimax designs for estimating slopes in a trigonometric regression model. Communications in Statistics: Theory and Methods, 49(21):5332–5341, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2019.1617881.

Haq:2022:AIB

[HA22] Abdul Haq and Shareen Akhtar. Auxiliary information based maximum EWMA and DEWMA charts with variable sampling intervals for process mean and variance. Communications in Statistics: Theory and Methods, 51(12):3985–4005, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2020.1805766.

Hussain:2020:SNC

[HAA^{+}20] Sajid Hussain, Rashid Ahmed, Muhammad Aslam, Azhar Shah, and H. M. Kashif Rasheed. Some new construction of circular weakly balanced repeated measurements designs in periods of two different sizes. Communications in Statistics: Theory and Methods, 49(9):2253–2263, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1570263.

Husseiny:2024:IMO

[HABE24] I. A. Husseiny, M. A. Alawady, H. M. Barakat, and M. A. Abd Elgawad. Information measures for order statistics and their concomitants from Cambanis bivariate family. Communications in Statistics: Theory and Methods, 53(3):865–881, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2093909.

Haq:2023:EFP

[HAK23] Abdul Haq, Mohsin Abbas, and Manzoor Khan. Estimation of finite population distribution function in a complex survey sampling. Communications in Statistics: Theory and Methods, 52(8):2574–2596, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1955386.

Hamura:2023:BPE

[Ham23] Yasuyuki Hamura. Bayesian point estimation and predictive density estimation for the binomial distribution with a restricted probability parameter. Communications in Statistics: Theory and Methods, 52(11):3767–3794, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1980046.

Hamura:2024:SPP

[Ham24] Yasuyuki Hamura. Short proof of posterior robustness: an illustration of basic ideas in a simple case. Communications in Statistics: Theory and Methods, 53(20):7298–7310, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2263113.

Han:2020:BEP

[Han20] Ming Han. E-Bayesian estimation and its E-posterior risk of the exponential distribution parameter based on complete and Type I censored samples. Communications in Statistics: Theory and Methods, 49(8):1858–1872, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1565837.

Han:2021:SEL

[Han21a] Ming Han. A study on the effect of the loss function on Bayesian estimation and posterior risk of binomial distribution. Communications in Statistics: Theory and Methods, 50(18):4386–4399, 2021. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2020.1719160.

Hanagal:2021:CPS

[Han21b] David D. Hanagal. Correlated positive stable frailty models. Communications in Statistics: Theory and Methods, 50(23): 5617–5633, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2020.1736305.

Han:2023:TLH

[Han23a] Ming Han. Take a look at the hierarchical Bayesian estimation of parameters from several different angles. Communications in Statistics: Theory and Methods, 52(21):7718–7730, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2022.2056752.

Hanagal:2023:CPS

[Han23b] David D. Hanagal. Compound Poisson shared frailty models based on additive hazards. Communications in Statistics: Theory and Methods, 52(17):6287–6309, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2027453.

Hanagal:2024:CCP

[Han24] David D. Hanagal. Correlated compound Poisson frailty models based on reversed hazard rate. Communications in Statistics: Theory and Methods, 53(4):1312–1330, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2098336.

Haq:2020:NEC

[Haq20] Abdul Haq. A nonparametric EWMA chart with auxiliary information for process mean. Communications in Statistics: Theory and Methods, 49(5):1232–1247, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1554140.

Haq:2022:AMC

[Haq22] Abdul Haq. Adaptive MEWMA charts for univariate and multivariate simple linear profiles. Communications in Statistics: Theory and Methods, 51(16):5383–5411, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1839100.

Haddad:2020:LAN

[HB20] Soraya Haddad and Karima Belaide. Local asymptotic normality for long-memory process with strong mixing noises.

Communications in Statistics: Theory and Methods, 49(12): 2817–2830, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2019.1584306.

Hazra:2024:OPS

[HBH24] Nil Kamal Hazra, Ghobad Barmalzan, and Ali Akbar Hosseinzadeh. Ordering properties of the second smallest and the second largest order statistics from a general semiparametric family of distributions. Communications in Statistics: Theory and Methods, 53(1):328–345, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2077964.

Hidayat:2021:RCE

[HBOR21] Rahmat Hidayat, I. Nyoman Budiantara, Bambang W. Otok, and Vita Ratnasari. The regression curve estimation by using mixed smoothing spline and kernel (MsS-K) model. Communications in Statistics: Theory and Methods, 50(17):3942– 3953, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2019.1710201.

Hsu:2022:SEM

[HC22] Szu-Yuan Hsu and Jeng-Tung Chiang. Suppression and enhancement in multiple linear regression: a viewpoint from the perspective of a semipartial correlation coefficient. Communications in Statistics: Theory and Methods, 51(7):2057–2072, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2020.1759094.

Hao:2024:GRE

[HCCD24] Bingguang Hao, Yuelin Che, Luyuan Chen, and Yong Deng. A generalized Rényi entropy to measure the uncertainty of a random permutation set. Communications in Statistics: Theory and Methods, 53(23):8543–8555, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2292973.

He:2021:NPI

[HCCM21] Ting He, Frank P. A. Coolen, and Tahani Coolen-Maturi. Nonparametric predictive inference for American option pricing based on the binomial tree model. Communications in Statistics: Theory and Methods, 50(20):4657–4684, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1764040.

Hu:2023:SDT

[HCCzW23] Ping Hu, Mengru Chen, Shu Chen, and Zhong zhi Wang. On strong deviation theorems concerning array of dependent random sequence. Communications in Statistics: Theory and Methods, 52(9):3098–3107, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1967396.

Hussain:2020:SRR

[HCH20] Zawar Hussain, Salman A. Cheema, and Ishtiaq Hussain. A stratified randomized response model for sensitive characteristics using non identical trials. Communications in Statistics: Theory and Methods, 49(1):99–115, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1530791.

Huang:2021:GWM

[HCL21] Chi-Jui Huang, Jen-Hsiang Chen, and Shin-Li Lu. Generally weighted moving average control charts using repetitive sampling. Communications in Statistics: Theory and Methods, 50(2):297–310, 2021. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2019.1634212.

Han:2022:QET

[HCL22] Qi Han, Zhihe Chen, and Ziqiang Lu. Quantum entropy in terms of local quantum Bernoulli noises and related properties. Communications in Statistics: Theory and Methods, 51(12):4210–4220, 2022. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2020.1812654.

Hsu:2022:AEG

[HCLW22] Yu-Sheng Hsu, Pei-Chun Chen, Ming-Yung Lee, and Cheng-Hsun Wu. Applications of an extended geometric Brownian motion degradation model. Communications in Statistics: Theory and Methods, 51(7):2139–2153, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1764039.

Hongyan:2020:FAN

[HCX20] Dui Hongyan, Zhang Chi, and Xu Xin. Failure analysis of network nodes and edges in scale-free networks. Communications in Statistics: Theory and Methods, 49(15):3635–3649, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2019.1703136.

Hu:2020:MAJ

[HCZ20] Guozhi Hu, Weihu Cheng, and Jie Zeng. Model averaging by jackknife criterion for varying-coefficient partially linear models. Communications in Statistics: Theory and Methods, 49(11):2671–2689, 2020. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2019.1580736.

He:2021:EMG

[HDZL21] Jie He, Xiaogang Duan, Shumei Zhang, and Hui Li. Estimation of marginal generalized linear model with subgroup auxiliary information. Communications in Statistics: Theory and Methods, 50(4):837–855, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1642490.

He:2023:OBA

[He23] Lei He. Objective Bayesian analysis of accelerated degradation models based on Wiener process with correlation. Communications in Statistics: Theory and Methods, 52(8):2666–2681, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2021.1957111.

Hu:2024:ESI

[HF24] Zhishui Hu and Qunqiang Feng. The enhanced strong invariance principle for the elephant random walk. Communications in Statistics: Theory and Methods, 53(3):834–847, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2092749.

Hu:2024:FRL

[HFGZ24] Feng Hu, Yanan Fu, Miaomiao Gao, and Zhaojun Zong. Further results on laws of large numbers for the array of random variables under sub-linear expectation. Communications in Statistics: Theory and Methods, 53(17):6076–6101, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2239400.

Hui:2021:GEM

[HFJ21] Jianan Hui, James M. Flegal, and Alicia Johnson. Geometric ergodicity of a more efficient conditional Metropolis– Hastings algorithm. Communications in Statistics: Theory and Methods, 50(19):4528–4547, 2021. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2020.1719418.

Hussain:2021:SNG

[HGAJ21] Sajid Hussain, Jigneshkumar Gondaliya, Rashid Ahmed, and Rida Jabeen. Some new generators to obtain efficient circular weakly balanced repeated measurements designs. Communications in Statistics: Theory and Methods, 50(15):3713–3730, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2019.1710200.

He:2024:SDR

[HGY24] Xinya He, Ailing Gu, and Haixiang Yao. Stochastic differential reinsurance and investment games with delay under VaR constraints. Communications in Statistics: Theory and Methods, 53(4):1479–1515, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2103149.

Hong:2023:NSS

[HHKC23] Hyunsook Hong, Seokyoung Hahn, Ho Kim, and Yunhee Choi. Nomogram for sample size calculation in assessing validity of a new method based on a regression line. Communications in Statistics: Theory and Methods, 52(16):5900–5909, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.2023182.

Han:2023:QME

[HHKL23] Qi Han, Yanan Han, Yaxin Kou, and Ziqiang Lu. Quantum mutual entropy in terms of local quantum Bernoulli noises. Communications in Statistics: Theory and Methods, 52(3): 515–522, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2021.1916532.

Hu:2022:BPI

[HHLG22] Jun-Ming Hu, Hong-Zhong Huang, Yan-Feng Li, and Hui-Ying Gao. Bayesian prior information fusion for power law process via evidence theory. Communications in Statistics: Theory and Methods, 51(14):4921–4939, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1828464.

Hossain:2021:RBM

[HHW21] Shahadut Hossain, Zahirul Hoque, and Jacek Wesolowski. Response-based multiple imputation method for minimizing the impact of covariate detection limit in logistic regression. Communications in Statistics: Theory and Methods, 50(2): 371–386, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2019.1635699.

He:2022:DSM

[HJ22] Xuyang He and Yuexiang Jiang. Data sharpening method in regression confidence band. Communications in Statistics: Theory and Methods, 51(5):1349–1366, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1760887.

Hamura:2020:BPD

[HK20] Yasuyuki Hamura and Tatsuya Kubokawa. Bayesian predictive distribution for a Poisson model with a parametric restriction. Communications in Statistics: Theory and Methods, 49(13):3257–3266, 2020. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2019.1586943.

Huang:2021:BCM

[HK21] A. Huang and A. S. I. Kim. Bayesian Conway–Maxwell– Poisson regression models for overdispersed and underdispersed counts. Communications in Statistics: Theory and Methods, 50(13):3094–3105, 2021. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2019.1682162.

Huang:2021:GBA

[HKL21] Longlong Huang, Karen Kopciuk, and Xuewen Lu. A group bridge approach for component selection in nonparametric accelerated failure time additive regression model. Communications in Statistics: Theory and Methods, 50(6):1477–1501, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2019.1651861.

Hanusz:2020:DCA

[HKNW20] Zofia Hanusz, Mirosław Krzyśko, Rafał Nadulski, and Lukasz Waszak. Discriminant coordinates analysis for multivariate functional data. Communications in Statistics: Theory and Methods, 49(18):4506–4519, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1602650.

Hsieh:2022:TST

[HL22a] Jin-Jian Hsieh and Jyun-Peng Li. Two-sample test based on empirical likelihood ratio under semi-competing risks data. Communications in Statistics: Theory and Methods, 51(10): 3301–3311, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2020.1793363.

Hu:2022:OPS

[HL22b] Xiaoxiao Hu and Xiaohu Li. On ordering parallel systems with two components having Marshall/Olkin-type lifetime distribution. Communications in Statistics: Theory and Methods, 51(15):5318–5329, 2022. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2020.1838544.

Han:2023:CFV

[HL23] Yuecai Han and Nan Li. Calibrating fractional Vasicek model. Communications in Statistics: Theory and Methods, 52(13): 4429–4443, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2021.1994609.

Holm:2024:CID

[HL24a] Sture Holm and Rolf Larsson. Confidence intervals for distributional positions. Communications in Statistics: Theory and Methods, 53(7):2645–2660, 2024. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2022.2148468.

Hu:2024:ELS

[HL24b] Yan-Ping Hu and Han-Ying Liang. Empirical likelihood in single-index partially functional linear model with missing observations. Communications in Statistics: Theory and Methods, 53(3):882–908, 2024. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2022.2094413.

Hu:2022:RSA

[HLPL22] Linmin Hu, Sijia Liu, Rui Peng, and Zhaocai Liu. Reliability and sensitivity analysis of a repairable k-out-of-n: G system with two failure modes and retrial feature. Communications in Statistics: Theory and Methods, 51(9):3043–3064, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2020.1788083.

He:2021:CMC

[HM21] Zongda He and Zhonglian Ma. A computational method to compare spectral densities of independent periodically correlated time series. Communications in Statistics: Theory and Methods, 50(8):1745–1755, 2021. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2019.1652758.

Haque:2022:AAI

[HM22] Shovanur Haque and Kerrie Mengersen. Assessing the accuracy of individual link with varying block sizes and cutoff values using MaCSim approach. Communications in Statistics: Theory and Methods, 51(18):6182–6196, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1857771.

Hashempour:2024:DCP

[HM24] Majid Hashempour and Morteza Mohammadi. On dynamic cumulative past inaccuracy measure based on extropy. Communications in Statistics: Theory and Methods, 53(4):1294– 1311, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2022.2098335.

Hutson:2020:LES

[HMM20] Alan D. Hutson, Terry L. Mashtare, Jr., and Govind S. Mudholkar. Log-epsilon-skew normal: a generalization of the log-normal distribution. Communications in Statistics: Theory and Methods, 49(17):4197–4215, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1595655.

Haslett:2023:PBF

[HMP23] Stephen J. Haslett, Augustyn Markiewicz, and Simo Puntanen. Properties of BLUEs in full versus small linear models. Communications in Statistics: Theory and Methods, 52(21): 7684–7698, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2022.2052899.

Hamza:2020:ANC

[HMZ20] Daoudi Hamza, Boubaker Mechab, and Chikr Elmezouar Zouaoui. Asymptotic normality of a conditional hazard function estimate in the single index for quasi-associated data. Communications in Statistics: Theory and Methods, 49(3): 513–530, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2018.1549248.

Hosseini:2020:CRL

[HN20] S. Mohammad Hosseini and Ahmad Nezakati. Convergence rates in the law of large numbers for END linear processes with random coefficients. Communications in Statistics: Theory and Methods, 49(1):88–98, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1530790.

Hyodo:2021:STM

[HN21] Masashi Hyodo and Takahiro Nishiyama. Simultaneous testing of the mean vector and covariance matrix among k populations for high-dimensional data. Communications in Statistics: Theory and Methods, 50(3):663–684, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1639751.

Hajrajabi:2023:OFE

[HN23] A. Hajrajabi and P. Nabati. Optimal filtering equations in state space model of the two factors mean reverting Ornstein-uhlenbech process. Communications in Statistics: Theory and Methods, 52(20):7532–7542, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2048309.

Hassan:2023:CCQ

[HNR^{+}23] Jamshaidul Hassan, Khadija Noreen, H. M. Kashif Rasheed, Mahmood ul Hassan, and Rashid Ahmed. Construction of circular quasi Rees neighbor designs which can be converted into minimal circular balanced and strongly balanced neighbor designs. Communications in Statistics: Theory and Methods, 52(16):5587–5605, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.2012686.

He:2023:ARC

[HO23] Zhulin He and Yuyuan Ouyang. An asymptotic result of conditional logistic regression estimator. Communications in Statistics: Theory and Methods, 52(13):4729–4740, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1999978.

Haile:2024:BAT

[HO24] Mulubrhan G. Haile and David J. Olive. Bootstrapping ARMA time series models after model selection. Communications in Statistics: Theory and Methods, 53(23):8255–8270, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2023.2280546.

Holm:2023:GLM

[Hol23] Sture Holm. Generalized linear models for ordered categorical data. Communications in Statistics: Theory and Methods, 52(3):670–683, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03610926.2021.1921210.

Homei:2021:SLC

[Hom21] Hazhir Homei. The stochastic linear combination of Dirichlet distributions. Communications in Statistics: Theory and Methods, 50(10):2354–2359, 2021. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2019.1664588.

Hong:2020:CCT

[Hon20] Liang Hong. Continuous counterexamples for three dependence notions. Communications in Statistics: Theory and Methods, 49(19):4853–4858, 2020. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2019.1609517.

Hanagal:2020:CIG

[HP20] David D. Hanagal and Arvind Pandey. Correlated inverse Gaussian frailty models for bivariate survival data. Communications in Statistics: Theory and Methods, 49(4):845–863, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2018.1549256.

Hunt:2024:SGA

[HPW24] Richard Hunt, Shelton Peiris, and Neville Weber. Seasonal generalized AR models. Communications in Statistics: Theory and Methods, 53(3):1065–1080, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2100422.

Henry:2023:REP

[HR23] Guillermo Henry and Daniela Rodriguezs. Robust estimators in partly linear regression models on Riemannian manifolds. Communications in Statistics: Theory and Methods, 52(14): 4835–4851, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2013.775302.

Hyodo:2020:APC

[HS20] Masashi Hyodo and Nobumichi Shutoh. Asymptotic power comparison of T^{2}-type test and likelihood ratio test for a mean vector based on two-step monotone missing data. Communications in Statistics: Theory and Methods, 49(17):4270–4287, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2019.1597122.

Hu:2021:CRL

[HS21] Ze-Chun Hu and Wei Sun. Convergence rates in the law of large numbers and new kinds of convergence of random variables. Communications in Statistics: Theory and Methods, 50(18):4308–4323, 2021. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2020.1716248.

Holmes:2022:MLN

[HS22] John B. Holmes and Matthew R. Schofield. Moments of the logit-normal distribution. Communications in Statistics: Theory and Methods, 51(3):610–623, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1752723.

Hapuhinna:2024:BME

[HS24a] Nelum S. S. M. Hapuhinna and Junfeng Shang. A bootstrap method for estimation in linear mixed models with heteroscedasticity. Communications in Statistics: Theory and Methods, 53(11):4012–4036, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2170181.

Haq:2024:ECD

[HS24b] Abdul Haq and Ehsana Anum Syed. Enhanced CUSUM and dual CUSUM mean charts. Communications in Statistics: Theory and Methods, 53(6):1906–1941, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2116285.

Hussain:2022:ESP

[HSC22] Zawar Hussain, Sidra Shakeel, and Salman A. Cheema. Estimation of stigmatized population total: a new additive quantitative randomized response model. Communications in Statistics: Theory and Methods, 51(24):8741–8753, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1906431.

Han:2021:EVS

[HSG21] Cun Han, Xiaofei Sun, and Wenliang Gao. Estimation and variable selection for a class of quantile regression models with multiple index. Communications in Statistics: Theory and Methods, 50(1):180–202, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1633353.

Shi:2024:SAP

[hSsXfX24] Jian hua Shi, Jian sen Xu, and Jin feng Xu. Strong asymptotic properties of kernel smoothing estimation for NA random variables with right censoring. Communications in Statistics: Theory and Methods, 53(12):4531–4541, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2184189.

Hemavathi:2022:TDR

[HSV^{+}22] M. Hemavathi, Shashi Shekhar, Eldho Varghese, Seema Jaggi, Bikas Sinha, and Nripes Kumar Mandal. Theoretical developments in response surface designs: an informative review and further thoughts. Communications in Statistics: Theory and Methods, 51(7):2009–2033, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1944213.

Han:2021:EAS

[HSWZ21] Miao Han, Xuefeng Song, Wei Wang, and Shengwu Zhou. Empirical analysis of SH50ETF and SH50ETF option prices under regime-switching jump-diffusion models. Communications in Statistics: Theory and Methods, 50(9):2170–2187, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1662045.

He:2023:OPR

[HT23] Lei He and Shuling Tian. Optimal post-retirement consumption and portfolio choices with idiosyncratic individual mortality force and awareness of mortality risk. Communications in Statistics: Theory and Methods, 52(7):2259–2275, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1949473.

Hao:2024:RSA

[HT24] Zuolu Hao and Yu Tang. Revised schematic array. Communications in Statistics: Theory and Methods, 53(22):8097–8108, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2023.2279911.

Hooshangifar:2022:ODL

[HTP22] Mina Hooshangifar, Hooshang Talebi, and Davood Poursina. D-optimal design for logistic model based on more precise approximation. Communications in Statistics: Theory and Methods, 51(7):1975–1992, 2022. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2021.1957482.

Hu:2020:WSL

[Hu20a] Cheng Hu. Weak and strong laws of large numbers for sub-linear expectation. Communications in Statistics: Theory and Methods, 49(2):430–440, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1543771.

Hu:2020:PRM

[Hu20b] Xiaomi Hu. A pseudo restricted MLE under multivariate order restrictions and its algorithm. Communications in Statistics: Theory and Methods, 49(1):169–177, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1535072.

Hu:2022:DSS

[Hu22] Jun Hu. A double-sequential sampling scheme. Communications in Statistics: Theory and Methods, 51(18):6319–6333, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2020.1860225.

Hu:2024:MSO

[Hu24] Xiaomi Hu. Matrix spaces and ordinary least square estimators in linear models for random matrices. Communications in Statistics: Theory and Methods, 53(21):7723–7732, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2272004.

Huang:2020:ETN

[Hua20] Yue Huang. Estimation and testing of nonparametric hidden Markov model with application in stock market. Communications in Statistics: Theory and Methods, 49(24):5917–5929, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2019.1622731.

Huang:2023:DSH

[Hua23] Rong Huang. Detecting sparse heterogeneous mixtures in a two-sample problem. Communications in Statistics: Theory and Methods, 52(5):1333–1347, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1926510.

Hepworth:2020:EPG

[HW20a] Graham Hepworth and Stephen D. Walter. Estimation of proportions by group testing with retesting of positive groups. Communications in Statistics: Theory and Methods, 49(22): 5587–5597, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2019.1620280.

Hu:2020:ATE

[HW20b] Sigui Hu and Honglei Wang. Assurance test and its equivalent truncated sequential test. Communications in Statistics: Theory and Methods, 49(11):2623–2633, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1576897.

He:2022:EST

[HW22a] Jingmin He and Fangling Wu. Exact solutions of the two-side exit time problems for the Vasicek model. Communications in Statistics: Theory and Methods, 51(24):8625–8633, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1901921.

He:2022:SUR

[HW22b] Tian He and Lichun Wang. On seemingly unrelated regressions with uniform correlation error. Communications in Statistics: Theory and Methods, 51(16):5714–5727, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1845736.

Hu:2022:SMD

[HW22c] Xiaomi Hu and Yufei Wang. Stochastic monotonicity of a distribution family associated with matrix projections and its application. Communications in Statistics: Theory and Methods, 51(11):3694–3704, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1801734.

Huang:2022:SLL

[HW22d] Weihuan Huang and Panyu Wu. Strong law of large numbers under moment restrictions in sublinear expectation spaces. Communications in Statistics: Theory and Methods, 51(24): 8671–8683, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2021.1903504.

Hu:2023:RCD

[HW23a] Duni Hu and Hailong Wang. Reinsurance contract design with heterogeneous beliefs and learning. Communications in Statistics: Theory and Methods, 52(14):5026–5047, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.2001657.

Hu:2023:CCT

[HW23b] Rong Hu and Qunying Wu. Complete convergence theorems for arrays of row-wise extended negatively dependent random variables under sub-linear expectations. Communications in Statistics: Theory and Methods, 52(21):7669–7683, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2051050.

Han:2024:OPE

[HW24] Miao Han and Wei Wang. Option pricing with exchange rate risk under regime-switching multi-scale jumpdiffusion models. Communications in Statistics: Theory and Methods, 53(7):2329–2354, 2024. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2022.2129992.

Hu:2022:EPG

[HWH22] Yusang Hu, S. D. Walter, and Graham Hepworth. Estimating proportions by group retesting with unequal group sizes at each stage. Communications in Statistics: Theory and Methods, 51(24):8532–8552, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1900253.

Hyodo:2023:NTD

[HWNN23] Masashi Hyodo, Hiroki Watanabe, Shigekazu Nakagawa, and Tomoyuki Nakagawa. Normalizing transformation of Dempster type statistic in high-dimensional settings. Communications in Statistics: Theory and Methods, 52(22):8096–8113, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2022.2056749.

Huang:2024:PPC

[HWW24] Wenyang Huang, Huiwen Wang, and Shanshan Wang. A pseudo principal component analysis method for multidimensional open-high-low-close data in candlestick chart. Communications in Statistics: Theory and Methods, 53(10): 3472–3498, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2022.2155787.

Huang:2020:QRB

[HWWY20] Zongyuan Huang, Haiyang Wang, Zhen Wu, and Zhiyong Yu. Quadratic reflected BSDEs and related obstacle problems for PDEs. Communications in Statistics: Theory and Methods, 49(3):567–589, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1543778.

Hu:2023:SIV

[HWZX23] Yuping Hu, Yilun Wang, Liying Zhang, and Liugen Xue. Statistical inference of varying-coefficient partial functional spatial autoregressive model. Communications in Statistics: Theory and Methods, 52(14):4960–4980, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1992438.

Hu:2022:ERI

[HX22] Shuwen Hu and Jianwen Xu. An efficient and robust inference method based on empirical likelihood in longitudinal data analysis. Communications in Statistics: Theory and Methods, 51(4):994–1010, 2022. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2020.1757110.

Hu:2023:RBE

[HX23] Guikai Hu and Xinhai Xiao. Robust Bayesian estimator in a normal model with uncertain hierarchical priors. Communications in Statistics: Theory and Methods, 52(3):567–582, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1917615.

Hou:2022:PEU

[HXCW22] Gege Hou, Ancha Xu, Fengjing Cai, and You-Gan Wang. Parameter estimation for univariate Skew-Normal distribution based on the modified empirical characteristic function. Communications in Statistics: Theory and Methods, 51(22):7897– 7910, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2021.1883655.

Hou:2023:DWL

[HY23] Chia-Ding Hou and Ti-Sung Yang. Distribution of weighted Lancaster’s statistic for combining independent or dependent P -values, with applications to human genetic studies. Communications in Statistics: Theory and Methods, 52(20):7442– 7454, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2022.2046088.

Huang:2024:CMC

[HYWZ24] Haiwu Huang, Yuan Yuan, Wei Wang, and Hongguo Zeng. On complete moment convergence for the maximal weighted sums of NSD random variables. Communications in Statistics: Theory and Methods, 53(10):3779–3796, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2163368.

He:2024:REE

[HYZ24] Ping He, Yiping Yang, and Peixin Zhao. Robust estimation with exponential squared loss for partially linear panel data model with fixed effects. Communications in Statistics: Theory and Methods, 53(15):5638–5656, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2226274.

Hosseini:2021:CBB

[HZ21] Reyhaneh Hosseini and Mahmoud Zarepour. A consistent Bayesian bootstrap for chi-squared goodness-of-fit test using a Dirichlet prior. Communications in Statistics: Theory and Methods, 50(8):1756–1773, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1653919.

Hu:2022:PLL

[HZ22] Hongchang Hu and Zhen Zeng. Penalized Lq-likelihood estimators and variable selection in linear regression models. Communications in Statistics: Theory and Methods, 51(17): 5957–5970, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2020.1850794.

Harfouche:2023:MBC

[HZZA23] Lynda Harfouche, Yasmina Ziane, Nabil Zougab, and Smail Adjabi. Multiplicative bias correction for inverse gamma and beta prime kernel density estimators. Communications in Statistics: Theory and Methods, 52(17):6088–6102, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.2025248.

Imed:2021:EPB

[IAM21] Bouzida Imed, Masmoudi Afif, and Zitouni Mouna. Estimation parameters for the Binomial q-distribution. Communications in Statistics: Theory and Methods, 50(21):5101–5113, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2020.1725825.

Ibrahim:2020:RMS

[IHDR20] Firas Ibrahim, Ali Hajj Hassan, Jacques Demongeot, and Mustapha Rachdi. Regression model for surrogate data in high dimensional statistics. Communications in Statistics: Theory and Methods, 49(13):3206–3227, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1586940.

Koike:2021:ABI

[iK21a] Ken ichi Koike. Attainments of the Bayesian information bounds. Communications in Statistics: Theory and Methods, 50(11):2696–2709, 2021. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2019.1676445.

Ito:2021:EBL

[IK21b] Tsubasa Ito and Tatsuya Kubokawa. Empirical best linear unbiased predictors in multivariate nested-error regression models. Communications in Statistics: Theory and Methods, 50(10):2224–2249, 2021. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2019.1662048.

Koike:2022:ACS

[iK22] Ken ichi Koike. Asymptotic comparison of some Bayesian information bounds. Communications in Statistics: Theory and Methods, 51(3):599–609, 2022. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2020.1752722.

Ivanov:2021:LDC

[IKM21] Alexander Ivanov, Yuriy Kozachenko, and Kateryna Moskvychova. Large deviations of the correlogram estimator of the random noise covariance function in the nonlinear regression model. Communications in Statistics: Theory and Methods, 50(18):4236–4254, 2021. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2020.1713369.

Ianevych:2022:TCH

[IKT22] Tetiana O. Ianevych, Yuriy V. Kozachenko, and Viktor B. Troshki. On test for checking hypothesis on expectation and covariance function of stochastic process. Communications in Statistics: Theory and Methods, 51(2):356–367, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1749280.

Izadi:2021:TEA

[IM21] Muhyiddin Izadi and Sirous Fathi Manesh. Testing exponentiality against a trend change in mean time to failure in age replacement. Communications in Statistics: Theory and Methods, 50(14):3358–3370, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1702693.

Imada:2020:MCP

[Ima20] Tsunehisa Imada. Multiple comparison procedures for finding non-maximum normal means. Communications in Statistics: Theory and Methods, 49(16):4073–4090, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1595652.

Ikeda:2021:LSE

[INKS21] Yuki Ikeda, Ryumei Nakada, Tatsuya Kubokawa, and Muni S. Srivastava. Linear shrinkage estimation of the variance of a distribution with unknown mean. Communications in Statistics: Theory and Methods, 50(9):2039–2047, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1657457.

Ikeda:2023:LSE

[INKS23] Yuki Ikeda, Ryumei Nakada, Tatsuya Kubokawa, and Muni S. Srivastava. Linear shrinkage estimation of highdimensional means. Communications in Statistics: Theory and Methods, 52(13):4444–4460, 2023. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2021.1994610.

Inoue:2024:DRS

[Ino24] Kiyoshi Inoue. Distributions of runs and scans in multistate Markov exchangeable sequences. Communications in Statistics: Theory and Methods, 53(8):3040–3061, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2150061.

Ishii:2020:CUS

[Ish20] Aki Ishii. A classifier under the strongly spiked eigenvalue model in high-dimension, low-sample-size context. Communications in Statistics: Theory and Methods, 49(7):1561–1577, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2018.1528365.

Ivanoff:2022:ICP

[IW22] B. Gail Ivanoff and Neville C. Weber. Incorporating a changepoint estimator when bootstrapping the empirical distribution of a stationary process. Communications in Statistics: Theory and Methods, 51(9):2765–2782, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1780260.

Ishihara:2023:TPC

[IY23] Takuma Ishihara and Kouji Yamamoto. A testing procedure in clinical trials with multiple binary endpoints. Communications in Statistics: Theory and Methods, 52(2):273–282, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1912766.

Jodra:2022:IMD

[JA22] Pedro Jodrá and Mohd Arshad. An intermediate muth distribution with increasing failure rate. Communications in Statistics: Theory and Methods, 51(23):8310–8327, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1892133.

Jelloul:2023:ASR

[JAE23] Allal Jelloul, Nabil Azouagh, and Said El Melhaoui. Aligned signed-rank tests of a linear autoregressive model against an exponential autoregressive one. Communications in Statistics: Theory and Methods, 52(22):7944–7965, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2052898.

Jafaraghaie:2022:PFP

[Jaf22] Razieh Jafaraghaie. Prediction of finite population parameters using parametric model under some loss functions. Communications in Statistics: Theory and Methods, 51(4):863–882, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2020.1801736.

Jamali:2021:CMS

[JAJ21] Dariush Jamali, Mehdi Amiri, and Ahad Jamalizadeh. Comparison of the multivariate skew-normal random vectors based on the integral stochastic ordering. Communications in Statistics: Theory and Methods, 50(22):5215–5227, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1740934.

Janicki:2020:PBR

[Jan20] Ryan Janicki. Properties of the beta regression model for small area estimation of proportions and application to estimation of poverty rates. Communications in Statistics: Theory and Methods, 49(9):2264–2284, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1570266.

Jasinski:2020:ABR

[Jas20] Krzysztof Jasiński. Asymptotic behavior of the ratio of weak k-th records. Communications in Statistics: Theory and Methods, 49(1):16–26, 2020. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2018.1535074.

Jasinski:2021:APS

[Jas21] Krzysztof Jasiński. Asymptotic properties of the spacings of k-th records from discrete distributions. Communications in Statistics: Theory and Methods, 50(24):5679–5691, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1738488.

Jeyadurga:2022:ODM

[JB22a] P. Jeyadurga and S. Balamurali. Optimal designing of multiple deferred (dependent) state repetitive group sampling plan for variables inspection. Communications in Statistics: Theory and Methods, 51(13):4417–4433, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1814815.

Jornsatian:2022:ZOI

[JB22b] Chanakarn Jornsatian and Winai Bodhisuwan. Zero-one inflated negative binomial — beta exponential distribution for count data with many zeros and ones. Communications in Statistics: Theory and Methods, 51(24):8517–8531, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1898642.

Joshi:2022:IAS

[JB22c] Neeraj Joshi and Sudeep R. Bapat. On improved accelerated sequential estimation of the mean of an inverse Gaussian distribution. Communications in Statistics: Theory and Methods, 51(17):6127–6143, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1854304.

Joshi:2023:ASO

[JBJ23] Chaitanya Joshi, Paul T. Brown, and Stephen Joe. On approximating the shape of one-dimensional posterior marginals using the low discrepancy points. Communications in Statistics: Theory and Methods, 52(16):5568–5586, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.2012577.

Jabeen:2024:MCG

[JBR^{+}24] Rida Jabeen, Zahid Bashir, H. M. Kashif Rasheed, Jigneshkumar Gondaliya, and Rashid Ahmed. Minimal circular generalized strongly balanced and their conversion into circular balanced and strongly balanced repeated measurements designs. Communications in Statistics: Theory and Methods, 53(4): 1440–1458, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2022.2101664.

Jhong:2022:ARS

[JBSK22] Jae-Hwan Jhong, Kwan-Young Bak, Jae-Kyung Shin, and Ja-Yong Koo. Additive regression splines with total variation and non negative garrote penalties. Communications in Statistics: Theory and Methods, 51(22):7713–7736, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1879860.

Jiang:2020:CLA

[JC20] Shu Jiang and Richard J. Cook. Composite likelihood for aggregate data from clustered multistate processes under intermittent observation. Communications in Statistics: Theory and Methods, 49(12):2913–2930, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1584310.

Jerbi:2022:HMB

[JC22] Yacin Jerbi and Samira Chaabene. Hybrid method based on neural networks and Monte Carlo simulation in view of a tradeoff between accuracy and computational time. Communications in Statistics: Theory and Methods, 51(4):909–918, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2020.1818782.

Jiang:2022:SCB

[JCY22] Jiakun Jiang, Li Cai, and Lijian Yang. Simultaneous confidence band for the difference of regression functions of two samples. Communications in Statistics: Theory and Methods, 51(11):3556–3572, 2022. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2020.1800039.

Javier:2023:UTP

[JG23] Marvin Javier and Kaushik Ghosh. Uniform k-tuple partially rank-ordered set sampling. Communications in Statistics: Theory and Methods, 52(7):2338–2355, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1952266.

Jiang:2024:SIP

[JH24] Zhiqiang Jiang and Zhensheng Huang. Single-index partially functional linear quantile regression. Communications in Statistics: Theory and Methods, 53(5):1838–1850, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2116282.

Ji:2020:PPE

[JJ20] Qinghua Ji and Zheng Ji. Penalized proportion estimation for non parametric mixture of regressions. Communications in Statistics: Theory and Methods, 49(7):1537–1560, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1473614.

Jain:2020:TCC

[JKD20] Kanchan Jain, Harmanpreet Singh Kapoor, and Isha Dewan. Test for comparing complete expectations of life of two groups. Communications in Statistics: Theory and Methods, 49(8): 1960–1974, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2019.1568478.

Juric:2020:AMW

[JKP20] Visnja Jurić, Tomasz J. Kozubowski, and Mihael Perman. An asymmetric multivariate Weibull distribution. Communications in Statistics: Theory and Methods, 49(18):4394–4412, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2019.1599949.

Jiang:2022:TBE

[JL22a] Yingying Jiang and Baokun Li. Tail behavior and extremes of the beta-normal distribution. Communications in Statistics: Theory and Methods, 51(14):4991–5002, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1830294.

Jo:2022:MTE

[JL22b] Minyoung Jo and Sangyeol Lee. Mean targeting estimation for integer-valued time series with application to change point test. Communications in Statistics: Theory and Methods, 51(16):5549–5565, 2022. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2020.1843054.

Ji:2024:SOK

[JL24a] Shaolin Ji and Haodong Liu. Solvability of one kind of forward-backward stochastic difference equations. Communications in Statistics: Theory and Methods, 53(16):5853–5870, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2023.2235444.

Junmei:2024:EPG

[JL24b] Zhou Junmei and Li Liqin. Estimating parameters of the gamma distribution easily and efficiently. Communications in Statistics: Theory and Methods, 53(17):6197–6205, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2241097.

Jin:2021:ADT

[JLG21] Linghui Jin, Yanyan Liu, and Lisha Guo. Asymptotic distribution theory on pseudo semiparametric maximum likelihood estimator with covariates missing not at random. Communications in Statistics: Theory and Methods, 50(12):2918–2929, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2019.1678639.

Jha:2024:IEH

[JMK24] Brijesh Kumar Jha, Ajaya Kumar Mahapatra, and Suchandan Kayal. Improved estimators of hazard rate from a selected exponential population. Communications in Statistics: Theory and Methods, 53(13):4927–4943, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2198624.

Jose:2022:ESS

[JMM22] Joby K. Jose, Drisya M., and Manoharan M. Estimation of stress-strength reliability using discrete phase type distribution. Communications in Statistics: Theory and Methods, 51(2):368–386, 2022. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2020.1749663.

Johnston:2020:GOP

[JMQ20] Carla Johnston, James McDonald, and Kramer Quist. A generalized ordered Probit model. Communications in Statistics: Theory and Methods, 49(7):1712–1729, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1565780.

Jose:2024:PTS

[JMSG24] Joby K. Jose, Drisya M., Kulathinal Sangita, and Sebastian George. Phase-type stress-strength reliability models under progressive type-II right censoring. Communications in Statistics: Theory and Methods, 53(23):8498–8524, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2292968.

Jing:2022:AEF

[JPJB22] Haojie Jing, Jiangyan Peng, Zhiquan Jiang, and Qian Bao. Asymptotic estimates for finite-time ruin probability in a discrete-time risk model with dependence structures and CMC simulations. Communications in Statistics: Theory and Methods, 51(11):3761–3786, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1801740.

Jiang:2023:UAR

[JPZ23] Zhiquan Jiang, Jiangyan Peng, and Lei Zou. Uniform asymptotics for ruin probabilities of a time-dependent renewal risk model with dependence structures and stochastic returns. Communications in Statistics: Theory and Methods, 52(13): 4553–4577, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2021.1995754.

Josey:2023:PBL

[JRB^{+}23] Kevin P. Josey, Brandy M. Ringham, Anna E. Barón, Margaret Schenkman, Katherine A. Sauder, Keith E. Muller, Dana Dabelea, and Deborah H. Glueck. Power for balanced linear mixed models with complex missing data processes. Communications in Statistics: Theory and Methods, 52(1):46– 64, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2021.1909732.

Janamiri:2023:RGE

[JRCB23] Fariba Janamiri, Abdolrahman Rasekh, Alireza Chaji, and Babak Babadi. Ridge-GME estimation in linear mixed models. Communications in Statistics: Theory and Methods, 52(15):5115–5132, 2023. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2021.2003402.

Jain:2024:PAM

[JRD24] Vidyottama Jain, Raina Raj, and S. Dharmaraja. Performability analysis of a MMAP[2]/PH [2]/S model with PH retrial times. Communications in Statistics: Theory and Methods, 53(8):2868–2887, 2024. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2022.2150053.

Joukar:2021:PEE

[JRM21] Abolfazl Joukar, Masoumeh Ramezani, and S. M. T. K. Mir-Mostafaee. Parameter estimation for the exponential-Poisson distribution based on ranked set samples. Communications in Statistics: Theory and Methods, 50(3):560–581, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1639745.

Jayalekshmi:2024:BLT

[JRN24] S. Jayalekshmi, G. Rajesh, and N. Unnikrishnan Nair. Bivariate Laplace transform of residual lives and their properties. Communications in Statistics: Theory and Methods, 53(2): 505–523, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2022.2085874.

Jayakumar:2021:EIP

[JS21] K. Jayakumar and K. K. Sankaran. Exponential intervened Poisson distribution. Communications in Statistics: Theory and Methods, 50(13):3063–3093, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1682161.

Jin:2023:EGS

[JS23a] Jin Jin and Liuquan Sun. Estimation in a general semiparametric hazards regression model with missing covariates. Communications in Statistics: Theory and Methods, 52(9): 3070–3097, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2021.1967395.

Jose:2023:OAC

[JS23b] Jitto Jose and E. I. Abdul Sathar. An ordered approach on cumulative extropy measures for information analysis. Communications in Statistics: Theory and Methods, 52(5):1512–1532, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2021.1928706.

Jafari:2022:SLL

[JSB22] Mehdi Jafari, Hamid Reza Nili Sani, and Abolghasem Bozorgnia. Strong laws of large numbers for WAPND Banach-valued random elements. Communications in Statistics: Theory and Methods, 51(9):3008–3017, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1786703.

Jena:2024:EPS

[JT24] Pravash Jena and Manas Ranjan Tripathy. Estimating powers of the scale parameters under order restriction for two shifted exponential populations with a common location. Communications in Statistics: Theory and Methods, 53(18):6614–6648, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2023.2248533.

Jamshidi:2022:PUM

[JTS22] Babak Jamshidi, Parisa Torkaman, and Azad Sheikhi. Pólya urn models in the presence of additional information. Communications in Statistics: Theory and Methods, 51(21):7433– 7451, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2021.1872644.

Jung:2020:ORP

[Jun20] Yoonsuh Jung. Optimal regression parameter-specific shrinkage by plug-in estimation. Communications in Statistics: Theory and Methods, 49(18):4490–4505, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1602649.

John:2024:FTR

[JV24] Majnu John and Sujit Vettam. A few theoretical results for Laplace and arctan penalized ordinary least squares linear regression estimators. Communications in Statistics: Theory and Methods, 53(13):4819–4840, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2195033.

Jia:2024:CCC

[JW24a] Chengcheng Jia and Qunying Wu. Complete convergence and complete integral convergence for randomly weighted sums under the sublinear expectations. Communications in Statistics: Theory and Methods, 53(13):4790–4804, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2195029.

Jia:2024:CCW

[JW24b] Chengcheng Jia and Qunying Wu. Complete convergence for weighted sums of widely negative dependent random variables under the sub-linear expectations. Communications in Statistics: Theory and Methods, 53(14):5018–5040, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2203283.

Jiang:2022:LAB

[JWW22] Jie Jiang, Lichun Wang, and Liqun Wang. Linear approximate Bayes estimator for regression parameter with an inequality constraint. Communications in Statistics: Theory and Methods, 51(6):1531–1548, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1890125.

Jiang:2024:ABE

[JWW24] Jie Jiang, Lichun Wang, and Liqun Wang. Approximate Bayesian estimator for the parameter vector in linear models with multivariate t distribution errors. Communications in Statistics: Theory and Methods, 53(7):2516–2534, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2138438.

Wang:2024:CAR

[jWWDjW24] Zi jian Wang, Yi Wu, Yue Du, and Xue jun Wang. L_{r} convergence for arrays of rowwise m-extended negatively dependent random variables. Communications in Statistics: Theory and Methods, 53(20):7370–7383, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2263600.

Jiang:2024:NRS

[JWWW24] Zhe Jiang, Yan Wu, Min Wang, and Liucang Wu. A novel residual subsampling method for skew-normal mode regression model with massive data. Communications in Statistics: Theory and Methods, 53(16):5972–5988, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2238860.

Jiang:2024:CML

[JZ24] Haosheng Jiang and Chongqi Zhang. Construction of mixed-level fractional factorial split-plot designs with combined minimum aberration. Communications in Statistics: Theory and Methods, 53(21):7560–7581, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2269442.

Jiang:2024:ODS

[JZC24] Haosheng Jiang, Chongqi Zhang, and Jiali Chen. K-optimal designs for the second-order Scheffé polynomial model. Communications in Statistics: Theory and Methods, 53(22):8127– 8139, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2023.2279914.

Jurkova:2022:UEC

[JZKH22] Veronika Jurková, Ivan Zezula, Daniel Klein, and Ondrej Hutník. Unbiased estimator of correlation coefficient. Communications in Statistics: Theory and Methods, 51(1):95–115, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2020.1743314.

Jiang:2020:TEG

[JZWT20] Qing Jiang, Xun Zhang, Min Wu, and Xingwei Tong. Testing economic “genetic pleiotropy” for Box–Cox linear model. Communications in Statistics: Theory and Methods, 49(19): 4804–4818, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2019.1609036.

Kunjunni:2020:C

[KA20] Sajana O. Kunjunni and Sajesh T. Abraham. S_{n} covariance. Communications in Statistics: Theory and Methods, 49(24): 6133–6138, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2019.1628275.

Kumar:2022:BIM

[KA22] Jitendra Kumar and Varun Agiwal. Bayesian inference for merged panel autoregressive model. Communications in Statistics: Theory and Methods, 51(18):6197–6217, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1858101.

Karamikabir:2021:NEG

[KAAH21] Hamid Karamikabir, Mahmoud Afshari, Morad Alizadeh, and G. G. Hamedani. A new extended generalized Gompertz distribution with statistical properties and simulations. Communications in Statistics: Theory and Methods, 50(2):251–279, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2019.1634209.

Kifayat:2022:MPC

[KAC22] Tanveer Kifayat, Muhammad Aslam, and Salman Arif Cheema. The Maxwell paired comparison model under Bayesian paradigm using informative priors. Communications in Statistics: Theory and Methods, 51(2):301–312, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1748198.

Kheyri:2022:ECR

[KAJ^{+}22] A. Kheyri, M. Amini, H. Jabbari, A. Bozorgnia, and A. Volodin. Exponential convergence rates for the kernel bivariate distribution function estimator under NSD assumption with application to hydrology data. Communications in Statistics: Theory and Methods, 51(12):4042–4054, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1808900.

Kang:2020:TSF

[Kan20] Joonsung Kang. Two-stage false discovery rate in microarray studies. Communications in Statistics: Theory and Methods, 49(4):894–908, 2020. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2018.1554122.

Krieger:2024:DCR

[KASK24] Abba M. Krieger, David Azriel, Michael Sklar, and Adam Kapelner. Design choices in randomization tests that affect power. Communications in Statistics: Theory and Methods, 53(9):3276–3291, 2024. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2022.2152286.

Kharazmi:2022:IGF

[KB22a] Omid Kharazmi and Narayanaswamy Balakrishnan. Information generating function for order statistics and mixed reliability systems. Communications in Statistics: Theory and Methods, 51(22):7846–7855, 2022. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2021.1881123.

Kumar:2022:SCR

[KB22b] Tanuj Kumar and Rakesh Kumar Bajaj. On some characterization results of R̃-norm dynamic survival entropies. Communications in Statistics: Theory and Methods, 51(22):7963– 7979, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2021.1884880.

Kharazmi:2023:CRC

[KB23a] Omid Kharazmi and Narayanaswamy Balakrishnan. Cumulative and relative cumulative residual information generating measures and associated properties. Communications in Statistics: Theory and Methods, 52(15):5260–5273, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.2005100.

Kimouche:2023:SAG

[KB23b] Karima Kimouche and Abdelouahab Bibi. Spectral analysis for GARCH processes through a bilinear representation. Communications in Statistics: Theory and Methods, 52(19):6840– 6856, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2022.2034019.

Kamislik:2024:MBA

[KBKK24] Asli Bektas Kamislik, Feyrouz Baghezze, Tulay Kesemen, and Tahir Khaniyev. Moment-based approximations for stochastic control model of type (s, S). Communications in Statistics: Theory and Methods, 53(21):7505–7516, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2268765.

Kim:2023:RSO

[KBMD23] Jinseog Kim, Gaurab Bhattacharyya, Sabyasachi Mukherjee, and Rabindra Nath Das. Robust second-order rotatable invariably designs for some lifetime distributions. Communications in Statistics: Theory and Methods, 52(9):2819–2835, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2021.1961001.

Kalita:2020:RSQ

[KC20] Chandi Ram Kalita and Gautam Choudhury. An M^{X}/(\frac{G_{1}}{G_{2}} )/1 repeated service queue with N-policy and setup time subject to server’s breakdown and delayed repair. Communications in Statistics: Theory and Methods, 49(23):5709–5737, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1620783.

Kayan:2021:SCC

[KÇ21a] Emine Kayan and Rifat Çolak. df-statistical convergence in connection with a modulus in metric spaces. Communications in Statistics: Theory and Methods, 50(10):2270–2280, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1663212.

Kundu:2021:OPL

[KC21b] Amarjit Kundu and Shovan Chowdhury. Ordering properties of the largest order statistics from Kumaraswamy-G models under random shocks. Communications in Statistics: Theory and Methods, 50(6):1502–1514, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1651862.

Kou:2023:MWD

[KC23] Junke Kou and Kaili Cui. Multivariate wavelet density estimation for strong mixing stratified size-biased sample. Communications in Statistics: Theory and Methods, 52(6):1888–1904, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2021.1941111.

Kundu:2023:OPS

[KCB23] Amarjit Kundu, Shovan Chowdhury, and Narayanaswamy Balakrishnan. Ordering properties of the smallest and largest lifetimes in Gompertz–Makeham model. Communications in Statistics: Theory and Methods, 52(3):643–669, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1919898.

Kou:2021:MWE

[KCG21] Junke Kou, Jia Chen, and Huijun Guo. MISE of wavelet estimators for regression derivatives with biased strong mixing data. Communications in Statistics: Theory and Methods, 50(14):3436–3452, 2021. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2019.1704007.

Koc:2024:JKL

[KD24] Tuba Koç and Emre Dünder. Jackknife Kibria–Lukman estimator for the beta regression model. Communications in Statistics: Theory and Methods, 53(21):7789–7805, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2273206.

Kang:2023:MMI

[KdXYC23] Qingqing Kang, Guo dong Xing, Shanchao Yang, and Zhiyong Chen. Maximal moment inequalities for partial sums of ρ-mixing random variables with application to conditional value-at-risk estimator. Communications in Statistics: Theory and Methods, 52(5):1457–1471, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1928203.

Khalili:2020:EFM

[KE20] Atefeh Khalili and Farzad Eskandari. Estimating finite mixture of continuous trees using penalized mutual information. Communications in Statistics: Theory and Methods, 49(20): 4974–4987, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2019.1609519.

Khademnoe:2024:RSR

[KE24] Omid Khademnoe and Hadi Emami. Restricted Stein-rule estimation in ultrastructural linear measurement error models. Communications in Statistics: Theory and Methods, 53(7): 2459–2476, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2022.2135380.

Kelkinnama:2022:SCC

[Kel22] Maryam Kelkinnama. On stochastic comparisons of coherent systems with two different types of components. Communications in Statistics: Theory and Methods, 51(10):3253–3268, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2020.1792496.

Kerns:2020:BPI

[Ker20] Lucy Kerns. Benchmark profile and inferences for jointexposure quantal data in quantitative risk assessment. Communications in Statistics: Theory and Methods, 49(11):2713– 2727, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2019.1580740.

Kamal:2023:WLR

[KERAE23] Kholoud S. Kamal, Abd El-Raheem M. Abd El-Raheem, and Ehab F. Abd-Elfattah. Weighted log-rank tests for lefttruncated data: Saddlepoint p-values and confidence intervals. Communications in Statistics: Theory and Methods, 52(12): 4103–4113, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2021.1986534.

Ketzaki:2023:MBC

[KF23] Eleni Ketzaki and Nikolaos Farmakis. A matrix based computational method of the Gini index. Communications in Statistics: Theory and Methods, 52(17):5923–5941, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.2024233.

Kudraszow:2024:RAR

[KF24] Nadia L. Kudraszow and M. Victoria Fasano. Robust approaches to redundancy analysis. Communications in Statistics: Theory and Methods, 53(2):607–626, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2087882.

Kaur:2020:CIS

[KG20] Parneet Kaur and Davinder Kumar Garg. Construction of incomplete Sudoku square and partially balanced incomplete block designs. Communications in Statistics: Theory and Methods, 49(6):1462–1474, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1563177.

Kumar:2021:APM

[KG21] Nitin Kumar and U. C. Gupta. Analysis of a population model with batch Markovian arrivals influenced by Markov arrival geometric catastrophes. Communications in Statistics: Theory and Methods, 50(13):3137–3158, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1682166.

Krishna:2022:JTI

[KG22] Hare Krishna and Rajni Goel. Jointly type-II censored Lindley distributions. Communications in Statistics: Theory and Methods, 51(1):135–149, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1743316.

Kumar:2023:ABM

[KG23] Nitin Kumar and U. C. Gupta. Analysis of BMAP/MSP/1 queue with MAP generated negative customers and disasters. Communications in Statistics: Theory and Methods, 52(12): 4283–4309, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2021.1990953.

KazemiNia:2021:DBG

[KGG21] AliAkbar KazemiNia, Bahram Sadeghpour Gildeh, and Zainab Abbasi Ganji. The design of Bayesian generalized likelihood ratio control chart for monitoring the normal process mean. Communications in Statistics: Theory and Methods, 50(6):1400–1415, 2021. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2019.1651338.

Krishnankutty:2020:NPE

[KGS20] Baby Alpettiyil Krishnankutty, Rajesh Ganapathy, and Paduthol Godan Sankaran. Non-parametric estimation of copula based mutual information. Communications in Statistics: Theory and Methods, 49(6):1513–1527, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1563180.

Karavarsamis:2020:TSA

[KH20a] Natalie Karavarsamis and Richard M. Huggins. Two-stage approaches to the analysis of occupancy data i: the homogeneous case (analysis of occupancy data). Communications in Statistics: Theory and Methods, 49(19):4751–4761, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1607385.

Kim:2020:SEA

[KH20b] Tae Yoon Kim and Sun Young Hwang. Slow-explosive AR(1) processes converging to random walk. Communications in Statistics: Theory and Methods, 49(9):2094–2109, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1568486.

Kumar:2020:SAG

[KH20c] C. Satheesh Kumar and S. Harisankar. On some aspects of a general class of Yule distribution and its applications. Communications in Statistics: Theory and Methods, 49(12):2887– 2897, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2019.1584308.

Kurata:2020:CRM

[KH20d] Sumito Kurata and Etsuo Hamada. On the consistency and the robustness in model selection criteria. Communications in Statistics: Theory and Methods, 49(21):5175–5195, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1615093.

Killeen:2021:BTL

[KH21] Timothy J. Killeen and Thomas P. Hettmansperger. A bivariate test for location based on simplicial depth. Communications in Statistics: Theory and Methods, 50(17):3954–3971, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2019.1710205.

Khattree:2020:CCB

[Kha20] Ravindra Khattree. On construction of constant block-sum partially balanced incomplete block designs. Communications in Statistics: Theory and Methods, 49(11):2585–2606, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1576895.

Khattree:2022:CEC

[Kha22] Ravindra Khattree. On construction of equireplicated constant block-sum designs. Communications in Statistics: Theory and Methods, 51(13):4434–4450, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1814816.

Khardani:2024:BPA

[Kha24] Salah Khardani. A Bernstein polynomial approach to the estimation of a distribution function and quantiles under censorship model. Communications in Statistics: Theory and Methods, 53(16):5673–5686, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2228948.

Kouakou:2024:EZI

[KHD24] Konan Jean Geoffroy Kouakou, Ouagnina Hili, and Jean-François Dupuy. Estimation of zero-inflated bivariate Poisson regression with missing covariates. Communications in Statistics: Theory and Methods, 53(20):7216–7243, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2262637.

Kuiti:2022:NSP

[KHF22] Mithu Rani Kuiti, Nil Kamal Hazra, and Maxim Finkelstein. A note on the stochastic precedence order between component redundancy and system redundancy for k-out-of-n systems. Communications in Statistics: Theory and Methods, 51(15): 5003–5011, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2020.1831539.

Kroupova:2020:KER

[KHK20] Monika Kroupová, Ivana Horová, and Jan Kolácek. Kernel estimation of regression function gradient. Communications in Statistics: Theory and Methods, 49(1):135–151, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1532518.

Kordestani:2020:MMS

[KHSS20] Moslem Kordestani, Farid Hassanvand, Yaser Samimi, and Hamid Shahriari. Monitoring multivariate simple linear profiles using robust estimators. Communications in Statistics: Theory and Methods, 49(12):2964–2989, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1584314.

Khodadadipour:2022:EEC

[KHVBRm22] M. Khodadadipour, A. Hadi-Vencheh, M. H. Behzadi, and M. Rostamy-malkhalifeh. Efficiency evaluation with crossefficiency in the presence of undesirable outputs in stochastic environment. Communications in Statistics: Theory and Methods, 51(22):7691–7712, 2022. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2021.1879859.

Kou:2024:UAS

[KHZ24] Junke Kou, Qinmei Huang, and Hao Zhang. Uniform almost sure convergence rate of wavelet estimator for regression model with mixed noise. Communications in Statistics: Theory and Methods, 53(13):4805–4818, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2195032.

Khalil:2021:NFA

[KIA^{+}21] Alamgir Khalil, Muhammad Ijaz, Kashif Ali, Wali Khan Mashwani, Muhammad Shafiq, Poom Kumam and Wiyada Kumam. A novel flexible additive Weibull distribution with real-life applications. Communications in Statistics: Theory and Methods, 50(7):1557–1572, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1732658.

Kiapour:2022:BEE

[Kia22] Azadeh Kiapour. Bayesian estimation of the expected queue length of a system M/M/1 with certain and uncertain priors. Communications in Statistics: Theory and Methods, 51(15): 5310–5317, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2020.1838543.

Kim:2020:MDE

[Kim20] Jiwoong Kim. Minimum distance estimation in linear regression with strong mixing errors. Communications in Statistics: Theory and Methods, 49(6):1475–1494, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1563178.

Kim:2021:EPC

[Kim21] Sungsu Kim. Estimation of population column proportions of 1’s in survey type designs. Communications in Statistics: Theory and Methods, 50(23):5554–5560, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1734827.

Kittaneh:2021:CAE

[Kit21] O. A. Kittaneh. The conditional average entropies. Communications in Statistics: Theory and Methods, 50(24):6256–6263, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2020.1742356.

Kumar:2023:MMQ

[KJ23] Anshul Kumar and Madhu Jain. M/M/1 queue with bi-level network process and bi-level vacation policy with balking. Communications in Statistics: Theory and Methods, 52(15): 5502–5526, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2021.2012197.

Kumar:2024:PAT

[KJ24] C. Satheesh Kumar and Rosmi Jose. Properties and applications of two-tailed quasi-Lindley distribution. Communications in Statistics: Theory and Methods, 53(22):8140–8154, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2023.2279915.

Kim:2024:NEE

[KJAZ24] Hyoung-Moon Kim, Yu-Hyeong Jang, Barry C. Arnold, and Jun Zhao. New efficient estimators for the Weibull distribution. Communications in Statistics: Theory and Methods, 53(13):4576–4601, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03610926.2023.2179880.

Kundu:2020:SSR

[KJKC20] Piyali Kundu, Nabakumar Jana, Somesh Kumar, and Kashinath Chatterjee. Stress-strength reliability estimation for exponentially distributed system with common minimum guarantee time. Communications in Statistics: Theory and Methods, 49(14):3375–3396, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1586948.

Kazemitabar:2020:MCD

[KK20] Jalil Kazemitabar and Javad Kazemitabar. Measuring the conformity of distributions to Benford’s Law. Communications in Statistics: Theory and Methods, 49(14):3530–3536, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2019.1590599.

Kim:2021:AAI

[KK21] Jeankyung Kim and Hyune-Ju Kim. Applications of asymptotic inference in segmented line regression. Communications in Statistics: Theory and Methods, 50(23):5585–5606, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1734835.

Khalaj:2022:DCS

[KK22] Fereshteh Khalaj and Mehran Khalaj. Developed cosine similarity measure on belief function theory: an application in medical diagnosis. Communications in Statistics: Theory and Methods, 51(9):2858–2869, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1782935.

Kayal:2023:ORE

[KK23a] Suchandan Kayal and Amarjit Kundu. Ordering results on extremes of inverse Kumaraswamy random samples. Communications in Statistics: Theory and Methods, 52(9):2998–3011, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2021.1966468.

Khalaj:2023:IDM

[KK23b]  Mehran Khalaj and Fereshteh Khalaj.  An improve-

ment decision-making method by similarity and belief function theory. Communications in Statistics: Theory and Methods, 52(7):2240–2258, 2023. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2021.1949472.

Kumar:2023:TNT

[KK23c] Sanjay Kumar and Nand Kumar. Testing normality in the time series of EMP indices: an application and powercomparison of alternative tests. Communications in Statistics: Theory and Methods, 52(2):364–377, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1914097.

Kocyigit:2024:ITA

[KK24] Eda Gizem Koçyigit and Cem Kadilar. Information theory approach to ranked set sampling and new sub-ratio estimators. Communications in Statistics: Theory and Methods, 53(4): 1331–1353, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2022.2100910.

Kubicka:2021:TPR

[KKKM21] Ewa M. Kubicka, Grzegorz Kubicki, Małgorzata Kuchta, and Michał Morayne. Tail probabilities of a random walk on an interval. Communications in Statistics: Theory and Methods, 50(9):2161–2169, 2021. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2019.1662044.

Kissita:2023:GMC

[KKNN23] Gabriel Kissita, Christian D’Aquin Nzouzi Kibangou, Léonard Niéré, and Guy Martial Nkiet. The Generalized Multiple CO-inertia Analysis (GMCOA). Communications in Statistics: Theory and Methods, 52(22):7861–7885, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2051049.

Khalid:2024:EQR

[KKS24a] Memoona Khalid, Hina Khana, and Javid Shabbir. Exponentially quantile regression-ratio-type estimators for robust mean estimation. Communications in Statistics: Theory and Methods, 53(19):7069–7086, 2024. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2023.2258426.

Kharrati-Kopaei:2024:AUT

[KKS24b] Mahmood Kharrati-Kopaei and Zahra Shenavari. Analyzing unreplicated two-level factorial designs by combining multiple tests. Communications in Statistics: Theory and Methods, 53(13):4680–4695, 2024. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2023.2185752.

Krishnamoorthy:2020:PIH

[KL20] Kalimuthu Krishnamoorthy and Shanshan Lv. Prediction intervals for hypergeometric distributions. Communications in Statistics: Theory and Methods, 49(6):1528–1536, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1563181.

Kim:2022:HCN

[KL22a] Min-Ho Kim and Jea-Young Lee. How to construct a nomogram for hypertension using complex sampling data from Korean adults. Communications in Statistics: Theory and Methods, 51(8):2357–2367, 2022. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2020.1774057.

Kong:2022:OEL

[KL22b] Xiaoyu Kong and Yuhua Lü. Optimal excess-of-loss reinsurance and investment with stochastic factor process. Communications in Statistics: Theory and Methods, 51(24):8705–8727, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2021.1904989.

Kang:2021:PPB

[KLK21] Sang Gil Kang, Woo Dong Lee, and Yongku Kim. Posterior propriety of bivariate Lomax distribution under objective priors. Communications in Statistics: Theory and Methods, 50(9):2201–2209, 2021. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2019.1662049.

Kang:2022:OBI

[KLK22] Sang Gil Kang, Woo Dong Lee, and Yongku Kim. Objective Bayesian inference for quantile ratios in normal models. Communications in Statistics: Theory and Methods, 51(15):5085– 5111, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2020.1833220.

Kang:2023:OPC

[KLK23] Sang Gil Kang, Woo Dong Lee, and Yongku Kim. Objective priors for common correlation coefficient in bivariate normal populations. Communications in Statistics: Theory and Methods, 52(7):2124–2143, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1945630.

Kang:2024:AOP

[KLK24] Sang Gil Kang, Woo Dong Lee, and Yongku Kim. Application of objective priors for the multivariate Lomax distribution. Communications in Statistics: Theory and Methods, 53(7): 2307–2328, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2022.2126945.

Kadankova:2023:FNH

[KLS23] Tetyana Kadankova, Nikolai Leonenko, and Enrico Scalas. Fractional non-homogeneous Poisson and Pólya–Aeppli processes of order k and beyond. Communications in Statistics: Theory and Methods, 52(8):2682–2701, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1958228.

Ke:2022:RQC

[KLSZ22] Jau-Chuan Ke, Tzu-Hsin Liu, Siping Su, and Zhe-George Zhang. On retrial queue with customer balking and feedback subject to server breakdowns. Communications in Statistics: Theory and Methods, 51(17):6049–6063, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1852432.

Kattumannil:2021:FNP

[KM21] Sudheesh K. Kattumannil and Deemat C. Mathew. A family of non-parametric tests for decreasing mean time to failure with censored data. Communications in Statistics: Theory and Methods, 50(1):203–215, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1634207.

Kaombe:2023:DID

[KM23a] Tsirizani M. Kaombe and Samuel O. M. Manda. Detecting influential data in multivariate survival models. Communications in Statistics: Theory and Methods, 52(11):3910–3926, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2021.1982983.

Kharfouchi:2023:EDG

[KM23b] Soumia Kharfouchi and Wafa Mili. Estimating 2-D GARCH models by quasi-maximum of likelihood. Communications in Statistics: Theory and Methods, 52(17):6275–6286, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2027452.

Kemda:2024:JML

[KM24a] Lionel Establet Kemda and Michael Murray. Joint modeling of the longitudinal student mark and the competing events of degree completion and academic dropout. Communications in Statistics: Theory and Methods, 53(11):3992–4011, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2170180.

Khariton:2024:DFM

[KM24b] Yakov Khariton and Nitis Mukhopadhyay. Distributionfree minimum risk point estimation of the mean under powered absolute error loss plus cost of sampling: Illustrations with cancer data. Communications in Statistics: Theory and Methods, 53(7):2617–2644, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2147796.

Kitani:2023:DSI

[KMH23] Masato Kitani, Hidetoshi Murakami, and Hiroki Hashiguchi. The distribution of the sum of independent and non identically generalized Lindley random variables. Communications in Statistics: Theory and Methods, 52(8):2597–2609, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1955387.

Kayal:2022:SCP

[KN22] Suchandan Kayal and Phalguni Nanda. Stochastic comparisons of parallel systems with generalized Kumaraswamy-G components. Communications in Statistics: Theory and Methods, 51(14):4712–4738, 2022. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2020.1821889.

Kaban:2024:INL

[KNCK24] Puspita Anggraini Kaban, Bahrul Ilmi Nasution, Rezzy Eko Caraka, and Robert Kurniawan. Implementing night light data as auxiliary variable of small area estimation. Communications in Statistics: Theory and Methods, 53(1):310–327, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2022.2077963.

Kammoo:2024:LLT

[KNL24] Punyapat Kammoo, Kritsana Neammanee, and Kittipong Laipaporn. The local limit theorem for general weighted sums of Bernoulli random variables. Communications in Statistics: Theory and Methods, 53(13):4918–4926, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2198623.

Ko:2020:NCM

[Ko20] Mi-Hwa Ko. A note on complete moment convergence for coordinatewise negatively associated random vectors in Hilbert spaces. Communications in Statistics: Theory and Methods, 49(7):1780–1791, 2020. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2019.1565833.

Khan:2023:RME

[KO23a] Manzoor Khan and Jake Olivier. Regression to the mean: Estimation and adjustment under the bivariate normal distribution. Communications in Statistics: Theory and Methods, 52(19):6972–6990, 2023. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2022.2037645.

Ko:2023:CMC

[Ko23b] Mi-Hwa Ko. The complete moment convergence for coordinatewise pairwise negatively quadrant dependent random vectors in Hilbert space. Communications in Statistics: Theory and Methods, 52(23):8467–8477, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2064500.

Ko:2023:CPW

[Ko23c] Mi-Hwa Ko. On convergence properties for weighted sums of coordinatewise ANA random vectors in Hilbert spaces. Communications in Statistics: Theory and Methods, 52(17):6136– 6146, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2022.2026395.

Ko:2023:SLB

[Ko23d] Mi-Hwa Ko. Some limiting behavior of the maximum of the partial sum for asymptotically negatively associated random vectors in Hilbert space. Communications in Statistics: Theory and Methods, 52(11):3598–3611, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1977957.

Kobayashi:2024:NPM

[Kob24] Yasuyuki Kobayashi. New precise model of studentized principal components. Communications in Statistics: Theory and Methods, 53(2):487–504, 2024. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2022.2084110.

Koc:2021:RTE

[Koç21] Haydar Koç. Ratio-type estimators for improving mean estimation using Poisson regression method. Communications in Statistics: Theory and Methods, 50(20):4685–4691, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1777307.

Kohansal:2020:BCE

[Koh20] Akram Kohansal. Bayesian and classical estimation of R = P (X < Y) based on Burr type XII distribution under hybrid progressive censored samples. Communications in Statistics: Theory and Methods, 49(5):1043–1081, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1554126.

Kong:2022:DSA

[Kon22] Yong Kong. Distributions of successions of arbitrary multisets. Communications in Statistics: Theory and Methods, 51(6): 1693–1705, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2020.1767137.

Kebaili:2021:ARD

[KOTAG21] Siham Kebaili, Megdouda Ourbih-Tari, Abdelouhab Aloui, and Sofia Guebli. Adaptive refined descriptive sampling algorithm for dependent variables using Iman and Conover method in Monte Carlo simulation. Communications in Statistics: Theory and Methods, 50(15):3632–3644, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1709645.

Kou:2023:ISR

[Kou23] Jiyao Kou. Identifying the support of rectangular signals in Gaussian noise. Communications in Statistics: Theory and Methods, 52(10):3262–3289, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1970771.

Kim:2021:MAR

[KP21] Sungsu Kim and Thelge Buddika Peiris. Meta analysis of regression: a review and new approach with application to linear-circular regression model. Communications in Statistics: Theory and Methods, 50(12):2723–2731, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1679183.

Kayal:2024:ESP

[KP24] Suchandan Kayal and Lakshmi Kanta Patra. Estimating the scale parameters of several exponential distributions under order restriction. Communications in Statistics: Theory and Methods, 53(23):8484–8497, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2292967.

Kim:2024:URT

[KPH^{+}24] Seul Gee Kim, Cheolyong Park, Sun-Young Hwang, Jeongcheol Ha, Inho Park, and Tae Yoon Kim. Unit root tests and their challenges. Communications in Statistics: Theory and Methods, 53(8):2840–2847, 2024. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2022.2150050.

Kim:2021:DAC

[KR21] Sungsu Kim and Md Monjurul Islam Rifat. Diagnostic analysis of a circular–circular regression model using asymmetric or asymmetric bi-modal circular errors. Communications in Statistics: Theory and Methods, 50(12):2848–2858, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1676448.

Klimova:2022:TFR

[KR22] Anna Klimova and Tamás Rudas. Testing the fit of relational models. Communications in Statistics: Theory and Methods, 51(23):8264–8282, 2022. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2021.1891437.

Kumar:2023:GZI

[KR23a] C. Satheesh Kumar and Rakhi Ramachandran. A generalization to zero-inflated hyper-Poisson distribution: Properties and applications. Communications in Statistics: Theory and Methods, 52(20):7289–7302, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2043378.

Kumar:2023:BAL

[KR23b] C. Satheesh Kumar and A. Riyaz. On bivariate alpha logarithmic series distribution and its properties. Communications in Statistics: Theory and Methods, 52(24):8875–8883, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2081708.

Kabaila:2024:CIG

[KR24a] Paul Kabaila and Nishika Ranathunga. Confidence intervals in general regression models that utilize uncertain prior information. Communications in Statistics: Theory and Methods, 53(17):6266–6284, 2024. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2023.2243528.

Katebi:2024:CAD

[KR24b] M. S. Mehdi Katebi and Abdur Rahim. A combined adaptive double sampling and variable sampling interval control chart for monitoring three-level products. Communications in Statistics: Theory and Methods, 53(19):7031–7053, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2256439.

Kumar:2024:SDA

[KRCS24] Nirpeksh Kumar, Athanasios C. Rakitzis, Subha Chakraborti, and Tribhuvan Singh. Statistical design of ATS-unbiased charts with runs rules for monitoring exponential time between events. Communications in Statistics: Theory and Methods, 53(3):815–833, 2024. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2022.2092143.

Kreider:2024:ES

[Kre24] Greg Kreider. Expected spacing. Communications in Statistics: Theory and Methods, 53(23):8286–8296, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2281265.

Khan:2020:OMC

[KRS20] M. G. M. Khan, Mahmood A. Rashid, and Sushita Sharma. An optimal multivariate cluster sampling design. Communications in Statistics: Theory and Methods, 49(24):6121–6128, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2019.1626430.

Kundu:2020:GIE

[KS20] Chanchal Kundu and Shivangi Singh. On generalized interval entropy. Communications in Statistics: Theory and Methods, 49(8):1989–2007, 2020. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2019.1568480.

Kulkarni:2021:COR

[KS21] Leena Kulkarni and Sanjeev Sabnis. Construction of optimal reliability test plans for binary type multi-state strongly coherent systems. Communications in Statistics: Theory and Methods, 50(12):2780–2800, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1676444.

Krishnamoorthy:2022:RPS

[KS22] Kalimuthu Krishnamoorthy and Arvind K. Shah. A report on the paper “Sungsu Kim. 2019. The probable error in the hypothesis test of normal means using a small sample. Communications in Statistics — Theory and Methods. DOI: 10.1080/03610926.2019.1703135”. Communications in Statistics: Theory and Methods, 51(10):3459–3461, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1767786.

Kang:2024:HML

[KS24a] Xiong Kang and Brian Yi Sun. Hyper Markov law in undirected graphical models with its applications. Communications in Statistics: Theory and Methods, 53(20):7262–7280, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2023.2263111.

Kaur:2024:NPD

[KS24b] Manpreet Kaur and Amit Srivastava. Negation of a probability distribution: an information theoretic analysis. Communications in Statistics: Theory and Methods, 53(17):6252–6265, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2023.2242986.

Kumar:2024:EEH

[KS24c] C. Satheesh Kumar and A. S. Satheenthar. An extended exponential hyper-Poisson distribution: Properties and applications. Communications in Statistics: Theory and Methods, 53(12):4311–4333, 2024. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2023.2178261.

Kwon:2024:SPC

[KS24d] Oh Kang Kwon and Stephen Satchell. Statistical properties of co-quantiles and their applications to momentum spillovers. Communications in Statistics: Theory and Methods, 53(20): 7332–7351, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2023.2263116.

Khan:2023:EBE

[KSA23] Nida Khan, Said Farooq Shah, and Syed Muhammad Asim. Efficient Bayes estimators of sensitive proportion with simple and mixture priors using direct and indirect responses. Communications in Statistics: Theory and Methods, 52(20):7502– 7531, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2022.2048308.

Kumar:2023:TSS

[KSK^{+}23] B. Krishna Kumar, R. Sankar, R. Navaneetha Krishnan, R. Rukmani, and Alexander Rumnyantsev. Transient and steady-state analysis of hybrid arrivals of single and batch customers queueing systems with switch-off period. Communications in Statistics: Theory and Methods, 52(10):3382–3413, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2021.1973501.

Khalil:2021:NAO

[KSM^{+}21] Alamgir Khalil, Salahuddin, Wali Khan Mashwani, Muhammad Shafiq, Saima Hassan, and Wiyada Kumam. New advanced outliers detection tests. Communications in Statistics: Theory and Methods, 50(7):1640–1655, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1741630.

Kumar:2021:IPF

[KSS21] Abhishek Kumar, Ajeet Kumar Singh, and V. K. Singh. Investigating the performance of a family of exponential-type estimators in presence of measurement error. Communications in Statistics: Theory and Methods, 50(12):2958–2977, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2019.1682167.

Kubokawa:2020:SEL

[KSY20] Tatsuya Kubokawa, William E. Strawderman, and Ryota Yuasa. Shrinkage estimation of location parameters in a multivariate skew-normal distribution. Communications in Statistics: Theory and Methods, 49(8):2008–2024, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1568481.

Kobayashi:2021:IUB

[KT21] Kensho Kobayashi and Hidekazu Tanaka. Improvement of the Ushakov bound. Communications in Statistics: Theory and Methods, 50(24):5929–5940, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1737878.

Kock:2024:HMT

[KT24] Ned Kock and Augustine Tarkom. How many times until a coincidence becomes a pattern? The case of yield curve inversions preceding recessions and the magical number 7. Communications in Statistics: Theory and Methods, 53(16):5785– 5792, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2023.2232908.

Kayal:2020:ERM

[KTDW20] Tanmay Kayal, Yogesh Mani Tripathi, Sanku Dey, and Shuo-Jye Wu. On estimating the reliability in a multicomponent stress-strength model based on Chen distribution. Communications in Statistics: Theory and Methods, 49(10):2429–2447, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2019.1576886.

Khalaj:2020:NDC

[KTMKS20] Mehran Khalaj, Reza Tavakkoli-Moghaddam, Fereshteh Khalaj, and Ali Siadat. New definition of the cross entropy based on the Dempster–Shafer theory and its application in a decision-making process. Communications in Statistics: Theory and Methods, 49(4):909–923, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1554123.

Khatun:2022:HTI

[KTP22] Habiba Khatun, Manas Ranjan Tripathy, and Nabendu Pal. Hypothesis testing and interval estimation for quantiles of two normal populations with a common mean. Communications in Statistics: Theory and Methods, 51(16):5692–5713, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1845735.

Kumar:2022:DGM

[KU22] A. N. Kumar and N. S. Upadhye. On discrete Gibbs measure approximation to runs. Communications in Statistics: Theory and Methods, 51(5):1488–1513, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1765256.

Kumar:2024:PMM

[Kum24] Nitin Kumar. A population model with Markovian arrival process and binomial correlated catastrophes. Communications in Statistics: Theory and Methods, 53(20):7181–7196, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2023.2261059.

Kundu:2023:CRP

[Kun23] Chanchal Kundu. On cumulative residual (past) extropy of extreme order statistics. Communications in Statistics: Theory and Methods, 52(16):5848–5865, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.2021238.

Kundu:2024:WLS

[Kun24] Debasis Kundu. Weighted least squares: a robust method of estimation for sinusoidal model. Communications in Statistics: Theory and Methods, 53(16):5935–5953, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2238362.

Kurita:2022:RUP

[Kur22] Takamitsu Kurita. Revealing unnoticed properties of super exogeneity in a cointegrated vector autoregression. Communications in Statistics: Theory and Methods, 51(23):8349–8370, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2021.1894448.

Kurata:2024:RMS

[Kur24] Sumito Kurata. On robustness of model selection criteria based on divergence measures: Generalizations of BHHJ divergence-based method and comparison. Communications in Statistics: Theory and Methods, 53(10):3499–3516, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2155788.

Kyriakoussis:2023:CNH

[KV23] Andreas Kyriakoussis and Malvina Vamvakari. Correction notice to “Heine process as a q-analog of the Poisson process-waiting and interarrival times”. Communications in Statistics: Theory and Methods, 52(14):5113, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1891440.

Kariya:2022:AAS

[KW22] Natsuki Kariya and Sumio Watanabe. Asymptotic analysis of singular likelihood ratio of normal mixture by Bayesian learning theory for testing homogeneity. Communications in Statistics: Theory and Methods, 51(17):5873–5888, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1849721.

Kassahun:2023:RBT

[KWB23] Taddesse Kassahun, Eshetu Wencheko, and Arne C. Bathke. A rank-based test for monotone trends in k-sample multivariate problems. Communications in Statistics: Theory and Methods, 52(3):864–885, 2023. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2021.1931333.

Kang:2021:RMB

[KWC21] Yao Kang, Dehui Wang, and Jianhua Cheng. Risk models based on copulas for premiums and claim sizes. Communications in Statistics: Theory and Methods, 50(10):2250–2269, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2019.1662443.

Kwon:2022:AEP

[Kwo22] Oh Kang Kwon. Analytic expressions for the positive definite and unimodal regions of Gram–Charlier series. Communications in Statistics: Theory and Methods, 51(15):5064–5084, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2020.1833219.

Kang:2020:EBA

[KWY20] Yao Kang, Dehui Wang, and Kai Yang. Extended binomial AR(1) processes with generalized binomial thinning operator. Communications in Statistics: Theory and Methods, 49(14): 3498–3520, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2019.1589519.

Kuang:2023:DTM

[KX23] Yongxin Kuang and Junshan Xie. Distributed testing on mutual independence of massive multivariate data. Communications in Statistics: Theory and Methods, 52(15):5332–5348, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2021.2006232.

Liu:2023:AID

[LA23] Bowen Liu and Malwane M. A. Ananda. Analyzing insurance data with an exponentiated composite inverse gamma-Pareto model. Communications in Statistics: Theory and Methods, 52(21):7618–7631, 2023. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2022.2050399.

Lakaour:2022:USS

[LAAA^{+}22] Lamia Lakaour, Djamil Aissani, Karima Adel-Aissanou, Kamel Barkaoui, and Sofiane Ziani. An unreliable single server retrial queue with collisions and transmission errors. Communications in Statistics: Theory and Methods, 51(4):1085–1109, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2020.1758943.

Labouriau:2023:BSF

[Lab23] R. Labouriau. On the bias of the score function of finite mixture models. Communications in Statistics: Theory and Methods, 52(13):4461–4467, 2023. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2021.1995429.

LaMotte:2020:FTI

[LaM20] Lynn Roy LaMotte. A formula for Type III sums of squares. Communications in Statistics: Theory and Methods, 49(13): 3126–3136, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2019.1586933.

LaMotte:2024:TAE

[LaM24] Lynn R. LaMotte. Testing ANOVA effects: a resolution for unbalanced models. Communications in Statistics: Theory and Methods, 53(22):7860–7870, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2274808.

Lavine:2024:PVD

[Lav24] Michael Lavine. P-values don’t measure evidence. Communications in Statistics: Theory and Methods, 53(2):718–726, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2022.2091783.

Linke:2022:INW

[LB22] Yuliana Linke and Igor Borisov. Insensitivity of nadaraya-Watson estimators to design correlation. Communications in Statistics: Theory and Methods, 51(19):6909–6918, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1876884.

Letmathe:2024:EES

[LBF24] Sebastian Letmathe, Jan Beran, and Yuanhua Feng. An extended exponential SEMIFAR model with application in R. Communications in Statistics: Theory and Methods, 53(22): 7914–7926, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2023.2276049.

Laxmi:2023:RIC

[LBG23] P. Vijaya Laxmi, E. Girija Bhavani, and Andwilile Abrahamu George. Retention of impatient customers in a multi-server Markovian queueing system with optional service and working vacations. Communications in Statistics: Theory and Methods, 52(15):5195–5212, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.2004427.

Li:2020:NMM

[LC20] Zhongping Li and Lirong Cui. Numerical method for means of linear Hawkes processes. Communications in Statistics: Theory and Methods, 49(15):3681–3697, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1713374.

Laumen:2022:SLT

[LC22a] Benjamin Laumen and Erhard Cramer. Stage life testing with random stage changing times. Communications in Statistics: Theory and Methods, 51(12):3934–3959, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1805764.

Liu:2022:NRM

[LC22b] Yi Liu and Xiaolin Chen. A new robust model-free feature screening method for ultra-high dimensional right censored data. Communications in Statistics: Theory and Methods, 51(6):1857–1875, 2022. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2020.1769672.

Lee:2024:ESD

[LC24] Pei-Hsi Lee and Chao-Yu Chou. An economic-statistical design of synthetic Tukey’s control chart with Taguchi’s asymmetric loss functions under log-normal distribution. Communications in Statistics: Theory and Methods, 53(21):7604– 7623, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2023.2269448.

Lu:2020:MDR

[LCF20] Dawei Lu, Xiaoyan Chen, and Jinghai Feng. Moderate deviations for the random weighted sums of WUOD random variables with consistently varying tails. Communications in Statistics: Theory and Methods, 49(3):531–551, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1543775.

Liu:2024:ART

[LCF24] Yang Liu, Zhenlong Chen, and Ke-Ang Fu. Asymptotics for the random time ruin probability with non stationary arrivals and Brownian perturbation. Communications in Statistics: Theory and Methods, 53(9):3337–3349, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2153227.

Ling:2021:CMO

[LCNB21] M. H. Ling, P. S. Chan, H. K. T. Ng, and N. Balakrishnan. Copula models for one-shot device testing data with correlated failure modes. Communications in Statistics: Theory and Methods, 50(16):3875–3888, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1725827.

Lu:2021:CDC

[LCW21] Shuiyun Lu, Xiaolin Chen, and Hong Wang. Conditional distance correlation sure independence screening for ultrahigh dimensional survival data. Communications in Statistics: Theory and Methods, 50(8):1936–1953, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1657454.

Luo:2021:AEC

[LCWZ21] Jing Luo, Qianqian Chen, Zhenghong Wang, and Laala Zeyneb. Approximate estimation in a class of directed networks. Communications in Statistics: Theory and Methods, 50(21):4963–4976, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03610926.2019.1565841.

Liu:2024:NUM

[LCYL24] Zhe Liu, Yu Cao, Xiangli Yang, and Lusi Liu. A new uncertainty measure via belief Rényi entropy in Dempster–Shafer theory and its application to decision making. Communications in Statistics: Theory and Methods, 53(19):6852–6868, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2023.2253342.

Liu:2022:LDE

[LD22] Qun Liu and Zhishan Dong. Large deviations for empirical measures of generalized random graphs. Communications in Statistics: Theory and Methods, 51(8):2676–2687, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1779748.

Lu:2022:CCR

[LDZ^{+}22] Zhengliang Lu, Shengnan Ding, Fei Zhang, Rui Wang, and Xuejun Wang. The consistency and convergence rate for the nearest neighbor density estimator based on ϕ-mixing random samples. Communications in Statistics: Theory and Methods, 51(3):669–684, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1752727.

LeGuevel:2021:GFT

[Le 21] R. Le Guével. Goodness-of-fit test for multistable Lévy processes. Communications in Statistics: Theory and Methods, 50(8):1807–1837, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03610926.2019.1653922.

Lee:2020:IPR

[Lee20] Dong-Hee Lee. An inspection procedure for radio frequency repeaters using a multiple linear regression method. Communications in Statistics: Theory and Methods, 49(13):3137–3152, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2019.1586934.

Lee:2021:GSM

[Lee21] Hyunju Lee. A general shock model for modelling coupled lives and its application to life insurance. Communications in Statistics: Theory and Methods, 50(12):2859–2876, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1676449.

Lee:2022:ABC

[Lee22a] Jin Lee. Asymptotic behavior of cross spectral density estimator at the zero frequency in the presence of degeneracy. Communications in Statistics: Theory and Methods, 51(3):840– 851, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2020.1758142.

Lee:2022:GSC

[Lee22b] Se Yoon Lee. Gibbs sampler and coordinate ascent variational inference: a set-theoretical review. Communications in Statistics: Theory and Methods, 51(6):1549–1568, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1921214.

Lee:2023:MDR

[Lee23] Hyunju Lee. Mixture discrete reversed hazard rate and its main properties. Communications in Statistics: Theory and Methods, 52(22):8053–8068, 2023. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2022.2056747.

Lefebvre:2022:FPP

[Lef22] Mario Lefebvre. First-passage problems for diffusion processes with state-dependent jumps. Communications in Statistics: Theory and Methods, 51(9):2908–2918, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1784433.

Lemonte:2022:CMM

[Lem22] Artur J. Lemonte. Covariance matrix of maximum likelihood estimators in censored exponential regression models. Communications in Statistics: Theory and Methods, 51(6):1765– 1777, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2020.1767142.

Lopez:2024:ROF

[LEM24] Romario A. Conto López, Alexander A. Correa Espinal, and Olga C. Úsuga Manco. Run orders in factorial designs: a literature review. Communications in Statistics: Theory and Methods, 53(13):4557–4575, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2185472.

Leulmi:2021:LLE

[Leu21] Sara Leulmi. Local linear estimation of the conditional quantile for censored data and functional regressors. Communications in Statistics: Theory and Methods, 50(14):3286–3300, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2019.1692033.

Lin:2020:CCS

[LF20] Yiwei Lin and Xinwei Feng. Complete convergence and strong law of large numbers for arrays of random variables under sublinear expectations. Communications in Statistics: Theory and Methods, 49(23):5866–5882, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1625924.

Li:2023:SGT

[LF23] Xinying Li and Shengjun Fan. L^{p} solutions of general time interval BSDEs with generators satisfying a p-order weak stochastic-monotonicity condition. Communications in Statistics: Theory and Methods, 52(16):5650–5676, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.2016833.

Liu:2022:SIB

[LG22] Shuhan Liu and Wenhao Gui. Statistical inference for bathtub-shaped distribution based on generalized progressive hybrid censored data. Communications in Statistics: Theory and Methods, 51(23):8149–8172, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1889602.

Li:2024:RLM

[LG24a] Yu-Shan Li and Chao-Yu Guo. Random logistic machine (RLM): Transforming statistical models into machine learning approach. Communications in Statistics: Theory and Methods, 53(21):7517–7525, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2268767.

Liu:2024:UAN

[LG24b] Xijun Liu and Qingwu Gao. Uniform asymptotics for a nonstandard compound renewal risk model with dependence structures and stochastic return on investments. Communications in Statistics: Theory and Methods, 53(2):641–665, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2088792.

Lobato:2024:CPE

[LG24c] Adriana Laura López Lobato and Martha Lorena Avendaño Garrido. Convergence of parameter estimation of a Gaussian mixture model minimizing the Gini index of dissimilarity. Communications in Statistics: Theory and Methods, 53(17): 6030–6037, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2023.2239396.

Liu:2020:LDS

[LGL20] Xijun Liu, Qingwu Gao, and Ming Liu. Large deviations for the stochastic present value of aggregate claims in the nonstandard compound renewal risk model with widely upper orthant dependent claims. Communications in Statistics: Theory and Methods, 49(13):3073–3093, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1586931.

Liu:2021:ILF

[LGLL21] Fuxiang Liu, Wenzhang Guo, Ran Liu, and Jun Liu. Improved load forecasting model based on two-stage optimization of gray model with fractional order accumulation and Markov chain. Communications in Statistics: Theory and Methods, 50(11):2659–2673, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1674873.

Liu:2022:MOO

[LGX^{+}22] Xiaoping Liu, Bin Guo, Lijian Xia, Xiao Tian, and Lijie Zhang. Multi-objective optimization design of accelerated degradation test based on Wiener process. Communications in Statistics: Theory and Methods, 51(5):1426–1443, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1764043.

Liu:2022:ROD

[LGZ22] Pengqi Liu, Lucy L. Gao, and Julie Zhou. R-optimal designs for multi-response regression models with multi-factors. Communications in Statistics: Theory and Methods, 51(2):340– 355, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2020.1748655.

Li:2021:BBL

[LH21] Yi Li and Zongyi Hu. Bayesian bent line quantile regression model. Communications in Statistics: Theory and Methods, 50(17):3972–3987, 2021. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2019.1710750.

Li:2022:CRS

[LH22] Yingqiu Li and Xulan Huang. A.s. convergence rate for a supercritical branching processes with immigration in a random environment. Communications in Statistics: Theory and Methods, 51(3):826–839, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1756330.

Li:2021:NOS

[LHJ21] Weiwei Li, Jan Hannig, and Corbin D. Jones. A note on optimal sampling strategy for structural variant detection using optical mapping. Communications in Statistics: Theory and Methods, 50(20):4763–4777, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1723638.

Li:2022:NRK

[LHLD22] Yi Li, Zongyi Hu, Jiaqi Liu, and Jingjing Deng. A note on regression kink model. Communications in Statistics: Theory and Methods, 51(23):8246–8263, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1890780.

Liu:2020:RAI

[LHS20] Yeqian Liu, Tao Hu, and Jianguo Sun. Regression analysis of interval-censored failure time data with cured subgroup and mismeasured covariates. Communications in Statistics: Theory and Methods, 49(1):189–202, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1535075.

Lee:2022:SES

[LHS22] Gi-Sung Lee, Ki-Hak Hong, and Chang-Kyoon Son. A stratified estimation of a sensitive attribute by using negative binomial and negative hypergeometric distribution as randomization devices. Communications in Statistics: Theory and Methods, 51(20):7148–7171, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1871020.

Li:2023:NSE

[LHS23] Fan Li, Hwanhee Hong, and Elizabeth A. Stuart. A note on semiparametric efficient generalization of causal effects from randomized trials to target populations. Communications in Statistics: Theory and Methods, 52(16):5767–5798, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.2020291.

Lee:2020:SES

[LHSK20] Gi-Sung Lee, Ki-Hak Hong, Chang-Kyoon Son, and Jong-Min Kim. A stratified estimation of a sensitive character by using geometric distribution. Communications in Statistics: Theory and Methods, 49(13):3184–3205, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1586939.

Li:2024:RAO

[LHWK24] Jing Li, Linmin Hu, Yuyu Wang, and Jia Kang. Reliability analysis and optimization design of a repairable kout-of-n retrial system with two failure modes and preventive maintenance. Communications in Statistics: Theory and Methods, 53(15):5524–5552, 2024. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2023.2222317.

Liu:2022:DIR

[LHXZ22] Juan Liu, Ya Huang, Xuyan Xiang, and Jieming Zhou. On a discrete interaction risk model with delayed claims and randomized dividends. Communications in Statistics: Theory and Methods, 51(15):5241–5257, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1836221.

Li:2020:NSM

[Li20] Yong Li. A new stochastic mixed Liu estimator in linear regression model. Communications in Statistics: Theory and Methods, 49(3):726–737, 2020. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2018.1549250.

Li:2021:JMT

[Li21a] Fei Li. The Jacobians of matrix transformation about singular random matrices and its applications. Communications in Statistics: Theory and Methods, 50(20):4718–4732, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1723634.

Li:2021:SEA

[Li21b] Jun Li. Simple and efficient adaptive two-sample tests for high-dimensional data. Communications in Statistics: Theory and Methods, 50(19):4428–4447, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1716253.

Li:2022:HDD

[Li22] Junlin Li. High-dimensional dynamic systems identification with additional constraints. Communications in Statistics: Theory and Methods, 51(15):5204–5225, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1836219.

Li:2023:ELS

[Li23a] Jinyu Li. Empirical likelihood for special self-exciting threshold autoregressive models with heavy-tailed errors. Communications in Statistics: Theory and Methods, 52(16):5814–5835, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2021.2020842.

Li:2023:OTC

[Li23b] Sheng Li. Optimal time-consistent investment-reinsurance strategy for state-dependent risk aversion with delay and common shocks. Communications in Statistics: Theory and Methods, 52(15):5294–5331, 2023. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2021.2005804.

Li:2024:ARP

[Li24] Jinzhu Li. Asymptotic ruin probabilities for a two-dimensional risk model with dependent claims and stochastic return. Communications in Statistics: Theory and Methods, 53(16):5773– 5784, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2023.2232906.

Ligtvoet:2023:NBF

[Lig23] Rudy Ligtvoet. A note on the Bayes factor for small interval hypotheses. Communications in Statistics: Theory and Methods, 52(14):5060–5067, 2023. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2021.2002361.

Lin:2020:SOT

[Lin20] Jianxi Lin. Second order tail behaviour of randomly weighted heavy-tailed sums and their maxima. Communications in Statistics: Theory and Methods, 49(11):2648–2670, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1576900.

Lin:2021:SOA

[Lin21] Jianxi Lin. Second order asymptotics for ruin probabilities of the delayed renewal risk model with heavy-tailed claims. Communications in Statistics: Theory and Methods, 50(5):1200– 1209, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2019.1648828.

Liu:2022:SPE

[Liu22] Changbiao Liu. Semi-parametric estimation of multinomial choice model with unobserved heterogeneity. Communications in Statistics: Theory and Methods, 51(24):8644–8656, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2021.1901923.

Liu:2024:PMM

[Liu24a] Yifan Liu. Penalized Mallow’s model averaging. Communications in Statistics: Theory and Methods, 53(20):7417–7435, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2023.2264995.

Liu:2024:ALV

[Liu24b] Yu Liu. Adaptive lasso variable selection method for semiparametric spatial autoregressive panel data model with random effects. Communications in Statistics: Theory and Methods, 53(6):2122–2140, 2024. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2022.2119088.

Lv:2024:FEO

[LJ24] Wujun Lv and Pingping Jiang. Foreign exchange option pricing under the 4/2 stochastic volatility model with CIR interest rates. Communications in Statistics: Theory and Methods, 53(7):2670–2687, 2024. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2023.2279921.

Luo:2021:MSP

[LJL^{+}21] Ronghua Luo, Jiakun Jiang, Wei Lan, Chengliang Yan, and Yue Ding. A multi-step procedure to determine the number of factors in large approximate factor models. Communications in Statistics: Theory and Methods, 50(17):3988–3999, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2019.1710752.

Lu:2023:CIP

[LJLW23] Hezhi Lu, Hua Jin, Yuan Li, and Zhining Wang. Confidence intervals for a Poisson parameter with background. Communications in Statistics: Theory and Methods, 52(19):6794–6805, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2022.2033268.

Lee:2022:ASS

[LK22a] Oesook Lee and Jooyoung Kim. Asymptotics for semi-strong augmented GARCH(1,1) model. Communications in Statistics: Theory and Methods, 51(23):8093–8109, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1887894.

Lun:2022:OLE

[LK22b] Zhixin Lun and Ravindra Khattree. An overlooked lomaxexponential connection. Communications in Statistics:

Theory and Methods, 51(1):26–28, 2022. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2020.1734836.

Lv:2022:FCI

[LK22c] Shanshan Lv and K. Krishnamoorthy. Fiducial confidence intervals for proportions in finite populations: Oneand two-sample problems. Communications in Statistics: Theory and Methods, 51(12):4179–4195, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1811341.

Liu:2024:PII

[LK24] Zhe Liu and Rui Kang. Pharmacokinetics with intravenous infusion of two-compartment model based on Liu process. Communications in Statistics: Theory and Methods, 53(14):4975– 4990, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2023.2198626.

Laksaci:2023:NPE

[LKA^{+}23] Ali Laksaci, Zoulikha Kaid, Mohamed Alahiane, Idir Ouassou, and Mustapha Rachdi. Non parametric estimations of the conditional density and mode when the regressor and the response are curves. Communications in Statistics: Theory and Methods, 52(13):4659–4674, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1998831.

Low:2024:GBM

[LKK24] Victor Jian Ming Low, Wooi Chen Khoo, and Hooi Ling Khoo. A generalized Burr mixture autoregressive models for modeling non linear time series. Communications in Statistics: Theory and Methods, 53(19):6832–6851, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2252121.

Laksaci:2022:SRK

[LKS22] Ali Laksaci, Salah Khardani, and Sihem Semmar. Semirecursive kernel conditional density estimators under random censorship and dependent data. Communications in Statistics: Theory and Methods, 51(7):2116–2138, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1764038.

Li:2020:IDN

[LL20] Na Li and Xuhua Liu. Inference of the derivative of nonparametric curve based on confidence distribution. Communications in Statistics: Theory and Methods, 49(11):2607–2622, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2019.1576896.

Li:2023:TVA

[LL23a] Jiyanglin Li and Tao Li. Time-varying additive model with autoregressive errors for locally stationary time series. Communications in Statistics: Theory and Methods, 52(11):3848– 3878, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2021.1980803.

Lin:2023:LPM

[LL23b]  Jianhua Lin and Xiaohu Li.  Lack-of-partial-memory

and aging properties of multivariate generalized Marshall– Olkin distributions. Communications in Statistics: Theory and Methods, 52(5):1564–1579, 2023. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2021.1931328.

Liu:2023:BAS

[LL23c] Chang-Sheng Liu and Han-Ying Liang. Bayesian analysis in single-index quantile regression with missing observation. Communications in Statistics: Theory and Methods, 52(20): 7223–7251, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2022.2042027.

Liu:2023:VBF

[LL23d] Guoqing Liu and Wenbo V. Li. Variance bounds for functions of unimodal random variable. Communications in Statistics: Theory and Methods, 52(14):4765–4772, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2012.718842.

Li:2024:ODS

[LL24a] Jingwei Li and Guoxin Liu. Optimal dividend and stopping problems for two-dimensional compound Poisson risk model. Communications in Statistics: Theory and Methods, 53(12): 4515–4530, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2023.2184188.

Liu:2024:EUD

[LL24b] Yang Liu and Baoding Liu. Estimation of uncertainty distribution function by the principle of least squares. Communications in Statistics: Theory and Methods, 53(21):7624–7641, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2023.2269451.

Liu:2024:MUM

[LL24c] Yang Liu and Baoding Liu. A modified uncertain maximum likelihood estimation with applications in uncertain statistics. Communications in Statistics: Theory and Methods, 53(18): 6649–6670, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2023.2248534.

Lu:2024:HOE

[LLG24] Yingyin Lu, Xin Liao, and Jinhui Guo. Higher-order expansions of sample range from general error distribution. Communications in Statistics: Theory and Methods, 53(12):4498– 4514, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2023.2184187.

Liu:2024:LTD

[LLH^{+}24] Jinyang Liu, Sheng Li, Yong He, Boping Tian, and Li Deng. The Legendre transform-dual-asymptotic solution for optimal investment strategy with random incomes. Communications in Statistics: Theory and Methods, 53(23):8329–8347, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2281896.

Li:2022:OIS

[LLL22] Danping Li, Xiaotao Liu, and Hailong Liu. Optimal investment strategy for a family with a random household expenditure under the CEV model. Communications in Statistics: Theory and Methods, 51(17):5993–6007, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1851718.

Liu:2023:ODS

[LLL23] Shanshan Liu, Zhaoyang Liu, and Guoxin Liu. Optimal dividend strategy for the dual model with surplusdependent expense. Communications in Statistics: Theory and Methods, 52(3):543–566, 2023. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2021.1917614.

Li:2022:ELP

[LLQ22] Yinghua Li, Yuan Li, and Yongsong Qin. Empirical likelihood for panel data models with spatial errors. Communications in Statistics: Theory and Methods, 51(9):2838–2857, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2020.1780449.

Lu:2020:SIT

[LLS20] Dawei Lu, Qing Liu, and Xinmei Shen. Some improved tail bounds for the sum of variables with geometric distribution. Communications in Statistics: Theory and Methods, 49(22): 5427–5435, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2019.1618476.

Li:2022:COS

[LLSZ22] Min Li, Min-Qian Liu, Fasheng Sun, and Dong Zhang. Construction of optimal supersaturated designs via generalized Hadamard matrices. Communications in Statistics: Theory and Methods, 51(8):2565–2579, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1777309.

Leung:2023:RAL

[LLW23] Man Fung Leung, Yiqi Lin, and Nicolas Wicker. A remark about a learning risk lower bound. Communications in Statistics: Theory and Methods, 52(24):8785–8793, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2076114.

Luo:2022:AUR

[LLWA22] Jing Luo, Tour Liu, Jing Wu, and Sailan Waleed Ahmed Ali. Asymptotic in undirected random graph models with a noisy degree sequence. Communications in Statistics: Theory and Methods, 51(3):789–810, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1755870.

Li:2021:PCO

[LLWL21] Cuixiang Li, Huili Liu, Mengna Wang, and Wenhan Li. The pricing of compound option under variance gamma process by FFT. Communications in Statistics: Theory and Methods, 50(24):6122–6136, 2021. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2020.1740268.

Li:2021:ERF

[LLX21] Chuan-Quan Li, You-Wu Lin, and Qing-Song Xu. An enhanced random forest with canonical partial least squares for classification. Communications in Statistics: Theory and Methods, 50(18):4324–4334, 2021. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2020.1716249.

Li:2024:SDQ

[LLXW24] Linhong Li, Liwei Liu, Wei Xu, and Zhen Wang. Stochastic decomposition in a queueing-inventory system with batch demands, randomized order policy and multiple vacations. Communications in Statistics: Theory and Methods, 53(12):4407– 4427, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2023.2179886.

Li:2020:LOC

[LLZ20] Zhi-Ming Li, Ming-Ming Li, and Sheng-Li Zhao. Lowerorder confounding information of inverse Yates-order twolevel designs. Communications in Statistics: Theory and Methods, 49(4):924–941, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03610926.2018.1554124.

Li:2024:WEO

[LLZ24] Junpeng Li, Guanghui Li, and Chongqi Zhang. Weighted Refficiency optimal design for experiments with mixture. Communications in Statistics: Theory and Methods, 53(4):1241– 1256, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2022.2096901.

Luo:2020:NKE

[LM20] Zhongde Luo and Haizhen Meng. Nonparametric kernel estimation of expected shortfall under negatively associated sequences. Communications in Statistics: Theory and Methods, 49(11):2749–2769, 2020. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2019.1584303.

Lu:2022:CCI

[LM22] Dawei Lu and Yao Meng. Complete and complete integral convergence for arrays of row wise widely negative dependent random variables under the sub-linear expectations. Communications in Statistics: Theory and Methods, 51(9):2994–3007, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2020.1786585.

Lmoudden:2023:BEP

[LM23a] Aziz Lmoudden and Éric Marchand. Bayesian estimation and prediction for certain type of mixtures. Communications in Statistics: Theory and Methods, 52(2):309–334, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1913185.

Lu:2023:CCI

[LM23b] Dawei Lu and Yao Meng. Complete and complete integral convergence for weighted sums of arrays of rowwise widely negative dependent random variables under the sub-linear expectations. Communications in Statistics: Theory and Methods, 52(4):1273–1286, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1934027.

Li:2021:EVS

[LMH21] Rui Li, Shuchuan Mu, and Ruili Hao. Estimation and variable selection for partially linear additive models with measurement errors. Communications in Statistics: Theory and Methods, 50(6):1416–1445, 2021. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2019.1651858.

Li:2024:ABA

[LMH24] Xin Li, Mingzhi Mao, and Gang Huang. The asymptotic behaviors for autoregression quantile estimates. Communications in Statistics: Theory and Methods, 53(15):5486–5506, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2023.2221357.

Liu:2023:UIC

[LMLS23] Guanfu Liu, Rongji Mu, Yang Liu, and Zhimei Sheng. Using an inequality constraint to increase the power of the homogeneity tests for a two-sample problem with a mixture structure. Communications in Statistics: Theory and Methods, 52(18):6475–6486, 2023. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2022.2031216.

Louati:2020:NTS

[LMM20] Mahdi Louati, Afif Masmoudi, and Farouk Mselmi. The normal tempered stable regression model. Communications in Statistics: Theory and Methods, 49(2):500–512, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1554121.

Loubes:2020:RCC

[LMS20] Jean-Michel Loubes, Clément Marteau, and Maikol Solís. Rates of convergence in conditional covariance matrix with nonparametric entries estimation. Communications in Statistics: Theory and Methods, 49(18):4536–4558, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1602652.

Lyu:2022:UOP

[LMS22] Jianping Lyu, Yong Ma, and Wei Sun. A unified option pricing model with Markov regime-switching double stochastic volatility, stochastic interest rate and jumps. Communications in Statistics: Theory and Methods, 51(15):5112–5123, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2020.1833221.

Luo:2023:NAD

[LMZ23] Jing Luo, Xiaohui Ma, and Lewei Zhou. A note on asymptotic distributions in a directed network model with degree heterogeneity and homophily. Communications in Statistics: Theory and Methods, 52(16):5703–5715, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.2016835.

Lorvand:2020:LDP

[LNP20] Hamed Lorvand, Alireza Nematollahi, and Mohammad Hossien Poursaeed. Life distribution properties of a new δshock model. Communications in Statistics: Theory and Methods, 49(12):3010–3025, 2020. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2019.1584316.

Liebenberg:2022:NGF

[LNWA22] Shawn Carl Liebenberg, Joseph Ngatchou-Wandji, and James Samuel Allison. On a new goodness-of-fit test for the Rayleigh distribution based on a conditional expectation characterization. Communications in Statistics: Theory and Methods, 51(15):5226–5240, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1836220.

Li:2024:RMW

[LNXC24] Yan Li, Jiaqi Niu, Mengxue Xing, and Jinzhi Chen. Reliability modeling of weighted-k-out-of-n: G system under multiple failure modes with dependent components. Communications in Statistics: Theory and Methods, 53(13):4880–4897, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2196594.

Lyhagen:2023:RPC

[LO23] Johan Lyhagen and Petra Ornstein. Robust polychoric correlation. Communications in Statistics: Theory and Methods, 52(10):3241–3261, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03610926.2021.1970770.

Loecher:2022:UVI

[Loe22] Markus Loecher. Unbiased variable importance for random forests. Communications in Statistics: Theory and Methods, 51(5):1413–1425, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03610926.2020.1764042.

Liu:2021:URA

[LOL21] Jiaqi Liu, Zujun Ou, and Hongyi Li. Uniform row augmented designs with multi-level. Communications in Statistics: Theory and Methods, 50(15):3491–3504, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1705979.

Liew:2022:PSH

[LOT22] Kian Wah Liew, Seng Huat Ong, and Kian Kok Toh. The Poisson-stopped Hurwitz–Lerch zeta distribution. Communications in Statistics: Theory and Methods, 51(16):5638–5652, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2020.1844238.

Loukissas:2020:UAB

[Lou20] Fotios Loukissas. Uniform asymptotic behavior of tail probability of maxima in a time-dependent renewal risk model. Communications in Statistics: Theory and Methods, 49(24): 6112–6120, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2019.1626429.

Loumponias:2023:CTK

[Lou23] Kostas Loumponias. Colored Tobit Kalman filter. Communications in Statistics: Theory and Methods, 52(5):1533–1547, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2021.1929321.

Leonenko:2020:CPC

[LP20] Nikolai N. Leonenko and Ivan Papić. Correlation properties of continuous-time autoregressive processes delayed by the inverse of the stable subordinator. Communications in Statistics: Theory and Methods, 49(20):5091–5113, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1612918.

Li:2023:LEN

[LPFL23] Yongming Li, Weicai Pang, Ziqing Feng, and Naiyi Li. On the linearly extended negative quadrant dependent random variables and its inequalities. Communications in Statistics: Theory and Methods, 52(24):8696–8711, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2068600.

Li:2021:ELM

[LQ21] Yinghua Li and Yongsong Qin. Empirical likelihood for moving average models. Communications in Statistics: Theory and Methods, 50(15):3661–3676, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1710196.

Luo:2021:NAD

[LQWW21] Jing Luo, Hong Qin, Weifeng Wang, and Jun Wang. A note on asymptotic distributions in a network model with degree heterogeneity and homophily. Communications in Statistics: Theory and Methods, 50(21):5022–5033, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1645852.

Li:2022:EEL

[LRWZ22] Danping Li, Ximin Rong, Yajie Wang, and Hui Zhao. Equilibrium excess-of-loss reinsurance and investment strategies for an insurer and a reinsurer. Communications in Statistics: Theory and Methods, 51(21):7496–7527, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1873379.

Li:2020:SPI

[LS20a] Yuying Li and Kristina P. Sendova. A surplus process involving a compound Poisson counting process and applications. Communications in Statistics: Theory and Methods, 49(13): 3238–3256, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2019.1586942.

Longo:2020:SOI

[LS20b] Michele Longo and Gabriele Stabile. Sub-optimal investment for insurers. Communications in Statistics: Theory and Methods, 49(17):4298–4312, 2020. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2019.1599020.

Lawson:2022:NGR

[LS22a] Nuanpan Lawson and Pachitjanut Siripanich. A new generalized regression estimator and variance estimation for unequal probability sampling without replacement for missing data. Communications in Statistics: Theory and Methods, 51(18): 6296–6318, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2020.1860224.

Lee:2022:ESA

[LS22b] Gi-Sung Lee and Chang-Kyoon Son. An estimation of a sensitive attribute using adjusted kuk’s randomization device with stratified unequal probability sampling. Communications in Statistics: Theory and Methods, 51(1):1–25, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1821890.

Liu:2023:ABE

[LS23] Xiangdong Liu and Mingyun Sun. Adaptive bridge estimator for Cox model with a diverging number of parameters. Communications in Statistics: Theory and Methods, 52(7):2291– 2308, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2021.1951292.

Li:2022:FRG

[LSP22] Mingming Li, Huafei Sun, and Linyu Peng. Fisher–Rao geometry and Jeffreys prior for Pareto distribution. Communications in Statistics: Theory and Methods, 51(6):1895–1910, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2020.1771593.

Liu:2022:CSK

[LSSW22] Naiqi Liu, Kunyang Song, Yuping Song, and Xiaochen Wang. Convoluted smoothed kernel estimation for drift coefficients in jump-diffusion models. Communications in Statistics: Theory and Methods, 51(21):7354–7389, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1872641.

Liu:2022:RBE

[LSY22] Wanrong Liu, Jun Sun, and Jing Yang. Rank-based estimation in varying coefficient partially functional linear regression models. Communications in Statistics: Theory and Methods, 51(1):212–225, 2022. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2020.1747079.

Lian:2024:NMT

[LT24] Lian Lian and Zhongda Tian. A novel multivariate time series combination prediction model. Communications in Statistics: Theory and Methods, 53(7):2253–2284, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2124522.

Liang:2024:NPM

[LTR24] Jinwen Liang, Maozai Tian, and Yaohua Rong. Non parametric maximin aggregation for data with inhomogeneity. Communications in Statistics: Theory and Methods, 53(22):8109– 8126, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2023.2279913.

Li:2024:NIC

[LTTL24] Fan Li, Zizhong Tian, Zibo Tian, and Fan Li. A note on identification of causal effects in cluster randomized trials with post-randomization selection bias. Communications in Statistics: Theory and Methods, 53(5):1825–1837, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2116281.

Li:2021:TEC

[LTW21] Qian Li, Xiangyong Tan, and Liming Wang. Testing for error correlation in partially functional linear regression models. Communications in Statistics: Theory and Methods, 50(3): 747–761, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2019.1642492.

Li:2024:ECR

[LTW24] Yingqiu Li, Xinping Tang, and Hesong Wang. Exact convergence rate in central limit theorem for a supercritical branching process with immigration in a random environment. Communications in Statistics: Theory and Methods, 53(23):8412– 8427, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2023.2288792.

Liu:2023:SIP

[LTY23] Ming Liu, Yongge Tian, and Ruixia Yuan. Statistical inference of a partitioned linear random-effects model. Communications in Statistics: Theory and Methods, 52(4):1251–1272, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2021.1926509.

Lim:2023:FCE

[LVB23] Munwon Lim, Brani Vidakovic, and Suk Joo Bae. Fault classification via energy based features of two-dimensional image data. Communications in Statistics: Theory and Methods, 52(11):3939–3959, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03610926.2021.1982986.

Li:2020:CIC

[LW20] Jie Li and Qunying Wu. Complete integral convergence for arrays of row-wise extended independent random variables under sub-linear expectations. Communications in Statistics: Theory and Methods, 49(22):5613–5626, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1620954.

Lu:2021:CCC

[LW21a] Dawei Lu and Jialu Wang. Complete convergence and complete moment convergence for maximal randomly weighted sums of widely orthant-dependent random variables with applications. Communications in Statistics: Theory and Methods, 50(4):763–791, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03610926.2019.1640879.

Luo:2021:VSS

[LW21b] Guowang Luo and Mixia Wu. Variable selection for semiparametric varying-coefficient spatial autoregressive models with a diverging number of parameters. Communications in Statistics: Theory and Methods, 50(9):2062–2079, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1659367.

Liu:2022:HWM

[LW22a] Lianyi Liu and Lifeng Wu. Holt-winters model with grey generating operator and its application. Communications in Statistics: Theory and Methods, 51(11):3542–3555, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1797804.

Lu:2022:CMC

[LW22b] Dawei Lu and Lina Wang. Complete moment convergence for the widely orthant dependent linear processes with random coefficients. Communications in Statistics: Theory and Methods, 51(3):811–825, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1756328.

Lu:2022:FOI

[LW22c] Feilong Lu and Dehui Wang. First-order integer-valued autoregressive process with Markov-switching coefficients. Communications in Statistics: Theory and Methods, 51(13):4313– 4329, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2020.1813302.

Liu:2023:OPM

[LW23] Peng Liu and Guanjun Wang. Optimal preventive maintenance policies for products with multiple failure modes after geometric warranty expiry. Communications in Statistics: Theory and Methods, 52(24):8794–8813, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2076115.

Liu:2021:MLE

[LWD^{+}21] Xiufang Liu, Dehui Wang, Dianliang Deng, Jianhua Cheng, and Feilong Lu. Maximum likelihood estimation of the DDRCINAR(p) model. Communications in Statistics: Theory and Methods, 50(24):6231–6255, 2021. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2020.1741627.

Lin:2020:FAW

[LWK20] Hung-Mo Lin, John M. Williamson, and Hae-Young Kim. Firth adjustment for Weibull current-status survival analysis. Communications in Statistics: Theory and Methods, 49(18): 4587–4602, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2019.1606241.

Liu:2020:MWS

[LWKQ20] Baoliang Liu, Yanqing Wen, Shugui Kang, and Qingan Qiu. A multiple warm standby repairable system under N-policy with multiple vacations following Markovian arrival process. Communications in Statistics: Theory and Methods, 49(15):3609– 3634, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2019.1710758.

Lou:2024:NAS

[LWS24] Yichen Lou, Peijie Wang, and Jianguo Sun. A new approach for semi-parametric regression analysis of bivariate intervalcensored outcomes from case-cohort studies. Communications in Statistics: Theory and Methods, 53(15):5405–5420, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2220850.

Li:2024:NPE

[LWT24] Yue Li, Yunyan Wang, and Mingtian Tang. Non parametric estimation of transition density for second-order diffusion processes. Communications in Statistics: Theory and Methods, 53(16):5840–5852, 2024. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2023.2234521.

Ling:2024:OSP

[LWW24] Xiaoliang Ling, Meng Wang, and Yinzhao Wei. An optimal screening policy for heterogeneous items with minimal repairs. Communications in Statistics: Theory and Methods, 53(16): 5793–5813, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2023.2233150.

Li:2023:NBA

[LWY^{+}23] Han Li, Haoyu Wang, Kai Yang, Jie Sun, and Yan Liu. A nonparametric Bayesian analysis for meningococcal disease counts based on integer-valued threshold time series models. Communications in Statistics: Theory and Methods, 52(22):8223– 8240, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2022.2059683.

Liu:2024:SEL

[LWYJ24] Rong Liu, Chunjie Wang, Yujing Yao, and Zhezhen Jin. Smoothed empirical likelihood for optimal cut point analysis. Communications in Statistics: Theory and Methods, 53(17): 6299–6314, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2023.2244096.

Liu:2023:LBA

[LWYL23] Jiaqi Liu, Kang Wang, Di Yuan, and Jianjun Li. Lower bounds of the average mixture discrepancy for row augmented designs with mixed four- and five-level. Communications in Statistics: Theory and Methods, 52(19):6716–6733, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2032752.

Li:2021:ATC

[LX21] Yan Ling Li and Gen Qi Xu. Analysis of two components parallel repairable system with vacation. Communications in Statistics: Theory and Methods, 50(10):2429–2450, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1670847.

Liu:2022:SFS

[LX22a] Jinshan Liu and Qiang Xia. Some finite sample results for a system of seemingly unrelated regression equations. Communications in Statistics: Theory and Methods, 51(11):3629– 3644, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2020.1800738.

Liu:2022:NMF

[LX22b] Zili Liu and Zikang Xiong. Non-marginal feature screening for additive hazard model with ultrahigh-dimensional covariates. Communications in Statistics: Theory and Methods, 51(6): 1876–1894, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2020.1770288.

Long:2022:CFP

[LX22c] Yuqi Long and Xingzhong Xu. Classification by fiducial predictive density functions. Communications in Statistics: Theory and Methods, 51(15):5187–5203, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1836218.

Liu:2023:RE

[LX23a] Jiali Liu and Fuyuan Xiao. Renyi extropy. Communications in Statistics: Theory and Methods, 52(16):5836–5847, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.2020843.

Liu:2023:GBA

[LX23b] Xiaomei Liu and Naiming Xie. Grey-based approach for estimating Weibull model and its application. Communications in Statistics: Theory and Methods, 52(21):7601–7617, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2050397.

Liu:2023:SAR

[LXLL23] Jiamin Liu, Wangli Xu, Hongmei Lin, and Heng Lian. Sketched approximation of regularized canonical correlation analysis. Communications in Statistics: Theory and Methods, 52(19):6960–6971, 2023. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2022.2037644.

Liao:2020:JDE

[LXW20] Xin Liao, Qian Xiong, and Zhichao Weng. Joint distributional expansions of maxima and minima from skew-normal samples.

Communications in Statistics: Theory and Methods, 49(24): 5930–5947, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2019.1623256.

Liu:2021:EAF

[LXWY21] Jinping Liu, Xiuli Xu, Shuo Wang, and Dequan Yue. Equilibrium analysis of the fluid model with two types of parallel customers and breakdowns. Communications in Statistics: Theory and Methods, 50(24):5792–5805, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1737124.

Li:2020:FSS

[LXZ20] Yuying Li, Xiaozhu Xu, and Haifeng Zhang. The further study of semimodules over commutative semirings. Communications in Statistics: Theory and Methods, 49(20):4929–4950, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1609516.

Li:2022:ELB

[LXZ22] Wanbin Li, Liugen Xue, and Peixin Zhao. Empirical likelihood based inference for varying coefficient panel data models with fixed effect. Communications in Statistics: Theory and Methods, 51(14):4973–4990, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1828924.

Li:2021:SFP

[LY21] Yunxia Li and Caiyun Ying. Semi-functional partial linear spatial autoregressive model. Communications in Statistics: Theory and Methods, 50(24):5941–5954, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1738485.

Linlin:2022:SIP

[LY22a] Gui Linlin and Liu Yang. Statistical inference for the partially linear single-index model of panel data with serially correlated error structure. Communications in Statistics: Theory and Methods, 51(18):6334–6366, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1860226.

Liu:2022:AGG

[LY22b] Gang Liu and Hong Yin. Admissibility in general Gauss– Markov model with respect to an ellipsoidal constraint under weighted balanced loss. Communications in Statistics: Theory and Methods, 51(4):1054–1066, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1758140.

Lu:2023:IDL

[LY23] Dawei Lu and Jingcai Yang. Identifying the distribution of linear combinations of gamma random variables via Stein’s method. Communications in Statistics: Theory and Methods, 52(7):2116–2123, 2023. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2021.1945104.

Leng:2024:BOD

[LY24] Wei Leng and Juliang Yin. Bayesian Φ_{q}-optimal designs for multi-factor additive non linear models with heteroscedastic errors. Communications in Statistics: Theory and Methods, 53(23):8428–8440, 2024. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2023.2288805.

Li:2020:DTR

[LYC20] Jiahui Li, Kam Chuen Yuen, and Mi Chen. A discrete-time risk model with Poisson ARCH claim-number process. Communications in Statistics: Theory and Methods, 49(16):3965– 3984, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2019.1594296.

Liu:2024:POT

[LYD24] Zaiming Liu, Xiaofeng Yang, and Hua Dong. Poissonian occupation times of refracted Lévy processes with applications. Communications in Statistics: Theory and Methods, 53(21):

7659–7677, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2023.2271589.

Li:2022:NER

[LYL22] Jing Li, Wenguang Yu, and Chaolin Liu. Nonparametric estimation of ruin probability by complex Fourier series expansion in the compound Poisson model. Communications in Statistics: Theory and Methods, 51(15):5048–5063, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1831542.

Liu:2021:NMV

[LYLY21] Yue Liu, Aijun Yang, Jinguan Lin, and Jingjing Yao. A new method of valuing American options based on Brownian models. Communications in Statistics: Theory and Methods, 50(20):4809–4821, 2021. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2020.1725053.

Liu:2024:AGM

[LYY24] Yufei Liu, Qingqing Ye, and Junnai Yan. Analysis of a GI^{X} /M/1 queue with two-stage vacation policy using shift operator method. Communications in Statistics: Theory and Methods, 53(19):6732–6761, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2250488.

Lv:2023:CCW

[LYZ^{+}23] Yutan Lv, Yunbao Yao, Jun Zhou, Xiaoqin Li, Ruiqi Yang, and Xuejun Wang. Complete consistency for the weighted least squares estimators in semiparametric regression models. Communications in Statistics: Theory and Methods, 52(22): 7797–7818, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2022.2050400.

Li:2020:KME

[LZ20a] Yongming Li and Yong Zhou. The Kaplan–Meier estimator and hazard estimator for censored END survival time observations. Communications in Statistics: Theory and Methods, 49(11):2690–2702, 2020. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2019.1580737.

Liu:2020:LIL

[LZ20b] Tianze Liu and Yong Zhang. Law of the iterated logarithm for error density estimators in nonlinear autoregressive models. Communications in Statistics: Theory and Methods, 49(5): 1082–1098, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2018.1554129.

Li:2021:NAA

[LZ21] Huajiang Li and Hong Zhou. A new approach to address multiplicity in hypothesis testing with constraints. Communications in Statistics: Theory and Methods, 50(1):35–60, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1628989.

Li:2023:ODM

[LZ23a] Chang Li and Chongqi Zhang. D-optimal designs for mixture experiments with various correlation structures. Communications in Statistics: Theory and Methods, 52(24):8828–8843, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2022.2076117.

Liu:2023:LDP

[LZ23b] Wei Liu and Yong Zhang. Large deviation principle for linear processes generated by real stationary sequences under the sub-linear expectation. Communications in Statistics: Theory and Methods, 52(16):5727–5741, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.2018462.

Luo:2023:NAR

[LZ23c] Lin Luo and Hui Zhao. A new approach to regression analysis of linear transformation model with interval-censored data. Communications in Statistics: Theory and Methods, 52(15): 5470–5482, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2021.2012195.

Luo:2024:ODM

[LZ24] Jiacheng Luo and Chongqi Zhang. R-optimal designs for mixture models in two types of regions. Communications in Statistics: Theory and Methods, 53(5):1851–1863, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2116283.

Lian:2020:PSK

[LZG20a] Ruiting Lian, Changle Zhou, and Ben Goertzel. The probabilistic support Kendall correlation and its transitivity properties. Communications in Statistics: Theory and Methods, 49(2):485–499, 2020. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2018.1543776.

Liang:2020:ANC

[LZG20b] Han-Ying Liang, Hong-Bing Zhou, and Qiu-Li Guo. Asymptotic normality of conditional density estimation under truncated, censored and dependent data. Communications in Statistics: Theory and Methods, 49(22):5371–5391, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1619769.

Li:2020:SCP

[LZH20] Bo Li, Huiming Zhang, and Jiao He. Some characterizations and properties of COM–Poisson random variables. Communications in Statistics: Theory and Methods, 49(6):1311–1329, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2018.1563164.

Liu:2020:OIP

[LZL20] Bing Liu, Ming Zhou, and Peng Li. Optimal investment and premium control for insurers with ambiguity. Communications in Statistics: Theory and Methods, 49(9):2110–2130, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1568487.

Liu:2020:AEC

[LZLL20] Yang Liu, Lin Zhu, Guanfu Liu, and Huapeng Li. Abundance estimation with a categorical covariate subject to missing in continuous-time capture-recapture studies. Communications in Statistics: Theory and Methods, 49(20):4919–4928, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1609039.

Liu:2023:BEL

[LZW23] Binxia Liu, Hui Zhao, and Chunjie Wang. Bayesian empirical likelihood of linear regression model with current status data. Communications in Statistics: Theory and Methods, 52(20): 7323–7333, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2022.2044491.

Ling:2024:SCL

[LZW24] Xiaoliang Ling, Jiaojiao Zhang, and Yinzhao Wei. Stochastic comparisons of linear degradation and failure time models with dependent failure modes. Communications in Statistics: Theory and Methods, 53(1):346–364, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2077965.

Mansourvar:2023:EHR

[MA23] Zahra Mansourvar and Majid Asadi. On the estimation of hazard rate in mixed populations with its application. Communications in Statistics: Theory and Methods, 52(21):7564–7575, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2022.2048858.

Maatouk:2023:FDA

[Maa23] H. Maatouk. Finite-dimensional approximation of Gaussian processes with linear inequality constraints and noisy observations. Communications in Statistics: Theory and Methods, 52(22):8018–8037, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03610926.2022.2055768.

Mohammed:2022:IFT

[MAAZ22] M. B. Mohammed, M. B. Adam, N. Ali, and H. S. Zulkafli. Improved frequency table’s measures of skewness and kurtosis with application to weather data. Communications in Statistics: Theory and Methods, 51(3):581–598, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1752386.

MartinAndres:2022:OTA

[MÁH22] Antonio Martín Andrés, Maria Álvarez Hernández, and Inmaculada Herranz Tejedor. One-tailed asymptotic inferences for the relative risk: a comparison of 63 inference methods. Communications in Statistics: Theory and Methods, 51(5): 1330–1348, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2020.1760299.

Mahmood:2024:RCC

[Mah24]  Ehab A. Mahmood.  Robust circular-circular correla-

tion coefficient. Communications in Statistics: Theory and Methods, 53(6):2034–2042, 2024. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2022.2117561.

Matula:2020:ESL

[MAK20] Przemysław Matuła, André Adler, and Paweł Kurasiński. On exact strong laws of large numbers for ratios of random variables and their applications. Communications in Statistics: Theory and Methods, 49(13):3153–3167, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1586935.

Makigusa:2024:TST

[Mak24] N. Makigusa. Two-sample test based on maximum variance discrepancy. Communications in Statistics: Theory and Methods, 53(15):5421–5438, 2024. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2023.2220851.

Mao:2020:HDT

[Mao20] Guangyu Mao. On high-dimensional tests for mutual independence based on pearson’s correlation coefficient. Communications in Statistics: Theory and Methods, 49(14):3572–3584, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2019.1593459.

Matsui:2021:AML

[Mat21] Muneya Matsui. Asymptotics of maximum likelihood estimation for stable law with continuous parameterization. Communications in Statistics: Theory and Methods, 50(15):3695– 3712, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2019.1710199.

Matsuura:2023:BEP

[Mat23] Shun Matsuura. Bayes estimator of process capability index C_{pk} with a specified prior mean. Communications in Statistics: Theory and Methods, 52(7):2215–2227, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1947508.

Mansouri:2024:RAO

[MAT24a] Akbar Mansouri and Akbar Alem-Tabriz. Redundancy allocation optimizing in the satellite attitude determination and control system based on the exact solution algorithm. Communications in Statistics: Theory and Methods, 53(4):1516–1534, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2022.2104873.

Mathew:2024:QCD

[Mat24b] Angel Mathew. Quantile cumulative distribution function and its applications. Communications in Statistics: Theory and Methods, 53(11):4194–4206, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2176716.

Mohebbi:2024:MSS

[MAT24c] Mahboubeh Mohebbi, Amirhossein Amiri, and Ali Reza Taheriyoun. Monitoring the structure of social networks based on exponential random graph model. Communications in Statistics: Theory and Methods, 53(10):3742–3757, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2163366.

Meng:2020:DRM

[MB20] Qingbin Meng and Junna Bi. On the dividends of the risk model with Markovian barrier. Communications in Statistics: Theory and Methods, 49(5):1272–1280, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1563175.

Maqsood:2021:ESL

[MB21a] Arfa Maqsood and S. M. Aqil Burney. Extended stochastic Laspeyres price index model with autocorrelated errors. Communications in Statistics: Theory and Methods, 50(14):3352– 3357, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2019.1702692.

Merahi:2021:ETF

[MB21b] Fateh Merahi and Abdelouahab Bibi. Evolutionary transfer functions solution for continuous-time bilinear stochastic processes with time-varying coefficients. Communications in Statistics: Theory and Methods, 50(22):5189–5214, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1726390.

Mohammed:2020:VFF

[MBA20] Siti Aisyah Mohammed, Mohd Aftar Abu Bakar, and Noratiqah Mohd Ariff. Volatility forecasting of financial time series using wavelet based exponential generalized autoregressive conditional heteroscedasticity model. Communications in Statistics: Theory and Methods, 49(1):178–188, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1535073.

Mohammedi:2024:ANN

[MBLB24] Mustapha Mohammedi, Salim Bouzebda, Ali Laksaci, and Oussama Bouanani. Asymptotic normality of the k-NN single index regression estimator for functional weak dependence data*. Communications in Statistics: Theory and Methods, 53(9):3143–3168, 2024. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2022.2150823.

Ma:2021:LBD

[MBLY21] Ming Ma, Li Na Bian, Hua Liu, and Jian Hua Ye. Lifetime behavior of discrete Markov chain censored δ shock model. Communications in Statistics: Theory and Methods, 50(5):1019– 1035, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2019.1650184.

Maadani:2022:VOS

[MBR22] S. Maadani, G. R. Mohtashami Borzadaran, and A. H. Rezaei Roknabadi. Varentropy of order statistics and some stochastic comparisons. Communications in Statistics: Theory and Methods, 51(18):6447–6460, 2022. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2020.1861299.

Mutis:2023:RSF

[MBSS23] Muge Mutis, Ufuk Beyaztas, Gulhayat Golbasi Simsek, and Han Lin Shang. A robust scalar-on-function logistic regression for classification. Communications in Statistics: Theory and Methods, 52(23):8538–8554, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2065018.

Mukherjee:2020:WDL

[MC20] Diganta Mukherjee and Prabir Chaudhury. Why does “last week” reporting give higher estimates than “last month”? Communications in Statistics: Theory and Methods, 49(8): 1873–1893, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2019.1565838.

Martinez-Camblor:2023:AUO

[MC23] Pablo Martínez-Camblor. About the use of the overlap coefficient in the binary classification context. Communications in Statistics: Theory and Methods, 52(19):6767–6777, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2032754.

McIntosh:2023:CSS

[McI23] Matthew J. McIntosh. Conservative sample size for multiple regression models. Communications in Statistics: Theory and Methods, 52(15):5527–5533, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.2012198.

Ma:2023:OVU

[MCL23] Yong Ma, Li Chen, and Jianping Lyu. Option valuation under double exponential jump with stochastic intensity, stochastic interest rates and Markov regime-switching stochastic volatility. Communications in Statistics: Theory and Methods, 52(7):2043–2056, 2023. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2021.1944214.

Mukhopadhyay:2021:DGM

[MDB^{+}21] Soumalya Mukhopadhyay, Amlan Jyoti Das, Ayanendranath Basu, Aditya Chatterjee, and Sabyasachi Bhattacharya. Does the generalized mean have the potential to control outliers? Communications in Statistics: Theory and Methods, 50(8): 1709–1727, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2019.1652320.

Minkah:2023:REP

[MdWG23] Richard Minkah, Tertius de Wet, and Abhik Ghosh. Robust estimation of Pareto-type tail index through an exponential regression model. Communications in Statistics: Theory and Methods, 52(2):479–498, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1916530.

Mizutani:2020:PRP

[MDZN20] Satoshi Mizutani, Wenjie Dong, Xufeng Zhao, and Toshio Nakagawa. Preventive replacement policies with products update announcements. Communications in Statistics: Theory and Methods, 49(15):3821–3833, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1710762.

Marzougue:2021:RBJ

[ME21] Mohamed Marzougue and Mohamed El Otmani. Reflected BSDEs with jumps and two rcll barriers under stochastic Lipschitz coefficient. Communications in Statistics: Theory and Methods, 50(24):6049–6066, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1738491.

Mefleh:2023:SCP

[Mef23] Aline Mefleh. Simple change point model in heteroscedastic extremes. Communications in Statistics: Theory and Methods, 52(20):7455–7464, 2023. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2022.2046089.

Mahadik:2023:SSC

[MG23] Shashibhushan B. Mahadik and Dadasaheb G. Godase. The SPRT sign chart for process location. Communications in Statistics: Theory and Methods, 52(7):2276–2290, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1949474.

Michael:2024:CBT

[MG24] Haben Michael and Musie Ghebremichael. A correction to Begg’s test for publication bias. Communications in Statistics: Theory and Methods, 53(21):7678–7698, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2271590.

Misra:2022:ISP

[MGC22] Amit Kumar Misra, Vaishali Gupta, and Ruby Chanchal. Inactivity stochastic precedence order. Communications in Statistics: Theory and Methods, 51(8):2590–2609, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1777430.

Mabitsela:2022:NRB

[MGK22] Lesedi Mabitsela, Calisto Guambe, and Rodwell Kufakunesu. A note on representation of BSDE-based dynamic risk measures and dynamic capital allocations. Communications in Statistics: Theory and Methods, 51(6):1791–1810, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1768405.

Martin-Gonzalez:2023:EDP

[MGK23] Ehyter M. Martín-González and Ekaterina T. Kolkovska. Expected discounted penalty function and asymptotic dependence of the severity of ruin and surplus prior to ruin for two-sided Lévy risk processes. Communications in Statistics: Theory and Methods, 52(23):8566–8583, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2065302.

Musleh:2022:KBE

[MHS22] Rola Musleh, Amal Helu, and Hani Samawi. Kernel-based estimation of P (X < Y) when X and Y are dependent random variables based on progressive type II censoring. Communications in Statistics: Theory and Methods, 51(8):2368–2384, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2020.1774058.

Meng:2023:EFP

[MHZJ23] Shuyu Meng, Zhensheng Huang, Jing Zhang, and Zhiqiang Jiang. Estimation on functional partially linear single index measurement error model. Communications in Statistics: Theory and Methods, 52(13):4741–4763, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1999979.

Mirakhmedov:2023:APC

[Mir23] Sherzod M. Mirakhmedov. On the asymptotic properties of a certain class of goodness-of-fit tests associated with multinomial distributions. Communications in Statistics: Theory and Methods, 52(3):602–622, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1918168.

Mohri:2024:GRE

[MiT24] Hiroaki Mohri and Jun ichi Takeshita. Graph reliability evaluation via random K-out-of-N systems. Communications in Statistics: Theory and Methods, 53(3):1024–1034, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2099897.

Murugeswari:2024:EOD

[MJB24] N. Murugeswari, P. Jeyadurga, and S. Balamurali. Evaluation and optimal designing of a two-level skip-lot sampling plan for resubmitted lots. Communications in Statistics: Theory and Methods, 53(6):2043–2066, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2117986.

Mondal:2020:JTI

[MK20] Shuvashree Mondal and Debasis Kundu. On the joint Type-II progressive censoring scheme. Communications in Statistics: Theory and Methods, 49(4):958–976, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1554128.

Mashwani:2021:HDE

[MKG^{+}21] Wali Khan Mashwani, Abdullah Khan, Atila Göktas, Yuksel Akay Unvan, Ozgur Yaniay, and Abdelouahed Hamdi. Hybrid differential evolutionary strawberry algorithm for realparameter optimization problems. Communications in Statistics: Theory and Methods, 50(7):1685–1698, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1783559.

Ma:2024:MAD

[MKM24] Yuyan Ma, Masato Kitani, and Hidetoshi Murakami. On modified Anderson–Darling test statistics with asymptotic properties. Communications in Statistics: Theory and Methods, 53(4):1420–1439, 2024. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2022.2101121.

Mirezi:2023:MMV

[MKÖ23] Buatikan Mirezi, Selahattin Kaçiranlar, and Nimet Özbay. A minimum matrix valued risk estimator combining restricted and ordinary least squares estimators. Communications in Statistics: Theory and Methods, 52(5):1580–1590, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1934032.

Medeiros:2021:LBI

[ML21] Francisco M. C. Medeiros and Artur J. Lemonte. Likelihoodbased inference in censored exponential regression models. Communications in Statistics: Theory and Methods, 50(13): 3214–3233, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2019.1700281.

Midya:2023:SDM

[ML23] Vishal Midya and Jiangang Liao. Systematic deviation in mean of log Bayes factor: Implication and application. Communications in Statistics: Theory and Methods, 52(10):3209– 3218, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2021.1970768.

Mottonen:2024:RBM

[MLSS24] Jyrki Möttönen, Tero Lähderanta, Janne Salonen, and Mikko J. Sillanpää. Reducing bias and mitigating the influence of excess of zeros in regression covariates with multioutcome adaptive LAD-lasso. Communications in Statistics: Theory and Methods, 53(13):4730–4744, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2189059.

Ma:2024:MSP

[MLW24] Jinping Ma, Xiaoling Liu, and Huanle Wen. Method of solving the problem of non-response error in dual sampling frame. Communications in Statistics: Theory and Methods, 53(3): 953–964, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2022.2097695.

Masioti:2024:NSE

[MLWSPS24] Marina Masioti, Connie S. N. Li-Wai-Suen, Luke A. Prendergast, and Amanda Shaker. A note on switching eigenvalues under small perturbations. Communications in Statistics: Theory and Methods, 53(20):7311–7325, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2263114.

Mi:2021:OIP

[MLZ21] Hui Mi, Lei Li, and Quanxin Zhu. Optimal investment problem with complete memory on an infinite time horizon. Communications in Statistics: Theory and Methods, 50(3):711– 724, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2019.1640877.

Majumder:2020:CET

[MM20] Priyanka Majumder and Murari Mitra. On a class exhibiting trend change in average failure rate. Communications in Statistics: Theory and Methods, 49(8):1975–1988, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1568479.

Malela-Majika:2024:DID

[MMG24] Jean-Claude Malela-Majika and Marien A. Graham. Design and implementation of distribution-free phase-II charting schemes based on unconditional run-length percentiles. Communications in Statistics: Theory and Methods, 53(1):276– 293, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2022.2077961.

Malik:2020:CLV

[MMLBP20] Waqas Ahmed Malik, Carles Marco-Llorca, Kenneth Berendzen, and Hans-Peter Piepho. Choice of link and variance function for generalized linear mixed models: a case study with binomial response in proteomics. Communications in Statistics: Theory and Methods, 49(17):4313–4332, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1599021.

Malela-Majika:2020:DFP

[MMRMG20] J. C. Malela-Majika, E. M. Rapoo, A. Mukherjee, and M. A. Graham. Distribution-free precedence schemes with a generalized runs-rule for monitoring unknown location. Communications in Statistics: Theory and Methods, 49(20):4996–5027, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2019.1612914.

Mukherjee:2023:EPC

[MMS23] Indrani Mukherjee, Sudhansu S. Maiti, and Vijay Vir Singh. On estimation of the PDF and the CDF of the one-parameter polynomial exponential family of distributions. Communications in Statistics: Theory and Methods, 52(1):104–120, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1910302.

Malela-Majika:2022:OSP

[MMSC22] Jean-Claude Malela-Majika, Sandile C. Shongwe, and Philippe Castagliola. One-sided precedence monitoring schemes for unknown shift sizes using generalized 2-of -(h + 1) and w-of-w improved runs-rules. Communications in Statistics: Theory and Methods, 51(9):2803–2837, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1780448.

Miao:2020:LTI

[MMZ20] Yu Miao, Jianyong Mu, and Shuili Zhang. Limit theorems for identically distributed martingale difference. Communications in Statistics: Theory and Methods, 49(6):1435–1445, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2018.1563174.

Manandhar:2021:HBM

[MN21a] Binod Manandhar and Balgobin Nandram. Hierarchical Bayesian models for continuous and positively skewed data from small areas. Communications in Statistics: Theory and Methods, 50(4):944–962, 2021. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2019.1645853.

Marchand:2021:BME

[MN21b] Éric Marchand and Theodoros Nicoleris. On Bayes minimax estimators for a normal mean with an uncertain constraint^{+}. Communications in Statistics: Theory and Methods, 50(8): 1873–1883, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2019.1656251.

Mahlooji:2023:ESP

[MN23a] Negin Mahlooji and Nader Nematollahi. Estimation of the scale parameter of a selected gamma population with unequal shape parameters under Stein loss function. Communications in Statistics: Theory and Methods, 52(18):6573–6596, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2032167.

Mbaeyi:2023:DAM

[MN23b] George Chinanu Mbaeyi and Chijioke Joel Nweke. Discriminant analysis with mixed non normal variables. Communications in Statistics: Theory and Methods, 52(1):39–45, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1908563.

Mirniam:2023:AEB

[MN23c] A. S. Mirniam and A. R. Nematollahi. On accelerating the EM-based algorithms for the VAR(1) models with multivariate generalized scaled t-distributed innovations. Communications in Statistics: Theory and Methods, 52(13):4414–4428, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2021.1994608.

Moheghi:2022:PPI

[MNA22] H. R. Moheghi, R. Noorossana, and O. Ahmadi. Phase I and phase II analysis of linear profile monitoring using robust estimators. Communications in Statistics: Theory and Methods, 51(5):1252–1269, 2022. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2020.1758724.

Mahdavi-Nasab:2020:WCA

[MNAM20] Narges Mahdavi-Nasab, Mostafa Abouei Ardakan, and Mohammad Mohammadi. Water cycle algorithm for solving the reliability-redundancy allocation problem with a choice of redundancy strategies. Communications in Statistics: Theory and Methods, 49(11):2728–2748, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1580741.

Mun:2020:DRM

[MO20] Jungwon Mun and Minkyung Oh. Diagnostics for repeated measurements in nonlinear mixed effects models. Communications in Statistics: Theory and Methods, 49(20):5045–5059, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2019.1612916.

Moustaaid:2022:EDG

[MO22] Jabrane Moustaaid and Idir Ouassou. Estimation of the drift of a Gaussian process under balanced loss function. Communications in Statistics: Theory and Methods, 51(23):8225–8245, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2021.1890779.

Motalebi:2023:MSN

[MOAF23] Narges Motalebi, Mohammad Saleh Owlia, Amirhossein Amiri, and Mohammad Saber Fallahnezhad. Monitoring social networks based on zero-inflated Poisson regression model. Communications in Statistics: Theory and Methods, 52(7): 2099–2115, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2021.1945103.

Modak:2023:NMA

[Mod23] Soumita Modak. A new measure for assessment of clustering based on kernel density estimation. Communications in Statistics: Theory and Methods, 52(17):5942–5951, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2032168.

Mohamed:2022:CRT

[Moh22a] Mohamed Said Mohamed. On cumulative residual Tsallis entropy and its dynamic version of concomitants of generalized order statistics. Communications in Statistics: Theory and Methods, 51(8):2534–2551, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1777306.

Mohammadi:2022:THM

[Moh22b] Shapour Mohammadi. A test of harmful multicollinearity: a generalized ridge regression approach. Communications in Statistics: Theory and Methods, 51(3):724–743, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1754855.

Mohamed:2023:SPO

[Moh23] Zarai Mohamed. Some properties of 2-orthogonal polynomials associated with inverse Gaussian distribution. Communications in Statistics: Theory and Methods, 52(5):1348–1355, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2021.1926511.

Monnez:2024:SAE

[Mon24] Jean-Marie Monnez. Stochastic approximation of eigenvectors and eigenvalues of the Q-symmetric expectation of a random matrix. Communications in Statistics: Theory and Methods, 53(5):1669–1683, 2024. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2022.2107225.

Morimoto:2024:ESS

[Mor24] Julian Morimoto. The effect of sample size and missingness on inference with missing data. Communications in Statistics: Theory and Methods, 53(9):3292–3311, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2152287.

Motoyama:2023:RCA

[Mot23] Hitoshi Motoyama. Rate of convergence of the asymptotic normality of sample quantiles from a finite population. Communications in Statistics: Theory and Methods, 52(3):886–895, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2021.1931334.

Motoyama:2024:EGC

[Mot24] Hitoshi Motoyama. Extended Glivenko–Cantelli theorem for simple random sampling without replacement from a finite population. Communications in Statistics: Theory and Methods, 53(16):5924–5934, 2024. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2023.2238233.

Markiewicz:2021:LPS

[MP21a] Augustyn Markiewicz and Simo Puntanen. Linear prediction sufficiency in the misspecified linear model. Communications in Statistics: Theory and Methods, 50(21):4977–4996, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1584311.

Masirevic:2021:NFC

[MP21b] Dragana Jankov Masirević and Tibor K. Pogány. On new formulae for cumulative distribution function for McKay Bessel distribution. Communications in Statistics: Theory and Methods, 50(1):143–160, 2021. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2019.1632898.

Marcondes:2020:SHM

[MPM20] D. Marcondes, C. Peixoto, and A. C. Maia. A survey of a hurdle model for heavy-tailed data based on the generalized lambda distribution. Communications in Statistics: Theory and Methods, 49(4):781–808, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1549251.

Meitz:2023:MAM

[MPS23] Mika Meitz, Daniel Preve, and Pentti Saikkonen. A mixture autoregressive model based on Student’s t-distribution. Communications in Statistics: Theory and Methods, 52(2):499– 515, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2021.1916531.

Mahdizadeh:2023:ESP

[MQK23] Z. Mahdizadeh, M. Naghizadeh Qomi, and Shahjahan Khan. Estimation of the slope parameter in a linear regression model under a bounded loss function. Communications in Statistics: Theory and Methods, 52(16):5799–5813, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.2020292.

Mohammadi:2021:SMG

[MR21] Ali Mohammadi and Mohammad Reza Salehi Rad. Solution of an M/G/1 queueing model with k sequential heterogeneous service steps and vacations using the Tauberian property. Communications in Statistics: Theory and Methods, 50(14): 3235–3248, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2019.1691232.

Muhammad:2020:EAL

[MRD20] Faqir Muhammad, Muhammad Riaz, and Hassan Dawood. An effective approach to linear calibration estimation with its applications. Communications in Statistics: Theory and Methods, 49(21):5154–5174, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1615092.

Muhammad:2022:BLB

[MRDD22] Faqir Muhammad, Muhammad Riaz, Hassan Dawood, and Hussain Dawood. On best linear and Bayesian linear predictor in calibration. Communications in Statistics: Theory and Methods, 51(11):3669–3693, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1801733.

Maladan:2021:NFQ

[MS21] Dileep Kumar Maladan and Paduthol Godan Sankaran. A new family of quantile functions and its applications. Communications in Statistics: Theory and Methods, 50(18):4216–4235, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2020.1713368.

Mathai:2023:DCS

[MS23] A. M. Mathai and Nicy Sebastian. On distributions of covariance structures. Communications in Statistics: Theory and Methods, 52(20):7370–7384, 2023. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2022.2045022.

Maleki:2023:EGM

[MSEB23] Mohammad Reza Maleki, Bahareh Shamseddin, Hossein Eghbali, and Aliasghar Bazdar. The effect of gauge measurement errors on double sampling X̄ control chart. Communications in Statistics: Theory and Methods, 52(8):2702–2717, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1958848.

Ma:2020:SEC

[MSM20] Huanhuan Ma, Yan Sun, and Yu Miao. Some extensions of the classical law of large numbers. Communications in Statistics: Theory and Methods, 49(13):3228–3237, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1586941.

Ma:2023:RPC

[MSW23] Ming Ma, Ailing Shi, and Miaomiao Wang. Reliability of Poisson censored δ-shock model. Communications in Statistics: Theory and Methods, 52(23):8501–8514, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2064502.

Maleki:2023:MEC

[MSY23] Mohammad Reza Maleki, Ali Salmasnia, and Sahand Yousefi. Multivariate ELR control chart with estimated mean vector and covariance matrix. Communications in Statistics: Theory and Methods, 52(24):8814–8827, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2076116.

Mishra:2022:BAA

[MU22] Rahul Kumar Mishra and Satyanshu K. Upadhyay. Bayes analysis of abridged age specific fertility pattern using parametric models. Communications in Statistics: Theory and Methods, 51(16):5505–5533, 2022. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2020.1843052.

Mukerjee:2024:SAB

[Muk24] Rahul Mukerjee. A statistical approach to broken stick problems. Communications in Statistics: Theory and Methods, 53(15):5629–5637, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03610926.2023.2224909.

Murat:2022:HTA

[Mur22] Naci Murat. A hybrid transformation approach for common scaling on various type Likert scales in Bayesian structural equation modeling. Communications in Statistics: Theory and Methods, 51(5):1217–1231, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1853774.

Melo:2020:IEM

[MVLMA20] Tatiane F. N. Melo, Tiago M. Vargas, Artur J. Lemonte, and Germán Moreno-Arenas. On improved estimation in multivariate Dirichlet regressions. Communications in Statistics: Theory and Methods, 49(23):5765–5777, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1620955.

MacOduol:2020:AGS

[MvSK20] Brenda V. Mac’Oduol, Paul J. van Staden, and Robert A. R. King. Asymmetric generalizations of symmetric univariate probability distributions obtained through quantile splicing. Communications in Statistics: Theory and Methods, 49(18): 4413–4429, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2019.1601219.

Meng:2024:CAN

[MW24a] Bing Meng and Qunying Wu. Convergence of asymptotically negatively associated random variables with random coefficients. Communications in Statistics: Theory and Methods, 53(8):2961–2976, 2024. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2022.2150058.

Motegi:2024:NEA

[MW24b] Kaiji Motegi and Sejun Woo. A note on the exponentiation approximation of the birthday paradox. Communications in Statistics: Theory and Methods, 53(18):6417–6426, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2245086.

Ma:2024:ECC

[MWLS24] Yuqing Ma, Peijie Wang, Shuwei Li, and Jianguo Sun. Estimation of complier causal treatment effects under the additive hazards model with interval-censored data. Communications in Statistics: Theory and Methods, 53(10):3547–3567, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2155791.

Meng:2022:CCC

[MWW22] Bing Meng, Dingcheng Wang, and Qunying Wu. Complete convergence and complete moment convergence for weighted sums of extended negatively dependent random variables. Communications in Statistics: Theory and Methods, 51(12): 3847–3863, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2020.1804587.

Meng:2023:CAA

[MWW23] Bing Meng, Dingcheng Wang, and Qunying Wu. Convergence of asymptotically almost negatively associated random variables with random coefficients. Communications in Statistics: Theory and Methods, 52(9):2931–2945, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1963457.

Miao:2022:APL

[MZ22] Yu Miao and Yuhang Zhen. Asymptotic properties of LS estimator in nonlinear functional EV models. Communications in Statistics: Theory and Methods, 51(21):7575–7606, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1875242.

Mao:2023:CBH

[MZ23] Guangcai Mao and Jing Zhang. Confidence bands for hazard rate function based on a debiased estimation. Communications in Statistics: Theory and Methods, 52(15):5246–5259, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2021.2005098.

Ma:2020:OIS

[MZR20] Jie Ma, Hui Zhao, and Ximin Rong. Optimal investment strategy for a DC pension plan with mispricing under the Heston model. Communications in Statistics: Theory and Methods, 49(13):3168–3183, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1586938.

Nahla:2023:CIG

[NA23] Ben Salah Nahla and Masmoudi Afif. The confidence interval of q-Gaussian distributions. Communications in Statistics: Theory and Methods, 52(10):3511–3525, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1974482.

Naveed:2022:NSP

[NAAQ22] Khalid Naveed, Muhammad Amin, Saima Afzal, and Muhammad Qasim. New shrinkage parameters for the inverse Gaussian Liu regression. Communications in Statistics: Theory and Methods, 51(10):3216–3236, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1791339.

Nagakura:2020:FRV

[Nag20] Daisuke Nagakura. Further results on the vecd operator and its applications. Communications in Statistics: Theory and Methods, 49(10):2321–2338, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1570265.

Nahla:2022:GCP

[Nah22] Ben Salah Nahla. Generalized convolution product of an infinitely divisible distribution and a Bernoulli distribution. Communications in Statistics: Theory and Methods, 51(20): 7297–7304, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2021.1879863.

Najarzadeh:2021:TIH

[Naj21] D. Najarzadeh. Testing independence in high-dimensional multivariate normal data. Communications in Statistics: Theory and Methods, 50(14):3421–3435, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1702699.

Najari:2022:NSP

[NAJ22a] Nader Najari, Mohammad Q. Vahidi Asl, and Abdollah Jalilian. Neyman–Scott process with skew-normal clusters. Communications in Statistics: Theory and Methods, 51(14):4692– 4711, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2020.1819324.

Najarzadeh:2022:OPT

[Naj22b] D. Najarzadeh. An optimal projection test for zero multiple correlation coefficient in high-dimensional normal data. Communications in Statistics: Theory and Methods, 51(4):1011– 1028, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2020.1757111.

Nakashima:2021:EAF

[Nak21] Eiji Nakashima. Estimation of accelerated failure time due to exposure risk using piecewise-exponential model in cohort follow-up study. Communications in Statistics: Theory and Methods, 50(3):725–737, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1640878.

Naderi:2022:NWL

[NBM22] Habib Naderi, Fakhreddine Boukhari, and Przemysław Matuła. A note on the weak law of large numbers for weighted negatively superadditive dependent random variables. Communications in Statistics: Theory and Methods, 51(21):7465– 7475, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2021.1873377.

Nafii:2024:NPM

[NBM24] Adil Nafii, Taoufik Bouezmarni, and Mhamed Mesfioui. Non parametric multivariate distribution estimation under right censoring. Communications in Statistics: Theory and Methods, 53(19):6785–6798, 2024. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2023.2251624.

Ngailo:2023:AMP

[NC23] Edward Kanuti Ngailo and Furaha Chuma. Approximation of misclassification probabilities in linear discriminant analysis based on repeated measurements. Communications in Statistics: Theory and Methods, 52(23):8388–8407, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2062605.

Nie:2024:TCS

[NCC24] Gaoqin Nie, Xingjiang Chen, and Hao Chang. Timeconsistent strategies between two competitive DC pension plans with the return of premiums clauses and salary risk. Communications in Statistics: Theory and Methods, 53(22): 7807–7828, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2023.2273207.

Nanda:2024:NBD

[NCGB24] Asok K. Nanda, Shovan Chowdhury, Sanjib Gayen, and Subarna Bhattacharjee. A new bivariate distribution with uniform marginals. Communications in Statistics: Theory and Methods, 53(19):6918–6943, 2024. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2023.2253944.

Naqvi:2022:SSR

[NCM22] Sameen Naqvi, Ping Shing Chan, and Deepa Bhatt Mishra. System signatures: a review and bibliometric analysis. Communications in Statistics: Theory and Methods, 51(7):1993– 2008, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2021.1937653.

Nguyen:2023:APD

[NCNM23] TrungTin Nguyen, Faicel Chamroukhi, Hien D. Nguyen, and Geoffrey J. McLachlan. Approximation of probability density functions via location-scale finite mixtures in Lebesgue spaces. Communications in Statistics: Theory and Methods, 52(14): 5048–5059, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2021.2002360.

Nguyen:2021:ARC

[ND21] Van Trinh Nguyen and Jean-François Dupuy. Asymptotic results in censored zero-inflated Poisson regression. Communications in Statistics: Theory and Methods, 50(12):2759–2779, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2019.1676442.

Negro:2024:SDT

[Neg24] L. Negro. Sample distribution theory using Coarea Formula. Communications in Statistics: Theory and Methods, 53(5): 1864–1889, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2022.2116284.

Nesrine:2024:TFA

[Nes24] Kara-Terki Nesrine. Tests in functional autoregressive processes via local asymptotic normality condition. Communications in Statistics: Theory and Methods, 53(23):8556–8570, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2023.2293641.

Newer:2023:SPV

[New23] Haidy A. Newer. Saddle-point p-values and confidence intervals based on log-rank tests when dependent subunits of clustered survival data are randomized by random allocation design. Communications in Statistics: Theory and Methods, 52(12):4072–4082, 2023. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2021.1986532.

Nanda:2022:EDD

[NGC22] Asok K. Nanda, Sanjib Gayen, and Shovan Chowdhury. Errors due to departure from independence in exponential series system. Communications in Statistics: Theory and Methods, 51(12):4162–4178, 2022. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2020.1811340.

Nakagawa:2020:RBI

[NH20] Tomoyuki Nakagawa and Shintaro Hashimoto. Robust Bayesian inference via γ-divergence. Communications in Statistics: Theory and Methods, 49(2):343–360, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1543765.

Nievergelt:2020:UOP

[Nie20] Yves Nievergelt. Uniqueness of one-parameter exponential curves fitted by non-linear least-squares to non-negative data in monotone non-increasing blocks. Communications in Statistics: Theory and Methods, 49(24):6129–6132, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1628274.

Naderi:2024:JWL

[NJMM24] Habib Naderi, Mehdi Jafari, Przemysław Matuła, and Morteza Mohammadi. On the Jajte weak law of large numbers for exchangeable random variables. Communications in Statistics: Theory and Methods, 53(9):3226–3234, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2150827.

Ng:2022:UAH

[NK22] Tin Lok James Ng and Kwok-Kun Kwong. Universal approximation on the hypersphere. Communications in Statistics: Theory and Methods, 51(24):8694–8704, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1904988.

Nagase:2023:CSC

[NK23] Mario Nagase and Yutaka Kano. Cyclic structural causal model with unobserved confounder effect. Communications in Statistics: Theory and Methods, 52(2):335–345, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1913186.

Nida:2024:DHC

[NKK^{+}24] Hafiza Nida, Muhammad Kashif, Muhammad Imran Khan, Liaquat Ahmad, and Muhammad Aslam. Design of Hotelling T^{2} control chart using various covariance structures. Communications in Statistics: Theory and Methods, 53(16):5828– 5839, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2023.2234520.

Nkou:2023:SCK

[Nko23] Emmanuel De Dieu Nkou. Strong consistency of kernel method for sliced average variance estimation. Communications in Statistics: Theory and Methods, 52(21):7586–7600, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2022.2049821.

Ng:2022:DL

[NLL22] Chi Tim Ng, Woojoo Lee, and Youngjo Lee. In defense of LASSO. Communications in Statistics: Theory and Methods, 51(9):3018–3042, 2022. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2020.1788080.

Liang:2024:ADS

[nLSZ24] Shu na Liang, Fei Sun, and Qi Zhang. Adaptive distributed support vector regression of massive data. Communications in Statistics: Theory and Methods, 53(9):3365–3382, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2153604.

Nasrollahzadeh:2023:STN

[NMF23] Sh. Nasrollahzadeh, M. Bameni Moghadam, and R. Farnoosh. A Shewhart-type nonparametric multivariate depth-based control chart for monitoring location. Communications in Statistics: Theory and Methods, 52(20):7385–7404, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2045023.

Naseri:2022:INL

[NMT22] Parisa Naseri, Hamid Alavi Majd, and Seyyed Mohammad Tabatabaei. Integrated nested Laplace approximation method for hierarchical Bayesian inference of spatial model with application to functional magnetic resonance imaging data. Communications in Statistics: Theory and Methods, 51(8):2414– 2437, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2020.1776327.

Nesaee:2021:DVS

[NN21] Marziyeh Nesaee and Mohammad Saber Fallah Nezhad. Designing variables sampling plans based on the yield index S_{pk}. Communications in Statistics: Theory and Methods, 50(3): 507–520, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2019.1639742.

Ngailo:2024:ARE

[NN24] Edward Kanuti Ngailo and Innocent Ngaruye. Asymptotic results for expected probability of misclassifications in linear discriminant analysis with repeated measurements. Communications in Statistics: Theory and Methods, 53(6):1942–1963, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2022.2116286.

Naseri:2020:SPC

[NNS20] Hassan Naseri, S. Esmaeil Najafi, and Abbas Saghaei. Statistical process control (SPC) for short production run with Cauchy distribution, a case study with corrected numbers approach. Communications in Statistics: Theory and Methods, 49(4):879–893, 2020. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2018.1549260.

Noughabi:2022:CRE

[Nou22] Hadi Alizadeh Noughabi. Cumulative residual entropy applied to testing uniformity. Communications in Statistics: Theory and Methods, 51(12):4151–4161, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1811339.

Noughabi:2023:TUB

[Nou23] Hadi Alizadeh Noughabi. Testing uniformity based on negative cumulative extropy. Communications in Statistics: Theory and Methods, 52(14):4998–5009, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.2001015.

Novikov:2024:GSH

[Nov24] Andrey Novikov. Group sequential hypothesis tests with variable group sizes: Optimal design and performance evaluation. Communications in Statistics: Theory and Methods, 53(16): 5744–5760, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2023.2231155.

Nandy:2023:SJD

[NP23] N. Nandy and S. Pradhan. Stationary joint distribution of a discrete-time group-arrival and batch-size-dependent service queue with single and multiple vacation. Communications in Statistics: Theory and Methods, 52(9):3012–3046, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1966469.

Najafi:2021:SMC

[NS21a] Hoorieh Najafi and Abbas Saghaei. Statistical monitoring for change detection of interactions between nodes in networks: With a case study in financial interactions network. Communications in Statistics: Theory and Methods, 50(20):4900–4911, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2020.1725830.

Narjis:2021:BAO

[NS21b] Ghulam Narjis and Javid Shabbir. Bayesian analysis of optional unrelated question randomized response models. Communications in Statistics: Theory and Methods, 50(18):4203– 4215, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2020.1713367.

Narjis:2021:EPR

[NS21c] Ghulam Narjis and Javid Shabbir. An efficient partial randomized response model for estimating a rare sensitive attribute using Poisson distribution. Communications in Statistics: Theory and Methods, 50(1):1–17, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1628992.

Nevzorov:2021:SRM

[NS21d] V. B. Nevzorov and A. Stepanov. Some results on maximum of expected sums of sequential minima. Communications in Statistics: Theory and Methods, 50(6):1362–1369, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1649431.

Narjis:2022:TSU

[NS22a] Ghulam Narjis and Javid Shabbir. A two-stage unrelated question randomized response model for estimating the rare sensitive parameter under Poisson distribution. Communications in Statistics: Theory and Methods, 51(6):1836–1856, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2020.1769671.

Natarajan:2022:ANR

[NS22b] Balasubramaniam Natarajan and Weixing Song. Adaptive nonparametric regression on finite support. Communications in Statistics: Theory and Methods, 51(2):427–447, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1749855.

Nair:2022:BIQ

[NSD22] N. Unnikrishnan Nair, P. G. Sankaran, and M. Dileepkumar. Bayesian inference in quantile functions. Communications in Statistics: Theory and Methods, 51(14):4877–4889, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1827430.

Nair:2021:SAR

[NSG21] N. Unnikrishnan Nair, S. M. Sunoj, and Rajesh G. Some aspects of reversed hazard rate and past entropy. Communications in Statistics: Theory and Methods, 50(9):2106–2116, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2019.1659970.

Najafabadi:2022:JML

[NSG22] Mojtaba Zeinali Najafabadi, Ehsan Bahrami Samani, and Mojtaba Ganjali. Joint modeling of longitudinal count and timeto-event data with excess zero using accelerated failure time model: an application with CD4 cell counts. Communications in Statistics: Theory and Methods, 51(20):7243–7263, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1872635.

Nair:2023:CRE

[NSG23] N. Unnikrishnan Nair, S. M. Sunoj, and Rajesh G. Cumulative residual entropy of equilibrium distribution of order n. Communications in Statistics: Theory and Methods, 52(3):851– 863, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2021.1931332.

Nasuda:2023:ABD

[NSH23] Ryo Nasuda, Koki Shimizu, and Hiroki Hashiguchi. Asymptotic behavior of the distributions of eigenvalues for beta-Wishart ensemble under the dispersed population eigenvalues. Communications in Statistics: Theory and Methods, 52(22): 7840–7860, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2022.2050404.

Nili-Sani:2022:KFT

[NSJ22] H. R. Nili-Sani and M. Jafari. The Kolmogrov–Feller type weak law of large numbers for APND random variables. Communications in Statistics: Theory and Methods, 51(24):8634– 8643, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2021.1901922.

Naderi:2021:COD

[NSM21] M. Hossein Naderi, Asghar Seif, and M. Bameni Moghadam. Constrained optimal design of X̄ control chart for correlated data under Weibull Shock Model with multiple assignable causes. Communications in Statistics: Theory and Methods, 50(10):2322–2353, 2021. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2019.1664587.

Nair:2020:CSB

[NSN20] N. Unnikrishnan Nair, S. M. Sunoj, and Vipin N. On characterizations of some bivariate continuous distributions by properties of higher-degree bivariate stop-loss transforms. Communications in Statistics: Theory and Methods, 49(2):403–420, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2018.1543768.

Nevzorov:2022:LPS

[NSS22] Valery Borisovich Nevzorov, Valeriia Savinova, and Alexei Stepanov. Linear prediction of sums of sequential minima. Communications in Statistics: Theory and Methods, 51(16): 5446–5454, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2020.1841237.

Naatjes:2024:URT

[NSS24] Daryan Naatjes, Stephen A. Sedory, and Sarjinder Singh. An unbiased regression type estimator of proportion in randomized response sampling by using analysis of variance mechanism. Communications in Statistics: Theory and Methods, 53(14):5210–5217, 2024. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2023.2214296.

Nair:2023:BRE

[NSSR23] N. Unnikrishnan Nair, Silpa Subhash, S. M. Sunoj, and G. Rajesh. Bivariate residual entropy function: a quantile approach. Communications in Statistics: Theory and Methods, 52(24): 8611–8635, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2022.2067334.

Nadeb:2020:PPS

[NT20a] Hossein Nadeb and Hamzeh Torabi. Preservation properties of stochastic orders by transformation to the transmuted-G model. Communications in Statistics: Theory and Methods, 49(17):4333–4346, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03610926.2019.1601220.

Nagamani:2020:IEQ

[NT20b] Nadiminti Nagamani and Manas Ranjan Tripathy. Improved estimation of quantiles of two normal populations with common mean and ordered variances. Communications in Statistics: Theory and Methods, 49(19):4669–4692, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1604964.

Nadeb:2022:NRS

[NT22] Hossein Nadeb and Hamzeh Torabi. New results on stochastic comparisons of finite mixtures for some families of distributions. Communications in Statistics: Theory and Methods, 51(10):3104–3119, 2022. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2020.1788082.

Nath:2023:SAC

[NTDB23] Jagannath Nath, Binod Chandra Tripathy, Piyali Debnath, and Baby Bhattacharya. Strongly almost convergence in sequences of complex uncertain variables. Communications in Statistics: Theory and Methods, 52(3):714–729, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1921802.

Noor:2022:HEW

[NuAMA22] Surria Noor, Muhammad Noor ul Amin, Muhammad Mohsin, and Azaz Ahmed. Hybrid exponentially weighted moving average control chart using Bayesian approach. Communications in Statistics: Theory and Methods, 51(12):3960–3984, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2020.1805765.

Naqvi:2020:ORI

[NZZ20] Sameen Naqvi, Yiying Zhang, and Peng Zhao. Ordering results for individual risk model with dependent location–scale claim severities. Communications in Statistics: Theory and Methods, 49(4):942–957, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1554127.

Okur:2021:SHH

[OA21] Nurgul Okur and Rovshan Aliyev. Some Hermite–Hadamard type integral inequalities for multidimensional general preinvex stochastic processes. Communications in Statistics: Theory and Methods, 50(14):3338–3351, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1696976.

Ozkale:2023:BCI

[ÖA23] M. Revan Özkale and Hüsniye Altuner. Bootstrap confidence interval of ridge regression in linear regression model: a comparative study via a simulation study. Communications in Statistics: Theory and Methods, 52(20):7405–7441, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2045024.

Oner:2021:REC

[ÖB21] Yüksel Öner and Hasan Bulut. A robust EM clustering approach: ROBEM. Communications in Statistics: Theory and Methods, 50(19):4587–4605, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1722840.

Ohaegbulem:2022:APR

[OC22a] Emmanuel Uchenna Ohaegbulem and Polycarp E. Chigbu. Assessing the percent rotatability of the central composite designs. Communications in Statistics: Theory and Methods, 51(21):7665–7682, 2022. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2021.1876885.

Ohaegbulem:2022:MOC

[OC22b] Emmanuel Uchenna Ohaegbulem and Polycarp Emeka Chigbu. A measure of orthogonality for the central composite designs. Communications in Statistics: Theory and Methods, 51(9):2710–2724, 2022. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2021.1982984.

Oliveira:2023:LBB

[OCSV23] Kessys L. P. Oliveira, Bruno S. Castro, Helton Saulo, and Roberto Vila. On a length-biased Birnbaum–Saunders regression model applied to meteorological data. Communications in Statistics: Theory and Methods, 52(19):6916–6935, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2037642.

Omidi:2023:GFT

[OFH23] Fatemeh Omidi, Vahid Fakoor, and Arezou Habibirad. Goodness-of-fit test based on information criterion for interval censored data. Communications in Statistics: Theory and Methods, 52(3):830–850, 2023. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2021.1931331.

Ohishi:2021:COG

[OFO^{+}21] M. Ohishi, K. Fukui, K. Okamura, Y. Itoh, and H. Yanagihara. Coordinate optimization for generalized fused Lasso. Communications in Statistics: Theory and Methods, 50(24):5955– 5973, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2021.1931888.

OGorman:2023:GNS

[O’G23] Thomas W. O’Gorman. A generalized normal scores test that increases the power of a test of significance for a coefficient in a linear model. Communications in Statistics: Theory and Methods, 52(12):4216–4228, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1987471.

Ogasawara:2020:AEF

[Oga20a] Haruhiko Ogasawara. Alternative expectation formulas for real-valued random vectors. Communications in Statistics: Theory and Methods, 49(2):454–470, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1543773.

Ogasawara:2020:ACM

[Oga20b] Haruhiko Ogasawara. Asymptotic cumulants of the minimum phi-divergence estimator for categorical data under possible model misspecification. Communications in Statistics: Theory and Methods, 49(10):2448–2465, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1576888.

Ogasawara:2020:AEC

[Oga20c] Haruhiko Ogasawara. An asymptotic equivalence of the cross-data and predictive estimators. Communications in Statistics: Theory and Methods, 49(3):755–768, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1549258.

Ogasawara:2020:EMC

[Oga20d] Haruhiko Ogasawara. The echelon Markov and Chebyshev inequalities. Communications in Statistics: Theory and Methods, 49(7):1578–1591, 2020. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2018.1530359.

Ogasawara:2020:MMM

[Oga20e] Haruhiko Ogasawara. The multivariate Markov and multiple Chebyshev inequalities. Communications in Statistics: Theory and Methods, 49(2):441–453, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1543772.

Ogasawara:2021:IMC

[Oga21] Haruhiko Ogasawara. Improvements of the Markov and Chebyshev inequalities using the partial expectation. Communications in Statistics: Theory and Methods, 50(1):116–131, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2019.1630438.

Ogasawara:2022:UNR

[Oga22] Haruhiko Ogasawara. Unified and non-recursive formulas for moments of the normal distribution with stripe truncation. Communications in Statistics: Theory and Methods, 51(19): 6834–6862, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2020.1867742.

Ogasawara:2024:MDM

[Oga24] Haruhiko Ogasawara. The multivariate t-distribution with multiple degrees of freedom. Communications in Statistics: Theory and Methods, 53(1):144–169, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2076122.

Oh:2023:GEP

[Oh23] Jin H. Oh. Graphical evaluation of prediction capability when the number of noise variable increases in robust parameter design. Communications in Statistics: Theory and Methods, 52(18):6561–6572, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2032166.

Oh:2021:SGR

[OHS21] Jungtaek Oh, Hee-Jin Hwang, and Key-Il Shin. A study on the generalized ratio-type estimator based on the multiple regression estimator. Communications in Statistics: Theory and Methods, 50(24):6151–6166, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1740270.

Ohyama:2020:SIA

[Ohy20] Tetsuji Ohyama. Statistical inference of agreement coefficient between two raters with binary outcomes. Communications in Statistics: Theory and Methods, 49(10):2529–2539, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1576894.

Ohyama:2021:SIG

[Ohy21] Tetsuji Ohyama. Statistical inference of Gwet’s AC_{1} coefficient for multiple raters and binary outcomes. Communications in Statistics: Theory and Methods, 50(15):3564–3572, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2019.1708397.

Oliazadeh:2021:NSC

[OIF21] Farzaneh Oliazadeh, Anis Iranmanesh, and Vahid Fakoor. A note on the strong consistency of nonparametric estimation of Shannon entropy in length-biased sampling. Communications in Statistics: Theory and Methods, 50(24):5779–5791, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2020.1737123.

Ozawa:2020:ETC

[OK20] Kazuhiro Ozawa and Shinji Kuriki. On A-efficient treatmentcontrol designs constructed by cyclic and generalized cyclic designs. Communications in Statistics: Theory and Methods, 49(24):6096–6111, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03610926.2019.1626428.

Okumura:2023:BRM

[Oku23] Hidenori Okumura. Bias reduction and model selection in misspecified models. Communications in Statistics: Theory and Methods, 52(8):2751–2765, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1959613.

Oumaima:2024:SPG

[OMS24] Ben Mrad Oumaima, Afif Masmoudi, and Yousri Slaoui. Some properties of q-Gaussian distributions. Communications in Statistics: Theory and Methods, 53(17):6315–6337, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2244097.

Oda:2021:HDB

[OMYF21] Ryoya Oda, Yoshie Mima, Hirokazu Yanagihara, and Yasunori Fujikoshi. A high-dimensional bias-corrected AIC for selecting response variables in multivariate calibration. Communications in Statistics: Theory and Methods, 50(14):3453–3476, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2019.1705978.

Oladugba:2023:RDS

[ON23] Abimbola V. Oladugba and Ogechukwu C. Nwanonobi. Robustness of definitive screening composite designs to missing observations. Communications in Statistics: Theory and Methods, 52(15):5349–5363, 2023. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2021.2006715.

Ossai:2024:EMS

[ONM^{+}24] Everestus O. Ossai, Uchenna C. Nduka, Mbanefo S. Madukaife, Akaninyene U. Udom, and Samson O. Ugwu. An extended Markov-switching model approach to latent heterogeneity in departmentalized manpower systems. Communications in Statistics: Theory and Methods, 53(19):6957–6976, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2023.2255322.

Onwuamaeze:2021:OPV

[Onw21] Charity Uchenna Onwuamaeze. Optimal prediction variance properties of some central composite designs in the hypercube. Communications in Statistics: Theory and Methods, 50(8): 1911–1924, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2019.1656746.

Ommane:2020:ARE

[OO20] Younes Ommane and Idir Ouassou. Adaptive ridge estimator in a linear regression model with spherically symmetric error under constraint. Communications in Statistics: Theory and Methods, 49(1):1–15, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1532006.

Olive:2022:PIG

[ORH22] David J. Olive, Rasanji C. Rathnayake, and Mulubrhan G. Haile. Prediction intervals for GLMs, GAMs, and some survival regression models. Communications in Statistics: Theory and Methods, 51(22):8012–8026, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1887238.

Oryshchenko:2020:EMI

[Ory20] Vitaliy Oryshchenko. Exact mean integrated squared error and bandwidth selection for kernel distribution function estimators. Communications in Statistics: Theory and Methods, 49(7):1603–1628, 2020. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2018.1563182.

Oliveira:2016:MHL

[OSX^{+}16] José Oliveira, Jéssica Santos, Cleber Xavier, Daniele Trindade, and Gauss M. Cordeiro. The McDonald halflogistic distribution: Theory and practice. Communications in Statistics: Theory and Methods, 45(7):2005–2022, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2013.873131. See correction [AE23c].

Ong:2021:NCN

[OTL21] Seng-Huat Ong, Kian-Kok Toh, and Yeh-Ching Low. The non-central negative binomial distribution: Further properties and applications. Communications in Statistics: Theory and Methods, 50(2):329–344, 2021. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2019.1634817.

Orjuela:2021:BEM

[OV21] Lizet Viviana Romero Orjuela and Sergio Alejandro Calderón Villanueva. Bayesian estimation of a multivariate TAR model when the noise process follows a Student-t distribution. Communications in Statistics: Theory and Methods, 50(11):2508– 2530, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2019.1669807.

Oladugba:2024:PVC

[OYA24] Abimibola Victoria Oladugba, Brenda Mbouamba Yankam, and Oluchukwu Chukwuemeka Asogwa. Prediction variance capability of orthogonal uniform composite designs and orthogonal array composite designs in the spherical region. Communications in Statistics: Theory and Methods, 53(5): 1611–1623, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2022.2107222.

Oladugba:2024:FDS

[OYO24] Abimibola Victoria Oladugba, Brenda Mbouamba Yankam, and Uchenna Charity Onwuamaeze. Fraction of design space plots for evaluating orthogonal composite designs. Communications in Statistics: Theory and Methods, 53(20):7405–7416, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2023.2264993.

Ozkale:2021:IOI

[Özk21] M. Revan Özkale. Identification of outlying and influential data with principal components regression estimation in binary logistic regression. Communications in Statistics: Theory and Methods, 50(3):609–630, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1639749.

Pascual:2021:MBI

[PA21] Francis G. Pascual and Sherzod B. Akhundjanov. Monitoring a bivariate INAR(1) process with application to hepatitis a. Communications in Statistics: Theory and Methods, 50(5): 1036–1058, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2019.1645856.

Pekalp:2020:PEM

[PAAA20] Mustafa Hilmi Pekalp, Ömer Altindag, Özgür Acar, and Halil Aydogdu. Plug-in estimators for the mean value and variance functions in delayed renewal processes. Communications in Statistics: Theory and Methods, 49(19):4693–4711, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1604965.

Pu:2020:EAF

[PAC20] Xiao Pu and Ery Arias-Castro. An EM algorithm for fitting a mixture model with symmetric log-concave densities. Communications in Statistics: Theory and Methods, 49(1):78–87, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2018.1530789.

Pang:2022:BVO

[PAdlPK22] Guodong Pang, Demissie Alemayehu, Victor de la Peña, and Michael J. Klass. On the bias and variance of odds ratio, relative risk and false discovery proportion. Communications in Statistics: Theory and Methods, 51(19):6883–6908, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1867744.

Park:2021:WGC

[Par21] Sangun Park. Weighted general cumulative entropy and a goodness of fit for normality. Communications in Statistics: Theory and Methods, 50(20):4733–4742, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1723635.

Pareto:2023:NLC

[Par23] Adriano Pareto. A new look at the correlation coefficient: Correlation as the difference-sum ratio of SSEs. Communications in Statistics: Theory and Methods, 52(9):2852–2859, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2021.1961153.

Pirmohammadi:2022:LEB

[PB22] Shima Pirmohammadi and Hamid Bidram. On the Liu estimator in the beta and Kumaraswamy regression models: a comparative study. Communications in Statistics: Theory and Methods, 51(24):8553–8578, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1900254.

Panwar:2024:OBA

[PB24] M. S. Panwar and Vikas Barnwal. Objective Bayesian analysis of Marshall–Olkin bivariate Weibull distribution with partial information. Communications in Statistics: Theory and Methods, 53(15):5331–5352, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2219418.

Peng:2022:MVA

[PC22] Xingchun Peng and Fenge Chen. Mean-variance asset-liability management with inside information. Communications in Statistics: Theory and Methods, 51(7):2281–2302, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1772982.

Pereira:2020:CRE

[PCdP^{+}20] André G. C. Pereira, Viviane S. M. Campos, André L. S. de Pinho, Carla A. Vivacqua, and Roberto T. G. de Oliveira. On the convergence rate of the elitist genetic algorithm based on mutation probability. Communications in Statistics: Theory and Methods, 49(4):769–780, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1528361.

Pang:2021:NIS

[PCZ21] Tianxiao Pang, Terence Tai-Leung Chong, and Danna Zhang. Non identification of structural change in non stationary AR(1) models. Communications in Statistics: Theory and Methods, 50(18):4145–4166, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1711125.

Pang:2024:HDT

[PCZ24] Shanqi Pang, Mengqian Chen, and Xiao Zhang. Hamming distances of tight orthogonal arrays. Communications in Statistics: Theory and Methods, 53(17):6012–6029, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2239395.

Prevedello:2022:DCS

[PD22] Giulio Prevedello and Ken R. Duffy. Discrete convolution statistic for hypothesis testing. Communications in Statistics: Theory and Methods, 51(12):4097–4118, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1811336.

Pakdaman:2023:IEF

[PD23] Zohreh Pakdaman and Mahdi Doostparast. Influence of environmental factors on stress-strength systems with replaceable components. Communications in Statistics: Theory and Methods, 52(18):6487–6503, 2023. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2022.2031217.

Pan:2024:CVR

[PD24] Lipeng Pan and Yong Deng. Complex-valued Rényi entropy. Communications in Statistics: Theory and Methods, 53(3): 926–937, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2022.2094963.

Peng:2022:EUM

[Pen22] Xingchun Peng. Expected utility maximization for an insurer with investment and risk control under inside information. Communications in Statistics: Theory and Methods, 51(4): 1029–1053, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2020.1757716.

Peng:2024:OIW

[Pen24] Ling Peng. Oracle inequalities for weighted group lasso in high-dimensional Poisson regression model. Communications in Statistics: Theory and Methods, 53(19):6891–6917, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2253940.

Papatsouma:2020:ASD

[PF20] Ioanna Papatsouma and Nikolaos Farmakis. Approximating symmetric distributions via sampling and coefficient of variation. Communications in Statistics: Theory and Methods, 49 (1):61–77, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2018.1529244.

Pizzinga:2023:DKF

[PF23] Adrian Pizzinga and Marcelo Fernandes. Diffuse Kalman filtering with linear constraints on the state parameters. Communications in Statistics: Theory and Methods, 52(24):8884– 8893, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2022.2084109.

Pradhan:2022:SQS

[PG22] S. Pradhan and U. C. Gupta. Stationary queue and server content distribution of a batch-size-dependent service queue with batch Markovian arrival process: BMAP/G_{n}^{(a,b)} /1. Communications in Statistics: Theory and Methods, 51(13):4330– 4357, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2020.1813304.

Pielaszkiewicz:2021:MTW

[PH21] Jolanta Pielaszkiewicz and Thomas Holgersson. Mixtures of traces of Wishart and inverse Wishart matrices. Communications in Statistics: Theory and Methods, 50(21):5084–5100, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2019.1691733.

Pan:2024:DPE

[PH24] Lu Pan and Jianwei Hu. Differentially private estimation in a class of bipartite graph models. Communications in Statistics: Theory and Methods, 53(18):6477–6496, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2246090.

Pandher:2023:EEM

[PHBV23] S. S. Pandher, S. Hossain, K. Budsaba, and A. Volodin. Efficient estimation method for generalized ARFIMA models. Communications in Statistics: Theory and Methods, 52(23): 8515–8537, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2022.2064503.

Pho:2024:ZIL

[Pho24] Kim-Hung Pho. Zero-inflated logit probit model: a novel model for binary data. Communications in Statistics: Theory and Methods, 53(18):6580–6599, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2248325.

Panahi:2024:OPP

[PJB24] Abed Hossein Panahi, Habib Jafari, and Ghobad Saadat Kia (Barmalzan). Ordering properties of parallel and series systems with a general lifetime family of distributions for independent components under random shocks. Communications in Statistics: Theory and Methods, 53(21):7582–7603, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2269445.

Park:2020:RDB

[PK20a] Chun Gun Park and Inyoung Kim. Robust differencebased outlier detection. Communications in Statistics: Theory and Methods, 49(22):5553–5577, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1620278.

Patra:2020:FRR

[PK20b] Arijit Patra and Chanchal Kundu. Further results on residual life and inactivity time at random time. Communications in Statistics: Theory and Methods, 49(5):1261–1271, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1563170.

Park:2022:DBM

[PK22a] Junhui Park and Seung-Ho Kang. The determination of biosimilarity margin and the assessment of biosimilarity for an (m + 1)-arm parallel design. Communications in Statistics: Theory and Methods, 51(2):387–403, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1749664.

Park:2022:HME

[PK22b] Myung Hyun Park and Joseph H. T. Kim. Hierarchical mixture-of-experts models for count variables with excessive zeros. Communications in Statistics: Theory and Methods, 51(12):4072–4096, 2022. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2020.1811335.

Patra:2022:SCA

[PK22c] Arijit Patra and Chanchal Kundu. Stochastic comparisons and ageing properties of RLRT (ITRT) based on variance residual life. Communications in Statistics: Theory and Methods, 51(13):4276–4295, 2022. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2020.1812655.

Patra:2020:IEF

[PKK20] Lakshmi Kanta Patra, Somesh Kumar, and B. M. Golam Kibria. Improved estimation of a function of scale parameter of a doubly censored exponential distribution. Communications in Statistics: Theory and Methods, 49(9):2049–2064, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1568482.

Pilz:2022:NSS

[PKK22] Maximilian Pilz, Samuel Kilian, and Meinhard Kieser. A note on the shape of sample size functions of optimal adaptive two-stage designs. Communications in Statistics: Theory and Methods, 51(6):1911–1918, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1776875.

Panja:2023:DSO

[PKP23] Arindam Panja, Pradip Kundu, and Biswabrata Pradhan. Dispersive and star ordering of sample extremes from dependent random variables following the proportional odds model. Communications in Statistics: Theory and Methods, 52(19): 6936–6959, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2022.2037643.

Pathak:2022:AEB

[PKSS22] Anurag Pathak, Manoj Kumar, Sanjay Kumar Singh, and Umesh Singh. Assessing the effect of E-Bayesian inference for Poisson inverse exponential distribution parameters under different loss functions and its application. Communications in Statistics: Theory and Methods, 51(17):5763–5805, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1847293.

Priyanka:2021:CEQ

[PKT21] Kumari Priyanka, Ajay Kumar, and Pidugu Trisandhya. Calibration estimators for quantitative sensitive mean estimation under successive sampling. Communications in Statistics: Theory and Methods, 50(6):1341–1361, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1649430.

Peng:2023:NSM

[PLLC23] Qing Peng, Xin Lai, Liu Liu, and Jing Chen. A new standardized mortality ratio method for hospital quality evaluation. Communications in Statistics: Theory and Methods, 52(8): 2482–2492, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2021.1955381.

Pan:2023:CCI

[PLSW23] Yingli Pan, Zhan Liu, Guangyu Song, and Sha Wei. Casecohort and inference for the proportional hazards model with covariate adjustment. Communications in Statistics: Theory and Methods, 52(13):4379–4399, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1996607.

Pal:2021:ODP

[PM21] Manisha Pal and Nripes Kumar Mandal. Optimum designs for parameter estimation in mixture experiments with group synergism. Communications in Statistics: Theory and Methods, 50(9):2001–2014, 2021. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2019.1657455.

Prajapat:2024:MAE

[PMMK24] Kiran Prajapat, Shuvashree Mondal, Sharmishtha Mitra, and Debasis Kundu. Meta-analysis of exponential lifetime data from Type-I hybrid censored samples. Communications in Statistics: Theory and Methods, 53(11):3973–3991, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2169048.

Popinski:2020:LSO

[Pop20] Waldemar Popiński. Least squares orthogonal polynomial regression estimation for irregular design. Communications in Statistics: Theory and Methods, 49(3):631–647, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1549244.

Posa:2024:PDF

[Pos24] Donato Posa. Positive definite functions, stationary covariance functions, and Bochner’s theorem: Some results and a critical overview. Communications in Statistics: Theory and Methods, 53(15):5612–5628, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2223780.

Poursaeed:2020:SMM

[Pou20] Mohammad Hossein Poursaeed. On δ-shock model in a multi-state system. Communications in Statistics: Theory and Methods, 49(7):1761–1767, 2020. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2019.1565784.

Poursaeed:2021:RSE

[Pou21] Mohammad Hossein Poursaeed. A run shock-erosion model. Communications in Statistics: Theory and Methods, 50(5): 1228–1239, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2019.1649425.

Panga:2022:ASL

[PP22a] Zacarias Panga and Luísa Pereira. On the almost sure limit theorem for the joint version of maxima and minima of non-stationary random fields. Communications in Statistics: Theory and Methods, 51(10):3408–3418, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1823001.

Parasharam:2022:HSG

[PP22b] Gadre Mukund Parasharam and Nisha Padappa. Hybrid synthetic and group runs charts using alternating the charting statistic to monitor the mean vector of a multivariate process. Communications in Statistics: Theory and Methods, 51(21): 7607–7630, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2021.1875243.

Peng:2023:STE

[PP23] Kai Peng and Cheng Peng. Semiparametric tests for equality of two independent variances. Communications in Statistics: Theory and Methods, 52(9):3047–3069, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1966470.

Parui:2021:EBD

[PPM21] Shyamsundar Parui, Rajender Parsad, and B. N. Mandal. Efficient block designs for incomplete factorial experiments for two factors with unequal block sizes. Communications in Statistics: Theory and Methods, 50(11):2531–2545, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1670848.

Paige:2021:SSI

[PR21] Robert L. Paige and Noroharivelo V. Randrianampy. Small sample inference for exponential survival times with heavy right-censoring. Communications in Statistics: Theory and Methods, 50(3):521–539, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03610926.2019.1639743.

Prakash:2021:BAD

[Pra21] G. Prakash. A Bayesian approach to degradation modeling and reliability assessment of rolling element bearing. Communications in Statistics: Theory and Methods, 50(23):5453– 5474, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2020.1734826.

Priam:2020:VGM

[Pri20] Rodolphe Priam. Visualization of generalized mean estimators using auxiliary information in survey sampling. Communications in Statistics: Theory and Methods, 49(18):4468–4489, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2019.1601224.

Priyanka:2024:ABG

[Pri24] Kumari Priyanka. Analyzing the behavior of general class of estimators of population mean in presence of correlated measurement errors. Communications in Statistics: Theory and Methods, 53(6):2219–2235, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2122842.

Popovic:2023:ITC

[PRZ23] Bozidar V. Popović, Miroslav M. Ristić, and Konstantinos Zografos. Integral transformation of a copula function. Communications in Statistics: Theory and Methods, 52(18):6341– 6354, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2022.2028837.

Piradl:2021:RMD

[PS21] Sajjad Piradl and Ali Shadrokh. Robust minimum distance estimation of a linear regression model with correlated errors in the presence of outliers. Communications in Statistics: Theory and Methods, 50(23):5488–5498, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1734831.

Pawlas:2022:MOR

[PS22] Piotr Pawlas and Dominik Szynal. On a model of ordering random variables — the (i, k)-th record values. Communications in Statistics: Theory and Methods, 51(4):895–908, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1842891.

Pal:2023:EFP

[PS23a] Sanghamitra Pal and Purnima Shaw. Estimation of finite population distribution function of sensitive variable*. Communications in Statistics: Theory and Methods, 52(4):1318–1331, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2021.1934030.

Pant:2023:MPI

[PS23b] Himani Pant and S. B. Singh. Modeling periodically inspected k/r-out-of-n system. Communications in Statistics: Theory and Methods, 52(11):3895–3909, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1982982.

Pillai:2023:RAS

[PS23c] Thulasyammal R. Pillai and Mahendran Shitan. Retracted Article: Some properties of the generalized autoregressive moving average (GARMA(1, 2; δ, 1)) model. Communications in Statistics: Theory and Methods, 52(14):??, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.

1080/03610926.2013.851240. See retraction notice [Ano23c].

PoloGonzalez:2023:VAH

[PS23d] Mayo Luz Polo González and Ernesto Javier San Martín Gutiérrez. Value added in hierarchical linear mixed models with error in variables. Communications in Statistics: Theory and Methods, 52(22):7984–8001, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2055071.

Panda:2024:ODL

[PS24] Mahesh Kumar Panda and Rushi Prasad Sahoo. Roptimal designs for linear log contrast model with mixture experiments. Communications in Statistics: Theory and Methods, 53(7):2355–2368, 2024. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2022.2129993.

Psarrakos:2022:RBI

[Psa22] Georgios Psarrakos. Relations between integrated tails and moments based on the deficit at ruin in the renewal risk model. Communications in Statistics: Theory and Methods, 51(21): 7631–7651, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2021.1875244.

Pezzott:2020:NIM

[PSLLN20] George Lucas Moraes Pezzott, Luis Ernesto Bueno Salasar, José Galvão Leite, and Francisco Louzada-Neto. A note on identifiability and maximum likelihood estimation for a heterogeneous capture-recapture model. Communications in Statistics: Theory and Methods, 49(21):5273–5293, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1615628.

Phuong:2022:DUE

[PT22] Cao Xuan Phuong and Le Thi Hong Thuy. Deconvolution of P(X < Y) with unknown error distributions. Communications in Statistics: Theory and Methods, 51(17):5889–5912, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2020.1849722.

Pajari:2021:PEQ

[PTM21] Matti Pajari, Maria Tikanmäki, and Lasse Makkonen. Probabilistic evaluation of quantile estimators. Communications in Statistics: Theory and Methods, 50(14):3319–3337, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1696975.

Prasangika:2023:DSL

[PTYZ23] K. D. Prasangika, Wan Tang, Zeng Yao, and Guoxin Zuo. Double smoothing local linear estimation in nonlinear time series. Communications in Statistics: Theory and Methods, 52(5):1385–1399, 2023. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2021.1927096.

Puig:2024:LPT

[Pui24] Pedro Puig. Letter on the paper “On the two-parameter Bell–Touchard discrete distribution”. Communications in Statistics: Theory and Methods, 53(1):446–447, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2050398. See [CLMA20].

Prataviera:2023:REM

[PVC^{+}23] Fábio Prataviera, Roberto Vila, Vicente G. Cancho, Edwin M. M. Ortega, and Gauss M. Cordeiro. Reparameterized extended Maxwell regression: Properties, estimation and application. Communications in Statistics: Theory and Methods, 52(20):7252–7270, 2023. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2022.2042561.

Park:2023:SCC

[PW23] Chanseok Park and Min Wang. A study on the g and h control charts. Communications in Statistics: Theory and Methods, 52(20):7334–7349, 2023. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2022.2044492.

Pararai:2017:EPL

[PWLO17] Mavis Pararai, Gayan Warahena-Liyanage, and Broderick O. Oluyede. Exponentiated power Lindley–Poisson distribution: Properties and applications. Communications in Statistics: Theory and Methods, 46(10):4726–4755, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1076473. See correction [AE23d].

Peng:2020:SGW

[PWM20] Xiaozhi Peng, Hecheng Wu, and Ling Ma. A study on geographically weighted spatial autoregression models with spatial autoregressive disturbances. Communications in Statistics: Theory and Methods, 49(21):5235–5251, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1615507.

Peng:2023:SMT

[PX23] Weicai Peng and Xinyue Xi. Shannon-McMillan theorem and strong law of large numbers for Markov chains indexed by generalized spherically symmetric trees. Communications in Statistics: Theory and Methods, 52(8):2562–2573, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1955385.

Pan:2020:AMN

[PY20] Lu Pan and Ting Yan. Asymptotics in the β-model for networks with a differentially private degree sequence. Communications in Statistics: Theory and Methods, 49(18):4378–4393, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2019.1599023.

Pang:2023:SSO

[PYZS23] Shanqi Pang, Rong Yan, Xiao Zhang, and Jinhui Shen. Schematic saturated orthogonal arrays obtained by using the expansive replacement method. Communications in Statistics: Theory and Methods, 52(8):2434–2447, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1953074.

Provost:2022:GPF

[PZAH22] Serge B. Provost, Hossein Zareamoghaddam, S. Ejaz Ahmed, and Hyung-Tae Ha. The generalized Pearson family of distributions and explicit representation of the associated density functions. Communications in Statistics: Theory and Methods, 51(16):5590–5606, 2022. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2020.1843680.

Pu:2024:GLS

[PZY24] Tong Pu, Yiying Zhang, and Chuancun Yin. Generalized location-scale mixtures of elliptical distributions: Definitions and stochastic comparisons. Communications in Statistics: Theory and Methods, 53(11):3851–3875, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2165407.

Pang:2020:HDS

[PZZ20] Shanqi Pang, Xiao Zhang, and Qingjuan Zhang. The Hamming distances of saturated asymmetrical orthogonal arrays with strength 2. Communications in Statistics: Theory and Methods, 49(16):3895–3910, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1591452.

Qasim:2020:PSN

[QAO20] Muhammad Qasim, Muhammad Amin, and Talha Omer. Performance of some new Liu parameters for the linear regression model. Communications in Statistics: Theory and Methods, 49(17):4178–4196, 2020. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2019.1595654.

Qarmalah:2022:LMM

[Qar22] Najla M. Qarmalah. Localized mixture models for prediction with application. Communications in Statistics: Theory and Methods, 51(9):2725–2747, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1779296.

Ding:2021:SSD

[qDSzW21] Fang qing Ding, Jing Song, and Zhong zhi Wang. Some strong deviation theorems for arrays of arbitrarily dependent stochastic sequence. Communications in Statistics: Theory and Methods, 50(20):4692–4702, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1722843.

Qiu:2021:EIM

[QE21] Guoxin Qiu and Abbas Eftekharian. Extropy information of maximum and minimum ranked set sampling with unequal samples. Communications in Statistics: Theory and Methods, 50(13):2979–2995, 2021. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2019.1678640.

Qin:2023:SIL

[QG23] Xinyan Qin and Wenhao Gui. Statistical inference of Lomax distribution based on adaptive progressive Type-II hybrid censored competing risks data. Communications in Statistics: Theory and Methods, 52(22):8114–8135, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2056750.

Qin:2024:SIM

[QG24] Jiajie Qin and Jing Guan. Statistical inference of multistate transition model for longitudinal data with measurement error and heterogeneity. Communications in Statistics: Theory and Methods, 53(20):7453–7476, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2264997.

Qureshi:2021:GEE

[QH21] Muhammad Nouman Qureshi and Muhammad Hanif. Generalized estimator for the estimation of clustered population mean in adaptive cluster sampling. Communications in Statistics: Theory and Methods, 50(14):3262–3275, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1691235.

Qiao:2022:OST

[QH22] Lei Qiao and Dong Han. Optimal sequential tests for detection of changes under finite measure space for finite sequences of networks. Communications in Statistics: Theory and Methods, 51(19):6585–6600, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1864824.

Qayed:2023:HTS

[QH23] Abdullah Qayed and Dong Han. Homogeneity test of several high-dimensional covariance matrices for stationary processes under non-normality. Communications in Statistics: Theory and Methods, 52(8):2783–2798, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1960375.

Qin:2021:ELG

[Qin21] Yongsong Qin. Empirical likelihood and GMM for spatial models. Communications in Statistics: Theory and Methods, 50(18):4367–4385, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03610926.2020.1716252.

Quan:2020:SPI

[QLZZ20] Hui Quan, Xiaodong Luo, Tianyue Zhou, and Peng-Liang Zhao. Seamless phase II/III/IIIb clinical trial designs with different endpoints for different phases. Communications in Statistics: Theory and Methods, 49(22):5436–5454, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1618871.

Qiu:2022:RET

[QMS22] Zhiping Qiu, Huijuan Ma, and Jianhua Shi. Reweighting estimators for the transformation models with length-biased sampling data and missing covariates. Communications in Statistics: Theory and Methods, 51(13):4252–4275, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1812653.

Qureshi:2022:ERR

[QQS^{+}22] Saba Qureshi, Fiza Qureshi, Arjumand Bano Soomro, Fida Hussain Chandio, Sobia Shafaq Shah, and Ijaz Ur Rehman. Exchange rate risk and sectoral returns: a waveletbased MRA-EDCC GARCH analysis. Communications in Statistics: Theory and Methods, 51(7):2154–2182, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1772304.

Qiu:2024:WER

[QR24] Guoxin Qiu and Mohammad Z. Raqab. On weighted extropy of ranked set sampling and its comparison with simple random sampling counterpart. Communications in Statistics: Theory and Methods, 53(1):378–395, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2082478.

Qiao:2022:ODO

[QSCH22] YuLong Qiao, JinSheng Sun, Philippe Castagliola, and XueLong Hu. Optimal design of one-sided exponential EWMA charts based on median run length and expected median run length. Communications in Statistics: Theory and Methods, 51(9):2887–2907, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1782937.

Qian:2024:ETI

[QSP^{+}24] Chen Qian, Deo Kumar Srivastava, Jianmin Pan, Melissa M. Hudson, and Shesh N. Rai. Estimating transition intensity rate on interval-censored data using semi-parametric with EM algorithm approach. Communications in Statistics: Theory and Methods, 53(17):6038–6054, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2239397.

Qi:2022:DCT

[QT22] Liangwei Qi and Yu Tang. Direct construction of three-level designs with less β-aberration. Communications in Statistics: Theory and Methods, 51(24):8598–8616, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1900869.

Qu:2022:CMC

[Qu22] Xiaoming Qu. Complete moment convergence of movingaverage processes under END assumptions. Communications in Statistics: Theory and Methods, 51(10):3446–3458, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1767138.

Qin:2022:NPC

[QZ22] Xu Qin and Yu Q. Zhang. Non parametric covariance model with circular condition and its application. Communications in Statistics: Theory and Methods, 51(22):7819–7829, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1881121.

Rashid:2022:EMO

[RAA22] Fareeha Rashid, Atif Akbar, and Hafiz Muhammad Arshad. Effects of missing observations on predictive capability of augmented Box-behnken designs. Communications in Statistics: Theory and Methods, 51(20):7225–7242, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1872633.

Raissi:2024:DST

[Raï24] Hamdi Raïssi. On the dependence structure of the trade/no trade sequence of illiquid assets. Communications in Statistics: Theory and Methods, 53(8):2715–2729, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2148471.

Rao:2022:CPM

[Rao22] B. L. S. Prakasa Rao. Characterization of probability measures by linear functions of Q-independent random variables defined on a homogeneous Markov chain. Communications in Statistics: Theory and Methods, 51(18):6529–6534, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1862873.

Rao:2023:MLE

[Rao23] B. L. S. Prakasa Rao. Maximum likelihood estimation for stochastic differential equations driven by a mixed fractional Brownian motion with random effects. Communications in Statistics: Theory and Methods, 52(11):3816–3824, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1980048.

Rao:2024:NET

[Rao24] B. L. S. Prakasa Rao. Nonparametric estimation of trend for SDEs driven by a Gaussian process. Communications in Statistics: Theory and Methods, 53(17):6152–6159, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2240917.

Rao:2022:GMD

[RAS^{+}22] Gadde Srinivasa Rao, Muhammad Aslam, Rehan Ahmad Khan Sherwani, Muhammad Ahmed Shehzad, and Chi-Hyuck Jun. Generalized multiple dependent state sampling plans for coefficient of variation. Communications in Statistics: Theory and Methods, 51(20):6990–7005, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1869989.

Rajab:2023:UOC

[RASD23] Muhammad Rajab, Rashid Ahmed, Farrukh Shehzad, and Muhammad Daniyal. Universal optimality of circular balanced repeated measurements designs through method of cyclic shifts. Communications in Statistics: Theory and Methods, 52(19):7057–7068, 2023. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2022.2038628.

Rayner:2021:CUT

[RB21] J. C. W. Rayner and D. J. Best. Conditional and unconditional tests for the extended Stuart–Maxwell design. Communications in Statistics: Theory and Methods, 50(24):6137– 6150, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2020.1740269.

Rodrigues:2023:BME

[RBC^{+}23] Josemar Rodrigues, Yury R. Benites, Vicente G. Cancho, N. Balakrishnan, and Adriano K. Suzuki. Bayesian meta-elliptical multivariate regression models with fixed marginals on unit intervals. Communications in Statistics: Theory and Methods, 52(3):918–938, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1933531.

Reyes:2020:GMS

[RBCG20] Jimmy Reyes, Inmaculada Barranco-Chamorro, and Héctor W. Gómez. Generalized modified slash distribution with applications. Communications in Statistics: Theory and Methods, 49(8):2025–2048, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03610926.2019.1568484.

Riyahi:2024:CJS

[RBD24] H. Riyahi, M. Baratnia, and M. Doostparast. Cumulative α-Jensen–Shannon measure of divergence: Properties and applications. Communications in Statistics: Theory and Methods, 53(17):5989–6011, 2024. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2023.2238861.

Rodrigues:2020:FPF

[RBS20] Josemar Rodrigues, Jorge L. Bazán, and Adriano K. Suzuki. A flexible procedure for formulating probability distributions on the unit interval with applications. Communications in Statistics: Theory and Methods, 49(3):738–754, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1549254.

Rodrigues:2020:FAD

[RBSNB20] Josemar Rodrigues, Marcelo Bourguignon, Manoel Santos-Neto, and N. Balakrishnan. Fractional approaches for the distribution of innovation sequence of INAR(1) processes. Communications in Statistics: Theory and Methods, 49(9):2205– 2216, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2019.1568492.

Chen:2024:ASC

[rChZhN^{+}24] Meng ru Chen, Ya hui Zhu, Xiao hui Niu, Wei cai Peng, and Zhong zhi Wang. On almost sure convergence for double arrays of dependent random variables. Communications in Statistics: Theory and Methods, 53(3):938–952, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2095403.

Righetto:2020:CMA

[RFVR20] Ana Julia Righetto, Christel Faes, Yannick Vandendijck, and Paulo Justiniano Ribeiro Jr. On the choice of the mesh for the analysis of geostatistical data using R-INLA. Communications in Statistics: Theory and Methods, 49(1):203–220, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1536209.

Ren:2021:GFT

[RG21a] Junru Ren and Wenhao Gui. Goodness-of-fit test for Rayleigh distribution based on progressively type-II censored sample. Communications in Statistics: Theory and Methods, 50(16): 3851–3874, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2020.1869988.

Roy:2021:RFS

[RG21b] Achintya Roy and Nitin Gupta. Reliability function of kout-of-n system equipped with two cold standby components. Communications in Statistics: Theory and Methods, 50(24): 5759–5778, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2020.1737122.

Rivas:2022:EIM

[RG22] Luisa Rivas and Manuel Galea. On estimation and influence measures for the negative binomial regression model based on Q-function. Communications in Statistics: Theory and Methods, 51(7):1954–1974, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1942493.

Rueda:2023:NAP

[RG23] Javier Neira Rueda and Andres Carrión García. New approaches to parameter estimation with statistical censoring by means of the CEV algorithm: Characterization of its properties for high-performance normal processes. Communications in Statistics: Theory and Methods, 52(10):3557–3573, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1977323.

Roy:2024:SUC

[RG24] Achintya Roy and Nitin Gupta. A study on utilization of a cold standby component to enhance the mean residual life function of a coherent system. Communications in Statistics: Theory and Methods, 53(19):6977–6996, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2255323.

Rezaee:2024:BIS

[RGS24] Alireza Rezaee, Mojtaba Ganjali, and Ehsan Bahrami Samani. Bayesian inference in a sample selection model with multiple selection rules. Communications in Statistics: Theory and Methods, 53(12):4290–4310, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2178260.

Ren:2020:MPT

[RIA20] Aizhen Ren, Takashi Ishida, and Yutaka Akiyama. Mathematical proof of the third order accuracy of the speedy double bootstrap method. Communications in Statistics: Theory and Methods, 49(16):3950–3964, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1594295.

Rilstone:2024:RBH

[Ril24] Paul Rilstone. On the relationship between higher-order stochastic expansions, influence functions and U-statistics for M-estimators. Communications in Statistics: Theory and Methods, 53(6):2103–2121, 2024. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2022.2118543.

Rayner:2023:RFT

[RJ23] J. C. W. Rayner and Glen Livingston Jr. Relating the Friedman test adjusted for ties, the Cochran–Mantel–Haenszel mean score test and the ANOVA F test. Communications in Statistics: Theory and Methods, 52(12):4369–4378, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1994606.

Roy:2020:TMI

[RK20] Anuradha Roy and Daniel Klein. Testing of mean interval for interval-valued data. Communications in Statistics: Theory and Methods, 49(20):5028–5044, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1612915.

Ramtirthkar:2021:JEO

[RK21] Mukund Ramtirthkar and Mohan Kale. Joint estimation of the offspring mean and offspring variance of a second order branching process. Communications in Statistics: Theory and Methods, 50(6):1314–1324, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1649427.

Ramtirthkar:2022:MCB

[RK22] Mukund Ramtirthkar and Mohan Kale. Multiplicative controlled branching process with immigration. Communications in Statistics: Theory and Methods, 51(21):7683–7690, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1887237.

Raj:2024:DMI

[RK24] Sayooj Sunil Raj and David S. Kim. Destructive measurement instrument repeatability estimation using two item types with equal coefficient of variation. Communications in Statistics: Theory and Methods, 53(4):1358–1380, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2100912.

Raqab:2021:CZT

[RKAA21] Mohammad Z. Raqab, Debasis Kundu, and Fahimah A. Al-Awadhi. Compound zero-truncated Poisson normal distribution and its applications. Communications in Statistics: Theory and Methods, 50(13):3030–3050, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1679182.

Rabbi:2021:MSL

[RKK^{+}21] Fazli Rabbi, Salahuddin Khan, Alamgir Khalil, Wali Khan Mashwani, Muhammad Shafiq, Pinar Göktas, and Yuksel. Akay Unvan. Model selection in linear regression using paired bootstrap. Communications in Statistics: Theory and Methods, 50(7):1629–1639, 2021. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2020.1725829.

Rakhmawati:2022:PCG

[RKS22] Dwi Yuli Rakhmawati, Kwang-Jae Kim, and Sumiati. Performance comparison of generalized confidence interval and modified sampling distribution approaches for assessing one-sided capability indices with gauge measurement errors. Communications in Statistics: Theory and Methods, 51(3):685–700, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2020.1752729.

Rakhmawati:2023:PAD

[RL23] Dwi Yuli Rakhmawati and Junghye Lee. A product acceptance decision-making method based on process capability with considering gauge measurement errors. Communications in Statistics: Theory and Methods, 52(8):2646–2665, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1955929.

Rong:2023:ELS

[RLQ23] Jianrong Rong, Yan Liu, and Yongsong Qin. Empirical likelihood for spatial dynamic panel data models with spatial lags and spatial errors. Communications in Statistics: Theory and Methods, 52(18):6658–6683, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2032172.

Reschenhofer:2021:DLR

[RM21] Erhard Reschenhofer and Manveer Kaur Mangat. Detecting long-range dependence with truncated ratios of periodogram ordinates. Communications in Statistics: Theory and Methods, 50(15):3645–3660, 2021. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2019.1709646.

Rhoads:2024:RBO

[RMS^{+}24] Gregory Rhoads, Eric Marland, Jose Almer Sanqui, Michael Bossé, and William Bauldry. A relationship between orthogonal regression and the coefficient of determination under rotation of data sets. Communications in Statistics: Theory and Methods, 53(23):8348–8358, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2281897.

Ratibenyakool:2020:RCB

[RN20] Yuttana Ratibenyakool and Kritsana Neammanee. Rate of convergence of binomial formula for option pricing. Communications in Statistics: Theory and Methods, 49(14):3537–3556, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2019.1590600.

Ratibenyakool:2022:CTF

[RN22] Yuttana Ratibenyakool and Kritsana Neammanee. Convergence of trinomial formula for European option pricing. Communications in Statistics: Theory and Methods, 51(18):6227– 6249, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2020.1860221.

Rasheed:2022:SUC

[RNA^{+}22] Muhammad Rasheed, Khadija Noreen, Rashid Ahmed, M. H. Tahir, and Farrukh Jamal. Some useful classes of minimal weakly balanced neighbor designs in circular blocks of two different sizes. Communications in Statistics: Theory and Methods, 51(24):8822–8839, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1975135.

Rathnayake:2023:BSG

[RO23] Rasanji C. Rathnayake and David J. Olive. Bootstrapping some GLM and survival regression variable selection estimators. Communications in Statistics: Theory and Methods, 52(8):2625–2645, 2023. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2021.1955389.

Rajapaksha:2024:WTT

[RO24] Kosman W. G. D. H. Rajapaksha and David J. Olive. Wald type tests with the wrong dispersion matrix. Communications in Statistics: Theory and Methods, 53(6):2236–2251, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2124116.

Rosenblad:2022:MVB

[Ros22] Andreas Karlsson Rosenblad. The mean, variance, and bias of the OLS based estimator of the extremum of a quadratic regression model for small samples. Communications in Statistics: Theory and Methods, 51(9):2870–2886, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1782936.

Ramos:2021:TSO

[ROSL21] Héctor M. Ramos, Jorge Ollero, and Alfonso Suárez-Llorens. Two sensitivity orders applied to the comparison of ROC curves. Communications in Statistics: Theory and Methods, 50(8):1884–1896, 2021. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2019.1656744.

Rahmani:2020:HTB

[RR20] Hamid Rahmani and Mostafa Razmkhah. Homogeneity test based on ranked set samples. Communications in Statistics: Theory and Methods, 49(19):4771–4786, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1609034.

Rattihalli:2023:GGD

[RR23] R. N. Rattihalli and S. R. Rattihalli. A generalisation of geometric distribution. Communications in Statistics: Theory and Methods, 52(13):4400–4413, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1994607.

Rincon:2024:RPF

[RS24] Luis Rincón and David J. Santana. Ruin probability for finite negative binomial mixture claims via recurrence sequences. Communications in Statistics: Theory and Methods, 53(2): 557–124, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2022.2087091.

Rasekhi:2020:NMP

[RSA20] A. Rasekhi and S. M. Sadooghi-Alvandi. A note on the most powerful invariant test of Rayleigh against exponential distribution. Communications in Statistics: Theory and Methods, 49(19):4762–4770, 2020. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2019.1608245.

Rezaee:2021:GCJ

[RSG21] Fateme Rezaee, Ehsan Bahrami Samani, and Mojtaba Ganjali. Gaussian copula joint models to analysis mixed correlated longitudinal count and continuous responses. Communications in Statistics: Theory and Methods, 50(23):5499–5516, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2020.1734825.

Rajesh:2022:KEE

[RSR22] G. Rajesh, S. M. Sunoj, and Richu Rajesh. Kernel estimation of entropy function under length-biased sampling. Communications in Statistics: Theory and Methods, 51(24):8684–8693, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2021.1904987.

Raju:2024:RAN

[RSR24] David Chris Raju, S. M. Sunoj, and G. Rajesh. Results and applications of a new inaccuracy measure based on cumulative Tsallis entropy. Communications in Statistics: Theory and Methods, 53(12):4334–4353, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2179371.

Rasekhi:2021:BCI

[RSY21] Mahdi Rasekhi, Mohammad Mehdi Saber, and Haitham M. Yousof. Bayesian and classical inference of reliability in multicomponent stress-strength under the generalized logistic model. Communications in Statistics: Theory and Methods, 50(21):5114–5125, 2021. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2020.1726958.

Rih:2023:KCD

[RT23] Soumia Rih and Abdelkader Tatachak. Kernel conditional density and mode estimation for psi-weakly dependent observations. Communications in Statistics: Theory and Methods, 52(7):2072–2098, 2023. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2021.1944216.

Roy:2020:CMP

[RTB20] Sudip Roy, Ram C. Tripathi, and Narayanaswamy Balakrishnan. A Conway Maxwell Poisson type generalization of the negative hypergeometric distribution. Communications in Statistics: Theory and Methods, 49(10):2410–2428, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1576885.

Roozegar:2024:DST

[RTN24] Rasool Roozegar, Abdolsaleh Toghdori, and Saralees Nadarajah. The distribution of the sum of two dependent randomly weighted random variables with applications. Communications in Statistics: Theory and Methods, 53(6):1985–2002, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2022.2117558.

Roa:2023:LDL

[RTT23] Tania Roa, Soledad Torres, and Ciprian Tudor. Limit distribution of the least square estimator with observations sampled at random times driven by standard Brownian motion. Communications in Statistics: Theory and Methods, 52(11):3730– 3750, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2021.1980044.

Riaz:2022:EME

[RuAD22] Afshan Riaz, Muhammad Noor ul Amin, and Eralp Dogu. Effect of measurement error on joint monitoring of process mean and coefficient of variation. Communications in Statistics: Theory and Methods, 51(19):6863–6882, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1867743.

Rezaei:2020:ESS

[RYJ20] Amir Rezaei, Fatemeh Yousefzadeh, and Sarah Jomhoori. Estimation of stress-strength reliability for the multivariate SGPII distribution. Communications in Statistics: Theory and Methods, 49(16):3860–3881, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1593457.

Rong:2024:ORI

[RYZ24] Ximin Rong, Yiqi Yan, and Hui Zhao. Optimal reinsurance and investment problem with multiple risky assets and correlation risk for an insurer under the Ornstein–Uhlenbeck model. Communications in Statistics: Theory and Methods, 53(8): 2689–2714, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2022.2148470.

Ren:2023:SSB

[RZ23] Min Ren and Sheng-Li Zhao. Subdata selection based on orthogonal array for big data. Communications in Statistics: Theory and Methods, 52(15):5483–5501, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.2012196.

Singh:2021:BMT

[SA21] Saroja Kumar Singh and Sarat Kumar Acharya. Bernstein– von Mises theorem and Bayes estimation from single server queues. Communications in Statistics: Theory and Methods, 50(2):286–296, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03610926.2019.1634211.

Sheikhrabori:2022:MLE

[SA22] Raza Sheikhrabori and Majid Aminnayeri. Maximum likelihood estimation of the change point in stationary state of auto regressive moving average (ARMA) models, using SVD-based smoothing. Communications in Statistics: Theory and Methods, 51(22):7801–7818, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1881120.

Singh:2023:BIP

[SACQ23] Saroja Kumar Singh, Sarat Kumar Acharya, Frederico R. B. Cruz, and Roberto C. Quinino. Bayesian inference and prediction in an M/D/1 queueing system. Communications in Statistics: Theory and Methods, 52(24):8844–8864, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2076120.

Shojaee:2024:HRM

[SAF24] Omid Shojaee, Majid Asadi, and Maxim Finkelstein. On the hazard rate of α-mixture of survival functions. Communications in Statistics: Theory and Methods, 53(11):4062–4084, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2023.2172586.

Sanaullah:2021:NRR

[SAH21] Aamir Sanaullah, Azaz Ahmed, and Muhammad Hanif. A new robust ratio estimator with reference to nonnormal distribution. Communications in Statistics: Theory and Methods, 50(5):1099–1116, 2021. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2019.1646766.

Shah:2023:MLE

[SAH^{+}23] Muhammad Taqi Shah, Muhammad Azam, Muhammad Hanif, Muhammad Aslam, and Uzma Sherazi. Monitoring largest extreme observations using Frechet distribution based on weighted variance method. Communications in Statistics: Theory and Methods, 52(22):8136–8151, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2057544.

Sudheesh:2021:MTF

[SAK21] K. K. Sudheesh, G. Asha, and K. M. Jagathnath Krishna. On the mean time to failure of an age-replacement model in discrete time. Communications in Statistics: Theory and Methods, 50(11):2569–2585, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1672742.

Salamanca:2021:STM

[Sal21] Juan Jesús Salamanca. Symmetry tests for manifoldvalued random variables. Communications in Statistics: Theory and Methods, 50(1):61–72, 2021. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2019.1628990.

Sales:2024:SST

[Sal24] Adam C. Sales. Sequential specification tests to choose a model: a change-point approach. Communications in Statistics: Theory and Methods, 53(12):4354–4368, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2179879.

Shahzad:2021:EFM

[SANHS21] Usman Shahzad, Nadia H. Al-Noor, Muhammad Hanif, and Irsa Sajjad. An exponential family of median based estimators for mean estimation with simple random sampling scheme. Communications in Statistics: Theory and Methods, 50(20): 4890–4899, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2020.1725828.

Sharma:2020:OAD

[SAP20] Mahendra Kumar Sharma, Yabebal Ayalew, and Anshula Pandey. Orthogonal arrays and designs for partial triallel crosses. Communications in Statistics: Theory and Methods, 49(1):125–134, 2020. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2018.1532005.

Savas:2021:GIS

[Sav21] Ekrem Savas. On generalized invariant statistical convergence of weight g. Communications in Statistics: Theory and Methods, 50(8):1699–1708, 2021. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2019.1651863.

Sikolya:2020:ODP

[SB20] Kinga Sikolya and Sándor Baran. On the optimal designs for the prediction of complex Ornstein–Uhlenbeck processes. Communications in Statistics: Theory and Methods, 49(20): 4859–4870, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2019.1645855.

Samanta:2022:EAC

[SB22] S. K. Samanta and B. Bank. Extended analysis and computationally efficient results for the GI/M^{a,b} /1 queueing system. Communications in Statistics: Theory and Methods, 51(11): 3739–3760, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2020.1801739.

Stepien-Baran:2023:SEQ

[SB23a] Agnieszka Stepie’n-Baran. Sequential estimation of quantiles from delayed observations. Communications in Statistics: Theory and Methods, 52(6):1937–1945, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1942048.

Sujith:2023:AIG

[SB23b] P. Sujith and N. Balakrishna. Autoregressive inverse Gaussian process and the stochastic volatility modeling. Communications in Statistics: Theory and Methods, 52(10):3574–3580, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2021.1977324.

Singh:2021:CEP

[SBB21] G. N. Singh, D. Bhattacharyya, and A. Bandyopadhyay. Calibration estimation of population variance under stratified successive sampling in presence of random non response. Communications in Statistics: Theory and Methods, 50(19):4487– 4509, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2020.1719158.

Sattari:2022:SCF

[SBB22] Mostafa Sattari, Ghobad Barmalzan, and Narayanaswamy Balakrishnan. Stochastic comparisons of finite mixture models with generalized Lehmann distributed components. Communications in Statistics: Theory and Methods, 51(22):7767– 7782, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2021.1880592.

Singh:2024:RES

[SBB24] G. N. Singh, D. Bhattacharyya, and A. Bandyopadhyay. Robust estimation strategy for handling outliers. Communications in Statistics: Theory and Methods, 53(15):5311–5330, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2023.2218567.

Saadaoui:2023:LLE

[SBC23] Allal Saadaoui, Fadila Benaissa, and Abdelhak Chouaf. On the local linear estimation of a generalized regression function with spatial functional data. Communications in Statistics: Theory and Methods, 52(21):7752–7779, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2064499.

Slavtchova-Bojkova:2024:SMM

[SBHY24] Maroussia Slavtchova-Bojkova, Ollivier Hyrien, and Nikolay M. Yanev. Subcritical multitype Markov branching processes with immigration generated by Poisson random measures. Communications in Statistics: Theory and Methods, 53(14):5076–5091, 2024. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2023.2205972.

Seifollahi:2024:IRE

[SBK24] Solmaz Seifollahi, Hossein Bevrani, and Kaniav Kamary. Inequality restricted estimator for gamma regression: Bayesian approach as a solution to the multicollinearity. Communications in Statistics: Theory and Methods, 53(23):8297–8311, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2023.2281267.

Synowka-Bejenka:2021:NAL

[SBZ21] Ewa Synówka-Bejenka and Stefan Zontek. A note on admissibility of linear estimators in random models with a special structure. Communications in Statistics: Theory and Methods, 50(22):5422–5428, 2021. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2020.1734823.

Sills:2020:EDA

[SC20] Andrew V. Sills and Charles W. Champ. The exponential distribution analog of the Grubbs–Weaver method. Communications in Statistics: Theory and Methods, 49(8):1894–1903, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2019.1565839.

Shaw:2022:FIU

[SC22] Purnima Shaw and Arijit Chaudhuri. Further improvements on unrelated characteristic models in randomized response techniques. Communications in Statistics: Theory and Methods, 51(21):7305–7321, 2022. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2021.1872638.

Samawi:2023:MDA

[SCAK23] Hani Samawi, Ding-Geng Chen, Ferdous Ahmed, and Jing Kersey. Medical diagnostics accuracy measures and cutpoint selection: an innovative approach based on relative net benefit. Communications in Statistics: Theory and Methods, 52(14):5010–5025, 2023. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2021.2001016.

Singh:2024:CBE

[SCGB24] Saroja Kumar Singh, Frederico R. B. Cruz, Eriky S. Gomes, and Abhijit Datta Banik. Classical and Bayesian estimations of performance measures in a single server Markovian queueing system based on arrivals during service times. Communications in Statistics: Theory and Methods, 53(10):3517–3546, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2022.2155789.

Shestopaloff:2023:SMA

[SCP23] Konstantin Shestopaloff, Mayilee Canizares, and J. Denise Power. A sequential modeling approach for predicting clinical outcomes with repeated measures. Communications in Statistics: Theory and Methods, 52(20):7465–7478, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2047203.

Sharifipanah:2021:FFS

[SCPAV21] Najme Sharifipanah, Rahim Chinipardaz, Gholam Ali Parham, and Reinaldo Boris Arellano-Valle. Flexible families of symmetric and asymmetric distributions based on the two-piece skew normal distribution. Communications in Statistics: Theory and Methods, 50(10):2281–2305, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1664585.

Shanubhogue:2020:MCT

[SD20] Ashok Shanubhogue and Rajendra G. Desai. A method of constructing test for the parameter of family of lifetime distributions under Type-II censored sample. Communications in Statistics: Theory and Methods, 49(13):3267–3285, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1586944.

Sultana:2023:TRV

[SD23] Farha Sultana and Anup Dewanji. Tampered random variable modeling for multiple step-stress life test. Communications in Statistics: Theory and Methods, 52(15):5387–5406, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.2008440.

Singh:2024:DBE

[SDK^{+}24] Brijesh P. Singh, Utpal Dhar Das, Kadir Karakaya, Hassan S. Bakouch, and Badamasi Abba. Doubly bounded exponential model: Some information measures and estimation. Communications in Statistics: Theory and Methods, 53(22):7842– 7859, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2023.2273779.

Shih:2021:PCR

[SE21] Jia-Han Shih and Takeshi Emura. Penalized Cox regression with a five-parameter spline model. Communications in Statistics: Theory and Methods, 50(16):3749–3768, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1772305.

Sabbaghi:2022:RHD

[SE22] Azam Sabbaghi and Farzad Eskandari. A robust high dimensional estimation of a finite mixture of the generalized linear model. Communications in Statistics: Theory and Methods, 51(13):4451–4463, 2022. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2020.1815780.

Sanchez-Espigares:2021:MNT

[SEGMA21] José A. Sánchez-Espigares, Pere Grima, and Lluís Marco-Almagro. Mosaic normality test. Communications in Statistics: Theory and Methods, 50(23):5561–5573, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1734828.

Semeraro:2020:NMG

[Sem20] Patrizia Semeraro. A note on the multivariate generalized asymmetric Laplace motion. Communications in Statistics: Theory and Methods, 49(10):2339–2355, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1571609.

Shirazi:2024:NWB

[SF24a] Esmaeil Shirazi and Olivier P. Faugeras. A new wavelet-based estimation of conditional density via block threshold method. Communications in Statistics: Theory and Methods, 53(22): 8155–8165, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2023.2279917.

Sun:2024:MCP

[SF24b] Yang Sun and Xiangzhong Fang. A model calibration procedure for count response. Communications in Statistics: Theory and Methods, 53(12):4272–4289, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2177109.

Sun:2024:REF

[SF24c] Yang Sun and Xiangzhong Fang. Robust estimators of functional single index models for longitudinal data. Communications in Statistics: Theory and Methods, 53(19):6869–6890, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2023.2253939.

Sun:2022:ODCb

[SG22a] Xuebo Sun and Yingnan Guan. Optimal designs for collapsed homogeneous linear model. Communications in Statistics: Theory and Methods, 51(8):2303–2329, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1963777.

Sun:2022:ODCa

[SG22b] Xuebo Sun and Yingnan Guan. Optimal designs of collapsed Scheffé model. Communications in Statistics: Theory and Methods, 51(5):1155–1178, 2022. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2020.1870697.

Scotto:2023:EMI

[SG23] Manuel G. Scotto and Sónia Gouveia. On the extremes of the max-INAR(1) process for time series of counts. Communications in Statistics: Theory and Methods, 52(4):1136–1154, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2021.1923750.

Salvador:2024:AAB

[SG24] Sara Salvador and Riccardo Gatto. An algebraic analysis of the bimodality of the generalized von Mises distribution. Communications in Statistics: Theory and Methods, 53(10):3642– 3658, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2022.2158345.

Shabbir:2022:IGC

[SGM22] Javid Shabbir, Sat Gupta, and Saadia Masood. An improved general class of estimators for finite population mean in simple random sampling. Communications in Statistics: Theory and Methods, 51(11):3508–3520, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1797802.

Shabbir:2022:ICD

[SGN22] Javid Shabbir, Sat Gupta, and Ghulam Narjis. On improved class of difference type estimators for population median in survey sampling. Communications in Statistics: Theory and Methods, 51(10):3334–3354, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1795195.

Singh:2024:GCE

[SGTG24] Housila P. Singh, Anurag Gupta, Rajesh Tailor, and Neha Garg. A generalized class of estimators for the mean using multiauxiliary information in adaptive cluster sampling. Communications in Statistics: Theory and Methods, 53(22):7871– 7892, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2023.2274809.

Sun:2023:ASR

[SGW23] Huimin Sun, Bingzhen Geng, and Shijie Wang. Asymptotic sum-ruin probability for a bidimensional renewal risk model with subexponential claims. Communications in Statistics: Theory and Methods, 52(7):2057–2071, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1944215.

Sharifi:2023:RES

[SGZA23] Mani Sharifi, Pedram Pourkarim Guilani, Arash Zaretalab, and Abdolreza Abhari. Reliability evaluation of a system with active redundancy strategy and load-sharing time-dependent failure rate components using Markov process. Communications in Statistics: Theory and Methods, 52(13):4514–4533, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2021.1995433.

Susam:2023:GFI

[SH23] Selim Orhun Susam and Burcu Hudaverdi. A goodness-of fit improvement based on τ-preserving transformation for semiparametric family of copulas. Communications in Statistics: Theory and Methods, 52(21):7699–7708, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2052900.

Shariati:2021:RRC

[Sha21] Nima Shariati. Robust residual control chart for contaminated time series: a solution to the effects of outlier-driven parameter misestimation on the control chart performance. Communications in Statistics: Theory and Methods, 50(24):5738– 5758, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2020.1736303.

Shang:2022:GMM

[Sha22a] Wenpeng Shang. Generalized method of moment for casecohort under additive hazards model. Communications in Statistics: Theory and Methods, 51(10):3355–3381, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1795196.

Shanmugam:2022:DHL

[Sha22b] Ramalingam Shanmugam. Does how long observing correlate with upper record value? Fukushima nuclear disaster’s radiation levels are illustrated. Communications in Statistics: Theory and Methods, 51(13):4487–4499, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1815783.

Shaarawy:2023:BMF

[Sha23] Samir M. Shaarawy. Bayesian modeling and forecasting of vector autoregressive moving average processes. Communications in Statistics: Theory and Methods, 52(11):3795–3815, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2021.1980047.

Saidi:2024:ENR

[SHA24] Amel Saidi, Abdelghani Hamaz, and Ouerdia Arezki. Estimation in nonlinear random fields models of autoregressive type with random parameters. Communications in Statistics: Theory and Methods, 53(1):294–309, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2077962.

Sattari:2023:OSP

[SHB23] Mostafa Sattari, Abdin Haidari, and Ghobad Barmalzan. Orderings for series and parallel systems comprising heterogeneous new extended Weibull components. Communications in Statistics: Theory and Methods, 52(19):6778–6793, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2033267.

Shah:2020:CAD

[SHC20] Said Farooq Shah, Zawar Hussain, and Salman Arif Cheema. Combining answers to direct and indirect questions: An implementation of Kuk’s randomized response model. Communications in Statistics: Theory and Methods, 49(16):3933–3949, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2019.1593458.

Shenkman:2022:MID

[She22] Natalia Shenkman. Min-infinite divisibility of the bivariate Marshall–Olkin copulas. Communications in Statistics: Theory and Methods, 51(1):226–231, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1747080.

Sedighi:2023:MLS

[SHFD23] Tabassom Sedighi, Amin Hosseinian-Far, and Alireza Daneshkhah. Measuring local sensitivity in Bayesian inference using a new class of metrics. Communications in Statistics: Theory and Methods, 52(11):3581–3597, 2023. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2021.1977956.

Shi:2023:VPM

[Shi23a] Xuejun Shi. Viability property for multi-dimensional stochastic differential equation and its applications to comparison theorem. Communications in Statistics: Theory and Methods, 52(2):378–397, 2023. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2021.1914098.

Shirazi:2023:EEQ

[Shi23b] Esmaeil Shirazi. An empirical estimate of quantile density function in presence of censoring. Communications in Statistics: Theory and Methods, 52(8):2718–2734, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1959611.

Short:2023:BQP

[Sho23] Michael Short. On binomial quantile and proportion bounds: With applications in engineering and informatics. Communications in Statistics: Theory and Methods, 52(12):4183–4199, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2021.1986540.

Shoaee:2024:BIP

[Sho24] Shirin Shoaee. Bayesian inference for the parameters of mortality rate in the models of dependent lives with application in life insurance. Communications in Statistics: Theory and Methods, 53(3):995–1023, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2099719.

Shutoh:2021:ENT

[Shu21] Nobumichi Shutoh. Effect of nonnormality on tests for a mean vector with missing data under an elliptically contoured pattern-mixture model. Communications in Statistics: Theory and Methods, 50(19):4448–4469, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1716254.

Su:2023:PHC

[SHXW23] Xiaonan Su, Miao Han, Yu Xing, and Wei Wang. Pricing and hedging for correlation options with regime switching and common jump risk. Communications in Statistics: Theory and Methods, 52(18):6504–6524, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2031219.

Sills:2021:IPP

[Sil21] Andrew V. Sills. Integer partitions probability distributions. Communications in Statistics: Theory and Methods, 50(15): 3556–3563, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2019.1708396.

Sills:2023:RBD

[Sil23] Andrew V. Sills. A refinement of the binomial distribution using the quantum binomial theorem. Communications in Statistics: Theory and Methods, 52(2):294–308, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1912768.

Siray:2024:CRE

[Sir24] Gülesen Üstündag Siray. Consistent ridge estimation for replicated ultrastructural measurement error models. Communications in Statistics: Theory and Methods, 53(17):6116–6136, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2023.2239403.

Sharpe:2023:EPR

[SJ23] James Sharpe and Miguel A. Juárez. Estimation of the Pareto and related distributions — a reference-intrinsic approach. Communications in Statistics: Theory and Methods, 52(3): 523–542, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2021.1916826.

Shi:2021:RFC

[SJF21] Xuejun Shi, Ronglin Ji, and Qun Feng. Representation of filtration-consistent nonlinear expectation by gexpectation in general framework. Communications in Statistics: Theory and Methods, 50(24):5721–5737, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1734829.

Singh:2023:IIM

[SJP23] Garib Nath Singh, Ashok Kumar Jaiswal, and Awadhesh K. Pandey. Improved imputation methods for missing data in two-occasion successive sampling. Communications in Statistics: Theory and Methods, 52(6):2010–2029, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1944211.

Sun:2021:PAC

[SJT21] Yuqin Sun, Hong Jiang, and Yongge Tian. A prediction analysis in a constrained multivariate general linear model with future observations. Communications in Statistics: Theory and Methods, 50(2):345–357, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1634819.

Shoaee:2020:SAN

[SK20a] Shirin Shoaee and Esmaile Khorram. Survival analysis for a new compounded bivariate failure time distribution in shock and competing risk models via an EM algorithm. Communications in Statistics: Theory and Methods, 49(21):5123–5153, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2019.1614193.

Ko:2020:AAS

[sK20b] Feng shou Ko. An alternate approach for sample size determination in a multi-regional trial. Communications in Statistics: Theory and Methods, 49(4):997–1007, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1554133.

Ko:2020:AUA

[sK20c] Feng shou Ko. An approach to use of an adaptive procedure to clinical trials for molecularly heterogeneous subject selection at interim. Communications in Statistics: Theory and Methods, 49(2):421–429, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1543770.

Ko:2020:HMM

[sK20d] Feng shou Ko. How the mechanism of missing data on longitudinal biomarkers influences the survival analysis. Communications in Statistics: Theory and Methods, 49(23):5794–5809, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2019.1622724.

Singh:2020:BMN

[SK20e] Deepak Singh and Somesh Kumar. A bivariate mixture of negative binomial distributions and its applications. Communications in Statistics: Theory and Methods, 49(17):4162–4177, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2019.1595651.

Singh:2020:SIM

[SK20f] Garib Nath Singh and Mohd Khalid. Some imputation methods to compensate with non-response for estimation of population mean in two-occasion successive sampling. Communications in Statistics: Theory and Methods, 49(14):3329–3351, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2019.1586945.

Ko:2022:IAS

[sK22a] Feng shou Ko. The issue about sample size for survival analysis considering the interaction of unrecognized heterogeneity and treatment. Communications in Statistics: Theory and Methods, 51(19):6653–6666, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1864827.

Soodan:2022:SSQ

[SK22b] Bhavneet Singh Soodan and Rakesh Kumar. A single server queuing system with correlated reneging and feedback of served customers. Communications in Statistics: Theory and Methods, 51(18):6461–6475, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1861300.

Ko:2023:CMR

[sK23a] Feng shou Ko. Comparisons of a multi-regional trial for four or five regions by fixed effect model and random effect model about allocating sample size rationally into individual regions for a multi-regional trial. Communications in Statistics: Theory and Methods, 52(23):8241–8260, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2065019.

Ko:2023:IAE

[sK23b] Feng shou Ko. An issue about the efficacy for the time-to-event outcome based on accelerated failure time model with interaction of unrecognized heterogeneity and main effect. Communications in Statistics: Theory and Methods, 52(3):702– 713, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2021.1921213.

Ko:2024:STS

[sK24a] Feng shou Ko. The Simon’s two-stage design accounting for genetic heterogeneity. Communications in Statistics: Theory and Methods, 53(7):2661–2669, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2148469.

Simsek:2024:AML

[SK24b] Güven Simsek and Duru Karasoy. Applying machine learning techniques in survival analysis to the private pension system in Turkey. Communications in Statistics: Theory and Methods, 53(16):5706–5720, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2230329.

Singh:2024:WGE

[SK24c] Shivangi Singh and Chanchal Kundu. On weighted generalized entropy for double truncated distribution with applications. Communications in Statistics: Theory and Methods, 53(9): 3102–3122, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2022.2150821.

Shih:2023:CGP

[SKCE23] Jia-Han Shih, Yoshihiko Konno, Yuan-Tsung Chang, and Takeshi Emura. A class of general pretest estimators for the univariate normal mean. Communications in Statistics: Theory and Methods, 52(8):2538–2561, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1955384.

Sun:2023:MHN

[SKP23] Jingchao Sun, Maiying Kong, and Subhadip Pal. The Modified-Half-Normal distribution: Properties and an efficient sampling scheme. Communications in Statistics: Theory and Methods, 52(5):1591–1613, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1934700.

Skrobotov:2023:LRT

[Skr23] Anton Skrobotov. Likelihood ratio test for change in persistence. Communications in Statistics: Theory and Methods, 52(17):5952–5965, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03610926.2022.2055070.

Spangl:2023:MSS

[SKRR23] Bernhard Spangl, Norbert Kaiblinger, Peter Ruckdeschel, and Dieter Rasch. Minimal sample size in balanced ANOVA models of crossed, nested, and mixed classifications. Communications in Statistics: Theory and Methods, 52(6):1728–1743, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2021.1938126.

Smitha:2024:DCR

[SKS24] S. Smitha, Sudheesh K. Kattumannil, and E. P. Sreedevi. Dynamic cumulative residual entropy generating function and its properties. Communications in Statistics: Theory and Methods, 53(16):5890–5909, 2024. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2023.2235448.

Sun:2020:LIN

[SL20] Huihui Sun and Qiang Liu. Local influence of nonlinear mixed effects model based on M-estimation. Communications in Statistics: Theory and Methods, 49(21):5342–5355, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1618474.

Song:2024:CVP

[SL24a] Shiyu Song and Ying Lu. Computation of VaR for portfolios in intensity models. Communications in Statistics: Theory and Methods, 53(16):5910–5923, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2237221.

Sun:2024:RVS

[SL24b] Huihui Sun and Qiang Liu. Robust variable selection in semiparametric mixed effects longitudinal data models. Communications in Statistics: Theory and Methods, 53(3):1049–1064, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2022.2100421.

Slaoui:2022:RKR

[Sla22] Yousri Slaoui. Recursive kernel regression estimation under α-mixing data. Communications in Statistics: Theory and Methods, 51(24):8459–8475, 2022. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2021.1897842.

Slaoui:2023:TTS

[Sla23] Yousri Slaoui. Two-time-scale nonparametric recursive regression estimator for independent functional data. Communications in Statistics: Theory and Methods, 52(15):5213–5245, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2021.2004428.

Shi:2022:SGS

[SLF^{+}22] Zhiyan Shi, Cong Liu, Yan Fan, Dan Bao, and Yang Chen. Some generalized strong limit theorems for Markov chains in bi-infinite random environments. Communications in Statistics: Theory and Methods, 51(1):150–161, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1744655.

Sheng:2021:MPF

[SLS21] De-Lei Sheng, Danping Li, and Peilong Shen. Minimum probability function of crossing the upper regulatory threshold for asset-liability management. Communications in Statistics: Theory and Methods, 50(23):5530–5553, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1734824.

Saranya:2023:ISR

[SLS23] N. Saranya, A. Shophia Lawrence, and B. Sivakumar. An inventory system with replacement and refurbishment of failed items. Communications in Statistics: Theory and Methods, 52(17):5966–5988, 2023. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2021.2023184.

Saulo:2020:MBB

[SLV^{+}20] Helton Saulo, Jeremias Leão, Roberto Vila, Victor Leiva, and Vera Tomazella. On mean-based bivariate Birnbaum– Saunders distributions: Properties, inference and application. Communications in Statistics: Theory and Methods, 49(24): 6032–6056, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2019.1626425.

Shi:2022:NSQ

[SLX22] Jianhua Shi, Yutao Liu, and Jinfeng Xu. Nonparametric smoothed quantile difference estimation for lengthbiased and right-censored data. Communications in Statistics: Theory and Methods, 51(10):3237–3252, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1791340.

Sun:2023:LWA

[SLYF23] Jun Sun, Wanrong Liu, Jing Yang, and Jianglin Fang. Local Walsh-average regression for single index varying coefficient models. Communications in Statistics: Theory and Methods, 52(23):8306–8316, 2023. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2022.2060514.

Silva:2020:FBM

[SM20] Ivair R. Silva and Reinaldo Marques. Frequentist–Bayesian Monte Carlo testing. Communications in Statistics: Theory and Methods, 49(10):2356–2364, 2020. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2019.1571610.

Sadok:2022:NPS

[SM22] Ibrahim Sadok and Afif Masmoudi. New parametrization of stochastic volatility models. Communications in Statistics: Theory and Methods, 51(7):1936–1953, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1934031.

Smaga:2021:GLH

[Sma21] Lukasz Smaga. General linear hypothesis testing in functional response model. Communications in Statistics: Theory and Methods, 50(21):5068–5083, 2021. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2019.1691233.

Smaga:2024:PTL

[Sma24] Lukasz Smaga. Projection tests for linear hypothesis in the functional response model. Communications in Statistics: Theory and Methods, 53(4):1402–1419, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2101120.

Sharma:2022:HHT

[SMH22] Nidhi Sharma, Rohan Mishra, and Abdelouahed Hamdi. Hermite-Hadamard type integral inequalities for multidimensional general h-harmonic preinvex stochastic processes. Communications in Statistics: Theory and Methods, 51(19):6719– 6740, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2020.1865403.

Sharma:2024:SGC

[SMH24] Nidhi Sharma, Rohan Mishra, and Abdelouahed Hamdi. On strongly generalized convex stochastic processes. Communications in Statistics: Theory and Methods, 53(8):2908–2923, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2022.2150055.

Shongwe:2021:NSR

[SMMC21] Sandile Charles Shongwe, Jean-Claude Malela-Majika, and Philippe Castagliola. The new synthetic and runs-rules schemes to monitor the process mean of autocorrelated observations with measurement errors. Communications in Statistics: Theory and Methods, 50(24):5806–5835, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1737125.

Shongwe:2020:SSS

[SMMCM20] Sandile Charles Shongwe, Jean-Claude Malela-Majika, Philippe Castagliola, and Thapelo Molahloe. Side-sensitive synthetic and runs-rules charts for monitoring AR(1) processes with skipping sampling strategies. Communications in Statistics: Theory and Methods, 49(17):4248–4269, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1596284.

Seijas-Macias:2023:SKP

[SMOO23] Antonio Seijas-Macias, Amílcar Oliveira, and Teresa A. Oliveira. The skewness and kurtosis of the product of two normally distributed random variables. Communications in Statistics: Theory and Methods, 52(1):80–93, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1909734.

Sinha:2024:SPE

[SMP^{+}24] Arup K. Sinha, Lemuel Moye Iii, Linda B. Piller, Jose-Miguel Yamal, Carlos H. Barcenas, Jaejoon Song, and Barry R. Davis. Simultaneous population enrichment and endpoint selection in phase 3 randomized controlled trials: an adaptive group sequential design with two binary alternative primary endpoints.

Communications in Statistics: Theory and Methods, 53(10): 3728–3741, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2022.2163180.

Sapam:2021:LSD

[SMS21] Sobita Sapam, Nripes Mandal, and Bikas Sinha. Latin square designs with neighbor effects — Part II. Communications in Statistics: Theory and Methods, 50(14):3371–3379, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1702694.

Sivula:2023:UEV

[SMV23] Tuomas Sivula, Måns Magnusson, and Aki Vehtari. Unbiased estimator for the variance of the leave-one-out cross-validation estimator for a Bayesian normal model with fixed variance. Communications in Statistics: Theory and Methods, 52(16): 5877–5899, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2021.2021240.

Sadok:2023:IEA

[SMZ23] Ibrahim Sadok, Afif Masmoudi, and Mourad Zribi. Integrating the EM algorithm with particle filter for image restoration with exponential dispersion noise. Communications in Statistics: Theory and Methods, 52(2):446–462, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1915336.

Sathar:2023:SWD

[SN23] E. I. Abdul Sathar and R. Dhanya Nair. A study on weighted dynamic survival and failure extropies. Communications in Statistics: Theory and Methods, 52(3):623–642, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1919308.

Si:2020:DQR

[SNS20] Yuancheng Si, Saralees Nadarajah, and Xiaodong Song. On the distribution of quotient of random variables conditioned to the positive quadrant. Communications in Statistics: Theory and Methods, 49(10):2514–2528, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1576893.

Silva:2022:CCI

[SO22] Ivair R. Silva and Dionatan W. R. Oliveira. Confidencecredible intervals. Communications in Statistics: Theory and Methods, 51(9):2783–2802, 2022. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2020.1780447.

Savas:2022:SAT

[SP22] Rabia Savas and Richard F. Patterson. Summability approximation theorems of double random variables. Communications in Statistics: Theory and Methods, 51(24):8617–8624, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2021.1901920.

Seung:2023:MLR

[SP23] Hyunseok Seung and Sangun Park. Modified likelihood ratio tests for extreme value distributions. Communications in Statistics: Theory and Methods, 52(16):5742–5751, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.2018463.

Santos:2024:NEB

[SP24a] Mara C. T. Santos and Rodrigo R. Pescim. A new extension of the Burr XII distribution generated by odd log-logistic random variables. Communications in Statistics: Theory and Methods, 53(14):5003–5017, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2200560.

Sebastian:2024:ESD

[SP24b] Nicy Sebastian and T. Princy. Exact sampling distribution of the general case sample correlation matrix. Communications in Statistics: Theory and Methods, 53(18):6380–6393, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2023.2244622.

Spezia:2024:BPM

[Spe24] Luigi Spezia. Bayesian prior modeling in vector autoregressions via the Yule–Walker equations. Communications in Statistics: Theory and Methods, 53(14):5230–5247, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2214827.

Singh:2021:SIA

[SPS21] Garib Nath Singh, Awadhesh K. Pandey, and Anup Kumar Sharma. Some improved and alternative imputation methods for finite population mean in presence of missing information. Communications in Statistics: Theory and Methods, 50(19): 4401–4427, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2020.1713375.

Singh:2023:SDC

[SPS23] Arpan Singh, Meghendar Pal, and Satya Prakash Singh. Some design considerations for cluster randomized trials with binary responses. Communications in Statistics: Theory and Methods, 52(12):4310–4328, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1990954.

Singh:2020:IAE

[SPSS20] Garib N. Singh, Awadhesh K. Pandey, Chandraketu Singh, and Surbhi Suman. An improved alternative estimation procedure for current population mean in presence of positively and negatively correlated auxiliary variables in twooccasion rotation patterns. Communications in Statistics: Theory and Methods, 49(13):3094–3125, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1586932.

Shekari:2023:SCP

[SPZ23] Marzieh Shekari, Zohreh Pakdaman, and Hossein Zamani. Stochastic comparisons of parallel systems with heterogeneous exponentiated half logistic-F components. Communications in Statistics: Theory and Methods, 52(1):183–195, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1910842.

Shekari:2024:SCS

[SPZ24] Marzieh Shekari, Zohreh Pakdaman, and Hossein Zamani. Stochastic comparisons of series systems with heterogeneous Gompertz-G components. Communications in Statistics: Theory and Methods, 53(14):5146–5167, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2209228.

Sathar:2021:DSE

[SR21] E. I. Abdul Sathar and Dhanya Nair R. On dynamic survival extropy. Communications in Statistics: Theory and Methods, 50(6):1295–1313, 2021. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2019.1649426.

Sarabia:2020:BDT

[SRA20] José María Sarabia, A. Vincent Raja, and G. Asha. Bivariate distributions with transmuted conditionals: Models and applications. Communications in Statistics: Theory and Methods, 49(1):221–242, 2020. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2018.1536785.

Solci:2020:ESR

[SRSB20] Carlo Corrêa Solci, Valdério Anselmo Reisen, Alessandro José Queiroz Sarnaglia, and Pascal Bondon. Empirical study of robust estimation methods for PAR models with application to the air quality area. Communications in Statistics: Theory and Methods, 49(1):152–168, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1533970.

Shaghaghi:2020:PLR

[SS20a] M. Shaghaghi and A. Saghaei. PCA likelihood ratio test approach for attributed social networks monitoring. Communications in Statistics: Theory and Methods, 49(12):2869–2886, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2018.1491599.

Shen:2020:ELR

[sS20b] Pao sheng Shen. Empirical likelihood ratio confidence intervals in terms of cumulative hazard function for lefttruncated and right-censored data. Communications in Statistics: Theory and Methods, 49(18):4571–4586, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1604960.

Srivastav:2021:EBD

[SS21a] Sudesh K. Srivastav and Apurv Srivastav. Efficiency balanced designs for bootstrap simulations. Communications in Statistics: Theory and Methods, 50(19):4510–4527, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1719417.

Sunethra:2021:NMJ

[SS21b] A. A. Sunethra and M. R. Sooriyarachchi. A novel method for joint modeling of survival data and count data for both simple randomized and cluster randomized data. Communications in Statistics: Theory and Methods, 50(18):4180–4202, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2020.1713366.

Sagir:2022:OAT

[SS22a] Murat Sagir and Vedat Saglam. Optimization and analysis of a tandem queueing system with parallel channel at second station. Communications in Statistics: Theory and Methods, 51(21):7547–7560, 2022. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2021.1875240.

Salamwade:2022:RVS

[SS22b] R. L. Salamwade and D. M. Sakate. Robust variable selection via penalized MT-estimator in generalized linear models. Communications in Statistics: Theory and Methods, 51(22): 8053–8065, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2021.1887240.

Samanthi:2022:BKD

[SS22c] Ranadeera Gamage Madhuka Samanthi and Jungsywan Sepanski. On bivariate Kumaraswamy-distorted copulas. Communications in Statistics: Theory and Methods, 51(8):2477–2495, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2020.1777303.

Sarada:2022:SMR

[SS22d] Y. Sarada and S. Sangeetha. A δ-shock model for reengineering of a repairable supply chain using quasi renewal process. Communications in Statistics: Theory and Methods, 51(18): 6476–6501, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2020.1861466.

Saurabh:2022:NCL

[SS22e] Shyam Saurabh and Mithilesh Kumar Singh. A note on the construction of Latin square type designs. Communications in Statistics: Theory and Methods, 51(10):3434–3437, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1734837.

Saurabh:2022:SNR

[SS22f] Shyam Saurabh and Kishore Sinha. Some new resolvable group divisible designs. Communications in Statistics: Theory and Methods, 51(13):4509–4514, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1817487.

Sayyareh:2022:DFT

[SS22g] Abdolreza Sayyareh and Saba Sayyareh. A distributionfree tracking interval for model selection: application in the strength of brittle materials. Communications in Statistics: Theory and Methods, 51(16):5607–5637, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1843681.

Schilling:2022:NAP

[SS22h] Mark F. Schilling and Alyssa Stanley. A new approach to precise interval estimation for the parameters of the hypergeometric distribution. Communications in Statistics: Theory and Methods, 51(1):29–50, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1737879.

Sharifnia:2022:SAS

[SS22i] S. Golshid Sharifnia and Abbas Saghaei. A statistical approach for social network change detection: an ERGM based framework. Communications in Statistics: Theory and Methods, 51(7):2259–2280, 2022. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2020.1772981.

Shen:2022:CAM

[sS22j] Pao sheng Shen. The Cox–Aalen model for doubly censored data. Communications in Statistics: Theory and Methods, 51(23):8075–8092, 2022. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2021.1887241.

Singh:2022:STS

[SS22k] Garib Nath Singh and Chandraketu Singh. A stratified threestage randomized response model for estimation of rare sensitive parameter under Poisson approximation. Communications in Statistics: Theory and Methods, 51(19):6626–6652, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2020.1864826.

Saurabh:2023:LNP

[SS23a] Shyam Saurabh and Kishore Sinha. A list of new partially balanced designs. Communications in Statistics:

Theory and Methods, 52(23):8607–8610, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2059685.

Shat:2023:OTP

[SS23b] Helmi Shat and Rainer Schwabe. Optimal time plan in accelerated degradation testing. Communications in Statistics: Theory and Methods, 52(23):8408–8424, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2063892.

Song:2023:DRM

[SS23c] Zhan-Jie Song and Fu-Yun Sun. The dual risk model under a mixed ratcheting and periodic dividend strategy. Communications in Statistics: Theory and Methods, 52(10):3526–3540, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2021.1974483.

Shat:2024:EDA

[SS24a] Helmi Shat and Rainer Schwabe. Experimental designs for accelerated degradation tests based on linear mixed effects models. Communications in Statistics: Theory and Methods, 53(6):2154–2177, 2024. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2022.2121612.

Shi:2024:PEU

[SS24b] Yuxin Shi and Yuhong Sheng. The parameter estimations for uncertain regression model with autoregressive time series errors. Communications in Statistics: Theory and Methods, 53(13):4841–4856, 2024. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2023.2195034.

Smithson:2024:FCQ

[SS24c] Michael Smithson and Yiyun Shou. Flexible CDF-quantile distributions on the closed unit interval, with software and applications. Communications in Statistics: Theory and Methods, 53(11):3876–3898, 2024. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2023.2166352.

Saleem:2023:DNM

[SSA23] Sehar Saleem, Rehan Ahmad Khan Sherwani, and Muhammad Amin. Development of a new modified Hogg type adaptive scheme for multilevel models with diverse error distributions. Communications in Statistics: Theory and Methods, 52(8):2735–2750, 2023. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2021.1959612.

Sharifian:2021:JML

[SSG21] Nastaran Sharifian, Ehsan Bahrami Samani, and Mojtaba Ganjali. Joint modeling for longitudinal set-inflated continuous and count responses. Communications in Statistics: Theory and Methods, 50(5):1134–1160, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1646768.

Saeki:2023:TPC

[SSH23] Yuichiro Saeki, Takashi Seo, and Hiroto Hyakutake. Testing parallelism and confidence intervals of level difference in an intraclass correlation model with monotone missing data. Communications in Statistics: Theory and Methods, 52(17):6147– 6160, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2022.2026961.

Shen:2020:NAR

[sSL20] Pao sheng Shen and Yu-Hsing Lai. Nonparametric analysis of recurrent gap time data. Communications in Statistics: Theory and Methods, 49(13):3298–3312, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1588322.

Sun:2021:NPC

[SSM21] Jing Sun, Deepak Sakate, and Sunil Mathur. A nonparametric procedure for changepoint detection in linear regression. Communications in Statistics: Theory and Methods, 50(8):1925– 1935, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2019.1657453.

Singh:2020:SCE

[SSNP20] Dhirendra Singh, Bhupendra Veer Singh Sisodia, Nidhi, and Sandeep Pundir. Some calibration estimators for finite population mean in two-stage stratified random sampling. Communications in Statistics: Theory and Methods, 49(17):4234–4247, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2019.1596283.

Sanaullah:2020:USR

[SSS20] Aamir Sanaullah, Iram Saleem, and Javid Shabbir. Use of scrambled response for estimating mean of the sensitivity variable. Communications in Statistics: Theory and Methods, 49(11):2634–2647, 2020. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2019.1576898.

Salinas:2022:HOC

[SSS22] Veronica I. Salinas, Stephen A. Sedory, and Sarjinder Singh. Higher order calibrated estimator in two-stage sampling. Communications in Statistics: Theory and Methods, 51(10): 3164–3180, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2020.1790005.

Saurabh:2023:SRS

[SSS23] Shyam Saurabh, Kishore Sinha, and Mithilesh Kumar Singh. A survey of resolvable solutions of partially balanced designs. Communications in Statistics: Theory and Methods, 52(6): 1946–1962, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2021.1942049.

Song:2022:ASC

[SST22] Wenyi Song, Jiamin Shao, and Zhongquan Tan. Almost sure central limit theorems for the maxima of Gaussian functions. Communications in Statistics: Theory and Methods, 51(20): 7136–7147, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2020.1871019.

Su:2020:EGS

[SSW20] Wen Su, Benxuan Shi, and Yunyun Wang. Estimating the Gerber–Shiu function under a risk model with stochastic income by Laguerre series expansion. Communications in Statistics: Theory and Methods, 49(23):5686–5708, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1620782.

Song:2022:VRA

[SSZC22] Yuping Song, Zheng Sun, Qicheng Zhao, and Youyou Chen. Variance reduction approach for the volatility over a finite-time horizon. Communications in Statistics: Theory and Methods, 51(11):3521–3541, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1797803.

Sharma:2021:OPT

[ST21] Mahendra Kumar Sharma and Mekonnen Tadesse. Optimal partial triallel cross designs through diallel cross designs. Communications in Statistics: Theory and Methods, 50(15): 3547–3555, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2019.1705982.

Shao:2023:LPL

[ST23a] Jiamin Shao and Zhongquan Tan. On the limit properties of the last exit time and the first crossing point for the stationary dependent chi-sequences. Communications in Statistics: Theory and Methods, 52(4):1237–1250, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1926508.

Singh:2023:TSS

[ST23b] Housila P. Singh and Tanveer A. Tarray. Two-stage stratified partial randomized response strategies. Communications in Statistics: Theory and Methods, 52(14):4862–4893, 2023.

CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2013.804571.

Saber:2024:CNS

[ST24] Mohammad Mehdi Saber and Mehrdad Taghipour. A crucial note on stress-strength models: Wrong asymptotic variance in some published papers. Communications in Statistics: Theory and Methods, 53(7):2422–2429, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2134731.

Sue:2023:DSE

[STC23] Kuen-Liang Sue, Chih-Fong Tsai, and Andy Chiu. The data sampling effect on financial distress prediction by single and ensemble learning techniques. Communications in Statistics: Theory and Methods, 52(12):4344–4355, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1992439.

Sarkar:2024:EPG

[STK24] Mojammel Haque Sarkar, Manas Ranjan Tripathy, and Debasis Kundu. Estimating parameters from the generalized inverse Lindley distribution under hybrid censoring scheme. Communications in Statistics: Theory and Methods, 53(7): 2483–2515, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2022.2138436.

Shen:2023:ASD

[STW23] Guangjun Shen, Zheng Tang, and Jun Wang. Approximation of stochastic differential equations driven by subfractional Brownian motion at discrete time observation. Communications in Statistics: Theory and Methods, 52(1):1–18, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1901924.

Singh:2022:ECQ

[SU22] G. N. Singh and M. Usman. Estimation of coefficient of quartile deviation in case of missing data. Communications in Statistics: Theory and Methods, 51(22):8027–8052, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1887239.

Suda:2022:SSP

[Sud22] David Suda. Stability of sampling proposals for reducible diffusions over large time intervals. Communications in Statistics: Theory and Methods, 51(18):6166–6181, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1856876.

Sulewski:2021:TPP

[Sul21] Piotr Sulewski. Two-piece power normal distribution. Communications in Statistics: Theory and Methods, 50(11):2619– 2639, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2019.1674871.

Sulewski:2022:NDP

[Sul22] Piotr Sulewski. Normal distribution with plasticizing component. Communications in Statistics: Theory and Methods, 51(11):3806–3835, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03610926.2020.1837881.

Sun:2023:CMC

[Sun23] Yang Sun. Consistent model checking procedures for parametric regressions with missing response. Communications in Statistics: Theory and Methods, 52(1):208–226, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1910844.

Singh:2021:IPE

[SV21] Abhishek Singh and Gajendra K. Vishwakarma. Improved predictive estimation for mean using the Searls technique under ranked set sampling. Communications in Statistics: Theory and Methods, 50(9):2015–2038, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1657456.

Singh:2022:CEM

[SV22] Neha Singh and Gajendra K. Vishwakarma. Computing the effect of measurement errors on the use of auxiliary information under systematic sampling. Communications in Statistics: Theory and Methods, 51(20):6919–6934, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1868513.

Sathar:2021:EPI

[SVR21] E. I. Abdul Sathar, K. V. Viswakala, and G. Rajesh. Estimation of past inaccuracy measure for the right censored dependent data. Communications in Statistics: Theory and Methods, 50(6):1446–1455, 2021. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2019.1651859.

Song:2021:FRA

[SW21] Pingfan Song and Shaochen Wang. A further remark on the alternative expectation formula. Communications in Statistics: Theory and Methods, 50(11):2586–2591, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1672743. See remark [Far24c].

Su:2023:NLM

[SW23]  Meihong Su and Wenjian Wang.  A network Lasso

model for regression. Communications in Statistics: Theory and Methods, 52(6):1702–1727, 2023. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2021.1938125.

Shen:2024:CCM

[SW24] Aiting Shen and Rui Wang. Complete and complete moment convergence for arrays of rowwise asymptotically almost negatively associated random variables. Communications in Statistics: Theory and Methods, 53(10):3818–3832, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2164466.

Swanepoel:2024:CBS

[Swa24] Jan W. H. Swanepoel. Construction of bivariate symmetric and asymmetric copulas and its relationship to ratios of conditional hazard rate functions. Communications in Statistics: Theory and Methods, 53(14):5280–5296, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2218505.

Shu:2022:SLC

[SWHW22] Ming-Hung Shu, Chien-Wei Wu, Bi-Min Hsu, and To-Cheng Wang. Standardized lifetime-capability and warranty-return-rate-based suppliers qualification and selection with accelerated Weibull-life type II testing data. Communications in Statistics: Theory and Methods, 51(23):8186–8204, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1890124.

Sheng:2024:NRM

[SWK24] Danshu Sheng, Dehui Wang, and Yao Kang. A new RCAR(1) model based on explanatory variables and observations. Communications in Statistics: Theory and Methods, 53(7):2285– 2306, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2022.2125267.

Sun:2023:BAM

[SWLZ23] Yuqin Sun, Yawen Wang, Yan Li, and Wei Zhu. Bayesian analysis of multiple break-points threshold ARMA model with exogenous inputs. Communications in Statistics: Theory and Methods, 52(24):8677–8695, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2068030.

Song:2024:TSS

[SWW24] Quanhong Song, Lichun Wang, and Liqun Wang. Two-stage shrunken least squares estimator and its superiority. Communications in Statistics: Theory and Methods, 53(19):6717– 6731, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2023.2250487.

Sun:2020:NSA

[SWYF20] Bin Sun, Yuehua Wu, Wenzhi Yang, and Yuejiao Fu. A note on the semiparametric approach to dimension reduction. Communications in Statistics: Theory and Methods, 49(9):2295– 2304, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2019.1576887.

Sun:2023:HDE

[SX23] Gaoming Sun and Junshan Xie. High-dimensional Edgeworth expansion of LR statistic for testing block circular symmetry covariance structure and its errors. Communications in Statistics: Theory and Methods, 52(24):8636–8657, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2067877.

Song:2024:CIC

[SXS24] Kai Song, Xiaoyue Xie, and Jian Shi. Confidence interval construction in massive data sets. Communications in Statistics: Theory and Methods, 53(3):1035–1048, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2100420.

Shi:2024:GEE

[SXZ24] Zhiyan Shi, Xinyue Xi, and Qingpei Zang. The generalized entropy ergodic theorem for Markov chains indexed by a spherically symmetric tree. Communications in Statistics: Theory and Methods, 53(6):2178–2193, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2122839.

Silahtaroglu:2021:DAH

[SY21] Gökhan Silahtaroglu and Nevin Yilmaztürk. Data analysis in health and big data: a machine learning medical diagnosis model based on patients’ complaints. Communications in Statistics: Theory and Methods, 50(7):1547–1556, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1622728.

Shi:2023:ERB

[SY23] Yufeng Shi and Zhi Yang. Existence result for the BSDE with superquadratic growth. Communications in Statistics: Theory and Methods, 52(24):8902–8908, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2105363.

Singh:2021:GFE

[SYG21] Housila P. Singh, Anita Yadav, and Swarangi M. Gorey. A generalized family of estimators for estimating finite population mean in circular systematic sampling (CSS). Communications in Statistics: Theory and Methods, 50(3):631–662, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1639750.

Shen:2023:UAR

[SYL23] Xinmei Shen, Meng Yuan, and Dawei Lu. Uniform asymptotics for ruin probabilities of multidimensional risk models with stochastic returns and regular variation claims. Communications in Statistics: Theory and Methods, 52(19):6878– 6895, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2022.2034868.

Shi:2024:SLL

[SYM24] Jianan Shi, Zhenhong Yu, and Yu Miao. On the strong laws of large numbers for pairwise NQD random variables. Communications in Statistics: Theory and Methods, 53(13):4745–4754, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2023.2189498.

Syring:2023:RPI

[Syr23] Nicholas Syring. Robust posterior inference for Youden’s index cutoff. Communications in Statistics: Theory and Methods, 52(10):3193–3208, 2023. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2021.1969409.

Samawi:2020:RSS

[SYRV20] Hani M. Samawi, Lili Yu, Haresh Rochani, and Robert Vogel. Reducing sample size needed for Cox-proportional hazards model analysis using more efficient sampling method. Communications in Statistics: Theory and Methods, 49(6):1281–1298, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2018.1554141.

Shi:2021:EPB

[SYT21] Zhiyan Shi, Weiguo Yang, and Ying Tang. Equivalent properties for the bifurcating Markov chains indexed by a binary tree. Communications in Statistics: Theory and Methods, 50(24): 6264–6272, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2020.1742923.

Shinoda:2024:DMS

[SYT24] Satoru Shinoda, Kouji Yamamoto, and Sadao Tomizawa. Decomposition of measure from symmetry for analyzing collapsed ordinal square contingency tables. Communications in Statistics: Theory and Methods, 53(16):5814–5827, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2233152.

Shan:2023:OPU

[SYZS23] Yuanchuang Shan, Haoran Yi, Xuekang Zhang, and Huisheng Shu. Option pricing under a Markov-modulated Merton jump-diffusion dividend. Communications in Statistics: Theory and Methods, 52(5):1490–1506, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1928205.

Song:2020:SCP

[SZ20a] Mingzhu Song and Quanxin Zhu. The strong convergence properties of weighted sums for a class of dependent random variables. Communications in Statistics: Theory and Methods, 49(14):3455–3465, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1589516.

Song:2020:NER

[SZ20b] Shanshan Song and Yong Zhou. Nonparametric estimation of the ROC curve for length-biased and right-censored data. Communications in Statistics: Theory and Methods, 49(19): 4648–4668, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2019.1604963.

Sun:2024:JMA

[SZ24] Xianwen Sun and Lixin Zhang. Jackknife model averaging for additive expectile prediction. Communications in Statistics: Theory and Methods, 53(19):6799–6831, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2251625.

Song:2021:BAC

[SZGK21] Guofeng Song, Lixun Zhu, Ai Gao, and Lingzhu Kong. Blockwise AIC_{c} and its consistency properties in model selection. Communications in Statistics: Theory and Methods, 50(13): 3198–3213, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2019.1691734.

Sun:2021:BAB

[SZL21] Zhongyang Sun, Xin Zhang, and Ya-Nan Li. A BSDE approach for bond pricing under interest rate models with self-exciting jumps. Communications in Statistics: Theory and Methods, 50(14):3249–3261, 2021. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2019.1691234.

Sun:2021:BTS

[SZW21] Nan Sun, Junjiang Zhong, and Miin-Jye Wen. Bayesian twostage design for drug screening trials with switching hypothesis tests based on continuous endpoints. Communications in Statistics: Theory and Methods, 50(2):415–431, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1635702.

Thomas:2022:ESP

[TA22] P. Yageen Thomas and V. Anjana. Estimation of the scale parameter of a family of distributions using a newly derived minimal sufficient statistic. Communications in Statistics: Theory and Methods, 51(22):7929–7962, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1884721.

Tahmasebi:2020:WEG

[Tah20] Saeid Tahmasebi. Weighted extensions of generalized cumulative residual entropy and their applications. Communications in Statistics: Theory and Methods, 49(21):5196–5219, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1615094.

Tan:2022:THD

[Tan22a]  Xiangyong Tan.  Test for high dimensional partially

linear models. Communications in Statistics: Theory and Methods, 51(18):6423–6434, 2022. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2020.1861297.

Tanaka:2022:BSV

[Tan22b] Masahiro Tanaka. Bayesian singular value regularization via a cumulative shrinkage process. Communications in Statistics: Theory and Methods, 51(16):5566–5589, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1843055.

Tang:2022:UTA

[Tan22c] Han Tang. Uncertain threshold autoregressive model with imprecise observations. Communications in Statistics: Theory and Methods, 51(24):8776–8785, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1906433.

Tanoue:2024:CIS

[Tan24] Yuta Tanoue. Concentration inequality of sums of dependent subexponential random variables and application to bounds for value-at-risk. Communications in Statistics: Theory and Methods, 53(9):3123–3142, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2150822.

Tapp:2022:CES

[Tap22] Kristopher Tapp. Clustering and expected seat-share for district maps. Communications in Statistics: Theory and Methods, 51(22):7751–7766, 2022. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2021.1879862.

Tasias:2024:SOI

[Tas24] Konstantinos A. Tasias. Simultaneous optimization of inventory, maintenance, and quality for production systems subject to multiple mean and variance shifts. Communications in Statistics: Theory and Methods, 53(9):3078–3101, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2150522.

Tian:2022:MRP

[TB22] Linlin Tian and Lihua Bai. Minimizing ruin probability under the sparre Anderson model. Communications in Statistics: Theory and Methods, 51(6):1622–1636, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1931887.

Tiwari:2024:ACS

[TBCB24] Neeraj Tiwari, Jharna Banerjie, Girish Chandra, and Shailja Bhari. Adaptive cluster sampling based on balanced sampling plan excluding contiguous units. Communications in Statistics: Theory and Methods, 53(15):5364–5377, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2220448.

Torres:2020:EED

[TDLL20] Raúl Torres, Elena Di Bernardino, Henry Laniado, and Rosa E. Lillo. On the estimation of extreme directional multivariate quantiles. Communications in Statistics: Theory and Methods, 49(22):5504–5534, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1619770.

Temizsu:2022:SBO

[TEÇK22] Fatih Temizsu, Mikail Et, Muhammed Çinar, and Hacer Sengül Kandemir. On (λ, f)-statistical boundedness of order α. Communications in Statistics: Theory and Methods, 51(19):6766– 6776, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2020.1866207.

Teerapabolarn:2022:NUB

[Tee22] K. Teerapabolarn. A non-uniform bound on binomial approximation with w-functions. Communications in Statistics: Theory and Methods, 51(23):8391–8405, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1896733.

Teamah:2022:BRD

[TEHEG22] Abd El-Moneim A. Teamah, Mohamed Abd Allah El-Hadidy, and Maraw M. El-Ghoul. On bounded range distribution of a Wiener process. Communications in Statistics: Theory and Methods, 51(4):919–942, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1267766.

Teles:2023:TCH

[Tel23] Paulo Teles. Testing conditional heteroscedasticity with systematic sampling of time series. Communications in Statistics: Theory and Methods, 52(15):5427–5450, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.2008976.

Tenreiro:2024:PRT

[Ten24] Carlos Tenreiro. A Parzen–Rosenblatt type density estimator for circular data: exact and asymptotic optimal bandwidths. Communications in Statistics: Theory and Methods, 53(20): 7436–7452, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2023.2264996.

Tas:2024:DSU

[TG24] Cigdem Kosar Tas and Hüseyin Guler. Determining seasonal unit roots with bridge estimator: Monte Carlo evidence and an application to convergence hypothesis. Communications in Statistics: Theory and Methods, 53(16):5721–5743, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2231111.

Torabi:2023:MEL

[TGMS23] Mahmoud Torabi, Malay Ghosh, Jiyoun Myung, and Mark Steel. Measurement error in linear regression models with fat tails and skewed errors. Communications in Statistics: Theory and Methods, 52(15):5407–5426, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.2008442.

Tavakoli:2023:CCE

[TH23] Masoud Tavakoli and Ali Akbar Heydari. Control chart for exponential individual samples with adaptive sampling interval method based on economic statistical design: an extension of Costa and Rahim’s model. Communications in Statistics: Theory and Methods, 52(14):4993–5009, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1999980.

Tang:2024:ADT

[TH24] Fuquan Tang and Dong Han. The asymptotic distribution of a truncated sample mean for the extremely heavytailed distributions. Communications in Statistics: Theory and Methods, 53(14):5168–5185, 2024. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2023.2209231.

Theodossiou:2022:TST

[The22] Panayiotis Theodossiou. Truncated skewed type III generalized logistic distribution: risk measurement applications. Communications in Statistics: Theory and Methods, 51(5): 1379–1402, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2020.1764036.

Tajuddin:2022:FPN

[TIIB22] Razik Ridzuan Mohd Tajuddin, Noriszura Ismail, Kamarulzaman Ibrahim, and Shaiful Anuar Abu Bakar. A fourparameter negative binomial-Lindley distribution for modeling over and underdispersed count data with excess zeros. Communications in Statistics: Theory and Methods, 51(2): 414–426, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2020.1749854.

Tsukuma:2020:ECM

[TK20] Hisayuki Tsukuma and Tatsuya Kubokawa. Estimation of a covariance matrix in multivariate skew-normal distribution. Communications in Statistics: Theory and Methods, 49(5): 1174–1200, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2018.1554137.

Taheri:2024:CBF

[TK24] Marzieh Taheri and Manouchehr Kheradmandnia. Consecutive Bayes factor for the mean vector. Communications in Statistics: Theory and Methods, 53(10):3833–3849, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2165406.

Tzavelas:2022:SPF

[TKP22] George Tzavelas, Christina Koutropoulou, and Konstadinos Politis. Some properties of the failure rate function for mixtures of Erlang distributions. Communications in Statistics: Theory and Methods, 51(17):5850–5872, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1849720.

Tanis:2024:NZI

[TKP24] Caner Tanis, Haydar Koç, and Ahmet Pekgör. A new zeroinflated discrete Lindley regression model. Communications in Statistics: Theory and Methods, 53(12):4252–4271, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2177108.

Tian:2023:ORI

[TLL23] Linlin Tian, Guoqing Li, and You Lv. Optimal reinsurance and investment problem with the minimum capital deposit constraint. Communications in Statistics: Theory and Methods, 52(19):6751–6766, 2023. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2023.2179372.

Tran:2023:EPL

[TLLL23] Phuoc-Loc Tran, Truong-Nhat Le, Shen-Ming Lee, and Chin-Shang Li. Estimation of parameters of logistic regression with covariates missing separately or simultaneously. Communications in Statistics: Theory and Methods, 52(6):1981–2009, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2021.1943443.

Tian:2024:JEL

[TLN24] Weizhong Tian, Tingting Liu, and Wei Ning. Jackknife empirical likelihood for the mean of a zero-and-one inflated population. Communications in Statistics: Theory and Methods, 53(3):980–994, 2024. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2022.2098977.

Trang:2021:DRM

[TLRB21] Bui Thi Thien Trang, Jean-Michel Loubes, Laurent Risser, and Patricia Balaresque. Distribution regression model with a Reproducing Kernel Hilbert Space approach. Communications in Statistics: Theory and Methods, 50(9):1955–1977, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2019.1658782.

Toker:2023:RED

[TÖ23] Selma Toker and Nimet Özbay. Restricted estimation of distributed lag model from a Bayesian point of view. Communications in Statistics: Theory and Methods, 52(11):3927–3938, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2021.1982985.

Thuy:2024:NEC

[TP24a] Le Thi Hong Thuy and Cao Xuan Phuong. Nonparametric estimations for the cumulative distribution functions of random effects in a linear mixed-effects model. Communications in Statistics: Theory and Methods, 53(10):3659–3687, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2158347.

Tzavelas:2024:FRC

[TP24b] George Tzavelas and Konstadinos Politis. The failure rate for the convolution of two distributions, one of which has bounded support. Communications in Statistics: Theory and Methods, 53(13):4717–4729, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2186729.

Thavaneswaran:2020:GVR

[TPF20] Aerambamoorthy Thavaneswaran, Alex Paseka, and Julieta Frank. Generalized value at risk forecasting. Communications in Statistics: Theory and Methods, 49(20):4988–4995, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2019.1610443.

Tripathi:2022:ESP

[TPM22] Yogesh Mani Tripathi, Constantinos Petropoulos, and Amulya Kumar Mahto. Estimating the scale parameter of an exponential distribution under progressive type II censoring. Communications in Statistics: Theory and Methods, 51(19): 6777–6791, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2020.1866609.

Tripathi:2020:QEP

[TPS20] Yogesh Mani Tripathi, Constantinos Petropoulos, and Tanmay Sen. Quantile estimation for a progressively censored exponential distribution. Communications in Statistics: Theory and Methods, 49(16):3919–3932, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1593456.

Trong:2020:NER

[TPV20] Dang Duc Trong, Cao Xuan Phuong, and Tran Quoc Viet. Nonparametric estimation of random effects densities in a linear mixed-effects model with Fourier-oscillating noise density. Communications in Statistics: Theory and Methods, 49(24): 5988–6015, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2019.1625923.

Tang:2024:AEL

[TQ24] Jie Tang and Yongsong Qin. Adjusted empirical likelihood for probability density functions under strong mixing samples. Communications in Statistics: Theory and Methods, 53(18): 6449–6461, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2023.2246088.

Triantafyllou:2021:WTR

[Tri21] Ioannis S. Triantafyllou. Wilcoxon-type rank-sum control charts based on progressively censored reference data. Communications in Statistics: Theory and Methods, 50(2):311– 328, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2019.1634816.

Triacca:2024:GNC

[Tri24] Umberto Triacca. Granger non causality and predictor spaces. Communications in Statistics: Theory and Methods, 53(22): 7908–7913, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2023.2276045.

Theodossiou:2022:MPR

[TS22] Panayiotis Theodossiou and Christos Savva. Market price of risk estimation: Does distribution matter? Communications in Statistics: Theory and Methods, 51(21):7413–7432, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1872643.

Tsai:2022:SRC

[Tsa22] Shin-Fu Tsai. Some results on the construction of sliced orthogonal arrays of parallel-flats type. Communications in Statistics: Theory and Methods, 51(3):569–580, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1822406.

Tsionas:2024:BIU

[Tsi24] Mike Tsionas. Bayesian inference using least median of squares and least trimmed squares in models with independent or correlated errors and outliers. Communications in Statistics: Theory and Methods, 53(16):5761–5772, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2232905.

Tahmasebi:2022:NCE

[TT22] Saeid Tahmasebi and Abdolsaeed Toomaj. On negative cumulative extropy with applications. Communications in Statistics: Theory and Methods, 51(15):5025–5047, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1831541.

Thiagarajan:2023:CBI

[TT23] S. A. Thiagarajan and S. Thobias. Classical and Bayesian inference procedures of a three unit cold standby system with dependent structure. Communications in Statistics: Theory and Methods, 52(12):4247–4258, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1988106.

Ta:2020:ASC

[TTL20] Son Cong Ta, Cuong Manh Tran, and Dung Van Le. On the almost sure convergence for sums of negatively superadditive dependent random vectors in Hilbert spaces and its application. Communications in Statistics: Theory and Methods, 49(11): 2770–2786, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2019.1584304.

Ta:2023:CWS

[TTLT23] Son Cong Ta, Cuong Manh Tran, Dung Van Le, and Chien Van Ta. On the convergence for weighted sums of Hilbert-valued coordinatewise pairwise NQD random variables and its application. Communications in Statistics: Theory and Methods, 52(23):8371–8387, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2062604.

Tian:2020:SRF

[TW20] Yongge Tian and Jie Wang. Some remarks on fundamental formulas and facts in the statistical analysis of a constrained general linear model. Communications in Statistics: Theory and Methods, 49(5):1201–1216, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1554138.

Tang:2022:NPB

[TWZ22] Mingtian Tang, Yunyan Wang, and Qingqing Zhan. Non parametric bias reduction of diffusion coefficient in integrated diffusion processes. Communications in Statistics: Theory and Methods, 51(18):6435–6446, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1861298.

Tan:2021:DOP

[TY21] Jiyang Tan and Senlin Yuan. A dividend optimization problem with constraint of survival probability in a Markovian environment model. Communications in Statistics: Theory and Methods, 50(15):3522–3546, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1705981.

Tao:2023:MLE

[TY23] Ye Tao and Juliang Yin. Maximum likelihood estimation for quantile autoregression models with Markovian switching. Communications in Statistics: Theory and Methods, 52(22): 7915–7943, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2022.2051052.

Tang:2021:LTL

[TZ21] Zhiqiang Tang and Yong Zhang. Limit theorems for linear processes generated by ρ-mixing sequence. Communications in Statistics: Theory and Methods, 50(17):4081–4095, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1710761.

Tang:2022:PEF

[TZ22a] Shiyi Tang and Jiali Zheng. Penalized estimation in finite mixture of ultra-high dimensional regression models. Communications in Statistics: Theory and Methods, 51(17):5971–5992, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2020.1851717.

Tang:2022:CMC

[TZ22b] Zhiqiang Tang and Yong Zhang. Complete moment convergence for the linear processes with random coefficients generated by a class of random variables. Communications in Statistics: Theory and Methods, 51(21):7652–7664, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1876238.

Tengteng:2023:ENM

[TZ23] Xu Tengteng and Riquan Zhang. Estimation of the nonparametric mean and covariance functions for multivariate longitudinal and sparse functional data. Communications in Statistics: Theory and Methods, 52(18):6616–6639, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2032170.

Tian:2024:AMM

[TZ24] Ruiling Tian and Yao Zhang. Analysis of M/M/1 queueing systems with negative customers and unreliable repairers. Communications in Statistics: Theory and Methods, 53(21): 7491–7504, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2023.2265000.

Tang:2024:LDP

[TZL24] Zhiqiang Tang, Yong Zhang, and Wei Liu. Limit distribution for products of sums of partial sums of linear processes generated by NSD sequence. Communications in Statistics: Theory and Methods, 53(4):1535–1545, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2104874.

Tang:2022:AEL

[TZYY22] Xinrong Tang, Peixin Zhao, Yiping Yang, and Weiming Yang. Adjusted empirical likelihood inferences for varying coefficient partially non linear models with endogenous covariates. Communications in Statistics: Theory and Methods, 51(4):953– 973, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2020.1747078.

Tengteng:2024:ECM

[TZZ24] Xu Tengteng, Xiuzhen Zhang, and Riquan Zhang. Estimation of a clustering model for non Gaussian functional data. Communications in Statistics: Theory and Methods, 53(18):6462– 6476, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2023.2246089.

Noor-ul-Amin:2021:MTE

[uA21] Muhammad Noor ul Amin. Memory type estimators of population mean using exponentially weighted moving averages for time scaled surveys. Communications in Statistics: Theory and Methods, 50(12):2747–2758, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1670850.

Noor-ul-Amin:2022:AIB

[uAK22] Muhammad Noor ul Amin and Waqas Kazmi. Auxiliary information based joint monitoring control chart using generalized likelihood ratio test statistic. Communications in Statistics: Theory and Methods, 51(8):2438–2460, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1776328.

Noor-ul-Amin:2021:ECC

[uATH21] Muhammad Noor ul Amin, Muhammad Tayyab, and Muhammad Hanif. Efficient control charts for monitoring the process mean using different paired double ranked set sampling designs. Communications in Statistics: Theory and Methods, 50(10):2211–2223, 2021. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2019.1662047.

Utkin:2021:NBB

[UC21] Lev V. Utkin and Frank P. A. Coolen. A new boostingbased software reliability growth model. Communications in Statistics: Theory and Methods, 50(24):6167–6194, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1740736.

Udom:2021:MHM

[UE21] Akaninyene Udo Udom and Ukobong Gregory Ebedoro. On multinomial hidden Markov model for hierarchical manpower systems. Communications in Statistics: Theory and Methods, 50(6):1370–1386, 2021. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2019.1650185.

Umali:2024:FOS

[UET24] Lester Charles A. Umali, Richard B. Eden, and Timothy Robin Y. Teng. A first-order Stein characterization for absolutely continuous bivariate distributions. Communications in Statistics: Theory and Methods, 53(18):6695–6716, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2250485.

Unsal:2021:NPD

[ÜFR21] Mehmet Güray Ünsal, Daniel Friesner, and Robert Rosenman. New posterior distributions for the incidence of inefficiency in DEA scores. Communications in Statistics: Theory and Methods, 50(8):1774–1780, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1653920.

Unal:2021:IFE

[ÜK21] Ceren Ünal and Cem Kadilar. Improved family of estimators using exponential function for the population mean in the presence of non-response. Communications in Statistics: Theory and Methods, 50(1):237–248, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1634818.

Uranga:2022:MRI

[UMA22] Rolando Uranga, Geert Molenberghs, and Sira Allende. A multiple regression imputation method with application to sensitivity analysis under intermittent missingness. Communications in Statistics: Theory and Methods, 51(15):5146–5161, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2020.1834581.

Udom:2023:CRS

[UN23] Akaninyene Udo Udom and Chijioke Joel Nweke. Convergence results for stochastic convex feasibility problem using random Mann and simultaneous projection iterative algorithms in Hilbert space. Communications in Statistics: Theory and Methods, 52(12):4329–4343, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1990956.

Unvan:2021:CDS

[Ünv21a] Yüksel Akay Ünvan. A credit default swap application by using quantile regression technique. Communications in Statistics: Theory and Methods, 50(7):1573–1586, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1711126.

Unvan:2021:EEC

[Ünv21b] Yüksel Akay Ünvan. The effects of “economic crisis” and “financial crisis” searches conducted in Google on dollar/ TL parity and a causality analysis. Communications in Statistics: Theory and Methods, 50(7):1587–1598, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1678645.

Unvan:2021:IBU

[Ünv21c] Yüksel Akay Ünvan. Impacts of Bitcoin on USA, Japan, China and Turkey stock market indexes: Causality analysis with value at risk method (VAR). Communications in Statistics: Theory and Methods, 50(7):1599–1614, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1678644.

Unvan:2021:MBA

[Ünv21d] Yüksel Akay Ünvan. Market basket analysis with association rules. Communications in Statistics: Theory and Methods, 50(7):1615–1628, 2021. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2020.1716255.

Vamvakari:2020:MDD

[Vam20] Malvina Vamvakari. On multivariate discrete q-distributions — a multivariate q-Cauchy’s formula. Communications in Statistics: Theory and Methods, 49(24):6080–6095, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1626427.

Vasconcelos:2022:PAP

[VCOB22] Julio C. S. Vasconcelos, Gauss M. Cordeiro, Edwin M. M. Ortega, and Marco A. M. Biaggioni. The parametric and additive partial linear regressions based on the generalized odd log-logistic log-normal distribution. Communications in Statistics: Theory and Methods, 51(11):3480–3507, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1795681.

Vasconcelos:2024:PPL

[VCOS24] Julio Cezar S. Vasconcelos, Gauss M. Cordeiro, Edwin M. M. Ortega, and Helton Saulo. Parametric and partially linear regressions for agricultural economy data. Communications in Statistics: Theory and Methods, 53(6):2067–2091, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2117987.

vandenHeuvel:2022:BZI

[vdHvDZ22] Edwin R. van den Heuvel, Stephan A. W. van Driel, and Zhuozhao Zhan. A bivariate zero-inflated Poisson control chart: Comments and corrections on earlier results. Communications in Statistics: Theory and Methods, 51(10):3438– 3445, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2020.1736304.

vanderMerwe:2021:BFA

[vdMGSM21] Abraham J. van der Merwe, Piet C. N. Groenewald, Morné R. Sjölander, and Johan H. Meyer. A Bayesianfrequentists approach for detecting outliers in a one-way variance components model. Communications in Statistics: Theory and Methods, 50(23):5652–5677, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1737127.

Visagie:2023:CCB

[VdWMB23] I. J. H. Visagie, D. J. de Waal, S. L. Makgai, and A. Bekker. Contributions to the class of beta-generated distributions. Communications in Statistics: Theory and Methods, 52(4): 1101–1117, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2021.1923748.

Vasquez:2021:PSC

[VE21] Alejandro R. Vásquez and Gabriel Escarela. Parametric and semiparametric copula-based models for the regression analysis of competing risks. Communications in Statistics: Theory and Methods, 50(12):2831–2847, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1676447.

Verdier:2020:ELB

[Ver20] Ghislain Verdier. An empirical likelihood-based CUSUM for on-line model change detection. Communications in Statistics: Theory and Methods, 49(8):1818–1839, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1565834.

vanHeerden:2023:TSB

[vHP23] Carl J. L. van Heerden and Charl Pretorius. Tests for symmetry based on the integrated empirical process. Communications in Statistics: Theory and Methods, 52(13):4675–4691, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2021.1998832.

Vidmar:2022:LEU

[Vid22] Matija Vidmar. On laws exhibiting universal ordering under stochastic restart. Communications in Statistics: Theory and Methods, 51(5):1290–1305, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1759640.

Volf:2023:MDR

[VK23a] Petr Volf and Tomás Kourim. A model of discrete random walk with history-dependent transition probabilities. Communications in Statistics: Theory and Methods, 52(15):5173– 5186, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2021.2004425.

Volovskiy:2023:CLB

[VK23b] Grigoriy Volovskiy and Udo Kamps. Comparison of likelihood-based predictors of future Pareto and Lomax record values in terms of Pitman closeness. Communications in Statistics: Theory and Methods, 52(6):1905–1922, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1941112.

Vanderperre:2022:GPS

[VM22] Edmond J. Vanderperre and Stanislav S. Makhanov. On Gaver’s parallel system supervised by a safety unit: The global recovery time. Communications in Statistics: Theory and Methods, 51(20):6935–6946, 2022. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2020.1868514.

Vogel:2022:GM

[Vog22] Richard M. Vogel. The geometric mean? Communications in Statistics: Theory and Methods, 51(1):82–94, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1743313.

Vinayaka:2024:PBB

[VPM^{+}24] Vinayaka, Rajender Parsad, B. N. Mandal, Sukanta Dash, Vinaykumar L. N. Mukesh Kumar, and D. R. Singh. Partially balanced bipartite block designs. Communications in Statistics: Theory and Methods, 53(19):6777–6784, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2251623.

Vasquez:2022:UVD

[VRB22] Jonathan K. J. Vasquez, Josemar Rodrigues, and N. Balakrishnan. A useful variance decomposition for destructive Waring regression cure model with an application to HIV data. Communications in Statistics: Theory and Methods, 51(20): 6978–6989, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2020.1869782.

vonRosen:2020:SAE

[vRvR20] Tatjana von Rosen and Dietrich von Rosen. Small area estimation using reduced rank regression models. Communications in Statistics: Theory and Methods, 49(13):3286–3297, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2019.1586946.

Ventosa-Santaularia:2022:SMR

[VSVVLRV22] Daniel Ventosa-Santaulària, J. Eduardo Vera-Valdés, Katarzyna Lasak, and Ricardo Ramírez-Vargas. Spurious multivariate regressions under fractionally integrated processes. Communications in Statistics: Theory and Methods, 51(7):2034–2056, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2020.1758945.

Vinaykumar:2024:MRP

[VVHK24] L. N. Vinaykumar, Cini Varghese, Mohd Harun, and Sayantani Karmakar. Minimally replicated PBIB designs for multi-environmental trials. Communications in Statistics: Theory and Methods, 53(13):4696–4716, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2185753.

Vilca:2024:TRM

[VVS^{+}24] Filidor Vilca, Roberto Vila, Helton Saulo, Luis Sánchez, and Jeremias Leão. Theoretical results and modeling under the discrete Birnbaum–Saunders distribution. Communications in Statistics: Theory and Methods, 53(5):1745–1759, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2110843.

Walker:2022:PNL

[Wal22]  Stephen G. Walker.  On a property of a non-local

moment prior. Communications in Statistics: Theory and Methods, 51(11):3799–3805, 2022. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2020.1804590.

Walker:2023:IDM

[Wal23] Stephen G. Walker. On infinitely divisible multivariate gamma distributions. Communications in Statistics: Theory and Methods, 52(13):4484–4490, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1995431.

Wan:2021:IPP

[Wan21a] Fei Wan. An interpretation of the properties of the propensity score in the regression framework. Communications in Statistics: Theory and Methods, 50(9):2096–2105, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1659369.

Wang:2021:NUR

[Wan21b] Qiuping Wang. A note on undirected random graph models parameterized by the strengths of vertices. Communications in Statistics: Theory and Methods, 50(22):5380–5398, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1728332.

Wang:2022:MPE

[Wan22a] Chendi Wang. Modified Poisson estimators for grouped and right-censored counts. Communications in Statistics: Theory and Methods, 51(6):1588–1604, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1926512.

Wang:2022:IET

[Wan22b] Lei Wang. Identifiability and estimation of two-sample data with nonignorable missing response. Communications in Statistics: Theory and Methods, 51(20):7073–7087, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1871015.

Wang:2023:PEP

[Wan23a] Chao Wang. Parameter estimation for power functionlognormal composite distribution. Communications in Statistics: Theory and Methods, 52(9):2966–2982, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1965622.

Wang:2023:LMD

[Wan23b] Haixu Wang. Large and moderate deviations for a discretetime marked Hawkes process. Communications in Statistics: Theory and Methods, 52(17):6037–6062, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.2024236.

Wang:2023:CAB

[Wan23c] Jun Wang. Consistency of the adaptive Bayesian estimator of reliability based on sequential experiments. Communications in Statistics: Theory and Methods, 52(24):8749–8762, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2022.2070765.

Wan:2023:SBD

[WBM^{+}23] Shu-Mei Wan, Narayanaswamy Balakrishnan, Monica Mayeni Manurung, Kwang-Hwa Chang, and Chien-Hua Wu. Semiblinded design in clinical trials. Communications in Statistics: Theory and Methods, 52(20):7165–7183, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2042024.

Wang:2020:THD

[WC20a] Siyang Wang and Hengjian Cui. Test for high dimensional regression coefficients of partially linear models. Communications in Statistics: Theory and Methods, 49(17):4091–4116, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2019.1594293.

Wang:2020:ERT

[WC20b] Yuebao Wang and Dongya Cheng. Elementary renewal theorems for widely dependent random variables with applications to precise large deviations. Communications in Statistics: Theory and Methods, 49(14):3352–3374, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1586947.

Wu:2020:EBP

[WC20c] Cheng-Hsun Wu and Kai-Jiun Chang. Error bounds for the perturbation solution of the transition density under a multi-factor CIR term structure model with weak mean-reversion effect. Communications in Statistics: Theory and Methods, 49(21):5294–5311, 2020. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2019.1617879.

Wang:2022:LRO

[WC22] Jiantian Wang and Bin Cheng. Likelihood ratio ordering for parallel systems with exponential components. Communications in Statistics: Theory and Methods, 51(19):6741–6759, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2020.1866205.

Wang:2024:CBA

[WC24] Qiao Wang and Zhongheng Cai. A composite Bayesian approach for quantile curve fitting with non-crossing constraints. Communications in Statistics: Theory and Methods, 53(20): 7119–7143, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2023.2259524.

Wang:2022:SBT

[WCT22] Jin Wang, Javier Cabrera, and Kwok Leung Tsui. A simulation-based tree method for building linear models with interactions. Communications in Statistics: Theory and Methods, 51(2):404–413, 2022. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2020.1749665.

Wei:2022:PPE

[WCX22] Yunfei Wei, Daijun Chen, and Shifeng Xiong. Projection pursuit emulation for many-input computer experiments. Communications in Statistics: Theory and Methods, 51(17):6078– 6090, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2020.1853772.

Wang:2023:EBS

[WCY23] Liyuan Wang, Zhiping Chen, and Peng Yang. Equilibrium behavioral strategy for a DC pension plan with piecewise linear state-dependent risk tolerance. Communications in Statistics: Theory and Methods, 52(7):2309–2337, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1952265.

Wang:2020:RTW

[WD20] Tao Wang and Nicholas DeVogel. A revisit to two-way factorial ANOVA with mixed effects and interactions. Communications in Statistics: Theory and Methods, 49(19):4618–4635, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2019.1604961.

Wang:2022:ORI

[WDH^{+}22] Yijun Wang, Yingchun Deng, Ya Huang, Jieming Zhou, and Xuyan Xiang. Optimal reinsurance-investment policies for insurers with mispricing under mean-variance criterion. Communications in Statistics: Theory and Methods, 51(16):5653– 5680, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2020.1844239.

Wu:2020:OAA

[WDLW20] Sang Wu, Yinghui Dong, Wenxin Lv, and Guojing Wang. Optimal asset allocation for participating contracts with mortality risk under minimum guarantee. Communications in Statistics: Theory and Methods, 49(14):3481–3497, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1589518.

Wang:2024:APW

[WDW24] Xuejun Wang, Xin Deng, and Yi Wu. Asymptotic properties of the wavelet estimator in non parametric regression model with martingale difference errors. Communications in Statistics: Theory and Methods, 53(9):3312–3336, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2152288.

Wei:2021:PES

[Wei21] Chao Wei. Parameter estimation for stochastic Lotka–Volterra model driven by small Lévy noises from discrete observations. Communications in Statistics: Theory and Methods, 50(24): 6014–6023, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2020.1738489.

Wei:2023:LSE

[Wei23] Chao Wei. Least squares estimation for discretely observed Ornstein–Uhlenbeck process driven by small stable noises. Communications in Statistics: Theory and Methods, 52(12): 4138–4150, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2021.1986537.

Wang:2023:SCE

[WF23] Boyang Wang and Rui Fang. Stochastic comparisons on extreme order statistics from observations associated by FGM copula. Communications in Statistics: Theory and Methods, 52(10):3492–3510, 2023. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2021.1974481.

Wilson:2020:KDE

[WG20] Christopher M. Wilson and Patrick Gerard. Kernel density estimation for hierarchical data. Communications in Statistics: Theory and Methods, 49(6):1495–1512, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1563179.

Wei:2021:ABE

[WGX21] Shengxue Wei, Xiaoli Gan, and Guodong Xing. Asymptotic behavior of expected shortfall for portfolio loss under bivariate dependent structure. Communications in Statistics: Theory and Methods, 50(1):132–142, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1630439.

Wang:2023:OIM

[WH23a] Xuhui Wang and Lei Hu. Optimal investment meanfield and N-player games with memory effect and relative performance competition. Communications in Statistics: Theory and Methods, 52(5):1472–1489, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1928204.

Wei:2023:CAM

[WH23b] Linhai Wei and Yijun Hu. Capital allocation with multivariate risk statistics with positive homogeneity and subadditivity. Communications in Statistics: Theory and Methods, 52(18): 6684–6694, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2022.2032173.

Wu:2023:FST

[WHCC23] Yuh-Jenn Wu, Li-Syuan Hong, Li-Hsueh Cheng, and Li-Chu Chien. Forecasting short-term mortality trends using Bernstein polynomials. Communications in Statistics: Theory and Methods, 52(8):2417–2433, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1952432.

Webber:2022:NPD

[WHMA22] James Webber, Erika Hussey, Eric Miller, and Shuchin Aeron. On non-parametric density estimation on linear and non-linear manifolds using generalized Radon transforms. Communications in Statistics: Theory and Methods, 51(23):8406–8426, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2021.1897143.

Wu:2023:DWL

[WHW23] Shuo-Jye Wu, Syuan-Rong Huang, and Jie-Huei Wang. Determination of warranty length for one-shot devices with Rayleigh lifetime distribution. Communications in Statistics: Theory and Methods, 52(5):1400–1416, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1927755.

Wichitaksorn:2020:AMV

[Wic20] Nuttanan Wichitaksorn. Analyzing multiple vector autoregressions through matrix-variate normal distribution with two covariance matrices. Communications in Statistics: Theory and Methods, 49(8):1801–1817, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1565832.

Wijesuriya:2024:NSV

[Wij24] Uditha Amarananda Wijesuriya. A novel sample variance formula and Sv-plot3 for testing hypotheses. Communications in Statistics: Theory and Methods, 53(17):6137–6151, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2239965.

Wada:2020:OTP

[WIN20] Naoya Wada, Kodo Ito, and Toshio Nakagawa. Optimal training plans on physical performance considering supercompensation. Communications in Statistics: Theory and Methods, 49(15):3761–3771, 2020. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2020.1722845.

Wiroonsri:2023:CIU

[Wir23] Nathakhun Wiroonsri. Concentration inequalities using approximate zero bias couplings with applications to Hoeffding’s statistic under the Ewens distribution. Communications in Statistics: Theory and Methods, 52(19):6734–6750, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2032753.

Wu:2023:SDL

[WJC23] Shu-Fei Wu, Jyun-Jhe Jheng, and Wei-Tsung Chang. Sampling design for the lifetime performance index of exponential lifetime distribution under progressive type I interval censoring. Communications in Statistics: Theory and Methods, 52(8):2766–2782, 2023. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2021.1959933.

Wang:2023:IBE

[WJL23] Zhining Wang, Hua Jin, and Hezhi Lu. An IM-based efficient test for non inferiority of the odds ratio between two independent binomial proportions. Communications in Statistics: Theory and Methods, 52(4):1217–1236, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1926507.

Wang:2022:GCB

[WJS22] Kangning Wang, Haotian Jin, and Xiaofei Sun. Gaussian copula based composite quantile regression in semivarying models with longitudinal data. Communications in Statistics: Theory and Methods, 51(4):1110–1132, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1758944.

Wu:2022:IVC

[WjWbK22] Fan Wu, Guan jun Wang, and Xin bing Kong. Inference on volatility curve at high frequencies via functional data analysis. Communications in Statistics: Theory and Methods, 51(19):6683–6700, 2022. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2020.1864829.

Wang:2021:WCQ

[WJXF21] Jiang-Feng Wang, Wei-Jun Jiang, Fang-Yin Xu, and Wu-Xin Fu. Weighted composite quantile regression with censoring indicators missing at random. Communications in Statistics: Theory and Methods, 50(12):2900–2917, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1678638.

Wu:2023:GDB

[WK23]  Jibo Wu and B. M. Golam Kibria.  A generalized

difference-based mixed two-parameter estimator in partially linear model. Communications in Statistics: Theory and Methods, 52(17):6008–6017, 2023. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2021.2024234.

Waseem:2021:GET

[WKS21] Zara Waseem, Hina Khan, and Javid Shabbir. Generalized exponential type estimator for the mean of sensitive variable in the presence of non-sensitive auxiliary variable. Communications in Statistics: Theory and Methods, 50(14):3477–3488, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2019.1708399.

Wang:2022:ELV

[WL22a] Bao-Hua Wang and Han-Ying Liang. Empirical likelihood in varying-coefficient quantile regression with missing observations. Communications in Statistics: Theory and Methods, 51(1):267–283, 2022. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2020.1747629.

Wu:2022:SPS

[WL22b] Xinlin Wu and Haixiong Li. A satisficing policy of the secretary problem: theory and simulation. Communications in Statistics: Theory and Methods, 51(18):6151–6165, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1856875.

Wang:2023:EQR

[WL23a] Yingjie Wang and Xinsheng Liu. Extreme quantile regression for tail single-index varying-coefficient models. Communications in Statistics: Theory and Methods, 52(9):2860–2881, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2021.1961154.

Wu:2023:PML

[WL23b] Weisan Wu and Shaoting Li. Penalized maximum likelihood estimator for finite multivariate skew normal mixtures. Communications in Statistics: Theory and Methods, 52(23):8280– 8305, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2022.2060513.

Wei:2020:GMG

[WLL^{+}20] Zhengyuan Wei, Suping Li, Qiao Li, Yucan Yu, and Xiaoyang Zheng. Gamma mixture of generalized error distribution. Communications in Statistics: Theory and Methods, 49(19): 4819–4833, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2019.1609037.

Wei:2023:SNR

[WLL23] Yinzhao Wei, Sanyang Liu, and Xiaoliang Ling. Some new results on redundancy allocation in series systems. Communications in Statistics: Theory and Methods, 52(3):788–805, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2021.1931329.

Wei:2024:ROS

[WLLB24] Yinzhao Wei, Sanyang Liu, Xiaoliang Ling, and Narayanaswamy Balakrishnan. Reliability optimization for series and parallel systems with series subsystem and comprising of dependent components under random shock environment. Communications in Statistics: Theory and Methods, 53(4):1188–1211, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2022.2093911.

Wang:2023:FRB

[WLS23] Guanjun Wang, Peng Liu, and Lijuan Shen. Failure rate-based models for systems subject to random shocks. Communications in Statistics: Theory and Methods, 52(11):3982–4000, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2021.1983602.

Wu:2021:EVS

[WLT21] Liucang Wu, Shuangshuang Li, and Ye Tao. Estimation and variable selection for mixture of joint mean and variance models. Communications in Statistics: Theory and Methods, 50(24):6081–6098, 2021. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2020.1738493.

Wang:2020:TLB

[WLZ20] Yanfei Wang, Zhiming Li, and Runchu Zhang. Threelevel blocked regular designs with general minimum lower order confounding. Communications in Statistics: Theory and Methods, 49(10):2498–2513, 2020. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2019.1576891.

Wang:2023:JSN

[WLZ^{+}23a] Zhen Wang, Liwei Liu, Yiqiang Q. Zhao, Linhong Li, and Wei Xu. Joining strategies of noncooperative and cooperative in a single server retrial queue with N-policy and multiple server vacations. Communications in Statistics: Theory and Methods, 52(4):1076–1100, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1923747.

Wen:2023:RAU

[WLZ^{+}23b] Yanqing Wen, Baoliang Liu, Zhiqiang Zhang, Shugui Kang, Qingan Qiu, and Haiyan Shi. Reliability analysis for uncertain competing failure degradation system with a change point. Communications in Statistics: Theory and Methods, 52(8): 2461–2481, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2021.1954196.

Wen:2024:MAR

[WLZ^{+}24] Yanqing Wen, Baoliang Liu, Zhiqiang Zhang, Haiyan Shi, and Shugui Kang. Modeling and analysis for a repairable system with multi-state components under Kmixed redundancy strategy. Communications in Statistics: Theory and Methods, 53(2):748–764, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2158346.

Wang:2021:AFT

[WM21] Kaiyong Wang and Yanzhu Mao. Asymptotics of the finitetime ruin probability of dependent risk model perturbed by diffusion with a constant interest rate. Communications in Statistics: Theory and Methods, 50(4):932–943, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1643888.

Wang:2022:TBS

[WM22] Li Wang and Ting Ma. Tail bounds for sum of gamma variables and related inferences. Communications in Statistics: Theory and Methods, 51(4):853–862, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1756329.

Wang:2022:UAR

[WQSG22] Shijie Wang, Huan Qian, Huimin Sun, and Bingzhen Geng. Uniform asymptotics for ruin probabilities of a non standard bidimensional perturbed risk model with subexponential claims. Communications in Statistics: Theory and Methods, 51(22):7871–7884, 2022. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2021.1882496.

Wang:2021:MAL

[WQZL21] Ning Wang, Linyi Qian, Nan Zhang, and Zehui Liu. Modelling the aggregate loss for insurance claims with dependence. Communications in Statistics: Theory and Methods, 50(9):2080– 2095, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2019.1659368.

Wang:2021:OIP

[WRZL21] Yajie Wang, Ximin Rong, Hui Zhao, and Danping Li. Optimal investment problem between two insurers with value-added service. Communications in Statistics: Theory and Methods, 50(8):1781–1806, 2021. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2019.1653921.

Wang:2021:ROI

[WRZW21] Peiqi Wang, Ximin Rong, Hui Zhao, and Yajie Wang. Robust optimal insurance and investment strategies for the government and the insurance company under mispricing phenomenon. Communications in Statistics: Theory and Methods, 50(4):993–1017, 2021. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2019.1646765.

Wu:2020:NIM

[WS20] Hongping Wu and Ang Shan. Nonparametric inference on mean residual life function with length-biased rightcensored data. Communications in Statistics: Theory and Methods, 49(9):2065–2079, 2020. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2019.1568483.

Wang:2022:AST

[WS22] Suping Wang and Zhanjie Song. Average sampling theorem for the homogeneous random fields in a reproducing kernel subspace of mixed Lebesgue space. Communications in Statistics: Theory and Methods, 51(8):2580–2589, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1777310.

Wang:2023:CMC

[WS23] Rui Wang and Aiting Shen. Complete moment convergence for moving average process based on m-WOD random variables. Communications in Statistics: Theory and Methods, 52(19): 7041–7056, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2022.2037649.

Wang:2021:LSE

[WSG21] Qingbo Wang, Guangjun Shen, and Zhenlong Gao. Least squares estimator for Ornstein–Uhlenbeck processes driven by small fractional Lévy noises. Communications in Statistics: Theory and Methods, 50(8):1838–1855, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1653923.

Weems:2023:FBD

[WSL23] Kimberly S. Weems, Kimberly F. Sellers, and Tong Li. A flexible bivariate distribution for count data expressing data dispersion. Communications in Statistics: Theory and Methods, 52(13):4692–4718, 2023. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2021.1999474.

Wang:2020:ATI

[WSSY20] Jun Wang, Xianmei Song, Guangjun Shen, and Xiuwei Yin. Approximation to two independent Gaussian processes from a unique lévy process and applications. Communications in Statistics: Theory and Methods, 49(21):5220–5234, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1615095.

Wang:2022:OWE

[WSZ22] Jinru Wang, Wenhui Shi, and Xiaochen Zeng. Optimal wavelet estimators of the heteroscedastic pointspread effects and Gauss white noises model. Communications in Statistics: Theory and Methods, 51(5):1133–1154, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1862874.

Wang:2020:SAA

[WT20] Yudong Wang and Yincai Tang. Statistical analysis of accelerated temperature cycling test based on Coffin–Manson model. Communications in Statistics: Theory and Methods, 49(15): 3663–3680, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2019.1702697.

Wang:2024:AEM

[WTS24] Xiaochen Wang, Xiaolong Tang, and Yuping Song. Asymptotics of M-estimators for moderate deviations from a unit root model with possibly infinite variance. Communications in Statistics: Theory and Methods, 53(9):3169–3186, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2150824.

Wu:2020:GCQ

[WTT20] Yanke Wu, Maozai Tian, and Man-Lai Tang. General composite quantile regression: Theory and methods. Communications in Statistics: Theory and Methods, 49(9):2217–2236, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2019.1568493.

Wu:2020:ONE

[Wu20] Yougui Wu. Optimal nonparametric estimator of the area under ROC curve based on clustered data. Communications in Statistics: Theory and Methods, 49(6):1446–1461, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1563176.

Wu:2021:RAM

[Wu21] Yougui Wu. A robust adjustment to McNemar test when the data are clustered. Communications in Statistics: Theory and Methods, 50(6):1515–1529, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1651864.

Wu:2022:EGR

[Wu22a] Jibo Wu. Efficiency of the generalized restricted differencebased almost unbiased ridge estimator in partially linear model. Communications in Statistics: Theory and Methods, 51(5):1403–1412, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03610926.2020.1764041.

Wu:2022:PGM

[Wu22b] Shujin Wu. Poisson-gamma mixture processes and applications to premium calculation. Communications in Statistics: Theory and Methods, 51(17):5913–5936, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1850791.

Wu:2022:CDL

[Wu22c] Yougui Wu. Comparison of diagnostic likelihood ratios of two binary tests with case-control clustered data. Communications in Statistics: Theory and Methods, 51(11):3836–3846, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2021.1980805.

Wu:2022:RBG

[Wu22d] Yougui Wu. A revisit to Bennett’s goodness-of-fit statistic for comparing two predictive values. Communications in Statistics: Theory and Methods, 51(10):3419–3433, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1982987.

Wu:2023:EGD

[Wu23a] Jibo Wu. Efficiency of a generalized difference-based weighted mixed ridge estimator in partially linear model. Communications in Statistics: Theory and Methods, 52(13):4622–4635, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2021.1998533.

Wu:2023:SLT

[Wu23b] Qunying Wu. Strong limit theorems of weighted sums for extended negatively dependent random variables under sub-linear expectations. Communications in Statistics: Theory and Methods, 52(12):4356–4368, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1993259.

Wu:2023:MOS

[Wu23c] Shu-Fei Wu. A modified one stage multiple comparison procedure of exponential location parameters with the control under heteroscedasticity. Communications in Statistics: Theory and Methods, 52(7):2356–2364, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1952267.

Wu:2023:NTS

[Wu23d] Shu-Fei Wu. A note on the two-stage multiple comparison procedures with the average for exponential location parameters under heteroscedasticity. Communications in Statistics: Theory and Methods, 52(10):3219–3228, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1970769.

Wu:2024:BOS

[Wu24a] Shu-Fei Wu. A better one stage multiple comparison procedure of several treatment mean lifetimes with the control for exponential distributions under heteroscedasticity. Communications in Statistics: Theory and Methods, 53(21):7651–7658, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2023.2271106.

Wu:2024:SDT

[Wu24b] Yanhong Wu. Sequential detection of transient signal by moving likelihood ratio statistic in an exponential family. Communications in Statistics: Theory and Methods, 53(15):5460– 5485, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2023.2220922.

Wu:2024:PCR

[WV24] Qiang Wu and Paul Vos. Permutation confidence region for multiple regression and fidelity to asymptotic approximation. Communications in Statistics: Theory and Methods, 53(1):90– 112, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2022.2076119.

Wang:2020:SCP

[WW20] Kunpeng Wang and Xuejun Wang. Strong convergence properties for partial sums of asymptotically negatively associated random vectors in Hilbert spaces. Communications in Statistics: Theory and Methods, 49(22):5578–5586, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1620279.

Wang:2022:EAA

[WW22] Lu Wang and Lianming Wang. An EM algorithm for analyzing right-censored survival data under the semiparametric proportional odds model. Communications in Statistics: Theory and Methods, 51(15):5284–5297, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1837879.

Wang:2023:CIC

[WW23a] Li Wang and Qunying Wu. Complete integral convergence for weighted sums of widely negative dependent random variables under the sub-linear expectations. Communications in Statistics: Theory and Methods, 52(24):8763–8784, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2071448.

Wu:2023:FSK

[WW23b] Jian Wu and Huang Wei. A feasible spline-kernel estimate for short cross-sectional dependence panel data models. Communications in Statistics: Theory and Methods, 52(21):7555– 7563, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2022.2059512.

Wang:2024:CCC

[WW24] Li Wang and Qunying Wu. Complete convergence and complete integral convergence for weighted sums of widely negative dependent random variables under the sub-linear expectations. Communications in Statistics: Theory and Methods, 53(10):3599–3615, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03610926.2022.2158343.

Wang:2024:NBA

[WWC24] Zheqi Wang, Dehui Wang, and Jianhua Cheng. A new bivariate autoregressive model driven by logistic regression. Communications in Statistics: Theory and Methods, 53(1):419– 445, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2022.2069262.

Wang:2023:DCQ

[WWJ23] Zih-Huei Wang, Chien-Wei Wu, and Jhen-Jia Jhu. Design and construction of a quick-switching sampling system with a third-generation capability index. Communications in Statistics: Theory and Methods, 52(11):3633–3651, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1977959.

Wang:2022:APM

[WWL22] Qian Wang, Qiuping Wang, and Jing Luo. Asymptotics in a probit model for directed networks. Communications in Statistics: Theory and Methods, 51(11):3463–3479, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1795197.

Wang:2020:AIL

[WWWH20] Xinghui Wang, Huilong Wang, Hongrui Wang, and Shuhe Hu. Asymptotic inference of least absolute deviation estimation for AR(1) processes. Communications in Statistics: Theory and Methods, 49(4):809–826, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1549252.

Wang:2024:ECC

[WWWZ24] Miaomiao Wang, Min Wang, Xuejun Wang, and Fei Zhang. Equivalent conditions of convergence properties for m-ANA sequence and statistical applications. Communications in Statistics: Theory and Methods, 53(5):1547–1575, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2106575.

Wu:2022:FDC

[WWX22] Ruhao Wu, Bo Wang, and Aiping Xu. Functional data clustering using principal curve methods. Communications in Statistics: Theory and Methods, 51(20):7264–7283, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1872636.

Wang:2020:RBM

[WWZ20] Yunlong Wang, Zhaojun Wang, and Xuemin Zi. Rank-based multiple change-point detection. Communications in Statistics: Theory and Methods, 49(14):3438–3454, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1589515.

Wang:2022:ESI

[WWZ22] Tonghui Wang, Liming Wang, and Xian Zhou. Estimation in single-index varying-coefficient panel data model. Communications in Statistics: Theory and Methods, 51(12):3864–3885, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2020.1804589.

Wu:2020:KSR

[WX20] Zhen Wu and Zhenda Xu. A kind of stochastic recursive zero-sum differential game problem with double obstacles constraint. Communications in Statistics: Theory and Methods, 49(21):5356–5370, 2020. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2019.1618477.

Wang:2024:STP

[WXW24] Yanxin Wang, Jiaqing Xu, and Zhi Wang. A simple tuning parameter selection method for high dimensional regression. Communications in Statistics: Theory and Methods, 53(6): 2003–2020, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2022.2117559.

Wang:2021:RAT

[WXY21] Jinting Wang, Nan Xie, and Nan Yang. Reliability analysis of a two-dissimilar-unit warm standby repairable system with priority in use. Communications in Statistics: Theory and Methods, 50(4):792–814, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1642488.

Wang:2020:JRP

[WY20] Leyang Wang and Fengbin Yu. Jackknife resampling parameter estimation method for weighted total least squares. Communications in Statistics: Theory and Methods, 49(23):5810– 5828, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2019.1622725.

Wei:2023:VSM

[WY23] Gongding Wei and Mingyuan Yu. The variable selection methods and algorithms in the multiple linear model. Communications in Statistics: Theory and Methods, 52(17):6232–6240, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2022.2027449.

Wang:2024:SIS

[WYG24] Xinjing Wang, Tianrui Ye, and Wenhao Gui. Statistical inference for the step-stress model with competing risks from the Kumaraswamy distribution under progressive type-II censoring. Communications in Statistics: Theory and Methods, 53(23):8456–8483, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2291342.

Wang:2021:SLT

[WYJD21] Zhong-Zhi Wang, Shan-Shan Yang, Hai-Feng Jiang, and Fang-Qing Ding. On strong limit theorems for general information sources with an application to AEP. Communications in Statistics: Theory and Methods, 50(6):1387–1399, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1650186.

Wang:2021:CAO

[WYP21] Jing Wang, Rong-Xian Yue, and Shan-Qi Pang. Construction of asymmetric orthogonal arrays of high strength by juxtaposition. Communications in Statistics: Theory and Methods, 50(12):2947–2957, 2021. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2019.1679184.

Wang:2024:EVS

[WYW24a] Yue Wang, Xiaohui Yuan, and Chunjie Wang. Estimation and variable selection for single-index models with non ignorable missing data. Communications in Statistics: Theory and Methods, 53(14):4945–4974, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2198625.

Wywial:2024:ALR

[Wyw24b] Janusz L. Wywiał. On approximation of linear regression disturbance distribution. Communications in Statistics: Theory and Methods, 53(23):8271–8285, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2280539.

Wolf:2021:SDD

[WZ21] Jared Wolf and Hong Zhou. A simple data-driven fallback procedure for multiple comparisons. Communications in Statistics: Theory and Methods, 50(13):3179–3197, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1691231.

Wang:2020:SSC

[WZA20] Jijia Wang, Song Zhang, and Chul Ahn. Sample size calculation for count outcomes in cluster randomization trials with varying cluster sizes. Communications in Statistics: Theory and Methods, 49(1):116–124, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1532004.

Wang:2021:SSE

[WZA21] Jijia Wang, Song Zhang, and Chul Ahn. Sample size estimation for comparing rates of change in K-group repeated binary measurements studies. Communications in Statistics: Theory and Methods, 50(23):5607–5616, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1736302.

Wang:2021:SIV

[WZD21] Xiuli Wang, Peixin Zhao, and Haiyan Du. Statistical inferences for varying coefficient partially non linear model with missing covariates. Communications in Statistics: Theory and Methods, 50(11):2599–2618, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1674870.

Wang:2022:IPM

[WZL22a] Qian Wang, Chen Zhao, and Jing Luo. Inference in a probit model for affiliation networks. Communications in Statistics: Theory and Methods, 51(21):7528–7546, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1873381.

Wang:2022:WDN

[WZL^{+}22b] Qiuping Wang, Xiao Zhang, Jing Luo, Yang Ouyang, and Qian Wang. Weighted directed networks with a differentially private bi-degree sequence. Communications in Statistics: Theory and Methods, 51(2):285–300, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1747630.

Wang:2023:EIH

[WZL23] Yuan-Quan Wang, Ying-Ying Zhang, and Jia-Lei Liu. Expectation identity of the hypergeometric distribution and its application in the calculations of high-order origin moments. Communications in Statistics: Theory and Methods, 52(17): 6018–6036, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2021.2024235.

Wang:2020:RAS

[WZTW20] Liang Wang, Xuanjia Zuo, Yogesh Mani Tripathi, and Junyuan Wang. Reliability analysis for stress-strength model from a general family of truncated distributions under censored data. Communications in Statistics: Theory and Methods, 49(15):3589–3608, 2020. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2019.1710759.

Wang:2022:NEI

[WZZ22] Wensheng Wang, Hui Zhou, and Anwei Zhu. A nonparametric estimation for infectious diseases with latent period. Communications in Statistics: Theory and Methods, 51(19):6701– 6718, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2020.1865402.

Xue:2023:TE

[XD23] Yige Xue and Yong Deng. Tsallis eXtropy. Communications in Statistics: Theory and Methods, 52(3):751–762, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1921804.

Xu:2023:LBG

[XFP23] Long-Hao Xu, Kai-Tai Fang, and Jianxin Pan. Limiting behavior of the gap between the largest two representative points of statistical distributions. Communications in Statistics: Theory and Methods, 52(10):3290–3313, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1970772.

Xu:2022:JEL

[XFW22] Hong-Xia Xu, Guo-Liang Fan, and Jiang-Feng Wang. Jackknife empirical likelihood for the error variance in linear errorsin-variables models with missing data. Communications in Statistics: Theory and Methods, 51(14):4827–4840, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1824274.

Xiong:2023:ISR

[XH23] Juan Xiong and Wenqing He. Identification of survival relevant genes with measurement error in gene expression incorporated. Communications in Statistics: Theory and Methods, 52(15):5155–5172, 2023. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2021.2004424.

Xu:2024:WEN

[XH24] Junlian Xu and Lu Hao. Wavelet estimation of norms for a probability density with negatively dependent biased data. Communications in Statistics: Theory and Methods, 53(18): 6497–6512, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2023.2246607.

Xia:2020:ANB

[Xia20] Weixuan Xia. The average of a negative-binomial lévy process and a class of Lerch distributions. Communications in Statistics: Theory and Methods, 49(4):1008–1024, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1554135.

Xing:2023:PCR

[Xin23] Yang Xing. A posterior convergence rate theorem for general Markov chains. Communications in Statistics: Theory and Methods, 52(16):5910–5921, 2023. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2021.2023183.

Xavier:2021:SSS

[XJ21] Thomas Xavier and Joby K. Jose. A study of stressstrength reliability using a generalization of power transformed half-logistic distribution. Communications in Statistics: Theory and Methods, 50(18):4335–4351, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1716250.

Xiao:2022:OBE

[XL22] Yanting Xiao and Fuxiao Li. Orthogonality-based empirical likelihood inference for varying-coefficient partially nonlinear model with longitudinal data. Communications in Statistics: Theory and Methods, 51(4):1067–1084, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1758141.

Xie:2023:BET

[XL23a] Feng-Chang Xie and Xian-Ju Li. Bayesian estimation for the threshold stochastic volatility model with generalized hyperbolic skew Student’s t distribution. Communications in Statistics: Theory and Methods, 52(12):4053–4071, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1990952.

Xu:2023:EAD

[XL23b] Feng Xu and Xiongying Li. The effect of autocorrelation on the diagnostic procedures. Communications in Statistics: Theory and Methods, 52(10):3333–3349, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1971246.

Xing:2024:REI

[XL24] Xiaoyu Xing and Xiaofang Li. Robust equilibrium investmentreinsurance strategy for n competitive insurers with squareroot factor process. Communications in Statistics: Theory and Methods, 53(12):4469–4486, 2024. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2023.2184185.

Xu:2023:ARQ

[XLW23] Jia Xu, Liwei Liu, and Kan Wu. Analysis of a retrial queueing system with priority service and modified multiple vacations. Communications in Statistics: Theory and Methods, 52(17): 6207–6231, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2022.2027448.

Xiaowen:2020:QRP

[XLY^{+}20] Dai Xiaowen, Jin Libin, Tian Yuzhu, Tian Maozai, and Tang Manlai. Quantile regression for panel data models with fixed effects under random censoring. Communications in Statistics: Theory and Methods, 49(18):4430–4445, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1601221.

Xu:2021:LBR

[XM21] Shoufang Xu and Changlin Mei. Limit behaviors for ratios of order statistics from exponential distributions. Communications in Statistics: Theory and Methods, 50(24):5918–5928, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2020.1737877.

Xiong:2021:NPE

[XO21] Xianzhu Xiong and Meijuan Ou. Non parametric estimation of the conditional density function with rightcensored and dependent data. Communications in Statistics: Theory and Methods, 50(13):3159–3178, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1691230.

Xi:2021:CBM

[XP21] Daiqing Xi and Tianxiao Pang. Common breaks in means for panel data under short-range dependence. Communications in Statistics: Theory and Methods, 50(2):486–505, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1637000.

Xiong:2021:PAM

[XS21] Jingwei Xiong and Junfeng Shang. A penalized approach to mixed model selection via cross-validation. Communications in Statistics: Theory and Methods, 50(11):2481–2507, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1669806.

Xiao:2020:BIZ

[XTXW20] Xiang Xiao, Yincai Tang, Ancha Xu, and Guoqiang Wang. Bayesian inference for zero-and-one-inflated geometric distribution regression model using Pólya–Gamma latent variables. Communications in Statistics: Theory and Methods, 49(15): 3730–3743, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2019.1709647.

Xu:2022:OWE

[Xu22] Junlian Xu. The optimal wavelet estimation for biased density. Communications in Statistics: Theory and Methods, 51(24): 8728–8740, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2021.1905845.

Xiao:2022:PFL

[XW22] Piaoxuan Xiao and Guochang Wang. Partial functional linear regression with autoregressive errors. Communications in Statistics: Theory and Methods, 51(13):4515–4536, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1818097.

Xu:2023:NPM

[XW23] Meng Xu and Qiuping Wang. A network Poisson model for weighted directed networks with covariates. Communications in Statistics: Theory and Methods, 52(15):5274–5293, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.2005101.

Xie:2021:VSS

[XWCT21] Li Xie, Xiaorui Wang, Weihu Cheng, and Tian Tang. Variable selection for spatial autoregressive models. Communications in Statistics: Theory and Methods, 50(6):1325–1340, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2019.1649428.

Xu:2021:ASL

[XWH21] Feng Xu, Binhui Wang, and Yawen Hou. Almost sure local central limit theorem for the product of some partial sums with optimized weight. Communications in Statistics: Theory and Methods, 50(13):3051–3062, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1680694.

Xu:2024:FTR

[XWW24] Chenghao Xu, Kaiyong Wang, and Xinyi Wu. The finitetime ruin probability of a risk model with stochastic return and subexponential claim sizes*. Communications in Statistics: Theory and Methods, 53(6):2194–2204, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2122840.

Xing:2021:EPF

[XWY21] Yu Xing, Dingcheng Wang, and Xiaoping Yang. Equilibrium pricing of foreign exchange options under a discontinuous model with stochastic jump intensity. Communications in Statistics: Theory and Methods, 50(5):1059–1081, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1646763.

Xie:2023:RBP

[XWYZ23] Lin Xie, Wei Wang, Zhixin Yang, and Nan Zhang. Risk-based premium evaluation with jump diffusion process for PBGC. Communications in Statistics: Theory and Methods, 52(6): 1854–1869, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2021.1939381.

Xie:2023:EUC

[XXBD23] Lin Xie, Rui Xiong, Kevin Blot, and Jean-Marc Dessirier. Estimating upper confidence bounds for positive ratios of normal random variables. Communications in Statistics: Theory and Methods, 52(14):4894–4903, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2013.833241.

Xia:2024:ROR

[XXZ24] Zichao Xia, Wanwan Xia, and Zhenfeng Zou. Revisit optimal reinsurance under a new distortion risk measure. Communications in Statistics: Theory and Methods, 53(15):5657–5672, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2023.2226783.

Xiao:2020:NJM

[XY20] Yugu Xiao and Jing Yao. A note on joint mix random vectors. Communications in Statistics: Theory and Methods, 49(12):

3063–3072, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2019.1586937.

Xu:2021:CMC

[XY21] Xiu Xu and Jigao Yan. Complete moment convergence for randomly weighted sums of END sequences and its applications. Communications in Statistics: Theory and Methods, 50(12): 2877–2899, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2019.1678637.

Xia:2023:LTC

[XY23] Xiaoyu Xia and Litan Yan. Limit theorems for a class of integral functionals driven by fractional Brownian motion. Communications in Statistics: Theory and Methods, 52(3):583– 601, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2021.1917616.

Xie:2024:ISE

[XZ24] Jiayi Xie and Zhimin Zhang. Infinite series expansion of some finite-time dividend and ruin related functions. Communications in Statistics: Theory and Methods, 53(1):201–214, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2076124.

Xing:2020:PES

[XZL20] Xiaoyu Xing, Danfeng Zhao, and Bing Li. Parameter estimation for the skew Ornstein–Uhlenbeck processes based on discrete observations. Communications in Statistics: Theory and Methods, 49(9):2176–2188, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1568490.

Xiao:2024:IMP

[XZLB24] Hongmin Xiao, Miaomiao Zhao, Xiang Li, and Aiqin Bai. An innovation mortality prediction model with cohort effect. Communications in Statistics: Theory and Methods, 53(20): 7477–7489, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2023.2264998.

Xiao:2020:TCS

[XZR^{+}20] Helu Xiao, Zhongbao Zhou, Tiantian Ren, Yanfei Bai, and Wenbin Liu. Time-consistent strategies for multiperiod mean-variance portfolio optimization with the serially correlated returns. Communications in Statistics: Theory and Methods, 49(12):2831–2868, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1636999.

Xia:2022:VSP

[XZZ22] Yafeng Xia, Lirong Zhang, and Aiping Zhang. Variable selection for partially varying coefficient model based on modal regression under high dimensional data. Communications in Statistics: Theory and Methods, 51(1):232–248, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1747081.

Y:2024:CAA

[Y.24] Shenbagam R. Sarada Y. On a cost and availability analysis for software systems via phase type nonhomogeneous Poisson process. Communications in Statistics: Theory and Methods, 53(13):4658–4679, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2185473.

Yavarizadeh:2022:WRE

[YA22] Bahareh Yavarizadeh and S. Ejaz Ahmed. The weighted ridge estimation for linear mixed models with measurement error under stochastic linear mixed restrictions. Communications in Statistics: Theory and Methods, 51(6):1605–1621, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1927097.

Yamada:2020:NWH

[Yam20] Hiroshi Yamada. A note on Whittaker–Henderson graduation: Bisymmetry of the smoother matrix. Communications in Statistics: Theory and Methods, 49(7):1629–1634, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1563183.

Yamada:2022:PSD

[Yam22] Hiroshi Yamada. A pioneering study on discrete cosine transform. Communications in Statistics: Theory and Methods, 51(15):5364–5368, 2022. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2020.1838547.

Yamada:2024:HDA

[Yam24] Takayuki Yamada. High-dimensional asymptotic expansion of the null distribution for Schott’s test statistic for complete independence of normal random variables. Communications in Statistics: Theory and Methods, 53(3):909–925, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2094414.

Yang:2020:OIL

[Yan20] Zhaoqiang Yang (Yeung Chiu Keung). Optimal investment and life insurance strategies in a mixed jumpdiffusion framework. Communications in Statistics: Theory and Methods, 49(16):4002–4029, 2020. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2019.1594298.

Yang:2022:CAO

[Yan22a] Dong-Yuh Yang. Computational analysis and optimization of randomized control of N-policy for an M/G/1/K queue with starting failures. Communications in Statistics: Theory and Methods, 51(8):2461–2476, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1776874.

Yang:2022:MHT

[Yan22b] Lin Yang. Multiple hypothesis testing for Poisson processes with variable change-point intensity. Communications in Statistics: Theory and Methods, 51(3):744–766, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1754856.

Yanagisawa:2023:TES

[Yan23] Yukio Yanagisawa. Testing for the equivalence of several sets of time series and its multiple comparison procedure. Communications in Statistics: Theory and Methods, 52(9):3149–3164, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2021.1968902.

Yang:2024:ORI

[Yan24] Peng Yang. Optimal reinsurance-investment problem for two competitive or cooperative insurers under two investment patterns. Communications in Statistics: Theory and Methods, 53(8):3005–3039, 2024. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2022.2150060.

Yao:2022:BIU

[Yao22] Kai Yao. Bayesian inference with uncertain data of imprecise observations. Communications in Statistics: Theory and Methods, 51(15):5330–5341, 2022. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2020.1838545.

Yu:2023:SCR

[YC23] Yuncai Yu and Zhicheng Chen. Strong convergence rates of multiple change-point estimator for ρ-mixing sequence. Communications in Statistics: Theory and Methods, 52(13):4605– 4621, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2021.1998532.

Yang:2023:ESS

[YCC23] Hong-Ding Yang, Yung-Huei Chiou, and Chun-Shu Chen. Estimation and selection for spatial confounding regression models. Communications in Statistics: Theory and Methods, 52(3):939–955, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03610926.2021.1934025.

Yan:2020:OPB

[YCS20] HuaHui Yan, Qihong Chen, and HuiSheng Shu. Option pricing based on a regime switching dividend process. Communications in Statistics: Theory and Methods, 49(24):5964–5974, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2019.1625920.

Yu:2023:SHO

[YCS23] Qian Yu, Qiangqiang Chang, and Guangjun Shen. Smoothness of higher order derivative of self-intersection local time for fractional Brownian motion. Communications in Statistics: Theory and Methods, 52(10):3541–3556, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1977322.

Yang:2021:TCR

[YCW21] Peng Yang, Zhiping Chen, and Liyuan Wang. Timeconsistent reinsurance and investment strategy combining quota-share and excess of loss for mean-variance insurers with jump-diffusion price process. Communications in Statistics: Theory and Methods, 50(11):2546–2568, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1670849.

Yan:2022:ROI

[YCWZ22] Ming Yan, Zheng Cao, Ting Wang, and Shuhua Zhang. Robust optimal investment strategy of DC pension plans with stochastic salary and a return of premiums clause. Communications in Statistics: Theory and Methods, 51(22):7980–8011, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2021.1887236.

Yu:2022:WSC

[YCY22] Yang Yu, John T. Chen, and Arthur B. Yeh. Weighted step-down confidence procedures with applications to gene expression data. Communications in Statistics: Theory and Methods, 51(8):2343–2356, 2022. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2020.1772983.

Yu:2022:ICL

[YD22] Qiqing Yu and Junyi Dong. Identifiability conditions for the linear regression model under right censoring. Communications in Statistics: Theory and Methods, 51(1):116–134, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1743315.

Yu:2023:GRE

[YD23] Ran Yu and Yong Deng. A generalization of Rényi entropy for basic probability assignment. Communications in Statistics: Theory and Methods, 52(19):6991–7008, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2037646.

Yang:2024:OPD

[YD24a] Long Yang and Guohe Deng. Optimal periodic dividends with penalty payments under a diffusion model. Communications in Statistics: Theory and Methods, 53(21):7699–7710, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2272002.

Yin:2024:BFE

[YD24b] Chuancun Yin and Hua Dong. The Bessel function expression of characteristic function. Communications in Statistics: Theory and Methods, 53(22):8009–8025, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2278426.

Yuecaia:2024:LSE

[YD24c] Han Yuecaia and Zhang Dingwen. Least squares estimators for reflected Ornstein–Uhlenbeck processes. Communications in Statistics: Theory and Methods, 53(21):7746–7759, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2273204.

Yang:2024:ELL

[YDHY24] Wenzhi Yang, Ran Ding, Shuhe Hu, and Chunyu Yao. Exact laws of large numbers for k-th order statistics from the asymmetrical Cauchy distribution. Communications in Statistics: Theory and Methods, 53(8):2757–2770, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2148474.

Yalcin:2022:CBB

[YF22] Femin Yalcin and Ceki Franko. Computing Birnbaum and Barlow–Proschan importance measures using system signature. Communications in Statistics: Theory and Methods, 51(23):8446–8457, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03610926.2021.1986536.

Ye:2023:BIB

[YFW^{+}23] Rendao Ye, Bingni Fang, Zhongchi Wang, Kun Luo, and Wenting Ge. Bootstrap inference on the Behrens–Fishertype problem for the skew-normal population under dependent samples. Communications in Statistics: Theory and Methods, 52(11):3751–3766, 2023. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2021.1980045.

Yaseen:2020:SVE

[YG20] Abdallah S. A. Yaseen and Ahmed M. Gad. A stochastic variant of the EM algorithm to fit mixed (discrete and continuous) longitudinal data with nonignorable missingness. Communications in Statistics: Theory and Methods, 49(18):4446–4467, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2019.1601223.

Yilmaz:2023:WSC

[YGAK23] Emrah Yilmaz, Tuba Gulsen, Yavuz Altin, and Hikmet Koyunbakan. λ-Wijsman statistical convergence on time scales. Communications in Statistics: Theory and Methods, 52(15):5364–5378, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03610926.2021.2006716.

Yan:2024:CPE

[YGG24] Litan Yan, Rui Guo, and Han Gao. Convergence and parameter estimation of the linear weighted-fractional selfrepelling diffusion. Communications in Statistics: Theory and Methods, 53(7):2390–2421, 2024. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2022.2132828.

Yuan:2023:OIS

[YH23] Haili Yuan and Yijun Hu. Optimal investment strategies for an insurer with liquid constraint. Communications in Statistics: Theory and Methods, 52(7):2198–2214, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1945634.

Yu:2024:RSC

[YH24] Han Yu and Alan D. Hutson. A robust Spearman correlation coefficient permutation test. Communications in Statistics: Theory and Methods, 53(6):2141–2153, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2121144.

Yazdi:2020:AEP

[YHA20] Ahmad Ahmadi Yazdi, Ali Zeinal Hamadani, and Amirhossein Amiri. Addressing the effect of parameter estimation on phase II monitoring of multivariate multiple linear profiles via a new cluster-based approach. Communications in Statistics: Theory and Methods, 49(17):4117–4132, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1594303.

Yang:2024:MSA

[YHH24] Cheng-Han Yang, Ya-Hsuan Hsu, and Cheng-Hung Hu. Misspecification analyses and optimum degradation test plan for Wiener and inverse Gaussian processes. Communications in Statistics: Theory and Methods, 53(2):700–717, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2091782.

Yin:2024:NBE

[Yin24] Xuehua Yin. New bounds on entropies based on order statistics and Gini’s mean difference. Communications in Statistics: Theory and Methods, 53(19):7013–7030, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2256438.

Yuyun:2023:ATM

[YJZZ23] Wang Yuyun, Luo Jing, Xu Zhimeng, and Lewei Zhou. Asymptotics in the Thurstone model with an increasing items. Communications in Statistics: Theory and Methods, 52(22): 8038–8052, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2022.2056201.

Yu:2020:PCT

[YK20] GangHyok Yu and SongGuk Kim. Parameter change test for periodic integer-valued autoregressive process. Communications in Statistics: Theory and Methods, 49(12):2898–2912, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2019.1584309.

Yang:2023:IBE

[YK23] Wenyu Yang and Xiaoning Kang. An improved banded estimation for large covariance matrix. Communications in Statistics: Theory and Methods, 52(1):141–155, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1910839.

Yu:2020:TET

[YKH20] Jianqi Yu, Kalimuthu Krishnamoorthy, and Yafei He. Testing equality of two normal covariance matrices with monotone missing data. Communications in Statistics: Theory and Methods, 49(16):3911–3918, 2020. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2019.1591453.

Yu:2023:MBF

[YKW23] Jianqi Yu, Kalimuthu Krishnamoorthy, and Bin Wang. Multivariate Behrens–Fisher problem using means of auxiliary variables. Communications in Statistics: Theory and Methods, 52(17):6103–6110, 2023. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2022.2026392.

Ye:2023:UST

[YL23] Tingqing Ye and Baoding Liu. Uncertain significance test for regression coefficients with application to regional economic analysis. Communications in Statistics: Theory and Methods, 52(20):7271–7288, 2023. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2022.2042562.

Yuan:2022:SSP

[YLC^{+}22] Yefang Yuan, Xiaohui Liu, Yuting Chen, Yuxin Dong, and Yuzi Liu. On similarity of the sample projection depth contours and its application. Communications in Statistics: Theory and Methods, 51(7):1919–1935, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1802651.

Yang:2023:ADR

[YLJ23] Zhen Yang, Yihui Luan, and Jiming Jiang. The asymptotic distribution of robust maximum likelihood estimator with Huber function for the mixed spatial autoregressive model with outliers. Communications in Statistics: Theory and Methods, 52(18):6311–6340, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2027985.

Yang:2021:MRP

[YLL21] Guangren Yang, Hongmei Lin, and Heng Lian. Minimax rate in prediction for functional principal component regression. Communications in Statistics: Theory and Methods, 50(5): 1240–1249, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2019.1649429.

Yang:2024:SCP

[YLL^{+}24] Shanchao Yang, Shuyi Luo, Zhiyong Li, Jiaying Xie, and Xin Yang. Strong consistency of parameter estimation for the CIR integrated diffusion process with longspan high-frequency data. Communications in Statistics: Theory and Methods, 53(22):8041–8061, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2278429.

Yu:2024:RLA

[YLLZ24] Jinzhao Yu, Daoji Li, Lin Luo, and Hui Zhao. Reproducible learning for accelerated failure time models via deep knockoffs. Communications in Statistics: Theory and Methods, 53(18): 6544–6560, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2023.2247508.

Yu:2023:SEI

[YLW^{+}23] Qin Yu, Yang Li, Yumeng Wang, Yachong Yang, and Zemin Zheng. Scalable and efficient inference via CPE. Communications in Statistics: Theory and Methods, 52(5):1614–1633, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2021.1936044.

Yang:2024:CIM

[YLW24] Xiaowei Yang, Xinqiao Liu, and Haoyu Wei. Concentration inequalities of MLE and robust MLE. Communications in Statistics: Theory and Methods, 53(19):6944–6956, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2253945.

Yousfi:2022:ATP

[YM22] A. Yousfi and M. Merzougui. Adaptive test for periodicity in restrictive EXPAR(p) models. Communications in Statistics: Theory and Methods, 51(17):6064–6077, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1852433.

Yamaguchi:2024:IDB

[YM24] Hikaru Yamaguchi and Hidetoshi Murakami. Interpoint distance-based two-sample tests for functional data. Communications in Statistics: Theory and Methods, 53(8):2771–2791, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2022.2148475.

Yadav:2024:IFE

[YMG24] Subhash Kumar Yadav, Sheela Misra, and Sat Gupta. Improved family of estimators of population coefficient of variation under simple random sampling. Communications in Statistics: Theory and Methods, 53(2):727–747, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2091784.

Yue:2023:PPE

[YMZD23] Shengjie Yue, Chaoqun Ma, Xinwei Zhao, and Chao Deng. Pricing power exchange options with default risk, stochastic volatility and stochastic interest rate. Communications in Statistics: Theory and Methods, 52(5):1431–1456, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1928202.

Yang:2021:LSE

[YN21a] Xiangfeng Yang and Yaodong Ni. Least-squares estimation for uncertain moving average model. Communications in Statistics: Theory and Methods, 50(17):4134–4143, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1713373.

Yuan:2021:TDS

[YN21b] Mingao Yuan and Yehong Nan. Test dense subgraphs in sparse uniform hypergraph. Communications in Statistics: Theory and Methods, 50(20):4743–4762, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1723637.

Yonghint:2024:PAE

[YN24] N. Yonghint and K. Neammanee. Poisson approximation for the expectation of call function with application in collateralized debt obligation. Communications in Statistics: Theory and Methods, 53(14):5265–5279, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2215359.

Yonghint:2022:PAL

[YNC22] Nat Yonghint, Kritsana Neammanee, and Nattakarn Chaidee. Poisson approximation for locally dependent CDO. Communications in Statistics: Theory and Methods, 51(7):2073–2081, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2020.1759638.

Yo:2022:AEM

[Yo22] Sheena Yo. Asymptoic efficiency of M.L.E. using prior survey in multinomial distributions. Communications in Statistics: Theory and Methods, 51(3):701–723, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1753077.

Yankam:2023:ROU

[YO23] Brenda Mbouamba Yankam and Abimibola Victoria Oladugba. Robustness of orthogonal uniform composite designs against missing data. Communications in Statistics: Theory and Methods, 52(5):1369–1384, 2023. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2021.1927095.

Yoshida:2021:EVI

[Yos21] Takuma Yoshida. Extreme value inference for quantile regression with varying coefficients. Communications in Statistics: Theory and Methods, 50(3):685–710, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1639752.

Yoshida:2023:APS

[Yos23] Takuma Yoshida. Asymptotics for penalized spline estimators in quantile regression. Communications in Statistics: Theory and Methods, 52(14):4815–4834, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2013.765477.

Youndje:2023:CKQ

[You23a] É. Youndjé. L^{2} consistency of the kernel quantile estimator. Communications in Statistics: Theory and Methods, 52(17): 6111–6125, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2022.2026393.

Younso:2023:CFP

[You23b] Ahmad Younso. Consistency of the frequency polygon estimators of density mode for strongly mixing processes. Communications in Statistics: Theory and Methods, 52(7):2182–2197, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2021.1945633.

Youssef:2024:UKL

[You24] Noha Youssef. The use of the Karhunen Loève expansion in the design of computer experiments. Communications in Statistics: Theory and Methods, 53(18):6394–6416, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2245085.

Yang:2020:OCI

[YP20] Mo Yang and Borek Puza. Optimal confidence intervals for the geometric parameter. Communications in Statistics: Theory and Methods, 49(3):590–606, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1549242.

Yadav:2024:EME

[YP24] Vinay Kumar Yadav and Shakti Prasad. Exponential method of estimation in sampling theory under robust quantile regression methods. Communications in Statistics: Theory and Methods, 53(17):6285–6298, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2243529.

Yang:2023:BIQ

[YPD23] Kai Yang, Bo Peng, and Xiaogang Dong. Bayesian inference for quantile autoregressive model with explanatory variables. Communications in Statistics: Theory and Methods, 52(9): 2946–2965, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2021.1964529.

Yan:2021:NFE

[YPXZ21] Xingyu Yan, Xiaolong Pu, Xiaolei Xun, and Yingchun Zhou. A new functional estimation procedure for varying coefficient models. Communications in Statistics: Theory and Methods, 50(5):1117–1133, 2021. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2019.1646767.

Yavarizadeh:2020:EPL

[YRB20] Bahareh Yavarizadeh, Abdolrahman Rasekh, and Babak Babadi. Estimation of parameters in linear mixed measurement error models with stochastic linear restrictions. Communications in Statistics: Theory and Methods, 49(23):5853– 5865, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2019.1622730.

Yaqub:2020:EPD

[YS20] Mazhar Yaqub and Javid Shabbir. Estimation of population distribution function involving measurement error in the presence of non response. Communications in Statistics: Theory and Methods, 49(10):2540–2559, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1580738.

Yamada:2022:HDA

[YSF22] Takayuki Yamada, Tetsuro Sakurai, and Yasunori Fujikoshi. High-dimensional asymptotic results for EPMCs of W - and Z-rules. Communications in Statistics: Theory and Methods, 51(8):2385–2413, 2022. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2020.1774060.

Shi:2022:PSI

[ySLZ22] Xiang yu Shi, Bo Liang, and Qi Zhang. Post-selection inference of generalized linear models based on the lasso and the elastic net. Communications in Statistics: Theory and Methods, 51(14):4739–4756, 2022. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2020.1821892.

Yang:2022:RAC

[YSS22a] Dian Yang, Nicholas Sun, and Jianguo Sun. Regression analysis of clustered interval-censored failure time data with cure fraction and informative cluster size. Communications in Statistics: Theory and Methods, 51(22):7737–7750, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1879861.

Yennum:2022:ISC

[YSS22b] Niharika Yennum, Stephen A. Sedory, and Sarjinder Singh. Improved strategy to collect sensitive data by using negative binomial and negative hypergeometric distribution as randomization devices. Communications in Statistics: Theory and Methods, 51(8):2640–2658, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1777565.

Yasmeen:2020:VEA

[YT20] Uzma Yasmeen and Mary Thompson. Variance estimation in adaptive cluster sampling. Communications in Statistics: Theory and Methods, 49(10):2485–2497, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1576890.

Yabuno:2022:NES

[YT22] Mashu Yabuno and Hidekazu Tanaka. A note on estimation of a shape parameter in a Pareto distribution. Communications in Statistics: Theory and Methods, 51(21):7561–7574, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2021.1875241.

Yu:2023:UBE

[YTT23] Miaomiao Yu, Jianfang Tang, and Yinghui Tang. UMVUEs and Bayes estimators for various performance measures on a Poisson queue with discouraged arrivals. Communications in Statistics: Theory and Methods, 52(13):4468–4483, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1995430.

Yu:2020:SIV

[Yu20] Qian Yu. Statistical inference for Vasicek-type model driven by self-similar Gaussian processes. Communications in Statistics: Theory and Methods, 49(2):471–484, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1543774.

Yu:2022:BAF

[Yu22a] I-Tang Yu. A Bayesian approach to factor screening in life tests. Communications in Statistics: Theory and Methods, 51(6):1778–1790, 2022. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2020.1768270.

Yu:2022:BNH

[Yu22b] I-Tang Yu. Bayesian non-homogeneous cumulative probability models for ordinal data from designed experiments. Communications in Statistics: Theory and Methods, 51(17):6008–6020, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2020.1851719.

Yu:2022:NMM

[Yu22c] Wei Yu. A new method for multi-sample high-dimensional covariance matrices test based on permutation. Communications in Statistics: Theory and Methods, 51(13):4476–4486, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2020.1815782.

Yuan:2021:SMM

[Yua21] Kang Fang Yuan. Semiparametric mean model with non linear time effect of the covariate for clustered recurrent events with terminal events. Communications in Statistics: Theory and Methods, 50(11):2640–2658, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1674872.

Yuan:2024:PTP

[Yua24] Mingao Yuan. Phase transitions in a power-law uniform hypergraph. Communications in Statistics: Theory and Methods, 53(4):1257–1276, 2024. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2022.2097265.

Yu:2020:NMS

[YW20] Conglian Yu and Xiyang Wang. A new model selection procedure for finite mixture regression models. Communications in Statistics: Theory and Methods, 49(18):4347–4366, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1601222.

Yan:2022:CLV

[YW22] Rongfang Yan and Junrui Wang. Component level versus system level at active redundancies for coherent systems with dependent heterogeneous components. Communications in Statistics: Theory and Methods, 51(6):1724–1744, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1767140.

Yang:2024:ESP

[YW24] Yang Yang and Lichun Wang. Estimation of structural parameters in balanced Bühlmann credibility model with correlation risk. Communications in Statistics: Theory and Methods, 53(8):2945–2960, 2024. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2022.2150057.

Yuan:2020:VSS

[YWL20] Xiaohui Yuan, Yue Wang, and Tianqing Liu. Variable selection for semiparametric random-effects conditional density models with longitudinal data. Communications in Statistics: Theory and Methods, 49(4):977–996, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1554130.

Yin:2022:NCS

[YWS22] Chuancun Yin, Yeshunying Wang, and Xiuyan Sha. A new class of symmetric distributions including the elliptically symmetric logistic. Communications in Statistics: Theory and Methods, 51(13):4537–4558, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1818098.

Yang:2024:HOR

[YXP24] Xi Yang, Qian Xiong, and Zuoxiang Peng. Higher-order representation of Karamata theorem. Communications in Statistics: Theory and Methods, 53(8):2744–2756, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2148473.

Yang:2024:IEM

[YXS^{+}24] Xiaoyu Yang, Liyang Xie, Jiaxin Song, Bingfeng Zhao, and Yuan Li. On interval estimation methods for the location parameter of the Weibull distribution: an application to alloy material fatigue failure data. Communications in Statistics: Theory and Methods, 53(17):6240–6251, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2242984.

Yu:2021:ENE

[YXZ21] Wei Yu, Wangli Xu, and Lixing Zhu. Estimating the number of equal components for two high-dimensional mean vectors. Communications in Statistics: Theory and Methods, 50(19): 4617–4638, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2020.1722842.

Yan:2023:ASC

[YXZ23] Hui Yan, Lifeng Xu, and Shaoyi Zhang. Almost sure central limit theorem for non-stationary Markov chains. Communications in Statistics: Theory and Methods, 52(15):5187–5194, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2021.2004426.

Yang:2021:RCA

[YY21] Yuehan Yang and Hu Yang. Rates of convergence of the adaptive elastic net and the post-selection procedure in ultra-high dimensional sparse models. Communications in Statistics: Theory and Methods, 50(1):73–94, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1628991.

Yevkin:2024:RAP

[YY24] Glib Yevkin and Olexandr Yevkin. On regression analysis with Padé approximants. Communications in Statistics: Theory and Methods, 53(22):8026–8040, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2278428.

Yu:2020:LTR

[YZ20] Xianye Yu and Mingbo Zhang. Limit theorems related to the integral functionals of one dimensional fractional Brownian motion. Communications in Statistics: Theory and Methods, 49(24):5908–5916, 2020. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2019.1622727.

Ye:2021:JRV

[YZ21] Xu-Guo Ye and Yan-Yong Zhao. Jump-robust volatility estimation using dynamic dual-domain integration method. Communications in Statistics: Theory and Methods, 50(5):1250– 1273, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2019.1650183.

Yu:2023:LLL

[YZ23] Haichao Yu and Yong Zhang. Law of the logarithm and law of the iterated logarithm for a class of random variables with non-identical distributions. Communications in Statistics: Theory and Methods, 52(22):8169–8183, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2059680.

Yang:2022:EAE

[YZC22] Jianfeng Yang, Ming Zhao, and Jing Chen. ELS algorithm for estimating open source software reliability with masked data considering both fault detection and correction processes. Communications in Statistics: Theory and Methods, 51(19): 6792–6817, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2020.1866610.

Yi:2023:OPR

[YZSS23] Haoran Yi, Xuekang Zhang, Yuanchuang Shan, and Huisheng Shu. Optimal portfolio and reinsurance with two differential risky assets. Communications in Statistics: Theory and Methods, 52(19):7094–7114, 2023. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2022.2039708.

Yuan:2021:SAN

[YZT21] Ao Yuan, Yizhao Zhou, and Ming T. Tan. Subgroup analysis with a nonparametric unimodal symmetric error distribution. Communications in Statistics: Theory and Methods, 50(17): 4000–4021, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2019.1710754.

Yu:2023:MRP

[YZW23] Han Yu, Yu Zhang, and Xikui Wang. Minimization of ruin probability with joint strategies of investment and reinsurance. Communications in Statistics: Theory and Methods, 52(15):5451–5469, 2023. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2021.2009870.

Yang:2024:SCN

[YZXY24] Shanchao Yang, Shi Zhang, Guodong Xing, and Xin Yang. Strong consistency of nonparametric kernel estimators for integrated diffusion process. Communications in Statistics: Theory and Methods, 53(8):2792–2815, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2148540.

Yan:2021:SRC

[YZZ21] Xingyu Yan, Yiying Zhang, and Peng Zhao. Standby redundancies at component level versus system level in series system. Communications in Statistics: Theory and Methods, 50(2):473–485, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03610926.2019.1635705.

Zamini:2024:BEB

[ZAF24] R. Zamini, M. Ajami, and V. Fakoor. Berry–Esseen bound for smooth estimator of distribution function under length-biased data. Communications in Statistics: Theory and Methods, 53(5):1800–1809, 2024. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2022.2112695.

Zamini:2024:KEM

[ZAG24] R. Zamini, M. Ajami, and S. Ghafouri. Kernel estimators for mean residual lifetime in length-biased sampling. Communications in Statistics: Theory and Methods, 53(22):7927–7941, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2023.2277129.

Zhang:2021:EMC

[ZAL21] Suhua Zhang, Chunxiang A., and Yongzeng Lai. Efficient multiple control variate method with applications to exotic option pricing. Communications in Statistics: Theory and Methods, 50(6):1275–1294, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1648829.

Zalzadeh:2022:DCC

[Zal22] Saeed Zalzadeh. On the dependency of the components in complex systems. Communications in Statistics: Theory and Methods, 51(3):624–635, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1752724.

Zaman:2020:MRE

[ZB20] Tolga Zaman and Hasan Bulut. Modified regression estimators using robust regression methods and covariance matrices in stratified random sampling. Communications in Statistics: Theory and Methods, 49(14):3407–3420, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1588324.

Zaman:2023:EFR

[ZB23a] Tolga Zaman and Hasan Bulut. An efficient family of robust-type estimators for the population variance in simple and stratified random sampling. Communications in Statistics: Theory and Methods, 52(8):2610–2624, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1955388.

Zemoul:2023:ABL

[ZB23b] Sara-Imane Zemoul and Youcef Berkoun. Asymptotic behavior of LSE estimator of an AR(1) coefficient with associated innovations. Communications in Statistics: Theory and Methods, 52(21):7780–7787, 2023. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2022.2071941.

Zandi:2024:SEZ

[ZBB24] Zahra Zandi, Hossein Bevrani, and Reza Arabi Belaghi. Shrinkage estimation in the zero-inflated Poisson regression model with right-censored data. Communications in Statistics: Theory and Methods, 53(13):4898–4917, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2196751.

Zhang:2021:RCE

[ZC21] Qiang Zhang and Ping Chen. Regression credibility estimator with two-level common effects. Communications in Statistics: Theory and Methods, 50(4):910–931, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1643887.

Zi:2022:RTE

[ZC22] Xuemin Zi and Hui Chen. Robust tests of the equality of two high-dimensional covariance matrices. Communications in Statistics: Theory and Methods, 51(10):3120–3141, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1788085.

Zhang:2023:CLI

[ZC23] Lvyun Zhang and Shouquan Chen. A class of location invariant estimators for heavy tailed distributions. Communications in Statistics: Theory and Methods, 52(3):896–917, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1931335.

Zhang:2024:EPD

[ZC24] Jun Zhang and Leyi Cui. Exponential parametric distortion nonlinear measurement errors models. Communications in Statistics: Theory and Methods, 53(5):1777–1799, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2111526.

Zhang:2022:NES

[ZCCT22] Yangchun Zhang, Jiaqi Chen, Bosen Cui, and Boping Tian. Nonparametric estimate of spectral density functions of sample covariance matrices generated by VARMA models. Communications in Statistics: Theory and Methods, 51(4):943– 952, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2020.1745842.

Zolfaghari:2023:NRB

[ZCE23] Parisa Zolfaghari, Rahim Chinipardaz, and Jafar Esmaily. The numerical reconcilability of Bayesian measure and p-value in interval hypotheses is not possible in general. Communications in Statistics: Theory and Methods, 52(4):1178–1189, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2021.1924785.

Zeng:2024:RWM

[ZCL24] Zhen Zeng, Lin Cong, and Xiangdong Liu. Random weighting method for M-test in linear model with dependent errors. Communications in Statistics: Theory and Methods, 53(4): 1381–1401, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2022.2101119.

Zhang:2020:FOR

[ZCT^{+}20] Shuxia Zhang, Xinrong Cong, Boping Tian, Yanpeng Li, and Mingjun Yao. The first-order random coefficient integer valued autoregressive process with the occasional level shift random noise based on dual empirical likelihood. Communications in Statistics: Theory and Methods, 49(12):2990–3009, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1584315.

Zeng:2023:EFM

[ZCW23] Xin Zeng, Xingyun Cao, and Liucang Wu. Estimation for finite mixture of mode regression models using skew-normal distribution. Communications in Statistics: Theory and Methods, 52(20):7479–7501, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2048026.

Zhang:2020:EPS

[ZCZ20a] Zhiyi Zhang, Chen Chen, and Jialin Zhang. Estimation of population size in entropic perspective. Communications in Statistics: Theory and Methods, 49(2):307–324, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1536786.

Zhao:2020:ETR

[ZCZ20b] Xu Zhao, Weihu Cheng, and Pengyue Zhang. Extreme tail risk estimation with the generalized Pareto distribution under the peaks-over-threshold framework. Communications in Statistics: Theory and Methods, 49(4):827–844, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1549253.

Zhou:2020:TET

[ZCZJ20] Yuejin Zhou, Ting Chen, Qianjin Zhao, and Tao Jiang. Testing the equality of two double-parameter exponential distributions via overlap coefficient. Communications in Statistics: Theory and Methods, 49(5):1248–1260, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1563169.

Zhou:2023:BEM

[ZD23] Qianli Zhou and Yong Deng. Belief eXtropy: Measure uncertainty from negation. Communications in Statistics: Theory and Methods, 52(11):3825–3847, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1980049.

Zhu:2022:OID

[ZDW22] Shunqing Zhu, Yinghui Dong, and Sang Wu. Optimal investment of DC pension plan with two VaR constraints. Communications in Statistics: Theory and Methods, 51(6):1745–1764, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2020.1767141.

Zhang:2024:CAO

[ZDY24] Tian-Fang Zhang, Yingxing Duan, and Jian-Feng Yang. Construction of asymmetric orthogonal arrays of high strength via generator matrix. Communications in Statistics: Theory and Methods, 53(17):6206–6223, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2242982.

Zhao:2022:EDS

[ZDZ22] Yongxia Zhao, Hua Dong, and Wei Zhong. Equilibrium dividend strategies for spectrally negative Lévy processes with time value of ruin and random time horizon. Communications in Statistics: Theory and Methods, 51(14):4757–4780, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2020.1822407.

Zeinal:2023:ETT

[Zei23] Amir Zeinal. The extended two-type parameter estimator in linear regression model. Communications in Statistics: Theory and Methods, 52(2):463–478, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1916528.

Zeng:2020:CMC

[Zen20a] Guoping Zeng. On the confusion matrix in credit scoring and its analytical properties. Communications in Statistics: Theory and Methods, 49(9):2080–2093, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1568485.

Zeng:2020:SLR

[Zen20b] Xingyuan Zeng. Spectrum of large random inner-product kernel matrices generated from l_{p} ellipsoids. Communications in Statistics: Theory and Methods, 49(19):4636–4647, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1604962.

Zeng:2022:GTV

[Zen22] Guoping Zeng. A graphic and tabular variable deduction method in logistic regression. Communications in Statistics: Theory and Methods, 51(16):5412–5427, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1839499.

Zeng:2023:APC

[Zen23] Guoping Zeng. On the analytical properties of category encodings in logistic regression. Communications in Statistics: Theory and Methods, 52(6):1870–1887, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1939382.

Zeng:2024:WLE

[Zen24] Xiaoqiang Zeng. Whittle likelihood estimation in INAR(1) process. Communications in Statistics: Theory and Methods, 53(17):6177–6196, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03610926.2023.2241093.

Zeybek:2020:IRD

[Zey20] Melis Zeybek. Interval robust design under contaminated and non normal data. Communications in Statistics: Theory and Methods, 49(22):5406–5418, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1710198.

Zhang:2024:TLS

[ZFDT24] Jingke Zhang, Beibei Fan, Umer Daraz, and Yu Tang. Threelevel saturated orthogonal arrays with less β-wordlength pattern. Communications in Statistics: Theory and Methods, 53(14):5218–5229, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03610926.2023.2209680.

Zhang:2024:TSM

[ZFZ24] Jun Zhang, Zhenghui Feng, and Yue Zhou. Testing symmetry of model errors for non linear multiplicative distortion measurement error models. Communications in Statistics: Theory and Methods, 53(18):6427–6448, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2245639.

Zhang:2021:SIL

[ZG21] Yue Zhang and Wenhao Gui. Statistical inference for the lifetime performance index of products with Pareto distribution on basis of general progressive type II censored sample. Communications in Statistics: Theory and Methods, 50(16):3790– 3808, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2020.1801735.

Zhao:2023:NRP

[ZG23] Yunfeng Zhao and Jing Guan. A new robust parameter estimation approach for multinomial categorical response data with outliers and mismeasured covariates. Communications in Statistics: Theory and Methods, 52(17):6179–6206, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2027447.

Zhao:2022:VSL

[ZGC22] Mingtao Zhao, Yuzhao Gao, and Yuehua Cui. Variable selection for longitudinal varying coefficient errorsin-variables models. Communications in Statistics: Theory and Methods, 51(11):3713–3738, 2022. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2020.1801738.

Zhao:2023:OBP

[ZGCZ23] Peixin Zhao, Haogeng Gan, Suli Cheng, and Xiaoshuang Zhou. Orthogonality based penalized GMM estimation for variable selection in partially linear spatial autoregressive models. Communications in Statistics: Theory and Methods, 52(6):1676–1691, 2023. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2021.1937652.

Zhao:2020:ACA

[ZGQN20] Xufeng Zhao, Chen Gao, Cunhua Qian, and Toshio Nakagawa. Approximate calculations of age-based random replacement times. Communications in Statistics: Theory and Methods, 49(15):3808–3820, 2020. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2019.1710203.

Zhu:2024:ODS

[ZH24] Xiaoyuan Zhu and Honghua Hao. A-optimal design for the special cubic mixture model. Communications in Statistics: Theory and Methods, 53(3):1081–1090, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2100907.

Zhang:2020:FRL

[Zha20a] Yong Zhang. Further research on limit theorems for selfnormalized sums. Communications in Statistics: Theory and Methods, 49(2):385–402, 2020. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2018.1543767.

Zohrevand:2020:ACK

[ZHA20b] Younes Zohrevand, Reza Hashemi, and Majid Asadi. An adjusted cumulative Kullback–Leibler information with application to test of exponentiality. Communications in Statistics: Theory and Methods, 49(1):44–60, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1529243.

Zhang:2021:ERM

[Zha21a] Lili Zhang. The Erlang(n) risk model with two-sided jumps and a constant dividend barrier. Communications in Statistics: Theory and Methods, 50(24):5899–5917, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1737712.

Zhao:2021:GFC

[Zha21b] Haibing Zhao. Generalized FWER control procedures for testing multiple hypotheses. Communications in Statistics: Theory and Methods, 50(22):5399–5410, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1728555.

Zhao:2022:BCI

[Zha22] Haibing Zhao. Bayesian confidence intervals with more power to estimate the minimum effect and determine the sign. Communications in Statistics: Theory and Methods, 51(3):636– 648, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2020.1752725.

Zhang:2024:PLA

[Zha24a] Jun Zhang. Partial linear additive distortion measurement errors models. Communications in Statistics: Theory and Methods, 53(1):232–259, 2024. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2022.2076126.

Zhang:2024:SLL

[Zha24b] Zhao-Ang Zhang. Strong law of large numbers for linear processes under sublinear expectation. Communications in Statistics: Theory and Methods, 53(6):2205–2218, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2122841.

Zhang:2021:CCE

[ZHQ21] Shuili Zhang, Tiantian Hou, and Cong Qu. Complete consistency for the estimator of nonparametric regression model based on martingale difference errors. Communications in Statistics: Theory and Methods, 50(2):358–370, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1635160.

Zhang:2023:CPC

[ZHQ23] Shuili Zhang, Tiantian Hou, and Cong Qu. Consistency of the P-C estimator in non parametric regression model based on m-END errors. Communications in Statistics: Theory and Methods, 52(4):1190–1201, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1924786.

Zhu:2024:REM

[ZHS^{+}24] Chao Zhu, David W. Hosmer, Jim Stankovich, Karen Wills, and Leigh Blizzard. Refinements on the exact method to solve the numerical difficulties in fitting the log binomial regression model for estimating relative risk. Communications in Statistics: Theory and Methods, 53(23):8359–8375, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2284674.

Zhu:2022:CAG

[Zhu22a] Yunlong Zhu. Covariate-adjusted Gaussian graphical model estimation with false discovery rate control. Communications in Statistics: Theory and Methods, 51(4):974–993, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1752385.

Zhu:2022:HTH

[Zhu22b] Yunlong Zhu. Hypothesis testing for high-dimensional multivariate regression with false discovery rate control. Communications in Statistics: Theory and Methods, 51(21):7476–7495, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2021.1873378.

Zaman:2023:RTO

[ZIZ23] Qamruz Zaman, Muhammad Ijaz, and Tolga Zaman. A randomization tool for obtaining efficient estimators through focus group discussion in sensitive surveys. Communications in Statistics: Theory and Methods, 52(10):3414–3428, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1973502.

Zhang:2023:PQR

[ZJF23] Yuanqing Zhang, Jiayuan Jiang, and Yaqin Feng. Penalized quantile regression for spatial panel data with fixed effects. Communications in Statistics: Theory and Methods, 52(4): 1287–1299, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2021.1934028.

Zaman:2021:NCE

[ZK21] Tolga Zaman and Cem Kadilar. New class of exponential estimators for finite population mean in two-phase sampling. Communications in Statistics: Theory and Methods, 50(4): 874–889, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2019.1643480.

Zhang:2024:VSH

[ZK24] Yunxi Zhang and Soeun Kim. Variable selection for high-dimensional incomplete data using horseshoe estimation with data augmentation. Communications in Statistics: Theory and Methods, 53(12):4235–4251, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2177107.

Zhang:2020:SLL

[ZL20] Ning Zhang and Yuting Lan. A strong law of large numbers for independent random variables under nonadditive probabilities. Communications in Statistics: Theory and Methods, 49(21):5252–5272, 2020. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2019.1615508.

Zhang:2021:MMA

[ZL21] Yangchun Zhang and Pengfei Liu. Median-of-means approach for repeated measures data. Communications in Statistics: Theory and Methods, 50(17):3903–3912, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1710204.

Zhang:2022:PDB

[ZL22a] Tianqi Zhang and Bingqing Li. Pricing double-barrier option with processes depending on various states of the economy. Communications in Statistics: Theory and Methods, 51(23): 8296–8309, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2021.1891439.

Zhao:2022:ABC

[ZL22b] Jing Zhao and Shenghong Li. Asymptotic behavior and calibration of short-time option prices under the normal tempered stable model. Communications in Statistics: Theory and Methods, 51(16):5428–5445, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1839774.

Zhao:2022:QQP

[ZL22c] Qiang Zhao and Jiajuan Liang. A Q–Q plot for detecting non-multinormality based on a normal characterization and the S–W statistic. Communications in Statistics: Theory and Methods, 51(5):1367–1378, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1761983.

Zhou:2022:CPE

[ZL22d] Hong-Bing Zhou and Han-Ying Liang. Change point estimation in regression model with response missing at random. Communications in Statistics: Theory and Methods, 51(20): 7101–7119, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2020.1871017.

Zhang:2023:CDB

[ZL23a] Guoyi Zhang and Yan Lu. Comparison of difference based variance estimators for partially linear models. Communications in Statistics: Theory and Methods, 52(23):8454–8466, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2022.2064498.

Zhang:2023:SLS

[ZL23b] Xiao Zhang and Yiming Liu. Sparse Laplacian shrinkage for nonparametric transformation survival model. Communications in Statistics: Theory and Methods, 52(20):7184–7205, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2022.2042025.

Zhang:2020:IAR

[ZLB20] Chao Zhang, Tao Liu, and Guanghan Bai. An improved algorithm for reliability bounds of multistate networks. Communications in Statistics: Theory and Methods, 49(15):3772–3791, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2020.1752728.

Zhou:2020:ESC

[ZLCW20] Meng Zhou, Liwei Liu, Xudong Chai, and Zhen Wang. Equilibrium strategies in a constant retrial queue with setup time and the N-policy. Communications in Statistics: Theory and Methods, 49(7):1695–1711, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1565779.

Zhang:2024:SCE

[ZLF24] Jing Zhang, Aoshuang Li, and Rui Fang. Stochastic comparisons on extremes of Burr type XII samples associated with Archimedean copula and heterogeneous shape parameters. Communications in Statistics: Theory and Methods, 53(2): 786–811, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2022.2092146.

Zhang:2020:WCE

[ZLH20] Yu Zhang, Xinsheng Liu, and Hongchang Hu. Weak consistency of M-estimator in linear regression model with asymptotically almost negatively associated errors. Communications in Statistics: Theory and Methods, 49(11):2800–2816, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1584307.

Zhang:2021:ODA

[ZLLS21] Min Zhang, Yong Li, Jun Lu, and Lei Shi. Outlier detection and accommodation in meta-regression models. Communications in Statistics: Theory and Methods, 50(8):1728–1744, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2019.1652321.

Zheng:2023:HDS

[ZLLZ23] Zemin Zheng, Lei Liu, Yang Li, and Ni Zhao. Highdimensional statistical inference via DATE. Communications in Statistics: Theory and Methods, 52(1):65–79, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1909733.

Zhu:2022:SCB

[ZLNS22] Hong Zhu, Yu Lan, Jing Ning, and Yu Shen. Semiparametric copula-based regression modeling of semi-competing risks data. Communications in Statistics: Theory and Methods, 51(22):7830–7845, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03610926.2021.1881122.

Zhang:2024:CSN

[ZLPZ24] Qingjuan Zhang, Yuan Li, Shanqi Pang, and Xingfa Zhang. On the construction of some new asymmetric orthogonal arrays. Communications in Statistics: Theory and Methods, 53(2):475–486, 2024. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2022.2083166.

Zhang:2023:EGT

[ZLS23a] Rongmao Zhang, Qimeng Liu, and Jianhua Shi. Estimation of generalized threshold autoregressive models. Communications in Statistics: Theory and Methods, 52(18):6456–6474, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2022.2029896.

Zong:2023:DGR

[ZLS23b] Yiwei Zong, Ffion Loring, and William F. Scott. On the distribution of Gini’s rank association index. Communications in Statistics: Theory and Methods, 52(21):7788–7796, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2071942.

Zhang:2021:EDG

[ZLW21] Lingyue Zhang, Dawei Lu, and Xiaoguang Wang. The essential dependence for a group of random vectors. Communications in Statistics: Theory and Methods, 50(24):5836–5872, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2020.1737128.

Zhao:2020:VCC

[ZLZ20a] Qian Zhao, Peng Li, and Jie Zhang. Valuation of contingent claims with stochastic interest rate and mortality driven by Lévy processes. Communications in Statistics: Theory and Methods, 49(14):3421–3437, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1589514.

Zhou:2020:SCC

[ZLZ20b] Wenhui Zhou, Na Liu, and Zhibin Zheng. A synthetic control chart for monitoring the small shifts in a process mean based on an attribute inspection. Communications in Statistics: Theory and Methods, 49(9):2189–2204, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1568491.

Zhang:2024:RHD

[ZLZZ24] Jie Zhang, Yang Li, Ni Zhao, and Zemin Zheng. L_{0}- regularization for high-dimensional regression with corrupted data. Communications in Statistics: Theory and Methods, 53(1):215–231, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03610926.2022.2076125.

Zhang:2024:ROD

[ZM24] Wanlu Zhang and Hui Meng. Robust optimal dynamic reinsurance policies under the mean-RVaR premium principle. Communications in Statistics: Theory and Methods, 53(1): 113–143, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2022.2076121.

Zhou:2021:EQR

[ZMLH21] Yuejin Zhou, Chi Ma, Dequan Li, and Fengyang He. Estimation in quantile regression models with jump discontinuities. Communications in Statistics: Theory and Methods, 50(22): 5248–5261, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2020.1728556.

Zhang:2024:SSD

[ZN24] Qiu-Hu Zhang and Yi-Qing Ni. A sample size-dependent prior strategy for bridging the Bayesian-frequentist gap in point null hypothesis testing. Communications in Statistics: Theory and Methods, 53(21):7733–7745, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2273202.

Zheng:2020:PEN

[ZOD20] Junjun Zheng, Hiroyuki Okamura, and Tadashi Dohi. A phase expansion for non-Markovian availability models with timebased aperiodic rejuvenation and checkpointing. Communications in Statistics: Theory and Methods, 49(15):3712–3729, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2019.1708400.

Zhang:2022:EIM

[ZP22] Yi Zhang and Weiquan Pan. Estimation and inference for mixture of partially linear additive models. Communications in Statistics: Theory and Methods, 51(8):2519–2533, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1777305.

Zou:2024:UAR

[ZPJY24] Lei Zou, Jiangyan Peng, Zhiquan Jiang, and Ruonan Yang. Uniform asymptotics for ruin probabilities of a delayed renewal risk model with one-sided linear dependence and stochastic returns. Communications in Statistics: Theory and Methods, 53(5):1624–1652, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2107223.

Zhu:2023:IAB

[ZPL23] Bin Zhu, Yongzhen Pei, and Changguo Li. An improved approximate Bayesian computation scheme for parameter inference based on a recalibration post-processing method. Communications in Statistics: Theory and Methods, 52(9):2917– 2930, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2021.1963456.

Zhang:2023:LBE

[ZQH23] Shuili Zhang, Cong Qu, and Tiantian Hou. Limit behaviors of the estimator of non parametric regression model based on extended negatively dependent errors. Communications in Statistics: Theory and Methods, 52(24):8712–8724, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2069263.

Zhi:2023:VDC

[ZQX23] Kangquan Zhi, Xiaosong Qian, and Ayu Xie. Valuation of k-th-to-default credit-linked notes with counterparty risk in a reduced-form model. Communications in Statistics: Theory and Methods, 52(8):2514–2537, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1955383.

Zhao:2020:CLI

[ZS20a] Qian Zhao and Tak Kuen Siu. Consumption-leisure-investment strategies with time-inconsistent preference in a life-cycle model. Communications in Statistics: Theory and Methods, 49(24):6057–6079, 2020. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2019.1626426.

Zhao:2020:P

[ZS20b] Xufeng Zhao and Jingyuan Shen. Preface. Communications in Statistics: Theory and Methods, 49(15):3585–3588, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1776049.

Zeebari:2023:LAD

[ZS23] Zangin Zeebari and Ghazi Shukur. On the least absolute deviations method for ridge estimation of sure models. Communications in Statistics: Theory and Methods, 52(14):4773–4791, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2012.755203.

Zhang:2024:UDS

[ZS24] Lixia Zhang and Xuguang Song. Ultrahigh dimensional single index model estimation via refitted cross-validation. Communications in Statistics: Theory and Methods, 53(12):4369– 4394, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2023.2179881.

Zeng:2022:MSR

[ZSCH22] Xianfu Zeng, Haiyan Song, Yanhong Chen, and Yijun Hu. Multivariate shortfall risk statistics with scenario analysis. Communications in Statistics: Theory and Methods, 51(3): 649–668, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2020.1752726.

Zhou:2022:EFL

[ZSCL22] Jie Zhou, Guohao Shen, Xuan Chen, and Yuanyuan Lin. Efficient fused learning for distributed imbalanced data. Communications in Statistics: Theory and Methods, 51(5):1306–1317, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2020.1759641.

Zhang:2024:SSG

[ZSL24] Xuekang Zhang, Huisheng Shu, and Dajun Liu. Stability of stochastic Gilpin–Ayala model driven by α-stable process under regime switching. Communications in Statistics: Theory and Methods, 53(10):3459–3471, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2154611.

Zhou:2023:WEN

[ZSNX23] Xingcai Zhou, Hao Shen, Beibei Ni, and Yingzhi Xu. Wavelet-L_{1}-estimation for non parametric location-scale models under a general dependence framework. Communications in Statistics: Theory and Methods, 52(10):3361–3381, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1972312.

Zhong:2021:SCM

[ZSY21] Pingping Zhong, Yuesheng Song, and Weiguo Yang. Strong convergence of ρ̃-mixing random sequences. Communications in Statistics: Theory and Methods, 50(11):2592–2598, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1673415.

Zhang:2024:SSS

[ZSYD24] Jie Zhang, Siyu Shao, Kai Yang, and Xiaogang Dong. A statistical study for some classes of first-order mixed generalized binomial autoregressive models. Communications in Statistics: Theory and Methods, 53(14):5057–5075, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2205046.

Zhong:2021:CSD

[ZSYM21] Pingping Zhong, Zhiyan Shi, Weiguo Yang, and Fan Min. A class of strong deviation theorems for the random fields associated with bifurcating Markov chains indexed by a binary tree. Communications in Statistics: Theory and Methods, 50(5): 1210–1227, 2021. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2019.1648830.

Zhu:2021:SBD

[ZT21] Runyu Zhu and Dejian Tian. L^{p}(1 < p < 2) solutions of backward doubly stochastic differential equations with locally monotone coefficients. Communications in Statistics: Theory and Methods, 50(8):1856–1872, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1654611.

Zheng:2024:MNS

[ZT24] Shengchao Zheng and Zhongquan Tan. On the maxima of non stationary random fields subject to missing observations. Communications in Statistics: Theory and Methods, 53(18): 6339–6361, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2023.2244098.

Zhang:2020:TSC

[ZW20] Erhua Zhang and Jilin Wu. Testing for structural changes in linear regressions with time-varying variance. Communications in Statistics: Theory and Methods, 49(20):4889–4918, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2019.1609038.

Zhan:2022:SLL

[ZW22a] Zhouting Zhan and Qunying Wu. Strong laws of large numbers for weighted sums of extended negatively dependent random variables under sub-linear expectations. Communications in Statistics: Theory and Methods, 51(5):1197–1216, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1873380.

Zhao:2022:OLE

[ZW22b] Qian Zhao and Jiaqin Wei. Open-loop equilibrium strategy for mean-variance asset-liability management with margin requirements. Communications in Statistics: Theory and Methods, 51(13):4296–4312, 2022. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2020.1812656.

Zambom:2023:TIB

[ZW23] Adriano Z. Zambom and Qing Wang. Testing independence between discrete random variables. Communications in Statistics: Theory and Methods, 52(3):956–971, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1934026.

Zhang:2024:ROP

[ZW24a] Qiang Zhang and Lijun Wu. Robust optimal proportional reinsurance and investment problem for an insurer with delay and dependent risks. Communications in Statistics: Theory and Methods, 53(1):34–65, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2072516.

Zhu:2024:HRE

[ZW24b] Yanling Zhu and Kai Wang. Heterogeneous robust estimation with the mixed penalty in high-dimensional regression model. Communications in Statistics: Theory and Methods, 53(8): 2730–2743, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2022.2148472.

Zhong:2022:STN

[ZWCP22] Junjiang Zhong, Miin-Jye Wen, Siu Hung Cheung, and Wai-Yin Poon. Simultaneous tests of non inferiority and superiority in three-arm clinical studies with heterogeneous variance. Communications in Statistics: Theory and Methods, 51(1): 249–266, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2020.1747082.

Zou:2023:JEL

[ZWFZ23] Yuye Zou, Chengxin Wu, Guoliang Fan, and Riquan Zhang. Jackknife empirical likelihood of error variance for partially linear varying-coefficient model with missing covariates. Communications in Statistics: Theory and Methods, 52(6):1744– 1766, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2021.1938128.

Zhai:2024:ODT

[ZWLF24] Yi Zhai, Chengci Wang, Hui-Yi Lin, and Zhide Fang. Doptimal designs for two-variable logistic regression model with restricted design space. Communications in Statistics: Theory and Methods, 53(11):3940–3957, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2167056.

Zhou:2022:VMP

[ZWLG22] Congjin Zhou, Guojing Wang, Liang Liu, and Jie Guo. Valuation of mortgage pass-through securities with partial prepayment risk. Communications in Statistics: Theory and Methods, 51(15):5124–5145, 2022. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2020.1833222.

Zhang:2024:SEE

[ZWT^{+}24] Zhenzhong Zhang, Xiaofeng Wang, Jinying Tong, Tiandao Zhou, and Zhenjiang Qin. Some explicit expressions for GBM with Markovian switching and parameter estimations. Communications in Statistics: Theory and Methods, 53(3):1091– 1121, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2022.2100908.

Zhao:2024:NFB

[ZWW24] Bo Zhao, Shuying Wang, and Chunjie Wang. A new frailtybased GEE approach of the informatively case K intervalcensored failure time data. Communications in Statistics: Theory and Methods, 53(18):6527–6543, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2247505.

Zhang:2024:WCN

[ZWWW24] Lu Zhang, Rui Wang, Min Wang, and Xuejun Wang. Weak consistency for the nonparametric kernel regression estimator based on negatively associated random errors. Communications in Statistics: Theory and Methods, 53(10):3581–3598, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2022.2158342.

Zheng:2022:SIH

[ZWYL22] Zemin Zheng, Yue Wang, Yugang Yu, and Yang Li. Scalable inference for high-dimensional precision matrix. Communications in Statistics: Theory and Methods, 51(23):8205–8224, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2021.1890778.

Wu:2020:IGI

[zWZ20] Wen ze Wu and Tao Zhang. An improved gray interval forecast method and its application. Communications in Statistics: Theory and Methods, 49(5):1120–1131, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1554132.

Zuo:2022:RMS

[ZX22] Kai Zuo and Mei Xiao. A repairable multi-state system with a general α-series process and an order-replacement policy. Communications in Statistics: Theory and Methods, 51(20): 7021–7037, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2020.1869991.

Zhao:2023:MSV

[ZX23] Peixin Zhao and Liugen Xue. Modified see variable selection for linear instrumental variable regression models. Communications in Statistics: Theory and Methods, 52(14):4852–4861, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2013.777739.

Zeng:2024:HDE

[ZX24] Zhen Zeng and Feng Xu. Huber–Dutter estimation of linear models with dependent errors. Communications in Statistics: Theory and Methods, 53(23):8441–8455, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2023.2290980.

Zhou:2021:VBA

[ZXLT21] Shirong Zhou, Ancha Xu, Yongqiang Lian, and Yincai Tang. Variational Bayesian analysis for Wiener degradation model with random effects. Communications in Statistics: Theory and Methods, 50(16):3769–3789, 2021. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1846747.

Zhang:2020:BRP

[ZXSZ20] Ying-Ying Zhang, Yu-Han Xie, Wen-He Song, and Ming-Qin Zhou. The Bayes rule of the parameter in (0, 1) under Zhang’s loss function with an application to the beta-binomial model. Communications in Statistics: Theory and Methods, 49(8): 1904–1920, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2019.1565840.

Zhou:2024:IHS

[ZXYJ24] Xiaojian Zhou, Dan Xiao, Jieyao Yu, and Ting Jiang. Incremental Huber-support vector regression based online robust parameter design. Communications in Statistics: Theory and Methods, 53(8):2924–2944, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2150056.

Zhao:2020:SPO

[ZY20] Qian Qian Zhao and Won Young Yun. A sampling plan for one-shot systems considering destructive inspection. Communications in Statistics: Theory and Methods, 49(15):3744– 3760, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2020.1719159.

Zhu:2023:ORP

[ZY23] Dan Zhu and Chuancun Yin. Optimal reinsurance policy under a new distortion risk measure. Communications in Statistics: Theory and Methods, 52(12):4151–4164, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1986538.

Zhao:2024:RCC

[ZY24a] Xiaofeng Zhao and Mingao Yuan. Robustness of clustering coefficients. Communications in Statistics: Theory and Methods, 53(20):7144–7180, 2024. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2023.2259525.

Zhu:2024:IEE

[ZY24b] Lin Zhu and Feifei Yan. Improving estimation efficiency for multivariate failure time data with auxiliary covariates. Communications in Statistics: Theory and Methods, 53(1):260– 275, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2022.2077960.

Zhou:2023:CCM

[ZYC23] Jinyu Zhou, Jigao Yan, and Dongya Cheng. Complete convergence for maximum of weighted sums of WNOD random variables and its application. Communications in Statistics: Theory and Methods, 52(22):8184–8206, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2059681.

Zhang:2024:LOD

[ZYC24] Min-Jue Zhang, Rong-Xian Yue, and Xue-Ping Chen. Locally D-optimal designs for spline measurement error models with estimated knots. Communications in Statistics: Theory and Methods, 53(10):3713–3727, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2161823.

Zhang:2020:SBL

[ZYJ20] Zhengcheng Zhang, Yonghong Yang, and Xiujie Ji. On stochastic behaviors of load-sharing parallel systems. Communications in Statistics: Theory and Methods, 49(22):5392– 5405, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2019.1642491.

Zhou:2020:OPC

[ZYNX20] Lei Zhou, Hisashi Yamamoto, Taishin Nakamura, and Xiao Xiao. Optimization problems for consecutive-2-out-of-n:G system. Communications in Statistics: Theory and Methods, 49(15):3792–3807, 2020. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2020.1772980.

Zhang:2021:PEC

[ZYS21] Xuekang Zhang, Haoran Yi, and Huisheng Shu. Parameter estimation for certain nonstationary processes driven by α-stable motions. Communications in Statistics: Theory and Methods, 50(1):95–104, 2021. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2019.1630436.

Zhang:2023:SIE

[ZYW23] Chunxiu Zhang, Ping Yu, and Xiaofeng Wang. Statistical inference in EV linear model. Communications in Statistics: Theory and Methods, 52(2):346–363, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1914096.

Zhong:2020:SLL

[ZYY20a] Pingping Zhong, Weiguo Yang, and Jie Yang. Strong law of large numbers of the delayed sums for Markov Chains indexed by a Cayley tree. Communications in Statistics: Theory and Methods, 49(9):2285–2294, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1571611.

Zhou:2020:SPG

[ZYY20b] Qi Zhou, Xue Yang, and Ziyang Yang. Some properties of general minimum lower-order confounding designs. Communications in Statistics: Theory and Methods, 49(8):1921–1932, 2020. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2019.1565842.

Zhao:2020:ELB

[ZYZ20] Peixin Zhao, Yiping Yang, and Xiaoshuang Zhou. Empirical likelihood based estimation for a class of functional coefficient ARCH-M models. Communications in Statistics: Theory and Methods, 49(5):1217–1231, 2020. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1554139.

Zhou:2024:NNC

[ZYZ24] Wenhui Zhou, Ziyu Ye, and Zhibin Zheng. A new np_{CEV} chart for monitoring process mean shifts based on an attribute inspection. Communications in Statistics: Theory and Methods, 53(2):666–686, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2089356.

Zhang:2021:NMS

[ZZ21] Jianling Zhang and Zhongzhan Zhang. Nonparametric multisamples test for simple stochastic ordering against unrestricted alternative. Communications in Statistics: Theory and Methods, 50(21):5160–5169, 2021. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
03610926.2020.1726389.

Zuo:2023:NFM

[ZZ23] Kai Zuo and Hang Zuo. A new fractional modified exponential curve model and its applications. Communications in Statistics: Theory and Methods, 52(20):7206–7222, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2042026.

Zhao:2022:LSE

[ZZGL22] Huiyan Zhao, Chongqi Zhang, Yu Guo, and Sheng Lin. Least squares estimator for a class of subdiffusion processes. Communications in Statistics: Theory and Methods, 51(15):5342– 5363, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2020.1838546.

Zhang:2023:NIM

[ZZK23] Jiayue Zhang, Fukang Zhu, and Naushad Mamode Khan. A new INAR model based on Poisson-BE2 innovations. Communications in Statistics: Theory and Methods, 52(17):6063– 6077, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2021.2024571.

Zhang:2022:OIR

[ZZRH22] Yongtao Zhang, Hui Zhao, Ximin Rong, and Kai Han. Optimal investment and reinsurance problem toward joint interests of the insurer and the reinsurer under default risk. Communications in Statistics: Theory and Methods, 51(19):6535–6558, 2022. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2020.1862872.

Zhang:2024:EBE

[ZZW^{+}24] Ying-Ying Zhang, Yuan-Yu Zhang, Ze-Yu Wang, Ya Sun, and Ji Sun. The empirical Bayes estimators of the variance parameter of the normal distribution with a conjugate inverse gamma prior under Stein’s loss function. Communications in Statistics: Theory and Methods, 53(1):170–200, 2024. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2022.2076123.

Zhao:2023:EQR

[ZZY^{+}23] Weihua Zhao, Xiaoyu Zhang, Kam Chuen Yuen, Rui Li, and Heng Lian. Estimation in quantile regression models for correlated data with diverging number of covariates and large cluster sizes. Communications in Statistics: Theory and Methods, 52(4):1012–1038, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.1922701.

Zheng:2020:NAE

[ZZZ20] Yanqiao Zheng, Xiaobing Zhao, and Xiaoqi Zhang. A novel approach to estimate the Cox model with temporal covariates and application to medical cost data. Communications in Statistics: Theory and Methods, 49(18):4520–4535, 2020.

CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2019.1602651.

Zheng:2022:QRM

[ZZZ22] Yanqiao Zheng, Xiaobing Zhao, and Xiaoqi Zhang. Quantile regression for massive data with network-induced dependence, and application to the new York statewide planning and research cooperative system. Communications in Statistics: Theory and Methods, 51(9):2962–2993, 2022. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2020.1786120.

Zhou:2024:MAD

[ZZZ24] Yuliang Zhou, Shengli Zhao, and Qianqian Zhao. Minimum aberration 4^{1}2^{n} designs via secondary complementary sets. Communications in Statistics: Theory and Methods, 53(11): 4153–4171, 2024. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03610926.2023.2174787.